A Complete Bibliography of Publications in Annals of Applied Probability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.26

Title word cross-reference

(max, +) [BS96]. 1/2 [LS17b, LS19]. 2 [BCOR16, BD20, BNS13, LT22, RW97]. 2D [Bal19]. 2 × 2 [Man93]. 3 [CS16a, FHLN22, GLS18]. 3x + 1 [LW92]. 4 [BM96]. [0, 1] [BJM10b]. ^{d} [CT11]. A · /M/K [Ana94]. α [AS97, DF20, ERY95, KR06, STZ14]. Arch(1) [BK01]. b [Sch10]. β [AN22, Dre00, LP23]. C^{1} [DP20]. · [Ana93]. ·/M/1 [MP95]. c|mu [vM95]. d [GTW22, LP17b]. D = 2 [DN91]. ϵ [BCD17]. f^{q} [JKM07]. G [Nut13, Jia08]. G/G/1 [BPT98, AB14]. G/GI/∞ [RT15]. G_{n,p} [JŁTV12]. H = 0 [HN22]. ∞ [DM08a]. K [FGG14, COMR22, FLR22a, JL08, Yao97]. κ [JKS18]. L [CCM06, TBUA23]. L^{2} [CHS17, Cho09, DM94, LW22]. L^{p} [Lud08]. Λ [BBL14, DK19, EV12, EFPS17, GIMM18, Hén15, KR23]. M/M/1 [Fla97]. M/M/N + M [ABP15]. M_{1} [PW10]. R [CT11]. R^{d} [CQ97, MSW97]. S^{2} × R [MRV21]. Z [ESTZ13]. Z^{2} [ADH15, BGV20]. Z^{d} [AC03, BZ15, BCP^{+}22, FN93, FN94, Zer98]. n [HLOST16]. N [CF18, Fis17, ZRH15, FP93, KPR10, Mar16, PS17a, Smi14]. NK [ES02]. n log n [PS17a]. p [Lal00]. P H/P H/1 [Bla96]. π [HR04]. q [BC16]. R [FKM96, Alv03, CK00, DK92]. σ [BB20c, JMRS09]. t [AS23]. T^{2} [PZ11]. Θ(n^{2} log n) [Jon06b]. U [DKP22, PSZ14, RR97a]. U(∞) [BBO15]. Ξ [CS18].

-ary [GTW22, Sch10]. -chains [KR06]. -coalescent [Hén15]. -coalescents [BBL14, DK19, GIMM18]. -color [GLS18]. -coloring [BCOR16]. -convergence [LW22]. -copulas [LP17b]. -core [JL08]. -cores [AS23]. -Cube [FP93, Mar16]. -CUSUM [ZRH15]. -cut [JKS18]. -cutoffs [CHS17]. -D [BNS13]. -diversity [DF20]. -ensembles [TBUA23]. -excessive [Alv03]. -expectations [Jia08, Nut13]. -fair [STZ14]. -fields [BB20c]. -finite [JMRS09]. -Fleming [EV12, KR23]. -Fleming-Viot [EFPS17]. -Hahn [BC16]. -linear [BS96]. -martingale [JKM07]. -martingales [BJM10b]. -mer [FLR22a]. -mixing [Dre00]. -norm [Cho09]. -player [CF18, Fis17]. -positivity [FKM96]. -process [FGG14]. -processes [DKP22]. -Quantile [ERY95]. -record [Yao97]. -reversal [CCM06]. -SAT [COMR22]. -Scan [DK92, CK00]. -shuffles [Lal00]. -simplex [Smi14]. -space [CS16a]. -sphere [PS17a]. -stable [AS97, BM96]. -statistics [PSZ14, RR97a]. -Strong [BCD17]. -tuple [KPR10]. -variations [Lud08].

/M/K [Ana93].

1 [CLW94, OCBG11]. 11 [Web01a].

20 [BvdHH17]. 23 [LO17]. 24 [Tou18]. 2D [Gar09].

=2 [vEK08].

A.S. [FP95a]. abandonment [KR12a, TW09]. Absence [HHK21, LS95]. Absolute [Tak93, Vit91]. Absolutely [DS95, AS97, CFY05]. Absorbing [BGR22, Gos01, HL17]. absorption [CF18]. Abstract [CD99]. Accelerating [HHMS93, HHMS05]. acceleration [EP10, MW98]. access [CBM^{+}21, SS18]. accessible [HM14]. Accuracy [KLS06]. ACFs [RSX99]. Achievable [BPT94]. Achieving [IV17]. Achlioptas [MS15, RW12]. acknowledgment [NV06]. across [ABT^{+}11a, ABT^{+}11b]. activity [BdHNT22, JT18, TKH09, TT11]. actual [Bla96]. Acyclic [LC93, BPT98, KLSY04]. adaptation [CH19, Kel13, YEC10]. Adapted [BLO23, BBBW22]. Adaptive [BCKP99, BDK06, ELM22, FG20, HHH09, ŁRR13, AM06, Atc10, BBC17a, BJKT16, BGG^{+}16, Cal97, CG16, CL07a, CGL^{+}15, DL16, DG21, GVR17, JL09, SV10, Zha16a]. add [TLC93]. add-with-carry [TLC93]. Addendum [BK17b, KP98, LP17a]. addictive [Yu15]. Additive [Col02, Ber02, CDFM20, GH22, HMGR00, JKO09, JKW11, Miy04]. adiabatic [JLZ21]. adjacency [DJ10]. adjacent [NN19]. adjusted [ZHC06]. adjustment [Mei09]. adsorption [CGZ07, Sud08]. advection [Gau98]. advection-diffusion [Gau98]. adversarial [CGL^{+}15]. adverse [NZ15]. advice [BEZ20]. affected [HRS99]. affecting [Lan15]. Affine [CFMT11, DFS03, JLP19, KRSW19, AA13, FJR20, KRM15]. after [CKL22]. Age [BK95, SJ05]. Age-Structured [BK95, SJ05]. agents [HP23]. ages [GT99]. aggregation [AAK17, FP18, TV18]. aging [BČ05, BF05, Guy07]. ahead [EKT07]. AIDS [Ish93]. AIMD [GRZ04]. Airy [DL22]. AKT [BL21]. Aldous [CD11, HS19a]. Algebraic [CW03a, BM05]. Algorithm [DFP93, FP95a, FP95b, Ing94, BRTP18, BP97, Cer96, CMY03, DH13b, DH18, DM06a, DRVZ17, DM17, EKT07, FMS96, Fil98, Ger11, Gu22, HSV14, JR14, JLM15, KMPT10, LPT04, MZ14, MPS12, PP23, PST12, SV10, SS12, SS18, Zha16a]. Algorithms [Fra02, AM06, AV15, BB22, BLM15, Béd07, Bel11, BH99, BGvdHK15, BB03, BKR22, BBFM03, BGG^{+}16, BDS20, Cal97, CMSS15, CL22, CBM^{+}21, CC98a, CAP20, CGL^{+}15, DOS19, Der11, Ebe14, Egl05, ELM22, Fri16, GGR97, GRZ04, Hub15, LC03, LW22, MMS20, Mic02, MR05, MR10, MP06, MV06, NR06, NRY12, NR04]. Alignment [DR13, Cha05, Han06a]. Alignment-free [DR13]. Allee [Bor12, Sch05a]. allele [Pap00]. Alleles [Mor92, JKK03]. Allocation [EK94, BM12]. allocations [AGGL10]. Allowing [AW94, HPTvD95, Zha95, DK99a]. Almost [BDL16, GL14, HHK20, Jor02, LV10, Nak11, Ber97a, GK16, vdHMS08]. along [Mat05]. Alpha [RWF13]. Alpha-diversity [RWF13]. alphabet [HT12]. alternating [PR98a]. Alternative [CH91, OT22, Sch05a]. alternatives [LL17]. ambit [CSS22]. amenability [TT13]. amenable [SSV18]. American [AK05a, BZ16, CKL22, DFT03, GY04, JM02, Lam98, LV03, Myn92]. among [FJ19, LN06a, PS20]. amongst [AV16]. Analysis [Aal92, ACLW95, AFRT06, BCTV07, DH91, DH92, DH93a, DHN00, DFH13, EOT05, FRST94, FP95b, Goe06a, HSV07, KL01, MW94, Ngu93, OC11, Pal11, PRW95, Rhe94, SFR16, ARS18, Ala03, AGK11, AD20, AK05c, BGHM10, BKPR06, BJ22, BRTP18, BCD17, CJL21, CL22, CG16, CMT21, CS02, DR13, Dec98, DG05, DH23, DG14, DKPRB20, DM17, EGK22, GHR03, Har98, HL00, IM02, KM13, KS06b, KLRS11, KT03, LR12, MR02, MR05, PP08b, QH21, QH22, RS03a, RV15, RS07, WH17]. Analytic [BH00, Kah08, BL21, IH23, KPS98]. analytical [FKP23]. analyze [MW13]. ancestor [ER10]. ancestral [CHS22, DFPR22, LPS21]. Andersen [BMX17]. Anderson [BC14a, CDS23, GK00, GS18, MPdS19, vdHMS08]. Andjel [Mou01]. Animals [CGGK93, GK94, Lee93, Alb09]. anisotropic [CM03, She02]. Ann [Tou18, Web01a]. Ann. [BvdHH17, LO17, Ano99, Ano02, Ano03]. Annealed [Can19, Ton08]. Annealing [FFS93, Fra02, CC98a, CC98b, Már97, Pel98]. annihilating [BEM07, PSY15]. annihilation [JL22, PS05]. annulus [BBW04]. Anomalous [BFM11, CZ18]. Antithetic [GS14, ST21, BKN22]. antivoter [RR97a]. anytime [BS21]. appear [DS07b]. appearing [PY98]. Appl [Ano99, Ano02, Ano03, BvdHH17, LO17, Tou18, Web01a]. Application [Ber92, Ber94, DH91, Del98, Fil91, Guy07, KO92, LN13, Mor92, NS19, Pèn05, ZZ02, BT05, Bha99, BT96, BN19, CW99, CS11a, CG17, CFJ16, CGL^{+}15, CSS22, DG07, GHLT14, GPR17, GR97, HK13, HNS11, HSX22, JT18, Jia19, JPV99, LO04, LL12, LLS08, LP08, MV20, Owa18, Pel98, PDG14, RR14, Sch97b, TKH09, TT11, WPRS21, WH17, dBG12]. Applications [CLW94, CMP94, CHO08, Dol10, Eic95, El 09, Fin94, Gos01, KLR91, LS20, Lig95, Miy04, Per94, Pha02, Tak92, AC20, ATV15, ABT^{+}11a, ABT^{+}11b, ABKR18, AKP04, BCHL98, BHZ02, BBM07, BCP18, BT08, Bha16, BJR16, BM97, CGR20, CL15, CP19b, CST05, CL17, CAP20, CW13, De 11, DFS03, EK09, FK22, FMMP08, GR21, Gol13, GK16, HR04, HS99, HH18, JSW19, KMS17, Keb05, KY10, LRM15, LP17b, MM03, MP98, NY23, Niu97, OTV12, QH22, RR19, RR97a, TSM19, YZ07, Zha16a, Zha16c]. applied [Ebe14]. Appriximated [MRRS02]. Approach [AB92, Dai95, HS93a, Roo94, Ser94, BPG19, BCPG22, BCFP18, BT22a, BK16a, BK17b, BG02, BF22a, BF08, BG05a, BG05b, CRT17, CGZ07, CZ16, CH19, CL11, CSX23, CC98b, CKHL06, Con22, CFJ16, CI11, CLR06, EMO10, FHH23, Fou00, FRLS21, GHLT14, Gué03, GZ08, Haa10, HIP06, INPY13, JKP13, JM12, JY17, Käl22, KPS98, KT03, LLC18, LC22, MWZZ15, May09, Nag12, RSS16, SZZ23, TKH09, VZ19]. approaches [KMZ17]. approaching [Gri16]. approximants [JS12]. Approximate [BKW08, GS02, CRV06, EPW06, MH21, PS05]. Approximating [ADS14, CX02, DS19, HK16, HR97, Mar19, RR19, Clé23, LST23, SD05]. Approximation [Ban19, BG93, BR04, BK95, GGO23, Hub15, JNP21, JM02, KX95, KLR91, KLST93, Loh92, MW94, RWW95, RS95, Roo94, See20, Xu19, Yam95, ATV15, APW08, BHZ02, BC01, BH99, BF12, BCP18, Ber97a, BH01, BH03b, BT96, BG96, CV21, CPV23, CEK12, CS11a, CT23, CSX23, CKHL06, CG17, DFFN21, Erh00, FS16, FJL18, FJ17, Fri16, GHR03, GM19, GPR17, GM13, Gol04, GP10, Gro04, GS05, HHMG19, HLT21, HR09, HMGR00, HT12, HLN16, HPS03, JMP21, Jel99, KKLW09, KLP15, KT04, KL04, LRSY19, LP23, Lam98, LP13b, LP17a, LNS21, LO04, LW19, LN05, Mån99, MV20, Mek15, MP06, MG02, MG04, MGY23, NT20, NY23, OC11, OTV12, PP12, Pel98, PVY20, PT15, Röl07, SY23, SW12, SZZ21, STT19, TY16]. approximation [Tan14, Tor16, Yos12, dBG12]. Approximations [ABT92, Che95, DK92, DZ16, EW01, GN91, RSM09, AV16, BKPR06, BJ22, BG06a, BD17, BG06b, BF18, CL03, CS00b, CCK23, Cla96, CJK19, Cou08, DK08b, DK08c, Dre00, ELM22, FK22, FM04, GZ06, Gur14, KR12a, KKP14, Kel16, Kif06a, Kif06b, KL02, KS05, LO13, LO17, LR13, PY98, RZ99, Sab16, Sch05b, TKH09, Tsu19, YY18, Zha16b, ZW08, MP98]. approximative [FR11]. Arbitrage [BZ16, BN15b, DH93b, LS95, Mil94, Ben12, BT13, CE10, FK10, FK11, FKR18, GHLT14, GR15, JPS09, Oha09, Str05]. arbitrage-free [Str05]. arbitrary [GLS18]. arc [Fus00, Tak96]. arc-sine [Fus00, Tak96]. ARCH [CmHP04, SZ06]. Archimedes [BCP11]. Arcs [CMP94]. Arctic [FV21]. Area [Hsi94, GS14]. areas [PW96]. ARIMA [Ino02]. arising [APS19, DP09, DSS09, HSV07, KP16, MD18]. Armed [KM95, GW00, GZ09, KM98, KR96, LPT04, TV12]. array [CK00, HHR96, LRdH98]. arrays [ACH97]. arrival [LRT03]. arrivals [CQ97]. ary [GTW22, Sch10]. Asexual [AR02, Kel13]. aspects [Grü14]. asset [GHK11, HH08, Pul14]. assets [DFT03]. assignment [Jár17]. assisted [BE23]. Associated [JKK02, AR19a, AMS06a, BK00a, Fen07, ILP15, Le15, Zha05]. Asymmetric [FR92, AC03, BC16, BFG13, Ber97b, ÇD16, DMP22, GLM17, GKS04, JL22, LL20a]. Asymptotic [ACLW95, Ald91, AK05c, Ata05b, AG14, BYZ00, BLW11, BKPR06, BP05a, Bäu00, BH03a, Ber10, BRTP18, BMM21, CRT17, CJL21, CD11, CJY15, DL08, DF06, DDSJ08, DS93, ESS93, EMO10, EF21, FW22, GGS03, GM12a, GW92, GM95, Haa10, HYTC20, HL21, Hsi94, HDP22, HL00, Ino02, JJ15, JL08, Jel99, JM03, KR12a, KL01, Kne00, KLST93, LR99, MP09, MR93, Ngu17, PGZ07, PR94, Pop04, Rei95, Sad98, She02, Sta97, Ste93, TD17, Yan05, YRF16, YEC10, ZHC06, AR20, AMR04, APP08, BGHM10, BN17, BF10, BH13, BEZ20, BW01, BBC17b, BB03, BG06b, CSS98, Coh96, De 11, DG08, DL18a, DK19, DRS16, Har98, HLN16, JS10, Jia04, KR12b, KR21, Ker12, KL04, KS99b, Kum00, KT03, KPR10, LLS08, Mag00, PP08b, PY98, RS07, TY16, Yat09]. Asymptotically [Ala03, ABB23, AC17, RT14, Ver16, ARS17, BG05b, BM96]. Asymptotics [BH05, BC02, Cho06, HM09, HNS10, Kni12, KS03b, McD99, Yuk99, AG06, AZ16, Asm98, BM19a, BOC22, BvdH12, BBDW22, BFW21, BGR22, CGZ14, DK99a, DLS03, FFK12, FM01, FM05a, FM05b, FGP21, GK00, HK22, HKV20, MV16, Mey06, OWZ97, Pau02, PRR13, ZBM04, ZBD05]. Atlas [BFK05, IPB^{+}11]. ATM [IMQ93]. atoms [HN19]. attachment [BOC22, BJN18, CS13a, CCL13, JM15, JW21, JK22, PRR13]. Attraction [CL09a]. Attractive [KV01, AC03]. attractor [DSZ15]. Augmentation [Ros93]. auto [JWB^{+}14, WJB^{+}15]. auto-cross [JWB^{+}14, WJB^{+}15]. autocorrelation [Ino02]. autocovariance [YY22]. Automata [FR92, BS23, FKM96]. automaton [HL21]. Autoregressive [BK01, Ros95]. Average [Cal97, DFP93, GR06, Jaś07, KL04, BH99, CCHH05, CSS22, GK21, HSZ17, KR96, Wu09]. average-overtaking [KR96]. average-reward [KR96]. averaged [BMRS23, FMP00]. averages [Che08, DL18a, FKR96, HRS99, HI09]. Averaging [CPR95, LN13, Cer09, MP06, QS23b]. Avoiding [Jac02]. aware [IH23]. Axelrod [Lan12, LS13]. Aysmptotic [Gos01]. Azéma [KTT17]. Aztec [AJvM22, BCJ22, CJY15, FV21].

back [FKK^{+}01]. background [BES04, BGZ22]. Backward [Ant93, BCFP18, CC16, CFJ16, DI10, MPST02, BQ19, Bec06, CDET13, ÇD16, CC14, CC98b, CE10, CM14, CM96, DM06a, DEH23, DMP96, FP15, GLW05, HNS11, KX22, MZ02, MWZZ15, MPZ13, MPZ21, Moy15, QX18]. Backward-Forward [Ant93]. bacterial [BHP10]. Bakry [JY17]. balance [BFJ06]. Balanced [BM12, Haj96, LM15, SV94]. balancing [AR19b, GTZ20, KPL03, SY13]. Ballistic [DS17, JL22]. Balls [KL91, LM05, Sel95, BGL02]. Band [LR21]. Band-limited [LR21]. bandit [GZ09, KR96, LPT04]. Bandits [KM95, Web92, GW00, KM98, TV12, Ver16]. bandwidth [Bra10, FW22, GW09, KKLW09, KW04, STZ14]. bandwidth-sharing [KW04, STZ14]. bank [BCKWB16]. barcodes [Owa18]. Barndorff [RS06]. Barrier [BLSW91, BL02, Bar20, CW13, GMO15, Han06b, Loe08]. barriers [Kah08]. Barron [CPV23]. barycenters [KSS21]. Barycentric [ABLT22]. Based [ACLW95, HK92, ACG17, BBC17a, Bel11, BJ22, CBM^{+}21, DOS19, DPR09, DKPRB20, DRVZ17, GW00, GK96, GLW05, JS96, Jia15, KS06b, MSZ19, MR17, MG04, Owe02, QH22, San10, vdHHKR16]. Basic [MR08]. basis [GY04, LT18]. Bättig [Dur99a, Dur99b]. Bayesian [KPS98, LPW14, AKLT22, BGHK20, DPS08, LRS17, LPW08, MV06, WHN07]. Bayesian/analytic [KPS98]. be [AL05, LPT04, Pit99, RSX99]. Beardwood [AS16b]. beat [FW99b, RvH15]. Beating [Mos01, RZ18]. Beckner [JY17]. become [Sch01]. bees [ABLT22]. behave [EG18]. Behavior [BM01, Gos01, GKS04, PR94, dGvZ93, Asm98, BP05a, BGT22, BGZ97, BBL^{+}97, BPT98, CMT21, CD11, CST05, CW03b, DF06, EMO10, EF21, FW22, GLS18, GM12b, Gri16, Haa10, HW96, HW07, Imh05, Ino02, KK01, LR99, LV10, MM10, OC21, Pap98, Sad98, SCZ10, Sta97, YEC10]. Behaviour [Gre94, Per94, Fox16, HL21, Kar15, MP23]. Belief [COMR22, KS14, MNS16]. Bellman [QZ23, GMŚ19, Nut12]. beneficial [BL16b]. benefit [Che13]. benefit-to-cost [Che13]. Benes [KLS95]. Bermudan [EKT07]. Bernoulli [BGT22, BMM21, EHW16, GQ03, GINR09, MLNW22, Ser94]. Bernstein [CHS22]. Berry [DT05, DF20, Gol13, HK22, Klo19, LRP17, LS09, LLS08]. Bessel [DH13b, FJ17, JP17]. best [DGR09]. Beta [Ker12, MP23]. Beta-coalescent [Ker12]. Betti [Owa18]. Between [HR94, AJKH14, AS10, BL16b, BS19, BBL14, BNT19, BKW08, CS00a, Chi05, CTZ04, DK99a, Fis17, HHSZ15, LSZ13, LRS17, MWZ07, MR17]. Beyond [EGP16, Ton08, Yan20, Chi15, CS17b, El 09, KRSW19]. bi [CCR23]. bi-dimensional [CCR23]. bias [BGvdHK15, ELM22, GR97, TD17]. Biased [Jon06a, Cha97, PY18, SSY19, Win08]. biasing [ELM22]. Bienayme [QS94, KR19]. Bienayme-Galton [QS94]. bifurcating [Guy07, PDG14, PD22]. Bilinear [IT99, DR96]. billiard [Pèn05]. billiards [Eva01]. Binary [CDJH01, DF95a, DF95b, Mah94, VG95, Bla09, DL18b, GS05, Grü09, HM09, MN03, MW08, PP04, Pem09, WM04]. Binomial [CX02, DK08b, DK08c, GP10, HDP22, Kif06a, Kif06b, LRP17, Lam98]. binomial-exhaustive [HDP22]. Bins [BGL02, LM05]. biodiversity [FK00b]. biological [AS10]. Biology [Gos01]. bipartite [BdHNT22]. biphased [Pia99]. birefringent [dBG12]. Birth [YH93, BBL^{+}97, CA18, Che18, DSC06, Erh00, ESU10, KST04, Lan15, Web01a, Web01b]. birth-and-death [ESU10, Web01a, Web01b]. birth-death [Erh00]. Birthday [KMPT10]. bisections [MP20]. Bisexual [AR02, AR96]. bistability [MP14]. bit [DM23b]. Bjerknes [LTVR14]. Black [Çet18, DK08a, DMY95, ET09, Loe18]. Blackwell [CvH10]. block [EG18, GLM18, MNS16]. blocking [BS17b]. Blow [Cho09, HLS19a]. Blow-up [Cho09]. blow-ups [HLS19a]. Blume [EOT05]. Board [Ano16a, Ano23a, Ano23b]. boards [BHMW16]. Bochner [JY17]. bodies [BR04]. body [JKM15]. Bolthausen [DK19]. Boltzmann [Bal19, Fou15, Gen23, Hey23, Xu18]. Bond [Pen93, CS00a, DP05, ET05, EFdS21, FN17, Taf11]. book [Swa18]. books [BHQ17, BJR16, HSH^{+}13]. Boolean [FP18, HLS16, LPZ17, MR94b]. Bootstrap [GHPS15, GS20, JŁTV12, CCK23, FKKM18, GH08, GHS21]. Border [TV18]. borrow [TLC93]. bottlenecks [CPSJ22]. bottom [Goe06a]. bottom- [Goe06a]. Bouchaud [BČ05, MPdS19]. bouncy [DPBCD21, DGM20]. bound [ABA20, BK00a, BG06b, GTZ20, Gol13, HS07, Jia12, Kel13, KMPT10, Lez98, LLS08]. boundaries [GNS23, KW07]. Boundary [ET11, FM94, Hol01, Loa92, RS95, BGV20, BL12b, BR08, CPV23, CL03, CCCS11, CI11, DHL18, DH18, DL18b, EF21, Fer15, FPZ05, GLM17, KRSS23, LST23, Mül18, Pes19]. boundary-driven [BL12b]. Bounded [Bec06, HW92b, ELM^{+}16, Kol17, MT99, Mon22, CDS23]. Bounding [KPL03, Hub04, RR08b]. Bounds [AR20, CS11b, CGGK93, DS91, Fil91, GN91, Har13, Ing94, Las02, Las04, MT94a, OW92, AJKH14, ABDW21, Bal19, BBC23, CRT17, CT23, CDS09, DRZ16, DMR04, Ebe14, EI17, FK22, FM16, FW99b, GHLT14, GH21, GGO23, HK22, HHMG19, HL20, HL22, JSTV04, Klo19, KM08, LRP17, LP04, LP17b, MMS20, Mei09, MRW18, MGY23, PP23, Sch05b, Ton08, YR23]. Box [Sel95]. Bradley [CDL17, HYTC20]. braid [MM07]. brain [CW99]. branch [DK15]. Branching [Ath94, AV95, BZ15, BLSW91, Big95, BNT92, Che01, CJ94, EFPS17, HLT21, Jag99, JR92, Jof93, KS93a, KZ94, McD95, NV04, Olo96, QS94, ABF13, ABK12, ABH17, BBC23, Ban08, BBL14, BPZ07, Ber10, Big12, BD07, BEM07, BH17, BH19, BD15, BR13, CA18, Cla18, CRT23, Coh96, CW03b, DR13, DJ12, DMM17, DKM17, ER10, FL22, GHH07, GLP23, GPW09, GK16, HMT22, HW07, ILP15, Jag97, Jon97, KLZ23, KS05, Lam05, LP13a, LSZ13, LYZ19, MPP17, Nak11, NV03, NV06, PSY15, Pia06, Pop04, SS15, YY09, Yos08, Yu07a]. Breaking [ABL21, CDG23]. bridge [CS00a]. Bridges [FM05a]. Broadcasting [ABDLV22, EKPS00]. broader [Har03a]. Brownian [ABT^{+}11a, Har03b, Har06, ABLT22, ABT^{+}11b, ABK12, Arm10, AGP95, AHS05, AC19, ABL21, AG09, BB20b, BB23, Bar20, BM13, BC15, BH17, BDH10, Bra11a, BR08, BK98, BCP11, Che96, Che08, CS06, CD99, CK03, CS17b, DH92, Das95, DeB04, DR08, ERY95, EFdS21, EPZ20, EF21, EP22, EFPS17, EEH14, EHW20, FK00a, FH98, FSW15, FK99b, FK00b, FJ19, GMO15, GHP13, GZ00, HN22, HW92a, HV97, Har00, Har03a, HW05, HLN21b, IK10, IT99, JR92, Kah08, KS93a, KW07, KS16, KL99, KLZ23, Leh23, Lej16, LT19, MNG09, MR00, MRRS02, MR08, MG04, Ngu93, PP08a, PS14, PW96, Pov95, RS15, RSS16, SV94, Tak93, Tak95, TW07, VZ19, Wik01]. Brunet [BZ18]. brute [BS21]. brute-force [BS21]. BSDE [JKP13, PZ19].

BSDE-decomposition [JKP13]. BSDEs [MPZ21, BVLT20, BL14, BEH18, CEK12, CR16, CM08, ESX22, GIO^{+}17, IRR12, JMSS12, KTPZ15, KLP15, KP16, LdRS18, PT15, Ric11, Zha04, Zha05]. Bubbles [ET09, KKN15, LY22]. Bucy [DT18]. Buffer [KLS95, DR98, GK09, HRS97, JM03, MR06a]. buffet [BCPR15]. built [YA15]. Bulk [He20]. Bures [KSS21]. Burgers [BFS21, BT96, FSW95, HW94, MSW97]. business [CTZ04]. buttons [FKK^{+}01]. Buying [DM92].

caching [Jel99]. Cahn [RYZ21]. Calculation [CS95, PRW95, CL09b]. calculus [BPG19, BCPG22, CST22, Dec98, HNS11, KL99]. calibration [GL21]. call [DFT03]. calls [LM15]. Camassa [CDG23]. Can [BT17, RvH15, BL16b, BCDS15, JKK03, LPT04, RSX99]. Canadization [KP03]. Canadized [AKP04]. cancellation [JMRS09]. cancer [Dur13, FL13, FGLS23, MD18]. canonical [BBC17a, CL07a, Har00, Har03b, Har06]. Cantelli [LZL21, SSV18]. capacitary [SS19]. Capacitated [Rhe94]. Capacitive [Gar09]. Capacity [BW22a, IMQ93, CDGL22, HM09, LPZ17, Löw98]. caps [Ott13]. carcinogenesis [DSS09]. Card [APW08, Ciu98, MNP14, Wil04]. card-cyclic-to-random [MNP14]. cardboard [BT17]. Caricature [Ana91]. Carlo [ABL12, AV15, ADNR21, Atc10, BH20a, BM22, BK15b, BKN22, BCJ14, BJKT16, BRSS17, BREZ20, CL07b, CL11, DD10, DPBCD21, Der11, DL16, DRZ16, DGMO11, DMO14, DG21, DG95, Egl05, EKT07, FHY92, GS14, GM13, GLW05, HLT21, HR04, HJK13, KKPvS11, MS21, MV06, Sad96, WKRS19, WPRS21]. carry [TLC93]. Cascade [LC93]. Cascades [HW92b, LR00, BJM10a, PRSS20]. Case [Ath94, AV95, AR96, AAK17, BK15a, BTZ04, Bud02, Bur07, CL22, Cha97, CL04, Cla22, DM15, HSV07, HP15, LT19, MMPP08, NST20, OCBG11, Sab16, Tom21]. casino [DFH13]. cat [LR12]. catalyst [BR13]. catalytic [BN97, DKM17, FK99b, FK00b, GPW09]. Cauchy [BZ10, HLT21, KR19]. causal [BL11]. cautionary [Atc10]. cavity [KM11]. Cayley [AP17, Gol16]. cdf [JM08]. CDMA [BS07]. cell [CS99a, Dur13, Gup12, dGvZ93]. cells [AS23, Ban08]. cellular [BS23, Guy07, HL21]. censoring [IM02]. Centered [Ste93, Mar08]. Central [Ata08, AB93, Bal21, BR19, BK15b, BKN22, BCPR15, Cao21, CQ97, CSS22, DG99, DL16, DKT91, Fér20, GR08, GVR17, HSS06, KKP14, Muk22, NS22, PZ08, PZ11, PD22, PY01, Yos08, Zha16a, BC22a, CCK23, CES23, DM08a, DGR09, DR10, FK21, FRT03, GK16, HSH^{+}13, HLS16, KL96, Lee97, Mer07, Nak11, Pèn05, Sei09, ZY96, Hwa96]. centrality [BB22]. Certain [Hog93, PRW95, ABW07, CS11b, FRZ04, HLN07, Kah08, She02]. Chain [Din95, DF95a, DF95b, AV15, Ass97, ABL21, BYZ00, DD10, DHN00, DRZ16, Eth96, HR04, JKO09, Jia15, KKP14, KL02, LR12, MMPP07, MR02, MV06, QH21, TKH09, WH17, Yun98]. Chains [BK92, DS91, DSC93, DGLM10, Fil91, Gos01, JR02, Kar07, MT94a, MT94b, Mor92, NP95, Per94, Rog94, ATV15, BCKP99, Bax05, BBC17b, BK16b, BRS09, BN19, CCM06, CCHH05, CW03a, CHS17, CBM18, CvH10, CE10, CK07a, CK07b, CT01, CGL^{+}15, DD09, DSC96, DSC06, DS05, DMR04, DRS16, DW05a, Erh00, Fil98, FK13, FMMP08, FW99b, GHL03, Guy07, Han06a, HPP22, HKL^{+}19, Hub04, JSTV04, JJQ16, JY17, KZ09, KM13, KMPT10, Klo19, KR06, LP04, Lez98, LN13, MW98, MLNW22, PDG14, PD22, RR06, RR08b, SCZ10, SS19, TvH12, VK21, WW20, WL16, Wil04, YZ07, ZY96]. chainsaw [ABBH14]. Change [BJN18, FLÓ19, BBC23, DPS08, IM10]. Change-point [FLÓ19]. change-set [IM10]. changed [MAL14]. changes [CFY05]. changing [RZ23]. channel [Ala03, Aus08, HM09]. channels [CN11]. Chaos [DR09, MD01, Bar20, BE23, BL14, BM19b, CST22, CF17, CP18, CF16a, CF16b, DP19, DT18, HN22, HP20, JM08, JSW20, Lac22, MSSZ20, Tou14, Tou18, Xu18]. Character [Pin92]. characteristic [CX16, JT18, LP23, TK02, Wik01]. Characteristics [Jou02, ARL08]. Characterization [GdH93, KR14, Str05, BDM02, BFRT18, BF05, CT04, CCHH05, LM21, Sch21, Žit09]. Characterizations [BPT94, BR08]. characters [BBO15]. Cheeger [FW99b, MP20]. Chemical [Blo92, ADE18, LW19, MP14, PP15]. Chen [Xia97]. Chernoff [HK22, Lez98]. Chernoff-type [HK22, Lez98]. Chi [GPR17]. Chi-square [GPR17]. Choice [AGGL10, RR91, RR94, ASCDH09, BR01, MP22a]. Choice-memory [AGGL10]. choices [AGSC02, BBD22, LM05]. chromosomal [BP12]. Chromosome [LPS21, CCM06]. CIR [MAL14]. Circle [CGS93, CMP94, Arm10, LvZ04, May09]. Circuit [Ana91]. circulant [DM23b]. circular [AN22, AEK18, Woo12]. circulation [JJQ16]. circumscribed [BR04]. city [Ken11]. claim [KT03]. Claims [CK93, Sch92, ARS17, BF04b, CT04, Cha99, DM06b, JB07, Jan01, KK96, Pul14]. Class [KZ94, MZ91, Mor92, Alv03, ABP15, AB22, AMR04, Cal97, CPV23, CRV06, CK07a, CK07b, DJ12, DS05, FKM96, FJL18, GH21, GR09, HRS97, JMP21, JK14, MG05, Niu97, RU08, SZ06, Wal09]. classes [CL09b]. Classical [BS05, JKM15, Yuk96]. classification [CPV23, HNS22]. classifiers [CPV23, KPL03]. cleaning [BGBP23]. clock [RZ18, Žit05]. clocks [JZ11]. clones [FGLS23]. Closed [DH93a, LL23, SV94, SXY21, ASCDH09, CS16a, HW96, Kum00, LRM15, Lac20, MY96, WZ20]. closed-form [ASCDH09]. Closed-loop [LL23, SXY21, Lac20, WZ20]. Closing [RS91, RZ08]. closure [GK23, Kar13]. clouds [Owa18]. CLT [Ale96, HLN08, RR97a]. CLTs [LS20]. clues [Pit99]. Cluster [CR18, BL06, BLZ11, BZ15, BGV20, BF96, CAP20, Dar21, GG11, GJ18, JKM15, LPZ17, Tei09, Yat09]. cluster-index [Yat09]. Clustering [Vys08, Blo13, DS18, GSS13, JM15, Mån99]. Clusters [CDN02, Gu22]. Coagualtion [EW01]. Coagulation [DGM06, Arm10, Ber02, Fou22, HH18, Nor99, SSW06, Wag05]. coagulation-fragmentation [HH18]. Coalescence [BES04, FRT14, HLR22, LL13, SD05, SS15]. coalescent [BRSSJ19, BCKWB16, CPSJ22, DK15, DDSJ08, DK19, HJR20, Hén15, JS10, Ker12, LP13a, MR17, Nor99, SJ05]. coalescents [BBL14, DK19, DGP07, FRLS21, GIMM18, MP23]. Coalescing [HMT22]. coalitions [BEM08]. coefficient [Blo13, CG17, DFG22, KP04a, Mei09, MGY23]. Coefficients [XS92a, XS92b, BFO18, Bha99, Bro99, CJK19, De 11, DK08a, GM13, HHMG19, HL00, HJK12, JM15, LM06, QZ23, Roi07, RU08, Sab16, SXY21]. Coexistence [BEM07, BN97, CDL09, DN97, Dur09, DZ15, FS22, FLR22b, GM05, HN18, HK21, Hof05, CD06, DH13a]. coins [NP05]. Cointegrated [PR98b]. collaboratively [BCDS15]. collapse [HV97, KKLW09, KLS06, SW12, Sto04]. collapsing [BCC^{+}23]. collect [LvZ04]. Collision [Bha16, GIMM18, KMPT10]. collisionless [BF22a]. collisions [FSW15, IK10, Nor16]. colonies [Ber10]. color [BT22a, GLS18, Mat05]. colored [DAM10, FJL18]. coloring [BCOR16, Bha16, BDM17, TVVY12]. Colorings [FN93, FN94, DGJ06, HV06, SZ17]. colors [LPS21, LM02]. comb [BEZ20]. combinations [KPL03]. Combinatorial [Cra18, ABT00, Cao21, FH98, Hwa96, Hwa98, MP21, NR04, Yuk96]. combinatorics [BDM17]. coming [Ban19, BRSSJ19, Fou22]. Common [Ale94a, Ale95, Rhe95, BS19, BC22b, CS22, CK00, DG13, ER10, Jin19, KNRS22, LL23, MP22b]. Communication [CMP94, RR00]. communications [BS07, PGZ07]. Community [VAC15, BH20b, Cos16]. commutator [DFG22]. Commuting [ESU10]. compact [BCP18]. company [AM10]. Comparing [FW99b]. Comparison [DSC93, FK13, CM08, LMT12]. Comparisons [CE10, Fil13, LL17]. compartment [PP15]. compatibility [EHW16]. compensator [GZ08]. compensatory [PSW12]. Competing [Gla93, Shk11, BEM07, CD06, FS22, Gou07, Hof05, KS16, LL20b]. Competition [DS93, DN97, FMP09, LN06a, NP99, vdHHKR16]. competitive [BP97, BF02]. Complete [AG93, BG06a, ARS17, AK05c, BD07, CvH10, CP14, GK03, KS14, LV10]. complete-like [LV10]. completely [Ber97b, HK16]. Completeness [Bät99, AL17]. complex [BJM10a, BJM10b, CR22, FM18, Lac22, Ona08]. complexes [BKS17, FIMS22, MM03, YA15]. Complexity [YR23, DR13, HLR22, QH22]. component [BN17, CDG23, FM18, HR07, Jan08, JW21, Jos14, Pit08]. Components [MR94b, ABT00, Hol98]. Composition [BNT92]. Compound [Aal92, BC01, Erh00, Roo94, Gap05, LM21]. compressed [DM23b]. compressible [BFH20, KK04]. Computable [KM08, LMT96, MT94a, Bax05]. Computation [ACLW95, Pan08, TV03]. Computational [GO19, PY01]. computations [BBM07, MW07]. Computer [BE23]. Computer-assisted [BE23]. Computing [GK21, CL11, Mek15]. Concave [FKP94a, FKP94b, BW22b, FK99c, MS21]. concavity [MMS20]. Concentration [BB92, DR11, El 09, FRZ04, MR12, PS14, CH19, Klo19, LC22]. Condensation [KR19, MY96, DMM17]. Condition [LS95, YH93, Che96, DY18, Fou15, PZ19, Sas18, Sha15, YY18]. Conditional [AKLT22, AG09, FLP13, Gos01, NZ20, Pov95, BS14, BC14b, BLM23, Che08, GRS00, LPW08, LPW14, LX14, TvH12]. conditionally [KMK10b]. Conditioned [FMN^{+}16, BDW23, HSV11]. Conditions [Hol01, RWW95, WSH09, BGV20, CFS18, CE10, Cou08, DFMS04, DMO14, ET11, Fed14, Gri16, KS03a, KLS11, LRM15, Las02, Las04, Mül18, PD22, SS18]. conductance [BZ10, CKW21]. Cone [CS13b]. Cone-constrained [CS13b]. Confidence [DM92]. configuration [AGvdHdH18, BN17, Bal21, BR19, RZ23, vdHvLS21, vdHOC18]. configurations [Cou10, DH07, DFPR22]. confined [CGZ14]. conic [GNP17]. Conjecture [DMY95, HS19a, Yu17b]. Connected [CH91]. connection [Fis17, LNS21, Mån99]. connections [BDM17, JSW20, MR06a]. Connectivity [DDM11, ELM^{+}16, Pen16, BEMT21, CW99, DDMT12]. consensus [BG21]. Conservation [Gos01, Jou02, BJ22, DWZZ20]. Conservative [Myk00, RYZ21, BGJ18, Hey23, Mor05]. Conserving [Che95]. considerations [Alb09]. Consistency [BBD99, DS18, Mas95, San10, BGHK20, Nor16, Oha09]. Consistent [GRS08, Ser94, JS96]. Consol [DMY95]. Constant [AB92, Jai93, vdBK93, BGT22, BH99, Bla96, BP97, Bro99, CC98a, DHL23, FKP23, Hub15, Mar02, PY98, Tra23, XS92b]. constants [JSTV04, SS19, Wu09]. Constrained [CK92, CK93, HSX22, Blo15, CS13b, DG14, GTZ20, GS11, HM09, Käl22, LW19, MT13, Sha20]. constraint [Sch13]. constraints [BdHNT22, BR06, CTZ04, EPQ01, KK96, KX22, LŽ13, Sha20]. constructed [DJ12]. Construction [CC93, CW13, FL22, CG17, CFF02, CDE^{+}17]. constructions [RR97a]. Consumption [HP92, SS94, XS92a, XS92b, BR01, BP05b, DK08a, Mon22, MS20, Yu15, Yu17a]. Contact [BLZ11, LS17a, BZ15, CD06, CDL09, FEvdD16, GM12a, JLM19, Kro99, Rem08, vdB11]. contact-processes [BZ15]. contagion [BLY22, DRST09]. Contagious [FKR17, AK18]. contamination [KLS97]. content [MN97]. Contents [Ano16b, Ano23c, Ano23d]. contests [FH16]. contingency [Bla09]. Contingent [CK93, ARS17, CT04, Cha99, DM06b, Jan01, KK96, KT03]. Continuity [PW10, Pes19, Ala03, KRSW19]. Continuous [BK15a, BD19a, BS12, BCL06, BZ91, DS95, DR91, EK94, GM08, GHL03, HSZ17, KS92, LZ06, MS21, Mas95, RR91, RR94, AS97, Ass97, BBC23, BDMT11, CFY05, CF11, CS13b, DW05b, FL22, GK00, GIO^{+}17, GR06, GS11, HJR20, HJK12, IRR12, ILP15, JKM15, JJQ16, KM98, Kle03, KS05, LYZ19, LM05, MP98, Pia06, PS17b, PSZ14, PY16, PVY20, RW12, Yan05, vEK08]. continuous-discrete [KS05]. continuous-state [FL22, ILP15, LYZ19]. Continuous-time [BK15a, BD19a, BS12, BCL06, GHL03, LZ06, CS13b, GR06, GS11, HJR20, JJQ16, Pia06, PS17b]. Continuum [BBW04, Cla22, Pen93, Pen96, Ale96, BBFM03, BDW22, Gou09]. contraction [NRS21]. Control [AB14, DZ94, DY95, DO94, HS93b, HS93a, KO92, KS99a, KLST93, RT92, Sto03, AC20, ABP15, AP16, AB22, AHS05, Ata05a, Ata05b, AC17, BVLT20, BCFP18, BK17a, BB20c, BBC19, BET05, BMR08, BLM17, BG05b, BR06, BR08, CL22, CST05, CC16, CS23, Cla18, CFJ16, CGK^{+}23, DPRZ19, FG18, FS02, FPRW18, FP15, GHLT14, GHK11, HV97, HH08, HP23, HMGR00, HSX22, IS04, JMP21, Jac02, Jaś07, JMSS12, Lef04, Mey06, Sch13, SXY21, Tan14, VPV08]. control-dependent [Lef04]. controllability [AMS06b, Sha15]. controllable [GHR03]. Controlled [BQ19, BG12, ATV15, AZ16, ABW07, BGZ97, BT19, BMN14, BG06b, BR13, BF18, CCHH05, CBM18, CC16, GHL03, IH23, YZ19]. controller [HHSZ15]. controls [BR06, CDGL22, Dje23, LRZ06, MP22b, SXY21, WZ20]. converge [CK03]. Convergence [Ale93, Ale94b, Ale94a, Ale95, AV15, Asm92b, BJM10a, BT96, BG12, CL22, Che01, CDMR12, DN94, Dar23, DM94, DF95b, FFS93, Fil91, FP95a, HP92, IR01, Ing94, JR02, JLZ21, KRSS23, Kes93, KM13, KT04, KL02, LdRS18, Mal03, Már97, ML16, MT94a, MT94b, MP06, MP95, NST20, Owa22, Pov95, RY94, Rhe95, Ros93, SX23, Wan23, AKH02, BB20b, BB23, BJM10b, Bax05, Béd07, Bel11, BF22a, BJKT16, BD07, BREZ20, BH17, But14, CW03a, Chi07, CJK19, CP14, DGL23, DPBCD21, Dje23, DKP22, DGR09, DR10, Dol10, DFMS04, DMR04, Duf16, DW05b, DM17, ELM22, FS16, Fil13, Fon10, GG13a, GvdHL20, Gau98, GGR97, Ger11, GG97, GH08, GK16, GK07b, HMY21, Hey19, HLN16, HPS03, HJK12, Jon97, Jor02, KZ09, LR19, Lac20]. convergence [LL23, LRW23, LMT12, LY17, Li18, LW22, LMT96, MR02, MNG09, MR06a, MMN22, Num00, Oli09, PP12, Pel98, Pèn05, Pit99, PW04, QH21, QH22, RR98a, RR98b, RS23, Yos12]. Convergences [Che22b]. Convergent [LC03]. converging [Die15, Jia15]. conversion [FS22]. Convex [BK17a, CY21, CK92, Hsi94, Lei08, vM95, BM22, BR04, Con22, KPL03, LŽ13, LR18]. Convexity [Jia08, ET09]. Convolution [Ber92, Gri13, Gri16, JCLP21]. convolutions [MRW18, Yu17b]. cooling [CC98a]. cooperation [Che13]. cooperative [BH99, BF12, SS15]. Coordinate [Fis96, RT92, EPZ20, EP22]. coordinates [LT18]. copolymers [BG05a, CGZ07, Ton08]. copulas [LP17b]. core [DM08b, JL08]. cores [AS23]. corporation [CTZ04]. corrected [BG06a, GLM18]. Correction [Ale95, Ana94, BH03b, Bra94a, CD03, Dur99a, FN94, FKP94a, Har03b, Har06, LS19, PS22, Web01a, Der11]. Corrections [ABT^{+}11a, NV06]. corrector [Dar21, MN17]. Correlated [CDFO13, RS22, AZ14b, Bud02, FK16, JSW19, Löw98, MRV16]. Correlation [HPP22, Ser94, BBD99, CW99, El 09, Jia04, Jia19, LR06, RZ08, WXY23, YM22]. correlations [CR22, EKR23, Häg99, MS11, OC21]. Corridor [Fus00]. Corrigendum [LO17, MPZ21, LP17a]. Cost [Kus95, TY93, BF10, CCHH05, Che13, DS15, HSZ17, Jel99, Pem09]. Costs [DPT01, DZ01, Kus95, SS94, SSC95, vM95, ADGS98, BDG16, BD19b, BT00, BT13, CS16b, Dol13, Gua02, GRS08, JB07, KMK10a, LS97, Wee98, Yu17a]. Coulomb [BCF18, BGZ22, Tur18]. countable [BT13, TSM19]. countably [BL08, BH19]. counter [Taf11]. counterexample [AS16b, Hey23]. Counterexamples [GW97, VK21, BK00b, DI10]. counterparty [DP15].

Counting [BG01, dLS97]. counts [Erh00, FLR22a, Gol13, YS96]. Coupled [GS09, BZ08, BFS21, EKR23]. Coupling [Asm92b, BREZ20, HV06, Jag97, BBL14, BF22a, ČT16, Hob98, HN19, MGY23, RR97a]. couplings [BG96, BK00b, BKS20, CRX21, Röl07]. covariance [BLW11, BH13, BGBP23, DY18, FJ22, HMY21, HLS19b, JWB^{+}14, LS16a, NY16, PY14, SX23, WJB^{+}15]. Covariances [VG95]. Cover [Bel12, KS14]. coverage [GP10, Pia05]. Covering [BBFM03, FP93]. Cox [CW93, DGC20, FSW95]. crack [BK98]. Cramér [BH22, DM05, Gri16, Pau02, SZZ21]. Cramér-like [Pau02]. Cramér-type [SZZ21]. Cramér’s [DG18]. Crank [HLN21b]. credit [BJR08, BLY22, ÇJPY04, DP15]. Crested [DD09]. crisscross [BG05b]. Criteria [MT94b]. criterion [Dai96, HSZ17, JR14]. Critical [BU18, Cer15, DS93, Fox16, HH19, LV03, MM03, AZ10, BR13, Bur07, Chi04, CS21, Cla22, CW19, FW22, GS23, ILP15, Jos14, KR19, Maj06, MZ05, Pop04, PW04, RŞ18]. criticality [CP19a, EMO10]. Critically [AS09]. cross [BGBP23, JWB^{+}14, WJB^{+}15]. cross-covariance [BGBP23]. Crossing [BCS19, Loa92, PW23, CL03, CCCS11, Kah08, LSZ97, LS18, MZ05, Yu17b]. crossing-over [LSZ97]. Crossings [ASG93, BD12, AS97]. Crossover [BdHNT22]. CRT [BM13]. Crump [LSZ13]. Crystal [GW93, AMS06a, GL18]. Cubature [CM19, CM14, LL12]. Cube [FP93, Jai93, RT92, Mar16]. Cucker [CDP18]. culture [Lan12]. Curie [BR17a, CS11a, FMP00]. Curie-Weiss [BR17a]. currents [BL12b, Sas18]. curse [RvH15]. curtain [HN19]. curvature [BD20, FM16, MS18, RSM09]. curvatures [MV16]. curve [FV21]. Customer [Ngu94]. CUSUM [ZRH15]. cut [BM13, DSC06, Die15, JKS18]. cut-offs [DSC06]. cut-tree [BM13]. Cutoff [BCMR21, CS21, CPS16, NN19, OT23, BH20b, Fou00, LL20a, Tra23]. cutoffs [CHS17]. cuts [MP20]. Cutting [ABBH14]. Cycle [JJQ16, BUV11]. cycles [BBK^{+}11, BSZ20, BKS17, CC98b, HH19]. Cyclic [BS23, PR94, CMT21, HL21, MNP14, NN19]. cylinder [BEMT21, Win08].

d [AW05, vEK08, BD20, BNS13, FHLN22, LT22]. damping [CLP16]. Darwinian [Leh23]. Dassios [ERY95]. Data [AG93, Cha93, CG92, GN91, Ros93, ASJ20, CLW16, DR98, JT18, OQR16, RS98, RS03b]. Dawson [Sch13]. deadline [DLS01, KLSY04, KLS06]. deadlines [GK07a]. Death [CW93, YH93, BBL^{+}97, CA18, Che18, DSC06, Erh00, ESU10, KST04, Lan15, Web01a, Web01b]. Decay [BLP13, Ber97b, BCP^{+}22, GQ03, HPP22, Rhe00, Zer98]. decentralised [SS18]. deciding [GK09]. Decision [Gla93, GR06, GS11, JWW11, MP22b]. Decisions [DZ01]. decomposable [JSTV04]. Decomposing [JR92]. Decomposition [FMP95, GGLO13, GM12b, JKP13, MR02]. Decompositions [JSW19, HPŠV04]. decoupling [AFKP20, FI20]. decreasing [MG05, Tri15]. deep [Gen23]. Default [GSS13, BJR08, EEH14, MAL14]. defaultable [DSS96]. defaults [JKP13]. Degenerate [DP09, KO92, BFO18, FRT03, Zha05]. Degree [BGvdHK15, BM17, Blo13, PRR13, AL05, CDS11, CS13a, Gol13, GLM18, Jan08, Jos14, OC21, Pit08, RW97]. degree-corrected [GLM18]. degrees [BvdHH10, BvdHH17, BS22, DS22]. Delaunay [EN18]. Delay [RS01, vM95, GTZ20, MSSZ20, Sto03]. delayed [DI10]. deleted [RR97b]. Deletions [AW94, Zha95]. delocalized [AZ14a, HK21]. delta [Myk00]. demand [Tou00]. Demands [BZ91]. demes [HP20]. Dense [AR16, RZ23, AABR22, BS17a, Ste99]. densest [AS16a]. Densities [Ber92, BRTP18, BHdS^{+}20, De 11, LRS17, NRY12, VK21]. Density [Dev92b, KZ94, BMJ06, EF21, FK99b, JB07, LPP15]. Density-Dependent [KZ94]. Departures [GW91, MP95]. Dependence [MS93a, BB20b, BJN18, BKT22, CSS22, EHW16, HRS99, Jag97, Jag99, KS96, LS09]. dependencies [Olo96]. Dependent [Aka95, Das95, GSvdB98, HS93b, HS93a, Kel93, KLR91, KZ94, KLS95, MT94b, Yam95, AS10, BPR22, BGR18, BCFP18, BHQ17, BT22b, BD15, CX97, CS18, Col09, CGK^{+}23, DKP22, EK09, FL96, GH21, GL21, HL00, IT12, Lef04, LT18, MP98, MS00, O’N97, PP08a, PTZ20, Ren16, RS98, RR97a, Roi07, Tan14, VZ19]. depinning [AZ10]. Deposition [GW22, PY02, PS05]. depth [GS05]. Derivation [BFS21]. derivatives [CDET13, GL21, JNP21, LR18]. derived [CK00, DGP07, LPSX23, LPW08, LPW14, YR23]. Derrida [BZ18]. descent [BCH22]. description [FM04]. designs [Zha16a]. detect [JKK03]. Detecting [Com97, CZ18]. detection [BJN18, DS15, DPS08, ELS17, EPZ20, EP22, FLÓ19, Guy07, IM10, JP17, PGZ07, Pes14, Sta15, VAC15]. Determinant [Jia19, AJvM22]. determinantal [BH20a, TBUA23]. Determinants [Vit91]. Determining [Cha97]. Deterministic [DY95, HLN07, ADNR21, Aus08, BCH22, BR17a, BKR22, CBM^{+}21, Cos16, CDMR12, FHLN22, LW22, dSDG10]. Deviation [Ath94, AV95, Mor92, PDG14, AB14, AC17, BFG13, BdHM23, CGM09, DS22, DAM10, DWZZ20, DR98, FH98, Kar07, LLLZ18, Mey06, NV03, PS17b, PW97, SZZ21]. Deviations [IR01, RS01, Too02, ADE18, AH98a, AH98b, ABDW21, BPR22, BMRS23, BBRZ20, Ben96, BPT98, BNS21, BL12b, BG05a, BGL02, BG05b, BFW21, CT11, CP19b, Chi05, Chi07, Chi15, CS13a, CZ03, CN23, DGC20, Dd04, DS05, DGN05, DW98, Fen07, FG08, FM05b, FN17, GHR03, GW97, GO00, Hei05, HIP06, HIP08, Hey23, HLMS05, Hwa96, Hwa98, IR00, JKM15, Maj06, MvU05, Nag12, NV04, NV06, Pap98, Pap00, PDG14, PPZ22, Pia99, PY98, Rey18, Sad96, Wis01, Yan20, YS22, Zhu15]. Diaconis [Yu17b]. diagonal [HMY21]. diagram [Owa22]. diagrams [HST18, ST98]. diameter [AL15]. diamond [BCJ22, CJY15, Cla22, FV21]. dice [MLNW22]. Dictator [PPM18]. Difference [Gre94, Xia97]. different [BČ05]. Differentiability [IRR12, KS06a]. Differential [Ant93, Fin94, HL01, KX95, MPST02, Wor95, BNT19, BCD17, BdRGL20, BS13, BLM23, CDET13, CC14, CST22, Che22b, CKHL06, CE10, CFJ16, DI10, DZ16, DEH23, DMP96, EKR23, FP22, GLW05, HLTT17, HMGR00, HNS11, HLN21b, HJK13, LS15, Laz04, LØP04, MZ02, MPZ13, MPZ21, MG02, NY23, Pan08, PTZ20, QX18, QZ23, Ren16, See20, Sei09, Zha12, Zha16b, Zha16c].

Diffusion [Blo92, BK95, BL01, BG06b, BF18, CV21, Çet18, EP10, EK92, FP18, Gro04, Gur14, KX95, KO92, KLR91, Kot92, KLST93, MPS12, Puh15, Yam95, ZDZ11, ADGS98, AJKH14, AAK17, AP16, Ata05a, AG14, BM19a, BM20, BEG00, BS05, Bha99, BG06a, Blo15, BL12b, BD17, Cer09, CL07a, CFY05, Cla18, CG17, CJK19, CDE^{+}17, CL09b, DHT10, DG14, FK99a, Gau98, GW22, GHH07, HSZ17, HLT21, JS17, KKLW09, KKP14, KS99a, Kel16, KL04, KL02, Kum00, Le15, Lef04, LM06, LdRS15, MP98, Már97, MT99, PP12, Per00, PST12, PY16, PVY20, RR03, RR14, Sab16, Web01a, Web01b, YY18, Zha16c, dBG12, dSY05, CLP16]. Diffusion-approximation [CV21]. Diffusions [HHMS93, Alv03, AB22, ABW07, BPR22, BC02, BGR18, BS14, BR05, BDKR19, CCCS11, Che18, CK03, CLSF18, De 11, DL10, DFM16, DP09, FRSW22, Goe06b, GDVM19, HSV11, HL17, HI05, HIP06, HIP08, HLN16, HP20, HHMS05, JR16, KR14, LST23, LS14, LS16b, MSSZ20, Pes19, RWF13, RU08]. diffusive [RS23]. Diffusivity [CF09]. digital [DG07]. dilute [Nor16]. diluted [CP18]. Dimension [AN22, BB23, HR94, Kar10b, Ser94, BB20b, BČ05, BBD99, CGK^{+}23, DPBCD21, GL18, JPS^{+}22, LLS08, PZ11]. Dimension-free [BB23, DPBCD21]. Dimensional [AGP95, CMP94, FMP95, GdH93, INPY13, Jou02, Kot95, Pov95, AJKH14, AAK17, AJO14, APP08, BC02, BL12a, BJ22, BKR22, BS17b, BDKR19, BR15, BGZ22, CDN02, CF09, CX97, CCR23, CG17, CGR09, CP08, Cox10, DH23, DI17, DR08, FRT14, FK21, Fon10, Fou00, FJ17, GS14, GHP13, GLM17, GS20, GZZ15, HL21, HNS22, HS21b, HJ23, Kab12, LS13, LM06, LC22, MM03, Mél00, Mor05, PP08a, PW21, Pes19, QS23a, QS23b, RR06, RYZ21, Roi07, SYY20, Tom21, Tur18, Vys08, YS96, YZ19, YM22, vdB11, vdBN17]. dimensionality [KPL03, RvH15]. Dimensions [BL01, HW92b, Tal92, BT05, BRS09, BCJ14, BDH10, CCK23, DRVZ17, Ebe14, GHS21, HSV14, LM02, MPS12, Pen96, PST12]. dioecious [Yu07a]. Diophantine [APW08]. diploid [LN09]. Dirac [CH19]. Directed [LC93, Viv23, Alb09, BOC22, Cla22, DFK12, FKP23, FJ19, GS23, vdHOC18, CCR23]. Directional [Zer98, LR18, NW22]. Dirichlet [ABF13, AZ14b, CDE^{+}17, DF06, DGM06, Fen07, FG08, FRSW22, GR21, GMŚ19, JLP08, JKK02]. disasters [FJ19]. discarding [MW13]. discipline [DLS01, Sto03]. Disciplines [Che95]. disconnection [Win08]. discontinuous [AH98b, ESX22, LM06, LS17b, LS19, MGY23, NRY12, NY23, RU08]. discount [KR96]. Discounted [GS11, GW00]. discovering [DR10]. Discrepancy [DRZ16, GH21, BDKT20]. Discrete [CEK12, FMP95, GN91, HPTvD95, JY17, KLP15, Lig95, Mag00, PR94, Tan14, AJvM22, AK05c, BCKP99, Bha16, BR15, BC18, BP05b, BN15b, BK00a, BG03, Bud02, CL07a, DH23, DW05b, ER08, FM16, FZ02, GOP03, GJKS15, GO12, Har98, JJQ16, KM98, KMPT10, KP96, KP98, KS05, LP04, Mar08, MG04, QS23b, RS05, SS18, Win08]. Discrete-review [Mag00, Har98]. Discrete-Time [HPTvD95, PR94, CEK12, Tan14, BP05b, BN15b, GOP03, JJQ16, RS05, SS18]. Discretionary [DZ94, HHSZ15]. Discretization [AGP95, Fuk11, AA13, BZ08, CM14, LdRS15, RT14]. discrets [Mic96]. Discussion [JLZ21]. disease [CS99a]. Disk [Hsi94]. Disorder [CP18, GL18, BDK06, BMRS23, BvdH12, BM19b, ELS17, Gap05, KQ18, Sez10, ZRH15]. Disordered [Ton08, AZ14a, CSZ17, Cla22, GKS04, Lac17, LS14, MZ14]. Dispersion [CS02, ABT^{+}11a, ABT^{+}11b, BDKT20]. dissemination [Lan12]. dissipation [BFM11, HS21a]. Distance [Ale94b, JK22, MR17, BBBW22, BvdH12, BNT19, Clé23, JJ15, SS06, Wan23, Xu19]. Distance-based [MR17]. Distance-Minimizing [Ale94b]. Distances [LRS17, CM05, FJR20, MN03, vdHOC18]. distortion [Coh04, DK99a, XZ13]. distributed [BHP10, KP96, KP98, YS96]. Distribution [Aal92, BK01, Che01, Das95, EK92, Hsi94, JKK02, Kab12, MS91, MN03, MR93, RT15, Tak93, AR20, APP15, AW05, BH22, BGvdHK15, BM04, BMM21, BFJ06, Bor16, BK00a, BC14b, CM05, CFJ00, DF06, DGLM10, EEH14, Fen07, FG08, FMS96, Fla97, GR21, HV06, HLN16, JJ15, Jan08, Jel99, Jia15, JWB^{+}14, Käl22, Ker12, KLZ98, KP04a, Kuz10, LRdH98, LS16a, LLS08, PP12, PW96, PTZ17, Sad98, SSV18, TK02, TT14, Yat09, Yu07a]. distribution-constrained [Käl22]. Distribution-valued [RT15, SSV18]. Distributional [Fil13, KL91, TSM19, AB05, PS97]. Distributions [Ald91, Ber94, CG92, CS95, FKP94a, FKP94b, JLP08, MS93a, MS93b, O’C91, ARL08, Asm98, BBP22, Béd07, BCP18, BN19, CPS11, CR18, CFS18, DJ10, DFPR22, DGM06, El 09, FS14, FKM96, FRSW22, FK99c, FW22, GW09, GS05, GK07b, Hub15, Hwa96, Hwa98, JKM15, JS12, Jia04, KR12a, KR14, Lac03, LR13, LX14, MP99, MS21, ML16, PGZ07, RR19, RR98a, SYY20, TvH12, WSH09, YY22, Yu07b]. Divergence [HJK13, ABDW21, DKPRB20, Feh23]. divergence-free [Feh23]. divergent [Bra11a]. Diverse [KS16]. Diversity [Ngu94, BHP10, DF20, RWF13]. dividend [APP07, APP15, AM10, BMX17, CTZ04, DE17, DFT03, LV03, Loe08, VPV08]. dividend-paying [DFT03, LV03]. dividing [Ban08]. divisible [AS04]. DNA [Lac03]. Do [JT10]. document [GW09]. Does [Sel95, FW99b]. domain [CT11]. domains [CF17, CI11, DHESC21, FV21, KW07, LR18, MW07, SV10, TK02]. dominated [GKS04]. domination [MN97]. domino [CJY15]. Donsker [FK16]. Donsker-Varadhan [FK16]. Doob [ABP^{+}13]. double [BG02, CST05, FRT14, MPdS19]. double-exponential [MPdS19]. double-well [BG02]. doubly [CM08, ELS17]. doubly-stochastic [ELS17]. Dovetail [BD92, Ciu98]. down [ABBH14, Ban19, BRSSJ19, Fou22, HNS10]. down-side [HNS10]. Downside [Nag12]. drainage [GRS04, RSS16]. drawdowns [Mei09]. Drift [AHS05, Cha92, Das95, ERY95, Pov95, DFMS04, EP22, FG20, Feh23, FZ03, GTW22, HI05, HIP06, LS17b, MS00, MGY23, QH21, Zha16c, LS19]. drifting [EPZ20]. drifts [PP08a]. Driven [KX95, KM95, APS19, AR16, BS96, BL12b, BMR08, Cha99, Clé23, CFJ16, DH04, Der11, DL16, FP22, GLM17, HLN21b, LT19, LØP04, MSSZ20, MR08, Pan08, TKH09, Yan05, Zha12, Zha16b, DK08a]. drivers [LdRS15, Ric11]. droplets [Wüt06]. Dual [BTZ04, DPT01, STZ13, Bel13, BT22b, QS23b, Žit09]. Duality [CS18, CK92, CS16b, Mon22, XS92a, XS92b, BK17a, BD19a, BN15b, BFM17, CGR20, CKHL06, CHS22, DGM06, EPS09]. Dubins [MP01, CDS09]. Dubins-freedman [MP01]. duplication [DP09, HLR22]. Duration [dGvZ93, FMS96, Sud08]. during [SD05]. dyadic [BFM11]. Dynamic [Ana91, BGZ97, BW01, Cra16, DW05a, DSW07, EK94, GM95, HV97, HS93b, KMZ17, MS05, MR10, NP02, RR94, vM95, ATV15, Ass97, AGvdHdH18, BBT17, BdHM23, DRST09, EKT07, EPQ01, IH23, Käl22, KM11, OT22, PS17b, Rem08, San10]. Dynamical [FIMS22, Ser94, BJ20, FS14, RŞ18]. Dynamics [BCF18, ER10, FL13, MSZ19, BBC17a, BHQ17, BF12, BGV20, BCP^{+}22, CL07a, CMT21, CF11, Cop22, DLZ21, DGP12, DdHJN17, FIKP98, FMP00, GJ18, HS07, HNS22, HI09, HX22, Imh05, ILP15, KM11, Leh23, LS22, MS21, NY23, OT23, RR19, RZ23, SZ17, TVVY12]. Dynkin [BY17, DMP22, FG18].

each [FJ22]. Earliest [KLSY04, DLS01, KLS06]. Earliest-deadline-first [KLSY04, DLS01, KLS06]. earthworms [BBPS15]. Ecological [BPZ07, HNS22]. Economic [Fin94, Col09]. economies [Num00]. Eden [AP17]. EDF [KLRS11]. edge [Bor16, DY18, FJ22, JM12, KS14, LNS21, Pen97, RZ23]. edge-changing [RZ23]. edge-marked [LNS21]. edges [CAP20]. Edgeworth [KMS17, PY16, PVY20]. Editorial [Ano16a, Ano23a, Ano23b]. Effect [BP12, Bor12, Sch05a]. Effective [ABBL09, Klo19, SJ05]. efficiency [Atc10, LZ06]. Efficient [ABL12, BG08, Bla09, BK00b, CL07b, DG95, JKW11, Gu22, LX14]. eigenfunctions [KZ09]. eigenmatrices [BLW11]. Eigenvalue [BP18, Fil91, Mos01, BB20a, Bor16, CN11, FW99b, HK17, He20, LS16a, PTZ17, SX23, SYY20, YY22]. Eigenvalues [DS91, CJL21, Chi16, GS18, KPT^{+}16, Ona08, WJB^{+}15, YM22]. Eigenvector [CR22]. eigenvectors [YM22]. Einstein [MZ14]. elastic [Nor16]. elasticity [KS99b]. election [FMS96]. electrical [PS16]. Elementary [JSTV04, CFJ16]. Elimination [CW16]. ellipse [MN97]. Elliptic [Pap98, BFO18, EKR23, GO12, Gu22, LLLZ18, Ren16]. elliptical [El 09]. Embedded [QS94, BMJ06, DJ12]. Embedding [JM02, HK13, KTT17]. embeddings [CHO08, DGPR19]. Emergence [JW21, Aus08, CMT21, MM10]. Emery [EOT05, JY17, Kar13]. emigration [Ber10]. emissions [CDET13]. Empirical [Eic95, MS93a, TSM19, Bel13, BH01, BH03b, BNS21, BL12b, BGL02, DAM10, Hei05, JT18, MRV21, ML16, TT14, Wan23]. encountering [ARL08]. end [Sto03]. end-to-end [Sto03]. endowment [Žit05]. endowments [Bar19, HK04, Yu17a]. energy [BK06, Gen23, Hey23, JNP21, KP04b, NS22, QS23a, Yos12]. enlargements [LR14]. enough [LL17]. enriched [WPRS21]. ensemble [BB20a, CR22, DH23, LP23, PTZ17, RS14, DT18]. ensembles [AN22, FT17, TBUA23, Tur18]. enterprise [MLNW22]. entries [HMY21, Jia04, LR06]. Entropic [FM16]. Entropy [BT19, BCP^{+}22, HS21a, Fou15, GQ03, Lei08, RS06]. Entropy-controlled [BT19]. enumeration [Alb09]. envelope [BD15]. Environment [CJ94, BEM08, CDL17, Col09, ESTZ13, GM12a, Goe06b, GLP23, GK23, KR23, KK04, LN06a, LPSX23, LS22, Nak11, RS03a, Rem08, WM04, YS22, Yos08]. environmental [CF16a, GL23]. environments [BMRS23, BL16b, CZ18, Coh96, FL22, Jon97, PPZ22]. Epidemic [Rei95, And98, BBT17, BL08, BFHM20, Cla96, DR09, FP22, O’N97, PP22]. Epidemics [BMST97, BST14, DN91, Ish93, Bal21, FLR22b, GSvdB98, LL20b, Nea06]. Equality [AZ10, Yos12]. Equation [CC93, EW01, HW94, Kot95, AKH02, Arm10, BBC17a, Bal19, BFS21, Ber02, BT96, BM19b, BS13, BLM23, BKT22, CV21, CDS23, Çet18, CC98b, CFJ16, DH18, DGR09, DR10, ET09, ET11, Fer15, Fou00, Fou15, FJ17, FJR20, GR08, GMO15, Gau98, GT19, Gué03, Haa10, HL17, HLS19a, HHK20, Hey23, HKV20, Kli19, Mél00, MM01, Muk22, Nor99, Nut12, Pan08, RYZ21, Xu18]. Equations [Ant93, Cho02, Fin94, Gol91, Gre94, HL01, KX95, MPST02, Wor95, AB05, ABW07, Aus08, BNT19, BCD17, BdRGL20, CC99, CDET13, Cer09, CC14, Che22b, Che22a, Cho06, Cho09, CKHL06, CE10, Col09, CDV14, CI11, DD10, DI10, DZ16, DEH23, DGPR19, DMP96, EW03, EGK22, EKR23, FHLN22, Fon10, FJL18, FP22, GS09, GO12, GLW05, GMŚ19, HLTT17, HMGR00, HNS11, HLN21b, HJK13, JMP21, JCLP21, KY10, KLP15, KL02, KNRS22, LS15, Loe18, LØP04, MZ02, Mal03, MPZ13, MPZ21, Mor05, MR08, Mou20, MG02, QX18, QZ23, Ren16, See20, Sei09, TV03, WZ20, Zha12, Zha16b, Zha16c, dRST19]. Equidistribution [MP22a]. equilateral [CS16a]. equilibria [DFM18, Dje23, Lac20, Num00]. Equilibrium [Haa05, Rey18, SSY19, BPG19, BC21, BCPG22, BGZ97, BF22a, BPZ07, BBD22, BLP13, ÇD16, DDM11, DGR09, DR10, ESX22, GO00, LY17, Mal03, Oli09]. equity [BFK05]. Equivalence [Asm92a, CW93, FT17, TK02]. Equivalent [CFY05, Gri13, Gri16, HV97, LM21]. equivalents [HLN07]. Erdos [BdHM23, Cao21, Gol13, KKM06, RŞ18]. Ergodic [Ana93, Ana94, ABP15, AP16, CBM18, HW07, KR21, PP12, Yca93, Yuk96, AS16b, Bax05, BFO18, BC22a, BKS20, CT01, DL18a, DRZ16, Gur14, KM03, LST23, PD22, RS03a, SSV18, TV12]. Ergodicity [AR19a, APS19, BBC17b, BRZ19, CT91, TvH12, AM06, Ber97b, CvH10, DGM20, FR05, FJR20, KR06, SZ06, SV10]. Erlang [MW94]. Erratum [BCPG22, BvdHH17, HL22, Tou18]. Error [AGK11, AGP95, DK08c, Ebe14, Kif06a, Kif06b, Lam98, LR13, PRW95, BJ22, Cer15, CG17, FK22, Fuk11, GO12, HHMG19, HJ23, HLN16, KPL03, MWZ07, MGY23, OC11, Yan05].

Errors [BK01, BJR16]. escape [FMN^{+}16]. Escaping [DSZ15]. Esscher [EPS09]. Esseen [DT05, DF20, Gol13, HK22, Klo19, LRP17, LS09, LLS08]. Establishing [AV16]. estimate [DM05, DG18, GO12, HS99, LW22, NRS21, Nor16]. Estimates [DM94, CN23, CM08, Dar21, De 11, DHESC21, DK08c, HS21a, HJ23, Kif06a, Kif06b, Lam98, MP09, MW08]. Estimating [AG93, Ass97, BBBW22, ES02, KPS98, RS91, RZ08, ADRS23]. Estimation [CLP16, CG92, Ser94, AR21, BJR16, CPV23, CCL16, DFdH04, DG21, EFT07, FKR96, FL03, GS14, HLR22, HKL^{+}19, JT18, KM08, LS18, MWZ07, MR17, RS98, Sad96, TT11, Wu09, ZHC06]. estimation-adjusted [ZHC06]. Estimator [BG95, Mas95, BBD99, HL00, San10]. estimators [Chi04, DFM16, FJ22, FRZ04, LPW08, LPW14]. Eternal [Ber02]. Euclidean [Ale94b, Ale96, Jai93, Lee97, Pen96, RY94, Rhe93]. Euler [AJKH14, BK15b, CC93, CC99, CX16, CG17, DL16, FG20, HLN16, HJK13, LT19, Mal03, NT20, PVY20, Sab16, TK02]. European [CKL22, JM02]. evaluated [KPS98]. evaluation [CL07b]. event [BSZ15, BBDW22, BG08, BDS20, CDV14, EK09]. Events [HPTvD95, NP95, BFW21, DG05, DKPRB20, GSS13, GS19]. Evolution [CL07a, Rei95, BBC23, Che13, FKKM18, Leh23, MP22a, Swa18]. Evolutionarily [HRW08]. Evolutionary [AGD94, Lan15, BST04, CMT21, Che18, KL04, MD18]. evolutions [JK22]. Evolving [BS17a, JLM19, Shk11]. Ewens [ABT92, DKT91, Fen07, GR21, GS02, HNNZ13, Jia15, JKK02, Tsu19]. Exact [AA13, APP08, BR05, CGR20, Che01, DM92, FM05a, JS17, KPR10, OQR16, PS05, ZBM04, Ala03, AV16, Chi07, FM01, JS12, MMX21]. examinations [Ber96]. Examples [DH13a, BK00b, Taf11]. exceedance [CL07b]. Exceedances [DKZ94]. exceeding [FPZ05]. Exceptional [RŞ18, vdBN17]. excessive [Alv03]. exchange [Ass98, CK19]. exchangeability [FHH23]. Exchangeable [DGP07, FRLS21, BFRT18, CS11a, FK22, FLP13, Fou22, GK96, Röl07]. excitable [GLS18]. exciting [HX22, KS96]. Exclusion [FR92, Fil91, BC16, CM03, FN17, GNS23, GLM17, GKS04, HMT22, LL20a, Mor06, RS15]. excursion [Adl00, CX16, LSZ13, MV16]. excursions [ABL12, LR19, TW07]. exhaustive [FL96, HDP22, MMPV06]. exhibits [BČ05]. Existence [AC03, BEMT21, BR06, CCCS11, CK12, DS95, DFFN21, DS07a, Fou00, HK23, Hol98, HP92, TY11, HM14, MM01, Num00]. Exit [AKP04, APP08, KZ94, APT20, BPG19, BC21, BCPG22, CP19b, CC98b, HI05, PW21]. exotic [GL21]. expanding [DF16, LY22]. expansion [BEG00, Bor16, BL14, CDGL22, GZ08, KT03, PY16, PVY20, TY16]. expansions [BS96, BH00, CL17, DPR09, KMS17, Kne00, LPP15, MW98]. expectation [NZ15]. expectations [ADRS23, CL09b, Jia08, Nut13]. Expected [ASG93, Vit91, BS15, FS99, Sud08]. expert [BEZ20]. Explicit [BT00, DP15, PS97, SV94, GK03, HH08, HJK12, LN06a, LN09, LdRS18, LW22]. explicitly [NP99]. exploitation [CS23]. Exploration [CS23, Fra02, BBDW22]. Exploration/Selection [Fra02]. Explosion [Wag05]. Exponent [Pin92, AZ10, CmHP04]. Exponential [Bar19, BT08, Ber97b, EG18, GQ03, GHH07, IS04, JM91, KRM15, TY93, AMR04, BBS11, BL02, FJR20, Gap05, GRZ04, JKM07, KM08, Lac17, LX14, LMT96, MS05, MPdS19, RY13, RR19, YRF16, Žit09]. exponentially [DG07, Dur13, Gur14]. Exponents [HW92b, BH00, BC14a, CL14, JT18, Ngu17]. expose [CD02, CD03]. Extended [BH17, TBUA23]. extension [Mou01, Sch97b]. Extensions [ABBJ94, AA13]. Extinction [AR02, BH19, KKN95, KL01, LL20b, PSY15, AR96, Bor12, Eth04, HN18]. extinctions [Sch05a]. extrapolation [Keb05, OTV12]. extrema [BM19a, Kuz10]. Extremal [BGZ22, HW94, Per94, Pin92, RS14, ABK12, Asm98, BDM17, BH17, Chi16, CAP20, KSS20, Pes14, Yun98]. Extreme [BM01, KPSC10, Niu97, BBO15, CJL21, CR18, DFdH04, Erh00, HR07, Owa18, WJB^{+}15]. Extremes [BD98, AZ14b, EI17, HHR96].

F [CDS23]. F-KPP [CDS23]. F. [CDS09]. face [GRS08]. face-lifting [GRS08]. Factor [Hog93, ET11, HXZ22]. Factorization [Rog94, Kuz10]. factory [MLNW22]. Failure [HS93b, HS93a, KLST93]. Failure-Prone [KLST93]. fair [Bra10, KKLW09, KW04, STZ14]. fairness [Mas07, Wal09]. Families [Ald91, Gla93, MP99]. Family [JM91, FRSW22, Kuz10, LPP15]. Fast [DRVZ17, JPS^{+}22, JM12, KLS95, MS99, MH21, NP05, CP19b, ES03, FFK12, Ger11, GK07b, JLM19, LS15, LS22, LN13, Mar01, RR19, YS22]. fastest [FK13]. fastest-mixing [FK13]. fault [Jel99]. FBSDE [LdRS15]. FBSDEs [BZ08, CCD19]. feasible [Ass97]. feature [BFRT18]. features [ES02]. fed [GZ00, ZBM04]. feedback [Ciu98, HLS19a]. Feedforward [HW92a, CS00b, Maj06]. ferromagnetic [Cou08, ST98]. Feynman [DT05, DPR09, MD01, PS16]. fibers [dBG12]. Field [Ana91, AS10, AZ14b, AW05, BJ22, BvdH12, BLM17, CF18, CDGL22, CS22, CL15, CZ16, CL22, CRT23, Dar23, DM23a, DR11, DFFN21, DEH23, Dje23, DS17, EOT05, EP10, FHH23, Fis17, FHLN22, FK99b, FMP00, Gra09, GLWZ22, JLM15, KSS20, Kol22, Lac20, LL23, LS14, LS16b, LP21, MR06a, MP22b, NST20, PSY15, Rey18, RS23, Tra23, YS22]. Fields [BFP93, BG93, BG95, FMP95, FHY92, FSW95, KGK95, RWW95, Adl00, ABL12, AFKP20, BB20c, BS19, BDW17, BD20, CW99, CRV06, CX16, CSS22, DFG22, FI20, Gol16, HN22, JS96, JSW19, KM96, LS15, LX14, MRV21, SSV18, Ste99, TK02, Tou14, Tou18]. FIFO [Bra94a, Bra94b, Bra94c, BLP13, GO98]. file [FW22, KPSC07]. filter [vH09]. Filtering [HL01, AC20, Ber97a, CvH07, CD99, CDFO13, CLR06, Del98, DG99, DGN05, INPY13, KLS97]. Filters [MP01, WH93, Bud02, CL04, DT18, DH23, JMRS09, KS05, LO04, OC11, RvH15, Whi13]. Filtration [Lar14, BJJN21, GZ08]. filtrations [LR14]. Finance [Pha02, DFS03, LP17b]. Financial [BR17b, BP05b, ÇD16, CTZ04, JB07, Kar10a, Kle03, Kle06, Pul14, RS05].

Finding [KMS22, BB22]. Finetti [Loe08]. Finite [BG93, DM08b, DKM17, KR96, Ros93, SY94, WH17, BS15, BFHM20, Ber97b, BvdHH10, BvdHH17, BF18, CDGL22, CL22, CCHH05, CvH07, Cla18, CS02, DE17, DL18a, DSC96, DHESC21, EMO10, GM19, HT12, HH18, Jac02, JR14, JMRS09, JM03, Kar07, Lez98, MH21, Mic02, Mor05, Mou20, QS23b, ZY96]. finite-alphabet [HT12]. finite-dimensional [Mor05, QS23b]. finite-fuel [CDGL22, Jac02]. Finite-length [WH17]. finite-rank [Mou20]. finite-size [EMO10]. Finiteness [DR92, Fou15]. fire [DIRT15, FLR22b]. First [Ale93, BGT22, BvdHH10, BvdHH17, CGZ14, Cha92, ESS93, FP93, FN93, FN94, GK95a, KS93a, Kes93, LT19, RS01, Rhe95, vdBK93, AK05a, Asm98, ADH15, BT12, CS22, CT11, CZ03, CN23, DH13a, DHL18, DP15, DFK97, DWZZ20, DLS01, EK09, EJ16, EHW20, FS22, GM05, GM08, Gri13, Hof08, KSS20, KLSY04, KLS06, KPR10, Mar02, Mar16, McD99, Mül18, PW97, Xia97, ZS09, dLS97]. First-order [BGT22, CGZ14, LT19, DWZZ20, Mül18, dLS97]. First-Passage [Ale93, FP93, Kes93, vdBK93, ADH15, BGT22, BT12, CZ03, CN23, DH13a, DHL18, DP15, EJ16, GM05, GM08, Mar16]. first-passage-time [PW97, ZS09]. Fisher [CS18, CPSJ22, Cer15, Che18, CDE^{+}17, GR21, JS17, Pap98, Pap00]. fitness [BCS19, ES02, HM14]. FitzHugh [EGK22, LS18, LP21]. Fixation [LS13, CS18, Hén15]. Fixed [Ana93, Ana94, BD19b, BKT22, CDV14]. fixed-point [BKT22]. FKG [VG95]. flat [BG19, JR14, OQR16]. Fleming [EV12, KR23, DGP12, DK99b, EFPS17, WHN07]. flexible [SY13]. flip [JK17]. Flow [CC93, AS01, CT11, KP97, MP03, MW13, RS03a, Sta97]. flows [BB20c, Ban19, BGHK20, CX97, CS02, FK99a, Gar09, Ken11, LS00, PR98a, RS23, ZBM04]. fluctuating [BL16b]. Fluctuation [BBC23, CG16, KKP12, KP03]. Fluctuations [CLW16, DM23a, FV21, Kot92, BBD22, BF22b, ESTZ13, GLM17, He20, JJQ16, KQ18, LS16b, MRV21, NY16]. Fluid [Che95, Dai95, DR02, GW09, KR10, Kel93, KW04, Kot95, Mey95, MW13, RR03, Rog94, SW12, Bra99, Bra11a, Dai96, FW22, GHR03, GH05, GPW02, HRS99, JM03, Kel96, LN13, Mag00, PR98a, PW04, RS03a, ZBM04, ZBD05]. fluid-scale [Mag00]. fluids [TY11, Yos12]. Föllmer [Lar14]. force [BS21, ELM22]. forced [Fon10, MSW97]. Forcing [Lef91, DWZZ20]. forest [CCR23, DZ15, FLR22b, JPV99]. Forgetting [DGLM10]. Fork [Ngu94]. Fork-Join [Ngu94]. Form [HW92a, ASCDH09, BFRT18, BZ19, Kel96, LC03]. Formalism [AGD94, MMN22]. formation [KPSC10, PS97, Yu15, Yu17a]. forms [BMM21, HS99]. Formula [ABT92, DKT91, JKK02, BBM07, BM96, CN23, EPW06, FG18, Fen07, GR21, Gen23, GNP17, JS10, KY10, Sas18, Tsu19]. Formulae [Aka95, FLP13, GS02, IKKM15, PS16]. formulas [Adl00, CGR20, CD02, CD03, OQR16]. Formulation [DPT01, DO94, KTPZ15, BTZ04, BT22b, CL15, HLN21a, PTZ20, STZ13]. formulations [HV97]. Forward [Ant93, BS14, CDET13, ÇD16, CM96, DM06a, DMP96, MWZZ15, Žit09]. forward-backward [CM96, DMP96, MWZZ15]. forward-reverse [BS14]. Fourier [BL21]. fourth [FK22]. Fractal [HR94, BEMT21]. fractals [Cla22]. fraction [Hei05]. Fractional [FK00a, MRRS02, Oha09, AG09, BC09, CS06, CD99, CS17b, Duf16, GZ00, HN22, HLN16, HLN21b, IT99, Ino02, JR16, JNP21, LS22, LT19, Lud08, MNG09, VZ19]. fracture [KL99]. fragile [PRSS20]. fragmentation [BHK11, DGM06, Fou22, Haa05, Haa10, HH18, Wag05]. fragmentation-coagulation [Fou22]. fragmentations [MP22a]. fragments [MP22a]. framework [BGHM10, BF08, BZ19, CSX23, DM06b, MS11]. Fréchet [Le15, LP17b]. Free [Kle03, BB23, BB15, BS22, BR08, BFM17, DR13, DPBCD21, Feh23, Fer15, FMMP08, Gen23, HHJ17, JLM19, KRSS23, Kle06, LP17b, NS22, QS23a, Str05, vdHvLS21]. free-boundary [Fer15]. freedman [MP01]. freezing [MRV16]. Freidlin [dRST19]. frequencies [CA18, PY18, TT14, YY09]. frequency [CS18, DK19, JT18, MV16]. frequency-dependent [CS18]. frictions [GR15]. Fringe [Ald91]. frog [AMP02, GTW22]. frogs [BC22b, HJJ16]. Front [Cha93, BF10, Jel99]. fronts [CDS23, MSW97]. Frost [Nea06]. frozen [vdBN17]. FSDE [Zha05]. fuel [CDGL22, Jac02, Sch13]. full [Che08, KTT17]. Fully [CH91, FTW11, GZZ15, KLP15, Ren16, See20]. Function [FP95a, ABB23, ABH17, AJO14, APP15, Bal19, BEG00, BC18, BG06b, CM23, DP20, DI17, EMO10, GK21, Har13, Hey23, IS04, OWZ97, TT14, Wik01, Zer98]. Functional [CES23, DGC20, DKP22, DKT91, HLMS05, LP08, PP22, PW97, BHQ17, CW99, DM08a, DPR09, DJN08, GK16, LPZ17, Sch10, Zha12, dRST19]. functionally [MSZ19]. Functionals [Ale94b, BMS02, RY94, Rhe93, YH93, BC09, GRZ04, HLN07, HK13, HLS16, JLP08, JKO09, LRP17, LRSY19, LR19, LO13, LO17, LZL21, PP08b, PY16, Sch04, ST21, Web01a, Web01b]. Functions [BGT01, RR91, APP15, BFRT18, BW22b, BG12, CL17, DG13, EN18, GOP03, GY04, Ino02, Jor02, Kar07, KPS98, KS99b, KS06a, LNS21, LR99, Mar19, Mey06, Pap98, PPM18]. fund [CK19]. fundamental [Pul14]. future [SSX14]. fuzzy [Häg99].

G [CLW94, GG13b, OCBG11, Ree09, DHT10]. gain [BP97]. Galerkin [CJK19, Yos12]. Galton [KR19, Als93, AR96, AR02, BDMT11, BS12, BM13, Chi04, Die15, HJR20, Mar08, QS94]. Gambler [DHESC21]. Gambling [FH16, Eth96]. Game [GK95b, BY17, CS22, DK08b, DK08c, Dol10, Dol13, HHSZ15, Kif06a, Kif06b, Lac20, MD18, MPZ21, NST20]. games [BHMW16, BL16a, BCZ19, BF12, BMN14, CF18, CDGL22, CL15, CZ16, CL22, Che18, CRT23, DDM11, DFM18, DMP22, DFFN21, Dje23, EV06, Fis17, GRS^{+}16, HRW08, Kol22, LL23, Lan15, LC03, NZ15, PTZ20, QZ23, SV02]. gamma [KQ18, Yu17b]. gap [Cha05, FIKP98, HS19a, KM18, Klo19, NRS21]. gaps [AG12, HSV14, HJ23]. GARCH [BH03b, BH01, BHK05, BCL06]. gas [Alb09, BF22a, BCF18, BG19, Nor16, Pèn05, Tur18, Vys08]. gases [BFG13, BGZ22]. Gaussian [ABL12, AZ14b, AW05, BHZ02, BY05, BPY09, BDW17, BM19b, CX97, CX16, Der11, DFG22, Ebe14, FHH23, GNP17, Har13, HJ23, HPS03, HHMS93, JvL07, JKK02, Kab10, KM96, KP97, Lac22, LP23, LS15, LX14, LS16b, MRV16, MvU05, May09, Mek15, NY16, Ros95, RWF13, TK02, Tor16]. Gaussianization [CN11]. Gaussians [AK05b]. gelation [SSW06]. GEM [PY18]. gene [CDMR12, DFPR22, DP09, GT99]. Genealogical [DG05, DK99b]. Genealogies [DF16, EV12, MD01, ZCD05]. Genealogy [GPW09, Pop04]. General [Ald91, Big95, CJ94, Col02, CHS22, DR92, FP95a, GHK11, KMS17, KS92, MS93b, Rei95, Sch92, ABA20, ADGS98, BC02, BGV20, BLM23, CLW16, DK19, DGMO11, Eva01, FW22, GM12b, GW09, Han06b, HS07, HP23, HQR96, HPŠV04, Jaś07, JMSS12, KKM04, LS16a, LYZ19, LM09, MMPP08, MS11, MPZ21, NR04, RS06, SYY20, Tie98, XS92a]. Generalisations [QS94]. generalists [LN06a]. generalization [BT22a, FJ17, KPL03, Tak96]. Generalized [BG95, BST04, BKS20, GM95, HPTvD95, Laz04, MD94, Taf11, vM95, BPY09, BK06, BGG^{+}16, CF09, CC98a, CCL13, CF16b, FG18, GZ06, HW05, HNNZ13, Lei08, Mic02, Moy15, RR03, RR08a, Sil96, Sto04]. generally [GO98]. generated [CDV14, IRR12, MSZ19]. generating [CM23]. Generation [EH95, DH23, GMO15, Žit09]. generator [Pet91]. Generators [MZ91, DI10, DM02, ESX22, TLC93]. generic [Hol98]. Genetic [MD01, Cer96, EPW06]. Genetics [EK92, Mor92, AdHR21, DR08, Dur13, FH98]. genome [BHP10, BCTV07]. genomic [Pia05]. geodesic [GO98, MS18]. Geodesics [Hof08, Ken11, ADH15, DS19, KMS22]. Geography [Cou10]. Geometric [DS91, DGM20, Ing94, MT94a, AABR22, BBK^{+}11, BY05, BKS17, BDL16, DS18, GRK05, HLS16, IT12, LR19, Leh23, MP21, MS15, MP20, PY03, Pen16, SZ06, SY23]. Geometrical [AB92, AB93]. geometrically [Bax05, KP96, KP98, KM03]. geometries [QS23b]. Geometry [PY01, BCJ22, BD20, BBFM03, CS16a, CW99, Yuk15]. Gerber [APP15]. GI [DM08a, DHT10, GG13b, Ree09]. GI/G/n [GG13b]. giant [BN17, JW21]. Gibbs [LPW14, BFP93, BR15, CRV06, CGR09, CG92, Com97, ER08, FLP13, FFS93, Fis96, FSW95, HK21, HK23, Hub15, Ing94, KP04b, ŁRR13, LPW08, MMN22, MV18, Muk22, PY18, RR98a, RR06, Smi14, vEK08]. Gibbs-type [FLP13]. Gibbsian [Mas95]. Ginibre [BB20a, CR22, Gol10, PTZ17, RS14]. Gittins [Tsi94, Web92]. Given [Lef91, BHS12, CDS11, DS22, GHLT14, HLN08, HLOST16, Jos14, KTT17]. glass [CP18, DLZ21, JKS18]. Glassy [MRV16]. Glauber [DdHJN17, RR19, SZ17, TVVY12]. Glivenko [LZL21, SSV18]. Glivenko-Cantelli [LZL21]. Global [CC93, DP20, DIRT15, FHLN22, Fra02, RS23, Cal97, CGZ14]. Global-in-time [RS23]. globally [Mar08]. globe [HS19b]. GOE [JM12]. golden [GHP13]. Goods [FP95b]. gossip [CD11]. gradient [CK12, HK21, HK23, Sas18, TD17, vEK08]. grain [Hei05]. Gram [HLN08]. granular [Mal03]. Graph [CH91, LC93, And98, AR16, Bha16, BDM17, Cra16, DDMT12, DFK12, FKP23, GS23, Gol13, Jan08, JŁTV12, Jos14, KS14, Pit08, RŞ18, RW97, vdHHKR16]. graph-based [vdHHKR16]. graphical [AL05]. Graphon [AdHR21]. Graphon-valued [AdHR21]. Graphs [PY01, Wor95, ABA20, ACD15, AABR22, AL15, AS16a, AK18, AP17, BST14, BN17, Bal21, BBK^{+}11, BS17a, BH20b, BFHM20, BP12, BvdHH10, BBS11, BvdHH17, BBDW22, BS22, BM17, Blo13, BdHM23, BDL16, BDW22, BMW19, Can19, Cao21, CDS11, Che13, Cop22, DGJ^{+}19, DS18, DM10, DS22, DJ10, DAM10, DdHJN17, EG18, FKR17, FKKM18, GRK05, GLS18, GSvdB98, HS07, HMQS20, IT12, JL08, Jár17, JPV99, JK22, KKM06, KMS22, KM18, LS17a, LV10, MP21, MV18, MP20, Nea06, PRR13, Pen00, PS05, Pen16, RS22, RY13, RR19, San08, Sly08, Web01a, Web01b, WXY23, YRF16]. grass [DZ15]. Greedy [CGGK93, DFP93, GK94, Lee93, CW19, GM08]. Greek [EFT07]. Green [Sas18, Zer98]. Greenberg [BS23, FGG93]. Greenberg-Hastings [BS23]. GREM [FGG14]. GREM-like [FGG14]. grid [HR97, LR13]. grids [SBF19]. Griffiths [EOT05]. group [KPSC10]. groups [SSV18]. growing [BB22, BBC23, BBD22, BBO15, Dur13, JK22, LM96, LZ98, MM07, RS22, Sab16]. grows [Tsu19]. Growth [BPT22, Big95, GK94, GW93, Lig95, dGvZ93, AMS06a, BHL96, BH13, BDKT20, CFS18, CQ97, DH13a, FS99, GHH07, Gou07, GM08, GG97, HHK20, Hof05, HQR96, Jag99, JW21, KR12b, KR21, Kar10b, Lam05, LdRS15, LdRS18, Mül15, OQR16, Ric11]. GUE [JM12]. guessing [Ciu98]. guide [BC01]. guided [MW10]. Gyöngy [DGL23].

habit [Yu15, Yu17a]. Hack [RSS16]. Hahn [BC16]. half [ADH20, EF21, HL17, OQR16]. half-flat [OQR16]. half-line [HL17]. half-plane [ADH20, EF21]. Halfin [ABP15, AP16, AG12, BM19a, BM20, GG13a, GG13b, Ree09]. Halton [AS16b]. Hamilton [BBK^{+}11, Che22a, Mou20, QZ23]. Hamiltonian [BRSS17, BREZ20, CLP16, DPBCD21, MS21]. Hammersley [AS16b]. Hamming [GHPS15, GS20]. hard [Bal19, Bor16, GK21, HPP22, Hey19]. harder [BL16b]. Harmonic [NV03, Pia06, Pin92, BGJ18, GL18]. Harmonisable [ASG93]. Harris [CT01, Dai95, RR06]. Hastings [BS23, BPT22, BRS09, BR17a, Ebe14, FGG93, HSV14, HHH09, HJ23, Tie98, VK21]. having [BS19, GM19]. Hawkes [DFH16, Gra21, JR15, JR16, Tor16, Zhu15]. Hazard [RS95, BJR08, PP08b]. heat [BM19b, DH18, FJL18, GT19, Muk22]. heavily [GPW02]. Heavy [AK05c, BBP22, BD01, CPR95, DN94, GK07a, Har98, KLRS11, Ngu93, TW09, YH93, Yam95, AMR04, Ata05b, AMS06b, AC17, BM20, BBC19, BW01, BG08, BG05b, BG12, CBM^{+}21, CSS98, DR96, DLS01, FPZ05, GZ06, Gro04, HW96, HRS99, HL97, JR16, JM03, KLSY03, KLSY04, Kum00, Lim01, LW14, MS00, OCBG11, PR10, RR03, RR08a, RT15, RR00, Sha15, Sto04, Viv23, ZBM04, vdHMS08]. heavy-tailed [BG08, DR96, FPZ05, JM03, MS00, ZBM04]. Heavy-Traffic [CPR95, TW09, BM20, KLSY03, KLSY04, LW14, RT15]. hedge [CS99b]. Hedging [CK93, CM96, Dol13, DR91, GR15, Sch92, SSC95, BZ16, BS05, Bec06, BT22b, CT04, CK19, Dol10, ER18, GL14, HS19b, HKK06, JMSS12, LS97, Myk00, Owe02, Pul14, RT14]. Height [BB92, Cha92, Mah94, MS93b, DH06, PS22]. heights [DFK97]. Heisenberg [CM03]. Hellinger [Clé23]. Hermitian [GS18]. Hessenberg [BH19]. Heston [ER18]. hetergeneous [JM03]. Heterogeneity [Aal92, FHH23]. Heterogeneous [BK92, Fin94, Ngu94, DDMT12, LN06a]. heteropolymer [BdH99, Wüt06]. heteroscedasticity [GRS00]. heterozygosity [SS08]. heuristic [DTW22]. Hidden [DTW22, DGLM10, KGK95, CvH10, CL14, DGMO11, FL03, KR06, MR06b, Pes14, WL16, vH09]. hierarchical [FRT14, Gol04, RS03b]. High [BKR22, BF02, CL17, FK21, LL12, MV16, RS91, SYY20, ABL12, AA13, BRS09, BCJ14, CCK23, DRVZ17, Ebe14, FPZ05, GZZ15, HS21b, JT18, Kab12, LLS08, MPS12, Pen96, PST12, QS23a, RZ08, TT14]. High-dimensional [BKR22, FK21, SYY20, GZZ15, HS21b, Kab12, QS23a]. High-frequency [MV16, JT18]. high-order [AA13]. High-risk [BF02]. higher [BLO23, Yun98]. higher-order [Yun98]. Hilbert [FG18, LO04, NP01]. Hilliard [RYZ21]. histogram [JR14]. hitchhiking [EPW06]. Hitting [DH13b, Miy04, RS95, BC02, NS19, San08, Sil96]. HIV [Ber92, Ber94]. HIV-Latency [Ber92]. HJB [ABW07, KLP15]. Hodgkin [Aus08]. Hoeffding [LP04, LP17b]. Hölder [Mar19]. Holm [CDG23]. homing [Lef04]. Homogeneity [FHH23]. Homogeneous [DM20, FMP95, Bal19, BHK11, CFS18, CCL16, EH11, Fou15, Jag99]. Homogenization [CX97, DFG22, Feh23, GO12, INPY13, MN17, Muk22, PS16, ST10, YZ19]. homogenized [GM13]. homology [MW07]. honest [LR14]. hook [AS23]. Hopf [BE23, Kuz10, KKPvS11, Rog94]. Hopfield [BM01, Löw98]. hopping [CF09]. Horizon [HP92, NP02, CL22, CvH07, Cla18, DE17, RZ18]. Host [Ish95, BLZ11, LN06b, Sch97a]. host-mutualist [LN06b]. host-parasite [Sch97a]. host-pathogen [LN06b]. host-symbiont [BLZ11]. Hotelling [PZ11]. Household [FP22]. Hull [Hsi94]. hulls [CY21]. Hunt [ZZ02]. Huxley [Aus08]. Hybrid [IPB^{+}11, CDMR12, RR98b, ZBD05]. Hydrodynamic [Bar20, CF17, LT22, Per00, AR19b, LPSX23, MP23, Nor99]. hyperbolic [Bro99, FM18, KM18, LP23]. hypercube [GH21, Yan20]. hypergraph [BCOR16]. hypergraphs [DN05, DM08b]. hyperplane [HSS06]. Hypocoercivity [ADNR21]. Hypoelliptic [LS18, HSV11]. Hypoellipticity [BT05]. hypothesis [JLZ21, WH17].

i.i.d [Béd07]. i.i.d. [AL05, DH13a, Yat09]. identically [YS96]. identification [JMRS09]. identities [KKP12]. IDLA [CCR23]. idleness [CBM^{+}21]. IGF [MD18]. IGF-II [MD18]. Ignatov [Yao97]. II [AAK17, AV95, BK15a, CL22, FS02, GK94, HSV07, HHK20, Hwa98, MD18, Rhe94, XS92b, ZCD05]. II-the [CL22]. IID [Sel95]. III [CN11, CLP16].

Imaginary [JSW20]. Immediate [JK17]. immigration [BS17b, BR13, Haa05]. immunization [Sta03]. impact [BD19a, BMRS23, BT22b, CHMK20, KP16, LRS17, Loe18]. impedance [PS16]. implementing [MR06a]. Implicit [Gol91, Whi91, NT20]. Importance [BFP93, Col02, DSZ15, MP99, ADRS23, BCKP99, Bla09, CL07b, CD18, DL10, DW05a, DSW07, GW97, JL09, LP10]. importation [BBT17]. impossibility [FLR22a, GLM18]. Improved [LP17b, MRW18, HV06]. Impulse [DY95, CS23, MR00]. including [GH21, Muk22]. inclusion [BYZ00]. inclusions [NY23]. income [MS20]. Incomplete [AG93, HIM05, HK04, KS99b, KS03a, KS06a, MS20, Owe02, Sch01, Yu15]. incompressible [CX97, KP97]. inconsistent [HP23, KMZ17, NZ20]. Increasing [MD01, Sep97a, BG00, DG07, GS19, HT12, Ste99]. increment [Alb04]. increments [BBRZ20, CX16, Das96, FZ03, GMO15, HKK06, KMK10b, TT14]. Indefinite [SXY21]. independence [AKLT22, LLS08, Yao97]. Independent [AB92, Ben96, CHL21, Che22b, CX02, Das96, DS07a, Gol16, HV06, HHH09, HKK06, Jin19, Jor02, KMK10b, Sep97a, Wik01, YS96]. Index [Ano99, Ano02, Ano03, GW00, HR94, Tsi94, Web92, DFdH04, RS98, Yat09, Yun98]. Index-based [GW00]. indexability [GHK11]. indexable [Ver16]. indexed [BW22a, BDMT11, IM02]. Indian [BCPR15]. Indices [Gla93]. indifference [BK15a, Bec06, MS05]. Individual [BL06, BBC17a]. individual-based [BBC17a]. Indivisible [FP95b]. Induced [CBM^{+}21, SSW06, BG02, BE23, HRS99, LP21, MP14]. industrial [BP05b]. Inequalities [Cha94, HK92, MN97, vdBK93, ABP^{+}13, AGSC02, BT08, CF07, CHO08, DR11, DSC96, FK13, FK99c, GLWZ22, HS23, JM08, JY17, Mar02, PDG14, SS19, WW20]. Inequality [Din95, Lee93, VG95, Kar07]. infection [BHL96, O’N97]. infections [BS15]. infectives [BBT17]. Inference [Whi91, Che22a, KSS21, MV06, Mou20]. Infinite [HP92, Mor92, MP95, NP02, ABGK12, BT22a, BT05, BLZ11, Chi15, CvH07, CGK^{+}23, DH13a, DKM17, GGLO13, Gol16, HSV14, Haj96, HH19, HMQS20, Hol98, HJ23, JKK03, LPZ17, PR98b, TKH09, Tei09, ZCD05]. infinite-alleles [JKK03]. infinite-dimension [CGK^{+}23]. infinite-dimensional [HJ23]. Infinitely [FF94, AS04]. infinity [Ban19, BRSSJ19, Fou22]. Influence [Gla93]. Information [MP03, MR17, NP02, ABA20, BB20c, BLY22, BEM08, ÇJPY04, ÇD16, DMP22, HLN08, IV17, Laz04, MWZ07, PRSS20, PPM18, SSX14, WW20]. information-percolation [ABA20]. Information-theoretic [MR17, HLN08]. informed [AL17]. ingful [KPSC10]. ingress [MW13]. inhomogeneous [AEK18, BBC17b, BS22, BDW22, CQ97, DMR04, FKKM18, HS21a, JKS18, SCZ10]. inhomogènes [Mic96]. Initial [DH18, BHL96, CFS18, CLW16, DH07, DGLM10, Fed14, FH16, Fou15, OQR16]. Initial-boundary [DH18]. initiation [FGLS23]. inner [DHESC21]. innovations [FKR96, PR98b]. Input [KW91, AS04, HM09, HRS97, IT99, KLSY03]. input-constrained [HM09]. Inputs [Kel93, CBM18, Kel96, MvU05, Wis01]. inscribed [EN18]. Insensitivity [HPTvD95, Mas07]. insider [KX22]. Inspection [BENP91]. Instability [Bra94a, Bra94b, Bra94c, GH05, SY13, Dai96, DHV04, MP17]. Insurance [Pha02, AM10, EK09, GM12b, Gri16, KKM04, LM09, Tou00]. integers [BK16b, CNV22, CW19, PRS23]. Integrability [TY93]. Integral [Tak93, DM02, Fer15, GMO15, PW10]. Integrals [PRW95, Ste93, Ste95, BD20, CST05, Fuk11, LR13, LX14, Wik01]. integrands [Taf11]. integrate [DIRT15]. integrate-and-fire [DIRT15]. Integrated [BC09, BDW23, BHK05, JT18, KM96, LZL21]. Integration [BBM07, RWW95, DP05]. integro [Zha12]. integro-differential [Zha12]. intensities [MAL14, dLS97]. Intensity [CW93, GZ08, BGR18, CC16, CCL16]. Interacting [DD10, GR09, Mül15, ABL21, BPR22, BS05, BP18, BT96, BF18, CGR20, CF17, CT23, CF16a, Cop22, CK03, Del98, DG99, DT05, EW03, FM16, GL23, HP20, JK14, JM08, LRW23, LY17, PS14, Shk11, YS22]. Interaction [Fin94, FRST94, Ish95, ACG17, BCH22, CM19, GL23, LS16b, LP21, NS19]. interactions [BLZ11, DP19, LN06b, LS14, Muk22]. interactive [Sta97, dLS97]. interchange [HS21b]. Interest [Mil94, BT05, CT04, FK16]. interface [ABT^{+}11a, ABT^{+}11b, BCJ22, BdH99, CP09, vEK08]. interfaces [AS06, BG05a, DS17, FMP09, PW23, SSY19]. Interference [SBF19, BS07, PGZ07, PZ08]. interior [Nar16]. interlaced [LT22]. interlacements [Bel12, ČT16, Tei09, TT13]. Intermediate [Cox10, CW16, KQ18]. internal [DK15]. Interplay [CTZ04, LSZ13, HRS99]. interpoint [JJ15]. interpretation [ABP^{+}13, BKH15, ML16, Tom21]. interpretations [Ber02]. interruptible [Fil98]. intersection [BST14, Blo13]. interspecific [NP99]. intertemporal [BR01]. Interval [CFJ00, MP22a, Ber97b, FZ03, FPZ05, FRLS21, GW22, KS99a, KLZ98]. interval-partitions [FRLS21]. intervals [BG01, BDS20, RT96]. intrinsic [GNP17]. Invariance [BDW23, BDW17, Pia05, Ber97a, CKW21, Jia08, KW07, Rey18]. Invariant [CC93, GOP03, HS93a, PW04, BL11, Bud02, Bur07, HK23, JK14, Mül15, PP18, Pan08, Sch04, Kar10a]. Invasion [ABGK12]. Inventory [BZ91, SY94, HSZ17]. Inverse [JMRS09, CCCS11, DP15, DW98, EJ16, EHW20, RWF13, ZS09]. inverse-Gaussian [RWF13]. Inversion [CLW94]. inversions [BP12, Ber96]. Investment [DZ01, SS94, XS92a, XS92b, AM10, BMX17, BB20c, BC18, BP05b, BF02, DK08a, Fer15, FS02, GGS03, Gua02, HK04, JKP13, KS99b, KS03a, KS06a, LM09, MS20, Sch01, Sch02, VPV08, Wee98]. investments [Pau02]. investor [BK15a, CM96]. investors [AL17]. Inviscid [Jou02, BFM11]. ion [Aus08]. ion-channel [Aus08]. IPDSAW [CP19a]. irreversible [BB20c, Fer15]. Isaacs [QZ23]. ISE [BMJ06]. Ising [BN19, CDN02, Can19, Cou08, DM23a, DM10, EKPS00, GG11, GJ18, JSW20, PW23, ST98]. isometries [Pel10]. Isoperimetric [CF07]. isotonic [HK22]. Isotropic [LS15, MRV21]. Itô [BS05, BS13, BM96, TT14, Wik01]. items [LvZ04]. Iterated [DeB04, ABB23, Jor02, LR99, Wik01, dRST19]. iteration [BZ08]. Iterative [STT19].

Jackson [FM05b, GZ06, IR00, KST04, MY96, MS99, Mar01, MD94]. Jacobi [QZ23, Che22a, Mou20]. Jagers [LSZ13]. Jeulin [GZ08]. Jigsaw [BCDS15, GS17]. jitter [DG08]. Join [BM19a, FM01, Ngu94, Bra11b, BLP13, BFW21, BM20]. Join-the-shortest [BM19a, BM20]. join-the-shortest-queue [BFW21]. joins [RS07]. Joint [Wik01, HLR22]. Jordan [BB22]. JS [KS03b]. JS-queues [KS03b]. Jump [KW91, ADE18, BGR18, BS12, CFY05, CNV22, CLSF18, HX22, JT10, JT18, MSSZ20, PPZ22, TT11]. jump-diffusion [CFY05]. jump-diffusions [CLSF18]. Jumps [YH93, ASCDH09, AB22, Bec06, BF18, CKW21, CM08, FK16, Gap05, HK16, KTPZ15, KLP15, MAL14, Miy04, RT14]. Justifying [YY18].

Kac [CF16b, DT05, DPR09, DGR09, DR10, Fou00, GR08, Hey19, Hey23, JPS^{+}22, Jia12, MD01, Nor16, Oli09, PS16, PS17a]. Kaijser [KR06]. Kalikow [GGLO13]. Kalikow-type [GGLO13]. Kalman [DT18, DH23]. Karhunen [LT18]. Karlin [BM97]. Keller [FJ17, Tom21]. Kernel [EFT07, NP01, AJvM22, Ber02, HL00]. kernels [DKP22, Tie98]. Kesten [GLP23]. Kesten-Stigum [GLP23]. Khasminskii [Cer09]. Killed [EEH14, EHW20]. Killing [MV16]. Kimmel [Ban08]. kind [BT13]. kinetic [BL12a, CV21, EW03, LP21]. Kinetically [MT13, Sha20, Blo15]. Kingman [BRSSJ19, DK15]. Kipnis [GM13]. Kirkpatrick [GK21]. KMT [CG17]. Knudsen [BG19, Pèn05]. Kohonen [FP95a]. Kolmogorov [CS02, HN18]. KPP [CDS23, Kli19]. KPZ [BFS21, CFS18]. KPZ-Burgers [BFS21]. Kramers [HIP08]. Kreweras [BM05]. Krylov [DGL23]. Kubo [Sas18]. Kurtz [ZZ02]. Kutta [CC14].

labor [MS20]. Lack [ST98]. lacunary [Jaf00]. Ladder [Cha92, Kee94, MS93b]. Lagrangian [CI11, KK04]. Laguerre [WH93]. Lair [BD92]. Lamperti [Jam10]. Lamperti-type [Jam10]. Landau [Gué03, JR14]. landscape [Cer15]. landscapes [ES02, HM14, Kab12]. Langevin [DLZ21, DRVZ17, DM17, LW19, PP18, PP23, PST12]. Langevin-like [DLZ21]. Laplace [GQ03]. Laplacian [DJ10]. LARCH [BH03a]. Large [ADE18, Ath94, AV95, BPR22, BCZ19, Ben96, BFG13, BS22, BNS21, Blo92, BL12b, BGL02, BFW21, CP19b, Chi05, CS13a, CZ03, DRST09, DDMT12, Dd04, DKZ94, DS22, DS05, DAM10, DWZZ20, DW98, DR98, Feh23, FH98, Fen07, Fin94, FM05b, FK16, FN17, GHR03, GO00, Hei05, HIP08, Hey23, Hwa96, Hwa98, IR00, IR01, JKM15, Mey06, MP23, Mor92, PR94, PPZ22, Pia99, RS01, Too02, YS22, ZZ02, Zhu15, ARS18, AH98a, AH98b, ABDW21, BLW11, BK15a, BL08, BF10, BMRS23, BBRZ20, BCC^{+}23, BM13, BB15, BPT98, BR17b, BG05a, BC22a, BdHM23, BG05b, BF18, CT11, CPS11, Chi07, Chi15, CN23, CGM09, CM96, DN05, DF06, DGC20, DFH16, DM08b, DTW22, DL22, EV12, ER10, FM18, FT17, FRZ04, GSS13, GW97]. large [GR09, HLN07, HIP06, HDP22, HLMS05, KPT^{+}16, Kar07, KR11, Kle03, Kle06, LC22, Maj06, MY96, MvU05, MSW97, MR12, Nag12, NY16, NV03, NV04, NV06, Ngu17, PZ11, Pap98, Pap00, PY03, PY98, PW97, Rey18, RR97b, Sad96, SY13, SFR16, Wis01, Yan20, YS96, YY22, YM22]. Large-deviations [GHR03]. large-dimensional [LC22, YS96, YM22]. Large-scale [Feh23, MP23, MSW97]. Large-time [FK16]. large-tree [MR12]. larger [GL18]. Largest [RS01, BB20a, DY18, FM18, Jan08, Jia04, LS16a, LR06, Ona08, Pit08, PTZ17, SX23]. laser [BT17]. Last [BT19, FMS14, FKP23, KPR10]. Last-Passage [BT19, FMS14]. Latency [Ber92, Ber94]. Latin [GH21]. Lattice [Ana91, CGGK93, GK94, Lee93, Zha16b, BFG13, CMSS15, FP18, GS20, HSST23, LR21, Lan15, MZ05, Sep97a, TLC93]. Law [Blo92, Fin94, FM18, Jou02, KR11, Mah94, AEK18, BF10, BC22a, BT96, Cha05, Duf16, FH16, Fus00, HIP08, HLS19b, Jan08, JM12, Pit08, PS20, RR97b, RSS16, Sch04, Tak96, TY11, Woo12, Xu19, dRST19]. law-invariant [Sch04]. Laws [BL08, BBM07, BJ22, BMJ06, Chi04, DJ23, DWZZ20, Gau98, Goe06b, HR07, JM15, JMRS09, Jam10, KPR10, NR01, PY03, RS14, SX23]. LDP [FK16, dRST19]. LDP-correlated [FK16]. leader [FMS96]. leads [CBM^{+}21]. leagues [BC22a]. leap [AGK11]. Learning [AK05b, MR06b, BP97, CL22, CS23, Egl05, KMS22, LL17, LC03]. Least [Ser94, Jel99]. least-recently [Jel99]. Leave [Lef91]. Leaves [MS91]. left [HN19]. left-curtain [HN19]. Leland [Per03]. Length [Ale94a, Ale95, Fri91, BH12, CS21, DDSJ08, Jin19, Ker12, MR17, WH17, Yuk99]. lengths [AS23, DK15, DK19]. Lerch [JvL07]. Level [ASG93, AS97, BD20, Fri16, MP09]. Levels [LC93, BMST97, Bel12, Kab12]. leveraged [CK19]. Levin [DZ15]. Levitov [BPT22]. Levy [DKZ94, KM95, KW91, Kel93, AK05a, APS19, AKP04, APP07, APP15, BN15a, BBRZ20, Ber97b, BL02, BMS02, BDW22, Cha99, Clé23, Cra18, DP15, Der11, DL16, DM05, DK06, EK09, FLÓ19, Ger11, GS14, GM12b, Gri13, Gri16, HK16, Iva18, JKM07, Kel96, KKM04, Kuz10, KKPvS11, KKP12, Loe08, LØP04, LPP15, LP08, MRRS02, Pan08, RZ18, Shk11, TKH09, Zha12, DK08a]. Lévy-driven [APS19, Der11, DL16, DK08a]. Lévy-type [LPP15]. Lie [CL09b]. lifetime [EEH14]. lifetimes [DG07]. LIFO [Lim01]. Lifshitz [KKM06]. lifted [BR17a]. lifting [GRS08]. Light [Asm92a, BR93]. Like [Dev92b, FMP95, ACH97, BF05, CKW21, DM10, DLZ21, EG18, FGG14, LV10, SCZ10, Pau02]. Likelihood [BG95, Mas95, AR20, DFM16, DI17, DRS16, DFdH04, FL03]. Likely [Pit92, AL05]. Limit [Ana91, And98, AB93, BDMT11, BBO15, CPS11, CFS18, Chi04, Coh04, CH91, Dai95, DR96, DFK12, Dev92a, DKT91, FM94, Goe06b, Guy07, HR07, HST18, HS99, ILP15, JT18, JR15, JKO09, KL91, Kus95, LR00, Mey95, MP21, NR01, Owa18, Pen93, PY01, PY02, PY13, Per03, PSZ14, RY94, ST04, TT11, TT14, YH93, YZ07, Zha95, AR19b, Ata08, Aus08, AG09, BPR22, Bal21, BM19a, BM20, BR19, Bar20, BHQ17, BK15b, BKN22, BCPR15, BDM17, BJR16, BR17a, BC22a, BT96, BMJ06, BC14b, BNS13, Can19, Cao21, CP19a, CCK23, CQ97, CES23, CW03b, CSS22, Dai96, DH13a, Dar23, DM08a, DDMT12, DG99, DPBCD21, DL16, DGR09, DR10, DJN08, DFG22, EP10, FK21, FK00a, Fér20, FKM96, FSW15, FGG14, FRT03, FJ19, GR08]. limit [GVR17, GS23, GPW02, GK07a, Grü14, GK16, GL23, HSS06, HSH^{+}13, HLS16, HP20, Hwa96, Hwa98, HLS19b, JLM15, KKP14, KL96, KSS20, KM96, KM03, Lac20, LPSX23, Lee97, LT22, LR06, LW14, LS14, LN13, Mer07, MP23, MR12, MN17, Muk22, Nak11, NR04, NV04, NV06, Nor99, NST20, Ona08, PZ08, PZ11, PP22, PDG14, PD22, Pèn05, PS20, RR08a, Sch10, Sei09, Sep97b, WJB^{+}15, Yos08, ZY96, Zha16a]. Limitations [MV06, QH21]. Limited [BL01, AAK17, FP18, LR21, ZDZ11]. Limiting [ADH15, GLS18, JWB^{+}14, BF10, BFJ06]. Limits [BD01, BDW22, GK95a, JKK02, Ngu93, RS03b, ADGS98, AR16, BBP22, BY05, BPY09, BCC^{+}23, BK16b, BKR22, CBM^{+}21, CF17, CF11, Cra16, DHT10, DR02, ESTZ13, GW09, JR16, JS07, KR10, KR11, KR13, Kum00, LSZ13, LST23, MP98, MPS12, MW13, Per00, PP15, PST12, Puh15, RS15, RR03, RT15, RR14, SYY20, TW09, TSM19, ZDZ11, vdHMS08]. line [ABB23, FMS14, HR97, HL17, Hén15, Pel10]. lineage [Mar08]. lineages [SD05]. Linear [ABBJ94, BGT01, FM94, GK94, GM08, HKV20, JMP21, PP08b, ABB23, AR21, AKLT22, BS96, Ber97a, BdH99, Bro99, CRT17, CG17, DM06a, Duf16, FHH23, GO98, GJKS15, Gra21, HQR96, HPS03, JMRS09, JLP08, KY10, KT04, Loe18, Miy04, NY16, PZ08, QS23a, Roi07, SXY21, Zha12, dSY05]. Linear-quadratic [JMP21, SXY21]. linear-quadratic-Gaussian [FHH23]. Lines [BCH22]. Liouville [BFO18]. Lipschitz [FG20, GS19, HJK12, MV16]. Lipschitz-Killing [MV16]. Lipschitzian [HHMG19]. liquid [BCP11]. liquidation [HXZ22, PZ19]. Lists [Dev92a]. LMMSE [PGZ07]. Load [BT10, MR94a, MW94, AR19b, AH98a, BGY98, GTZ20, SY13]. loaded [AS09, GPW02]. Loading [KLS95, Maj06]. loads [Haj96]. Local [AEK18, AZ16, BCJ22, DS95, DJ23, Fin94, GvdHL20, Han06a, HLS19b, LRW23, NV04, NV06, Roo94, Tak95, AR21, ABT^{+}11a, ABT^{+}11b, BB23, BRS09, BD07, BMJ06, BKT22, Cha05, Chi15, CS06, Cou10, CHO08, Duf16, DR08, FPRW18, JT18, JNP21, KKN15, Lar14, Olo96, RvH15, Hwa98]. Locality [Cha93]. Localization [BM19b, Sch21]. localized [HK21]. Locally [Loa92, BBM07, BEM07, De 11, DM10, Eth04, FM04, HW07, PPZ22]. location [IT12]. location-dependent [IT12]. loci [MR17]. locus [JS12]. Loève [LT18]. Log [FKP94a, FKP94b, FK99c, AZ14b, DRS16, GK03, HS23, JSTV04, JSW19, KQ18, MRV16, MMS20, MS21]. Log-Concave [FKP94a, FKP94b, FK99c, MS21]. log-concavity [MMS20]. log-correlated [AZ14b, JSW19, MRV16]. log-gamma [KQ18]. log-likelihood [DRS16]. log-optimal [GK03]. Log-Sobolev [FK99c, HS23, JSTV04]. logarithm [KMPT10, dRST19]. Logarithmic [DSC96, DLS03, Kol17, ABT00, DH06, GLWZ22, LL13, RV15, SS19]. logarithms [Bha16]. logistic [Lam05]. Long [BC21, BC14b, Cop22, DMO14, FMS14, HSZ17, KK01, MRS01, AAK17, BPG19, BCPG22, BP12, BH03a, BDKT20, BJN18, BDS20, CFF02, FS99, FZ03, GRS00, GR12, HRS97, HRS99, Imh05, MS11, MH21, Muk22, MW13]. Long-range [FMS14, AAK17, BP12, MS11, MH21]. long-run [Imh05]. long-tailed [FZ03]. Long-term [DMO14, HSZ17, MW13]. Longest [Ale94a, Ale95, Fri91, HT12, Jin19, Pen97]. Longstaff [Ger11]. look [EKT07, GT19]. look-ahead [EKT07]. Lookback [DH93b, GHLT14]. Looking [FP17]. loop [AZ10, BU18, HH19, Lac20, LL23, MP22b, SXY21, WZ20]. Loss [CL14, CH91, GM95, Kel91, MR94a, MW94, ZZ02, Ala03, AFRT06, BGZ97, BMN14, FRT03, GO00, HL97, IS04, JK17, JM03, LRZ06, OT22]. losses [DRST09]. Lotka [BL16b, CP08, KK01, KL01, NP99]. Loud [DO91]. Low [BHdS^{+}20, BL01, RS91, Blo15, RZ08]. Lower [HHMG19, HL20, BH19, CRT17, GTZ20, HS07, MGY23, Ott13, HL22]. lowest [MZ05]. lozenge [Wil04]. LPP [CLW16]. LQ [HSX22]. Luckock [Swa18]. LUE [Bor16]. lunch [Kle03, Kle06]. Lundberg [Mei09, Gri16]. Lundberg-type [Mei09]. LWDF [Sto03]. ly [KPSC10]. Lyapounov [CmHP04]. Lyapunov [BH00, BGT01, BC14a, CL14, DG13, Ngu17, Pin92].

M [Ana93, GG13a, BD17, CLW94, DM08a, GG13a, OCBG11]. M/G/1 [CLW94, OCBG11]. M/GI/ [DM08a]. M/M/N [GG13a]. M/Ph/n [BD17]. machine [CL22, EL10]. machines [DFH13, Gen23, GW00]. Macroparasite [Ish95]. Macroscopic [ČT16, BSZ20, FM04]. Mafia [BEM08]. magnetization [EMO10]. major [CZ16]. Majority [KM11, BG21, Cha97]. Majorization [O’C91]. make [BL16b, BT17]. Makovian [GL01]. MaLa [BPS04]. Malliavin [BPG19, BCPG22, Dec98, HNS11, TY16]. Mallows [Jin19]. Malyshev [KS03b]. management [GHK11]. Mandelbrot [LR00]. manifolds [MS18, PY13]. MANOVA [FJ22]. manta [MM07]. Manufacturing [DH93a, KLST93]. Many [BFW21, DHT10, GW91, LW14, Sel95, Tal92, AR19a, APS19, AMR04, Ata05a, Ata05b, AMS06b, Ata08, BL08, BH19, CT23, EP10, FF94, HP20, JKM15, KR10, KR12a, KR11, KR13, LM02, PR10, Sha15, TW09, Wis01]. many-body [JKM15]. many-demes [HP20]. Many-server [BFW21, DHT10, LW14, AR19a, APS19, Ata08, KR10, KR12a, KR11, KR13, PR10, Sha15, TW09]. map [ABKR18]. mapped [Pap00]. mappings [Alv03, DM20, Pia05]. maps [BM04, LR18, May09, Sad98]. marginal [AV15]. marginally [CSZ17]. marginals [GHLT14, HLOST16, KTT17]. margins [KPL03]. marked [CFJ16, LNS21, Tou00]. marker [CK00, Kli19]. Market [BK92, Kle06, XS92a, XS92b, BK15a, Bät99, BHS12, DK08a, KS16, Loe18, Pal11, RS05]. Markets [FP95b, ARS17, BFK05, BR17b, BT13, ÇD16, DP05, HIM05, HK04, JB07, Kle03, Kle06, KS99b, KS03a, KS06a, MS20, Owe02, Pul14, Sch01, Str05, Yu15, Yu17a]. Markov [ABBH14, ADE18, AH98b, AV15, ATV15, ADNR21, APT20, Ass97, BYZ00, BCKP99, BDMT11, BG95, Bax05, BBC17b, BK16b, BM04, BFH20, BN19, But14, BKS20, CCHH05, CL03, CW03a, CHS17, CBM18, CSX23, Chi16, CvH10, CPT12, CE10, Col02, Col09, CL14, CK07a, CK07b, CGL^{+}15, DD09, DR02, DD10, DY95, DM94, DS91, DSC93, DSC96, DHN00, DRZ16, DS05, Din95, DF95a, DF95b, DMR04, DGLM10, DGMO11, DRS16, DW05a, Erh00, Eth96, Fil91, Fil98, FK13, FR05, FMMP08, FHY92, FL03, Fuh04, FW99b, GW93, Gla93, Gol16, Gos01, GHL03, GR06, GS11, Guy07, Haa10, Han06a, HPP22, HPTvD95, HR04, HKL^{+}19, JKO09, JR02, JSTV04, JS96, JJQ16, JY17, KKP14, Kar07, KZ09, KM13, KMPT10, Klo19, KR06, KM03, KL02, KGK95, LRT03, LP04, Lez98]. Markov [LR12, LW22, LN13, LMT96, MMPP07, MR02, MW98, MT94a, MT94b, Miy04, MLNW22, Mor92, MV06, MR06b, MP22b, NP95, PDG14, PD22, Per94, PPZ22, QH21, RR06, RR08b, Rog94, Roi07, SCZ10, SS19, SY98, Sil96, TKH09, TvH12, WW20, WPRS21, WH17, WL16, Wil04, YZ07, Yun98, ZY96, dSY05, dSDG10, vH09]. Markov-dependent [Col09, Roi07]. Markovian [ATV15, BZ08, BGT01, BK00b, ÇD16, DH04, DMP22, DGP07, EL10, FK99a, FZ02, FP15, Gur14, HP23, HMS04, PP96, PP22, Zhu15, dGvZ93]. Markowitz [CS13b]. Marriage [Pit92]. Martin [EF21]. Martingale [DS95, Jof93, ABP^{+}13, BVP22, GO19, HLOST16, HN19, IH23, JKM07, KX22, LM21, RS06, VZ19, ZW08]. martingale-equivalent [LM21]. martingales [AZ16, BJM10b, BT08, BMR08, Hob16, IRR12, KKN15, Lar14]. Maruyama [FG20, NT20]. MAs [RSX99]. Mass [Sch05a, Bor12]. master [WZ20]. match [BNS13, NR01]. matches [BKW08]. Matching [AW94, CGR09, DFP93, RT92, Rhe93, Tal92, Zha95, BL21, CRV06, Chi05, DFPR22, DS07a, MMX21, MP17]. matchings [ADRS23, OT23]. material [Pia99]. mating [AR96]. Matrices [AGD94, FHY92, Man93, BH22, BGBP23, CJL21, Chi16, CES23, CFMT11, DJ10, DY18, DJ23, DM23b, El 09, FK22, FRZ04, GM19, HLN07, HLN08, HK17, He20, HMY21, HNNZ13, HLS19b, Jia04, Kar10b, LS16a, LR06, NY16, Ngu17, Oli09, Ona08, PY14, SX23, Woo12, YY22, YM22]. Matrix [DGJ09, BLW11, Che22a, EGP16, Jia19, JWB^{+}14, KPT^{+}16, LS20, LLC18, LC22, Mou20, WJB^{+}15, dLS97]. matrix-valued [dLS97]. Maturity [BET05, LV03]. Max [DR93, JKS18, RR91, RR94, AB05, AKLT22, BH00, Bra10, ST04]. max-linear [AKLT22]. max-plus [BH00]. Max-Stability [RR94]. Max-Stable [DR93, RR91]. max-type [AB05]. Maxima [ACH97, ABH17, Adl00, BCHL98]. maximal [CT11, JB07, Pul14]. Maximally [BKS17]. Maximization [DPT01, Bar19, BH13, BZ16, BF08, BTZ04, FI20, HIM05, KMK10b, KR12b, KR21, Kni12, LŽ13, Nut12, RS05, Yu15, Žit05]. maximize [PPM18]. Maximizing [CS99b, HK13, DRS16]. Maximum [BG95, FL03, Mas95, AG06, AW05, BG06a, BG08, BG03, CMY03, DL08, DFdH04, EPQ01, FZ03, GZ00, Han06b, HLOST16, HK13, JJ15, KLZ23, Ott13, RW97, SW12, TK02, dTP09]. MaxWeight [Sto04]. may [Pit99, Sch01]. McGregor [BM97]. McKean [AKH02, AD20, BPR22, BLM23, CCD19, CDFM20, CGK^{+}23, CM19, DIRT15, FP22, HL17, HLS19a, KNRS22, STT19, TV03, Tom21, WZ20, dRST19]. MCMC [AM06, CGL^{+}15, DRVZ17, FMMP08, GK07b, ŁRR13, LW22, NR06, QH22, RR14, YR23]. MCMCs [AV16]. Me [HP23]. Mean [Ale94a, Ale95, Ana91, BD20, CDGL22, CRT23, DEH23, Dje23, DR91, JMSS12, LS14, MR06a, MP22b, Sch92, BJ22, BvdHH10, BvdH12, BvdHH17, BFJ06, BEH18, BdRGL20, BLM17, CF18, CS22, CL15, CZ16, CL22, CX16, Chi15, DOS19, DM23a, DR11, DFFN21, EOT05, FFK12, FHH23, Fis17, GM19, Gra09, GLWZ22, JLM15, KSS20, Kol22, KPSC10, Lac20, LL23, LMT12, LZ06, LS16b, LP21, LP08, MAL14, NST20, PSY15, Rey18, RS23, Tra23, YS22]. mean-based [DOS19]. Mean-field [CDGL22, DEH23, MP22b, BJ22, BvdH12, BLM17, CF18, DM23a, EOT05, FHH23, Gra09, JLM15, Kol22, LS16b, LP21, Rey18, RS23, Tra23, YS22]. mean-reverting [FFK12, MAL14]. Mean-Variance [DR91, Sch92, JMSS12, LZ06]. Means [Ale94b, DL10, FG18, Le15]. Measure [BB92, Del98, Din95, HS93a, Pov95, AN22, ABKR18, BNS21, BKT22, CFY05, DD10, DHS18, El 09, FRSW22, Fou15, Gol10, HNNZ13, JK14, KP04b, Lar14, LRS17, LC22, MV20, MRV21, MR12, PS14, Pan08, RS06, ST21]. Measure-valued [Del98, ABKR18, DD10, FRSW22, MV20]. measured [DL18b]. measurements [AR21]. Measures [CC93, DS95, KX95, Ass98, AC03, BY05, BR15, BL11, Bud02, Bur07, CST22, CRX21, CGR09, Com97, DAM10, Ebe14, ER08, FMN^{+}16, HK21, HK23, HS23, JMRS09, JLP08, JKM07, Jia08, Las02, Las04, Muk22, PP18, Pia99, PS17b, RSM09, Sch04, Wan23, vEK08]. Measuring [GDVM19, BNK12]. mechanics [ML16]. mechanism [Aus08]. Media [CF94, FM94, Bha99, CGM09, GS09, GLS18, Mal03, MS11, She02]. medium [CP09, SS18, Wüt06]. Memory [BM12, AGGL10, BH03a, CL14, CFF02, GRS00, Gra21, HRS97, JPS^{+}22, LN13]. memory-optimal [JPS^{+}22]. Mendelian [LN09]. mer [FLR22a]. mergers [KS16]. merges [FRT14]. Meromorphic [KKP12]. Mesoscopic [HK17, LS20]. mesoscopics [BDKT20]. message [BLM15]. metadynamics [JLZ21]. metapopulation [BP05a]. Metastability [DdHJN17, FGG93, JLM19, LY22, OT22, HMS04]. metastable [BCS19, FMN^{+}16, SS19]. meteors [BBPS15]. Method [AG93, DH93a, JM02, Jou02, Loh92, MPST02, Roo94, XS92a, XS92b, AKH02, BKH15, BK15b, BKN22, BD17, BN19, CCD19, CS11a, CS17a, CM14, ER08, FTW11, FG20, FK21, Fon10, FMMP08, GR21, GPR17, GLW05, GR97, HJK12, HJK13, IH23, KM11, Keb05, LS17b, LS19, Mél00, OTV12, ST21, TV03, Tri15, Xia97, Xu19, dSDG10]. Methods [DH91, DH92, DO94, EH95, FHY92, NP01, AGK11, BM22, BCJ14, BJKT16, Bro99, CL09b, DD10, DMO14, DMP96, GO19, HHK20, HLR22, HKV20, JLR03, ŁRR13, Nar16, QH22, TK02, WKRS19]. Metric [Grü14, But14, LO04, TSM19]. Metropolis [Béd07, BRS09, BRTP18, BR17a, Ebe14, FIKP98, GGR97, HSV14, HHH09, HJ23, Ing94, Jia15, JLM15, MZ14, MPS12, Mic02, NRY12, RR06, SV10, Tie98, VK21]. Metropolis-within-Gibbs [RR06]. Meyer [BB20c]. Meyer- [BB20c]. microscopic [FM04]. Microscopics [BDKT20]. microstructure [HX22]. migration [Bor12, Cox10]. Milstein [Yan05]. Mimicking [BS13, Hob16]. mimicry [LR21]. min [Bra10, Dd04]. Minimal [CS17a, JKM07, McD95, Ale96, CHS22, GTW22, KL96, Lee97, LvZ04, Pen96, Pen97, RS06, Yuk99]. minimality [CRSF23]. Minimax [ELS17, Gen23]. Minimising [RR14]. minimization [Nag12, Sto04]. Minimizing [Ale94b, JWW11, Pha02, HNS10, Sch02]. Minimum [AB92, Jai93, DFT03]. minor [CZ16]. minorization [QH21]. minors [CJL21]. Mises [BH22]. misleading [RSX99]. misspecification [Hob98]. Mixed [CS00a, GS05, BGHM10, CSS22, LM21, Mar19, ST04, VPV08]. mixes [Jon06b, LPS21, PS17a]. Mixing [AGvdHdH18, BBS11, GNS23, HKL^{+}19, Jia15, LL20a, LS18, MS21, MS12, MNP14, QS23b, Tra23, Wil04, BMST97, BGV20, CF07, Dre00, DGJ09, FK13, GK07b, GJ18, HS07, Lal00, LRdH98, MS18, Mor06, RR19, SZ17, TVVY12, WSH09]. mixture [HR04]. mixtures [AK05b, EG18, MR12]. mobile [Cla96, Sta15]. mobility [ST10]. modal [GK07b]. Mode [LSZ13, CS02, DH06]. Model [AB92, Ber92, BM01, BFM17, CvH07, CW93, Dev92b, DZ01, DS93, EK92, FP93, FGG93, KL01, KLS95, Kot92, LC93, MR94a, MW94, Mor92, dGvZ93, ADH20, AS04, AZ14a, AMP02, AS10, And98, AP17, Ata05a, AC17, BCH22, BMX17, BN17, Bal21, BK15a, Ban08, BFM11, BHL96, BEG00, BP05a, BR19, BZ10, Bar19, BS17a, BGY98, BK16a, BK17b, BČ05, BFHM20, BCPR15, BvdH12, BSZ15, BZ19, BJR08, BR17a, BNK12, BU18, BL12b, BK06, BC14a, BM05, BF05, BCS19, BN97, Bra99, BMW19, CDN02, Can19, Cer15, CDS23, CS11a, CDL17, CQ97, CS21, COMR22, CCL13, Cos16, Cou08, CS99a, CD02, CD03, CP08, Cox10, CP14, CPS16, Dai96, DRST09, DH07, DIRT15, DM06b, DR09, DZ15, EOT05, EKPS00, FK11, FS02, FMP00, FGG14, FM18]. model [Fox16, FEvdD16, FLR22b, FW22, GK21, GRS04, GK00, GVR17, GRS00, GS18, GM08, GLM18, GJ18, GL21, GTW22, Gup12, Häg99, HH19, HYTC20, HMQS20, HQR96, HD19, JK17, JS96, JS07, JKK03, JSW20, KK13, KW04, KP04a, Kne00, KP16, KP04b, LN06a, LN09, Lan12, LS13, LSZ97, LPZ17, LNS21, LS18, LTVR14, LP21, LN05, LN13, LP17b, MMPP07, MS12, MR06a, MPdS19, NP99, Nor16, O’N97, PP96, PR98a, Pia99, PW04, RR03, RR08a, RR97a, RZ23, RSS16, ST98, SSW06, Swa18, TV18, Tom21, Vys08, Yu07b, ZCD05, vdHMS08, vdHOC18]. Model-free [BFM17, LP17b]. Modeling [ÇJPY04, HLR22, Kar10a, Lac03, NS19]. Modelling [Aal92, BB20c, WH93]. Models [BZ91, CmHP04, Dai95, GW93, HW92a, Hog93, Ish93, Ish95, Jai93, JM91, KKN95, KLR91, LW92, Lig95, MR94b, Mey95, MD01, PQ01, Rog94, SV94, ADE18, AMS06a, AS09, ABDW21, AGvdHdH18, BBC17a, BHZ02, BH05, BFK05, BL08, BL12a, Bät99, BS23, BST04, BLZ11, BJN18, BBD22, Bha16, BCKWB16, Blo15, BP05b, BN15b, BD98, BN19, BF02, BKS20, CM03, CF09, CPSJ22, CDP18, CH19, CL11, CP18, CKW21, CFS18, Cla96, Cla22, CL14, CHS22, CGM09, DH13a, DM23a, DT05, DPR09, DR11, DM10, DFPR22, DK99b, DGMO11, DN97, DR08, Dur09, DP09, DM15, ER18, EG18, EEH14, FFK12, FZ02, FP22, FIMS22, FT17, FGP21, GR21, GM19, GQ03, GM05, Ger11, GP10, Gou07, Gou09, GG11, GPW09, Gur14, Har00, Har03b, Har06, HM14, Hei05]. models [HSZ17, HH08, Hof05, HX22, HLS16, IPB^{+}11, KMK10b, KKP14, KS16, LN06b, LP13b, LP17a, LM09, Löw98, MT13, MSZ19, MR06b, MNS16, Oha09, Pal11, PP22, PP15, PY18, RS03a, RS05, RR00, RS03b, RS06, RZ18, SZ06, Sch97a, Sha20, Sta97, SS08, SSY19, Ton08, Tri15, Wag05, WL16, ZHC06, ZBD05, vEK08, vH09, vdHvLS21]. Moderate [DGN05, FG08, LLLZ18, AB14, AC17, PDG14, SZZ21]. moderately [GL23, Xu18]. Modified [HS23, MW94, FM05b, LdRS18]. modularity [DS18]. molecular [HMS04]. Moment [DH93a, GK00, FK22, Las02, Las04, YY18]. Moments [ACLW95, BF04a, BC14a, Cha92, DR92, JSW20, SY94, Cho09, KRM15, NV03, NS22, Pia06, Pit99]. Monotone [Fri91, GRK05, BF04a, BG96, GZZ15, HK16, Tra23]. monotone-separable [BF04a]. Monotonicity [AJO14, DFK97, KM13, KS96, PZ19, ST98]. Monte [BREZ20, DPBCD21, MS21, ABL12, AV15, ADNR21, Atc10, BH20a, BM22, BK15b, BKN22, BCJ14, BJKT16, BRSS17, CL07b, CL11, DD10, Der11, DL16, DRZ16, DGMO11, DMO14, DG21, DG95, Egl05, EKT07, FHY92, GS14, GM13, GLW05, HLT21, HR04, HJK13, KKPvS11, MV06, Sad96, WKRS19, WPRS21]. Monte-Carlo [BM22]. Moran [DM15]. mortal [AGSC02]. mosaics [EN18]. most [ER10]. Motion [AGP95, DH92, Das95, ERY95, KP97, MRRS02, Pov95, Tak93, Tak95, ABT^{+}11a, ABT^{+}11b, ABK12, AC19, AG09, BC15, BH17, BDH10, Bra11a, BK98, Che96, CD99, CK03, CS17b, DeB04, DS17, EFdS21, EF21, EFPS17, EEH14, EHW20, FH98, FK99b, FK00b, GHP13, GZ00, IT99, Kah08, KL99, KLZ23, Leh23, Lej16, MG05, MG04]. Motions [KS93a, BB20b, BB23, Duf16, FK00a, FSW15, HLN21b, KW07, LT19, MR08, RS15, VZ19, Wik01]. motivated [KX22]. Mott [CF09]. mouse [LR12]. Move [Cha93, DF95a, DF95b, BF10, Jel99]. Move-to-Front [Cha93, BF10, Jel99]. Move-to-Root [DF95a, DF95b]. movements [Miy04]. moving [Che08, CSS22, DH13b, FKR96, MN22]. MRF [AG93]. Muller [PSW12]. Multi [Fri16, GS19, KM95, KM98, ABP15, AP16, AMR04, BFS21, BZ15, CP09, Chi04, DSS09, GS14, GW00, GR09, GK07b, KR96, Pap00, Wal09]. multi-allele [Pap00]. Multi-Armed [KM95, KM98, GW00, KR96]. multi-class [ABP15, GR09, Wal09]. multi-dimensional [GS14]. multi-interface [CP09]. Multi-level [Fri16]. multi-modal [GK07b]. multi-pool [AP16]. Multi-scale [GS19]. multi-species [BFS21]. multi-stage [DSS09]. multi-type [BZ15, Chi04]. Multiarmed [Web92]. Multiclass [BPT94, BGT01, Che95, DN94, Dai95, Mey95, AP16, APS19, AB14, AC17, CS00b, Dai96, GK09, HW96, Maj06, Sto03, YY18]. multicolor [Bha16, GGLO13]. Multidimensional [CLW94, Col02, DKZ94, RS15, BG00, BZ18, BCD17, BP97, CEK12, CDFO13, EPS09, HIP06, LS17b, LS19, LR13, MPP17, Pèn05, TY16]. Multifractal [BGHM10, HW92b, DJ12, Lud08]. multigraphon [RZ23]. multigraphon-valued [RZ23]. multilayered [FGLS23]. Multilevel [Der11, DL16, BK15b, BKN22, BGG^{+}16, CG16, GS14, HJK13, PP18, STT19, ST21]. multimodal [WSH09]. Multinomial [Loh92]. multipart [Ber96]. multiplayer [BC22a]. Multiple [BC18, HPTvD95, KLSY03, CK00, CHL21, DTW22, JKP13, KQRM11, LC03, MR06a, Pel98, PW23, Wik01, ZBM04]. Multiple-input [KLSY03]. Multiple-priors [BC18]. multiple-timescales [LC03]. multiplication [DJ23]. Multiplicative [LR00, BJM10a, BM19b, BDW22, DPRZ19, GH22, HN22, Har13, JSW20, Lac22, PP23, SW12]. multiplicity [BG03]. Multiscale [Bha99, EGK22, BKPR06, CW16, KKP14]. multiset [Fér20]. Multisource [DPS08]. multistopping [FR11]. Multitype [AV95, BNT92, CW03b, GG97, Nea06, Big12, CDL09, GLP23, Jon97, YY09]. multiuser [PGZ07]. Multivariate [BNS11, MS93a, RWW95, BDM02, BT00, EI17, FK22, HLMS05, Jia15, KZ09, KM13, SY23]. mutant [FGLS23]. Mutation [CA18, BCH22, BB03, DF06, DGP12, Tsu19]. mutation-selection [BCH22, BB03]. Mutations [DL18b, Dur13, GT99, PSW12]. mutual [MWZ07, PPM18]. mutualist [LN06b]. mutually [DKM17]. myself [HP23]. myths [CD02, CD03].

N [Ree09, DHT10, LvZ04, GG13a, BD17, GG13b]. Nagumo [EGK22, LS18, LP21]. Nash [ÇD16, DFM18, Dje23, Lac20]. natural [SY13]. Navier [BFH20, CI11, FHLN22, Fon10, HLN21a, Kot95, Mél00]. Navigation [Bor08]. Near [BHK05, BBD22, BR13, BC21, BdH99, LV03]. Near-integrated [BHK05]. Nearest [Yca93, YS96]. Nearly [CCK23, JR15, JR16, Pem09]. Necessary [KS03a, KLS11, LS95, Sha15, DY18]. needed [LvZ04]. negative [AKP04, APP07, FZ03, Loe08, MS00, RZ18, Sch01]. negatively [BK00a, GH21]. Neighbor [San08, YS96]. Neighborhood [FP95a]. nested [BRSSJ19, FRLS21, LRS17]. Net [FF94, Gar09]. Nets [Whi91]. Network [Bra10, CH91, GM95, MRRS02, AR19b, Ala03, ACG17, BS15, BGvdHK15, BJN18, BT10, Bra99, BG05b, CPV23, DHV04, DDMT12, FM05b, GO00, HW96, HW05, HMQS20, KKLW09, KW04, KST04, LRZ06, MW13, RSS16, SY98, ST10, Sto03]. network-based [ACG17]. networked [DIRT15]. Networks [Ana91, BPT94, BGT01, Bra94a, Bra94b, Bra94c, BD01, Che95, CMP94, DN94, Dai95, FRST94, HW92a, Kel93, Kel91, KS93b, MD94, Mey95, Ngu93, Ngu94, OW92, PR94, SV94, Yam95, ZZ02, ADE18, ARS18, AH98a, AKLT22, AFRT06, AP16, AK05c, BF04a, BKPR06, BOC22, Bäu00, BGZ97, BPT98, BSZ15, BC15, BBFM03, BdHNT22, Bra11b, BLP13, BCDS15, BG06b, BG12, CW16, CSS98, CS00b, Cra16, CDMR12, Dai96, DL08, DFH16, Dd04, DGM08, DSW07, FM05a, FM18, GH05, GZ06, GK09, GRS04, GGO23, GR09, GW09, Haj96, HV97, Har00, Har03a, Har03b, Har06, IR00, IS04, IV17, JM15, JLM19, KK13, Kel96, KLSY04, Kum00, LRW23, LW19, LLC18, Mag00, Maj06, MY96, MP98, MS99, Mar01, MSSZ20, OT22, PP15, SBF19, SS12, SW12]. networks [SWZ14, STZ14, SFR16, VAC15, Wal09, YY18]. Neumann [BQ19]. Neural [CPV23, Whi91, GGO23, LLC18, Tou14, Tou18]. neuronal [LS18, MSSZ20]. Neutral [EK92, Dur13, DM15]. neutron [HHK20, HKV20]. Newtonian [Bar20]. Neyman [Sch04]. Nicolson [HLN21b]. Nielsen [RS06]. Nielsen/Shephard [RS06]. No [BT13, Ciu98, JPS09, Oha09, CE10, GHLT14, KPT^{+}16, Mon22, SSC95, Web01a]. No-arbitrage [BT13, Oha09, CE10, GHLT14]. No-feedback [Ciu98]. nodal [MW07]. Node [Ana93, Ana94, GS05]. nodes [Sta15]. Noise [DO91, BG02, BGJ18, BCC^{+}23, BD12, Bud02, CS22, CF16a, CDFM20, CDFO13, DH04, DR96, DG08, DPRZ19, FJL18, GH22, GS09, GL23, HMGR00, JKW11, Kli19, KNRS22, LR19, LL23, LLLZ18, LØP04, LP21, MP22b, MGY23, PP23, Zha16b]. noise-induced [BG02]. Noisy [BG95, Ass97, BT12, CZ18, CPS16, JT18, WM04]. Non [MRV21, Ros95, dGvZ93, Béd07, DMP22, FP15, GS18, HHMG19, HP23, HJ23, LN09, LS16b, PP22]. Non-Gaussian [Ros95, HJ23, LS16b]. non-Hermitian [GS18]. non-i.i.d [Béd07]. non-Lipschitzian [HHMG19]. Non-Markovian [dGvZ93, DMP22, FP15, HP23, PP22]. non-Mendelian [LN09]. Non-universal [MRV21]. Nonabsorption [Gos01]. Nonasymptotic [DM17, MMS20]. noncolliding [NT20]. noncompact [MV20]. nonconserving [Hey23]. Nonconvex [YZ19]. nondegenerate [AG14, PSZ14, TvH12]. nondominated [BN15b, CC16]. Nonequilibrium [GLM17, Li18]. nonergodic [DGLM10, JK14]. Nonexistence [vEK08]. nonexpansive [Ban19]. Nonexponential [BVLT20]. Nonextensive [DMM17]. nonglobally [FG20, HJK12]. nongradient [BFG13]. nonhierarchical [BK06]. Nonhomogeneous [DM94, NP95, DS05, MW10, WL16]. nonindexable [Ver16]. Nonintersecting [TW07]. Nonlinear [BPT94, Cho09, EH95, HST91, HL01, KLS97, Kot92, Yan20, BDG16, BP05b, BPS04, CHMK20, CL03, CL04, CvH07, CNV22, CJK19, CD99, Del98, DG99, DD10, DW98, EW03, FTW11, FW99a, GZZ15, HSV07, HLT21, HJK13, INPY13, JKW11, KLP15, LPSX23, LO04, LYZ19, Loe18, MM01, MP06, MR08, NZ15, Ren16, SZ06, See20, ST21, Tor16, Zhu15, dBG12]. Nonlinearity [Cho02]. nonlocal [CH19, GMŚ19]. nonlocalised [LR19]. nonmixing [CL04]. nonmonotonic [Moy15]. Nonmonotonicity [LRZ06]. nonnegative [Zer98]. Nonnormal [Ber94, CS11a, SZZ21]. Nonparametric [AR21, CS23, Tod19, LPW08, LPW14]. nonparametrics [WHN07]. nonpolygonal [DH13a]. nonpositive [KLP15]. nonproduct [BRS09, Kel96]. nonreconstruction [SZ17]. Nonreversible [Ass98, Fil91, DHN00]. nonsingular [MR06b]. nonsmooth [BTZ04]. nonspherical [AK05b]. Nonstandard [CC93, HK22, Puh15]. Nonstationary [MW94, Niu97, ZY96]. nonsymmetric [Che22a].

Nontrivial [SSC95]. Nonuniform [IT12, OP00, JPV99, Sad98]. nonuniqueness [Nor99]. nonzero [DFM18, IV17]. nonzero-sum [DFM18]. Norm [Blo92, Cho09, Kar10b]. Normal [Ber94, Gol04, GP10, LRSY19, LR19, MR93, Pit99, BC09, BMR08, CT23, FK22, HHR96, JL09, LNS21, LC03, SY23, Yat09]. Normality [LZ98, JL08, MP09]. normalized [BT08, Coh04, DRS16, RWF13]. normalizing [Hub15]. norms [DGJ09, HPS03]. Note [Ale93, Alb09, Bor16, LS19, LPW14, Tie98, Web01a]. notions [GHK11]. Nuclear [KX95]. Nucleation [GS17, GW22, HQR96]. Null [AMS06b, Sha15]. Number [ASG93, EH95, GY04, MZ91, Pet91, ABT00, BCHL98, BS15, CDL17, CQ97, CX02, DFPR22, Fil13, FP17, HH18, KP96, KP98, MR17, TLC93]. Numberical [MPST02]. Numbers [Blo92, Fin94, Pet91, BL08, BF10, BC22a, FM18, KR11, Owa18, PY03, RR97b]. Numéraire [Kar10a]. Numéraire-invariant [Kar10a]. Numerical [CR16, CCD19, DH91, DH92, DMP96, DO94, Mor05, PRW95, Ric11, dSDG10, BdRGL20, CL22, FTW11, HNS11, HJK12, PS16, Zha04].

oblique [CEK12]. obliquely [EF21]. observability [vH09]. observables [MSZ19]. observation [CLP16, CDFO13, DS15]. observations [BLM17, BT12, KK04, Ste99]. Observed [CG92, KO92, PQ01, BCFP18, BGHK20, CLR06, DRS16]. observing [Mat05]. obstacle [GIO^{+}17]. obstacles [DS17, PS20]. occupancy [BGL02, EF21]. Occupation [ABT^{+}11a, ABT^{+}11b, Din95, Ben96, Web01a, Web01b]. Occupied [MR94b]. O’Connell [NS22]. ODE [FMMP08]. off [HMY21, JM03, MR17, PR98a, RS03b, ZBM04]. off-diagonal [HMY21]. Offered [MW94]. Offered-Load [MW94]. offs [DSC06]. Old [CDP18, CD02, CD03, NP05]. omega [Har13, RZ18]. on-off [PR98a]. on/off [RS03b]. One [AAK17, AGP95, BS17b, CMP94, DR08, GdH93, Jou02, PP08a, Pov95, RT92, Roi07, AZ10, AJKH14, BBC17a, BC02, BL12a, BCMR21, BČ05, BJ22, BJR16, BDKR19, BR15, BT17, CF09, CG17, CGR09, DH23, DI17, DM23b, FRT14, FSW15, FP17, Goe06b, GZ09, GLM17, HL21, JKK03, Kol17, LS13, LM06, Tur18, Vys08, YZ19]. one-armed [GZ09]. one-bit [DM23b]. One-Dimensional [AGP95, CMP94, GdH93, Jou02, Pov95, AAK17, BS17b, DR08, PP08a, Roi07, AJKH14, BC02, BL12a, BJ22, BDKR19, BR15, CF09, CG17, CGR09, DH23, DI17, FRT14, GLM17, HL21, LS13, LM06, Tur18, Vys08, YZ19]. one-sided [BCMR21, BJR16, FSW15]. onto [HR97]. Open [Yam95, APT20, AK05c, DeB04, GNS23, Har00, Har03b, Har06, HSH^{+}13, MP22b, SXY21]. open-loop [MP22b, SXY21]. opening [RZ08]. operating [KKLW09, KW04]. operator [Bät99, BDW17, DL22, TKH09]. operator-scaling [BDW17]. Operators [Too02, Mer07]. opinion [CF11]. opportunities [BP05b]. optical [dBG12]. Optimal [AFKP20, AM10, BMX17, BR01, BHRS23, BK16a, BK17b, BGBP23, BRS09, BP05b, BPS04, BR08, CT23, CPT12, Cla18, CGK^{+}23, Dar21, DOS19, DK08a, DO94, EPZ20, FK11, FS99, FHY92, GK95b, GK96, Gua02, HS93b, HS93a, HK92, HP92, HK04, IM10, Jan01, JKP13, JLM15, KTT17, KQRM11, LP04, LM09, MS20, MG04, MP20, NR06, NRY12, NZ15, Ott13, PQ01, PST12, Sch01, SWZ14, SS94, Tou00, XS92a, XS92b, XZ13, Yu17a, vdHvLS21, AJKH14, ARS17, AJO14, AC17, APP07, APP15, BVP22, BCFP18, Bel11, Bel13, BC18, BL21, BDL16, BG05b, BR06, CRT17, CCHH05, Çet18, CMY03, CCK23, CFJ16, CHO08, DP20, DW05b, Egl05, EV06, EJ16, FG18, FK10, FS02, FP15, GGS03, GGR97, GO12, GL14, GK03, GIO^{+}17, GO19, GL21]. optimal [HH08, HKK06, JPS^{+}22, Käl22, KS14, KMPT10, KS99b, KS03a, KS06a, KX22, Mon22, MNS16, MG02, NZ20, NW22, Owe02, Pem09, Pes19, RZ18, RT14, RU08, SXY21, TSM19, Ver16, VK21, Wee98, dSDG10]. Optimality [Cha93, KLST93, RS01, AK05c, AG12, AMR04, Ata05b, AG14, Bäu00, BEZ20, BW01, CW13, DL08, ELS17, GR15, GR06, Har98, HDP22, Jaś07, KR96, Kum00, Loe08, Mag00, Sto03]. optimisation [CS16b, CS17b]. Optimization [BPT94, CK92, Fra02, HL97, NP02, ABDW21, BM22, Bec06, Bel11, Bel13, BLY22, BT10, Cal97, Cao21, EPQ01, KMK10a, KMZ17, LLS08, MP21, Nar16, NZ20, PS17b, Yuk96]. Option [ADGS98, Aka95, KLR91, Kus95, Loe18, Myn92, SS93, SSC95, EH11, FK16, GY04, Hob98, JT03, Keb05, LV03, RZ99]. optional [Kar15]. Options [Das95, DH93b, AKP04, BZ16, BT22b, BL02, CKL22, CW13, CK19, CM96, DFT03, DK08b, DK08c, Dol10, Dol13, EKT07, Fus00, GHLT14, Kif06a, Kif06b, KP03, Lam98, LS97, MPZ21, Tod19]. Order [ABBJ94, BENP91, BB92, BL11, GK95b, LT18, Rhe00, Tak92, AA13, AV16, AD20, BGT22, BHQ17, BJR16, CS22, CGZ14, CL17, CM14, DWZZ20, Fla97, Gol16, HLS16, KTPZ15, LNS21, LS17b, LS19, LL12, LT19, MPZ13, MPZ21, Mül18, PZ19, PT15, RS03a, STZ13, Swa18, Yuk15, Yun98, dLS97]. Order-invariant [BL11]. Ordered [PY18, LMT96]. ordering [Pen00]. Orderings [MS93a]. orders [DM02]. Organizing [Cha93, BF96, Sad98]. Oriented [KSS20, Lig95, BGR22, GRS04]. origin [Jac02]. Orlicz [BF08]. Ornstein [CW93, DG13, SSV18]. Orthant [DH92, Che96, DG14, McD99]. oscillating [Fed14, Tri15]. Oscillator [Pin92]. oscillators [BGJ18, Cop22]. oscillatory [CM05]. other [BG21, BKS20, CCM06, CAP20, Pap98]. Outperforming [BHS12]. Output [FF94, KLS95]. overdominance [GS02]. overdominant [JKK03]. overflow [HRS97]. overhand [Jon06b]. overload [AS04]. Overrides [FH95]. Overshoot [Cha94]. Overshoots [DK06, KKM04]. overtaking [KR96].

Packing [PY02, RT96, BG01, CFJ00, Rhe00]. Padé [JS12]. PageRank [BOC22, GvdHL20]. PageRank’s [OC21]. painting [LPS21]. pair [CGZ14, Röl07]. pairs [CS11a, FK22]. pairwise [BCH22]. Palm [CRX21, ARL08, Gol10, LS00]. pancreatic [MD18]. paper [HL21]. parabolic [BC14a, CDS23, FTW11, GK00, GJKS15, GMŚ19, HL00, JKW11, Tom21, WZ20, vdHMS08]. parabolic-parabolic [Tom21]. paradigm [DL10]. Paradox [BK92, KMPT10, EL10]. Parallel [FFS93, Kel93, Ngu93, AG14, BW01, GW00, Har98, WSH09]. Parameter [BB20b, CG92, FKR96, BM20, BU18, CDE^{+}17, EFT07, FRSW22, JLP08, MMPP07, MMPP08, Tsu19]. Parameters [AG93, IR01, AW05, KM18, LW14]. Parametric [WH93]. parametrix [BKH15]. parasite [Sch97a]. parasites [Ban08]. parasitic [BHL96]. paraxial [GS09]. Parking [CMP94, DGJ^{+}19, GL01, PRS23]. Parrondo [EL10]. Part [PRW95, HSV07, MP98]. Partial [ABBJ94, BB92, NP02, Sta03, BLY22, BEM08, BNS13, BLM17, CLP16, ÇJPY04, DMP22, DZ16, DM23b, Ino02, LS15, LØP04, NR01, Ren16, See20, Zha12, Zha16b]. Partially [CG92, FFS93, KO92, PQ01, BCFP18, DRS16, NR06]. Particle [BL01, DGM20, EW01, NS19, Yca93, AKH02, AC03, BJ22, Ben96, BZ18, CGR20, CF16b, Del98, DG99, DT05, DG05, DPR09, DR11, DH23, DPBCD21, DGN05, EW03, FJ17, GGLO13, GLWZ22, GL23, Hey23, JK14, JS07, KS05, LO04, LT22, MG05, NT20, OC11, PP08a, PS14, RvH15, Shk11, SS15, STT19, TV03, Whi13, YS22, Yu07a]. particle-survivor [JS07]. Particles [BdRGL20, Che01, ABL21, Bar20, BT96, BGZ22, CGR20, CGZ14, CF16a, EP10, GKS04, HH18, IK10, JM08, KS16, LY17]. partition [Ber10, GK21]. Partitioning [ZZ02, Pen00]. partitions [AS23, DGP07, FLP13, Fér20, FRLS21, SD05]. partner [Fox16, FEvdD16]. parts [BBM07]. Passage [Ale93, BT19, FP93, FN93, FN94, GK95a, KS93a, Kes93, Rhe95, vdBK93, AK05a, Asm98, ADH15, BGT22, BvdHH10, BvdHH17, BT12, CT11, CZ03, CN23, DH13a, DHL18, DP15, DFK97, EK09, EJ16, EHW20, FS22, FMS14, FKP23, GM05, GM08, Gri13, Hof08, KSS20, KPR10, Mar02, Mar16, McD99, PW97, ZS09]. passing [BLM15]. passive [BGHK20, KK04]. Past [KLR91]. pasting [AK05a]. Path [Aka95, AZ14a, Das95, GM12b, GL21, IR01, JR92, BPR22, BCFP18, BBRZ20, BGvdHK15, BCD17, BT22b, Bra11a, BdHM23, BH12, CZ18, CST05, CGK^{+}23, Gri16, HSV07, HKL^{+}19, HDI16, Kel16, Maj06, MvU05, Mat05, Muk22, Pem09, PTZ20, Ren16, Tan14, VZ19, Wis01, dRST19]. Path-Dependent [Aka95, Das95, BPR22, BCFP18, BT22b, CGK^{+}23, PTZ20, Tan14]. pathogen [LN06b]. paths [ABKR18, BG02, DFM16, GY04, GM08, HM14]. Pathwise [AJKH14, AC20, CDFM20, CHO08, DFM16, Hey19, CG17, Fon10]. pattern [CRV06, PS97]. Patterns [HRS97, KL91, Coh96, Erh00, Fér20, Löw98]. paying [DFT03, LV03]. payment [AM10]. Payoffs [Lan15]. PCA [JLM20]. PDE [LLLZ18, dBG12]. PDEs [DM06a, FTW11, GZZ15, LdRS15, VZ19]. peacocks [KTT17]. peak [Cer15]. Pearcey [CM23]. Pearson [GPR17, Sch04]. penalized [JS96]. penalties [Cha05]. penalty [APP15, IS04, Lei08]. pendulum [FW99a]. Percolation [Ale93, BT19, FP93, FN93, FN94, Hol01, Kes93, Lig95, Pen93, vdBK93, ABA20, ABGK12, Ale96, ADH15, BBW04, Bal21, BGT22, BLZ11, BZ15, BB15, BvdHH10, BvdHH17, BBFM03, BCDS15, CT11, CS00a, CZ03, CN23, DH13a, DHL18, Dar21, FS22, FMS14, FKP23, GM05, GM08, Gou09, GS19, GH08, GHPS15, GS17, GS20, GHS21, Gu22, HHJ17, HSST23, HHK21, Hof08, Hol98, JŁTV12, KSS20, Mar02, Mar16, MZ05, Mül15, Pen96, Sta15, vdB11, vdBN17]. Perfect [CK07a, CK07b, CT01, ER18, Hub04, WL16, ADRS23, CFF02, CS99b, Fil98, HR04, MLNW22, OT23]. Performance [BPT94, BGT01, DFP93, MZ14, OW92, BM12, Cal97, HRS99, MSZ19]. performances [Žit09]. perimeter [BD20, BP18]. Periodic [BG05a, BC22b, BCJ22, Bha99, CGZ07, EP10, HI05, HIP06, PPZ22]. Periodicity [LP21, MMPV06]. periodogram [KM96]. Perlman [Yu17b]. permanent [FRZ04]. Permutation [Fri91, BM17, HNNZ13]. permutations [BUV11, BSZ20, BR15, BS20, Fér20, Jin19]. Perpetual [DH93b, KP03, EH11]. Perpetuities [HW09, HW10]. Perron [BL16a]. persistence [BB22, CS99a, HST18, Owa18, Owa22]. persistent [BKS17]. perspective [Leh23]. Perturbation [Ber94, Dec98, HLN21a, SZZ23]. perturbations [BB23, BJ20, CP14, Ebe14, Fed14, FW99a, KK01, KQ18, May09, MM10]. perturbed [AG06, BYZ00, BGJ18, BM04, BE23, CV21, CY21, FS14, FHLN22, HPŠV04, KLZ98, YZ07, ZY96]. Ph [BD17, DHT10]. Phase [AW94, CHL21, HSST23, HMS04, Jár17, MR94a, MMX21, O’C91, RY13, TVVY12, Tur18, BCOR16, BLM15, BEMT21, ČT16, Com97, EOT05, FMP09, Fou22, FKKM18, GNP17, JLM15, JL22, Lac17, LRZ06, MRV16, Pit08, RW12, ST98]. Phase-Type [O’C91]. phenomena [Wag05]. Philosopher [Yca93]. Phylogenetic [MR12, BFJ06, DR13, DI17]. phylogenies [MR06b]. phylogeny [FLR22a, MV06]. physical [Pia05]. Piecewise [BGT01, CC98a, DY95, ADNR21, BR17a, BKR22, Cos16, CDMR12, DG13, KW07, LW22, TK02, dSDG10]. pinned [BC02, GL18]. Pinning [AS06, AZ14a, Lac17, Ton08]. pioneering [MZ05]. Pitfalls [Taf11]. Pitman [DF20]. pivotal [MZ05]. Place [ESS93]. planar [BLMZ14, BCF18, BT10, Sep97a]. Plane [KO92, ADH20, BM05, EF21, MM01]. planted [MMX21]. plaquette [CS21]. player [CF18, DFM18, Fis17]. players [BEM08, CZ16]. plug [MP09]. plug-in [MP09]. plus [BH00]. Plymers [KV01]. Poincaré [FW99b, GLWZ22, JSTV04, JM08, SS19, Tra23]. Point [Ana93, Ana94, Eic95, HMY21, JM91, Mar01, BB07, BBC23, BH20a, Bät99, BS05, BJN18, BR15, Bor08, BG96, BKT22, CY21, CF07, Chi05, CCL16, CDV14, CFJ16, DTW22, FLÓ19, GP10, HH19, JS07, KLS11, KS96, LRM15, LRP17, LR21, LP13a, Nar16, PY13, Sch05b, Tou00, TBUA23, YA15]. points [AZ10, DS19, DSZ15, GH21, HDI16, KL96, Mån99, MG04, Sep97a]. pointwise [MG04, PD22]. Poisson [Aal92, ABGK12, ABF13, AS10, ACH97, AZ14b, ABT92, BS96, BB07, BG93, BC01, BDK06, Bor08, BG96, CR18, CQ97, CX02, CCL16, CDE^{+}17, Cou08, DF06, DK92, DGM06, DGPR19, EN18, ELS17, Erh00, FS16, Fen07, FG08, FP22, FRSW22, GR21, Gap05, GS19, GG97, GS05, Hei05, HSS06, HR09, HJJ16, IM10, JLP08, JS07, JKK02, KX95, LRPY22, LSZ97, LNS21, Loa92, LM21, Mån99, Pel10, Röl07, Roo94, Tsu19, YS96]. Poisson-driven [BS96]. Poisson-like [ACH97]. Poisson-skip [LSZ97]. Poisson-weighted [ABGK12]. Poissonian [ABK12, FJ19, Ken11, Pia99]. polarity [Gup12]. Policies [BENP91, HP92, AK05c, Bäu00, DL08, FL96, GK09, GW00, Har98, Mag00, RS07, SW12, STZ14, Ver16]. policy [AM10, BW01, Dd04, HDP22, KKLW09, KW04, VPV08]. Polish [GR06, GS11]. Pollard [KMPT10]. Polling [CPR95, KS92, MMPP08, FL96, HDP22, MM03, MMPV06, MMPP07]. Polluted [GHS21, GK23]. Pólya [BT22a, CPS11, CCL13, MV20, vdHHKR16]. polygons [HRW08]. Polyhedral [BPT94, LR18]. polymer [AZ10, CP09, FJ19, KQ18, MN22, NS22, Viv23]. polymers [AS06, Cla22]. polymorphism [AS10]. Polynomial [Cho02, CLSF18, JR02, LY17, KZ09, Li18, LdRS15, LdRS18, Mek15]. polynomials [BGZ22, LP23]. polytope [BLM15]. polytopes [BR04, EN18]. pool [AP16]. pooling [AK05c, BW01]. pools [SY13]. Population [BK95, BNT92, CW93, Dur13, EK92, Mor92, Ngu94, dGvZ93, AdHR21, BBT17, BL06, BCS19, Cer15, CMT21, Cla96, Eth04, ER10, FH98, FM04, FLR22b, Jag97, Kel13, LS16a, Sta97]. populations [BHP10, BEM07, CL07a, Dur13, DF16, HW07, SJ05]. porous [Bha99]. Portfolio [ASCDH09, CK92, CS17b, HXZ22, HP92, NP02, PQ01, SSC95, BHS12, BLY22, CS16b, DRST09, GK03, HSX22, KMK10a, LŽ13, PS17b]. Portfolios [CK93, GR12, LS95, CT04, ET05, GSS13, MSZ19, Taf11]. posedness [CDG23, FHLN22, KNRS22, MWZZ15]. position [AFKP20]. Positions [McD95, ARS17]. Positive [AMS06a, AGD94, BDH10, Dai95, DG13, Häg99, MS93a, BCOR16, BDW23, BK16b, Bra11a, BGZ22, Che96, CK07a, CK07b, CFMT11, FKR96, FPRW18, HLS19a, Kol17, MS18]. positivity [FKM96]. possibility [LS97]. possible [DGR09]. posterior [MMN22]. potential [Bal19, BG02, CDL17, MPdS19, YZ19, Zer98]. potentials [AS06, Gué03, MRV16, Xu18, vdHMS08]. Potts [Häg99, JKS18]. Power [JM15, HP15, Jan08, LM05, Nut12, PS20, TY11, TT11]. power-law [PS20]. Practical [DFMS04]. Precise [FGP21]. Precision [CS95]. Predator [Als93, Sch97a, Cos16]. Predator-Prey [Als93, Sch97a]. predicted [Hey23]. Predicting [Ste93, Ste95, Ste99]. Prediction [DR93, Ros95, BEZ20, WM04]. preference [Laz04]. preferences [Kar10a]. Preferential [BJN18, JW21, BOC22, CS13a, CCL13, JM15, JK22, PRR13]. premalignant [FGLS23]. premia [GR12]. prescribed [EEH14]. presence [APP15, CD99, DM06b, Dol13, EV12, HN19, LS97, MM10]. preserved [GO98]. pressure [DL08]. Prey [Als93, Cos16, Sch97a]. prey-predator [Cos16]. price [BD19a, BHQ17, BF04b, BT22b, CHMK20, DH04, EH11, GRS08, HP15, JT10, Kle03, KP16, LV03, Num00]. Prices [DG95, JM02, RS91, BK15a, CLR06, CS17b, EH11, JT03, KMK10a, KS06b, RZ99, RZ08]. Pricing [BJR08, Cha99, Das95, DH93b, KLR91, Myn92, SSC95, AL17, ARS17, DM06b, EKT07, FK16, GY04, GRS08, HS19b, Hob98, KK96, Keb05, Lei08, Loe18, Pul14]. principal [CJL21]. Principle [CPR95, BN15a, Ber97a, BdHM23, BCP11, Cer09, CKW21, EPS09, EPQ01, KW07, Pia05]. principles [AK05a, BDW23, BFG13, BDW17, DAM10, DWZZ20, FH98, PW97]. prior [LRS17]. priori [CM08]. priority [ABKR18, GK09, MvU05, Ver16]. priors [BC18, HJ23]. Probab [Tou18, Web01a]. Probab. [BvdHH17, LO17, Ano99, Ano02, Ano03]. Probabilistic [Bro99, FKP23, FH95, FP95b, HPTvD95, Jou02, MW07, Rhe94, BKH15, BL12a, Ber02, CL15, CZ16, CH19, Con22, FTW11, FKM96, Fou00, FM04, Gué03, JLR03, MR05, Tan14]. Probabilities [ACLW95, BMS02, GN91, Loa92, AR96, Asm98, BC02, CL03, CL07b, CL11, CDS09, Fuh04, GGS03, GW97, HPŠV04, Jel99, Kah08, KKM04, KS03b, LS00, Miy04, PW23, Sad96]. Probability [AR02, AB92, AB93, Din95, ESS93, BY05, BFRT18, BHS12, CSX23, CX16, CS99b, FPZ05, HNS10, JM03, KPT^{+}16, Lac03, MN97, PY03, PS20, San08, Sch02, SY23, WH17, XZ13]. Problem [Col02, DZ94, DPT01, DO94, KS93a, KZ94, LW92, Pit92, Rhe93, Rhe94, AS16a, AJO14, APP07, APP08, BB20c, BDK06, BQ19, BT00, BTZ04, BLM17, CP19b, CCCS11, CC98b, CvH10, Cla18, CRSF23, DS15, DP15, DE17, DH18, DW05b, DSS09, EJ16, ES03, EHW20, Gap05, GZ09, GK03, HH08, HI05, HK13, Jac02, JLR03, Jár17, KRSS23, KLS97, KQRM11, KS03a, KS06a, KX22, LRM15, Lef04, Loe08, MMX21, Mül18, Pit99, PZ19, Pul14, Rhe00, RZ18, Sch04, Sch13, Sez10, VPV08, ZRH15, ZS09]. problèmes [Mic96]. Problems [EK94, FM94, Hog93, Pha02, AKP04, BVLT20, BMX17, BK17a, BS05, Bel11, Bel13, BDM17, BF08, BET05, BMR08, BR06, BR08, Cao21, CC16, CS23, CPT12, CS13b, DP20, Del98, Dd04, FR11, FG18, Fer15, GK96, GMŚ19, Gu22, GO19, HLT21, HT12, HPS03, IH23, JMRS09, JP17, KMZ17, KR96, LST23, LM09, Ott13, Pen00, RU08, STZ13, SXY21, Tan14, Yuk96]. problems-Exploration [CS23]. procedure [JS96]. procedures [Ber97a]. Process [Als93, ABT92, BK01, CW93, CGS93, FF94, Fil91, GL01, KOOS91, Lef91, QS94, Yca93, dGvZ93, AR96, ADNR21, ABK12, APP07, APP08, APP15, BB07, BC16, BG01, BH01, BH03b, Ber10, BJR08, BRZ19, BNS21, Bor08, BGL02, BH17, BG21, BEMT21, BNS13, BS13, BGR22, BM96, BF96, CV21, CF07, CDL09, CM23, CD11, CA18, CSX23, CCL16, CG17, CNV22, CFJ16, CS99a, DR13, DDMT12, DJ12, DHS18, DM05, DG18, DLS03, EH11, EV12, EFPS17, FGG14, FN17, GNS23, GM12a, Gol10, Gra09, Gra21, GLP23, GM12b, Gri16, GKS04, GZ08, HMT22, HMY21, HS19a, HS21b, Hey19, IM10, Iva18, JLM19, KS99a, Kar13, KSS20, KR23, KST04, Kro99, LL20a, Lam05, LP13a, Le15, LT22, LYZ19, MRS01, Mar08, Miy04, MS15, Ott13]. process [Pan08, Pap98, Pap00, Per00, Pop04, QS23b, Rem08, RW12, Sch05b, SYY20, Tra23, WHN07, ZS09, dLS97, vdB11]. process-Monte [ADNR21]. Processes [ASG93, AR02, Ath94, AV95, BMS02, CT91, CJ94, CFF02, DR93, DK92, DKZ94, DY95, DM94, Eic95, GK95a, Gla93, HR94, HPTvD95, JM91, KM95, KW91, Loa92, Mas95, MP01, MS93b, NP01, RR91, RR94, SY94, Ste93, Ste95, Tak92, Wor95, YH93, ACD15, AS97, AA13, AH98b, Alb04, ACG17, AK05a, AMS06a, APT20, Asm98, Ass98, AdHR21, AKP04, BBBW22, BBC23, BDMT11, Ban19, BH20a, BNS11, BS12, BN15a, BBRZ20, BBL^{+}97, BS05, BST04, BBL14, BFS21, BLZ11, BZ15, Ber97b, BDKT20, BBDW22, Bha99, BM04, BD12, BL02, BG21, BH19, BCL06, BG96, BD15, BC09, BR13, But14, CM03, CS18, Cha99, CD06, CFY05, Chi04, Chi05, Chi07, CC98b, CPT12, Cla18, CS06, Coh96, CDV14, CS11b, CL09b, Cra18, CDMR12]. processes [CFMT11, DR02, Das96, DR96, DP15, DH13b, DGC20, DJ12, Del98, DFH16, DK99a, DMM17, DG13, DG14, DKP22, DTW22, DK06, DK99b, Dre00, DFS03, DW98, EK09, ELS17, ES03, ESU10, FRT14, FL22, FLÓ19, FPRW18, FR05, Fou22, FKKM18, FJ17, FEvdD16, FJR20, FL03, Gap05, GP10, GLM17, Gri13, GR06, GS11, Haa10, HK16, Har13, HS21a, HH18, HKK06, HPS03, HPŠV04, Ino02, ILP15, JLP19, Jag97, Jag99, JR15, JR16, JKM07, JM03, JP17, Jon97, JS07, Kab10, Kar15, KRM15, KRSW19, Kle03, KKM04, KM03, KS06b, KLS11, Kuz10, KKPvS11, KKP12, KS96, LRM15, LRP17, LRW23, LR21, LS17a, LSZ13, LS00, LNS21, LRT03, Lef04, LM06, LZL21, Loe08, LPP15, LW22, LMT96, LP08, LM21, MV20, Már97, Mar01]. processes [Mek15, MW10, Miy04, MW08, MP22b, Mül15, NV03, NV04, NV06, Olo96, PR98b, Pau02, PW21, Pel10, Pèn05, PS05, PY13, PPZ22, PY16, PVY20, PW97, RS15, RZ23, RW97, RWF13, San10, SY98, Shk11, Sil96, ST04, Sta03, TT11, Tor16, Tou00, TBUA23, WPRS21, Web01a, Web01b, YY09, YS96, Yao97, YA15, Zha12, Zhu15, dSDG10]. Processing [Ngu93, AHS05, AK05c, BBP22, DL08, Har00, Har03b, Har06, Puh15]. processor [GPW02, Gro04, GK07a, LSZ13, PW04, RR03, RR08a, ZDZ11]. processor-sharing [LSZ13]. Product [HW92a, BFRT18, GS20, RS03b]. product-form [BFRT18]. Production [KLST93]. Products [AGD94, Kar10b, Man93, Too02, DD09, HS21b, Kol17]. Profile [CDJH01, DH06, DJN08, HLN08, MG05, Sch10]. profiles [KMS17]. profit [Mon22]. progenitor [CK00]. Programmed [BOW95]. Programming [HS93b, CRT17, Käl22]. Progressive [LR14]. Prohibition [XS92a, XS92b]. prohibitions [Pul14]. Proliferating [Ban08]. Promiscuous [AR02, AR96]. Prone [KLST93]. Proof [ERY95, Tsi94, BL21, BE23, KR06, Mél00]. Propagation [CF16a, CRSF23, DP19, HP20, JM08, MSSZ20, MD01, Tou14, Tou18, Bar20, CST22, CF17, COMR22, CF16b, DT18, KS14, MNS16, OC11, Xu18]. properly [JB07]. Properties [KOOS91, Ste93, YM22, AZ14a, Alv03, AM06, AV15, BYZ00, BLW11, DT18, DR98, FKP23, GRK05, Gra21, HLS16, JT03, KST04, KS96, LNS21, LdRS18, Mas07, MR08, MN22, RR98b, San10, STZ14, She02, WKRS19, Whi13]. Property [HW94, GO98, JMRS09, YS96]. prophet [AGSC02]. Proportional [Wal09, BT13, Mas07]. Pseudo [BG95, Mas95, AV15, LR14]. pseudo-honest [LR14]. Pseudo-Likelihood [BG95, Mas95]. pseudo-marginal [AV15]. Pseudolikelihood [JM91, JS96]. Pseudorandom [EH95]. pseudotrajectories [BBC17b]. Pulse [CF94, EGK22]. pure [TT11]. pure-jump [TT11]. Put [JM02, AK05a, Lam98]. puzzle [BCDS15]. Pyramids [Mah94].

q [KP04a]. QNET [DH93a]. quadrant [APP08, PW21]. quadratic [BM04, BMM21, CR16, DG13, ESX22, FHH23, HS99, IRR12, JMP21, KP16, PP08b, Ric11, SXY21]. Quadrature [Kee94]. quadtrees [BNS13, NR01]. Qualitative [STZ14, LdRS18]. quality [GDVM19, KPSC10]. Quantifying [DGL23]. Quantile [Das95, ERY95, JZ11]. quantiles [Das96, LS09]. Quantitative [CF16b, DMR04, GM13, LRPY22, NRS21, CST22, YR23]. quantization [BP97, LP08, YRF16]. Quantum [Kol22, CM03, CN11, ML16]. quarter [BM05]. quartet [HLR22]. Quasi [FS14, AC03, BBL^{+}97, BCP18, CMY03, DM06a, DRZ16, FHH23, FMN^{+}16, FKM96, KLZ98, KST04, MSW97, SSV18, WKRS19, Zha12]. quasi-birth-and-death [BBL^{+}97, KST04]. quasi-exchangeability [FHH23]. quasi-linear [DM06a, Zha12]. quasi-Monte [DRZ16]. quasi-optimal [CMY03]. Quasi-stationary [FS14, AC03, BBL^{+}97, BCP18, FMN^{+}16, FKM96, KLZ98, WKRS19]. quasi-tilings [SSV18]. quasi-Voronoi [MSW97]. quasilocality [JK17]. Quasireversibility [HW92a]. Quenched [BMRS23, BC22a, CKW21, ESTZ13]. queries [BNS13, NR01]. Queue [CLW94, AR19a, AMR04, Ata08, AB14, AC19, BM19a, Bla96, Bra11b, BLP13, BFW21, CBM^{+}21, DM08a, DLS01, Fla97, FM01, GHR03, GG13a, GG13b, GO98, GZ00, GPW02, Gro04, GK07a, HRS99, JM03, Kne00, Lim01, OCBG11, PW04, Ree09, RT15, SWZ14, BM20]. queue-based [CBM^{+}21]. queue-size [SWZ14]. Queueing [AMS06b, BPT94, BGT01, Bra94a, Bra94b, Bra94c, BD01, Che95, DH91, DN94, Dai95, HW92a, Mey95, OW92, PR94, Rog94, SV94, Yam95, Ata05a, Ata05b, AS09, AC17, ABDW21, Bra99, CSS98, CS00b, Dai96, DHV04, DSW07, GH05, GK09, GTZ20, GHK11, HW96, Maj06, Miy04, PW10, PW21, SBF19, SS12, Wal09, YY18]. Queues [Asm92a, BR93, DR92, FF94, GW91, KW91, MP95, ABP15, APS19, ABKR18, BBP22, BPT98, DHT10, DLS01, GOP03, Gur14, HRS97, KR10, KR12a, KR11, KR13, KLSY03, KLS06, KLRS11, KS03b, LSZ13, LW14, MvU05, MP17, Puh15, PR10, Sep97b, TW09, Wis01, ZDZ11, ZBM04]. Queuing [SSX14]. Quick [Bra94c]. Quickest [EP22, JP17, Pes14, DS15]. QuickSort [Fil13].

Radial [CW93, BB07]. radii [IT12]. radius [EN18]. Radner [ESX22]. Random [Ald91, ABBJ94, AGD94, AW94, BFP93, BG93, BW22b, BS19, Ber94, BUV11, BSZ20, BLSW91, Big95, BGV20, BHdS^{+}20, BB92, BNT92, BOW95, CMSS15, ČT16, Cha92, CDS11, Che01, CKW21, Chi16, CF94, CS06, Col02, Col09, Dev92a, EN18, FKK^{+}01, FN93, FN94, FM94, FMP95, Fri91, FHY92, FSW95, GRS04, Gol91, GdH93, Gre94, GJ18, HST91, HW92b, Hsi94, HNNZ13, IM02, Jof93, KX95, KV01, KGK95, LC93, Mah94, Man93, MZ91, McD95, Nut13, PY02, Pet91, PRW95, QS94, RR91, RT92, RWW95, RW97, SD05, Tak92, Tal92, TT13, Too02, VG95, Vit91, Wor95, Zha95, ACD15, ABBL09, ABF13, ABDLV22, AABR22, Adl00, ABL12, AS06, AS10, AL15, AS16a, And98, AK18, AG06, ABH17, Ata08, AGvdHdH18, AS23, BCHL98, BW22a, BST14, BN17]. random [Bal21, BBK^{+}11, BB22, BH22, BCOR16, BDW23, BZ10, BJ20, BES04, BG00, BY05, BHMW16, BS17a, BMRS23, BBRZ20, Bel12, BH20b, BP12, BPZ07, BZ15, BCPR15, BB15, BRTP18, BvdHH10, BBS11, BvdHH17, BBDW22, BS22, BM17, BDW17, BD20, Big12, BS17b, BP18, BU18, BG06a, BG08, BEM07, Blo13, BKS17, BK06, BR04, BK00a, BdHM23, BNS13, BDL16, BDW22, BG03, BKW08, BT12, BGZ22, CJL21, CY21, Can19, CS16a, CF07, CBM^{+}21, CL03, CMY03, CRV06, CF17, CHL21, CRX21, CX16, CDL17, CFS18, CM05, CAP20, Coh04, COMR22, CN11, CK12, Cra16, CGM09, CSS22, CLR06, Dar23, DN05, DDM11, DS18, DS19, DDMT12, DM08b, DFK12, DH06, DJ10, DJ23, DAM10, DdHJN17, DS17, DGM06, Dre00, DL18b, DS07a, EGP16, El 09]. random [EG18, ESTZ13, EFdS21, FLP13, Fed14, FKR17, FH98, FH16, Fla97, FMP00, FZ03, FPZ05, FKP23, FKKM18, FIMS22, FW99a, Fuh04, GM19, GQ03, Gar09, GM12a, GS09, GGR97, GH21, GRK05, Goe06b, Gol16, GS23, GR97, Gol13, GGO23, GS19, GG11, GLP23, GSvdB98, GS05, GK16, HLN07, HLN08, HR97, HH19, Han06b, Har13, HRW08, He20, HR07, HHR96, HK04, HLMS05, HDI16, IKKM15, IT12, JLP08, JJ15, JL08, Jan08, JŁTV12, JvL07, JS96, Jia12, Jia15, Jon06a, JPV99, Jor02, Jos14, JL09, JLM15, Kab12, KMS17, KPT^{+}16, Kar15, Kar10b, KL96, KS14, KK01, KKM06, KM18, KR23, KP04a, Kol17, KP04b, LRM15, LS17a, LPSX23, LS20, LS15, LNS21, LST23, LMT12, LR99, LLLZ18, LR13, LX14, LLC18, LC22, LM15, Lud08, Mån99]. random [MM03, MN03, Mar19, MS11, MPS12, Mat05, MPP17, McD99, Mei09, MRW18, MS00, MW10, MP21, MNP14, MV18, MP20, NY16, Nak11, Nea06, NRY12, Ngu17, Num00, OP00, Oli09, OQR16, PP04, PP08b, PW21, PRR13, Pen97, Pen00, Pen16, PSY15, Pit08, PS20, QZ23, RY13, RR19, Rem08, RT96, RZ18, Sch21, Sha20, She02, Ste99, Sud08, SXY21, TV03, Tei09, TLC93, VAC15, Viv23, WM04, Win08, Woo12, WXY23, Wüt06, Yat09, YZ19, YRF16, YA15, Yos08, Yu17a, Yu17b, Yuk99, Zer98, Žit05, dBG12, vEK08]. random-access [CBM^{+}21]. Random-cluster [BGV20, GG11]. random-field [Dar23]. random-number [Pet91]. random-to-top [Jon06a]. randomization [BCFP18, BET05, EFT07]. Randomized [BK16a, BK17b, BRSS17, DPBCD21, FP15, LP13b, LP17a, Nar16, SS12, AR19b, ABH17, BH05, CDS23, FMS96, Nea06]. Randomly [EKR23, MM07, BLMZ14, BL16b, BM04, FS14, GVR17, HSST23, KLZ98, RS22, RR97b]. Range [BG93, MR93, AAK17, BW22a, BP12, BFS21, BDKT20, BJN18, CF09, Cox10, FMS14, GGLO13, Gra09, HRS99, HS19a, HS21a, MS11, MH21, Muk22, Tra23]. Rank [Ber96, KPSC07, BJ22, BLO23, MSZ19, Mou20, PP08a, PZ08]. rank-based [BJ22, MSZ19]. rank-dependent [PP08a]. Ranks [HK92, PS14]. Rapid [BGV20, MS18, DGJ09, WSH09, Yu07a]. rapidly [Fed14, GJ18, SZ17]. Rare [BBDW22, CDV14, GK95a, BG08, BFW21, BDS20, DG05, DKPRB20, Erh00]. rare-event [BG08]. Rarity [CAP20]. ratchet [PSW12]. Rate [Ale94a, Ale95, AKH02, BBC19, DMY95, Gra09, HS93b, HS93a, HLN16, Ing94, PP12, Pèn05, RS95, AHS05, BT05, BDKT20, BT96, CT04, CTZ04, DF06, DGR09, DKM17, FS99, Fon10, GG13a, GOP03, Hey23, HT12, Kel13, KT04, Lal00, LMT12, LS18, Li18, LW22, LP08, MR02, MP06, O’N97, Pia05, SX23, YEC10]. Rates [Ale93, Ale94b, Ath94, AV95, BG93, Che01, DF95b, JR02, KZ09, KLS95, KS05, MT94a, Mil94, RY94, Rhe95, Ros93, SY23, Ata08, BB20b, BHL96, Bax05, Bel11, BdHNT22, But14, BKS20, CJK19, CS02, DPBCD21, DR10, DFMS04, FS16, FK16, GHH07, Lan15, LMT96, Már97, MW98, MW13, NV03, PP08b, PRR13, Pel98, PW04, RR97a, RR98a, Tra23]. Ratio [AR02, AGSC02, AR20, BS07, FKM96, GHP13, PZ08]. rational [MM10]. ratios [PGZ07]. ray [MM07]. Rayleigh [HS23]. RBM [DH91]. RC [MP01]. RDEs [CP19b]. reaches [JR14]. Reaching [DGR09]. reactant [BR13]. Reaction [Blo92, Kot92, ADE18, BKPR06, BS05, BL12b, BN97, Bro99, CW16, Cer09, KK13, KL02, LW19, LdRS15, MP14, Per00, PP15]. Reaction-Diffusion [Kot92, KL02, LdRS15, Per00]. reaction-hyperbolic [Bro99]. Reactions [BL01]. Reading [RS07]. Real [DLS01, GS18, ABB23, BB20a, EPZ20, EP22, FJ22, KPT^{+}16, KLSY03, LS16a, PTZ17, RS14]. Real-time [DLS01, EPZ20, EP22, KLSY03]. realisability [LRM15]. realizability [KLS11]. Rearrangement [HW94]. rearrangements [BCTV07]. receiver [PZ08]. recently [Jel99]. Recolonization [KKN95]. recombinant [ADS14]. Recombination [JS10, DK99b, EV12, LPS21, LL12]. Reconstructing [DRR21, Mat05]. Reconstruction [DO94, MPP17, Mos01, DR13, DOS19, FLR22a, GLM18, HL20, LM02, MNS16, HL22]. record [Yao97]. recoveries [BL06]. Recovering [EH11]. Recovery [DN91, JK17, MRW18, MNS16]. Rectangle [DH91]. rectangles [AJvM22]. rectangular [BS20]. Recurrence [Dai95, PW21, Sil96, AMS06a, BDH10, CDN02, Che96, Col09, DFK97, DG13, FL13, GTW22, HI09, HJJ16, Pit99, RR06, Wyn99]. recurrent [Bra11a, CK07a, CK07b, CW19, DW05a]. recursions [Bur07, Kel16, Moy15, Roi07]. Recursive [DSS96, MS91, Pan08, Wu09, ABDLV22, AB05, DGM06, EPQ01, LP10, Mer07, NR04, PS22, SV02, Zha16a]. RED [MR06a]. redistribution [CV21]. Reduced [BZ19, SS93, PZ08]. Reduced-form [BZ19]. reducible [Big12]. reduction [HW05, INPY13, JPS^{+}22, KK13, Keb05]. Reed [Nea06]. Reference [Ish93]. References [Cha93]. Refinement [ZZ02]. Reflected [CM08, DH92, GIO^{+}17, QX18, BB23, BQ19, BC15, BdRGL20, DEH23, DM05, DG18, EF21, Han06b, KR14, LST23, MPZ13, MPZ21, Zha16b]. Reflecting [AGP95, BB20b, Bar20, BDH10, Bra11a, Che96, KW07]. reflection [BN15a, BEH18, MPZ21]. reflections [CEK12]. regarding [FKKM18]. Regeneration [WPRS21, CS11b, MMPP07, MMPP08]. Regeneration-enriched [WPRS21]. Regenerative [CFF02, GK95a, Gra21]. regime [AZ14a, ABP15, AP16, AG12, AB14, AG14, AC17, BM19a, BM20, GG13a, GG13b, GS09, HHK21, HSX22, HP20, LP23, MMPV06, Owa22, Ree09, SZ17]. regimes [BF10, BČ05, Ber10, CGM09, Gou09]. Region [Lef91]. Regression [HST91, BPS04, GLW05, HK22, QS23a, BS21]. regression-based [GLW05]. Regret [SS93]. Regular [Gre94, ACD15, BDM02, BS23, BL02, Can19, Che13, CS11b, DTW22, HK23, JMRS09, LS17a, LTVR14, TVVY12]. Regularity [BGR18, LRM15, LS15, DP20, Feh23, Fou00, JSW20, LP08]. Regularization [GH22]. regularly [DHS18, DM20, HLMS05, OCBG11]. regulated [AG09, BEM07, Eth04, FM04, HW07]. regulation [BD07]. regulatory [DS07b]. reinforced [ACG17, BS12, CNV22, CL09a, DMM17, GVR17, Sch21, vdHHKR16]. Reinforcement [LL17, CNV22, HHK21, KMS22, LC03]. reinsurance [LM09, Sch02]. reinsurance/investment [LM09]. Related [Das95, DO94, AS10, ACH97, BC15, ÇD16, CE10, Jia08, KM18, ŁRR13, MN97, NZ15, VZ19]. relation [BB20a, MZ14]. relations [MWZ07, Sil96]. Relationship [HR94, Wal09]. Relative [YY09]. Relaxation [CCM06, CM03, Gra09, Mic02]. relevant [CSZ17]. REM [BF05, FIKP98]. REM-like [BF05]. remaining [BBP22, Puh15]. remark [FW99a]. remarks [AK05a]. removal [O’N97]. removal-dependent [O’N97]. reneging [KR10, KLRS11]. Renewal [Bax05, Gol91, CGZ07, Chi07, Chi15, CL17, Fuh04, MW08, PS20, PY98, Sch97b, ST04]. renewal-theoretic [PY98]. Renewals [DG07]. Renormalization [CP08]. Rényi [ABDW21, BdHM23, Cao21, DKPRB20, Gol13, KKM06, RŞ18]. Repair [BENP91]. repeated [GRS^{+}16]. Repellent [GdH93]. repelling [BC21]. Replacement [CGS93, HS93b, HS93a, GM19]. replica [Gen23]. Replication [Kus95, ADGS98, BDG16, BD19b, BF04b, BT00, JLR03, Lac03]. replicator [HI09, Imh05, Leh23]. Representation [ERY95, MZ02, QS94, Zha05, CC16, FP15, Har00, Har03b, Har06, HLT21, HR04, KP04b, PW10, PY98, SXY21, TBUA23]. representations [BS14, GMŚ19]. Reproducing [NP01]. repulsion [CGZ14]. required [CD18]. Rescaled [CK03, Che22b]. reservation [HL97]. reservoir [GGO23]. resistance [ABBL09, QX18]. resolution [DSS96]. Resolvent [KO92]. resonance [BG02, HI05, HIP06]. Resource [BD15, AK05c, BW01, FRT03, GTZ20, HL97, RV15]. resources [CH19]. Respecting [CH91]. respondent [AR16]. respondent-driven [AR16]. rest [DSZ15]. restless [Ver16]. restrained [BPZ07]. restricted [FMN^{+}16, FV21, Gen23]. Result [CH91, CD06, FLR22a, FJL18, GLM18, KR06, Lac22, Mou01]. Results [Ger11, GM95, Mor92, AN22, AG14, APP08, BH03a, BB03, DDSJ08, DI10, DN97, ELM22, Har13, MN97, MMX21, PS97, PS05, Swa18, YR23, YZ07]. return [Pau02]. returns [BP05b]. reversal [CCM06]. reverse [BS14]. Reversible [DSC93, Din95, FHY92, Yca93, CHS17, Chi16, HKL^{+}19, Kar07, LP04, VPV08]. reverting [FFK12, MAL14]. review [AK05c, Har98, Mag00]. revisited [Ban08, DG18, GINR09, GZ09, HW09, HW10, Lan12, LP13b, LP17a, Yao97]. reward [Chi07, KR96]. Rewards [HK92, CS06]. rewiring [OT23]. Rho [KMPT10]. Riccati [BDKR19]. Ricci [FM16]. Richardson [ADH20, DH07, FP93, Hof05]. ridged [BRTP18]. Riemann [ABH17]. Riesz [CDS09, RS23]. Riesz-type [RS23]. riffle [ADS11]. Right [Kli19, DH04, GIO^{+}17]. right-continuous [GIO^{+}17]. Rigidity [Hol01, Hol98]. Rigorous [Swa18]. Risk [AB22, BNK12, BLY22, FS02, MS93b, NP02, Pha02, APP08, APP15, BF02, CCHH05, ÇJPY04, DP15, DK08b, EK09, GM12b, Gri16, GR12, HNS10, HPŠV04, IH23, Jaś07, Jia08, JKP13, KKM04, KS06b, Miy04, Mon22, NS19, Nag12, Pau02, Sch04, Sch97b]. risk-aware [IH23]. Risk-Sensitive [NP02, AB22, BLY22, FS02, CCHH05, Jaś07]. risk-tolerance [KS06b]. Robert [Dur99a, Dur99b]. Robin [LST23]. Robust [ABDW21, BH13, CK19, CDFO13, DM23b, EI17, FGP21, HS19b, JL09, KR12b, AL17, AC20, BY17, ELS17, KR21, Kni12, LC22, MNS16, Sch04]. Robustness [ZRH15, CvH07]. ROC [CDS09]. rock [HL21]. role [Bor12, DR10, MD18]. Romberg [Keb05]. Root [BB22, DF95a, DF95b, De 11, DFPR22, CW13, GMO15]. root-type [De 11]. Rooted [MS91]. rotated [MN97]. rotation [CDG23]. rotation-two-component [CDG23]. Rough [JR16, Kel16, BCJ22, BCD17, DFM16, ER18, FM05b, HLN21a, LT19, Pel10, VK21]. rough-smooth [BCJ22]. rounding [Lac17]. Routines [Kee94]. Routing [Ana91, CH91, Rhe94, AFRT06, FL96, LM15, OT22]. Rowlinson [JK17]. RSW [Ale96]. Ruin [Col02, ES03, HPŠV04, KKM04, CL17, Col09, DHESC21, Fuh04, GGS03, Gri16, Sch02, ZBD05]. ruinous [GM12b]. Rule [DF95a, vM95, ADS11, GHP13, LZ06, ZRH15]. Rules [GW92, Che13, ELM^{+}16, Fis96]. run [GR12, Imh05]. Runge [CC14]. runs [BC14b]. Russian [AKP04, DH93b, SS93].

Sacks [RWW95]. Saddlepoint [CL03, RZ99]. sale [HH08]. sales [Pul14]. Sample [BBRZ20, DR13, Das96, Gri16, HLR22, IR01, Maj06, MvU05, MR93, Ros93, Wis01, BLW11, BG00, BG02, BdHM23, BG03, CD18, GDVM19, HMY21, HKL^{+}19, HLS19b, JJ15, Jia04, Jia19, KP96, KP98, LS09, LS16a, LR06, Owa18, RSX99, SX23, Tsu19, YM22, Yuk99]. Sample-path [MvU05, BdHM23]. sample-paths [BG02]. Sampler [Ing94, DPBCD21, DHN00, Smi14, DGM20]. Samplers [FFS93, ŁRR13, RR98a, RR98b]. Samples [Sel95, Tal92, GH21, Mar19]. Sampling [ABT92, BFP93, Col02, DKT91, FKP94a, FKP94b, HSV11, JKK02, Ste93, Alb04, ADRS23, Ass97, AR16, BCKP99, BGvdHK15, BKR22, Bla09, CL07b, CD18, CT01, DL10, DW05a, DSW07, DSZ15, DGJ06, EPW06, Fen07, Fil98, Fis96, FK99c, GR21, GW97, GR97, HSV07, HV06, HR04, Hub04, JS10, JS12, JL09, LP10, MP99, MMS20, MW10, MLNW22, Nar16, NRS21, PY18, ST21, Tsu19, WL16]. Sanov [BVLT20]. saplings [DZ15]. sardine [DR08]. SAT [COMR22]. Satisfying [RWW95, Las02, Las04]. Saturation [GL01]. sausage [RSM09]. Scalar [Jou02, BJ22, BGHK20, CM19, DWZZ20, HHMG19]. scale [ARS18, ABDW21, BP12, BB15, BS22, CS21, Feh23, GS19, HHJ17, JLM19, KT04, LRPY22, Mag00, MP23, MP06, MSW97, vdHvLS21]. scale-free [BB15, BS22, HHJ17, JLM19, vdHvLS21]. scaled [TT14]. scales [KK13, vdBN17]. Scaling [CF11, DFG22, FSW15, Gau98, GL23, HW92b, LSZ13, MN17, MN22, PP15, BM19a, BBP22, BK16b, BR17a, BKR22, BDW17, BPS04, CP19a, DPBCD21, DM08b, FGG14, GGR97, GS23, GS17, JR16, JLM15, LR12, NR06, NRY12, PST12, Puh15, Sep97b, SWZ14, VK21, Yuk15]. scalings [BRS09, BF22b, DW98]. Scan [DK92, GN91, CK00, DGJ06, FS16]. Scenery [LM02, Mat05, MPP17]. schedules [CC98a]. Scheduling [AMR04, Ata05b, AG14, HPTvD95, OW92, RS01, vM95, Ata05a, BW01, GK09, Mag00, SS12, Sto04]. scheme [AJKH14, DKM17, GZZ15, HLN21b, KRSS23, LM06, LT19, Mek15, Moy15, Yan05, Zha04]. schemes [AA13, Atc10, BT22a, BdRGL20, CC14, CS13a, DL16, HLN16, KNRS22, LdRS18, Mal03, OTV12, See20]. Schilder [BVLT20]. Scholes [Çet18, DK08a, ET09, Loe18]. Schrödinger [BVLT20, Mor05, MR08]. Schwartz [Ger11]. scissors [HL21]. Score [AW94, Cha05, LMT12]. scoring [Ber96]. screened [KM08]. SDE [APS19, CC16, CM96, FP15]. SDEs [LS19, BCFP18, BQ19, Bec06, BCC^{+}23, Clé23, CGK^{+}23, CM14, CM19, De 11, Der11, DL16, FG20, FP15, GH22, GS14, HHMG19, HJK12, LS17b, LT19, MWZZ15, MG04, MGY23, STT19, TKH09, WZ20, Yan05]. Search [CDJH01, DF95a, DF95b, FH95, Grü14, Pem09, BF10, CMY03, DG07, DJN08, GS05, Grü09, Jel99, LM96, LZ98, MN03, PP04, TD17]. search-cost [BF10]. seas [ARL08]. seasons [CDL09]. Second [CM14, Gol16, HLS16, KTPZ15, LNS21, MPZ13, MPZ21, Mos01, PZ19, AD20, BT13, Cho09, MG05, PT15, RS03a, STZ13]. Second-order [Gol16, HLS16, KTPZ15, MPZ21, PZ19, PT15, RS03a]. Secondary [KW91]. securities [Bät99, DSS96]. Security [DG95]. See [Sel95]. seed [BCKWB16]. seed-bank [BCKWB16]. seeds [CQ97]. seen [KST04]. Segel [FJ17, Tom21]. segments [MRS01]. Select [KPSC07]. Selection [CGS93, Fra02, GK95b, BCH22, BEG00, BG00, BB03, BFH20, CS18, CHS22, DGP12, DK99b, EFPS17, Fis96, GK96, GS02, HSX22, JS96, JKK03, San08, SS08]. Selection-Replacement [CGS93]. selective [BG05a, DM11, SD05]. Self [BL12a, BK16b, Cha93, GdH93, KV01, RR00, AS04, BT08, BS17b, BP18, BF96, CS06, Haa10, HIP08, HX22, KS96, NP01, Sad98, Žit09]. Self- [GdH93]. Self-Attractive [KV01]. self-blocking [BS17b]. self-exciting [HX22, KS96]. self-generation [Žit09]. self-interacting [BP18]. self-normalized [BT08]. Self-Organizing [Cha93, BF96, Sad98]. Self-similar [BL12a, BK16b, RR00, AS04, CS06, Haa10, NP01]. self-stabilizing [HIP08]. sell [DH04]. Selling [XS92a, XS92b, dTP09]. Semi [AL17, GJKS15, HPTvD95, NT20, APT20, BZ16, Sil96]. Semi-discrete [GJKS15]. Semi-implicit [NT20]. semi-linear [GJKS15]. Semi-Markov [HPTvD95, APT20, Sil96]. Semi-static [AL17, BZ16]. semidefinite [CFMT11]. semigroup [BGR18, HKV20]. semigroups [AD20]. semilinear [Cho06, Wan23]. semimartingale [Che96, KW07, Kar13, Yu15]. Semimartingales [BJJN21, ASJ20, CS17b, EPS09, Gua02, JPS09, JMSS12, Nut12, PSZ14, TT14, Yan05]. sense [NP01]. sensing [DM23b]. Sensitive [NP02, AB22, BFJ06, BLY22, CCHH05, FS02, Jaś07, KR96]. sensitive-discount [KR96]. Sensitivity [BM20, DG14, DKPRB20, KS06b, BBM07]. separable [BF04a]. separated [AK05b]. Separation [DSC06, KK13]. sequence [AL05, BG00, BM17, CDS11, Jos14, LMT12, MR17, Pit08]. Sequences [AW94, Col02, Ros95, Zha95, AS16b, BH01, BH03a, BH03b, BHK05, BKW08, CK00, DS07b, LRdH98, Mer07, Sep97a, WM04]. Sequential [ADRS23, BG00, DGMO11, GW92, KPSC10, Ngu93, PY02, BCJ14, BJKT16, Bha16, CL11, DPS08, DKP22, DMO14, DG21, Jár17, Sud08, ZHC06]. Series [BR93, GW91, MP95, BS96, DHS18, Jaf00, Niu97, Sep97b]. Serve [AC19]. Server [Asm92a, DR92, KW91, KOOS91, AR19a, APS19, Ata08, BFW21, CW19, DHT10, KR10, KR12a, KR11, KR13, Kum00, LW14, PR10, Sha15, SY13, TW09]. servers [AMR04, Ata05a, Ata05b, AMS06b, AG14, BBC19, BW01, Har98]. Service [Bra94c, HPTvD95, KS92, Ata08, Fla97, FL96, KLSY04]. set [DeB04, DRS16, Fér20, IKKM15, IM02, IM10, MP09, Tei09]. set-indexed [IM02]. Sets [GK95a, Adl00, AK18, APT20, BR15, BL11, CY21, FKR17, Gol16, HV06, Kar13, KPSC07, LRM15, MV16, RR97b]. several [AGSC02, CH19]. Shabat [BB20a]. Shadow [CS17b, HP15, KMK10a]. Shape [DO94, Gou07, AMP02, BZ18, BP18, CP19a, DI17, GM12a]. Shaped [GM95, GO00]. shapes [BBO15, DH13a]. Share [BK92, DR98]. Sharing [MR94a, AH98a, Bra10, FT17, FW22, GPW02, Gro04, GK07a, GW09, KKLW09, KW04, LSZ13, PW04, RR03, RR08a, RV15, STZ14, ZDZ11]. Sharp [AK18, BB03, Che13, GG11, HK92, MGY23, Cer15, DR10, GRK05, JK17]. Sharpness [vdB11]. shattering [Haa10]. shear [BE23, Yos12]. shear-induced [BE23]. shelf [DFH13]. Shephard [RS06]. Sherrington [GK21]. shift [CGR09]. Shiu [APP15]. Shock [BF22b, MSW97, MG05]. Shocks [FR92, Tou00]. Shooting [BOW95]. Short [Tsi94, XS92a, XS92b, Pul14]. Short- [XS92a]. Short-Selling [XS92b]. Shortest [HDI16, AC19, BM19a, BBP22, BGvdHK15, Bra11b, BLP13, BFW21, FM01, Puh15, BM20]. Shortfall [Pha02, DK08b]. Shot [DO91, BD12, LR19]. shot-noise [LR19]. shrinkage [BJJN21, Lar14]. Shuffle [BD92, BCMR21, Jon06b, MNP14, NN19]. shuffles [CCM06, Ciu98, Goe06a, Lal00]. shuffling [APW08, ADS11, BCTV07, DFH13, Jon06a, Wil04]. sickle [CS99a]. sickle-cell [CS99a]. side [HNS10]. sided [BCMR21, BJR16, FSW15, MAL14]. sieve [GINR09]. Signal [WH93, BS07, Bud02, PGZ07, PZ08]. signal-noise [Bud02]. signal-to-interference [BS07, PGZ07, PZ08]. signals [GRS^{+}16, vH09]. signatures [BHRS23, BLO23]. silhouette [Grü09]. similar [AS04, BL12a, BK16b, CS06, EN18, Haa10, NP01, RR00]. Simple [Ben12, CW93, BL21, CM03, CMY03, Fou22, GNS23, GR97, HMT22, KR06, LL20a, MD18, Mat05, Mor06, NRS21, PS20, WH17]. Simplest [LST23]. simplex [CLSF18, Leh23, Smi14]. simplicial [FIMS22]. simplified [BB03]. Simulated [Fra02, CC98a, CC98b, Pel98, RR14, WSH09]. simulating [LM06, Mey06]. Simulation [AGP95, BS14, BL14, DL10, DG95, FHY92, MT99, AGK11, Bel11, BR05, BG08, BC15, BCD17, BDS20, CR16, CDV14, CFF02, CK07a, CK07b, CM14, GS14, HLTT17, JS17, JKW11, KPS98, KM08, KKPvS11, LS15, MH21, NP05, PP18, Ric11, San10, Wik01]. simulation-based [Bel11, San10]. Simulations [GW92, LX14]. simultaneous [BF18, Wik01]. Sinai [LPSX23, Pèn05]. Sinai-type [LPSX23]. sine [Fus00, Tak96]. Single [Asm92a, DR92, DTW22, ET11, FRT03, HS07, HKL^{+}19, HL97, QH21, Ath94]. single-factor [ET11]. Single-Server [Asm92a, DR92]. single-site [HS07]. single-step [QH21]. single/multiple [DTW22]. Singular [CDET13, DZ94, Wee98, AJvM22, BK17a, BB20c, BGBP23, BR06, CDGL22, CGZ14, DY18, GL23, HHSZ15, LS14, Muk22, NS19, Ona08, RS23, VPV08, Zha16c]. singular/switching [VPV08]. Singularly [YZ07, ABW07, BYZ00, ZY96]. SIR [Bal21, DDMT12]. SIS [LL20b]. site [CS00a, DK19, HS07, SSW06]. sites [MZ05, Sch21, ZCD05]. Situation [Yam95]. Size [LC93, AS23, BH99, CD18, CR18, DHL18, DM08b, DKP22, ELM^{+}16, EMO10, FW22, GW09, HMGR00, JKM15, PP04, PY18, SJ05, SWZ14, Tsu19]. size-biased [PY18]. size-dependent [DKP22]. sizes [GM19, Jos14, MZ05]. skew [AJvM22, ABT^{+}11a, ABT^{+}11b]. skew-Aztec [AJvM22]. Skip [Dev92a, FMMP08, LSZ97]. skip-free [FMMP08]. Skorokhod [ABKR18, CHO08, DGPR19, HK13, KTT17, LR18]. slab [DFK12]. slice [NRS21]. slot [EL10]. Slow [GH08, LS22, EFdS21, FN17, Ger11, GKS04, GK07b]. Slow-fast [LS22]. slowdown [AG14]. Smale [CDP18]. Small [Cha92, CST05, DG08, FFK12, GK07b, BJ20, BBL14, CHMK20, KK01, LLLZ18, San08]. Small-time [FFK12, BBL14]. Small-world [GK07b, San08]. smallest [Bor16]. Smoluchoski [EW01]. Smoluchowski [Arm10, Ber02, ILP15, Nor99]. Smooth [DM02, FK99b, BBM07, BCJ22, BD12, BR04, CM19, De 11, KW07, TK02]. smoothing [DGMO11]. Smoothness [De 11]. snakes [Mar08]. snapping [EFdS21, Lej16]. Sobolev [DSC96, FK99c, GLWZ22, HS23, JSTV04, SS19, Zha16c]. Social [FEvdD16, BCDS15, HMQS20]. societies [BD15]. soft [GK07a, Gué03, Pen16, Xu18]. Sojourn [KS93b]. sojourns [Erh00]. solid [MS12]. solid-on-solid [MS12]. soliton [DG08]. solitons [Fed14]. solute [Bha99]. solution [ASCDH09, AKH02, Bal19, BT00, BS13, CL22, CvH10, DP15, DR10, GR08, GK03, HH08, JCLP21, MM01, Mor05]. Solutions [CC93, Gol91, Gre94, HW92a, HL01, HW94, Pit92, SV94, BL12a, BS05, Bec06, Ber02, BNT19, CCCS11, Cho06, DFFN21, DGR09, EGK22, FR11, GMŚ19, Haa10, Hey23, HNS11, IKKM15, KK01, Kli19, Ren16, See20, TV03, TY11, WZ20, Zha05]. Solvability [JMP21, DIRT15]. solve [BCDS15, GLW05]. Solving [Bel13, Gué03, DD10, Gu22, Moy15]. Some [Aka95, Ale93, AK05a, BD01, FFS93, KL91, KS96, Lac03, LR06, MWZ07, OWZ97, PY01, AM06, ATV15, AV16, APT20, Atc10, BGG^{+}16, BR08, CP18, CPT12, DH06, DN97, ES02, FH98, FRST94, Gou07, Gou09, Gué03, KMZ17, KY10, KZ09, KM13, KLZ98, MR05, NP95, PDG14, Sta97, Whi13, Yuk96]. sophisticated [HP23]. sortie [Mic96]. Space [KX95, Mas95, NP01, Sta15, ADRS23, BVLT20, BF08, BM19b, CS16a, CST22, CM19, DGMO11, HV97, Jaś07, KKLW09, KL04, KLS06, LRPY22, LL12, LØP04, MWZ07, MT99, Rhe00, SW12, Sto04, Zha16b, dRST19]. Space-time [Sta15, LØP04, MT99, Zha16b]. Space-Valued [KX95]. Spaces [Ros93, BCP18, FG18, FRLS21, GR06, GS11, MV20, Muk22, TSM19, Tie98]. spacings [BPY09]. Spanning [AB92, Jai93, PP04, Ale96, BB07, CS17a, KL96, Lee97, Pen96, Pen97]. Sparre [BMX17]. sparse [AS16a, Blo13, Cao21, HYTC20, He20, HP20, HLS19b, LRW23, MRW18, Owa22, VAC15, Woo12]. sparsification [BDL16]. Spatial [BDKT20, BL01, DM15, JM15, JM91, MV18, ST10, Ste93, BD07, Che18, CCL16, CS99a, DH13a, DR09, Dur09, EV12, EFPS17, GSvdB98, Hof05, IM10, KR23, LN06b, LS16b, NP99, Pen00, PP15, Puh15, Sch97a, Sch05b, SSW06]. Spatially [LN09, BL06, CFS18, LN06a, MSSZ20]. SPDE [AR19a, HL00, KR13, LPSX23, SZZ23]. SPDEs [AR21, BQ19, BKS20, CJK19, GJKS15, HSV07, JKW11, Wan23]. Special [Ish93]. specialists [LN06a]. speciation [Yu07b]. Species [HLR22, AS10, BFS21, Bor12, CW16, DFPR22, MR17, Per00, PY18]. species-tree [DFPR22]. spectra [El 09, KKM06]. Spectral [BF05, DJ10, Eic95, FHY92, HSV14, HJ23, KM18, KM03, KST04, AN22, BH22, CJK19, DHS18, FIKP98, HS19a, JWB^{+}14, Klo19, May09, NY16, NRS21, DM94]. spectral-gap [HS19a]. spectrally [AKP04, APP07, FPRW18, Loe08, RZ18]. spectrum [AABR22, DK19, GIMM18]. Speed [BRSSJ19, GdH93, Kes93, CS11b, EMO10]. speeds [Big12, Kli19]. sphere [LS15, MV16, Pap00, PS17a]. spheres [DH13b, HPP22, Hey19]. spherically [JJ15]. spike [LS18]. spikes [CHL21]. Spiking [BCC^{+}23]. spin [ABA20, CP18, DLZ21, DGJ09, JKS18, JK17, MT13, MH21]. split [BH12, KL04]. splits [KS16]. Splitting [BDS20, BGG^{+}16, CG16]. spot [Tod19]. Spread [Big95, FGLS23, Pen93, BS15, BDKT20, DGM08]. Spread-Out [Pen93]. Spreading [Big12, CF94]. Square [BS20, TY93, CS21, De 11, GPR17, GS20, HSST23]. Squared [CW93]. Squares [Ser94]. Stability [BVP22, BM04, Bra11b, Che95, CL04, CmHP04, CGL^{+}15, DHV04, Fed14, FM05b, FL96, Kel96, LO04, Mas07, MD94, RS03a, RR94, SS18, Whi13, ATV15, BGR18, BLMZ14, BCJ14, Bra10, DT18, DFM16, DMO14, FM01, FMMP08, GK09, TvH12, vH09]. stabilization [LRPY22]. stabilized [Pal11]. stabilizing [HIP08, LRSY19]. Stable [ASG93, DR93, HST91, MRRS02, Pit92, RR91, AS97, Bra99, BM96, CKW21, CS06, Die15, Duf16, FPRW18, HRW08, JMRS09, RSX99, Xu19]. stable-like [CKW21]. stage [DSS09, Kro99]. Star [GM95, Dd04, GO00]. Star-Shaped [GM95, GO00]. starting [MG05]. State [DH91, DH92, HS93b, HS93a, KKLW09, KLS95, MP98, Mas95, MT94b, Sto04, Yam95, Ass97, BGR18, BHQ17, BC15, BD17, BR06, BF18, CvH07, DGMO11, FL22, FL96, GZ06, GG13b, Gur14, HV97, ILP15, Jaś07, KR96, KL04, KLS06, LYZ19, SW12, Tie98, ZY96]. State-Dependent [HS93b, HS93a, KLS95, MT94b, Yam95, MP98, BGR18, FL96]. states [BYZ00, CK12, FMN^{+}16, Li18, PW04]. static [AL17, BZ16]. Station [SV94, DHV04]. Stationarity [Asm92b, Fil91, SZ06, BB20b, CSS98, GG13a, KK04, MNG09]. Stationary [ASG93, Ana93, Ana94, AP17, BK01, DR92, DR93, Eic95, EK92, HR94, Kab10, Per94, Ros95, Yu07a, Yu07b, AS97, Alb04, AS16b, Asm98, Ass98, AC03, BBL^{+}97, BCP18, BPT22, BN19, CBM18, CX16, CSS22, Das96, DK99a, DFK97, FS14, FMN^{+}16, FKM96, FRSW22, HV06, Hei05, HKK06, KR12a, KR14, KLZ98, KP04a, KS03b, LS00, LRdH98, MRV21, NS22, PP12, Pèn05, RS03a, RR19, WKRS19, YA15, Yun98]. Statistic [GN91, BH22, GPR17, PZ11]. Statistical [CC93, HW94, JLM20, Keb05, KSS21, Egl05, TV03, TSM19]. Statistics [Tak92, AH98b, ABK12, AZ14b, BFJ06, CK00, CJY15, CN11, FS16, HLN08, HK17, LRS17, LT18, LC22, NY16, PSZ14, RR97a, SY23]. Staver [DZ15]. Stay [Lef91, BDW23]. Steady [BC15, DH91, DH92, GG13b, GW93, BD17, GZ06, Gur14, Li18]. Steady-State [DH91, DH92, BC15, GG13b, BD17, GZ06, Gur14]. Stefan [CRSF23, KRSS23, Mül18]. Stefan-type [Mül18]. Stein [BD17, BN19, CS11a, CS17a, CRX21, CCL16, ER08, FK21, GR21, GPR17, GR97, Loh92, Roo94, Xia97, Xu19]. Steiner [GNP17]. Steining [GNP17]. Step [HMGR00, BBC17a, BH99, GM19, QH21, Tri15]. Stepping [KKN95, CD02, CD03, Cox10, DR08, KRSS23, ZCD05]. Stepping-Stone [KKN95]. steps [BBW04, Jon06b, MS00, PS17a]. Sticky [HSH^{+}13, LY22, RS15]. Stigler [Swa18]. Stigum [GLP23]. stimuli [Sad98]. stirring [Yu07a]. Stochastic [Ant93, BL16a, BH99, BF12, BCP18, BG02, BENP91, BMN14, BZ91, BMR08, BKT22, CC99, CDP18, CMT21, Cho02, DZ94, DFK97, EW01, EW03, Eva01, Fin94, Fon10, FJ17, FJR20, FHY92, GM19, GW92, GK95b, GMŚ19, Gup12, HR94, HHK20, HL01, HKV20, HH18, Ish93, Ish95, KX95, KO92, KL99, KL01, KS93b, Kot95, LW92, LN06b, Lef91, LL13, LØP04, MPST02, MV20, NY23, PQ01, Pin92, QZ23, Rei95, Ste95, TY11, Yos12, Zha12, Zha16c, AS01, ARS18, ADS14, ACG17, AC20, ATV15, Asm98, AdHR21, Aus08, BBC17a, BF04a, BCFP18, BK17a, BB20c, Ban19, BFM11, BM22, Bäu00, BS14, Ber97a, BFG13, BvdH12, BNT19, BC15, BCD17, BET05, BCS19, BD98, BFH20, BdRGL20, BM19b, BS13, BLM17, BLM23, BG06b, BG12, Bur07]. stochastic [CDN02, CW16, CDET13, CLP16, Cer09, CP19b, CL11, CC14, Che18, Che22b, CDG23, CST05, CC16, Cho06, Cho09, CKHL06, CE10, Col09, CDV14, CFJ16, CI11, CS99a, CDMR12, DL08, Das96, DP05, DH18, DM06a, DK08a, DI10, DM02, DZ16, DEH23, DWZZ20, DMP96, DLS03, DL22, Dur09, EGK22, ELS17, EG18, EP10, ES03, FG18, Feh23, FFK12, Fer15, FJL18, FK16, FP22, Fri16, FGP21, Fuk11, GHLT14, GH05, GO12, GLW05, GR09, GT19, GLM18, HL17, HN18, HNS22, HLTT17, HHSZ15, HP23, HI05, HIP06, HI09, HMGR00, HQR96, HX22, HNS11, HLN21b, HSX22, HJK13, IT99, Imh05, JMP21, JCLP21, Jac02, JMSS12, KK13, Kar15, Kel96, KY10, KRSW19, KM13, Kne00, KT04, KL04, KL02, LS15, LP13b, LP17a, LS00, Laz04, Leh23]. stochastic [LS18, LZL21, LR13, LLS08, MZ02, Mag00, MP98, MRS01, MN97, MPZ13, MPZ21, MSSZ20, Mer07, MM10, MM01, MP06, Mor05, MR08, MN17, Moy15, Muk22, MG02, Mül18, MW13, Nor99, Pan08, Pau02, Pel98, PS16, PTZ20, QX18, Rhe00, RV15, RYZ21, RZ23, Sch97a, See20, ST10, SFR16, SXY21, TD17, TV03, Tan14, Tom21, Tri15, Vys08, Wag05, Yuk15, Zha16a, Zha16b, Žit05, DM15]. stochastic-Lagrangian [CI11]. Stochastically [MP14, BE23, CMY03, LMT96]. Stochastically-induced [MP14]. stock [DH04, EH11, LV03, dTP09]. stocks [Cha99]. Stokes [BFH20, CI11, FHLN22, Fon10, HLN21a, Kot95, Mél00]. Stone [KKN95, CD02, CD03, Cox10, DR08, ZCD05]. stopper [HHSZ15]. Stopping [DZ94, GW92, GK95b, HK92, AJO14, BHRS23, Bel11, Bel13, BK16a, BK17b, BR08, Çet18, CPT12, CHO08, DFM18, DP20, DW05b, Egl05, EV06, EJ16, GIO^{+}17, HH08, Käl22, KS99a, KQRM11, NZ15, NZ20, Ott13, Pes19, RZ18, RU08, ST04, XZ13, ZRH15, dSDG10]. stopping/optimal [HH08]. storage [AS04, Kne00, Löw98, MNG09, PP96]. strands [MM07]. strange [MRS01]. strategic [BBC19]. strategies [BZ16, HRW08, LM96, LZ98, RT14, Wee98]. strategy [AM10, BEZ20, GL14, Loe08, Per03]. Stratigraphy [LC93]. stream [HRS97]. streams [DR98]. strengths [BC22a]. stretched [HSST23, Lac17]. Strict [KKN15, Mar02, Lar14]. string [DR02, SY98]. Strong [BH01, BH03b, BCD17, CS00b, Chi15, Cla96, Fin94, HJK12, LP23, LN05, Mah94, WJB^{+}15, Yos12, BJ22, BM19b, Cha05, FR05, Grü14, HHMG19, HS23, HHK21, LS17b, LS19, LR06, MGY23]. strong-Rayleigh [HS23]. Strongly [HW92b, vdHHKR16, CL09a, LRdH98, MS21, Tri15]. Structural [Mas07]. Structure [BL01, CMSS15, CH91, DN05, FHY92, Hog93, Mil94, ABT00, CS13a, ET11, FZ02, HJR20, JL22, MSW97, Oha09, TLC93]. Structured [BK95, KOOS91, BL06, MSSZ20, SJ05]. Structures [Cha93, Dev92b, HHJ17, CHS22, FH98, Gol04, LPW08, LPW14, NR04, vdHvLS21]. Study [Dev92b, MR00, BEM08, DSS09, Fou00, MD18, NST20]. stymies [DPRZ19]. sub [DS17]. sub-ballistic [DS17]. Subadditive [Ale94b, BMS02, Rhe93]. subadditivity [Jia08]. Subcritical [Gou09, Jan08, Lac22, Pit08].

Subexponential [Asm98, ZBD05]. Subgeometric [But14, FR05, DFMS04]. subgraph [AS16a, vdHvLS21]. subject [CF16a]. Submodular [DFFN21]. suboptimal [AS09]. subpower [Pit08]. subpower-law [Pit08]. subsampled [MRW18]. Subsequence [Ale94a, Ale95, Fri91, HT12, Jin19]. Subsequences [Rhe95, BC22b]. Subsets [RT92, ADH15]. substitution [BR01]. substrate [LY22]. subtract [TLC93]. subtract-with-borrow [TLC93]. Subtrees [MS91]. successes [CX02]. Sufficient [LS95, Che96, Cou08, DY18, KS03a, KLS11]. sum [BK00a, DFM18, Har13, HHSZ15, PTZ20, ST04]. summands [RR97a]. Summation [Cha05]. Sums [CS95, Che22b, EK09, JS07, Kol17, LRdH98]. Super [BDG16, BD19b, BF04b, BT00, CK03, FK99b, FK00b, JLR03]. super-Brownian [FK99b, FK00b]. Super-replication [BDG16, BD19b, BT00, JLR03]. supercooled [CRSF23, KRSS23]. Supercritical [BB15, AR96, Coh96, GLP23, HHJ17, NV03, NV04, NV06]. Superextremal [RR94]. superhedging [BFM17]. superior [Pet91]. superlinear [GR15]. superlinearly [Sab16]. Supermarket [BMW19, BBD22, LN05, LN13]. supermodular [BF12]. superposition [PR98a]. Superprocesses [AS01, Sch13]. superreplication [BD19a, Hob98]. Superstar [BSZ15]. supOU [BNS11]. supplies [Pet91]. support [Che08]. supported [LR21]. suprema [Har13]. Supremum [Blo92, DLS03, Iva18, Mek15, MS00]. sure [Ber97a, GL14, GK16, HHK20, Jor02, LV10, Nak11, vdHMS08]. Surface [Yuk15, BN97, LS00]. surfaces [CAP20, Dar23, Pes14]. survey [AB05]. Survival [Aal92, BD07, Eth04, Lig95, Pov95, BL16b, BEM07, Bor12, IM02, PS20]. survivor [JS07]. SVD [BS21]. swap [BCTV07, BGR22]. swaps [BJR08, HK13]. sweep [SD05]. sweeps [DM11]. Swendsen [BCP^{+}22]. switch [Sto04]. Switched [Ana91, SW12, BLMZ14, SWZ14]. Switches [IMQ93]. Switching [DZ01, BL16b, BS19, BdHNT22, HSX22, VPV08, dSY05]. Switchover [CPR95, HDP22]. symbiont [BLZ11]. symbol [Fil13]. Symmetric [ASG93, AS97, CPSJ22, Fis17, Gen23, GS02, HM09, HMT22, JJ15]. symmetries [JJQ16]. symmetrized [JWB^{+}14, WJB^{+}15]. symmetry [CL09b]. sympatric [Yu07b]. symplectic [CS16a]. Synchronization [ACG17, ABA20, BG02]. System [FF94, Fin94, KLST93, KS92, Yca93, AHS05, BB20a, BW01, BD07, BFH20, Bro99, CDG23, DKM17, Eth96, Har98, HRS99, Hey23, HLN21a, JKM15, JM08, KK01, KP16, LY17, MMPP07, SS15]. Systematic [DGJ06, Ste93]. systemic [NS19]. Systems [CPR95, DH93a, KZ94, SV94, Ser94, SY13, ABB23, AC03, Ata05a, Ata05b, AMS06b, BS96, BS07, BCFP18, BJ20, BS05, BFO18, BLMZ14, Ben96, BZ18, BD17, BE23, BdRGL20, BMR08, BF18, CSZ17, CS22, CGR20, CLP16, ÇD16, CF11, CF16b, Del98, DG99, DGJ09, EW03, ESX22, FM16, FS14, FL96, FRT03, GGLO13, GGO23, GRS08, GLWZ22, GL23, HN18, HNS22, HMGR00, HDP22, HMS04, HL97, IT99, JK14, Kab10, KL04, LS22, MM03, MMPV06, MMPP08, MP14, MH21, MG02, NS19, NT20, PP08a, PS14, Rey18, Shk11, YS22, Yu07a]. Sznitman [DK19].

Table [Ano16b, Ano23c, Ano23d]. tables [Bla09, Eva01]. tacnode [AJvM22]. tagged [MP22a]. Tail [ACLW95, AG06, BM19a, Ber92, BK01, BMS02, DHS18, Gre94, NP95, RS98, dGvZ93, dSY05, CL11, CDS09, CN23, Dre00, EHW16, KP04a, LRdH98, WH17]. tail-dependence [EHW16]. tailed [BBP22, BG08, DR96, FZ03, FPZ05, JR16, JM03, MS00, Viv23, ZBM04, vdHMS08]. Tails [Gol91, BF04a, BZ10, BLP13, HRS99, HW09, HW10, KKM06, Kol17, Lac17, OCBG11, RR00, Sto03]. Take [Sel95]. Takens [BBD99]. takes [DvdH16]. tale [Atc10]. Tame [LS95]. tamed [KNRS22]. Tandem [Kel93, GHR03, KST04, MvU05]. Tangent [RS95]. TAP [QS23a]. Tar [CT91]. Target [BOW95, BL16a, Béd07, BMN14, NRY12, Pes14, STZ13, VK21]. targeting [AFKP20]. targets [BRS09, Pel98]. TASEP [BF22b, CLW16, MG05]. Tasks [Ngu93]. tau [AGK11]. tau-leap [AGK11]. Taylor [BS96]. TCP [MR06a]. team [FHH23]. technique [KKPvS11]. Techniques [Col02]. technology [VPV08]. temperature [BCOR16, Blo15, CDN02, DL22, FJ19]. tempering [RR14, WSH09]. template [Chi05]. tensor [ELM22, JLM20]. tensors [CHL21]. Term [Hog93, Mil94, CLP16, DMO14, ET11, FZ02, FS99, HSZ17, KK01, MW13, Oha09]. terms [PS22]. territories [Gou07]. Terry [CDL17, HYTC20]. tessellation [CR18, Gol10, MSW97]. tessellations [GG97, HSS06]. test [Cha05, LLS08]. Testing [WXY23, WH17]. Their [Eic95, AA13, BHZ02, Ber02, Cra16, JPV99, JM08, KKP12, Mal03, PS14, Tak92]. Theorem [DKT91, FM94, Kus95, YH93, Zha95, Ale96, AMP02, AS16b, Ata08, BHQ17, BFO18, BK15b, BKN22, BP18, BL21, BM97, BC22a, CP19a, CCK23, CP14, DGL23, DM08a, DG99, DT05, DGR09, DR10, DF20, DJN08, FRT03, GR08, GVR17, GM13, GLP23, GK16, GT19, GZ08, KL96, LS09, Lee97, Mer07, Muk22, Nak11, NR04, PZ08, PZ11, PD22, Pèn05, Per03, PS20, Pul14, Sch97b, Sch10, Sei09, Yao97, Yos08, ZY96, dBG12, Tsi94]. Theorems [AB93, DSC93, KL91, LR00, PY01, RY94, And98, AG09, BVLT20, BHZ02, Bal21, BDMT11, BR19, BCPR15, BDM17, Can19, Cao21, Chi15, CES23, CDS09, Coh04, CSS22, DFK12, DL16, FK21, FK22, FG18, Fér20, FKM96, Grü14, Guy07, HSS06, HST18, HS99, HSH^{+}13, HLS16, Hwa96, Hwa98, ILP15, JT18, JR15, JKO09, KKP14, KM03, LR06, LZL21, MZ02, Owa18, PP22, PDG14, PSZ14, SSV18, SZZ21, ST04, TT11, TT14, Yuk96, ZHC06, Zha16a]. theoretic [HLN08, MR17, PY98]. Theoretical [WKRS19, BEM08, CAP20, DH23, DM06b]. Theory [Dev92a, FP95b, Gol91, Mil94, PY02, Rog94, ZZ02, ADE18, BT05, Bax05, Ber97a, CGZ07, CBM18, CRX21, CQ97, CL17, CDFM20, CD99, CS16b, DR96, DP05, EGP16, ET05, Fuh04, HYTC20, HP23, JCLP21, KM03, KS03b, KP03, LSZ13, LS20, MP21, NV04, NV06, Niu97, PY13, Pet91, Sch97b, SSV18, SZZ23]. There [SSC95]. Thermalisation [BJ20]. thermodynamic [Dar23, MMN22]. thickening [Yos12]. thin [HW09, HW10]. Three [GHP13, LL17, SV94, BDH10, Fon10, GL18, GHS21, HNS22, MM07]. Three-dimensional [GHP13, Fon10, HNS22]. Three-Station [SV94]. threshold [BW01, BH20b, Cer15, CmHP04, DFM18, GG97]. Thresholds [BHL96, CZ18, DGM08, AK18, ELM^{+}16, GRK05, GG11, JLM20, Sly08]. throughout [SZ17]. throughput [AS09]. thrown [HR97]. thumb [ADS11]. tick [ASJ20]. Tie [ESS93]. Tight [GN91, RT92]. tiling [Wil04]. tilings [FV21, SSV18]. Time [BZ08, Ber92, Ber94, DR92, DM94, DR91, EK94, HPTvD95, HI09, Lig95, LdRS15, MAL14, NP02, PR94, Rhe95, SCZ10, Tak95, vdBK93, AS10, APT20, Ass97, BBP22, BBC23, BK15a, BD19a, BDMT11, BS12, BGT22, BBL14, BBS11, Bla96, BC18, BdHNT22, BP05b, BN15b, BCL06, Bud02, BGR22, BF22b, BDS20, BT12, CCM06, CM03, CF07, CX97, CEK12, CST05, CvH07, Cop22, CL09a, CHO08, CS13b, DH04, DHL23, DP15, DH13b, DI10, DHS18, DMR04, DLS01, DW05b, DR08, DSS09, EH11, ELS17, EPZ20, EP22, EHW20, ER10, FFK12, FZ02, Fla97, FK16, FZ03, FPZ05, GOP03, GHL03, GR06, GS11, HJR20, HP23, HI05, HIP06, HKL^{+}19, HL00, JWW11, JR14, JT03, JJQ16, Jia15, KK13, KMZ17, KM98, KRSS23, KLP15, KQRM11, KT04]. time [KLSY03, LL20a, LZ06, LvZ04, LW14, LL13, LØP04, LL20b, LM05, Mar02, MS12, MW98, Mek15, Mic02, MT99, MP06, Mor06, MNP14, Niu97, NZ20, Pia06, PS17b, Puh15, PW97, RS05, RZ18, RS23, SS18, Sta15, Tan14, TVVY12, Tra23, Web01a, Web01b, Wu09, Zha16b, ZS09]. time-average [Wu09]. Time-changed [MAL14]. time-dependent [AS10, CX97, HL00]. time-homogeneous [EH11]. time-horizon [RZ18]. time-inconsistent [HP23, KMZ17, NZ20]. time-inhomogeneous [DMR04]. time-periodic [HI05]. time-scales [KK13]. time-stepping [KRSS23]. time-varying [BdHNT22, LW14, MW98]. Times [Bra94c, CPR95, FP93, GK95a, KS93b, KS92, RS95, ABT^{+}11a, ABT^{+}11b, AGvdHdH18, BPG19, BC21, BCPG22, BK16a, BK17b, Ben96, Bha16, BdHNT22, CS06, CS11b, CLR06, DK99a, DFK97, Duf16, FPRW18, GNS23, Gri13, HIP06, HDP22, JNP21, Kar15, KS06a, LR14, McD99, NS19, PW21, RŞ18, Sil96, TW09, Wil04, Wyn99]. timescales [LC03]. timing [DG08, DSS96]. tire [BG19]. tissue [FGLS23]. Toeplitz [AJvM22]. together [JT10]. tolerance [KS06b]. tomography [PS16]. top [BCTV07, Goe06a, Jon06a]. top-swap [BCTV07]. Topical [Too02, Mer07]. topological [DP19]. topologies [BLO23]. Topology [MW10, Kar13, PW10, YA15]. Topology-guided [MW10]. Torus [Jai93, CC99, ČT16, EV12, GHPS15]. Total [Jia12, PP23, BS15, BD20, BH12, Clé23]. touch [BL02]. touch-and-out [BL02]. tournament [BCZ19]. trace [DOS19, HL20, HL22]. Tracer [Blo15, KK04]. traces [DRR21]. Tracking [BT12, Bäu00, CRT17, CLR06, Mag00]. Tracy [FJ22, HLS19b, JM12, KQ18, LS16a, Ona08, SX23]. Tracy-Widom [FJ22]. trade [MR17]. trade-off [MR17]. traded [CK19]. tradeoff [AGGL10]. tradeoffs [BM12]. trader [BR17b]. Trading [CHMK20, BK15a, BZ16, BJR08, KX22]. Traffic [Asm92a, BR93, BD01, CPR95, DN94, MRRS02, Ngu93, YH93, Yam95, AK05c, AMR04, Ata05b, AMS06b, AC17, BM20, BBP22, BBC19, BW01, BG05b, BG12, CBM^{+}21, CSS98, DLS01, GZ06, Gro04, GK07a, HW96, Har98, HL97, KLSY03, KLSY04, KLRS11, Kum00, Lim01, LW14, OCBG11, PR10, RR03, RR08a, RT15, Sha15, Sto04, TW09]. Trailing [BD92]. trajectorial [ABP^{+}13, Mél00]. trajectory [Mag00]. trans [RR06]. trans-dimensional [RR06]. Transaction [DPT01, Kus95, SS94, SSC95, ADGS98, BDG16, BD19b, BT00, BT13, CS16b, Dol13, Gua02, GRS08, JB07, KMK10a, LS97, Wee98, Yu17a]. transcendent [JvL07]. transfer [BGY98, CMT21]. Transform [CLW94, EPS09]. transformation [GR97]. transformations [Çet18, KLZ98, OP00]. Transience [Mey95, HI09, HJJ16]. Transient [CLW94, LRT03, AAK17, BYZ00, BD19a, ESTZ13, JLM15, MMPV06]. Transition [AW94, HIP06, LS16b, MR94a, BCOR16, BEG00, BCS19, ČT16, CDS23, CHL21, Com97, FMP09, Fou22, FKKM18, GL18, HSST23, Jár17, Lac17, OCBG11, TVVY12, vdB11]. transitions [AZ10, BLM15, BEMT21, EOT05, GNP17, HMS04, LRZ06, MMX21, RY13, RW12, SY98, Tur18]. Transitive [GK23, DGJ^{+}19]. Translated [Röl07]. translation [HK23, Jia08, Rey18]. Translations [Ber94]. transmission [DG08, RS03b]. transmissions [GR09]. Transport [WW20, AJKH14, BVP22, Bha99, CP19b, FK00a, GO19, GL21, HHK20, HKV20, KX22, NW22, TSM19]. Transport-information [WW20]. Transportation [TY93]. transposition [BCMR21, GQ03, Jia15, NN19]. trap [BF05, FGG14]. Trapping [BHMW16]. Travel [KS93b, LvZ04]. Traveling [BHK11, DM11, EGK22]. traveling-pulse [EGK22]. Travelling [CF94]. Tree [AB92, DPR09, DGP12, DK19, Jai93, JPV99, KOOS91, ABGK12, BB07, BW22a, BM13, BFJ06, DM10, Die15, DFPR22, DL18b, FGG14, Gol16, HLR22, HJJ16, LM96, LZ98, MZ14, MR05, MR10, MR12, MR17, PP04, Pem09, Pen97]. tree-growing [LM96, LZ98]. tree-indexed [BW22a]. tree-like [DM10]. Tree-Structured [KOOS91]. Tree-valued [DGP12]. Treelike [CSS98]. Trees [Ald91, CDJH01, DF95a, DF95b, MS91, Mos01, ABBL09, ABBH14, ABDLV22, ADS14, Ale96, BB22, BBC23, BDMT11, BS12, BS23, BM13, BB15, BGvdHK15, BU18, BMJ06, BG21, BH12, BNS13, BDW22, CS17a, CS11b, DRR21, DDSJ08, DG07, DH06, Die15, DFPR22, DGM06, DJN08, DZ15, EKPS00, GRS04, GT99, GS05, Grü09, Grü14, GK16, GTW22, HH19, HJR20, HK21, HK23, KMS17, KM11, Ker12, KL96, KR19, LP13a, Lee97, LTVR14, MN03, Mar08, MT13, MP03, OWZ97, PS22, PP04, Pen96, Sch10, Sly08, SZ17, SS06, TVVY12]. trials [CX02, GK96]. Triangle [VG95]. triangular [ACH97, CC98a, HHR96, MZ05]. triangulation [Yuk99]. triangulations [CMSS15]. Trie [Dev92b]. Trie-Like [Dev92b]. tries [CM05]. Trimmed [CS95]. triple [FRT14]. Trophic [LC93]. Trouble [Ngu94]. truncation [BS21]. trunk [HL97]. trusted [LPT04]. tube [Adl00, TK02]. tunable [Chi16]. tunneling [DM15]. tuple [KPR10].

Turbulence [FSW95, MSW97]. Turbulent [FK99a, FK00a]. Twitter [BSZ15]. Two [BL01, Cou08, DH93a, FMP95, Hog93, Kot95, Kum00, LW92, Pit99, RR98b, vdBN17, ADH20, AS10, AJO14, APP08, BMST97, BCJ22, BW01, BH20b, BČ05, BFJ06, BKW08, BGZ22, CDN02, CGR20, CX97, CDG23, CDE^{+}17, CP08, Cox10, DHV04, DFM18, DH07, DFT03, DFPR22, FS16, FRSW22, Fou00, FJ17, GM05, GS20, JLP08, JS12, Jin19, KT04, KS06a, Kro99, LRPY22, LPT04, LL17, LL20b, LM02, LM05, MM03, Mat05, Mél00, MAL14, MP06, PW21, Per00, Pes19, RYZ21, SSW06, SS06, TV12, Tom21, vdB11]. two-armed [LPT04, TV12]. two-color [Mat05]. two-community [BH20b]. Two-Dimensional [FMP95, Kot95, vdBN17, AJO14, APP08, BGZ22, CDN02, CX97, CP08, Cox10, Fou00, FJ17, GS20, MM03, Mél00, PW21, Pes19, RYZ21, Tom21, vdB11]. Two-Factor [Hog93]. two-locus [JS12]. Two-Moment [DH93a]. two-parameter [CDE^{+}17, FRSW22]. Two-Particle [BL01]. two-periodic [BCJ22]. two-player [DFM18]. two-scale [LRPY22]. Two-server [Kum00]. two-sided [MAL14]. two-site [SSW06]. two-species [Per00]. two-stage [Kro99]. two-station [DHV04]. two-time-scale [KT04, MP06]. two-times [KS06a]. two-type [ADH20, DH07, GM05]. Type [Aka95, Ath94, O’C91, ADH20, AB05, BZ15, BNT19, BL02, Cer09, CCD19, Chi04, DFM18, De 11, DH07, DIRT15, FLP13, GGLO13, GM05, GLP23, HK22, HSZ17, HIP08, Hof05, JCLP21, Jam10, LPSX23, Lez98, LPP15, Mei09, Mül18, RS23, SZZ21]. typed [GHH07]. types [BL08, BH19]. Typical [GSS13, vdHOC18]. typically [BS20].

Uhlenbeck [CW93, DG13]. ultimate [dTP09]. ultrametric [FRLS21]. Un- [CDS23]. Unadjusted [PP23, DM17]. unbalanced [RR08a]. Unbiased [HLTT17]. Unbiasedness [BGG^{+}16]. Unbounded [MR94b, Bar19, BF04b, BC18, DH07, Gra21, Mon22, SV10, Tra23, Žit05]. uncertainty [BDG16, Bar19, BH13, BK16a, BK17b, BZ19, BNK12, DM06b, DSS96, FK11, HXZ22, MPZ21, PZ19]. Unconstrained [LP10]. unconventional [HW96]. underlying [Aus08]. undershoots [DK06]. Understanding [BT22b]. Unified [AB92, Dai95, BF08, FHH23, MWZZ15]. Uniform [BJM10b, CC93, Chi07, FPRW18, Fuh04, GLWZ22, MW98, SV02, vH09, DT18, DS22, DHESC21, Duf16, HMGR00, HPS03, LO04, MG02, RR98a, SSV18]. uniformly [DRZ16, DW05a]. unifying [FRLS21]. unimodular [DGJ^{+}19]. unique [CvH10, JK14, KR06, Yu17b]. Uniqueness [Ana93, Ana94, HL01, MR94b, Sly08, Xu18, CCCS11, Hol98, KTPZ15, MM01, Nor99, Tei09, TY11]. unit [CLSF18]. Universal [BDM17, WM04, MRV21, MW08]. Universality [BLM15, CSZ17, DHL23, DLZ21, FRT14, PY14, Woo12, DY18, EGP16, Lac22]. universe [MSW97]. unlabeled [WXY23]. Unoriented [Mar16]. unreliable [ARS18, SFR16]. unstable [BPG19, BCPG22, Bra99, JR15, JR16, vH09]. up-crossing [LS18]. updating [Che18, NR06]. upon [Nea06]. Upper [Bal19, CT11, CGGK93, Kel13, Sch05b, Ott13]. ups [HLS19a]. upward [Miy04]. urn [BHZ02, BH05, BT22a, Bha16, CCL13, GM19, GVR17, LP13b, LP17a, ZHC06, BST04]. Urns [KL91, CPS11, vdHHKR16]. USC [RR91]. used [Fil13, Jel99]. user [BC01]. using [BZ16, DJ12, FK16, Hub04, JPS^{+}22, KMK10a, KM13, KMS22, LP13b, LP17a, LO04, Mar19, Röl07, WH93, Xia97]. utilities [EPQ01]. Utility [DPT01, FI20, HIM05, KMK10b, Owe02, Yu15, Žit05, Bar19, BZ16, Bec06, BF08, BC18, BTZ04, FS99, HP15, Kni12, KS99b, KS06b, LŽ13, Laz04, MS05, Nut12, RS05, Sch04]. utility-based [KS06b].

Vacant [MR94b, CFJ00, Tei09]. Vacations [KW91]. validation [MW07]. Validity [GZ06, GW92, KT03]. valley [BCS19]. valuation [DFT03, DSS96, MS05]. Value [BM01, DS93, FM94, Tak93, AJO14, BG06b, BG12, DP20, DMP22, DH18, DFdH04, EV06, KS06a, LST23, Niu97, SV02]. Valued [KX95, ABKR18, AdHR21, DR02, Del98, DD10, DGP12, FRSW22, MV20, RT15, RZ23, SSV18, dLS97]. values [BGBP23, CR18, DY18, GRS^{+}16]. Varadhan [FK16, GM13]. variable [CF09]. variable-range [CF09]. Variables [BOW95, HST91, Kee94, VG95, ACH97, BK00a, Dre00, HHR96, Jor02, Kol17, LR99, LN13, Yat09, Yu17b]. Variance [DG21, DR91, HKK06, RR08b, RS91, Sch92, BCHL98, BN17, Bla96, BFJ06, CW13, DL18a, HLN08, HK13, JMSS12, Keb05, Lef04, LZ06, PR98b, RR14, Wu09]. Variance-optimal [HKK06]. variates [Mey06]. Variation [Gre94, BDM02, Clé23, DTW22, JMRS09, Jia12, PP23]. Variational [AG93, GdH93, CN23, SS06]. variations [Lud08, TT11]. variety [BF22b]. various [HM14]. Varying [CJ94, BdHNT22, Coh96, DHS18, DM20, FL22, HLMS05, Jon97, LW14, MW98, OCBG11, Sab16, YS22]. vector [BS19, BP97, FHLN22, Kar07]. Vectors [Ber94, Col02, BCHL98, DM20, HR07, JJ15, JL09, LT18]. Vehicle [Rhe94, FT17]. vehicle-sharing [FT17]. velocity [CV21, Goe06b, IV17, KK04]. verifiable [DMO14]. Verification [FG18]. version [Arm10, BK06, GM13, HS19a, NP99]. Versus [AR02, Jai93, BL06, BP18, CX02, GY04, Sly08]. Vertex [CNV22, Pen00, BS12, Die15, FP17, Gol13, JW21, Sch21]. vertex-cut-tree [Die15]. Vertex-reinforced [CNV22]. vertices [HRW08]. Very [RSX99, HHK21, RR00, Viv23]. Via [BGT01, Dai95, Mey95, Alb09, AFKP20, ABDW21, Bel13, BCD17, BG96, CST22, CDS09, CDV14, DR13, DS19, DFM16, DGR09, FK22, Fil98, FI20, Haa10, HPP22, Hob98, HL97, JLR03, JMSS12, JM02, JY17, KLP15, LSZ13, MLNW22, Nag12, QS94, RR14, SS19, Sei09]. view [BL12a, Bät99, Har03a]. Viot [DGP12, DK99b, EV12, EFPS17, KR23, WHN07]. Viral [ACD15]. Virtual [ZZ02]. virus [BS15, DGM08]. Viscosity [Ren16, WZ20, See20]. viscous [BJ22]. Vlasov [AKH02, AD20, BPR22, BLM23, CCD19, CDFM20, CGK^{+}23, CM19, DIRT15, FP22, HL17, HLS19a, KNRS22, STT19, TV03, Tom21, WZ20, dRST19]. vol [Ano99, Ano02, Ano03]. Volatility [BJR16, FKR18, Hob98, JT03, PQ01, ADS14, BDG16, BD98, CLR06, FFK12, FK16, FGP21, HX22, JT10, JT18, Pal11, Tod19]. volatility-stabilized [Pal11]. Volterra [BL16b, CP08, JLP19, JMP21, JCLP21, KK01, KL01, NP99]. Volume [Ano99, Ano02, Ano03, Hei05, vdBN17]. volume-frozen [vdBN17]. volumes [GNP17]. Voronoi [CR18, Gol10, GG97, HR09, MSW97]. vortex [Fon10, Mél00]. Vorticity [Kot95, HLN21a]. vote [BG21]. Voter [SS08, BS17a, BLZ11, CP14, CPS16, HD19, SSY19]. VRRW [LV10]. vs [CS23].

Waiting [DR92, DS07b, Bla96, DK99a, DSS09, Fla97, TW09, Wyn99]. Waiting-Time [DR92]. Walk [BLSW91, Big95, BNT92, GdH93, Jof93, McD95, QS94, Tak92, AG06, BW22a, BRTP18, Big12, BG06a, BEM07, BHdS^{+}20, BT12, CF07, CMY03, DH13b, EFdS21, FZ03, FPZ05, GGR97, Han06b, JPS^{+}22, JvL07, Jia12, Jia15, JLM15, LST23, MPS12, Mat05, MPP17, McD99, Mei09, MS00, NRY12, Oli09, PW21, PSY15, PS17a, PS20, Win08]. Walks [BM05, Cha92, Che01, ACD15, ABF13, ABH17, AGvdHdH18, BDW23, BMRS23, BBRZ20, BPZ07, BZ15, BS17b, BP18, BG08, BHdS^{+}20, CS16a, ČT16, CL03, CF17, CL09a, ESTZ13, FKK^{+}01, Fuh04, GS19, GK16, HLMS05, Lud08, MM03, MS18, Nak11, Sch21, She02, Viv23, YZ19, Yos08]. Walsh [AC19]. Wang [BCP^{+}22, JR14]. WARM [HHK21]. Wasserstein [Xu19, BBBW22, BNT19, But14, FJR20, KSS21, NRS21, PP23, QH22, Wan23]. Wasserstein-1 [Xu19]. Wasserstein-based [QH22]. Wasserstein-type [BNT19]. wasted [Rhe00]. Watanabe [Sch13]. Watson [Als93, AR96, AR02, BDMT11, BS12, BM13, Chi04, Die15, HJR20, KR19, Mar08, QS94]. Wave [Cho02, ARL08, CDG23, Cho06, Cho09, GS09, MM01, SCZ10]. wave-breaking [CDG23]. wave-like [SCZ10]. wavelet [Jaf00]. waves [BHK11, DM11, MS11]. Weak [Asm92b, BN15a, Béd07, BJ22, BvdH12, CST22, CJK19, GGR97, LO13, LO17, MP95, Pel98, PY03, PT15, Pov95, Ser94, vdHMS08, BVP22, CL15, CKHL06, CG17, CSS22, Dol10, GvdHL20, GK16, JCLP21, LRW23, LS09, OTV12, PTZ20, TY16, TKH09, TY11]. weaker [LL20b]. Weakly [BGJ18, BFG13, BT96, BF18, GLM17, HP20, LL20a]. wealth [FS99, Kar13, KS06b, Sch01]. wealth-process [Kar13]. web [RSS16]. Weibull [BBRZ20]. weight [SW12]. Weighted [Dre00, PP18, RS01, ABGK12, AL15, FKP23, KL96, OWZ97, PS22, RR97a]. weights [BCPR15, BUV11, EFT07, JPV99, Pal11, RV15, TY16, TV03]. Weiss [BR17a, CS11a, FMP00, SSV18]. Well [CDG23, KNRS22, BG02, FHLN22, MWZZ15]. Well-posedness [CDG23, KNRS22, FHLN22, MWZZ15]. Wentzell [dRST19]. wetting [GL18]. Which [DL18a, HH08, HK23]. white [GS09, LØP04, Zha16b]. white-noise [GS09]. Whitt [ABP15, AP16, AG12, BM19a, BM20, GG13a, GG13b, Ree09]. Whittaker [Che22b]. Whittle [HS99]. whose [BH99, DH04]. wide [NP01]. wide-sense [NP01]. Widom [FJ22, HLS19b, JK17, JM12, KQ18, LS16a, Ona08, SX23]. Width [DH06, OWZ97]. Wiener [LO17, BVLT20, BL14, BM19b, CM19, Kuz10, KKPvS11, LO13, LL12, MWZ07, RSM09, Rog94, Sez10, ZRH15, ZS09]. Wigner [CES23, HK17]. Williams [LTVR14]. WIMPs [BBPS15]. window [PSY15]. winner [DvdH16]. winners [CDL17, KP96, KP98]. wireless [BS07, BBFM03, IV17]. Wishart [AA13, FK22, Ona08, SYY20]. within [AS10, BL06, RR06]. without [BSZ20, Fou00, GS14, Gua02, JW21, JPS09, MMS20]. Wold [FMP95]. Wold-Like [FMP95]. WoMS [DH13b]. Wong [GT19]. Woodroofe [GZ09]. word [BKW08]. words [BC22b]. Work [Che95]. Work-Conserving [Che95]. Workload [HW05, GZ00, HV97, Har00, Har03b, Har06, Sto04]. world [GK07b, San08]. worlds [CT23]. Wright [CS18, CPSJ22, Cer15, Che18, CDE^{+}17, GR21, JS17, Pap98, Pap00].

XORSAT [IKKM15].

Ylvisaker [RWW95]. Yor [GZ08, KTT17, NS22]. Young [Pia99].

zag [BNS21]. Zakai [GT19]. Zakharov [BB20a]. zealot [HD19]. Zero [CPR95, FJ19, JNP21, PTZ20, BS19, BFS21, CDN02, Goe06b, GR97, Gra09, HS19a, HS21a, HHSZ15, Tra23]. zero-range [BFS21, Gra09, HS21a, Tra23]. Zero-sum [PTZ20, HHSZ15]. zero-temperature [CDN02]. zeta [ABH17]. zig [BNS21]. zig-zag [BNS21]. zigzag [BRZ19]. Zonoids [Vit91]. Zooming [Iva18].

References

Ahdida:2013:EHO

[AA13] Abdelkoddousse Ahdida and Aurélien Alfonsi. Exact and highorder discretization schemes for Wishart processes and their affine extensions. Annals of Applied Probability, 23(3):1025– 1073, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-

             8737 (electronic). URL http://projecteuclid.org/euclid.
             aoap/1362684853.

Adhikari:2022:SDR

[AABR22] Kartick Adhikari, Robert J. Adler, Omer Bobrowski, and Ron Rosenthal. On the spectrum of dense random geometric graphs. Annals of Applied Probability, 32(3):1734–1773, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

             (electronic). URL https://projecteuclid.org/journals/
             annals-of-applied-probability/volume-32/issue-3/On-the-

spectrum-of-dense-random-geometric-graphs/10.1214/21-
AAP1720.full.

Amir:2017:ODL

[AAK17] Gideon Amir, Omer Angel, and Gady Kozma. One-dimensional long-range diffusion limited aggregation II: The transient case. Annals of Applied Probability, 27(3):1886–1922, June 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1500451246.

Aalen:1992:MHS

[Aal92] Odd O. Aalen. Modelling heterogeneity in survival analysis by the compound Poisson distribution. Annals of Applied Probability, 2(4):951–972, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005583.

Avram:1992:MST

[AB92] Florin Avram and Dimitris Bertsimas. The minimum spanning tree constant in geometrical probability and under the independent model: a unified approach. Annals of Applied Probability, 2(1):113–130, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005773.

Avram:1993:CLT

[AB93] Florin Avram and Dimitris Bertsimas. On central limit theorems in geometrical probability. Annals of Applied Probability, 3(4): 1033–1046, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005271.

Aldous:2005:SMT

[AB05] David J. Aldous and Antar Bandyopadhyay. A survey of max-type recursive distributional equations. Annals of Applied Probability, 15(2):1047–1110, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1115137969.

Atar:2014:CMQ

[AB14] Rami Atar and Anup Biswas. Control of the multiclass G/G/1 queue in the moderate deviation regime. Annals of Applied Probability, 24(5):2033–2069, 10 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Arapostathis:2022:RSC

[AB22] Ari Arapostathis and Anup Biswas. Risk-sensitive control for a class of diffusions with jumps. Annals of Applied Probability, 32 (6):4106–4142, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-6/Risk-sensitive-control-for-a-class-of-diffusions-
with-jumps/10.1214/21-AAP1758.full.

Abbe:2020:IPB

[ABA20] Emmanuel Abbe and Enric Boix-Adserà. An informationpercolation bound for spin synchronization on general graphs. Annals of Applied Probability, 30(3):1066–1090, June 2020. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/journals/annals-of-applied-
probability/volume-30/issue-3/An-information-percolation-
bound-for-spin-synchronization-on-general-graphs/10.
1214/19-AAP1523.full.

Alsmeyer:2023:ALI

[ABB23] Gerold Alsmeyer, Sara Brofferio, and Dariusz Buraczewski. Asymptotically linear iterated function systems on the real line. Annals of Applied Probability, 33(1):161–199, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Asymptotically-
linear-iterated-function-systems-on-the-real-line/10.
1214/22-AAP1812.full.

Addario-Berry:2014:CTM

[ABBH14] Louigi Addario-Berry, Nicolas Broutin, and Cecilia Holmgren. Cutting down trees with a Markov chainsaw. Annals of Applied Probability, 24(6):2297–2339, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1409058033.

Alon:1994:LER

[ABBJ94] Noga Alon, Bela Bollobas, Graham Brightwell, and Svante Janson. Linear extensions of a random partial order. Annals of Applied Probability, 4(1):108–123, February 1994. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005202.

Addario-Berry:2009:ERR

[ABBL09] Louigi Addario-Berry, Nicolas Broutin, and Gábor Lugosi. Effective resistance of random trees. Annals of Applied Probability, 19(3):1092–1107, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1245071020.

Addario-Berry:2022:BRR

[ABDLV22] Louigi Addario-Berry, Luc Devroye, Gábor Lugosi, and Vasiliki Velona. Broadcasting on random recursive trees. Annals of Applied Probability, 32(1):497–528, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-1/Broadcasting-on-random-recursive-trees/
10.1214/21-AAP1686.full.

Atar:2021:RBO

[ABDW21] Rami Atar, Amarjit Budhiraja, Paul Dupuis, and Ruoyu Wu. Robust bounds and optimization at the large deviations scale for queueing models via Rényi divergence. Annals of Applied Probability, 31(3):1061–1099, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-3/Robust-bounds-and-optimization-at-
the-large-deviations-scale-for/10.1214/20-AAP1613.full.

Addario-Berry:2013:PDB

[ABF13] Louigi Addario-Berry and Kevin Ford. Poisson–Dirichlet branching random walks. Annals of Applied Probability, 23 (1):283–307, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1359124387.

Addario-Berry:2012:IPP

[ABGK12] Louigi Addario-Berry, Simon Griffiths, and Ross J. Kang. Invasion percolation on the Poisson-weighted infinite tree. Annals of Applied Probability, 22(3):931–970, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1337347535.

Arguin:2017:MRR

[ABH17] Louis-Pierre Arguin, David Belius, and Adam J. Harper. Maxima of a randomized Riemann zeta function, and branching random walks. Annals of Applied Probability, 27(1):178–215, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1488790826.

Arguin:2012:PSE

[ABK12] Louis-Pierre Arguin, Anton Bovier, and Nicola Kistler. Poissonian statistics in the extremal process of branching Brownian motion. Annals of Applied Probability, 22(4):1693–1711, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1344614208.

Atar:2018:SMM

[ABKR18] Rami Atar, Anup Biswas, Haya Kaspi, and Kavita Ramanan. A Skorokhod map on measure-valued paths with applications to priority queues. Annals of Applied Probability, 28(1):418–481, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1520046092.

Adler:2012:EMC

[ABL12] Robert J. Adler, Jose H. Blanchet, and Jingchen Liu. Efficient Monte Carlo for high excursions of Gaussian random fields. Annals of Applied Probability, 22(3):1167–1214, June 2012. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1337347542.

Aurzada:2021:BCI

[ABL21] Frank Aurzada, Volker Betz, and Mikhail Lifshits. Breaking a chain of interacting Brownian particles. Annals of Applied Probability, 31(6):2585–2611, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-6/Breaking-a-chain-of-interacting-Brownian-
particles/10.1214/20-AAP1658.full.

Addario-Berry:2022:BBB

[ABLT22] Louigi Addario-Berry, Jessica Lin, and Thomas Tendron. Barycentric Brownian bees. Annals of Applied Probability, 32(4):

2504–2539, August 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
4/Barycentric-Brownian-bees/10.1214/21-AAP1738.full.

Acciaio:2013:TID

[ABP^{+}13] B. Acciaio, M. Beiglböck, F. Penkner, W. Schachermayer, and J. Temme. A trajectorial interpretation of Doob’s martingale inequalities. Annals of Applied Probability, 23(4):1494–1505, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1371834036.

Arapostathis:2015:ECM

[ABP15] Ari Arapostathis, Anup Biswas, and Guodong Pang. Ergodic control of multi-class M/M/N + M queues in the Halfin–Whitt regime. Annals of Applied Probability, 25(6):3511–3570, 12 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Arratia:1992:PPA

[ABT92] Richard Arratia, A. D. Barbour, and Simon Tavare. Poisson process approximations for the Ewens sampling formula. Annals of Applied Probability, 2(3):519–535, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005647.

Arratia:2000:NCL

[ABT00] Richard Arratia, A. D. Barbour, and Simon Tavaré. The number of components in a logarithmic combinatorial structure. Annals of Applied Probability, 10(2):331–361, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1019487347.

Appuhamillage:2011:COL

[ABT^{+}11a] Thilanka Appuhamillage, Vrushali Bokil, Enrique Thomann, Edward Waymire, and Brian Wood. Corrections: “Occupation and local times for skew Brownian motion with applications to dispersion across an interface”. Annals of Applied Probability, 21(5):2050–2051, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576617. See corrections [ABT^{+}11b].

Appuhamillage:2011:OLT

[ABT^{+}11b] Thilanka Appuhamillage, Vrushali Bokil, Enrique Thomann, Edward Waymire, and Brian Wood. Occupation and local times for skew Brownian motion with applications to dispersion across an interface. Annals of Applied Probability, 21(1):183–214, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1292598031. See corrections [ABT^{+}11a].

Atar:2007:HEC

[ABW07] Rami Atar, Amarjit Budhiraja, and Ruth J. Williams. HJB equations for certain singularly controlled diffusions. Annals of Applied Probability, 17(5/6):1745–1776, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1191419182.

Asselah:2003:EQS

[AC03] Amine Asselah and Fabienne Castell. Existence of quasistationary measures for asymmetric attractive particle systems on Z^{d}. Annals of Applied Probability, 13(4):1569–1590, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1069786511.

Atar:2017:AOC

[AC17] Rami Atar and Asaf Cohen. Asymptotically optimal control for a multiclass queueing model in the moderate deviation heavy traffic regime. Annals of Applied Probability, 27(5):2862–2906, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1509696036.

Atar:2019:SSQ

[AC19] Rami Atar and Asaf Cohen. Serve the shortest queue and Walsh Brownian motion. Annals of Applied Probability, 29(1):613–651, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1544000438.

Allan:2020:PSC

[AC20] Andrew L. Allan and Samuel N. Cohen. Pathwise stochastic control with applications to robust filtering. Annals of Applied Probability, 30(5):2274–2310, October 2020. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-5/Pathwise-stochastic-control-with-applications-
to-robust-filtering/10.1214/19-AAP1558.full.

Abdullah:2015:VPR

[ACD15] Mohammed Abdullah, Colin Cooper, and Moez Draief. Viral processes by random walks on random regular graphs. Annals of Applied Probability, 25(2):477–522, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1424355121.

Aletti:2017:SRS

[ACG17] Giacomo Aletti, Irene Crimaldi, and Andrea Ghiglietti. Synchronization of reinforced stochastic processes with a network-based interaction. Annals of Applied Probability, 27(6):3787–3844, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1513328714.

Anderson:1997:MPL

[ACH97] Clive W. Anderson, Stuart G. Coles, and Jürg Hüsler. Maxima of Poisson-like variables and related triangular arrays. Annals of Applied Probability, 7(4):953–971, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1043862420.

Abate:1995:AAT

[ACLW95] Joseph Abate, Gagan L. Choudhury, David M. Lucantoni, and Ward Whitt. Asymptotic analysis of tail probabilities based on the computation of moments. Annals of Applied Probability, 5 (4):983–1007, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004603.

Arnaudon:2020:SOA

[AD20] M. Arnaudon and P. Del Moral. A second order analysis of McKean–Vlasov semigroups. Annals of Applied Probability, 30 (6):2613–2664, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-6/A-second-order-analysis-of-McKeanVlasov-semigroups/
10.1214/20-AAP1568.full.

Agazzi:2018:LDT

[ADE18] Andrea Agazzi, Amir Dembo, and Jean-Pierre Eckmann. Large deviations theory for Markov jump models of chemical reaction networks. Annals of Applied Probability, 28(3):1821–1855, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1527840033.

Ahn:1998:ORT

[ADGS98] Hyungsok Ahn, Mohit Dayal, Eric Grannan, and Glen Swindle. Option replication with transaction costs: general diffusion limits. Annals of Applied Probability, 8(3):676–707, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1028903447.

Auffinger:2015:LGF

[ADH15] Antonio Auffinger, Michael Damron, and Jack Hanson. Limiting geodesics for first-passage percolation on subsets of Z^{2}. Annals of Applied Probability, 25(1):373–405, 02 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Ahlberg:2020:TTR

[ADH20] Daniel Ahlberg, Maria Deijfen, and Christopher Hoffman. The two-type Richardson model in the half-plane. Annals of Applied Probability, 30(5):2261–2273, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-5/The-two-type-Richardson-model-in-the-
half-plane/10.1214/19-AAP1557.full.

Athreya:2021:GVS

[AdHR21] Siva Athreya, Frank den Hollander, and Adrian Röllin. Graphon-valued stochastic processes from population genetics. Annals of Applied Probability, 31(4):1724–1745, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/Graphon-
valued-stochastic-processes-from-population-genetics/
10.1214/20-AAP1631.full.

Adler:2000:EST

[Adl00] Robert J. Adler. On excursion sets, tube formulas and maxima of random fields. Annals of Applied Probability, 10(1):1–74, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1019737664.

Andrieu:2021:HPD

[ADNR21] Christophe Andrieu, Alain Durmus, Nikolas Nüsken, and Julien Roussel. Hypocoercivity of piecewise deterministic Markov process-Monte Carlo. Annals of Applied Probability, 31(5):2478– 2517, October 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
5/Hypocoercivity-of-piecewise-deterministic-Markov-process-
Monte-Carlo/10.1214/20-AAP1653.full.

Alimohammadi:2023:SIS

[ADRS23] Yeganeh Alimohammadi, Persi Diaconis, Mohammad Roghani, and Amin Saberi. Sequential importance sampling for estimating expectations over the space of perfect matchings. Annals of Applied Probability, 33(2):799–833, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Sequential-
importance-sampling-for-estimating-expectations-over-
the-space-of/10.1214/22-AAP1834.full.

Assaf:2011:RTR

[ADS11] Sami Assaf, Persi Diaconis, and K. Soundararajan. A rule of thumb for riffle shuffling. Annals of Applied Probability, 21(3):843–875, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1307020385.

Akyildirim:2014:ASV

[ADS14] Erdinç Akyildirim, Yan Dolinsky, and H. Mete Soner. Approximating stochastic volatility by recombinant trees. Annals of Applied Probability, 24(5):2176–2205, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1403812373.

Alt:2018:LIC

[AEK18] Johannes Alt, László Erdős, and Torben Krüger. Local inhomogeneous circular law. Annals of Applied Probability, 28(1):148– 203, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1520046086.

Ankirchner:2020:OPT

[AFKP20] Stefan Ankirchner, Alexander Fromm, Thomas Kruse, and Alexandre Popier. Optimal position targeting via decoupling fields. Annals of Applied Probability, 30(2):644–672, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1591603218.

Antunes:2006:ALN

[AFRT06] Nelson Antunes, Christine Fricker, Philippe Robert, and Danielle Tibi. Analysis of loss networks with routing. Annals of Applied Probability, 16(4):2007–2026, November 2006. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1169065214.

Almeida:1993:VME

[AG93] Murilo P. Almeida and Basilis Gidas. A variational method for estimating the parameters of MRF from complete or incomplete data. Annals of Applied Probability, 3(1):103–136, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005510.

Araman:2006:TAM

[AG06] Victor F. Araman and Peter W. Glynn. Tail asymptotics for the maximum of perturbed random walk. Annals of Applied Probability, 16(3):1411–1431, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1159804986.

Awad:2009:CLT

[AG09] Hernan Awad and Peter Glynn. Conditional limit theorems for regulated fractional Brownian motion. Annals of Applied Probability, 19(6):2102–2136, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1259158768.

Ata:2012:OGH

[AG12] Baris Ata and Itai Gurvich. On optimality gaps in the Halfin– Whitt regime. Annals of Applied Probability, 22(1):407–455, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1328623704.

Atar:2014:SPS

[AG14] Rami Atar and Itai Gurvich. Scheduling parallel servers in the nondegenerate slowdown diffusion regime: Asymptotic optimality results. Annals of Applied Probability, 24(2):760–810, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1394465371.

Arnold:1994:EFP

[AGD94] Ludwig Arnold, Volker Matthias Gundlach, and Lloyd Demetrius. Evolutionary formalism for products of positive random matrices. Annals of Applied Probability, 4(3):859–901, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177004975.

Alon:2010:CMT

[AGGL10] Noga Alon, Ori Gurel-Gurevich, and Eyal Lubetzky. Choicememory tradeoff in allocations. Annals of Applied Probability, 20(4):1470–1511, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1279638792.

Anderson:2011:EAT

[AGK11] David F. Anderson, Arnab Ganguly, and Thomas G. Kurtz. Error analysis of tau-leap simulation methods. Annals of Applied Probability, 21(6):2226–2262, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1322057321.

Asmussen:1995:DES

[AGP95] Soren Asmussen, Peter Glynn, and Jim Pitman. Discretization error in simulation of one-dimensional reflecting Brownian motion. Annals of Applied Probability, 5(4):875–896, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004597.

Assaf:2002:RPI

[AGSC02] David Assaf, Larry Goldstein, and Ester Samuel-Cahn. Ratio prophet inequalities when the mortal has several choices. Annals of Applied Probability, 12(3):972–984, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1031863177.

Avena:2018:MTR

[AGvdHdH18] Luca Avena, Hakan Güldaş, Remco van der Hofstad, and Frank den Hollander. Mixing times of random walks on dynamic configuration models. Annals of Applied Probability, 28(4):1977–2002, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1533780265.

Alanyali:1998:LDL

[AH98a] Murat Alanyali and Bruce Hajek. On large deviations in load sharing networks. Annals of Applied Probability, 8(1):67–97, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1027961034.

Alanyali:1998:LDM

[AH98b] Murat Alanyali and Bruce Hajek. On large deviations of Markov processes with discontinuous statistics. Annals of Applied Probability, 8(1):45–66, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1027961033.

Ata:2005:DRC

[AHS05] Bariş Ata, J. M. Harrison, and L. A. Shepp. Drift rate control of a Brownian processing system. Annals of Applied Probability, 15(2):1145–1160, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1115137971.

Alfonsi:2014:POT

[AJKH14] A. Alfonsi, B. Jourdain, and A. Kohatsu-Higa. Pathwise optimal transport bounds between a one-dimensional diffusion and its Euler scheme. Annals of Applied Probability, 24(3):1049–1080, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1398258095.

Assing:2014:MVF

[AJO14] Sigurd Assing, Saul Jacka, and Adriana Ocejo. Monotonicity of the value function for a two-dimensional optimal stopping problem. Annals of Applied Probability, 24(4):1554–1584, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1400073657.

Adler:2022:STD

[AJvM22] Mark Adler, Kurt Johansson, and Pierre van Moerbeke. A singular Toeplitz determinant and the discrete tacnode kernel for skew-aztec rectangles. Annals of Applied Probability, 32(2): 1234–1294, April 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
2/A-singular-Toeplitz-determinant-and-the-discrete-tacnode-
kernel-for/10.1214/21-AAP1708.full.

Alili:2005:SRF

[AK05a] L. Alili and A. E. Kyprianou. Some remarks on first passage of Lévy processes, the American put and pasting principles. Annals of Applied Probability, 15(3):2062–2080, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1121433777.

Arora:2005:LMS

[AK05b] Sanjeev Arora and Ravi Kannan. Learning mixtures of separated nonspherical Gaussians. Annals of Applied Probability, 15(1A):69–92, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1106922321.

Ata:2005:HTA

[AK05c] Baris Ata and Sunil Kumar. Heavy traffic analysis of open processing networks with complete resource pooling: Asymptotic optimality of discrete review policies. Annals of Applied Probability, 15(1A):331–391, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1106922331.

Angel:2018:STC

[AK18] Omer Angel and Brett Kolesnik. Sharp thresholds for contagious sets in random graphs. Annals of Applied Probability, 28(2):1052–1098, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1523433631.

Akahori:1995:SFN

[Aka95] Jiro Akahori. Some formulae for a new type of path-dependent option. Annals of Applied Probability, 5(2):383–388, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177004769.

Antonelli:2002:RCP

[AKH02] Fabio Antonelli and Arturo Kohatsu-Higa. Rate of convergence of a particle method to the solution of the McKean–Vlasov equation. Annals of Applied Probability, 12(2):423–476, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1026915611.

Amendola:2022:CIM

[AKLT22] Carlos Améndola, Claudia Klüppelberg, Steffen Lauritzen, and Ngoc M. Tran. Conditional independence in max-linear Bayesian networks. Annals of Applied Probability, 32(1):1–45, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Conditional-
independence-in-max-linear-Bayesian-networks/10.1214/
21-AAP1670.full.

Avram:2004:EPS

[AKP04] F. Avram, A. E. Kyprianou, and M. R. Pistorius. Exit problems for spectrally negative Lévy processes and applications to (canadized) Russian options. Annals of Applied Probability, 14 (1):215–238, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1075828052.

Arratia:2005:HLD

[AL05] Richard Arratia and Thomas M. Liggett. How likely is an i.i.d. degree sequence to be graphical? Annals of Applied Probability, 15(1B):652–670, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1107271663.

Amini:2015:DWR

[AL15] Hamed Amini and Marc Lelarge. The diameter of weighted random graphs. Annals of Applied Probability, 25(3):1686–1727, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1427124140.

Acciaio:2017:SSC

[AL17] Beatrice Acciaio and Martin Larsson. Semi-static completeness and robust pricing by informed investors. Annals of Applied Probability, 27(4):2270–2304, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1504080033.

Alanyali:2003:AEA

[Ala03] Murat Alanyali. Asymptotically exact analysis of a loss network with channel continuity. Annals of Applied Probability, 13(4): 1474–1493, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1069786506.

Albin:2004:SSI

[Alb04] J. M. P. Albin. On sampling of stationary increment processes. Annals of Applied Probability, 14(4):2016–2037, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1099674087.

Albenque:2009:NED

[Alb09] Marie Albenque. A note on the enumeration of directed animals via gas considerations. Annals of Applied Probability, 19 (5):1860–1879, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1255699546.

Aldous:1991:AFD

[Ald91] David Aldous. Asymptotic fringe distributions for general families of random trees. Annals of Applied Probability, 1(2):228–266, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005936.

Alexander:1993:NSR

[Ale93] Kenneth S. Alexander. A note on some rates of convergence in first-passage percolation. Annals of Applied Probability, 3(1):81– 90, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005508.

Alexander:1994:RCMb

[Ale94a] Kenneth S. Alexander. The rate of convergence of the mean length of the longest common subsequence. Annals of Applied Probability, 4(4):1074–1082, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004903. See correction

[Ale95].

Alexander:1994:RCMa

[Ale94b] Kenneth S. Alexander. Rates of convergence of means for distance-minimizing subadditive Euclidean functionals. Annals of Applied Probability, 4(3):902–922, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004976.

Alexander:1995:CRC

[Ale95] K. S. Alexander. Correction: “The Rate of Convergence of the Mean Length of the Longest Common Subsequence”. Annals of Applied Probability, 5(1):327, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004844. See [Ale94a].

Alexander:1996:RTC

[Ale96] Kenneth S. Alexander. The RSW theorem for continuum percolation and the CLT for Euclidean minimal spanning trees. Annals of Applied Probability, 6(2):466–494, May 1996. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1034968140.

Alsmeyer:1993:GWP

[Als93] Gerold Alsmeyer. On the Galton–Watson predator-prey process. Annals of Applied Probability, 3(1):198–211, February 1993. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005515.

Alvarez:2003:PEM

[Alv03] Luis H. R. Alvarez. On the properties of r-excessive mappings for a class of diffusions. Annals of Applied Probability, 13(4): 1517–1533, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1069786509.

Andrieu:2006:EPS

[AM06] Christophe Andrieu and Éric Moulines. On the ergodicity properties of some adaptive MCMC algorithms. Annals of Applied Probability, 16(3):1462–1505, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1159804988.

Azcue:2010:OIP

[AM10] Pablo Azcue and Nora Muler. Optimal investment policy and dividend payment strategy in an insurance company. Annals of Applied Probability, 20(4):1253–1302, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1279638786.

Alves:2002:STF

[AMP02] O. S. M. Alves, F. P. Machado, and S. Yu. Popov. The shape theorem for the frog model. Annals of Applied Probability, 12(2): 533–546, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1026915614.

Atar:2004:SMC

[AMR04] Rami Atar, Avi Mandelbaum, and Martin I. Reiman. Scheduling a multi class queue with many exponential servers: asymptotic optimality in heavy traffic. Annals of Applied Probability, 14 (3):1084–1134, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1089736279.

Andjel:2006:PRP

[AMS06a] E. D. Andjel, M. V. Menshikov, and V. V. Sisko. Positive recurrence of processes associated to crystal growth models. Annals of Applied Probability, 16(3):1059–1085, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1159804975.

Atar:2006:QSM

[AMS06b] Rami Atar, Avi Mandelbaum, and Gennady Shaikhet. Queueing systems with many servers: Null controllability in heavy traffic. Annals of Applied Probability, 16(4):1764–1804, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1169065207.

Alberts:2022:DRS

[AN22] Tom Alberts and Raoul Normand. Dimension results for the spectral measure of the circular β ensembles. Annals of Applied Probability, 32(6):4642–4680, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-6/Dimension-results-for-the-spectral-
measure-of-the-circular-%ce%b2/10.1214/22-AAP1798.full.

Anantharam:1991:MFL

[Ana91] V. Anantharam. A mean field limit for a lattice caricature of dynamic routing in circuit switched networks. Annals of Applied Probability, 1(4):481–503, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005834.

Anantharam:1993:USE

[Ana93] V. Anantharam. Uniqueness of stationary ergodic fixed point for a · /M/K node. Annals of Applied Probability, 3(1):154–172, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1177005512. See correction [Ana94].

Anantharam:1994:CUS

[Ana94] V. Anantharam. Correction: “Uniqueness of Stationary Ergodic Fixed Point For A·/M/K Node”. Annals of Applied Probability, 4(2):607, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005075. See [Ana93].

Andersson:1998:LTR

[And98] Håkan Andersson. Limit theorems for a random graph epidemic model. Annals of Applied Probability, 8(4):1331–1349, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903384.

Anonymous:1999:VIA

[Ano99] Anonymous. Volume index, Ann. Appl. Probab., vol. 9 (1999). Annals of Applied Probability, 9(4):iii–vii, November 1999. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1029962875.

Anonymous:2002:VIA

[Ano02] Anonymous. Volume index, Ann. Appl. Probab., vol. 12 (2002). Annals of Applied Probability, 12(4):iii–ix, November 2002. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1037125872.

Anonymous:2003:VIA

[Ano03] Anonymous. Volume index, Ann. Appl. Probab., vol. 13 (2003). Annals of Applied Probability, 13(4):iii–x, November 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1069786515.

Anonymous:2016:EB

[Ano16a] Anonymous. Editorial board. Annals of Applied Probability, 26(6):??, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1481792584.

Anonymous:2016:TC

[Ano16b] Anonymous. Table of contents. Annals of Applied Probability, 26(6):??, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1481792583.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial board. Annals of Applied Probability, 33(3):??, June 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
3/Editorial-Board/aap333p1.full.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial board. Annals of Applied Probability, 33 (4):??, August 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
4/Editorial-Board/aap334p1.full.

Anonymous:2023:TCa

[Ano23c] Anonymous. Table of contents. Annals of Applied Probability, 33(3):??, June 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
3/Table-of-Contents/aap333c1.full.

Anonymous:2023:TCb

[Ano23d] Anonymous. Table of contents. Annals of Applied Probability, 33(4):??, August 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
4/Table-of-Contents/aap334c1.full.

Antonelli:1993:BFS

[Ant93] Fabio Antonelli. Backward-forward stochastic differential equations. Annals of Applied Probability, 3(3):777–793, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005363.

Arapostathis:2016:EDC

[AP16] Ari Arapostathis and Guodong Pang. Ergodic diffusion control of multiclass multi-pool networks in the Halfin–Whitt regime. Annals of Applied Probability, 26(5):3110–3153, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1476884313.

Antunovic:2017:SEM

[AP17] Tonći Antunović and Eviatar B. Procaccia. Stationary Eden model on Cayley graphs. Annals of Applied Probability, 27 (1):517–549, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1488790834.

Avram:2007:ODP

[APP07] Florin Avram, Zbigniew Palmowski, and Martijn R. Pistorius. On the optimal dividend problem for a spectrally negative Lévy process. Annals of Applied Probability, 17(1):156–180, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1171377180.

Avram:2008:EPT

[APP08] Florin Avram, Zbigniew Palmowski, and Martijn R. Pistorius. Exit problem of a two-dimensional risk process from the quadrant: Exact and asymptotic results. Annals of Applied Probability, 18(6):2421–2449, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1227708924.

Avram:2015:GSF

[APP15] F. Avram, Z. Palmowski, and M. R. Pistorius. On Gerber– Shiu functions and optimal dividend distribution for a Lévy risk process in the presence of a penalty function. Annals of Applied Probability, 25(4):1868–1935, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1432212432.

Arapostathis:2019:ELD

[APS19] Ari Arapostathis, Guodong Pang, and Nikola Sandrić. Ergodicity of a Lévy-driven SDE arising from multiclass manyserver queues. Annals of Applied Probability, 29(2):1070–1126, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1548298937.

Ascione:2020:ETO

[APT20] Giacomo Ascione, Enrica Pirozzi, and Bruno Toaldo. On the exit time from open sets of some semi-Markov processes. Annals of Applied Probability, 30(3):1130–1163, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-3/On-the-
exit-time-from-open-sets-of-some-semi/10.1214/19-AAP1525.
full.

Angel:2008:CSD

[APW08] Omer Angel, Yuval Peres, and David B. Wilson. Card shuffling and Diophantine approximation. Annals of Applied Probability, 18(3):1215–1231, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1211819799.

Alsmeyer:1996:BGW

[AR96] Gerold Alsmeyer and Uwe Rösler. The bisexual Galton–Watson process with promiscuous mating: extinction probabilities in the supercritical case. Annals of Applied Probability, 6(3):922–939, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968234.

Alsmeyer:2002:AVP

[AR02] Gerold Alsmeyer and Uwe Rösler. Asexual versus promiscuous bisexual Galton–Watson processes: The extinction probability ratio. Annals of Applied Probability, 12(1):125–142, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015961158.

Athreya:2016:DGL

[AR16] Siva Athreya and Adrian Röllin. Dense graph limits under respondent-driven sampling. Annals of Applied Probability, 26 (4):2193–2210, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1472745456.

Aghajani:2019:ESA

[AR19a] Reza Aghajani and Kavita Ramanan. Ergodicity of an SPDE associated with a many-server queue. Annals of Applied Probability, 29(2):994–1045, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1548298935.

Aghajani:2019:HLR

[AR19b] Reza Aghajani and Kavita Ramanan. The hydrodynamic limit of a randomized load balancing network. Annals of Applied Probability, 29(4):2114–2174, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1563869039.

Anastasiou:2020:BAD

[AR20] Andreas Anastasiou and Gesine Reinert. Bounds for the asymptotic distribution of the likelihood ratio. Annals of Applied Probability, 30(2):608–643, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1591603217.

Altmeyer:2021:NEL

[AR21] Randolf Altmeyer and Markus Reiß. Nonparametric estimation for linear SPDEs from local measurements. Annals of Applied Probability, 31(1):1–38, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/

volume-31/issue-1/Nonparametric-estimation-for-linear-
SPDEs-from-local-measurements/10.1214/20-AAP1581.full.

Aberg:2008:PDW

[ARL08] Sofia Aberg, Igor Rychlik, and M. Ross Leadbetter. Palm distributions of wave characteristics in encountering seas. Annals of Applied Probability, 18(3):1059–1084, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1211819794.

Armendariz:2010:BCV

[Arm10] Inés Armendáriz. Brownian coagulation and a version of Smoluchowski’s equation on the circle. Annals of Applied Probability, 20(2):660–695, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1268143436.

Anthropelos:2017:PCC

[ARS17] Michail Anthropelos, Scott Robertson, and Konstantinos Spiliopoulos. The pricing of contingent claims and optimal positions in asymptotically complete markets. Annals of Applied Probability, 27(3):1778–1830, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1500451244.

Aghajani:2018:LSA

[ARS18] Reza Aghajani, Philippe Robert, and Wen Sun. A large scale analysis of unreliable stochastic networks. Annals of Applied Probability, 28(2):851–887, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1523433626.

Adler:1997:LCA

[AS97] Robert Adler and Gennady Samorodnitsky. Level crossings of absolutely continuous stationary symmetric α-stable processes. Annals of Applied Probability, 7(2):460–493, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1034625340.

Adler:2001:SSF

[AS01] Robert J. Adler and Georgios Skoulakis. Superprocesses over a stochastic flow. Annals of Applied Probability, 11(2):488–543, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345302.

Albin:2004:OSM

[AS04] J. M. P. Albin and Gennady Samorodnitsky. On overload in a storage model, with a self-similar and infinitely divisible input. Annals of Applied Probability, 14(2):820–844, May 2004. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1082737113.

Alexander:2006:PPI

[AS06] Kenneth S. Alexander and Vladas Sidoravicius. Pinning of polymers and interfaces by random potentials. Annals of Applied Probability, 16(2):636–669, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1151592246.

Atar:2009:CLQ

[AS09] Rami Atar and Gennady Shaikhet. Critically loaded queueing models that are throughput suboptimal. Annals of Applied Probability, 19(2):521–555, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1241702240.

Amei:2010:TDP

[AS10] Amei Amei and Stanley Sawyer. A time-dependent Poisson random field model for polymorphism within and between two related biological species. Annals of Applied Probability, 20(5): 1663–1696, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1282747397.

Anantharam:2016:DSP

[AS16a] Venkat Anantharam and Justin Salez. The densest subgraph problem in sparse random graphs. Annals of Applied Probability, 26(1):305–327, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1452003240.

Arlotto:2016:BHH

[AS16b] Alessandro Arlotto and J. Michael Steele. Beardwood–Halton– Hammersley theorem for stationary ergodic sequences: a counterexample. Annals of Applied Probability, 26(4):2141–2168, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1472745454.

Ayyer:2023:SCH

[AS23] Arvind Ayyer and Shubham Sinha. The size of t-cores and hook lengths of random cells in random partitions. Annals of Applied Probability, 33(1):85–106, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-1/The-size-of-t-cores-and-hook-lengths-
of-random/10.1214/22-AAP1809.full.

Ait-Sahalia:2009:PCJ

[ASCDH09] Yacine Aït-Sahalia, Julio Cacho-Diaz, and T. R. Hurd. Portfolio choice with jumps: a closed-form solution. Annals of Applied Probability, 19(2):556–584, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1241702241.

Adler:1993:ENL

[ASG93] Robert J. Adler, Gennady Samorodnitsky, and Tamar Gadrich. The expected number of level crossings for stationary, harmonisable, symmetric, stable processes. Annals of Applied Probability, 3(2):553–575, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005437.

Ait-Sahalia:2020:TDS

[ASJ20] Yacine Aït-Sahalia and Jean Jacod. From tick data to semimartingales. Annals of Applied Probability, 30(6):2740–2768, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-6/From-
tick-data-to-semimartingales/10.1214/20-AAP1571.full.

Asmussen:1992:LTE

[Asm92a] Soren Asmussen. Light traffic equivalence in single-server queues. Annals of Applied Probability, 2(3):555–574, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005649.

Asmussen:1992:CWC

[Asm92b] Soren Asmussen. On coupling and weak convergence to stationarity. Annals of Applied Probability, 2(3):739–751, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005657.

Asmussen:1998:SAS

[Asm98] Søren Asmussen. Subexponential asymptotics for stochastic processes: extremal behavior, stationary distributions and first passage probabilities. Annals of Applied Probability, 8(2):354–374, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903531.

Assaf:1997:ESN

[Ass97] David Assaf. Estimating the state of a noisy continuous time Markov chain when dynamic sampling is feasible. Annals of Applied Probability, 7(3):822–836, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034801256.

Asselah:1998:NSM

[Ass98] Amine Asselah. Nonreversible stationary measures for exchange processes. Annals of Applied Probability, 8(4):1303–1311, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1028903382.

Atar:2005:DMS

[Ata05a] Rami Atar. A diffusion model of scheduling control in queueing systems with many servers. Annals of Applied Probability, 15 (1B):820–852, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1107271669.

Atar:2005:SCQ

[Ata05b] Rami Atar. Scheduling control for queueing systems with many servers: Asymptotic optimality in heavy traffic. Annals of Applied Probability, 15(4):2606–2650, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1133965774.

Atar:2008:CLT

[Ata08] Rami Atar. Central limit theorem for a many-server queue with random service rates. Annals of Applied Probability, 18(4):1548– 1568, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1216677131.

Atchade:2010:CTE

[Atc10] Yves F. Atchadé. A cautionary tale on the efficiency of some adaptive Monte Carlo schemes. Annals of Applied Probability, 20(3):841–868, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1276867299.

Athreya:1994:LDR

[Ath94] K. B. Athreya. Large deviation rates for branching processes — I. Single type case. Annals of Applied Probability, 4(3):779–790, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004971.

Andrieu:2015:SSC

[ATV15] Christophe Andrieu, Vladislav B. Tadić, and Matti Vihola. On the stability of some controlled Markov chains and its applications to stochastic approximation with Markovian dynamic. Annals of Applied Probability, 25(1):1–45, February 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1418740177.

Austin:2008:EDH

[Aus08] Tim D. Austin. The emergence of the deterministic Hodgkin– Huxley equations as a limit from the underlying stochastic ionchannel mechanism. Annals of Applied Probability, 18(4):1279– 1325, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1216677123.

Athreya:1995:LDR

[AV95] K. B. Athreya and A. N. Vidyashankar. Large deviation rates for branching processes. II. The multitype case. Annals of Applied Probability, 5(2):566–576, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004778.

Andrieu:2015:CPP

[AV15] Christophe Andrieu and Matti Vihola. Convergence properties of pseudo-marginal Markov chain Monte Carlo algorithms. Annals of Applied Probability, 25(2):1030–1077, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1424355137.

Andrieu:2016:ESO

[AV16] Christophe Andrieu and Matti Vihola. Establishing some order amongst exact approximations of MCMCs. Annals of Applied Probability, 26(5):2661–2696, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1476884300.

Arratia:1994:PTS

[AW94] Richard Arratia and Michael S. Waterman. A phase transition for the score in matching random sequences allowing deletions. Annals of Applied Probability, 4(1):200–225, February 1994. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005208.

Azais:2005:DMG

[AW05] Jean-Marc Azaïs and Mario Wschebor. On the distribution of the maximum of a Gaussian field with d parameters. Annals of Applied Probability, 15(1A):254–278, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1106922328.

Alexander:2010:ECP

[AZ10] Kenneth S. Alexander and Nikos Zygouras. Equality of critical points for polymer depinning transitions with loop exponent one. Annals of Applied Probability, 20(1):356–366, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1262962326.

Alexander:2014:PPD

[AZ14a] Kenneth S. Alexander and Nikos Zygouras. Path properties of the disordered pinning model in the delocalized regime. Annals of Applied Probability, 24(2):599–615, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1394465366.

Arguin:2014:PDS

[AZ14b] Louis-Pierre Arguin and Olivier Zindy. Poisson–Dirichlet statistics for the extremes of a log-correlated Gaussian field. Annals of Applied Probability, 24(4):1446–1481, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1400073654.

Armstrong:2016:LAC

[AZ16] Scott N. Armstrong and Ofer Zeitouni. Local asymptotics for controlled martingales. Annals of Applied Probability, 26 (3):1467–1494, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1465905009.

Bally:2019:UBF

[Bal19] Vlad Bally. Upper bounds for the function solution of the homogeneous 2D Boltzmann equation with hard potential. Annals of Applied Probability, 29(3):1929–1961, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1550566847.

Ball:2021:CLT

[Bal21] Frank Ball. Central limit theorems for SIR epidemics and percolation on configuration model random graphs. Annals of Applied Probability, 31(5):2091–2142, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-5/Central-limit-theorems-for-SIR-epidemics-
and-percolation-on-configuration/10.1214/20-AAP1642.
full.

Bansaye:2008:PPD

[Ban08] Vincent Bansaye. Proliferating parasites in dividing cells: Kimmel’s branching model revisited. Annals of Applied Probability, 18(3):967–996, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1211819791.

Bansaye:2019:ASP

[Ban19] Vincent Bansaye. Approximation of stochastic processes by nonexpansive flows and coming down from infinity. Annals of Applied Probability, 29(4):2374–2438, August 2019. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1563869046.

Bartl:2019:EUM

[Bar19] Daniel Bartl. Exponential utility maximization under model uncertainty for unbounded endowments. Annals of Applied Probability, 29(1):577–612, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1544000437.

Barnes:2020:HLP

[Bar20] Clayton L. Barnes. Hydrodynamic limit and propagation of chaos for Brownian particles reflecting from a Newtonian barrier. Annals of Applied Probability, 30(4):1582–1613, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/Hydrodynamic-
limit-and-propagation-of-chaos-for-Brownian-particles-
reflecting/10.1214/19-AAP1536.full.

Battig:1999:CSM

[Bät99] Robert Bättig. Completeness of securities market models — an operator point of view. Annals of Applied Probability, 9(2):529– 566, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962754.

Bauerle:2000:AOT

[Bäu00] Nicole Bäuerle. Asymptotic optimality of tracking policies in stochastic networks. Annals of Applied Probability, 10(4): 1065–1083, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487606.

Baxendale:2005:RTC

[Bax05] Peter H. Baxendale. Renewal theory and computable convergence rates for geometrically ergodic Markov chains. Annals of Applied Probability, 15(1B):700–738, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1107271665.

Bollobas:1992:HRP

[BB92] Bela Bollobas and Graham Brightwell. The height of a random partial order: Concentration of measure. Annals of Applied Probability, 2(4):1009–1018, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005586.

Bienvenue:2003:SAR

[BB03] A. Bienvenüe and J. Bérard. Sharp asymptotic results for simplified mutation-selection algorithms. Annals of Applied Probability, 13(4):1534–1568, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1069786510.

Baccelli:2007:RST

[BB07] Francois Baccelli and Charles Bordenave. The radial spanning tree of a Poisson point process. Annals of Applied Probability, 17(1):305–359, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1171377186.

Bertoin:2015:SPL

[BB15] Jean Bertoin and Gerónimo Uribe Bravo. Supercritical percolation on large scale-free random trees. Annals of Applied Probability, 25(1):81–103, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1418740179.

Baik:2020:LRE

[BB20a] Jinho Baik and Thomas Bothner. The largest real eigenvalue in the real Ginibre ensemble and its relation to the Zakharov– Shabat system. Annals of Applied Probability, 30(1):460–501, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1582621229.

Banerjee:2020:PDD

[BB20b] Sayan Banerjee and Amarjit Budhiraja. Parameter and dimension dependence of convergence rates to stationarity for reflecting Brownian motions. Annals of Applied Probability, 30(5): 2005–2029, October 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
5/Parameter-and-dimension-dependence-of-convergence-
rates-to-stationarity-for/10.1214/19-AAP1550.full.

Bank:2020:MIF

[BB20c] Peter Bank and David Besslich. Modelling information flows by Meyer-σ-fields in the singular stochastic control problem of irreversible investment. Annals of Applied Probability, 30(6):2923– 2962, December 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
6/Modelling-information-flows-by-Meyer-sigma--fields-
in-the/10.1214/20-AAP1577.full.

Banerjee:2022:RFA

[BB22] Sayan Banerjee and Shankar Bhamidi. Root finding algorithms and persistence of Jordan centrality in growing random trees. Annals of Applied Probability, 32(3):2180–2210, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Root-
finding-algorithms-and-persistence-of-Jordan-centrality-
in-growing/10.1214/21-AAP1731.full.

Banerjee:2023:DFL

[BB23] Sayan Banerjee and Brendan Brown. Dimension-free local convergence and perturbations for reflected Brownian motions. Annals of Applied Probability, 33(1):376–416, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Dimension-
free-local-convergence-and-perturbations-for-reflected-
Brownian-motions/10.1214/22-AAP1818.full.

Backhoff:2022:EPA

[BBBW22] Julio Backhoff, Daniel Bartl, Mathias Beiglböck, and Johannes Wiesel. Estimating processes in adapted Wasserstein distance. Annals of Applied Probability, 32(1):529–550, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Estimating-
processes-in-adapted-Wasserstein-distance/10.1214/21-
AAP1687.full.

Baar:2017:SIB

[BBC17a] Martina Baar, Anton Bovier, and Nicolas Champagnat. From stochastic, individual-based models to the canonical equation of adaptive dynamics in one step. Annals of Applied Probability, 27(2):1093–1170, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1495764375.

Benaim:2017:EIM

[BBC17b] Michel Benaïm, Florian Bouguet, and Bertrand Cloez. Ergodicity of inhomogeneous Markov chains through asymptotic pseudotrajectories. Annals of Applied Probability, 27(5):3004–3049, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1509696040.

Bayraktar:2019:RCU

[BBC19] Erhan Bayraktar, Amarjit Budhiraja, and Asaf Cohen. Rate control under heavy traffic with strategic servers. Annals of Applied Probability, 29(1):1–35, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1544000424.

Banerjee:2023:FBC

[BBC23] Sayan Banerjee, Shankar Bhamidi, and Iain Carmichael. Fluctuation bounds for continuous time branching processes and evolution of growing trees with a change point. Annals of Applied Probability, 33(4):2919–2980, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-4/Fluctuation-bounds-for-continuous-
time-branching-processes-and-evolution-of/10.1214/22-
AAP1881.full.

Borovkova:1999:CTE

[BBD99] S. Borovkova, R. Burton, and H. Dehling. Consistency of the Takens estimator for the correlation dimension. Annals of Applied Probability, 9(2):376–390, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962747.

Bhamidi:2022:NEF

[BBD22] Shankar Bhamidi, Amarjit Budhiraja, and Miheer Dewaskar. Near equilibrium fluctuations for supermarket models with growing choices. Annals of Applied Probability, 32(3):2083–2138, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Near-
equilibrium-fluctuations-for-supermarket-models-with-
growing-choices/10.1214/21-AAP1729.full.

Bhamidi:2022:REA

[BBDW22] Shankar Bhamidi, Amarjit Budhiraja, Paul Dupuis, and Ruoyu Wu. Rare event asymptotics for exploration processes for random graphs. Annals of Applied Probability, 32(2):1112–1178, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Rare-
event-asymptotics-for-exploration-processes-for-random-
graphs/10.1214/21-AAP1704.full.

Booth:2003:CAC

[BBFM03] Lorna Booth, Jehoshua Bruck, Massimo Franceschetti, and Ronald Meester. Covering algorithms, continuum percolation and the geometry of wireless networks. Annals of Applied Probability, 13(2):722–741, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1050689601.

Balogh:2011:HCR

[BBK^{+}11] József Balogh, Béla Bollobás, Michael Krivelevich, Tobias Müller, and Mark Walters. Hamilton cycles in random geometric graphs. Annals of Applied Probability, 21(3):1053–1072, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1307020391.

Bean:1997:QSB

[BBL^{+}97] N. G. Bean, L. Bright, G. Latouche, C. E. M. Pearce, P. K. Pollett, and P. G. Taylor. The quasi-stationary behavior of quasi-birth-and-death processes. Annals of Applied Probability, 7(1):134–155, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034625256.

Berestycki:2014:STC

[BBL14] Julien Berestycki, Nathanaël Berestycki, and Vlada Limic. A small-time coupling between Λ-coalescents and branching processes. Annals of Applied Probability, 24(2):449–475, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1394465362.

Bally:2007:IPF

[BBM07] Vlad Bally, Marie-Pierre Bavouzet, and Marouen Messaoud. Integration by parts formula for locally smooth laws and applications to sensitivity computations. Annals of Applied Probability, 17(1):33–66, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1171377176.

Borodin:2015:LSG

[BBO15] Alexei Borodin, Alexey Bufetov, and Grigori Olshanski. Limit shapes for growing extreme characters of U(∞). Annals of Applied Probability, 25(4):2339–2381, 08 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Banerjee:2022:HTS

[BBP22] Sayan Banerjee, Amarjit Budhiraja, and Amber L. Puha. Heavy traffic scaling limits for shortest remaining processing time queues with heavy tailed processing time distributions. Annals of Applied Probability, 32(4):2587–2651, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Heavy-
traffic-scaling-limits-for-shortest-remaining-processing-
time-queues/10.1214/21-AAP1741.full.

Billey:2015:MEW

[BBPS15] Sara Billey, Krzysztof Burdzy, Soumik Pal, and Bruce E. Sagan. On meteors, earthworms and WIMPs. Annals of Applied Probability, 25(4):1729–1779, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1432212429.

Bazhba:2020:SPL

[BBRZ20] Mihail Bazhba, Jose Blanchet, Chang-Han Rhee, and Bert Zwart. Sample path large deviations for Lévy processes and random walks with Weibull increments. Annals of Applied Probability, 30(6):2695–2739, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-6/Sample-path-large-deviations-for-L%
c3%a9vy-processes-and-random-walks/10.1214/20-AAP1570.
full.

Bhamidi:2011:MTE

[BBS11] Shankar Bhamidi, Guy Bresler, and Allan Sly. Mixing time of exponential random graphs. Annals of Applied Probability, 21 (6):2146–2170, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1322057318.

Ball:2017:EDP

[BBT17] Frank Ball, Tom Britton, and Pieter Trapman. An epidemic in a dynamic population with importation of infectives. Annals of Applied Probability, 27(1):242–274, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1488790828.

Balister:2004:CPS

[BBW04] Paul Balister, Béla Bollobás, and Mark Walters. Continuum percolation with steps in an annulus. Annals of Applied Probability, 14(4):1869–1879, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674081.

Barbour:2001:CPA

[BC01] A. D. Barbour and O. Chryssaphinou. Compound Poisson approximation: a user’s guide. Annals of Applied Probability, 11 (3):964–1002, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015345355.

Baldi:2002:AHP

[BC02] Paolo Baldi and Lucia Caramellino. Asymptotics of hitting probabilities for general one-dimensional pinned diffusions. Annals of Applied Probability, 12(3):1071–1095, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1031863181.

BenArous:2005:BME

[BČ05] Gérard Ben Arous and Jiří Černý. Bouchaud’s model exhibits two different aging regimes in dimension one. Annals of Applied Probability, 15(2):1161–1192, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1115137972.

Buchmann:2009:IFN

[BC09] Boris Buchmann and Ngai Hang Chan. Integrated functionals of normal and fractional processes. Annals of Applied Probability, 19(1):49–70, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1235140332.

Borodin:2014:MLE

[BC14a] Alexei Borodin and Ivan Corwin. Moments and Lyapunov exponents for the parabolic Anderson model. Annals of Applied Probability, 24(3):1172–1198, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1398258098.

Broniatowski:2014:LRU

[BC14b] Michel Broniatowski and Virgile Caron. Long runs under a conditional limit distribution. Annals of Applied Probability, 24(6): 2246–2296, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1409058032.

Blanchet:2015:SSS

[BC15] Jose Blanchet and Xinyun Chen. Steady-state simulation of reflected Brownian motion and related stochastic networks. Annals of Applied Probability, 25(6):3209–3250, December 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1443703773.

Barraquand:2016:HAE

[BC16] Guillaume Barraquand and Ivan Corwin. The q-Hahn asymmetric exclusion process. Annals of Applied Probability, 26(4):2304– 2356, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1472745460.

Blanchard:2018:MPO

[BC18] Romain Blanchard and Laurence Carassus. Multiple-priors optimal investment in discrete time for unbounded utility function. Annals of Applied Probability, 28(3):1856–1892, June 2018. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1527840034.

Bakhtin:2021:LET

[BC21] Yuri Bakhtin and Hong-Bin Chen. Long exit times near a repelling equilibrium. Annals of Applied Probability, 31(2): 594–624, April 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
2/Long-exit-times-near-a-repelling-equilibrium/10.1214/
20-AAP1599.full.

Borga:2022:QLL

[BC22a] Jacopo Borga and Benedetta Cavalli. Quenched law of large numbers and quenched central limit theorem for multiplayer leagues with ergodic strengths. Annals of Applied Probability, 32(6):4398–4425, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-6/Quenched-law-of-large-numbers-and-quenched-central-
limit-theorem/10.1214/22-AAP1790.full.

Bukh:2022:PWC

[BC22b] Boris Bukh and Christopher Cox. Periodic words, common subsequences and frogs. Annals of Applied Probability, 32(2): 1295–1332, April 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
2/Periodic-words-common-subsequences-and-frogs/10.1214/
21-AAP1709.full.

Bernardin:2023:SCL

[BCC^{+}23] Cédric Bernardin, Raphaël Chetrite, Reda Chhaibi, Joseph Najnudel, and Clément Pellegrini. Spiking and collapsing in large noise limits of SDEs. Annals of Applied Probability, 33(1):417– 446, February 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-

1/Spiking-and-collapsing-in-large-noise-limits-of-SDEs/
10.1214/22-AAP1819.full.

Blanchet:2017:SSM

[BCD17] Jose Blanchet, Xinyun Chen, and Jing Dong. ϵ-strong simulation for multidimensional stochastic differential equations via rough path analysis. Annals of Applied Probability, 27(1):275– 336, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1488790829.

Brummitt:2015:JPW

[BCDS15] Charles D. Brummitt, Shirshendu Chatterjee, Partha S. Dey, and David Sivakoff. Jigsaw percolation: What social networks can collaboratively solve a puzzle? Annals of Applied Probability, 25(4):2013–2038, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1432212435.

Bolley:2018:DPC

[BCF18] François Bolley, Djalil Chafaï, and Joaquín Fontbona. Dynamics of a planar Coulomb gas. Annals of Applied Probability, 28 (5):3152–3183, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1535443245.

Bandini:2018:BSO

[BCFP18] Elena Bandini, Andrea Cosso, Marco Fuhrman, and Huyên Pham. Backward SDEs for optimal control of partially observed path-dependent stochastic systems: a control randomization approach. Annals of Applied Probability, 28(3):1634–1678, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1527840029.

Baake:2022:LDD

[BCH22] Ellen Baake, Fernando Cordero, and Sebastian Hummel. Lines of descent in the deterministic mutation-selection model with pairwise interaction. Annals of Applied Probability, 32(4):2400–2447, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Lines-
of-descent-in-the-deterministic-mutationselection-model-
with-pairwise/10.1214/21-AAP1736.full.

Bai:1998:VNM

[BCHL98] Zhi-Dong Bai, Chern-Ching Chao, Hsien-Kuei Hwang, and Wen-Qi Liang. On the variance of the number of maxima in random vectors and its applications. Annals of Applied Probability, 8(3):886–895, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903455.

Beskos:2014:SSM

[BCJ14] Alexandros Beskos, Dan Crisan, and Ajay Jasra. On the stability of sequential Monte Carlo methods in high dimensions. Annals of Applied Probability, 24(4):1396–1445, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1400073653.

Beffara:2022:LGR

[BCJ22] Vincent Beffara, Sunil Chhita, and Kurt Johansson. Local geometry of the rough-smooth interface in the two-periodic Aztec diamond. Annals of Applied Probability, 32(2):974–1017, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Local-
geometry-of-the-rough-smooth-interface-in-the-two/10.
1214/21-AAP1701.full.

Baggerly:1999:AIS

[BCKP99] Keith Baggerly, Dennis Cox, Craig Kollman, and Rick Picard. Adaptive importance sampling on discrete Markov chains. Annals of Applied Probability, 9(2):391–412, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1029962748.

Blath:2016:NCS

[BCKWB16] Jochen Blath, Adrián González Casanova, Noemi Kurt, and Maite Wilke-Berenguer. A new coalescent for seed-bank models. Annals of Applied Probability, 26(2):857–891, April 2016. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1458651822.

Brockwell:2006:CTG

[BCL06] Peter Brockwell, Erdenebaatar Chadraa, and Alexander Lindner. Continuous-time GARCH processes. Annals of Applied Probability, 16(2):790–826, May 2006. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1151592251.

Bate:2021:COS

[BCMR21] Michael E. Bate, Stephen B. Connor, and Oliver Matheau-Raven. Cutoff for a one-sided transposition shuffle. Annals of Applied Probability, 31(4):1746–1773, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/journals/annals-of-applied-
probability/volume-31/issue-4/Cutoff-for-a-one-sided-
transposition-shuffle/10.1214/20-AAP1632.full.

Bapst:2016:PTP

[BCOR16] Victor Bapst, Amin Coja-Oghlan, and Felicia Raßmann. A positive temperature phase transition in random hypergraph 2coloring. Annals of Applied Probability, 26(3):1362–1406, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1465905006.

Burdzy:2011:APB

[BCP11] Krzysztof Burdzy, Zhen-Qing Chen, and Soumik Pal. Archimedes’ principle for Brownian liquid. Annals of Applied Probability, 21 (6):2053–2074, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1322057315.

Benaim:2018:SAQ

[BCP18] Michel Benaim, Bertrand Cloez, and Fabien Panloup. Stochastic approximation of quasi-stationary distributions on compact spaces and applications. Annals of Applied Probability, 28(4): 2370–2416, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1533780276.

Blanca:2022:EDS

[BCP^{+}22] Antonio Blanca, Pietro Caputo, Daniel Parisi, Alistair Sinclair, and Eric Vigoda. Entropy decay in the Swendsen–Wang dynamics on Z^{d}. Annals of Applied Probability, 32(2):1018–1057, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Entropy-

decay-in-the-SwendsenWang-dynamics-on-Zd/10.1214/21-
AAP1702.full.

Bakhtin:2022:EMC

[BCPG22] Yuri Bakhtin, Hong-Bin Chen, and Zsolt Pajor-Gyulai. Erratum: Malliavin calculus approach to long exit times from an unstable equilibrium. Annals of Applied Probability, 32(4):3199–3200, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Erratum-
-Malliavin-calculus-approach-to-long-exit-times-from/
10.1214/21-AAP1751.full. See [BPG19].

Berti:2015:CLT

[BCPR15] Patrizia Berti, Irene Crimaldi, Luca Pratelli, and Pietro Rigo. Central limit theorems for an Indian buffet model with random weights. Annals of Applied Probability, 25(2):523–547, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1424355122.

Bovier:2019:CFV

[BCS19] Anton Bovier, Loren Coquille, and Charline Smadi. Crossing a fitness valley as a metastable transition in a stochastic population model. Annals of Applied Probability, 29(6):3541–3589, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1578366321.

Bhatnagar:2007:ATS

[BCTV07] Nayantara Bhatnagar, Pietro Caputo, Prasad Tetali, and Eric Vigoda. Analysis of top-swap shuffling for genome rearrangements. Annals of Applied Probability, 17(4):1424–1445, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1186755245.

Bayraktar:2019:LTG

[BCZ19] Erhan Bayraktar, Jaksa Cvitanić, and Yuchong Zhang. Large tournament games. Annals of Applied Probability, 29(6):3695– 3744, December 2019. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
euclid.aoap/1578366324.

Bayer:1992:TDS

[BD92] Dave Bayer and Persi Diaconis. Trailing the dovetail shuffle to its lair. Annals of Applied Probability, 2(2):294–313, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005705.

Breidt:1998:ESV

[BD98] F. Jay Breidt and Richard A. Davis. Extremes of stochastic volatility models. Annals of Applied Probability, 8(3):664–675, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903446.

Bramson:2001:HTL

[BD01] Maury Bramson and J. G. Dai. Heavy traffic limits for some queueing networks. Annals of Applied Probability, 11(1):49–90, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
998926987.

Birkner:2007:SCC

[BD07] Matthias Birkner and Andrej Depperschmidt. Survival and complete convergence for a spatial branching system with local regulation. Annals of Applied Probability, 17(5/6):1777–1807, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1197909503.

Bierme:2012:CSS

[BD12] Hermine Biermé and Agnès Desolneux. Crossings of smooth shot noise processes. Annals of Applied Probability, 22(6):2240–2281, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1353695953.

Bruss:2015:RDB

[BD15] F. Thomas Bruss and Mitia Duerinckx. Resource dependent branching processes and the envelope of societies. Annals of Applied Probability, 25(1):324–372, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1418740188.

Braverman:2017:SMS

[BD17] Anton Braverman and J. G. Dai. Stein’s method for steadystate diffusion approximations of M/Ph/n + M systems. Annals of Applied Probability, 27(1):550–581, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1488790834.

Bank:2019:CTD

[BD19a] Peter Bank and Yan Dolinsky. Continuous-time duality for superreplication with transient price impact. Annals of Applied Probability, 29(6):3893–3917, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1578366329.

Bank:2019:SRF

[BD19b] Peter Bank and Yan Dolinsky. Super-replication with fixed transaction costs. Annals of Applied Probability, 29(2):739–757, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1548298929.

Bierme:2020:MGR

[BD20] Hermine Biermé and Agnès Desolneux. Mean geometry for 2D random fields: Level perimeter and level total curvature integrals. Annals of Applied Probability, 30(2):561–607, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1591603216.

Bank:2016:SRN

[BDG16] Peter Bank, Yan Dolinsky, and Selim Gökay. Super-replication with nonlinear transaction costs and volatility uncertainty. Annals of Applied Probability, 26(3):1698–1726, June 2016. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1465905016.

Biskup:1999:HNL

[BdH99] Marek Biskup and Frank den Hollander. A heteropolymer near a linear interface. Annals of Applied Probability, 9(3):668–687, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962808.

Bramson:2010:PRR

[BDH10] Maury Bramson, J. G. Dai, and J. M. Harrison. Positive recurrence of reflecting Brownian motion in three dimensions. Annals of Applied Probability, 20(2):753–783, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1268143439.

Braunsteins:2023:SPL

[BdHM23] Peter Braunsteins, Frank den Hollander, and Michel Mandjes. A sample-path large deviation principle for dynamic Erdős–Rényi random graphs. Annals of Applied Probability, 33(4):3278–3320, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-4/A-sample-
path-large-deviation-principle-for-dynamic-Erd%c5%91sR%
c3%a9nyi-random/10.1214/22-AAP1892.full.

Borst:2022:CTB

[BdHNT22] Sem Borst, Frank den Hollander, Francesca Romana Nardi, and Siamak Taati. Crossover times in bipartite networks with activity constraints and time-varying switching rates. Annals of Applied Probability, 32(6):4279–4314, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/Crossover-
times-in-bipartite-networks-with-activity-constraints-
and-time/10.1214/22-AAP1787.full.

Bayraktar:2006:APD

[BDK06] Erhan Bayraktar, Savas Dayanik, and Ioannis Karatzas. Adaptive Poisson disorder problem. Annals of Applied Probability, 16(3):1190–1261, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1159804980.

Bishop:2019:ODR

[BDKR19] A. N. Bishop, P. Del Moral, K. Kamatani, and B. Rémillard. On one-dimensional Riccati diffusions. Annals of Applied Probability, 29(2):1127–1187, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1548298938.

Bezborodov:2020:SGP

[BDKT20] Viktor Bezborodov, Luca Di Persio, Tyll Krueger, and Pasha Tkachov. Spatial growth processes with long range dispersion: Microscopics, mesoscopics and discrepancy in spread rate. Annals of Applied Probability, 30(3):1091–1129, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-3/Spatial-
growth-processes-with-long-range-dispersion--Microscopics-
mesoscopics/10.1214/19-AAP1524.full.

Broutin:2016:AOS

[BDL16] Nicolas Broutin, Luc Devroye, and Gábor Lugosi. Almost optimal sparsification of random geometric graphs. Annals of Applied Probability, 26(5):3078–3109, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1476884312.

Basrak:2002:CMR

[BDM02] Bojan Basrak, Richard A. Davis, and Thomas Mikosch. A characterization of multivariate regular variation. Annals of Applied Probability, 12(3):908–920, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1031863174.

Bhattacharya:2017:ULT

[BDM17] Bhaswar B. Bhattacharya, Persi Diaconis, and Sumit Mukherjee. Universal limit theorems in graph coloring problems with connections to extremal combinatorics. Annals of Applied Probability, 27(1):337–394, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1488790830.

Bansaye:2011:LTM

[BDMT11] Vincent Bansaye, Jean-François Delmas, Laurence Marsalle, and Viet Chi Tran. Limit theorems for Markov processes indexed by continuous time Galton–Watson trees. Annals of Applied Probability, 21(6):2263–2314, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1322057322.

Briand:2020:PSN

[BdRGL20] Philippe Briand, Paul-Éric Chaudru de Raynal, Arnaud Guillin, and Céline Labart. Particles systems and numerical schemes for mean reflected stochastic differential equations. Annals of Applied Probability, 30(4):1884–1909, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/Particles-
systems-and-numerical-schemes-for-mean-reflected-stochastic-
differential/10.1214/19-AAP1546.full.

Buijsrogge:2020:SAR

[BDS20] Anne Buijsrogge, Paul Dupuis, and Michael Snarski. Splitting algorithms for rare event simulation over long time intervals. Annals of Applied Probability, 30(6):2963–2998, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-6/Splitting-
algorithms-for-rare-event-simulation-over-long-time-
intervals/10.1214/20-AAP1578.full.

Bierme:2017:IPO

[BDW17] Hermine Biermé, Olivier Durieu, and Yizao Wang. Invariance principles for operator-scaling Gaussian random fields. Annals of Applied Probability, 27(2):1190–1234, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1495764377.

Broutin:2022:LMI

[BDW22] Nicolas Broutin, Thomas Duquesne, and Minmin Wang. Limits of multiplicative inhomogeneous random graphs and Lévy trees: The continuum graphs. Annals of Applied Probability, 32(4): 2448–2503, August 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
4/Limits-of-multiplicative-inhomogeneous-random-graphs-
and-L%c3%a9vy-trees/10.1214/21-AAP1737.full.

Bar:2023:IPI

[BDW23] Michael Bär, Jetlir Duraj, and Vitali Wachtel. Invariance principles for integrated random walks conditioned to stay positive. Annals of Applied Probability, 33(1):127–160, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Invariance-
principles-for-integrated-random-walks-conditioned-to-
stay-positive/10.1214/22-AAP1811.full.

Breden:2023:CAP

[BE23] Maxime Breden and Maximilian Engel. Computer-assisted proof of shear-induced chaos in stochastically perturbed Hopf systems. Annals of Applied Probability, 33(2):1052–1094, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Computer-
assisted-proof-of-shear-induced-chaos-in-stochastically-
perturbed/10.1214/22-AAP1841.full.

Becherer:2006:BSB

[Bec06] Dirk Becherer. Bounded solutions to backward SDEs with jumps for utility optimization and indifference hedging. Annals of Applied Probability, 16(4):2027–2054, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1169065215.

Bedard:2007:WCM

[Béd07] Mylène Bédard. Weak convergence of Metropolis algorithms for non-i.i.d. target distributions. Annals of Applied Probability, 17 (4):1222–1244, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1186755238.

Barbour:2000:TFE

[BEG00] A. D. Barbour, S. N. Ethier, and R. C. Griffiths. A transition function expansion for a diffusion model with selection. Annals of Applied Probability, 10(1):123–162, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1019737667.

Briand:2018:BMR

[BEH18] Philippe Briand, Romuald Elie, and Ying Hu. BSDEs with mean reflection. Annals of Applied Probability, 28(1):482–510, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1520046093.

Belomestny:2011:RCS

[Bel11] Denis Belomestny. On the rates of convergence of simulationbased optimization algorithms for optimal stopping problems. Annals of Applied Probability, 21(1):215–239, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1292598032.

Belius:2012:CLR

[Bel12] David Belius. Cover levels and random interlacements. Annals of Applied Probability, 22(2):522–540, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1333372006.

Belomestny:2013:SOS

[Bel13] Denis Belomestny. Solving optimal stopping problems via empirical dual optimization. Annals of Applied Probability, 23 (5):1988–2019, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1377696304.

Blath:2007:CLR

[BEM07] Jochen Blath, Alison Etheridge, and Mark Meredith. Coexistence in locally regulated competing populations and survival of branching annihilating random walk. Annals of Applied Probability, 17(5/6):1474–1507, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1191419173.

Braverman:2008:MTS

[BEM08] Mark Braverman, Omid Etesami, and Elchanan Mossel. Mafia: a theoretical study of players and coalitions in a partial information environment. Annals of Applied Probability, 18(3):825–846, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1211819786.

Broman:2021:FCP

[BEMT21] Erik I. Broman, Olof Elias, Filipe Mussini, and Johan Tykesson. The fractal cylinder process: Existence and connectivity phase transitions. Annals of Applied Probability, 31(5):2192–2243, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-5/The-
fractal-cylinder-process-Existence-and-connectivity-
phase-transitions/10.1214/20-AAP1644.full.

Benois:1996:LDO

[Ben96] O. Benois. Large deviations for the occupation times of independent particle systems. Annals of Applied Probability, 6 (1):269–296, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968074.

Bender:2012:SA

[Ben12] Christian Bender. Simple arbitrage. Annals of Applied Probability, 22(5):2067–2085, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1350067994.

Boland:1991:SOI

[BENP91] Philip J. Boland, Emad El-Neweihi, and Frank Proschan. Stochastic order for inspection and repair policies. Annals of Applied Probability, 1(2):207–218, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005934.

Berman:1992:TCD

[Ber92] Simeon M. Berman. The tail of the convolution of densities and its application to a model of HIV-latency time. Annals of Applied Probability, 2(2):481–502, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005712.

Berman:1994:PNR

[Ber94] Simeon M. Berman. Perturbation of normal random vectors by nonnormal translations, and an application to HIV latency time distributions. Annals of Applied Probability, 4(4):968–980, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004899.

Berman:1996:RIS

[Ber96] Simeon M. Berman. Rank inversions in scoring multipart examinations. Annals of Applied Probability, 6(3):992–1005, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968237.

Berger:1997:ASI

[Ber97a] Erich Berger. An almost sure invariance principle for stochastic approximation procedures in linear filtering theory. Annals of Applied Probability, 7(2):444–459, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034625339.

Bertoin:1997:EDE

[Ber97b] Jean Bertoin. Exponential decay and ergodicity of completely asymmetric Lévy processes in a finite interval. Annals of Applied Probability, 7(1):156–169, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034625257.

Bertoin:2002:ESS

[Ber02] Jean Bertoin. Eternal solutions to Smoluchowski’s coagulation equation with additive kernel and their probabilistic interpretations. Annals of Applied Probability, 12(2):547–564, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1026915615.

Bertoin:2010:ARP

[Ber10] Jean Bertoin. Asymptotic regimes for the partition into colonies of a branching process with emigration. Annals of Applied Probability, 20(6):1967–1988, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1287494551.

Barton:2004:CRB

[BES04] N. H. Barton, A. M. Etheridge, and A. K. Sturm. Coalescence in a random background. Annals of Applied Probability, 14(2):754– 785, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1082737110.

Bouchard:2005:MRS

[BET05] Bruno Bouchard, Nicole El Karoui, and Nizar Touzi. Maturity randomization for stochastic control problems. Annals of Applied Probability, 15(4):2575–2605, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1133965773.

Bayraktar:2020:AOC

[BEZ20] Erhan Bayraktar, Ibrahim Ekren, and Yili Zhang. On the asymptotic optimality of the comb strategy for prediction with expert advice. Annals of Applied Probability, 30(6):2517–2546, December 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
6/On-the-asymptotic-optimality-of-the-comb-strategy-
for-prediction/10.1214/20-AAP1565.full.

Burton:1996:SOC

[BF96] Robert M. Burton and William G. Faris. A self-organizing cluster process. Annals of Applied Probability, 6(4):1232–1247, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1035463330.

Bruss:2002:HRC

[BF02] F. Thomas Bruss and Thomas S. Ferguson. High-risk and competitive investment models. Annals of Applied Probability, 12 (4):1202–1226, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1037125860.

Baccelli:2004:MTM

[BF04a] François Baccelli and Serguei Foss. Moments and tails in monotone-separable stochastic networks. Annals of Applied Probability, 14(2):612–650, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1082737105.

Biagini:2004:SRP

[BF04b] Sara Biagini and Marco Frittelli. On the super replication price of unbounded claims. Annals of Applied Probability, 14 (4):1970–1991, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674085.

Bovier:2005:SCA

[BF05] Anton Bovier and Alessandra Faggionato. Spectral characterization of aging: The REM-like trap model. Annals of Applied Probability, 15(3):1997–2037, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1121433775.

Biagini:2008:UFU

[BF08] Sara Biagini and Marco Frittelli. A unified framework for utility maximization problems: An Orlicz space approach. Annals of Applied Probability, 18(3):929–966, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1211819790.

Barrera:2010:LMF

[BF10] Javiera Barrera and Joaquín Fontbona. The limiting move-to-front search-cost in law of large numbers asymptotic regimes. Annals of Applied Probability, 20(2):722–752, April 2010. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1268143438.

Benaim:2012:SAC

[BF12] Michel Benaïm and Mathieu Faure. Stochastic approximation, cooperative dynamics and supermodular games. Annals of Applied Probability, 22(5):2133–2164, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1350067997.

Budhiraja:2018:DAC

[BF18] Amarjit Budhiraja and Eric Friedlander. Diffusion approximations for controlled weakly interacting large finite state systems with simultaneous jumps. Annals of Applied Probability, 28 (1):204–249, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1520046087.

Bernou:2022:CAC

[BF22a] Armand Bernou and Nicolas Fournier. A coupling approach for the convergence to equilibrium for a collisionless gas. Annals of Applied Probability, 32(2):764–811, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/A-coupling-
approach-for-the-convergence-to-equilibrium-for-a/10.
1214/21-AAP1696.full.

Bufetov:2022:SFT

[BF22b] Alexey Bufetov and Patrik L. Ferrari. Shock fluctuations in TASEP under a variety of time scalings. Annals of Applied Probability, 32(5):3614–3644, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Shock-fluctuations-in-TASEP-under-
a-variety-of-time-scalings/10.1214/21-AAP1769.full.

Bertini:2013:LDP

[BFG13] Lorenzo Bertini, Alessandra Faggionato, and Davide Gabrielli. Large deviation principles for nongradient weakly asymmetric stochastic lattice gases. Annals of Applied Probability, 23(1):1– 65, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1359124381.

Breit:2020:MSS

[BFH20] Dominic Breit, Eduard Feireisl, and Martina Hofmanová. Markov selection for the stochastic compressible Navier–Stokes system. Annals of Applied Probability, 30(6):2547–2572, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-6/Markov-
selection-for-the-stochastic-compressible-NavierStokes-
system/10.1214/20-AAP1566.full.

Benjamini:2020:EMF

[BFHM20] Itai Benjamini, Luiz Renato Fontes, Jonathan Hermon, and Fábio Prates Machado. On an epidemic model on finite graphs. Annals of Applied Probability, 30(1):208–258, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1582621223.

Blum:2006:MVL

[BFJ06] Michael G. B. Blum, Olivier François, and Svante Janson. The mean, variance and limiting distribution of two statistics sensitive to phylogenetic tree balance. Annals of Applied Probability, 16(4):2195–2214, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1169065221.

Banner:2005:AME

[BFK05] Adrian D. Banner, Robert Fernholz, and Ioannis Karatzas. Atlas models of equity markets. Annals of Applied Probability, 15(4): 2296–2330, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965764.

Barbato:2011:ADS

[BFM11] David Barbato, Franco Flandoli, and Francesco Morandin. Anomalous dissipation in a stochastic inviscid dyadic model. Annals of Applied Probability, 21(6):2424–2446, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1322057326.

Burzoni:2017:MFS

[BFM17] Matteo Burzoni, Marco Frittelli, and Marco Maggis. Model-free superhedging duality. Annals of Applied Probability, 27(3):1452– 1477, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1500451228.

Bella:2018:LTE

[BFO18] Peter Bella, Benjamin Fehrman, and Felix Otto. A Liouville theorem for elliptic systems with degenerate ergodic coefficients. Annals of Applied Probability, 28(3):1379–1422, June 2018. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1527840022.

Baldi:1993:ISG

[BFP93] Paolo Baldi, Arnoldo Frigessi, and Mauro Piccioni. Importance sampling for Gibbs random fields. Annals of Applied Probability, 3(3):914–933, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005372.

Battiston:2018:CPF

[BFRT18] Marco Battiston, Stefano Favaro, Daniel M. Roy, and Yee Whye Teh. A characterization of product-form exchangeable feature probability functions. Annals of Applied Probability, 28(3):1423– 1448, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1527840023.

Bernardin:2021:DCK

[BFS21] C. Bernardin, T. Funaki, and S. Sethuraman. Derivation of coupled KPZ-Burgers equation from multi-species zero-range processes. Annals of Applied Probability, 31(4):1966–2017, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/Derivation-
of-coupled-KPZ-Burgers-equation-from-multi-species-zero/
10.1214/20-AAP1639.full.

Budhiraja:2021:MSA

[BFW21] Amarjit Budhiraja, Eric Friedlander, and Ruoyu Wu. Manyserver asymptotics for join-the-shortest-queue: Large deviations and rare events. Annals of Applied Probability, 31(5):2376–2419, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-5/Many-
server-asymptotics-for-join-the-shortest-queue--Large/
10.1214/20-AAP1650.full.

Barbour:1993:RPA

[BG93] A. D. Barbour and P. E. Greenwood. Rates of Poisson approximation to finite range random fields. Annals of Applied Probability, 3(1):91–102, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005509.

Barsky:1995:GMP

[BG95] David J. Barsky and Alberto Gandolfi. A generalized maximum pseudo-likelihood estimator for noisy Markov fields. Annals of Applied Probability, 5(4):1095–1125, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177004608.

Brown:1996:PAP

[BG96] Timothy C. Brown and Darryl Greig. Poisson approximation for point processes via monotone couplings. Annals of Applied Probability, 6(2):545–560, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968143.

Baryshnikov:2000:SSI

[BG00] Yuliy M. Baryshnikov and Alexander V. Gnedin. Sequential selection of an increasing sequence from a multidimensional random sample. Annals of Applied Probability, 10(1):258–267, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1019737672.

Baryshnikov:2001:CIP

[BG01] Yuliy Baryshnikov and Alexander Gnedin. Counting intervals in the packing process. Annals of Applied Probability, 11(3):863– 877, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1015345351.

Berglund:2002:SPA

[BG02] Nils Berglund and Barbara Gentz. A sample-paths approach to noise-induced synchronization: Stochastic resonance in a doublewell potential. Annals of Applied Probability, 12(4):1419–1470, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1037125869.

Bruss:2003:MMD

[BG03] F. Thomas Bruss and Rudolf Grübel. On the multiplicity of the maximum in a discrete random sample. Annals of Applied Probability, 13(4):1252–1263, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1069786498.

Bolthausen:2005:PCS

[BG05a] Erwin Bolthausen and Giambattista Giacomin. Periodic copolymers at selective interfaces: a large deviations approach. Annals of Applied Probability, 15(1B):963–983, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1107271674.

Budhiraja:2005:LDA

[BG05b] Amarjit Budhiraja and Arka Prasanna Ghosh. A large deviations approach to asymptotically optimal control of crisscross network in heavy traffic. Annals of Applied Probability, 15 (3):1887–1935, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433772.

Blanchet:2006:CCD

[BG06a] Jose Blanchet and Peter Glynn. Complete corrected diffusion approximations for the maximum of a random walk. Annals of Applied Probability, 16(2):951–983, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1151592256.

Budhiraja:2006:DAC

[BG06b] Amarjit Budhiraja and Arka Prasanna Ghosh. Diffusion approximations for controlled stochastic networks: An asymptotic bound for the value function. Annals of Applied Probability, 16 (4):1962–2006, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1169065213.

Blanchet:2008:ERE

[BG08] Jose Blanchet and Peter Glynn. Efficient rare-event simulation for the maximum of heavy-tailed random walks. Annals of Applied Probability, 18(4):1351–1378, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1216677125.

Budhiraja:2012:CSN

[BG12] Amarjit Budhiraja and Arka P. Ghosh. Controlled stochastic networks in heavy traffic: Convergence of value functions. Annals of Applied Probability, 22(2):734–791, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1333372012.

Burdzy:2019:KGF

[BG19] Krzysztof Burdzy and Carl-Erik Gauthier. Knudsen gas in flat tire. Annals of Applied Probability, 29(1):217–263, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1544000429.

Bramson:2021:MVP

[BG21] Maury Bramson and Lawrence F. Gray. The majority vote process and other consensus processes on trees. Annals of Applied Probability, 31(1):169–198, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-1/The-majority-vote-process-and-other-
consensus-processes-on-trees/10.1214/20-AAP1586.full.

Benaych-Georges:2023:OCS

[BGBP23] Florent Benaych-Georges, Jean-Philippe Bouchaud, and Marc Potters. Optimal cleaning for singular values of cross-covariance matrices. Annals of Applied Probability, 33(2):1095–1126, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Optimal-
cleaning-for-singular-values-of-cross-covariance-matrices/
10.1214/22-AAP1842.full.

Brehier:2016:USG

[BGG^{+}16] Charles-Edouard Bréhier, Maxime Gazeau, Ludovic Goudenège, Tony Lelièvre, and Mathias Rousset. Unbiasedness of some generalized adaptive multilevel splitting algorithms. Annals of Applied Probability, 26(6):3559–3601, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1481792593.

Borggaard:2020:BCF

[BGHK20] Jeff Borggaard, Nathan Glatt-Holtz, and Justin Krometis. On Bayesian consistency for flows observed through a passive scalar. Annals of Applied Probability, 30(4):1762–1783, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/On-Bayesian-
consistency-for-flows-observed-through-a-passive-scalar/
10.1214/19-AAP1542.full.

Bacry:2010:MAM

[BGHM10] Emmanuel Bacry, Arnaud Gloter, Marc Hoffmann, and Jean François Muzy. Multifractal analysis in a mixed asymptotic framework. Annals of Applied Probability, 20(5):1729–1760, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1282747399.

Bernardin:2018:WHO

[BGJ18] Cédric Bernardin, Patrícia Gonçalves, and Milton Jara. Weakly harmonic oscillators perturbed by a conservative noise. Annals of Applied Probability, 28(3):1315–1355, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1527840020.

Boucheron:2002:BBL

[BGL02] Stéphane Boucheron, Fabrice Gamboa, and Christian Léonard. Bins and balls; large deviations of the empirical occupancy process. Annals of Applied Probability, 12(2):607–636, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1026915618.

Bally:2018:RSS

[BGR18] Vlad Bally, Dan Goreac, and Victor Rabiet. Regularity and stability for the semigroup of jump diffusions with state-dependent intensity. Annals of Applied Probability, 28(5):3028–3074, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1535443241.

Bufetov:2022:ATA

[BGR22] Alexey Bufetov, Vadim Gorin, and Dan Romik. Absorbing time asymptotics in the oriented swap process. Annals of Applied Probability, 32(2):753–763, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-2/Absorbing-time-asymptotics-in-the-
oriented-swap-process/10.1214/21-AAP1695.full.

Bertsimas:2001:PMM

[BGT01] Dimitris Bertsimas, David Gamarnik, and John N. Tsitsiklis. Performance of multiclass Markovian queueing networks via piecewise linear Lyapunov functions. Annals of Applied Probability, 11(4):1384–1428, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1015345407.

Basdevant:2022:FOB

[BGT22] Anne-Laure Basdevant, Jean-Baptiste Gouéré, and Marie Théret. First-order behavior of the time constant in Bernoulli first-passage percolation. Annals of Applied Probability, 32(6): 4535–4567, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-6/First-order-behavior-of-the-time-constant-in-
Bernoulli-first/10.1214/22-AAP1795.full.

Blanca:2020:RCD

[BGV20] Antonio Blanca, Reza Gheissari, and Eric Vigoda. Randomcluster dynamics in Z^{2}: Rapid mixing with general boundary conditions. Annals of Applied Probability, 30(1):418–459, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1582621228.

Bhamidi:2015:DDS

[BGvdHK15] Shankar Bhamidi, Jesse Goodman, Remco van der Hofstad, and Júlia Komjáthy. Degree distribution of shortest path trees and bias of network sampling algorithms. Annals of Applied Probability, 25(4):1780–1826, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1432212430.

Beguin:1998:LTM

[BGY98] M. Beguin, L. Gray, and B. Ycart. The load transfer model. Annals of Applied Probability, 8(2):337–353, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1028903530.

Bean:1997:DEB

[BGZ97] N. G. Bean, R. J. Gibbens, and S. Zachary. Dynamic and equilibrium behavior of controlled loss networks. Annals of Applied Probability, 7(4):873–885, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1043862415.

Butez:2022:EPT

[BGZ22] Raphael Butez and David García-Zelada. Extremal particles of two-dimensional Coulomb gases and random polynomials on a positive background. Annals of Applied Probability, 32(1):392– 425, February 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
1/Extremal-particles-of-two-dimensional-Coulomb-gases-
and-random-polynomials/10.1214/21-AAP1682.full.

Benaim:1999:SAA

[BH99] Michel Benaïm and Morris W. Hirsch. Stochastic approximation algorithms with constant step size whose average is cooperative. Annals of Applied Probability, 9(1):216–241, February 1999. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1029962603.

Baccelli:2000:AEM

[BH00] François Baccelli and Dohy Hong. Analytic expansions of maxplus Lyapunov exponents. Annals of Applied Probability, 10 (3):779–827, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487510.

Berkes:2001:SAE

[BH01] István Berkes and Lajos Horváth. Strong approximation of the empirical process of GARCH sequences. Annals of Applied Probability, 11(3):789–809, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015345349. See correction

[Web01b].

Berkes:2003:ARL

[BH03a] István Berkes and Lajos Horváth. Asymptotic results for long memory LARCH sequences. Annals of Applied Probability, 13(2):641–668, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1050689598.

Berkes:2003:CSA

[BH03b] István Berkes and Lajos Horváth. Correction: “Strong approximation of the empirical process of GARCH sequences”. Annals of Applied Probability, 13(1):389, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1042765672. See [BH01].

Bai:2005:ARU

[BH05] Zhi-Dong Bai and Feifang Hu. Asymptotics in randomized urn models. Annals of Applied Probability, 15(1B):914–940, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1107271672.

Broutin:2012:TPL

[BH12] Nicolas Broutin and Cecilia Holmgren. The total path length of split trees. Annals of Applied Probability, 22(5):1745–1777, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1350067985.

Bayraktar:2013:RMA

[BH13] Erhan Bayraktar and Yu-Jui Huang. Robust maximization of asymptotic growth under covariance uncertainty. Annals of Applied Probability, 23(5):1817–1840, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1377696299.

Bovier:2017:ECE

[BH17] Anton Bovier and Lisa Hartung. Extended convergence of the extremal process of branching Brownian motion. Annals of Applied Probability, 27(3):1756–1777, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1500451241.

Braunsteins:2019:ELH

[BH19] Peter Braunsteins and Sophie Hautphenne. Extinction in lower Hessenberg branching processes with countably many types. Annals of Applied Probability, 29(5):2782–2818, October 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1571385622.

Bardenet:2020:MCD

[BH20a] Rémi Bardenet and Adrien Hardy. Monte Carlo with determinantal point processes. Annals of Applied Probability, 30(1):368– 417, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1582621227.

Ben-Hamou:2020:TCT

[BH20b] Anna Ben-Hamou. A threshold for cutoff in two-community random graphs. Annals of Applied Probability, 30(4):1824–1846, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/A-threshold-
for-cutoff-in-two-community-random-graphs/10.1214/19-
AAP1544.full.

Bao:2022:CMS

[BH22] Zhigang Bao and Yukun He. On Cramér–von Mises statistic for the spectral distribution of random matrices. Annals of Applied Probability, 32(6):4315–4355, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-6/On-Cram%c3%a9rvon-Mises-statistic-
for-the-spectral-distribution-of-random/10.1214/22-AAP1788.
full.

Bhattacharya:1999:MDP

[Bha99] Rabi Bhattacharya. Multiscale diffusion processes with periodic coefficients and an application to solute transport in porous media. Annals of Applied Probability, 9(4):951–1020, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962863.

Bhattacharya:2016:CTM

[Bha16] Bhaswar B. Bhattacharya. Collision times in multicolor urn models and sequential graph coloring with applications to discrete logarithms. Annals of Applied Probability, 26(6):3286– 3318, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1481792586.

Blondel:2020:RWR

[BHdS^{+}20] Oriane Blondel, Marcelo R. Hilário, Renato S. dos Santos, Vladas Sidoravicius, and Augusto Teixeira. Random walk on random walks: Low densities. Annals of Applied Probability, 30 (4):1614–1641, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/

issue-4/Random-walk-on-random-walks-Low-densities/10.
1214/19-AAP1537.full.

Berkes:2005:NIG

[BHK05] István Berkes, Lajos Horváth, and Piotr Kokoszka. Nearintegrated GARCH sequences. Annals of Applied Probability, 15(1B):890–913, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1107271671.

Berestycki:2011:TWH

[BHK11] J. Berestycki, S. C. Harris, and A. E. Kyprianou. Traveling waves and homogeneous fragmentation. Annals of Applied Probability, 21(5):1749–1794, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576608.

Barbour:1996:TIG

[BHL96] A. D. Barbour, J. A. P. Heesterbeek, and C. J. Luchsinger. Thresholds and initial growth rates in a model of parasitic infection. Annals of Applied Probability, 6(4):1045–1074, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1035463323.

Basu:2016:TGR

[BHMW16] Riddhipratim Basu, Alexander E. Holroyd, James B. Martin, and Johan Wästlund. Trapping games on random boards. Annals of Applied Probability, 26(6):3727–3753, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1481792598.

Baumdicker:2010:DDG

[BHP10] F. Baumdicker, W. R. Hess, and P. Pfaffelhuber. The diversity of a distributed genome in bacterial populations. Annals of Applied Probability, 20(5):1567–1606, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1282747394.

Bayer:2017:FLT

[BHQ17] Christian Bayer, Ulrich Horst, and Jinniao Qiu. A functional limit theorem for limit order books with state dependent price dynamics. Annals of Applied Probability, 27(5):2753–2806, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1509696034.

Bayer:2023:OSS

[BHRS23] Christian Bayer, Paul P. Hager, Sebastian Riedel, and John Schoenmakers. Optimal stopping with signatures. Annals of Applied Probability, 33(1):238–273, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-1/Optimal-stopping-with-signatures/10.
1214/22-AAP1814.full.

Bayraktar:2012:OMP

[BHS12] Erhan Bayraktar, Yu-Jui Huang, and Qingshuo Song. Outperforming the market portfolio with a given probability. Annals of Applied Probability, 22(4):1465–1494, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1344614201.

Bai:2002:GAT

[BHZ02] Z. D. Bai, Feifang Hu, and Li-Xin Zhang. Gaussian approximation theorems for urn models and their applications. Annals of Applied Probability, 12(4):1149–1173, November 2002. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1037125857.

Biggins:1995:GSG

[Big95] J. D. Biggins. The growth and spread of the general branching random walk. Annals of Applied Probability, 5(4):1008–1024, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004604.

Biggins:2012:SSR

[Big12] J. D. Biggins. Spreading speeds in reducible multitype branching random walk. Annals of Applied Probability, 22(5):1778–1821, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1350067986.

Barrera:2020:TSR

[BJ20] Gerardo Barrera and Milton Jara. Thermalisation for small random perturbations of dynamical systems. Annals of Applied Probability, 30(3):1164–1208, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-3/Thermalisation-for-small-random-perturbations-
of-dynamical-systems/10.1214/19-AAP1526.full.

Bencheikh:2022:WSE

[BJ22] Oumaima Bencheikh and Benjamin Jourdain. Weak and strong error analysis for mean-field rank-based particle approximations of one-dimensional viscous scalar conservation laws. Annals of Applied Probability, 32(6):4143–4185, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/Weak-
and-strong-error-analysis-for-mean-field-rank-based/
10.1214/21-AAP1776.full.

Bielecki:2021:SSF

[BJJN21] Tomasz R. Bielecki, Jacek Jakubowski, Monique Jeanblanc, and Mariusz Niewegłowski. Semimartingales and shrinkage of filtration. Annals of Applied Probability, 31(3):1376–1402, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-3/Semimartingales-
and-shrinkage-of-filtration/10.1214/20-AAP1621.full.

Beskos:2016:CAS

[BJKT16] Alexandros Beskos, Ajay Jasra, Nikolas Kantas, and Alexandre Thiery. On the convergence of adaptive sequential Monte Carlo methods. Annals of Applied Probability, 26(2):1111–1146, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1458651829.

Barral:2010:CCM

[BJM10a] Julien Barral, Xiong Jin, and Benoît Mandelbrot. Convergence of complex multiplicative cascades. Annals of Applied Probability, 20(4):1219–1252, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1279638785.

Barral:2010:UCC

[BJM10b] Julien Barral, Xiong Jin, and Benoît Mandelbrot. Uniform convergence for complex [0, 1]-martingales. Annals of Applied Probability, 20(4):1205–1218, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1279638784.

Bhamidi:2018:CPD

[BJN18] Shankar Bhamidi, Jimmy Jin, and Andrew Nobel. Change point detection in network models: Preferential attachment and long range dependence. Annals of Applied Probability, 28(1):35–78, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1520046083.

Bielecki:2008:PTC

[BJR08] Tomasz R. Bielecki, Monique Jeanblanc, and Marek Rutkowski. Pricing and trading credit default swaps in a hazard process model. Annals of Applied Probability, 18(6):2495–2529, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1227708926.

Bibinger:2016:VEU

[BJR16] Markus Bibinger, Moritz Jirak, and Markus Reiß. Volatility estimation under one-sided errors with applications to limit order books. Annals of Applied Probability, 26(5):2754–2790, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1476884303.

Benjamini:1992:MSP

[BK92] Yoav Benjamini and Abba M. Krieger. Market share paradox and heterogeneous chains. Annals of Applied Probability, 2(4): 1019–1023, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005587.

Bose:1995:DAA

[BK95] A. Bose and I. Kaj. Diffusion approximation for an agestructured population. Annals of Applied Probability, 5(1):140– 157, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004833.

Burdzy:1998:BMB

[BK98] Krzysztof Burdzy and Davar Khoshnevisan. Brownian motion in a Brownian crack. Annals of Applied Probability, 8(3):708–748, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903448.

Boutsikas:2000:BDS

[BK00a] Michael V. Boutsikas and Markos V. Koutras. A bound for the distribution of the sum of discrete associated or negatively associated random variables. Annals of Applied Probability, 10 (4):1137–1150, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487610.

Burdzy:2000:EMC

[BK00b] Krzysztof Burdzy and Wilfrid S. Kendall. Efficient Markovian couplings: examples and counterexamples. Annals of Applied Probability, 10(2):362–409, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487348.

Borkovec:2001:TSD

[BK01] Milan Borkovec and Claudia Klüppelberg. The tail of the stationary distribution of an autoregressive process with Arch(1) errors. Annals of Applied Probability, 11(4):1220–1241, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345401.

Bolthausen:2006:NVG

[BK06] Erwin Bolthausen and Nicola Kistler. On a nonhierarchical version of the generalized random energy model. Annals of Applied Probability, 16(1):1–14, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654278.

Bank:2015:MLI

[BK15a] Peter Bank and Dmitry Kramkov. A model for a large investor trading at market indifference prices. II: Continuous-time case.

Annals of Applied Probability, 25(5):2708–2742, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1438261052.

BenAlaya:2015:CLT

[BK15b] Mohamed Ben Alaya and Ahmed Kebaier. Central limit theorem for the multilevel Monte Carlo Euler method. Annals of Applied Probability, 25(1):211–234, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1418740184.

Belomestny:2016:OSU

[BK16a] Denis Belomestny and Volker Krätschmer. Optimal stopping under model uncertainty: Randomized stopping times approach. Annals of Applied Probability, 26(2):1260–1295, April 2016. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1458651832. See

addendum [BK17b].

Bertoin:2016:SSS

[BK16b] Jean Bertoin and Igor Kortchemski. Self-similar scaling limits of Markov chains on the positive integers. Annals of Applied Probability, 26(4):2556–2595, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1472745467.

Bank:2017:CDS

[BK17a] Peter Bank and Helena Kauppila. Convex duality for stochastic singular control problems. Annals of Applied Probability, 27 (1):485–516, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1488790833.

Belomestny:2017:AOS

[BK17b] Denis Belomestny and Volker Krätschmer. Addendum to “Optimal stopping under model uncertainty: Randomized stopping times approach”. Annals of Applied Probability, 27(2):1289– 1293, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1495764379. See [BK16a].

Bally:2015:PIP

[BKH15] Vlad Bally and Arturo Kohatsu-Higa. A probabilistic interpretation of the parametrix method. Annals of Applied Probability, 25(6):3095–3138, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1443703770.

BenAlaya:2022:CLT

[BKN22] Mohamed Ben Alaya, Ahmed Kebaier, and Thi Bao Tram Ngo. Central limit theorem for the antithetic multilevel Monte Carlo method. Annals of Applied Probability, 32(3):1970–2027, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Central-
limit-theorem-for-the-antithetic-multilevel-Monte-Carlo-
method/10.1214/21-AAP1726.full.

Ball:2006:AAM

[BKPR06] Karen Ball, Thomas G. Kurtz, Lea Popovic, and Greg Rempala. Asymptotic analysis of multiscale approximations to reaction networks. Annals of Applied Probability, 16(4):1925–1961, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1169065212.

Bierkens:2022:HDS

[BKR22] Joris Bierkens, Kengo Kamatani, and Gareth O. Roberts. Highdimensional scaling limits of piecewise deterministic sampling algorithms. Annals of Applied Probability, 32(5):3361–3407, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-5/High-
dimensional-scaling-limits-of-piecewise-deterministic-
sampling-algorithms/10.1214/21-AAP1762.full.

Bobrowski:2017:MPC

[BKS17] Omer Bobrowski, Matthew Kahle, and Primoz Skraba. Maximally persistent cycles in random geometric complexes. Annals of Applied Probability, 27(4):2032–2060, August 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1504080025.

Butkovsky:2020:GCE

[BKS20] Oleg Butkovsky, Alexei Kulik, and Michael Scheutzow. Generalized couplings and ergodic rates for SPDEs and other Markov models. Annals of Applied Probability, 30(1):1–39, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1582621218.

Burdzy:2022:SFP

[BKT22] Krzysztof Burdzy, Bartosz Kołodziejek, and Tvrtko Tadi’c. Stochastic fixed-point equation and local dependence measure. Annals of Applied Probability, 32(4):2811–2840, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Stochastic-
fixed-point-equation-and-local-dependence-measure/10.
1214/21-AAP1749.full.

Burden:2008:AWM

[BKW08] Conrad J. Burden, Miriam R. Kantorovitz, and Susan R. Wilson. Approximate word matches between two random sequences. Annals of Applied Probability, 18(1):1–21, February 2008. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1199890013.

Bramson:2001:SSL

[BL01] Maury Bramson and Joel L. Lebowitz. Spatial structure in low dimensions for diffusion limited two-particle reactions. Annals of Applied Probability, 11(1):121–181, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/998926989.

Boyarchenko:2002:BOT

[BL02] Svetlana Boyarchenko and Sergei Levendorskiĭ. Barrier options and touch-and-out options under regular Lévy processes of exponential type. Annals of Applied Probability, 12(4):1261–1298, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1037125863.

Belhadji:2006:IVC

[BL06] L. Belhadji and N. Lanchier. Individual versus cluster recoveries within a spatially structured population. Annals of Applied Probability, 16(1):403–422, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654292.

Barbour:2008:LLN

[BL08] A. D. Barbour and M. J. Luczak. Laws of large numbers for epidemic models with countably many types. Annals of Applied Probability, 18(6):2208–2238, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1227708917.

Brightwell:2011:OIM

[BL11] Graham Brightwell and Malwina Luczak. Order-invariant measures on causal sets. Annals of Applied Probability, 21(4):1493– 1536, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1312818843.

Bassetti:2012:SSS

[BL12a] Federico Bassetti and Lucia Ladelli. Self-similar solutions in one-dimensional kinetic models: a probabilistic view. Annals of Applied Probability, 22(5):1928–1961, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1350067990.

Bodineau:2012:LDE

[BL12b] Thierry Bodineau and Maxime Lagouge. Large deviations of the empirical currents for a boundary-driven reaction diffusion model. Annals of Applied Probability, 22(6):2282–2319, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1353695954.

Briand:2014:SBW

[BL14] Philippe Briand and Céline Labart. Simulation of BSDEs by Wiener chaos expansion. Annals of Applied Probability, 24 (3):1129–1171, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1398258097.

Bayraktar:2016:SPS

[BL16a] Erhan Bayraktar and Jiaqi Li. Stochastic Perron for stochastic target games. Annals of Applied Probability, 26(2):1082–1110, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1458651828.

Benaim:2016:LVR

[BL16b] Michel Benaïm and Claude Lobry. Lotka–Volterra with randomly fluctuating environments or “how switching between beneficial environments can make survival harder”. Annals of Applied Probability, 26(6):3754–3785, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1481792599.

Bobkov:2021:SFA

[BL21] Sergey G. Bobkov and Michel Ledoux. A simple Fourier analytic proof of the AKT optimal matching theorem. Annals of Applied Probability, 31(6):2567–2584, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-6/A-simple-Fourier-analytic-proof-of-
the-AKT-optimal-matching/10.1214/20-AAP1656.full.

Bladt:1996:VCA

[Bla96] Mogens Bladt. The variance constant for the actual waiting time of the P H/P H/1 queue. Annals of Applied Probability, 6(3):766– 777, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1034968227.

Blanchet:2009:EIS

[Bla09] Jose H. Blanchet. Efficient importance sampling for binary contingency tables. Annals of Applied Probability, 19(3):949–982, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1245071015.

Bayati:2015:UPP

[BLM15] Mohsen Bayati, Marc Lelarge, and Andrea Montanari. Universality in polytope phase transitions and message passing algorithms. Annals of Applied Probability, 25(2):753–822, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1424355130.

Buckdahn:2017:MFS

[BLM17] Rainer Buckdahn, Juan Li, and Jin Ma. A mean-field stochastic control problem with partial observations. Annals of Applied Probability, 27(5):3201–3245, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1509696045.

Buckdahn:2023:GCM

[BLM23] Rainer Buckdahn, Juan Li, and Jin Ma. A general conditional McKean–Vlasov stochastic differential equation. Annals of Applied Probability, 33(3):2004–2023, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-3/A-general-conditional-McKeanVlasov-
stochastic-differential-equation/10.1214/22-AAP1858.
full.

Benaim:2014:SPR

[BLMZ14] Michel Benaïm, Stéphane Le Borgne, Florent Malrieu, and Pierre-André Zitt. On the stability of planar randomly switched systems. Annals of Applied Probability, 24(1):292–311, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1389278726.

Blount:1992:LLN

[Blo92] Douglas Blount. Law of large numbers in the supremum norm for a chemical reaction with diffusion. Annals of Applied Probability, 2(1):131–141, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005774.

Bloznelis:2013:DCC

[Blo13] Mindaugas Bloznelis. Degree and clustering coefficient in sparse random intersection graphs. Annals of Applied Probability, 23 (3):1254–1289, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1362684860.

Blondel:2015:TDL

[Blo15] Oriane Blondel. Tracer diffusion at low temperature in kinetically constrained models. Annals of Applied Probability, 25 (3):1079–1107, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1427124124.

Bonnier:2023:ATH

[BLO23] Patric Bonnier, Chong Liu, and Harald Oberhauser. Adapted topologies and higher rank signatures. Annals of Applied Probability, 33(3):2136–2175, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-3/Adapted-topologies-and-higher-rank-
signatures/10.1214/22-AAP1862.full.

Bramson:2013:DTE

[BLP13] Maury Bramson, Yi Lu, and Balaji Prabhakar. Decay of tails at equilibrium for FIFO join the shortest queue networks. Annals of Applied Probability, 23(5):1841–1878, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1377696300.

Biggins:1991:BRW

[BLSW91] J. D. Biggins, Boris D. Lubachevsky, Adam Shwartz, and Alan Weiss. A branching random walk with a barrier. Annals of Applied Probability, 1(4):573–581, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005839.

Bai:2011:APE

[BLW11] Z. D. Bai, H. X. Liu, and W. K. Wong. Asymptotic properties of eigenmatrices of a large sample covariance matrix. Annals of Applied Probability, 21(5):1994–2015, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1319576615.

Bo:2022:RSC

[BLY22] Lijun Bo, Huafu Liao, and Xiang Yu. Risk-sensitive credit portfolio optimization under partial information and contagion risk. Annals of Applied Probability, 32(4):2355–2399, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Risk-
sensitive-credit-portfolio-optimization-under-partial-
information-and-contagion/10.1214/21-AAP1735.full.

Bertacchi:2011:CVP

[BLZ11] D. Bertacchi, N. Lanchier, and F. Zucca. Contact and voter processes on the infinite percolation cluster as models of hostsymbiont interactions. Annals of Applied Probability, 21(4): 1215–1252, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1312818835.

Burdzy:1996:IFA

[BM96] Krzysztof Burdzy and Andrzej Madrecki. Itô formula for an asymptotically 4-stable process. Annals of Applied Probability, 6(1):200–217, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968071.

Bohm:1997:KMT

[BM97] W. Böhm and S. G. Mohanty. On the Karlin–McGregor theorem and applications. Annals of Applied Probability, 7(2):314–325, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034625333.

Bovier:2001:EVB

[BM01] Anton Bovier and David M. Mason. Extreme value behavior in the Hopfield model. Annals of Applied Probability, 11(1):91–120, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
998926988.

Bhattacharya:2004:SDR

[BM04] Rabi Bhattacharya and Mukul Majumdar. Stability in distribution of randomly perturbed quadratic maps as Markov processes. Annals of Applied Probability, 14(4):1802–1809, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1099674078.

Bousquet-Melou:2005:WQP

[BM05] Mireille Bousquet-Mélou. Walks in the quarter plane: Kreweras’ algebraic model. Annals of Applied Probability, 15(2):1451–1491, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1115137982.

Benjamini:2012:BAM

[BM12] Itai Benjamini and Yury Makarychev. Balanced allocation: Memory performance tradeoffs. Annals of Applied Probability, 22(4):1642–1649, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1344614206.

Bertoin:2013:CTL

[BM13] Jean Bertoin and Grégory Miermont. The cut-tree of large Galton–Watson trees and the Brownian CRT. Annals of Applied Probability, 23(4):1469–1493, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1371834035.

Bhattacharya:2017:DSR

[BM17] Bhaswar B. Bhattacharya and Sumit Mukherjee. Degree sequence of random permutation graphs. Annals of Applied Probability, 27(1):439–484, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1488790832.

Banerjee:2019:JSQ

[BM19a] Sayan Banerjee and Debankur Mukherjee. Join-the-shortest queue diffusion limit in Halfin–Whitt regime: Tail asymptotics and scaling of extrema. Annals of Applied Probability, 29 (2):1262–1309, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1548298941.

Broker:2019:LGM

[BM19b] Yannic Bröker and Chiranjib Mukherjee. Localization of the Gaussian multiplicative chaos in the Wiener space and the stochastic heat equation in strong disorder. Annals of Applied Probability, 29(6):3745–3785, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1578366325.

Banerjee:2020:JSQ

[BM20] Sayan Banerjee and Debankur Mukherjee. Join-the-Shortest Queue diffusion limit in Halfin–Whitt regime: Sensitivity on the heavy-traffic parameter. Annals of Applied Probability, 30 (1):80–144, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1582621220.

Bartl:2022:MCM

[BM22] Daniel Bartl and Shahar Mendelson. On Monte-Carlo methods in convex stochastic optimization. Annals of Applied Probability, 32(4):3146–3198, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-4/On-Monte-Carlo-methods-in-convex-stochastic-
optimization/10.1214/22-AAP1781.full.

Bousquet-Melou:2006:DIL

[BMJ06] Mireille Bousquet-Mélou and Svante Janson. The density of the ISE and local limit laws for embedded trees. Annals of Applied Probability, 16(3):1597–1632, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1159804993.

Bhattacharya:2021:ADB

[BMM21] Bhaswar B. Bhattacharya, Somabha Mukherjee, and Sumit Mukherjee. Asymptotic distribution of Bernoulli quadratic forms. Annals of Applied Probability, 31(4):1548–1597, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/Asymptotic-
distribution-of-Bernoulli-quadratic-forms/10.1214/20-
AAP1626.full.

Bouchard:2014:STG

[BMN14] Bruno Bouchard, Ludovic Moreau, and Marcel Nutz. Stochastic target games with controlled loss. Annals of Applied Probability, 24(3):899–934, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1398258092.

Buckdahn:2008:SCP

[BMR08] Rainer Buckdahn, Jin Ma, and Catherine Rainer. Stochastic control problems for systems driven by normal martingales. Annals of Applied Probability, 18(2):632–663, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1206018200.

Bazaes:2023:QAL

[BMRS23] Rodrigo Bazaes, Chiranjib Mukherjee, Alejandro F. Ramírez, and Santiago Saglietti. Quenched and averaged large deviations for random walks in random environments: The impact of disorder. Annals of Applied Probability, 33(3):2210–2246, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/Quenched-
and-averaged-large-deviations-for-random-walks-in-random/
10.1214/22-AAP1864.full.

Braverman:2002:TPS

[BMS02] Michael Braverman, Thomas Mikosch, and Gennady Samorodnitsky. Tail probabilities of subadditive functionals of Lévy processes. Annals of Applied Probability, 12(1):69–100, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015961156.

Ball:1997:ETL

[BMST97] Frank Ball, Denis Mollison, and Gianpaolo Scalia-Tomba. Epidemics with two levels of mixing. Annals of Applied Probability, 7(1):46–89, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034625252.

Budhiraja:2019:SMG

[BMW19] Amarjit Budhiraja, Debankur Mukherjee, and Ruoyu Wu. Supermarket model on graphs. Annals of Applied Probability, 29 (3):1740–1777, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1550566841.

Bai:2017:ODI

[BMX17] Lihua Bai, Jin Ma, and Xiaojing Xing. Optimal dividend and investment problems under Sparre–Andersen model. Annals of Applied Probability, 27(6):3588–3632, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1513328709.

Bramson:1997:CCS

[BN97] Maury Bramson and Claudia Neuhauser. Coexistence for a catalytic surface reaction model. Annals of Applied Probability, 7(3):565–614, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034801245.

Bayraktar:2015:WRP

[BN15a] Erhan Bayraktar and Sergey Nadtochiy. Weak reflection principle for Lévy processes. Annals of Applied Probability, 25(6): 3251–3294, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1443703774.

Bouchard:2015:ADN

[BN15b] Bruno Bouchard and Marcel Nutz. Arbitrage and duality in nondominated discrete-time models. Annals of Applied Probability, 25(2):823–859, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355131.

Ball:2017:AVG

[BN17] Frank Ball and Peter Neal. The asymptotic variance of the giant component of configuration model random graphs. Annals of Applied Probability, 27(2):1057–1092, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1495764374.

Bresler:2019:SMS

[BN19] Guy Bresler and Dheeraj Nagaraj. Stein’s method for stationary distributions of Markov chains and application to Ising models. Annals of Applied Probability, 29(5):3230–3265, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1571385634.

Bion-Nadal:2012:RMU

[BNK12] Jocelyne Bion-Nadal and Magali Kervarec. Risk measuring under model uncertainty. Annals of Applied Probability, 22(1):213– 238, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1328623699.

Barndorff-Nielsen:2011:MSP

[BNS11] Ole Eiler Barndorff-Nielsen and Robert Stelzer. Multivariate supOU processes. Annals of Applied Probability, 21(1):140–182, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1292598030.

Broutin:2013:LPP

[BNS13] Nicolas Broutin, Ralph Neininger, and Henning Sulzbach. A limit process for partial match queries in random quadtrees and 2-D trees. Annals of Applied Probability, 23(6):2560–2603, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1382447698.

Bierkens:2021:LDE

[BNS21] Joris Bierkens, Pierre Nyquist, and Mikola C. Schlottke. Large deviations for the empirical measure of the zig-zag process. Annals of Applied Probability, 31(6):2811–2843, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-6/Large-
deviations-for-the-empirical-measure-of-the-zig-zag/
10.1214/21-AAP1663.full.

Bramson:1992:PCM

[BNT92] M. Bramson, P. Ney, and J. Tao. The population composition of a multitype branching random walk. Annals of Applied Probability, 2(3):575–596, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005650.

Bion–Nadal:2019:WTD

[BNT19] Jocelyne Bion-Nadal and Denis Talay. On a Wasserstein-type distance between solutions to stochastic differential equations. Annals of Applied Probability, 29(3):1609–1639, June 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1550566838.

Banerjee:2022:PAD

[BOC22]  Sayan Banerjee and Mariana Olvera-Cravioto.  PageR-

ank asymptotics on directed preferential attachment networks. Annals of Applied Probability, 32(4):3060–3084, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/PageRank-

asymptotics-on-directed-preferential-attachment-networks/
10.1214/21-AAP1757.full.

Bordenave:2008:NPP

[Bor08] Charles Bordenave. Navigation on a Poisson point process. Annals of Applied Probability, 18(2):708–746, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1206018202.

Borrello:2012:RAE

[Bor12] Davide Borrello. On the role of Allee effect and mass migration in survival and extinction of a species. Annals of Applied Probability, 22(2):670–701, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1333372010.

Bornemann:2016:NES

[Bor16] Folkmar Bornemann. A note on the expansion of the smallest eigenvalue distribution of the LUE at the hard edge. Annals of Applied Probability, 26(3):1942–1946, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905023.

Brightwell:1995:TSP

[BOW95] Graham Brightwell, Teunis J. Ott, and Peter Winkler. Target shooting with programmed random variables. Annals of Applied Probability, 5(3):834–853, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004707.

Bouton:1997:AMC

[BP97] Catherine Bouton and Gilles Pagès. About the multidimensional competitive learning vector quantization algorithm with constant gain. Annals of Applied Probability, 7(3):679–710, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034801249.

Barbour:2005:ABM

[BP05a] A. D. Barbour and A. Pugliese. Asymptotic behavior of a metapopulation model. Annals of Applied Probability, 15 (2):1306–1338, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1115137976.

Bouchard:2005:OCD

[BP05b] Bruno Bouchard and Huyên Pham. Optimal consumption in discrete-time financial models with industrial investment opportunities and nonlinear returns. Annals of Applied Probability, 15 (4):2393–2421, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965766.

Berestycki:2012:ESL

[BP12] Nathanaël Berestycki and Richard Pymar. Effect of scale on long-range random graphs and chromosomal inversions. Annals of Applied Probability, 22(4):1328–1361, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1344614197.

Biskup:2018:EVP

[BP18] Marek Biskup and Eviatar B. Procaccia. Eigenvalue versus perimeter in a shape theorem for self-interacting random walks. Annals of Applied Probability, 28(1):340–377, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1520046090.

Bakhtin:2019:MCA

[BPG19] Yuri Bakhtin and Zsolt Pajor-Gyulai. Malliavin calculus approach to long exit times from an unstable equilibrium. Annals of Applied Probability, 29(2):827–850, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1548298931. See erratum

[BCPG22].

Baldasso:2022:LDI

[BPR22] Rangel Baldasso, Alan Pereira, and Guilherme Reis. Large deviations for interacting diffusions with path-dependent McKean– Vlasov limit. Annals of Applied Probability, 32(1):665–695, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Large-
deviations-for-interacting-diffusions-with-path-dependent-
McKeanVlasov-limit/10.1214/21-AAP1692.full.

Breyer:2004:OSM

[BPS04] Laird Arnault Breyer, Mauro Piccioni, and Sergio Scarlatti. Optimal scaling of MaLa for nonlinear regression. Annals of Applied Probability, 14(3):1479–1505, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1089736293.

Bertsimas:1994:OMQ

[BPT94] Dimitris Bertsimas, Ioannis Ch. Paschalidis, and John N. Tsitsiklis. Optimization of multiclass queueing networks: Polyhedral and nonlinear characterizations of achievable performance. Annals of Applied Probability, 4(1):43–75, February 1994. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005200.

Bertsimas:1998:LDB

[BPT98] Dimitris Bertsimas, Ioannis Ch. Paschalidis, and John N. Tsitsiklis. On the large deviations behavior of acyclic networks of G/G/1 queues. Annals of Applied Probability, 8(4):1027–1069, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1028903373.

Berger:2022:GSH

[BPT22] Noam Berger, Eviatar B. Procaccia, and Amanda Turner. Growth of stationary Hastings–Levitov. Annals of Applied Probability, 32(5):3331–3360, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Growth-of-stationary-HastingsLevitov/
10.1214/21-AAP1761.full.

Baryshnikov:2009:GLG

[BPY09] Yu. Baryshnikov, Mathew D. Penrose, and J. E. Yukich. Gaussian limits for generalized spacings. Annals of Applied Probability, 19(1):158–185, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1235140336.

Bertacchi:2007:EER

[BPZ07] Daniela Bertacchi, Gustavo Posta, and Fabio Zucca. Ecological equilibrium for restrained branching random walks. Annals of Applied Probability, 17(4):1117–1137, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1186755234.

Bayraktar:2019:CRS

[BQ19] Erhan Bayraktar and Jinniao Qiu. Controlled reflected SDEs and Neumann problem for backward SPDEs. Annals of Applied Probability, 29(5):2819–2848, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1571385623.

Blaszczyszyn:1993:QSL

[BR93] B. Blaszczyszyn and T. Rolski. Queues in series in light traffic. Annals of Applied Probability, 3(3):881–896, August 1993. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005370.

Bank:2001:OCC

[BR01] Peter Bank and Frank Riedel. Optimal consumption choice with intertemporal substitution. Annals of Applied Probability, 11(3):750–788, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015345348.

Boroczky:2004:ASC

[BR04] Károly Böröczky and Matthias Reitzner. Approximation of smooth convex bodies by random circumscribed polytopes. Annals of Applied Probability, 14(1):239–273, February 2004. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1075828053.

Beskos:2005:ESD

[BR05] Alexandros Beskos and Gareth O. Roberts. Exact simulation of diffusions. Annals of Applied Probability, 15(4):2422–2444, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1133965767.

Budhiraja:2006:EOC

[BR06] Amarjit Budhiraja and Kevin Ross. Existence of optimal controls for singular control problems with state constraints. Annals of Applied Probability, 16(4):2235–2255, November 2006. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1169065223.

Budhiraja:2008:OSF

[BR08] Amarjit Budhiraja and Kevin Ross. Optimal stopping and free boundary characterizations for some Brownian control problems. Annals of Applied Probability, 18(6):2367–2391, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1227708922.

Budhiraja:2013:NCC

[BR13] Amarjit Budhiraja and Dominik Reinhold. Near critical catalyst reactant branching processes with controlled immigration. Annals of Applied Probability, 23(5):2053–2098, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1377696306.

Biskup:2015:GMP

[BR15] Marek Biskup and Thomas Richthammer. Gibbs measures on permutations over one-dimensional discrete point sets. Annals of Applied Probability, 25(2):898–929, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1424355133.

Bierkens:2017:PDS

[BR17a] Joris Bierkens and Gareth Roberts. A piecewise deterministic scaling limit of lifted Metropolis–Hastings in the Curie-Weiss model. Annals of Applied Probability, 27(2):846–882, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1495764368.

Blummel:2017:FML

[BR17b] Tilmann Blümmel and Thorsten Rheinländer. Financial markets with a large trader. Annals of Applied Probability, 27(6): 3735–3786, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1513328713.

Barbour:2019:CLT

[BR19] A. D. Barbour and Adrian Röllin. Central limit theorems in the configuration model. Annals of Applied Probability, 29(2):1046– 1069, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1548298936.

Bramson:1994:CIF

[Bra94a] Maury Bramson. Correction: “Instability of FIFO Queueing Networks”. Annals of Applied Probability, 4(3):952, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1177004979. See [Bra94b].

Bramson:1994:IFQa

[Bra94b] Maury Bramson. Instability of FIFO queueing networks. Annals of Applied Probability, 4(2):414–431, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005066. See correction

[Bra94a].

Bramson:1994:IFQb

[Bra94c] Maury Bramson. Instability of FIFO queueing networks with quick service times. Annals of Applied Probability, 4(3):693–718, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004967.

Bramson:1999:SQN

[Bra99] Maury Bramson. A stable queueing network with unstable fluid model. Annals of Applied Probability, 9(3):818–853, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962815.

Bramson:2010:NSU

[Bra10] Maury Bramson. Network stability under max–min fair bandwidth sharing. Annals of Applied Probability, 20(3):1126–1176, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1276867309.

Bramson:2011:PRR

[Bra11a] Maury Bramson. A positive recurrent reflecting Brownian motion with divergent fluid path. Annals of Applied Probability, 21(3):951–986, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1307020388.

Bramson:2011:SJS

[Bra11b] Maury Bramson. Stability of join the shortest queue networks. Annals of Applied Probability, 21(4):1568–1625, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1312818845.

Bou-Rabee:2020:CCH

[BREZ20] Nawaf Bou-Rabee, Andreas Eberle, and Raphael Zimmer. Coupling and convergence for Hamiltonian Monte Carlo. Annals of Applied Probability, 30(3):1209–1250, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-3/Coupling-and-convergence-for-Hamiltonian-
Monte-Carlo/10.1214/19-AAP1528.full.

Brooks:1999:PML

[Bro99] Elizabeth A. Brooks. Probabilistic methods for a linear reactionhyperbolic system with constant coefficients. Annals of Applied Probability, 9(3):719–731, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1029962811.

Beskos:2009:OSL

[BRS09] Alexandros Beskos, Gareth Roberts, and Andrew Stuart. Optimal scalings for local Metropolis–Hastings chains on nonproduct targets in high dimensions. Annals of Applied Probability, 19(3):863–898, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1245071013.

Bou-Rabee:2017:RHM

[BRSS17] Nawaf Bou-Rabee and Jesús María Sanz-Serna. Randomized Hamiltonian Monte Carlo. Annals of Applied Probability, 27 (4):2159–2194, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1504080029.

Blancas:2019:NKC

[BRSSJ19] Airam Blancas, Tim Rogers, Jason Schweinsberg, and Arno Siri-Jégousse. The nested Kingman coalescent: Speed of coming down from infinity. Annals of Applied Probability, 29(3):1808– 1836, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1550566843.

Beskos:2018:AAR

[BRTP18] Alexandros Beskos, Gareth Roberts, Alexandre Thiery, and Natesh Pillai. Asymptotic analysis of the random walk Metropolis algorithm on ridged densities. Annals of Applied Probability, 28(5):2966–3001, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1535443239.

Bierkens:2019:EZP

[BRZ19] Joris Bierkens, Gareth O. Roberts, and Pierre-André Zitt. Ergodicity of the zigzag process. Annals of Applied Probability, 29(4):2266–2301, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1563869043.

Baccelli:1996:TSE

[BS96] François Baccelli and Volker Schmidt. Taylor series expansions for Poisson-driven (max, +)-linear systems. Annals of Applied Probability, 6(1):138–185, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968069.

Becherer:2005:CSR

[BS05] Dirk Becherer and Martin Schweizer. Classical solutions to reaction–diffusion systems for hedging problems with interacting Itô and point processes. Annals of Applied Probability, 15(2):1111–1144, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1115137970.

Bai:2007:SIR

[BS07] Z. D. Bai and Jack W. Silverstein. On the signal-to-interference ratio of CDMA systems in wireless communications. Annals of Applied Probability, 17(1):81–101, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1171377178.

Basdevant:2012:CTV

[BS12] Anne-Laure Basdevant and Arvind Singh. Continuous-time vertex reinforced jump processes on Galton–Watson trees. Annals of Applied Probability, 22(4):1728–1743, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1344614210.

Brunick:2013:MIP

[BS13] Gerard Brunick and Steven Shreve. Mimicking an Itô process by a solution of a stochastic differential equation. Annals of Applied Probability, 23(4):1584–1628, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1371834039.

Bayer:2014:SFR

[BS14] Christian Bayer and John Schoenmakers. Simulation of forwardreverse stochastic representations for conditional diffusions. Annals of Applied Probability, 24(5):1994–2032, October 2014. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1403812368.

Bandyopadhyay:2015:ETN

[BS15] Antar Bandyopadhyay and Farkhondeh Sajadi. On the expected total number of infections for virus spread on a finite network. Annals of Applied Probability, 25(2):663–674, April 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1424355127.

Basu:2017:EVM

[BS17a] Riddhipratim Basu and Allan Sly. Evolving voter model on dense random graphs. Annals of Applied Probability, 27(2):1235–1288, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1495764378.

Birkner:2017:ODR

[BS17b] Matthias Birkner and Rongfeng Sun. One-dimensional random walks with self-blocking immigration. Annals of Applied Probability, 27(1):109–139, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1488790824.

Benaim:2019:RSB

[BS19] Michel Benaïm and Edouard Strickler. Random switching between vector fields having a common zero. Annals of Applied Probability, 29(1):326–375, February 2019. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1544000431.

Borga:2020:SPT

[BS20] Jacopo Borga and Erik Slivken. Square permutations are typically rectangular. Annals of Applied Probability, 30(5):2196– 2233, October 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
5/Square-permutations-are-typically-rectangular/10.1214/
19-AAP1555.full.

Bender:2021:SRA

[BS21] Christian Bender and Nikolaus Schweizer. “Regression anytime” with brute-force SVD truncation. Annals of Applied Probability, 31(3):1140–1179, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-3/Regression-anytime-with-brute-force-SVD-truncation/
10.1214/20-AAP1615.full.

Bhattacharjee:2022:LDS

[BS22] Chinmoy Bhattacharjee and Matthias Schulte. Large degrees in scale-free inhomogeneous random graphs. Annals of Applied Probability, 32(1):696–720, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-1/Large-degrees-in-scale-free-inhomogeneous-
random-graphs/10.1214/21-AAP1693.full.

Bello:2023:CCA

[BS23] Jason Bello and David J. Sivakoff. Cyclic cellular automata and Greenberg-Hastings models on regular trees. Annals of Applied Probability, 33(4):3080–3097, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-4/Cyclic-cellular-automata-and-GreenbergHastings-
models-on-regular-trees/10.1214/22-AAP1885.full.

Benaim:2004:GUM

[BST04] Michel Benaïm, Sebastian J. Schreiber, and Pierre Tarrès. Generalized URN models of evolutionary processes. Annals of Applied Probability, 14(3):1455–1478, August 2004. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1089736292.

Ball:2014:ERI

[BST14] Frank G. Ball, David J. Sirl, and Pieter Trapman. Epidemics on random intersection graphs. Annals of Applied Probability, 24(3):1081–1128, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1398258096.

Bhamidi:2015:TEN

[BSZ15] Shankar Bhamidi, J. Michael Steele, and Tauhid Zaman. Twitter event networks and the superstar model. Annals of Applied Probability, 25(5):2462–2502, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1438261046.

Betz:2020:RPM

[BSZ20] Volker Betz, Helge Schäfer, and Dirk Zeindler. Random permutations without macroscopic cycles. Annals of Applied Probability, 30(3):1484–1505, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-3/Random-permutations-without-macroscopic-cycles/
10.1214/19-AAP1538.full.

Bossy:1996:CRA

[BT96] Mireille Bossy and Denis Talay. Convergence rate for the approximation of the limit law of weakly interacting particles: application to the Burgers equation. Annals of Applied Probability, 6(3):818–861, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968229.

Bouchard:2000:ESM

[BT00] Bruno Bouchard and Nizar Touzi. Explicit solution to the multivariate super-replication problem under transaction costs. Annals of Applied Probability, 10(3):685–708, August 2000. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1019487506.

Baudoin:2005:HID

[BT05] Fabrice Baudoin and Josef Teichmann. Hypoellipticity in infinite dimensions and an application in interest rate theory. Annals of Applied Probability, 15(3):1765–1777, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1121433768.

Bercu:2008:EIS

[BT08] Bernard Bercu and Abderrahmen Touati. Exponential inequalities for self-normalized martingales with applications. Annals of Applied Probability, 18(5):1848–1869, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1225372953.

Bordenave:2010:LOP

[BT10] Charles Bordenave and Giovanni Luca Torrisi. Load optimization in a planar network. Annals of Applied Probability, 20(6): 2040–2085, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1287494554.

Burnashev:2012:TRW

[BT12] Marat V. Burnashev and Aslan Tchamkerten. Tracking a random walk first-passage time through noisy observations. Annals of Applied Probability, 22(5):1860–1879, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1350067988.

Bouchard:2013:NAS

[BT13] Bruno Bouchard and Erik Taflin. No-arbitrage of second kind in countable markets with proportional transaction costs. Annals of Applied Probability, 23(2):427–454, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1360682019.

Burdzy:2017:COM

[BT17] Krzysztof Burdzy and Tvrtko Tadić. Can one make a laser out of cardboard? Annals of Applied Probability, 27(4):1951–1991, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1504080023.

Berger:2019:ECL

[BT19] Quentin Berger and Niccolò Torri. Entropy-controlled lastpassage percolation. Annals of Applied Probability, 29(3):1878– 1903, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1550566845.

Bandyopadhyay:2022:NAP

[BT22a] Antar Bandyopadhyay and Debleena Thacker. A new approach to Pólya urn schemes and its infinite color generalization. Annals of Applied Probability, 32(1):46–79, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/A-new-
approach-to-P%c3%b3lya-urn-schemes-and-its-infinite/
10.1214/21-AAP1671.full.

Bouchard:2022:UDF

[BT22b] Bruno Bouchard and Xiaolu Tan. Understanding the dual formulation for the hedging of path-dependent options with price impact. Annals of Applied Probability, 32(3):1705–1733, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Understanding-
the-dual-formulation-for-the-hedging-of-path-dependent/
10.1214/21-AAP1719.full.

Bouchard:2004:DFU

[BTZ04] B. Bouchard, N. Touzi, and A. Zeghal. Dual formulation of the utility maximization problem: The case of nonsmooth utility. Annals of Applied Probability, 14(2):678–717, May 2004. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1082737107.

Bjornberg:2018:CPR

[BU18] Jakob E. Björnberg and Daniel Ueltschi. Critical parameter of random loop model on trees. Annals of Applied Probability, 28 (4):2063–2082, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1533780267.

Budhiraja:2002:IMD

[Bud02] A. Budhiraja. On invariant measures of discrete time filters in the correlated signal-noise case. Annals of Applied Probability, 12(3):1096–1113, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1031863182.

Buraczewski:2007:IMS

[Bur07] Dariusz Buraczewski. On invariant measures of stochastic recursions in a critical case. Annals of Applied Probability, 17 (4):1245–1272, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1186755239.

Butkovsky:2014:SRC

[But14] Oleg Butkovsky. Subgeometric rates of convergence of Markov processes in the Wasserstein metric. Annals of Applied Probability, 24(2):526–552, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1394465364.

Betz:2011:RPC

[BUV11] Volker Betz, Daniel Ueltschi, and Yvan Velenik. Random permutations with cycle weights. Annals of Applied Probability, 21 (1):312–331, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1292598036.

Bhamidi:2012:WDA

[BvdH12] Shankar Bhamidi and Remco van der Hofstad. Weak disorder asymptotics in the stochastic mean-field model of distance. Annals of Applied Probability, 22(1):29–69, February 2012. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1328623695.

Bhamidi:2010:FPP

[BvdHH10] Shankar Bhamidi, Remco van der Hofstad, and Gerard Hooghiemstra. First passage percolation on random graphs with finite mean degrees. Annals of Applied Probability, 20 (5):1907–1965, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1282747405. See erratum [BvdHH17].

Bhamidi:2017:EFP

[BvdHH17] Shankar Bhamidi, Remco van der Hofstad, and Gerard Hooghiemstra. Erratum: “First passage percolation on random graphs with finite mean degrees” [Ann. Appl. Probab. 20

(5) (2010) 1907–1965]. Annals of Applied Probability, 27(5): 3246–3253, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1509696046. See [BvdHH10].

Backhoff-Veraguas:2020:NSS

[BVLT20] Julio Backhoff-Veraguas, Daniel Lacker, and Ludovic Tangpi. Nonexponential Sanov and Schilder theorems on Wiener space: BSDEs, Schrödinger problems and control. Annals of Applied Probability, 30(3):1321–1367, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-3/Nonexponential-Sanov-and-Schilder-
theorems-on-Wiener-space--BSDEs/10.1214/19-AAP1531.full.

Backhoff-Veraguas:2022:SMO

[BVP22] J. Backhoff-Veraguas and G. Pammer. Stability of martingale optimal transport and weak optimal transport. Annals of Applied Probability, 32(1):721–752, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-1/Stability-of-martingale-optimal-transport-
and-weak-optimal-transport/10.1214/21-AAP1694.full.

Bell:2001:DSS

[BW01] S. L. Bell and R. J. Williams. Dynamic scheduling of a system with two parallel servers in heavy traffic with resource pooling: asymptotic optimality of a threshold policy. Annals of Applied Probability, 11(3):608–649, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015345343.

Bai:2022:CRT

[BW22a] Tianyi Bai and Yijun Wan. Capacity of the range of tree-indexed random walk. Annals of Applied Probability, 32(3):1557–1589, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Capacity-

of-the-range-of-tree-indexed-random-walk/10.1214/21-
AAP1689.full.

Baxendale:2022:RCF

[BW22b] Peter Baxendale and Ting-Kam Leonard Wong. Random concave functions. Annals of Applied Probability, 32(2):812–852, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Random-
concave-functions/10.1214/21-AAP1697.full.

Baryshnikov:2005:GLR

[BY05] Yu. Baryshnikov and J. E. Yukich. Gaussian limits for random measures in geometric probability. Annals of Applied Probability, 15(1A):213–253, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1106922327.

Bayraktar:2017:RDG

[BY17] Erhan Bayraktar and Song Yao. On the robust Dynkin game. Annals of Applied Probability, 27(3):1702–1755, June 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1500451239.

Badowski:2000:APS

[BYZ00] G. Badowski, G. Yin, and Q. Zhang. Asymptotic properties of a singularly perturbed Markov chain with inclusion of transient states. Annals of Applied Probability, 10(2):549–572, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1019487355.

Browne:1991:IMC

[BZ91] Sidney Browne and Paul Zipkin. Inventory models with continuous, stochastic demands. Annals of Applied Probability, 1 (3):419–435, August 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005875.

Bender:2008:TDM

[BZ08] Christian Bender and Jianfeng Zhang. Time discretization and Markovian iteration for coupled FBSDEs. Annals of Applied Probability, 18(1):143–177, February 2008. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1199890019.

Barlow:2010:RCM

[BZ10] Martin T. Barlow and Xinghua Zheng. The random conductance model with Cauchy tails. Annals of Applied Probability, 20(3):869–889, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1276867300.

Bertacchi:2015:BRW

[BZ15] Daniela Bertacchi and Fabio Zucca. Branching random walks and multi-type contact-processes on the percolation cluster of Z^{d}. Annals of Applied Probability, 25(4):1993–2012, 08 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Bayraktar:2016:AHU

[BZ16] Erhan Bayraktar and Zhou Zhou. Arbitrage, hedging and utility maximization using semi-static trading strategies with American options. Annals of Applied Probability, 26(6):3531–3558, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1481792592.

Berestycki:2018:SMB

[BZ18] Nathanaël Berestycki and Lee Zhuo Zhao. The shape of multidimensional Brunet–Derrida particle systems. Annals of Applied Probability, 28(2):651–687, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1523433619.

Biagini:2019:RFF

[BZ19] Francesca Biagini and Yinglin Zhang. Reduced-form framework under model uncertainty. Annals of Applied Probability, 29(4):2481–2522, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1563869048.

Cheek:2018:MFB

[CA18] David Cheek and Tibor Antal. Mutation frequencies in a birth– death branching process. Annals of Applied Probability, 28(6): 3922–3947, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1538985639.

Calvin:1997:APC

[Cal97] James M. Calvin. Average performance of a class of adaptive algorithms for global optimization. Annals of Applied Probability, 7(3):711–730, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034801250.

Can:2019:ALT

[Can19] Van Hao Can. Annealed limit theorems for the Ising model on random regular graphs. Annals of Applied Probability, 29 (3):1398–1445, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1550566834.

Cao:2021:CLT

[Cao21] Sky Cao. Central limit theorems for combinatorial optimization problems on sparse Erdős–Rényi graphs. Annals of Applied Probability, 31(4):1687–1723, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-4/Central-limit-theorems-for-combinatorial-
optimization-problems-on-sparse-Erd%c5%91sR%c3%a9nyi/
10.1214/20-AAP1630.full.

Cohen-Alloro:2020:REE

[CAP20] Omri Cohen-Alloro and Ron Peled. Rarity of extremal edges in random surfaces and other theoretical applications of cluster algorithms. Annals of Applied Probability, 30(5):2439–2464, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-5/Rarity-
of-extremal-edges-in-random-surfaces-and-other-theoretical/
10.1214/20-AAP1562.full.

Chen:2018:ETC

[CBM18] Yue Chen, Ana Bušić, and Sean Meyn. Ergodic theory for controlled Markov chains with stationary inputs. Annals of Applied Probability, 28(1):79–111, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1520046084.

Castiel:2021:IIL

[CBM^{+}21] Eyal Castiel, Sem Borst, Laurent Miclo, Florian Simatos, and Phil Whiting. Induced idleness leads to deterministic heavy traffic limits for queue-based random-access algorithms. Annals of Applied Probability, 31(2):941–971, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/Induced-
idleness-leads-to-deterministic-heavy-traffic-limits-
for-queue/10.1214/20-AAP1609.full.

Capinski:1993:EEU

[CC93] Marek Capinski and Nigel J. Cutland. The Euler equation: a uniform nonstandard construction of a global flow, invariant measures and statistical solutions. Annals of Applied Probability, 3(1):212–227, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005516.

Catoni:1998:PCT

[CC98a] Olivier Catoni and Cécile Cot. Piecewise constant triangular cooling schedules for generalized simulated annealing algorithms. Annals of Applied Probability, 8(2):375–396, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1028903532.

Chiang:1998:PEC

[CC98b] Tzuu-Shuh Chiang and Yunshyong Chow. On the problem of exit from cycles for simulated annealing processes — a backward equation approach. Annals of Applied Probability, 8(3):896–916, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903456.

Capinski:1999:SEE

[CC99] Marek Capiński and Nigel J. Cutland. Stochastic Euler equations on the torus. Annals of Applied Probability, 9(3):688–705, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962809.

Chassagneux:2014:RKS

[CC14] Jean-François Chassagneux and Dan Crisan. Runge–Kutta schemes for backward stochastic differential equations. Annals of Applied Probability, 24(2):679–720, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1394465369.

Choukroun:2016:BSR

[CC16] Sébastien Choukroun and Andrea Cosso. Backward SDE representation for stochastic control problems with nondominated controlled intensity. Annals of Applied Probability, 26(2):1208– 1259, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1458651831.

Chen:2011:EUS

[CCCS11] Xinfu Chen, Lan Cheng, John Chadam, and David Saunders. Existence and uniqueness of solutions to the inverse boundary crossing problem for diffusions. Annals of Applied Probability, 21(5):1663–1693, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576606.

Chassagneux:2019:NMF

[CCD19] Jean-François Chassagneux, Dan Crisan, and François Delarue. Numerical method for FBSDEs of McKean–Vlasov type. Annals of Applied Probability, 29(3):1640–1684, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1550566839.

Cavazos-Cadena:2005:COR

[CCHH05] Rolando Cavazos-Cadena and Daniel Hernández-Hernández. A characterization of the optimal risk-sensitive average cost in finite controlled Markov chains. Annals of Applied Probability, 15 (1A):175–212, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1106922326.

Chernozhukov:2023:NOC

[CCK23] Victor Chernozhukov, Denis Chetverikov, and Yuta Koike. Nearly optimal central limit theorem and bootstrap approximations in high dimensions. Annals of Applied Probability, 33(3):

2374–2425, June 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
3/Nearly-optimal-central-limit-theorem-and-bootstrap-
approximations-in-high/10.1214/22-AAP1870.full.

Collevecchio:2013:PAG

[CCL13] Andrea Collevecchio, Codina Cotar, and Marco LiCalzi. On a preferential attachment and generalized Pólya’s urn model. Annals of Applied Probability, 23(3):1219–1253, June 2013. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1362684859.

Clausel:2016:SEI

[CCL16] Marianne Clausel, Jean-François Coeurjolly, and Jérôme Lelong. Stein estimation of the intensity of a spatial homogeneous Poisson point process. Annals of Applied Probability, 26(3):1495– 1534, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1465905010.

Cancrini:2006:RTR

[CCM06] N. Cancrini, P. Caputo, and F. Martinelli. Relaxation time of L-reversal chains and other chromosome shuffles. Annals of Applied Probability, 16(3):1506–1527, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1159804989.

Chenavier:2023:BDD

[CCR23] Nicolas Chenavier, David Coupier, and Arnaud Rousselle. The bi-dimensional Directed IDLA forest. Annals of Applied Probability, 33(3):2247–2290, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-3/The-bi-dimensional-Directed-IDLA-forest/
10.1214/22-AAP1865.full.

Coutin:1999:ANF

[CD99] L. Coutin and L. Decreusefond. Abstract nonlinear filtering theory in the presence of fractional Brownian motion. Annals of Applied Probability, 9(4):1058–1090, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962865.

Cox:2002:SSM

[CD02] J. Theodore Cox and Richard Durrett. The stepping stone model: New formulas expose old myths. Annals of Applied Probability, 12(4):1348–1377, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1037125866. See correction

[CD03].

Cox:2003:CSS

[CD03] Theodore Cox and Richard Durrett. Correction: “The stepping stone model: new formulas expose old myths”. Annals of Applied Probability, 13(2):816, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1050689605. See [CD02].

Chan:2006:NCR

[CD06] Benjamin Chan and Richard Durrett. A new coexistence result for competing contact processes. Annals of Applied Probability, 16(3):1155–1165, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1159804978.

Chatterjee:2011:ABA

[CD11] Shirshendu Chatterjee and Rick Durrett. Asymptotic behavior of Aldous’ gossip process. Annals of Applied Probability, 21(6): 2447–2482, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1322057327.

Cetin:2016:MNE

[ÇD16] Umut Çetin and Albina Danilova. Markovian Nash equilibrium in financial markets with asymmetric information and related forward–backward systems. Annals of Applied Probability, 26 (4):1996–2029, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1472745450.

Chatterjee:2018:SSR

[CD18] Sourav Chatterjee and Persi Diaconis. The sample size required in importance sampling. Annals of Applied Probability, 28(2):1099–1135, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1523433632.

Costantini:2017:WFC

[CDE^{+}17] Cristina Costantini, Pierpaolo De Blasi, Stewart N. Ethier, Matteo Ruggiero, and Dario Spanò. Wright–Fisher construction of the two-parameter Poisson–Dirichlet diffusion. Annals of Applied Probability, 27(3):1923–1950, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1500451247.

Carmona:2013:SFB

[CDET13] René Carmona, François Delarue, Gilles-Edouard Espinosa, and Nizar Touzi. Singular forward–backward stochastic differential equations and emissions derivatives. Annals of Applied Probability, 23(3):1086–1128, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1362684855.

Coghi:2020:PMV

[CDFM20] Michele Coghi, Jean-Dominique Deuschel, Peter K. Friz, and Mario Maurelli. Pathwise McKean–Vlasov theory with additive noise. Annals of Applied Probability, 30(5):2355–2392, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-5/Pathwise-
McKeanVlasov-theory-with-additive-noise/10.1214/20-AAP1560.
full.

Crisan:2013:RFC

[CDFO13] D. Crisan, J. Diehl, P. K. Friz, and H. Oberhauser. Robust filtering: Correlated noise and multidimensional observation. Annals of Applied Probability, 23(5):2139–2160, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1377696308.

Chen:2023:WPW

[CDG23] Yong Chen, Jinqiao Duan, and Hongjun Gao. Well-posedness and wave-breaking for the stochastic rotation-two-component Camassa–Holm system. Annals of Applied Probability, 33(4): 2734–2785, August 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-

4/Well-posedness-and-wave-breaking-for-the-stochastic-
rotation-two/10.1214/22-AAP1877.full.

Campi:2022:MFG

[CDGL22] Luciano Campi, Tiziano De Angelis, Maddalena Ghio, and Giulia Livieri. Mean-field games of finite-fuel capacity expansion with singular controls. Annals of Applied Probability, 32(5): 3674–3717, October 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
5/Mean-field-games-of-finite-fuel-capacity-expansion-
with-singular/10.1214/21-AAP1771.full.

Chauvin:2001:PBS

[CDJH01] Brigitte Chauvin, Michael Drmota, and Jean Jabbour-Hattab. The profile of binary search trees. Annals of Applied Probability, 11(4):1042–1062, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015345394.

Chan:2009:CMC

[CDL09] B. Chan, R. Durrett, and N. Lanchier. Coexistence for a multitype contact process with seasons. Annals of Applied Probability, 19(5):1921–1943, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1255699548.

Chetrite:2017:NPW

[CDL17] Raphael Chetrite, Roland Diel, and Matthieu Lerasle. The number of potential winners in Bradley–Terry model in random environment. Annals of Applied Probability, 27(3):1372–1394, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1500451225.

Crudu:2012:CSG

[CDMR12] A. Crudu, A. Debussche, A. Muller, and O. Radulescu. Convergence of stochastic gene networks to hybrid piecewise deterministic processes. Annals of Applied Probability, 22(5):1822–1859, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1350067987.

Camia:2002:CRT

[CDN02] Federico Camia, Emilio De Santis, and Charles M. Newman. Clusters and recurrence in the two-dimensional zero-temperature stochastic Ising model. Annals of Applied Probability, 12(2):565– 580, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1026915616.

Cattiaux:2018:SCS

[CDP18] Patrick Cattiaux, Fanny Delebecque, and Laure Pédèches. Stochastic Cucker–Smale models: Old and new. Annals of Applied Probability, 28(5):3239–3286, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1535443248.

Clarkson:2009:RBT

[CDS09] Eric Clarkson, J. L. Denny, and Larry Shepp. ROC and the bounds on tail probabilities via theorems of Dubins and F. Riesz. Annals of Applied Probability, 19(1):467–476, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1235140346.

Chatterjee:2011:RGG

[CDS11] Sourav Chatterjee, Persi Diaconis, and Allan Sly. Random graphs with a given degree sequence. Annals of Applied Probability, 21(4):1400–1435, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1312818840.

Cerny:2023:BTF

[CDS23] Jirí Cerný, Alexander Drewitz, and Lars Schmitz. (Un-)bounded transition fronts for the parabolic Anderson model and the randomized f-KPP equation. Annals of Applied Probability, 33(3): 2342–2373, June 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
3/Un-bounded-transition-fronts-for-the-parabolic-Anderson-
model-and/10.1214/22-AAP1869.full.

Collamore:2014:RES

[CDV14] Jeffrey F. Collamore, Guoqing Diao, and Anand N. Vidyashankar. Rare event simulation for processes generated via stochastic fixed point equations. Annals of Applied Probability, 24(5):2143–2175, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1403812372.

Cohen:2010:CBS

[CE10] Samuel N. Cohen and Robert J. Elliott. Comparisons for backward stochastic differential equations on Markov chains and related no-arbitrage conditions. Annals of Applied Probability, 20 (1):267–311, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1262962324.

Chassagneux:2012:DTA

[CEK12] Jean-Francois Chassagneux, Romuald Elie, and Idris Kharroubi. Discrete-time approximation of multidimensional BSDEs with oblique reflections. Annals of Applied Probability, 22(3):971– 1007, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1337347536.

Cerf:1996:NGA

[Cer96] Raphaël Cerf. A new genetic algorithm. Annals of Applied Probability, 6(3):778–817, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1034968228.

Cerrai:2009:KTA

[Cer09] Sandra Cerrai. A Khasminskii type averaging principle for stochastic reaction–diffusion equations. Annals of Applied Probability, 19(3):899–948, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1245071014.

Cerf:2015:CPE

[Cer15] Raphaël Cerf. Critical population and error threshold on the sharp peak landscape for the Wright–Fisher model. Annals of Applied Probability, 25(4):1936–1992, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1432212433.

Cipolloni:2023:FCL

[CES23] Giorgio Cipolloni, László Erdős, and Dominik Schröder. Functional central limit theorems for Wigner matrices. Annals of Applied Probability, 33(1):447–489, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-1/Functional-central-limit-theorems-
for-Wigner-matrices/10.1214/22-AAP1820.full.

Cetin:2018:DTB

[Çet18] Umut Çetin. Diffusion transformations, Black–Scholes equation and optimal stopping. Annals of Applied Probability, 28 (5):3102–3151, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1535443244.

Clouet:1994:SPT

[CF94] J. F. Clouet and J. P. Fouque. Spreading of a pulse travelling in random media. Annals of Applied Probability, 4(4):1083–1097, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004904.

Caputo:2007:IIM

[CF07] Pietro Caputo and Alessandra Faggionato. Isoperimetric inequalities and mixing time for a random walk on a random point process. Annals of Applied Probability, 17(5/6):1707–1744, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1191419181.

Caputo:2009:DOD

[CF09] P. Caputo and A. Faggionato. Diffusivity in one-dimensional generalized Mott variable-range hopping models. Annals of Applied Probability, 19(4):1459–1494, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1248700624.

Como:2011:SLC

[CF11] Giacomo Como and Fabio Fagnani. Scaling limits for continuous opinion dynamics systems. Annals of Applied Probability, 21 (4):1537–1567, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1312818844.

Coghi:2016:PCI

[CF16a] Michele Coghi and Franco Flandoli. Propagation of chaos for interacting particles subject to environmental noise. Annals of Applied Probability, 26(3):1407–1442, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905007.

Cortez:2016:QPC

[CF16b] Roberto Cortez and Joaquin Fontbona. Quantitative propagation of chaos for generalized Kac particle systems. Annals of Applied Probability, 26(2):892–916, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1458651823.

Chen:2017:HLP

[CF17] Zhen-Qing Chen and Wai-Tong (Louis) Fan. Hydrodynamic limits and propagation of chaos for interacting random walks in domains. Annals of Applied Probability, 27(3):1299–1371, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1500451224.

Campi:2018:PGM

[CF18] Luciano Campi and Markus Fischer. N-player games and meanfield games with absorption. Annals of Applied Probability, 28 (4):2188–2242, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1533780271.

Comets:2002:PLM

[CFF02] Francis Comets, Roberto Fernández, and Pablo A. Ferrari. Processes with long memory: Regenerative construction and perfect simulation. Annals of Applied Probability, 12(3):921–943, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1031863175.

Coffman:2000:IPV

[CFJ00] E. G. Coffman, Jr., Leopold Flatto, and Predrag Jelenković. Interval packing: the vacant interval distribution. Annals of Applied Probability, 10(1):240–257, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1019737671.

Confortola:2016:BSD

[CFJ16] Fulvia Confortola, Marco Fuhrman, and Jean Jacod. Backward stochastic differential equation driven by a marked point process: An elementary approach with an application to optimal control. Annals of Applied Probability, 26(3):1743–1773, June 2016. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1465905018.

Cuchiero:2011:APP

[CFMT11] Christa Cuchiero, Damir Filipović, Eberhard Mayerhofer, and Josef Teichmann. Affine processes on positive semidefinite matrices. Annals of Applied Probability, 21(2):397–463, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1300800978.

Chhita:2018:LDK

[CFS18] S. Chhita, P. L. Ferrari, and H. Spohn. Limit distributions for KPZ growth models with spatially homogeneous random initial conditions. Annals of Applied Probability, 28(3):1573–1603, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1527840027.

Cheridito:2005:EAC

[CFY05] Patrick Cheridito, Damir Filipović, and Marc Yor. Equivalent and absolutely continuous measure changes for jump-diffusion processes. Annals of Applied Probability, 15(3):1713–1732, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1121433766.

Comets:1992:PEG

[CG92] Francis Comets and Basilis Gidas. Parameter estimation for Gibbs distributions from partially observed data. Annals of Applied Probability, 2(1):142–170, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005775.

Cerou:2016:FAA

[CG16] Frédéric Cérou and Arnaud Guyader. Fluctuation analysis of adaptive multilevel splitting. Annals of Applied Probability, 26 (6):3319–3380, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1481792587.

Clement:2017:AKC

[CG17] Emmanuelle Clément and Arnaud Gloter. An application of the KMT construction to the pathwise weak error in the Euler approximation of one-dimensional diffusion process with linear diffusion coefficient. Annals of Applied Probability, 27(4):2419– 2454, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1504080037.

Cox:1993:GLA

[CGGK93] J. Theodore Cox, Alberto Gandolfi, Philip S. Griffin, and Harry Kesten. Greedy lattice animals i: Upper bounds. Annals of Applied Probability, 3(4):1151–1169, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005277.

Cosso:2023:OCP

[CGK^{+}23] Andrea Cosso, Fausto Gozzi, Idris Kharroubi, Huyên Pham, and Mauro Rosestolato. Optimal control of path-dependent McKean–Vlasov SDEs in infinite-dimension. Annals of Applied Probability, 33(4):2863–2918, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-4/Optimal-control-of-path-dependent-
McKeanVlasov-SDEs-in-infinite-dimension/10.1214/22-AAP1880.
full.

Craiu:2015:SAM

[CGL^{+}15] Radu V. Craiu, Lawrence Gray, Krzysztof Latuszyński, Neal Madras, Gareth O. Roberts, and Jeffrey S. Rosenthal. Stability of adversarial Markov chains, with an application to adaptive MCMC algorithms. Annals of Applied Probability, 25(6): 3592–3623, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1443703783.

Cranston:2009:LDR

[CGM09] M. Cranston, D. Gauthier, and T. S. Mountford. On large deviation regimes for random media models. Annals of Applied Probability, 19(2):826–862, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1241702251.

Collet:2009:MSO

[CGR09] P. Collet, C. Giardina, and F. Redig. Matching with shift for one-dimensional Gibbs measures. Annals of Applied Probability, 19(4):1581–1602, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1248700628.

Carinci:2020:EFT

[CGR20] Gioia Carinci, Cristian Giardinà, and Frank Redig. Exact formulas for two interacting particles and applications in particle systems with duality. Annals of Applied Probability, 30(4):1934– 1970, August 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
4/Exact-formulas-for-two-interacting-particles-and-applications-
in-particle/10.1214/19-AAP1548.full.

Coffman:1993:SRP

[CGS93] E. G. Coffman, E. N. Gilbert, and P. W. Shor. A selectionreplacement process on the circle. Annals of Applied Probability, 3(3):802–818, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005365.

Caravenna:2007:RTA

[CGZ07] Francesco Caravenna, Giambattista Giacomin, and Lorenzo Zambotti. A renewal theory approach to periodic copolymers with adsorption. Annals of Applied Probability, 17(4):1362–1398, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1186755243.

Chafai:2014:FOG

[CGZ14] Djalil Chafaï, Nathael Gozlan, and Pierre-André Zitt. Firstorder global asymptotics for confined particles with singular pair repulsion. Annals of Applied Probability, 24(6):2371–2413, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1409058035.

Crametz:1991:LRR

[CH91] J.-P. Crametz and P. J. Hunt. A limit result respecting graph structure for a fully connected loss network with alternative routing. Annals of Applied Probability, 1(3):436–444, August 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005876.

Champagnat:2019:PAD

[CH19] Nicolas Champagnat and Benoit Henry. A probabilistic approach to Dirac concentration in nonlocal models of adaptation with several resources. Annals of Applied Probability, 29 (4):2175–2216, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1563869040.

Chang:1992:MFL

[Cha92] Joseph T. Chang. On moments of the first ladder height of random walks with small drift. Annals of Applied Probability, 2(3):714–738, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005656.

Chassaing:1993:OMF

[Cha93] Philippe Chassaing. Optimality of move-to-front for selforganizing data structures with locality of references. Annals of Applied Probability, 3(4):1219–1240, November 1993. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005280.

Chang:1994:IO

[Cha94] Joseph T. Chang. Inequalities for the overshoot. Annals of Applied Probability, 4(4):1223–1233, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177004913.

Chassaing:1997:DMB

[Cha97] Philippe Chassaing. Determining the majority: the biased case. Annals of Applied Probability, 7(2):523–544, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1034625343.

Chan:1999:PCC

[Cha99] Terence Chan. Pricing contingent claims on stocks driven by Lévy processes. Annals of Applied Probability, 9(2):504–528, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962753.

Chan:2005:STG

[Cha05] Hock Peng Chan. Summation test for gap penalties and strong law of the local alignment score. Annals of Applied Probability, 15(2):1492–1505, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1115137983.

Chen:1995:FAS

[Che95] Hong Chen. Fluid approximations and stability of multiclass queueing networks: Work-conserving disciplines. Annals of Applied Probability, 5(3):637–665, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004699.

Chen:1996:SCP

[Che96] Hong Chen. A sufficient condition for the positive recurrence of a semimartingale reflecting Brownian motion in an orthant. Annals of Applied Probability, 6(3):758–765, August 1996. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1034968226.

Chen:2001:ECR

[Che01] Xia Chen. Exact convergence rates for the distribution of particles in branching random walks. Annals of Applied Probability, 11(4):1242–1262, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015345402.

Cherny:2008:BMA

[Che08] Alexander Cherny. Brownian moving averages have conditional full support. Annals of Applied Probability, 18(5):1825–1830, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1225372951.

Chen:2013:SBC

[Che13] Yu-Ting Chen. Sharp benefit-to-cost rules for the evolution of cooperation on regular graphs. Annals of Applied Probability, 23(2):637–664, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1360682025.

Chen:2018:WFD

[Che18] Yu-Ting Chen. Wright–Fisher diffusions in stochastic spatial evolutionary games with death–birth updating. Annals of Applied Probability, 28(6):3418–3490, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1538985626.

Chen:2022:HJE

[Che22a] Hong-Bin Chen. Hamilton–Jacobi equations for nonsymmetric matrix inference. Annals of Applied Probability, 32(4):2540– 2567, August 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
4/HamiltonJacobi-equations-for-nonsymmetric-matrix-inference/
10.1214/21-AAP1739.full.

Chen:2022:CRW

[Che22b] Yu-Ting Chen. Convergences of the rescaled Whittaker stochastic differential equations and independent sums. Annals of Applied Probability, 32(4):2914–2966, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-4/Convergences-of-the-rescaled-Whittaker-
stochastic-differential-equations-and-independent/10.
1214/21-AAP1753.full.

Chi:2004:LLE

[Chi04] Zhiyi Chi. Limit laws of estimators for critical multi-type Galton–Watson processes. Annals of Applied Probability, 14(4): 1992–2015, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674086.

Chi:2005:LDT

[Chi05] Zhiyi Chi. Large deviations for template matching between point processes. Annals of Applied Probability, 15(1A):153–174, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1106922325.

Chi:2007:UCE

[Chi07] Zhiyi Chi. Uniform convergence of exact large deviations for renewal reward processes. Annals of Applied Probability, 17 (3):1019–1048, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1179839181.

Chi:2015:SRT

[Chi15] Zhiyi Chi. Strong renewal theorems with infinite mean beyond local large deviations. Annals of Applied Probability, 25(3):1513– 1539, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1427124135.

Chi:2016:RRM

[Chi16] Zhiyi Chi. Random reversible Markov matrices with tunable extremal eigenvalues. Annals of Applied Probability, 26(4):2257– 2272, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1472745458.

Chen:2021:PTRa

[CHL21] Wei-Kuo Chen, Madeline Handschy, and Gilad Lerman. Phase transition in random tensors with multiple independent spikes. Annals of Applied Probability, 31(4):1868–1913, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/Phase-
transition-in-random-tensors-with-multiple-independent-
spikes/10.1214/20-AAP1636.full.

Caye:2020:TSN

[CHMK20] Thomas Cayé, Martin Herdegen, and Johannes Muhle-Karbe. Trading with small nonlinear price impact. Annals of Applied Probability, 30(2):706–746, April 2020. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1591603220.

Chow:2002:SWE

[Cho02] Pao-Liu Chow. Stochastic wave equations with polynomial nonlinearity. Annals of Applied Probability, 12(1):361–381, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015961168.

Chow:2006:ASS

[Cho06] Pao-Liu Chow. Asymptotics of solutions to semilinear stochastic wave equations. Annals of Applied Probability, 16(2):757–789, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1151592250.

Cox:2008:PIL

[CHO08] A. M. G. Cox, David Hobson, and Jan Obłój. Pathwise inequalities for local time: Applications to Skorokhod embeddings and optimal stopping. Annals of Applied Probability, 18 (5):1870–1896, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1225372954.

Chow:2009:NSW

[Cho09] Pao-Liu Chow. Nonlinear stochastic wave equations: Blow-up of second moments in L^{2}-norm. Annals of Applied Probability, 19 (6):2039–2046, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1259158765.

Chen:2017:CRM

[CHS17] Guan-Yu Chen, Jui-Ming Hsu, and Yuan-Chung Sheu. The L^{2}- cutoffs for reversible Markov chains. Annals of Applied Probability, 27(4):2305–2341, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1504080034.

Cordero:2022:GSM

[CHS22] Fernando Cordero, Sebastian Hummel, and Emmanuel Schertzer. General selection models: Bernstein duality and minimal ancestral structures. Annals of Applied Probability, 32(3):1499–1556, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/General-
selection-models-Bernstein-duality-and-minimal-ancestral-
structures/10.1214/21-AAP1683.full.

Constantin:2011:SLA

[CI11] Peter Constantin and Gautam Iyer. A stochastic-Lagrangian approach to the Navier–Stokes equations in domains with boundary. Annals of Applied Probability, 21(4):1466–1492, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1312818842.

Ciucu:1998:NFC

[Ciu98] Mihai Ciucu. No-feedback card guessing for dovetail shuffles. Annals of Applied Probability, 8(4):1251–1269, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1028903379.

Cohn:1994:GBP

[CJ94] Harry Cohn and Peter Jagers. General branching processes in varying environment. Annals of Applied Probability, 4(1):184– 193, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005206.

Conus:2019:WCR

[CJK19] Daniel Conus, Arnulf Jentzen, and Ryan Kurniawan. Weak convergence rates of spectral Galerkin approximations for SPDEs with nonlinear diffusion coefficients. Annals of Applied Probability, 29(2):653–716, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1548298927.

Cai:2021:AAE

[CJL21] T. Tony Cai, Tiefeng Jiang, and Xiaoou Li. Asymptotic analysis for extreme eigenvalues of principal minors of random matrices. Annals of Applied Probability, 31(6):2953–2990, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-6/Asymptotic-
analysis-for-extreme-eigenvalues-of-principal-minors-
of-random/10.1214/21-AAP1668.full.

Cetin:2004:MCR

[ÇJPY04] Umut Çetin, Robert Jarrow, Philip Protter, and Yıldıray Yıldırım. Modeling credit risk with partial information. Annals of Applied Probability, 14(3):1167–1178, August 2004. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1089736281.

Chhita:2015:ADS

[CJY15] Sunil Chhita, Kurt Johansson, and Benjamin Young. Asymptotic domino statistics in the Aztec diamond. Annals of Applied Probability, 25(3):1232–1278, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1427124128.

Cvitanic:1992:CDC

[CK92] Jaksa Cvitanic and Ioannis Karatzas. Convex duality in constrained portfolio optimization. Annals of Applied Probability, 2(4):767–818, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005576.

Cvitanic:1993:HCC

[CK93] Jaksa Cvitanic and Ioannis Karatzas. Hedging contingent claims with constrained portfolios. Annals of Applied Probability, 3 (3):652–681, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005357.

Chen:2000:SSM

[CK00] Chingfer Chen and Samuel Karlin. r-scan statistics of a marker array in multiple sequences derived from a common progenitor. Annals of Applied Probability, 10(3):709–725, August 2000. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1019487507.

Cox:2003:RID

[CK03] J. Theodore Cox and Achim Klenke. Rescaled interacting diffusions converge to super Brownian motion. Annals of Applied Probability, 13(2):501–514, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1050689591.

Connor:2007:PSCa

[CK07a] Stephen B. Connor and Wilfrid S. Kendall. Perfect simulation for a class of positive recurrent Markov chains. Annals of Applied Probability, 17(3):781–808, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1179839174.

Connor:2007:PSCb

[CK07b] Stephen B. Connor and Wilfrid S. Kendall. Perfect simulation for a class of positive recurrent Markov chains. Annals of Applied Probability, 17(5/6):1808–1810, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1191419184.

Cotar:2012:ERG

[CK12] Codina Cotar and Christof Külske. Existence of random gradient states. Annals of Applied Probability, 22(4):1650–1692, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1344614207.

Cox:2019:RHO

[CK19] Alexander M. G. Cox and Sam M. Kinsley. Robust hedging of options on a leveraged exchange traded fund. Annals of Applied Probability, 29(1):531–576, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1544000436.

Clement:2006:DAW

[CKHL06] Emmanuelle Clément, Arturo Kohatsu-Higa, and Damien Lamberton. A duality approach for the weak approximation of stochastic differential equations. Annals of Applied Probability, 16(3):1124–1154, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1159804977.

Christensen:2022:AOE

[CKL22] Sören Christensen, Jan Kallsen, and Matthias Lenga. Are American options European after all? Annals of Applied Probability, 32(2):853–892, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/

issue-2/Are-American-options-European-after-all/10.1214/
21-AAP1698.full.

Chen:2021:RCM

[CKW21] Xin Chen, Takashi Kumagai, and Jian Wang. Random conductance models with stable-like jumps: Quenched invariance principle. Annals of Applied Probability, 31(3):1180–1231, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-3/Random-
conductance-models-with-stable-like-jumps--Quenched-
invariance/10.1214/20-AAP1616.full.

Chan:2003:SAN

[CL03] Hock Peng Chan and Tze Leung Lai. Saddlepoint approximations and nonlinear boundary crossing probabilities of Markov random walks. Annals of Applied Probability, 13(2):395–429, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1050689586.

Chigansky:2004:SNF

[CL04] Pavel Chigansky and Robert Liptser. Stability of nonlinear filters in nonmixing case. Annals of Applied Probability, 14(4): 2038–2056, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674088.

Champagnat:2007:EDP

[CL07a] Nicolas Champagnat and Amaury Lambert. Evolution of discrete populations and the canonical diffusion of adaptive dynamics. Annals of Applied Probability, 17(1):102–155, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1171377179.

Chan:2007:EIS

[CL07b] Hock Peng Chan and Tze Leung Lai. Efficient importance sampling for Monte Carlo evaluation of exceedance probabilities. Annals of Applied Probability, 17(2):440–473, April 2007. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1174323253.

Cotar:2009:ATS

[CL09a] Codina Cotar and Vlada Limic. Attraction time for strongly reinforced walks. Annals of Applied Probability, 19(5):1972–2007, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1255699550.

Craddock:2009:CEC

[CL09b] Mark Craddock and Kelly A. Lennox. The calculation of expectations for classes of diffusion processes by Lie symmetry methods. Annals of Applied Probability, 19(1):127–157, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1235140335.

Chan:2011:SMC

[CL11] Hock Peng Chan and Tze Leung Lai. A sequential Monte Carlo approach to computing tail probabilities in stochastic models. Annals of Applied Probability, 21(6):2315–2342, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1322057323.

Collet:2014:LMH

[CL14] Pierre Collet and Florencia Leonardi. Loss of memory of hidden Markov models and Lyapunov exponents. Annals of Applied Probability, 24(1):422–446, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278730.

Carmona:2015:PWF

[CL15] René Carmona and Daniel Lacker. A probabilistic weak formulation of mean field games and applications. Annals of Applied Probability, 25(3):1189–1231, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1427124127.

Clement:2017:HOE

[CL17] Dombry Clément and Rabehasaina Landy. High order expansions for renewal functions and applications to ruin theory. Annals of Applied Probability, 27(4):2342–2382, August 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1504080035.

Carmona:2022:CAM

[CL22] René Carmona and Mathieu Laurière. Convergence analysis of machine learning algorithms for the numerical solution of mean field control and games: II-the finite horizon case. Annals of Applied Probability, 32(6):4065–4105, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/Convergence-
analysis-of-machine-learning-algorithms-for-the-numerical-
solution/10.1214/21-AAP1715.full.

Clancy:1996:SAM

[Cla96] Damian Clancy. Strong approximations for mobile population epidemic models. Annals of Applied Probability, 6(3):883–895, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968231.

Claisse:2018:OCB

[Cla18] Julien Claisse. Optimal control of branching diffusion processes: A finite horizon problem. Annals of Applied Probability, 28(1):1– 34, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1520046082.

Clark:2022:CMD

[Cla22] Jeremy Thane Clark. Continuum models of directed polymers on disordered diamond fractals in the critical case. Annals of Applied Probability, 32(6):4186–4250, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/Continuum-
models-of-directed-polymers-on-disordered-diamond-fractals-
in/10.1214/22-AAP1783.full.

Clement:2023:HTV

[Clé23] Emmanuelle Clément. Hellinger and total variation distance in approximating Lévy driven SDEs. Annals of Applied Probability, 33(3):2176–2209, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-3/Hellinger-and-total-variation-distance-in-approximating-
L%c3%a9vy-driven-SDEs/10.1214/22-AAP1863.full.

Cattiaux:2016:ESD

[CLP16] Patrick Cattiaux, José R. León, and Clémentine Prieur. Estimation for stochastic damping Hamiltonian systems under partial observation. III. Diffusion term. Annals of Applied Probability, 26(3):1581–1619, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1465905012.

Cvitanic:2006:FAT

[CLR06] Jakša Cvitanić, Robert Liptser, and Boris Rozovskii. A filtering approach to tracking volatility from prices observed at random times. Annals of Applied Probability, 16(3):1633–1652, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804994.

Cuchiero:2018:PJD

[CLSF18] Christa Cuchiero, Martin Larsson, and Sara Svaluto-Ferro. Polynomial jump-diffusions on the unit simplex. Annals of Applied Probability, 28(4):2451–2500, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1533780278.

Choudhury:1994:MTI

[CLW94] Gagan L. Choudhury, David M. Lucantoni, and Ward Whitt. Multidimensional transform inversion with applications to the transient M/G/1 queue. Annals of Applied Probability, 4(3):719– 740, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004968.

Corwin:2016:FTL

[CLW16] Ivan Corwin, Zhipeng Liu, and Dong Wang. Fluctuations of TASEP and LPP with general initial data. Annals of Applied Probability, 26(4):2030–2082, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1472745451.

Cvitanic:1996:HOL

[CM96] Jakša Cvitanić and Jin Ma. Hedging options for a large investor and forward-backward SDE’s. Annals of Applied Probability, 6(2):370–398, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968136.

Caputo:2003:RTA

[CM03] Pietro Caputo and Fabio Martinelli. Relaxation time of anisotropic simple exclusion processes and quantum Heisenberg models. Annals of Applied Probability, 13(2):691–721, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1050689600.

Christophi:2005:ODD

[CM05] Costas A. Christophi and Hosam M. Mahmoud. The oscillatory distribution of distances in random tries. Annals of Applied Probability, 15(2):1536–1564, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1115137985.

Crepey:2008:RDR

[CM08] Stéphane Crépey and Anis Matoussi. Reflected and doubly reflected BSDEs with jumps: A priori estimates and comparison. Annals of Applied Probability, 18(5):2041–2069, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1225372960.

Crisan:2014:SOD

[CM14] Dan Crisan and Konstantinos Manolarakis. Second order discretization of backward SDEs and simulation with the cubature method. Annals of Applied Probability, 24(2):652–678, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1394465368.

Crisan:2019:CWS

[CM19] Dan Crisan and Eamon McMurray. Cubature on Wiener space for McKean–Vlasov SDEs with smooth scalar interaction. Annals of Applied Probability, 29(1):130–177, February 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1544000427.

Charlier:2023:GFP

[CM23] Christophe Charlier and Philippe Moreillon. On the generating function of the Pearcey process. Annals of Applied Probability, 33(4):3240–3277, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-4/On-the-generating-function-of-the-Pearcey-process/
10.1214/22-AAP1890.full.

Cline:2004:SLE

[CmHP04] Daren B. H. Cline and Huay min H. Pu. Stability and the Lyapounov exponent of threshold AR–ARCH models. Annals of Applied Probability, 14(4):1920–1949, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1099674083.

Coffman:1994:PAC

[CMP94] E. G. Coffman, C. L. Mallows, and Bjorn Poonen. Parking arcs on the circle with applications to one-dimensional communication networks. Annals of Applied Probability, 4(4):1098–1111, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004905.

Caputo:2015:RLT

[CMSS15] Pietro Caputo, Fabio Martinelli, Alistair Sinclair, and Alexandre Stauffer. Random lattice triangulations: Structure and algorithms. Annals of Applied Probability, 25(3):1650–1685, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1427124139.

Champagnat:2021:SAE

[CMT21] Nicolas Champagnat, Sylvie Méléard, and Viet Chi Tran. Stochastic analysis of emergence of evolutionary cyclic behavior in population dynamics with transfer. Annals of Applied Probability, 31(4):1820–1867, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-4/Stochastic-analysis-of-emergence-of-
evolutionary-cyclic-behavior-in-population/10.1214/20-
AAP1635.full.

Chassaing:2003:SQO

[CMY03] P. Chassaing, J. F. Marckert, and M. Yor. A stochastically quasioptimal search algorithm for the maximum of the simple random walk. Annals of Applied Probability, 13(4):1264–1295, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1069786499.

Collins:2011:GES

[CN11] Benoît Collins and Ion Nechita. Gaussianization and eigenvalue statistics for random quantum channels (III). Annals of Applied Probability, 21(3):1136–1179, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1307020394.

Cosco:2023:VFL

[CN23] Clément Cosco and Shuta Nakajima. A variational formula for large deviations in first-passage percolation under tail estimates. Annals of Applied Probability, 33(3):2103–2135, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/A-variational-
formula-for-large-deviations-in-first-passage-percolation/
10.1214/22-AAP1861.full.

Collevecchio:2022:VRJ

[CNV22] Andrea Collevecchio, Tuan-Minh Nguyen, and Stanislav Volkov. Vertex-reinforced jump process on the integers with nonlinear reinforcement. Annals of Applied Probability, 32(4):2671–2705, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Vertex-
reinforced-jump-process-on-the-integers-with-nonlinear-
reinforcement/10.1214/21-AAP1743.full.

Cohn:1996:APS

[Coh96] Harry Cohn. On the asymptotic patterns of supercritical branching processes in varying environments. Annals of Applied Probability, 6(3):896–902, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968232.

Cohort:2004:LTR

[Coh04] Pierre Cohort. Limit theorems for random normalized distortion. Annals of Applied Probability, 14(1):118–143, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1075828049.

Collamore:2002:IST

[Col02] J. F. Collamore. Importance sampling techniques for the multidimensional ruin problem for general Markov additive sequences of random vectors. Annals of Applied Probability, 12(1):382–421, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015961169.

Collamore:2009:RRE

[Col09] Jeffrey F. Collamore. Random recurrence equations and ruin in a Markov-dependent stochastic economic environment. Annals of Applied Probability, 19(4):1404–1458, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1248700623.

Comets:1997:DPT

[Com97] Francis Comets. Detecting phase transition for Gibbs measures. Annals of Applied Probability, 7(2):545–563, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1034625344.

Coja-Oghlan:2022:BPR

[COMR22] Amin Coja-Oghlan, Noela Müller, and Jean B. Ravelomanana. Belief propagation on the random k-SAT model. Annals of Applied Probability, 32(5):3718–3796, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Belief-propagation-on-the-random-k-
SAT-model/10.1214/21-AAP1772.full.

Conforti:2022:PAC

[Con22] Giovanni Conforti. A probabilistic approach to convex. Annals of Applied Probability, 32(2):932–973, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-2/A-probabilistic-approach-to-convex-
%cf%95-entropy-decay-for-Markov/10.1214/21-AAP1700.full.

Coppini:2022:LTD

[Cop22] Fabio Coppini. Long time dynamics for interacting oscillators on graphs. Annals of Applied Probability, 32(1):360–391, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Long-
time-dynamics-for-interacting-oscillators-on-graphs/
10.1214/21-AAP1680.full.

Costa:2016:PDM

[Cos16] Manon Costa. A piecewise deterministic model for a preypredator community. Annals of Applied Probability, 26(6): 3491–3530, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1481792591.

Coupier:2008:TSC

[Cou08] David Coupier. Two sufficient conditions for Poisson approximations in the ferromagnetic Ising model. Annals of Applied Probability, 18(4):1326–1350, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1216677124.

Coupier:2010:GLC

[Cou10] David Coupier. Geography of local configurations. Annals of Applied Probability, 20(3):806–840, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1276867298.

Cox:2010:IRM

[Cox10] J. Theodore Cox. Intermediate range migration in the twodimensional stepping stone model. Annals of Applied Probability, 20(3):785–805, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1276867297.

Cox:2008:RTD

[CP08] J. Theodore Cox and Edwin A. Perkins. Renormalization of the two-dimensional Lotka–Volterra model. Annals of Applied Probability, 18(2):747–812, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1206018203.

Caravenna:2009:PMI

[CP09] Francesco Caravenna and Nicolas Pétrélis. A polymer in a multi-interface medium. Annals of Applied Probability, 19(5): 1803–1839, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1255699544.

Cox:2014:CCT

[CP14] J. Theodore Cox and Edwin A. Perkins. A complete convergence theorem for voter model perturbations. Annals of Applied Probability, 24(1):150–197, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278723.

Chen:2018:DCS

[CP18] Wei-Kuo Chen and Dmitry Panchenko. Disorder chaos in some diluted spin glass models. Annals of Applied Probability, 28 (3):1356–1378, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1527840021.

Carmona:2019:STS

[CP19a] Philippe Carmona and Nicolas Pétrélis. A shape theorem for the scaling limit of the IPDSAW at criticality. Annals of Applied Probability, 29(2):875–930, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1548298933.

Cerrai:2019:LDF

[CP19b] Sandra Cerrai and Nicholas Paskal. Large deviations for fast transport stochastic RDEs with applications to the exit problem. Annals of Applied Probability, 29(4):1993–2032, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1563869035.

Coffman:1995:PSZ

[CPR95] E. G. Coffman, A. A. Puhalskii, and M. I. Reiman. Polling systems with zero switchover times: A heavy-traffic averaging principle. Annals of Applied Probability, 5(3):681–719, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177004701.

Chauvin:2011:LDL

[CPS11] Brigitte Chauvin, Nicolas Pouyanne, and Reda Sahnoun. Limit distributions for large Pólya urns. Annals of Applied Probability, 21(1):1–32, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1292598027.

Cox:2016:CNV

[CPS16] J. Theodore Cox, Yuval Peres, and Jeffrey E. Steif. Cutoff for the noisy voter model. Annals of Applied Probability, 26(2):917–932, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1458651824.

Casanova:2022:SCW

[CPSJ22] Adrián González Casanova, Verónica Miró Pina, and Arno Siri-Jégousse. The symmetric coalescent and Wright–Fisher models with bottlenecks. Annals of Applied Probability, 32(1):235–268, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/The-
symmetric-coalescent-and-WrightFisher-models-with-bottlenecks/
10.1214/21-AAP1676.full.

Cisse:2012:OSP

[CPT12] Mamadou Cissé, Pierre Patie, and Etienne Tanré. Optimal stopping problems for some Markov processes. Annals of Applied Probability, 22(3):1243–1265, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1337347544.

Caragea:2023:NNA

[CPV23] Andrei Caragea, Philipp Petersen, and Felix Voigtlaender. Neural network approximation and estimation of classifiers with classification boundary in a Barron class. Annals of Applied Probability, 33(4):3039–3079, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-4/Neural-network-approximation-and-estimation-
of-classifiers-with-classification-boundary/10.1214/
22-AAP1884.full.

Chiu:1997:CLT

[CQ97] S. N. Chiu and M. P. Quine. Central limit theory for the number of seeds in a growth model in R^{d} with inhomogeneous Poisson arrivals. Annals of Applied Probability, 7(3):802–814, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1034801254.

Chassagneux:2016:NSQ

[CR16] Jean-François Chassagneux and Adrien Richou. Numerical simulation of quadratic BSDEs. Annals of Applied Probability, 26 (1):262–304, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1452003239.

Chenavier:2018:CSD

[CR18] Nicolas Chenavier and Christian Y. Robert. Cluster size distributions of extreme values for the Poisson–Voronoi tessellation. Annals of Applied Probability, 28(6):3291–3323, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1538985623.

Crawford:2022:ECC

[CR22] Nicholas Crawford and Ron Rosenthal. Eigenvector correlations in the complex Ginibre ensemble. Annals of Applied Probability, 32(4):2706–2754, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-4/Eigenvector-correlations-in-the-complex-Ginibre-
ensemble/10.1214/21-AAP1746.full.

Crane:2016:DRN

[Cra16] Harry Crane. Dynamic random networks and their graph limits. Annals of Applied Probability, 26(2):691–721, April 2016. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1458651817.

Crane:2018:CLP

[Cra18] Harry Crane. Combinatorial Lévy processes. Annals of Applied Probability, 28(1):285–339, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1520046089.

Cuchiero:2023:PMS

[CRSF23] Christa Cuchiero, Stefan Rigger, and Sara Svaluto-Ferro. Propagation of minimality in the supercooled Stefan problem. Annals of Applied Probability, 33(2):1388–1418, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Propagation-

of-minimality-in-the-supercooled-Stefan-problem/10.1214/
22-AAP1850.full.

Cai:2017:ALB

[CRT17] Jiatu Cai, Mathieu Rosenbaum, and Peter Tankov. Asymptotic lower bounds for optimal tracking: a linear programming approach. Annals of Applied Probability, 27(4):2455–2514, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1504080038.

Claisse:2023:MFG

[CRT23] Julien Claisse, Zhenjie Ren, and Xiaolu Tan. Mean field games with branching. Annals of Applied Probability, 33(2):834–875, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Mean-
field-games-with-branching/10.1214/22-AAP1835.full.

Chazottes:2006:APM

[CRV06] Jean-Rene Chazottes, Frank Redig, and Evgeny Verbitskiy. On approximate pattern matching for a class of Gibbs random fields. Annals of Applied Probability, 16(2):670–684, May 2006. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1151592247.

Chen:2021:PTRb

[CRX21] Louis H. Y. Chen, Adrian Röllin, and Aihua Xia. Palm theory, random measures and Stein couplings. Annals of Applied Probability, 31(6):2881–2923, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-6/Palm-theory-random-measures-and-Stein-
couplings/10.1214/21-AAP1666.full.

Csorgo:1995:PCD

[CS95] Sandor Csorgo and Gordon Simons. Precision calculation of distributions for trimmed sums. Annals of Applied Probability, 5(3):854–873, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004708.

Cox:1999:SSP

[CS99a] J. Theodore Cox and Rinaldo B. Schinazi. A stochastic spatial process to model the persistence of sickle-cell disease. Annals of Applied Probability, 9(2):319–330, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962744.

Cvitanic:1999:MPP

[CS99b] Jakša Cvitanić and Gennady Spivak. Maximizing the probability of a perfect hedge. Annals of Applied Probability, 9(4): 1303–1328, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962873.

Chayes:2000:MPB

[CS00a] Lincoln Chayes and Roberto H. Schonmann. Mixed percolation as a bridge between site and bond percolation. Annals of Applied Probability, 10(4):1182–1196, November 2000. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1019487612.

Chen:2000:SAM

[CS00b] Hong Chen and Xinyang Shen. Strong approximations for multiclass feedforward queueing networks. Annals of Applied Probability, 10(3):828–876, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487511.

Cranston:2002:DRU

[CS02] Mike Cranston and Michael Scheutzow. Dispersion rates under finite mode Kolmogorov flows analysis. Annals of Applied Probability, 12(2):511–532, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1026915613.

Cohen:2006:RRF

[CS06] Serge Cohen and Gennady Samorodnitsky. Random rewards, fractional Brownian local times and stable self-similar processes. Annals of Applied Probability, 16(3):1432–1461, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1159804987.

Chatterjee:2011:NAS

[CS11a] Sourav Chatterjee and Qi-Man Shao. Nonnormal approximation by Stein’s method of exchangeable pairs with application to the Curie–Weiss model. Annals of Applied Probability, 21(2):464– 483, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1300800979.

Collevecchio:2011:BSR

[CS11b] Andrea Collevecchio and Tom Schmitz. Bounds on the speed and on regeneration times for certain processes on regular trees. Annals of Applied Probability, 21(3):1073–1101, June 2011. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1307020392.

Choi:2013:LDD

[CS13a] Jihyeok Choi and Sunder Sethuraman. Large deviations for the degree structure in preferential attachment schemes. Annals of Applied Probability, 23(2):722–763, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1360682028.

Czichowsky:2013:CCC

[CS13b] Christoph Czichowsky and Martin Schweizer. Cone-constrained continuous-time Markowitz problems. Annals of Applied Probability, 23(2):764–810, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1360682029.

Cantarella:2016:SGC

[CS16a] Jason Cantarella and Clayton Shonkwiler. The symplectic geometry of closed equilateral random walks in 3-space. Annals of Applied Probability, 26(1):549–596, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1452003247.

Czichowsky:2016:DTP

[CS16b] Christoph Czichowsky and Walter Schachermayer. Duality theory for portfolio optimisation under transaction costs. Annals of Applied Probability, 26(3):1888–1941, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905022.

Chatterjee:2017:MST

[CS17a] Sourav Chatterjee and Sanchayan Sen. Minimal spanning trees and Stein’s method. Annals of Applied Probability, 27(3):1588– 1645, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1500451236.

Czichowsky:2017:POB

[CS17b] Christoph Czichowsky and Walter Schachermayer. Portfolio optimisation beyond semimartingales: Shadow prices and fractional Brownian motion. Annals of Applied Probability, 27 (3):1414–1451, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1500451227.

Casanova:2018:DFW

[CS18] Adrián González Casanova and Dario Spanò. Duality and fixation in Ξ–Wright–Fisher processes with frequency-dependent selection. Annals of Applied Probability, 28(1):250–284, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1520046088.

Chleboun:2021:CSP

[CS21] Paul Chleboun and Aaron Smith. Cutoff for the square plaquette model on a critical length scale. Annals of Applied Probability, 31 (2):668–702, April 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
2/Cutoff-for-the-square-plaquette-model-on-a-critical-
length/10.1214/20-AAP1601.full.

Cardaliaguet:2022:FOM

[CS22] Pierre Cardaliaguet and Panagiotis E. Souganidis. On first order mean field game systems with a common noise. Annals of Applied Probability, 32(3):2289–2326, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-3/On-first-order-mean-field-game-systems-
with-a-common/10.1214/21-AAP1734.full.

Christensen:2023:NLI

[CS23] Sören Christensen and Claudia Strauch. Nonparametric learning for impulse control problems-exploration vs. exploitation. Annals of Applied Probability, 33(2):1369–1387, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Nonparametric-
learning-for-impulse-control-problemsExploration-vs-
exploitation/10.1214/22-AAP1849.full.

Chang:1998:TQN

[CSS98] Kuo-Hwa Chang, Richard F. Serfozo, and Władysław Szczotka. Treelike queueing networks: asymptotic stationarity and heavy traffic. Annals of Applied Probability, 8(2):541–568, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1028903538.

Curato:2022:CLT

[CSS22] Imma Valentina Curato, Robert Stelzer, and Bennet Ströh. Central limit theorems for stationary random fields under weak dependence with application to ambit and mixed moving average fields. Annals of Applied Probability, 32(3):1814–1861, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Central-
limit-theorems-for-stationary-random-fields-under-weak-
dependence/10.1214/21-AAP1722.full.

Cheridito:2005:STP

[CST05] Patrick Cheridito, H. Mete Soner, and Nizar Touzi. Small time path behavior of double stochastic integrals and applications to stochastic control. Annals of Applied Probability, 15 (4):2472–2495, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965769.

Chassagneux:2022:WQP

[CST22] Jean-François Chassagneux, Lukasz Szpruch, and Alvin Tse. Weak quantitative propagation of chaos via differential calculus on the space of measures. Annals of Applied Probability, 32(3): 1929–1969, June 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
3/Weak-quantitative-propagation-of-chaos-via-differential-
calculus-on-the/10.1214/21-AAP1725.full.

Chen:2023:PAF

[CSX23] Peng Chen, Qi-Man Shao, and Lihu Xu. A probability approximation framework: Markov process approach. Annals of Applied Probability, 33(2):1419–1459, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-2/A-probability-approximation-framework-
Markov-process-approach/10.1214/22-AAP1853.full.

Caravenna:2017:UMR

[CSZ17] Francesco Caravenna, Rongfeng Sun, and Nikos Zygouras. Universality in marginally relevant disordered systems. Annals of Applied Probability, 27(5):3050–3112, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1509696041.

Chen:1991:ETP

[CT91] Rong Chen and Ruey S. Tsay. On the ergodicity of tar(1) processes. Annals of Applied Probability, 1(4):613–634, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005841.

Corcoran:2001:PSE

[CT01] J. N. Corcoran and R. L. Tweedie. Perfect sampling of ergodic Harris chains. Annals of Applied Probability, 11(2):438–451, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345299.

Carmona:2004:CHP

[CT04] Rene Carmona and Michael Tehranchi. A characterization of hedging portfolios for interest rate contingent claims. Annals of Applied Probability, 14(3):1267–1294, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1089736285.

Cerf:2011:ULD

[CT11] Raphaël Cerf and Marie Théret. Upper large deviations for the maximal flow through a domain of R ^{d} in first passage percolation. Annals of Applied Probability, 21(6):2075–2108, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1322057316.

Cerny:2016:RWT

[ČT16] Jiří Černý and Augusto Teixeira. Random walks on torus and random interlacements: Macroscopic coupling and phase transition. Annals of Applied Probability, 26(5):2883–2914, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1476884307.

Chen:2023:OBN

[CT23] Louis H. Y. Chen and Lê Văn Thành. Optimal bounds in normal approximation for many interacting worlds. Annals of Applied Probability, 33(2):625–642, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-2/Optimal-bounds-in-normal-approximation-
for-many-interacting-worlds/10.1214/21-AAP1747.full.

Choulli:2004:IBD

[CTZ04] Tahir Choulli, Michael Taksar, and Xun Yu Zhou. Interplay between dividend rate and business constraints for a financial corporation. Annals of Applied Probability, 14(4):1810–1837, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1099674079.

Caillerie:2021:DAK

[CV21] Nils Caillerie and Julien Vovelle. Diffusion-approximation for a kinetic equation with perturbed velocity redistribution process. Annals of Applied Probability, 31(3):1299–1335, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-3/Diffusion-
approximation-for-a-kinetic-equation-with-perturbed-
velocity-redistribution/10.1214/20-AAP1619.full.

Chigansky:2007:MRF

[CvH07] Pavel Chigansky and Ramon van Handel. Model robustness of finite state nonlinear filtering over the infinite time horizon. Annals of Applied Probability, 17(2):688–715, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1174323261.

Chigansky:2010:CSB

[CvH10] Pavel Chigansky and Ramon van Handel. A complete solution to Blackwell’s unique ergodicity problem for hidden Markov chains. Annals of Applied Probability, 20(6):2318–2345, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1287494562.

Clifford:1993:ECP

[CW93] Peter Clifford and Gang Wei. The equivalence of the Cox process with squared radial Ornstein–Uhlenbeck intensity and the death process in a simple population model. Annals of Applied Probability, 3(3):863–873, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005368.

Cao:1999:GCF

[CW99] Jin Cao and Keith Worsley. The geometry of correlation fields with an application to functional connectivity of the brain. Annals of Applied Probability, 9(4):1021–1057, November 1999. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1029962864.

Chen:2003:ACM

[CW03a] Mu-Fa Chen and Ying-Zhe Wang. Algebraic convergence of Markov chains. Annals of Applied Probability, 13(2):604–627, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1050689596.

Cohn:2003:MBL

[CW03b] Harry Cohn and Qiao Wang. Multitype branching limit behavior. Annals of Applied Probability, 13(2):490–500, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1050689590.

Cox:2013:RBC

[CW13] Alexander M. G. Cox and Jiajie Wang. Root’s barrier: Construction, optimality and applications to variance options. Annals of Applied Probability, 23(3):859–894, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1362684848.

Cappelletti:2016:EIS

[CW16] Daniele Cappelletti and Carsten Wiuf. Elimination of intermediate species in multiscale stochastic reaction networks. Annals of Applied Probability, 26(5):2915–2958, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1476884308.

Cruise:2019:CGS

[CW19] James R. Cruise and Andrew R. Wade. The critical greedy server on the integers is recurrent. Annals of Applied Probability, 29(2):1233–1261, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1548298940.

Carmona:1997:HTD

[CX97] René A. Carmona and Lin Xu. Homogenization for timedependent two-dimensional incompressible Gaussian flows. Annals of Applied Probability, 7(1):265–279, February 1997. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1034625262.

Choi:2002:ANS

[CX02] K. P. Choi and Aihua Xia. Approximating the number of successes in independent trials: Binomial versus Poisson. Annals of Applied Probability, 12(4):1139–1148, November 2002. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1037125856.

Cheng:2016:MEC

[CX16] Dan Cheng and Yimin Xiao. The mean Euler characteristic and excursion probability of Gaussian random fields with stationary increments. Annals of Applied Probability, 26(2):722–759, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1458651818.

Calka:2021:CHP

[CY21] Pierre Calka and J. E. Yukich. Convex hulls of perturbed random point sets. Annals of Applied Probability, 31(4):1598–1632, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/Convex-
hulls-of-perturbed-random-point-sets/10.1214/20-AAP1627.
full.

Chow:2003:LDF

[CZ03] Yunshyong Chow and Yu Zhang. Large deviations in firstpassage percolation. Annals of Applied Probability, 13(4): 1601–1614, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1069786513.

Carmona:2016:PAM

[CZ16] René Carmona and Xiuneng Zhu. A probabilistic approach to mean field games with major and minor players. Annals of Applied Probability, 26(3):1535–1580, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1465905011.

Chatterjee:2018:TDA

[CZ18] Shirshendu Chatterjee and Ofer Zeitouni. Thresholds for detecting an anomalous path from noisy environments. Annals of Applied Probability, 28(5):2635–2663, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1535443230.

Dai:1995:PHR

[Dai95] J. G. Dai. On positive Harris recurrence of multiclass queueing networks: a unified approach via fluid limit models. Annals of Applied Probability, 5(1):49–77, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004828.

Dai:1996:FLM

[Dai96] J. G. Dai. A fluid limit model criterion for instability of multiclass queueing networks. Annals of Applied Probability, 6(3):751– 757, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1034968225.

Doku-Amponsah:2010:LDP

[DAM10] Kwabena Doku-Amponsah and Peter Mörters. Large deviation principles for empirical measures of colored random graphs. Annals of Applied Probability, 20(6):1989–2021, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1287494552.

Dario:2021:OCE

[Dar21] Paul Dario. Optimal corrector estimates on percolation cluster. Annals of Applied Probability, 31(1):377–431, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/Optimal-
corrector-estimates-on-percolation-cluster/10.1214/20-
AAP1593.full.

Dario:2023:CTL

[Dar23] Paul Dario. Convergence to the thermodynamic limit for random-field random surfaces. Annals of Applied Probability, 33(2):1173–1195, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-2/Convergence-to-the-thermodynamic-limit-for-random-
field-random-surfaces/10.1214/22-AAP1844.full.

Dassios:1995:DQB

[Das95] Angelos Dassios. The distribution of the quantile of a Brownian motion with drift and the pricing of related path-dependent options. Annals of Applied Probability, 5(2):389–398, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177004770.

Dassios:1996:SQS

[Das96] Angelos Dassios. Sample quantiles of stochastic processes with stationary and independent increments. Annals of Applied Probability, 6(3):1041–1043, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968241.

deBouard:2012:DAT

[dBG12] A. de Bouard and M. Gazeau. A diffusion approximation theorem for a nonlinear PDE with application to random birefringent optical fibers. Annals of Applied Probability, 22(6):2460–2504, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1353695959.

Delcoigne:2004:LDP

[Dd04] Franck Delcoigne and Arnaud de La Fortelle. Large deviations problems for star networks: The min policy. Annals of Applied Probability, 14(2):1006–1028, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1082737120.

DAngeli:2009:CPM

[DD09] Daniele D’Angeli and Alfredo Donno. Crested products of Markov chains. Annals of Applied Probability, 19(1):414–453, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1235140344.

DelMoral:2010:IMC

[DD10] Pierre Del Moral and Arnaud Doucet. Interacting Markov chain Monte Carlo methods for solving nonlinear measure-valued equations. Annals of Applied Probability, 20(2):593–639, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1268143434.

Dommers:2017:MGD

[DdHJN17] S. Dommers, F. den Hollander, O. Jovanovski, and F. R. Nardi. Metastability for Glauber dynamics on random graphs. Annals of Applied Probability, 27(4):2130–2158, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1504080028.

Daskalakis:2011:CER

[DDM11] Constantinos Daskalakis, Alexandros G. Dimakis, and Elchanan Mossel. Connectivity and equilibrium in random games. Annals of Applied Probability, 21(3):987–1016, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1307020389.

Decreusefond:2012:LGL

[DDMT12] Laurent Decreusefond, Jean-Stéphane Dhersin, Pascal Moyal, and Viet Chi Tran. Large graph limit for an SIR process in random network with heterogeneous connectivity. Annals of Applied Probability, 22(2):541–575, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1333372007.

Delmas:2008:ARL

[DDSJ08] Jean-François Delmas, Jean-Stéphane Dhersin, and Arno Siri-Jegousse. Asymptotic results on the length of coalescent trees. Annals of Applied Probability, 18(3):997–1025, June 2008. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1211819792.

DeMarco:2011:SAE

[De 11] Stefano De Marco. Smoothness and asymptotic estimates of densities for SDEs with locally smooth coefficients and applications to square root-type diffusions. Annals of Applied Probability, 21 (4):1282–1321, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1312818837.

DeAngelis:2017:DPF

[DE17] Tiziano De Angelis and Erik Ekström. The dividend problem with a finite horizon. Annals of Applied Probability, 27(6): 3525–3546, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1513328707.

DeBlassie:2004:IBM

[DeB04] R. Dante DeBlassie. Iterated Brownian motion in an open set. Annals of Applied Probability, 14(3):1529–1558, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1089736295.

Decreusefond:1998:PAM

[Dec98] L. Decreusefond. Perturbation analysis and Malliavin calculus. Annals of Applied Probability, 8(2):496–523, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1028903536.

Djehiche:2023:MFR

[DEH23] Boualem Djehiche, Romuald Elie, and Said Hamadène. Mean-field reflected backward stochastic differential equations. Annals of Applied Probability, 33(4):2493–2518, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-4/Mean-
field-reflected-backward-stochastic-differential-equations/
10.1214/20-AAP1657.full.

DelMoral:1998:MVP

[Del98] P. Del Moral. Measure-valued processes and interacting particle systems. application to nonlinear filtering problems. Annals of Applied Probability, 8(2):438–495, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1028903535.

Dereich:2011:MMC

[Der11] Steffen Dereich. Multilevel Monte Carlo algorithms for Lévydriven SDEs with Gaussian correction. Annals of Applied Probability, 21(1):283–311, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1292598035.

Devroye:1992:LTR

[Dev92a] Luc Devroye. A limit theory for random skip lists. Annals of Applied Probability, 2(3):597–609, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005651.

Devroye:1992:STL

[Dev92b] Luc Devroye. A study of trie-like structures under the density model. Annals of Applied Probability, 2(2):402–434, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005709.

Dobrow:1995:MCM

[DF95a] Robert P. Dobrow and James Allen Fill. On the Markov chain for the move-to-root rule for binary search trees. Annals of Applied Probability, 5(1):1–19, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004824.

Dobrow:1995:RCM

[DF95b] Robert P. Dobrow and James Allen Fill. Rates of convergence for the move-to-root Markov chain for binary search trees. Annals of Applied Probability, 5(1):20–36, February 1995. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004825.

Dawson:2006:ABP

[DF06] Donald A. Dawson and Shui Feng. Asymptotic behavior of the Poisson–Dirichlet distribution for large mutation rate. Annals of Applied Probability, 16(2):562–582, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1151592243.

Durrett:2016:GEP

[DF16] Rick Durrett and Wai-Tong (Louis) Fan. Genealogies in expanding populations. Annals of Applied Probability, 26(6):3456–3490, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1481792590.

Dolera:2020:BET

[DF20] Emanuele Dolera and Stefano Favaro. A Berry–Esseen theorem for Pitman’s α-diversity. Annals of Applied Probability, 30(2):847–869, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1591603224.

Drees:2004:MLE

[DFdH04] Holger Drees, Ana Ferreira, and Laurens de Haan. On maximum likelihood estimation of the extreme value index. Annals of Applied Probability, 14(3):1179–1201, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1089736282.

Dianetti:2021:SMF

[DFFN21] Jodi Dianetti, Giorgio Ferrari, Markus Fischer, and Max Nendel. Submodular mean field games: Existence and approximation of solutions. Annals of Applied Probability, 31(6):2538–2566, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-6/Submodular-
mean-field-games-Existence-and-approximation-of-solutions/
10.1214/20-AAP1655.full.

Duerinckx:2022:SLH

[DFG22] Mitia Duerinckx, Julian Fischer, and Antoine Gloria. Scaling limit of the homogenization commutator for Gaussian coefficient fields. Annals of Applied Probability, 32(2):1179–1209, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Scaling-
limit-of-the-homogenization-commutator-for-Gaussian-
coefficient-fields/10.1214/21-AAP1705.full.

Diaconis:2013:ACS

[DFH13] Persi Diaconis, Jason Fulman, and Susan Holmes. Analysis of casino shelf shuffling machines. Annals of Applied Probability, 23(4):1692–1720, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1371834042.

Delattre:2016:HPL

[DFH16] Sylvain Delattre, Nicolas Fournier, and Marc Hoffmann. Hawkes processes on large networks. Annals of Applied Probability, 26 (1):216–261, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1452003238.

Denardo:1997:SMS

[DFK97] Eric V. Denardo, Eugene A. Feinberg, and Offer Kella. Stochastic monotonicity for stationary recurrence times of first passage heights. Annals of Applied Probability, 7(2):326–339, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1034625334.

Denisov:2012:LTR

[DFK12] D. Denisov, S. Foss, and T. Konstantopoulos. Limit theorems for a random directed slab graph. Annals of Applied Probability, 22(2):702–733, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1333372011.

Diehl:2016:PSL

[DFM16] Joscha Diehl, Peter Friz, and Hilmar Mai. Pathwise stability of likelihood estimators for diffusions via rough paths. Annals of Applied Probability, 26(4):2169–2192, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1472745455.

DeAngelis:2018:NET

[DFM18] Tiziano De Angelis, Giorgio Ferrari, and John Moriarty. Nash equilibria of threshold type for two-player nonzero-sum games of stopping. Annals of Applied Probability, 28(1):112–147, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1520046085.

Douc:2004:PDC

[DFMS04] Randal Douc, Gersende Fort, Eric Moulines, and Philippe Soulier. Practical drift conditions for subgeometric rates of convergence. Annals of Applied Probability, 14(3):1353–1377, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1089736288.

Dyer:1993:APG

[DFP93] Martin Dyer, Alan Frieze, and Boris Pittel. The average performance of the greedy matching algorithm. Annals of Applied Probability, 3(2):526–552, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005436.

Disanto:2022:DUT

[DFPR22] Filippo Disanto, Michael Fuchs, Ariel R. Paningbatan, and Noah A. Rosenberg. The distributions under two species-tree models of the number of root ancestral configurations for matching gene trees and species trees. Annals of Applied Probability, 32(6):4426–4458, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-6/The-distributions-under-two-species-tree-models-
of-the-number/10.1214/22-AAP1791.full.

Duffie:2003:APA

[DFS03] D. Duffie, D. Filipović, and W. Schachermayer. Affine processes and applications in finance. Annals of Applied Probability, 13 (3):984–1053, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1060202833.

Detemple:2003:VAC

[DFT03] Jerome Detemple, Shui Feng, and Weidong Tian. The valuation of American call options on the minimum of two dividendpaying assets. Annals of Applied Probability, 13(3):953–983, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1060202832.

Duffie:1995:EMC

[DG95] Darrell Duffie and Peter Glynn. Efficient Monte Carlo simulation of security prices. Annals of Applied Probability, 5(4):897–905, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004598.

DelMoral:1999:CLT

[DG99] P. Del Moral and A. Guionnet. Central limit theorem for nonlinear filtering and interacting particle systems. Annals of Applied Probability, 9(2):275–297, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1029962742.

DelMoral:2005:GPA

[DG05] Pierre Del Moral and Josselin Garnier. Genealogical particle analysis of rare events. Annals of Applied Probability, 15(4): 2496–2534, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965770.

Dennert:2007:REI

[DG07] Florian Dennert and Rudolf Grübel. Renewals for exponentially increasing lifetimes, with an application to digital search trees. Annals of Applied Probability, 17(2):676–687, April 2007. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1174323260.

Debussche:2008:SNA

[DG08] Arnaud Debussche and Eric Gautier. Small noise asymptotic of the timing jitter in soliton transmission. Annals of Applied Probability, 18(1):178–208, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1199890020.

Dieker:2013:PRP

[DG13] A. B. Dieker and Xuefeng Gao. Positive recurrence of piecewise Ornstein–Uhlenbeck processes and common quadratic Lyapunov functions. Annals of Applied Probability, 23(4):1291–1317, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1371834029.

Dieker:2014:SAD

[DG14] A. B. Dieker and X. Gao. Sensitivity analysis for diffusion processes constrained to an orthant. Annals of Applied Probability, 24(5):1918–1945, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1403812366.

Doney:2018:CER

[DG18] R. A. Doney and Philip S. Griffin. Cramér’s estimate for the reflected process revisited. Annals of Applied Probability, 28(6): 3629–3651, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1538985631.

Du:2021:VEA

[DG21] Qiming Du and Arnaud Guyader. Variance estimation in adaptive sequential Monte Carlo. Annals of Applied Probability, 31(3):1021–1060, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-3/Variance-estimation-in-adaptive-sequential-Monte-
Carlo/10.1214/20-AAP1611.full.

Dean:2020:FLD

[DGC20] Justin Dean, Ayalvadi Ganesh, and Edward Crane. Functional large deviations for Cox processes and. Annals of Applied Probability, 30(5):2465–2490, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-5/Functional-large-deviations-for-Cox-
processes-and-mathitCox-G-infty/10.1214/20-AAP1563.full.

Dyer:2006:SSS

[DGJ06] Martin Dyer, Leslie Ann Goldberg, and Mark Jerrum. Systematic scan for sampling colorings. Annals of Applied Probability, 16(1):185–230, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1141654285.

Dyer:2009:MNR

[DGJ09] Martin Dyer, Leslie Ann Goldberg, and Mark Jerrum. Matrix norms and rapid mixing for spin systems. Annals of Applied Probability, 19(1):71–107, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1235140333.

Damron:2019:PTU

[DGJ^{+}19] Michael Damron, Janko Gravner, Matthew Junge, Hanbaek Lyu, and David Sivakoff. Parking on transitive unimodular graphs. Annals of Applied Probability, 29(4):2089–2113, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1563869038.

Dareiotis:2023:QCT

[DGL23] Konstantinos Dareiotis, Máté Gerencsér, and Khoa Lê. Quantifying a convergence theorem of Gyöngy and Krylov. Annals of Applied Probability, 33(3):2291–2323, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-3/Quantifying-a-convergence-theorem-
of-Gy%c3%b6ngy-and-Krylov/10.1214/22-AAP1867.full.

Douc:2010:FID

[DGLM10] Randal Douc, Elisabeth Gassiat, Benoit Landelle, and Eric Moulines. Forgetting of the initial distribution for nonergodic hidden Markov chains. Annals of Applied Probability, 20(5): 1638–1662, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1282747396.

Dong:2006:CFD

[DGM06] Rui Dong, Christina Goldschmidt, and James B. Martin. Coagulation–fragmentation duality, Poisson–Dirichlet distributions and random recursive trees. Annals of Applied Probability, 16(4):1733–1750, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1169065205.

Draief:2008:TVS

[DGM08] Moez Draief, Ayalvadi Ganesh, and Laurent Massoulié. Thresholds for virus spread on networks. Annals of Applied Probability, 18(2):359–378, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1206018191.

Durmus:2020:GEB

[DGM20] Alain Durmus, Arnaud Guillin, and Pierre Monmarché. Geometric ergodicity of the Bouncy Particle Sampler. Annals of Applied Probability, 30(5):2069–2098, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-5/Geometric-ergodicity-of-the-Bouncy-
Particle-Sampler/10.1214/19-AAP1552.full.

Douc:2011:SMC

[DGMO11] Randal Douc, Aurélien Garivier, Eric Moulines, and Jimmy Olsson. Sequential Monte Carlo smoothing for general state space hidden Markov models. Annals of Applied Probability, 21(6): 2109–2145, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1322057317.

Douc:2005:MDP

[DGN05] R. Douc, A. Guillin, and J. Najim. Moderate deviations for particle filtering. Annals of Applied Probability, 15(1B):587–614, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1107271661.

Dong:2007:EPD

[DGP07] Rui Dong, Alexander Gnedin, and Jim Pitman. Exchangeable partitions derived from Markovian coalescents. Annals of Applied Probability, 17(4):1172–1201, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1186755236.

Depperschmidt:2012:TVF

[DGP12] Andrej Depperschmidt, Andreas Greven, and Peter Pfaffelhuber. Tree-valued Fleming–Viot dynamics with mutation and selection. Annals of Applied Probability, 22(6):2560–2615, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1353695962.

Doring:2019:SEP

[DGPR19] Leif Döring, Lukas Gonon, David J. Prömel, and Oleg Reichmann. On Skorokhod embeddings and Poisson equations. Annals of Applied Probability, 29(4):2302–2337, August 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1563869044.

Dolera:2009:RBP

[DGR09] Emanuele Dolera, Ester Gabetta, and Eugenio Regazzini. Reaching the best possible rate of convergence to equilibrium for solutions of Kac’s equation via central limit theorem. Annals of Applied Probability, 19(1):186–209, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1235140337.

deGunst:1993:NMM

[dGvZ93] Mathisca C. M. de Gunst and Willem R. van Zwet. A non-Markovian model for Cell population growth: Tail behavior and duration of the growth process. Annals of Applied Probability, 3 (4):1112–1144, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005275.

Dai:1991:SSA

[DH91] J. G. Dai and J. M. Harrison. Steady-state analysis of RBM in a rectangle: Numerical methods and a queueing application. Annals of Applied Probability, 1(1):16–35, February 1991. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005979.

Dai:1992:RBM

[DH92] J. G. Dai and J. M. Harrison. Reflected Brownian motion in an orthant: Numerical methods for steady-state analysis. Annals of Applied Probability, 2(1):65–86, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005771.

Dai:1993:QMT

[DH93a] J. G. Dai and J. M. Harrison. The QNET method for twomoment analysis of closed manufacturing systems. Annals of Applied Probability, 3(4):968–1012, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005269.

Duffie:1993:APR

[DH93b] J. Darrell Duffie and J. Michael Harrison. Arbitrage pricing of Russian options and perpetual lookback options. Annals of Applied Probability, 3(3):641–651, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005356.

Dalang:2004:RTS

[DH04] Robert C. Dalang and M.-O. Hongler. The right time to sell a stock whose price is driven by Markovian noise. Annals of Applied Probability, 14(4):2176–2201, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1099674093.

Devroye:2006:WMP

[DH06] Luc Devroye and Hsien-Kuei Hwang. Width and mode of the profile for some random trees of logarithmic height. Annals of Applied Probability, 16(2):886–918, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1151592254.

Deijfen:2007:TTR

[DH07] Maria Deijfen and Olle Häggström. The two-type Richardson model with unbounded initial configurations. Annals of Applied Probability, 17(5/6):1639–1656, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1197909502.

Damron:2013:ENL

[DH13a] Michael Damron and Michael Hochman. Examples of nonpolygonal limit shapes in i.i.d. first-passage percolation and infinite coexistence in spatial growth models. Annals of Applied Probability, 23(3):1074–1085, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1362684854.

Deaconu:2013:HTB

[DH13b] Madalina Deaconu and Samuel Herrmann. Hitting time for Bessel processes—walk on moving spheres algorithm (WoMS).

Annals of Applied Probability, 23(6):2259–2289, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1382447688.

Deaconu:2018:IBV

[DH18] Madalina Deaconu and Samuel Herrmann. Initial-boundary value problem for the heat equation — a stochastic algorithm. Annals of Applied Probability, 28(3):1943–1976, June 2018. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1527840036.

DelMoral:2023:TAO

[DH23] Pierre Del Moral and Emma Horton. A theoretical analysis of one-dimensional discrete generation ensemble Kalman particle filters. Annals of Applied Probability, 33(2):1127–1172, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/A-theoretical-
analysis-of-one-dimensional-discrete-generation-ensemble-
Kalman/10.1214/22-AAP1843.full.

Diaconis:2021:GRE

[DHESC21] Persi Diaconis, Kelsey Houston-Edwards, and Laurent Saloff-Coste. Gambler’s ruin estimates on finite inner uniform domains. Annals of Applied Probability, 31(2):865–895, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/Gamblers-
ruin-estimates-on-finite-inner-uniform-domains/10.1214/
20-AAP1607.full.

Damron:2018:SBF

[DHL18] Michael Damron, Jack Hanson, and Wai-Kit Lam. The size of the boundary in first-passage percolation. Annals of Applied Probability, 28(5):3184–3214, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1535443246.

Damron:2023:UTC

[DHL23] Michael Damron, Jack Hanson, and Wai-Kit Lam. Universality of the time constant for. Annals of Applied Probability, 33(3): 1701–1731, June 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
3/Universality-of-the-time-constant-for-2D-critical-
first-passage/10.1214/22-AAP1808.full.

Diaconis:2000:ANM

[DHN00] Persi Diaconis, Susan Holmes, and Radford M. Neal. Analysis of a nonreversible Markov chain sampler. Annals of Applied Probability, 10(3):726–752, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487508.

Dombry:2018:TMS

[DHS18] Clément Dombry, Enkelejd Hashorva, and Philippe Soulier. Tail measure and spectral tail process of regularly varying time series. Annals of Applied Probability, 28(6):3884–3921, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1538985638.

Dai:2010:MSD

[DHT10] J. G. Dai, Shuangchi He, and Tolga Tezcan. Many-server diffusion limits for G / Ph / n + GI queues. Annals of Applied Probability, 20(5):1854–1890, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1282747403.

Dai:2004:SIT

[DHV04] J. G. Dai, John J. Hasenbein, and John H. Vande Vate. Stability and instability of a two-station queueing network. Annals of Applied Probability, 14(1):326–377, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1075828055.

Delong:2010:BSD

[DI10] Łukasz Delong and Peter Imkeller. Backward stochastic differential equations with time delayed generators—results and counterexamples. Annals of Applied Probability, 20(4):1512–1536, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1279638793.

Dinh:2017:SOD

[DI17] Vu Dinh and Frederick A. Matsen IV. The shape of the onedimensional phylogenetic likelihood function. Annals of Applied Probability, 27(3):1646–1677, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1500451237.

Dieuleveut:2015:VCT

[Die15] Daphné Dieuleveut. The vertex-cut-tree of Galton–Watson trees converging to a stable tree. Annals of Applied Probability, 25 (4):2215–2262, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1432212441.

Dinwoodie:1995:PIO

[Din95] I. H. Dinwoodie. A probability inequality for the occupation measure of a reversible Markov chain. Annals of Applied Probability, 5(1):37–43, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004826.

Delarue:2015:GSN

[DIRT15] François Delarue, James Inglis, Sylvain Rubenthaler, and Etienne Tanré. Global solvability of a networked integrate-and-fire model of McKean–Vlasov type. Annals of Applied Probability, 25(4):2096–2133, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1432212438.

Ding:2010:SDA

[DJ10] Xue Ding and Tiefeng Jiang. Spectral distributions of adjacency and Laplacian matrices of random graphs. Annals of Applied Probability, 20(6):2086–2117, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1287494555.

Decrouez:2012:CMP

[DJ12] Geoffrey Decrouez and Owen Dafydd Jones. A class of multifractal processes constructed using an embedded branching process. Annals of Applied Probability, 22(6):2357–2387, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1353695956.

Ding:2023:LLM

[DJ23] Xiucai Ding and Hong Chang Ji. Local laws for multiplication of random matrices. Annals of Applied Probability, 33(4):2981– 3009, August 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
4/Local-laws-for-multiplication-of-random-matrices/10.
1214/22-AAP1882.full.

Djete:2023:MFG

[Dje23] Mao Fabrice Djete. Mean field games of controls: On the convergence of Nash equilibria. Annals of Applied Probability, 33(4): 2824–2862, August 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
4/Mean-field-games-of-controls--On-the-convergence-of/
10.1214/22-AAP1879.full.

Drmota:2008:FLT

[DJN08] Michael Drmota, Svante Janson, and Ralph Neininger. A functional limit theorem for the profile of search trees. Annals of Applied Probability, 18(1):288–333, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1199890024.

Dembo:1992:PAS

[DK92] Amir Dembo and Samuel Karlin. Poisson approximations for r-scan processes. Annals of Applied Probability, 2(2):329–357, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005707.

Dembo:1999:AWT

[DK99a] Amir Dembo and Ioannis Kontoyiannis. The asymptotics of waiting times between stationary processes, allowing distortion. Annals of Applied Probability, 9(2):413–429, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1029962749.

Donnelly:1999:GPF

[DK99b] Peter Donnelly and Thomas G. Kurtz. Genealogical processes for Fleming–Viot models with selection and recombination. Annals of Applied Probability, 9(4):1091–1148, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1029962866.

Doney:2006:OUL

[DK06] R. A. Doney and A. E. Kyprianou. Overshoots and undershoots of Lévy processes. Annals of Applied Probability, 16(1):91–106, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1141654282.

Delong:2008:OIC

[DK08a] Łukasz Delong and Claudia Klüppelberg. Optimal investment and consumption in a Black–Scholes market with Lévy-driven stochastic coefficients. Annals of Applied Probability, 18(3):879– 908, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1211819788.

Dolinsky:2008:BAS

[DK08b] Yan Dolinsky and Yuri Kifer. Binomial approximations of shortfall risk for game options. Annals of Applied Probability, 18 (5):1737–1770, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1225372948.

Dolinsky:2008:EEB

[DK08c] Yan Dolinsky and Yuri Kifer. Error estimates for binomial approximations of game options. Annals of Applied Probability, 18(3):1271–1277, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1211819801.

Dahmer:2015:IBL

[DK15] Iulia Dahmer and Götz Kersting. The internal branch lengths of the Kingman coalescent. Annals of Applied Probability, 25 (3):1325–1348, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1427124130.

Diehl:2019:TLG

[DK19] Christina S. Diehl and Götz Kersting. Tree lengths for general Λ-coalescents and the asymptotic site frequency spectrum around the Bolthausen–Sznitman coalescent. Annals of Applied Probability, 29(5):2700–2743, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1571385620.

Doring:2017:FSS

[DKM17] Leif Döring, Achim Klenke, and Leonid Mytnik. Finite system scheme for mutually catalytic branching with infinite branching rate. Annals of Applied Probability, 27(5):3113–3152, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1509696042.

Dobler:2022:FCS

[DKP22] Christian Döbler, Mikołaj J. Kasprzak, and Giovanni Peccati. Functional convergence of sequential U-processes with sizedependent kernels. Annals of Applied Probability, 32(1):551–601, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Functional-
convergence-of-sequential-U-processes-with-size-dependent-
kernels/10.1214/21-AAP1688.full.

Dupuis:2020:SAR

[DKPRB20] Paul Dupuis, Markos A. Katsoulakis, Yannis Pantazis, and Luc Rey-Bellet. Sensitivity analysis for rare events based on Rényi divergence. Annals of Applied Probability, 30(4):1507–1533, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/Sensitivity-
analysis-for-rare-events-based-on-R%c3%a9nyi-divergence/
10.1214/19-AAP1468.full.

Donnelly:1991:FCL

[DKT91] Peter Donnelly, Thomas G. Kurtz, and Simon Tavare. On the functional central limit theorem for the Ewens sampling formula. Annals of Applied Probability, 1(4):539–545, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005837.

Dembo:1994:LEM

[DKZ94] Amir Dembo, Samuel Karlin, and Ofer Zeitouni. Large exceedances for multidimensional Levy processes. Annals of Applied Probability, 4(2):432–447, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005067.

Dai:2008:AOM

[DL08] J. G. Dai and Wuqin Lin. Asymptotic optimality of maximum pressure policies in stochastic processing networks. Annals of Applied Probability, 18(6):2239–2299, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1227708918.

Deaconu:2010:SDM

[DL10] Madalina Deaconu and Antoine Lejay. Simulation of diffusions by means of importance sampling paradigm. Annals of Applied Probability, 20(4):1389–1424, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1279638790.

Dereich:2016:MMC

[DL16] Steffen Dereich and Sangmeng Li. Multilevel Monte Carlo for Lévy-driven SDEs: Central limit theorems for adaptive Euler schemes. Annals of Applied Probability, 26(1):136–185, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1452003236.

Deligiannidis:2018:WEA

[DL18a] George Deligiannidis and Anthony Lee. Which ergodic averages have finite asymptotic variance? Annals of Applied Probability, 28(4):2309–2334, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1533780274.

Duchamps:2018:MRB

[DL18b] Jean-Jil Duchamps and Amaury Lambert. Mutations on a random binary tree with measured boundary. Annals of Applied Probability, 28(4):2141–2187, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1533780270.

Dumaz:2022:SAO

[DL22] Laure Dumaz and Cyril Labbé. The stochastic Airy operator at large temperature. Annals of Applied Probability, 32(6):4481– 4534, December 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
6/The-stochastic-Airy-operator-at-large-temperature/
10.1214/22-AAP1793.full.

deLima:1997:MVC

[dLS97] A. C. Pedroso de Lima and Pranab K. Sen. A matrix-valued counting process with first-order interactive intensities. Annals of Applied Probability, 7(2):494–507, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034625341.

Doytchinov:2001:RTQ

[DLS01] Bogdan Doytchinov, John Lehoczky, and Steven Shreve. Realtime queues in heavy traffic with earliest-deadline-first queue discipline. Annals of Applied Probability, 11(2):332–378, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1015345295.

Duffy:2003:LAS

[DLS03] Ken Duffy, John T. Lewis, and Wayne G. Sullivan. Logarithmic asymptotics for the supremum of a stochastic process. Annals of Applied Probability, 13(2):430–445, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1050689587.

Dembo:2021:ULL

[DLZ21] Amir Dembo, Eyal Lubetzky, and Ofer Zeitouni. Universality for Langevin-like spin glass dynamics. Annals of Applied Probability, 31(6):2864–2880, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-6/Universality-for-Langevin-like-spin-glass-dynamics/
10.1214/21-AAP1665.full.

Dalal:1992:BEC

[DM92] S. R. Dalal and C. L. Mallows. Buying with exact confidence. Annals of Applied Probability, 2(3):752–765, August 1992. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005658.

Deuschel:1994:CTN

[DM94] Jean-Dominique Deuschel and Christian Mazza. L^{2} convergence of time nonhomogeneous Markov processes: I. Spectral estimates. Annals of Applied Probability, 4(4):1012–1056, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004901.

Denuit:2002:SGI

[DM02] Michel Denuit and Alfred Müller. Smooth generators of integral stochastic orders. Annals of Applied Probability, 12(4): 1174–1184, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1037125858.

Doney:2005:CER

[DM05] R. A. Doney and R. A. Maller. Cramér’s estimate for a reflected Lévy process. Annals of Applied Probability, 15(2):1445–1450, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1115137981.

Delarue:2006:FBS

[DM06a] François Delarue and Stéphane Menozzi. A forward–backward stochastic algorithm for quasi-linear PDEs. Annals of Applied Probability, 16(1):140–184, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654284.

Denis:2006:TFP

[DM06b] Laurent Denis and Claude Martini. A theoretical framework for the pricing of contingent claims in the presence of model uncertainty. Annals of Applied Probability, 16(2):827–852, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1151592252.

Decreusefond:2008:FCL

[DM08a] Laurent Decreusefond and Pascal Moyal. A functional central limit theorem for the M/GI/∞ queue. Annals of Applied Probability, 18(6):2156–2178, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1227708915.

Dembo:2008:FSS

[DM08b] Amir Dembo and Andrea Montanari. Finite size scaling for the core of large random hypergraphs. Annals of Applied Probability, 18(5):1993–2040, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1225372959.

Dembo:2010:IML

[DM10] Amir Dembo and Andrea Montanari. Ising models on locally tree-like graphs. Annals of Applied Probability, 20(2):565–592, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1268143433.

Durrett:2011:TWS

[DM11] Rick Durrett and John Mayberry. Traveling waves of selective sweeps. Annals of Applied Probability, 21(2):699–744, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1300800986.

Durrett:2015:SMM

[DM15] Richard Durrett and Stephen Moseley. Spatial Moran models I. Stochastic tunneling in the neutral case. Annals of Applied Probability, 25(1):104–115, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1418740180.

Durmus:2017:NCA

[DM17] Alain Durmus and Éric Moulines. Nonasymptotic convergence analysis for the unadjusted Langevin algorithm. Annals of Applied Probability, 27(3):1551–1587, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1500451235.

Dyszewski:2020:HMR

[DM20] Piotr Dyszewski and Thomas Mikosch. Homogeneous mappings of regularly varying vectors. Annals of Applied Probability, 30 (6):2999–3026, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-6/Homogeneous-mappings-of-regularly-varying-vectors/
10.1214/20-AAP1579.full.

Deb:2023:FMF

[DM23a] Nabarun Deb and Sumit Mukherjee. Fluctuations in mean-field Ising models. Annals of Applied Probability, 33(3):1961–2003, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/Fluctuations-
in-mean-field-Ising-models/10.1214/22-AAP1857.full.

Dirksen:2023:ROB

[DM23b] Sjoerd Dirksen and Shahar Mendelson. Robust one-bit compressed sensing with partial circulant matrices. Annals of Applied Probability, 33(3):1874–1903, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-3/Robust-one-bit-compressed-sensing-
with-partial-circulant-matrices/10.1214/22-AAP1855.full.

Dereich:2017:NCR

[DMM17] Steffen Dereich, Cécile Mailler, and Peter Mörters. Nonextensive condensation in reinforced branching processes. Annals of Applied Probability, 27(4):2539–2568, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1504080040.

Douc:2014:LTS

[DMO14] Randal Douc, Eric Moulines, and Jimmy Olsson. Long-term stability of sequential Monte Carlo methods under verifiable conditions. Annals of Applied Probability, 24(5):1767–1802, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1403812361.

Douglas:1996:NMF

[DMP96] Jim Douglas, Jin Ma, and Philip Protter. Numerical methods for forward-backward stochastic differential equations. Annals of Applied Probability, 6(3):940–968, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968235.

DeAngelis:2022:VNM

[DMP22] Tiziano De Angelis, Nikita Merkulov, and Jan Palczewski. On the value of non-Markovian Dynkin games with partial and asymmetric information. Annals of Applied Probability, 32(3): 1774–1813, June 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
3/On-the-value-of-non-Markovian-Dynkin-games-with-partial/
10.1214/21-AAP1721.full.

Douc:2004:QBC

[DMR04] R. Douc, E. Moulines, and Jeffrey S. Rosenthal. Quantitative bounds on convergence of time-inhomogeneous Markov chains. Annals of Applied Probability, 14(4):1643–1665, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1099674073.

Duffie:1995:BCR

[DMY95] Darrell Duffie, Jin Ma, and Jiongmin Yong. Black’s consol rate conjecture. Annals of Applied Probability, 5(2):356–382, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004768.

Durrett:1991:ER

[DN91] R. Durrett and C. Neuhauser. Epidemics with recovery in D = 2. Annals of Applied Probability, 1(2):189–206, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005933.

Dai:1994:CMQ

[DN94] J. G. Dai and Vien Nguyen. On the convergence of multiclass queueing networks in heavy traffic. Annals of Applied Probability, 4(1):26–42, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005199.

Durrett:1997:CRS

[DN97] Rick Durrett and Claudia Neuhauser. Coexistence results for some competition models. Annals of Applied Probability, 7 (1):10–45, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034625251.

Darling:2005:SLR

[DN05] R. W. R. Darling and J. R. Norris. Structure of large random hypergraphs. Annals of Applied Probability, 15(1A):125–152, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1106922324.

Doney:1991:LSN

[DO91] R. A. Doney and George L. O’Brien. Loud shot noise. Annals of Applied Probability, 1(1):88–103, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005982.

Dupuis:1994:OCF

[DO94] Paul Dupuis and John Oliensis. An optimal control formulation and related numerical methods for a problem in shape reconstruction. Annals of Applied Probability, 4(2):287–346, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005063.

Dolinsky:2010:AWC

[Dol10] Yan Dolinsky. Applications of weak convergence for hedging of game options. Annals of Applied Probability, 20(5):1891–1906, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1282747404.

Dolinsky:2013:HGO

[Dol13] Yan Dolinsky. Hedging of game options with the presence of transaction costs. Annals of Applied Probability, 23(6): 2212–2237, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1382447686.

De:2019:OMB

[DOS19] Anindya De, Ryan O’Donnell, and Rocco A. Servedio. Optimal mean-based algorithms for trace reconstruction. Annals of Applied Probability, 29(2):851–874, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1548298932.

DeDonno:2005:TSI

[DP05] M. De Donno and M. Pratelli. A theory of stochastic integration for bond markets. Annals of Applied Probability, 15(4): 2773–2791, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965779.

Durrett:2009:DDA

[DP09] Rick Durrett and Lea Popovic. Degenerate diffusions arising from gene duplication models. Annals of Applied Probability, 19(1):15–48, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1235140331.

Davis:2015:ESI

[DP15] M. H. A. Davis and M. R. Pistorius. Explicit solution of an inverse first-passage time problem for Lévy processes and counterparty credit risk. Annals of Applied Probability, 25(5):2383–2415, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1438261044.

Degond:2019:PCT

[DP19] P. Degond and M. Pulvirenti. Propagation of chaos for topological interactions. Annals of Applied Probability, 29(4):2594–2612, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1563869051.

DeAngelis:2020:GRV

[DP20] T. De Angelis and G. Peskir. Global C^{1} regularity of the value function in optimal stopping problems. Annals of Applied Probability, 30(3):1007–1031, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-3/Global-C1-regularity-of-the-value-
function-in-optimal-stopping/10.1214/19-AAP1517.full.

Deligiannidis:2021:RHM

[DPBCD21] George Deligiannidis, Daniel Paulin, Alexandre Bouchard-Côté, and Arnaud Doucet. Randomized Hamiltonian Monte Carlo as scaling limit of the bouncy particle sampler and dimension-free convergence rates. Annals of Applied Probability, 31(6):2612– 2662, December 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
6/Randomized-Hamiltonian-Monte-Carlo-as-scaling-limit-
of-the-bouncy/10.1214/20-AAP1659.full.

DelMoral:2009:TBF

[DPR09] Pierre Del Moral, Frédéric Patras, and Sylvain Rubenthaler. Tree based functional expansions for Feynman–Kac particle models. Annals of Applied Probability, 19(2):778–825, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1241702250.

Ding:2019:WMN

[DPRZ19] Jian Ding, Yuval Peres, Gireeja Ranade, and Alex Zhai. When multiplicative noise stymies control. Annals of Applied Probability, 29(4):1963–1992, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1563869034.

Dayanik:2008:MBS

[DPS08] Savas Dayanik, H. Vincent Poor, and Semih O. Sezer. Multisource Bayesian sequential change detection. Annals of Applied Probability, 18(2):552–590, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1206018197.

Deelstra:2001:DFU

[DPT01] Griselda Deelstra, Huyên Pham, and Nizar Touzi. Dual formulation of the utility maximization problem under transaction costs. Annals of Applied Probability, 11(4):1353–1383, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1015345406.

Duffie:1991:MVH

[DR91] Darrell Duffie and Henry R. Richardson. Mean-variance hedging in continuous time. Annals of Applied Probability, 1(1):1–15, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005978.

Daley:1992:FWT

[DR92] D. J. Daley and T. Rolski. Finiteness of waiting-time moments in general stationary single-server queues. Annals of Applied Probability, 2(4):987–1008, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005585.

Davis:1993:PSM

[DR93] Richard A. Davis and Sidney I. Resnick. Prediction of stationary max-stable processes. Annals of Applied Probability, 3(2):497– 525, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005435.

Davis:1996:LTB

[DR96] Richard A. Davis and Sidney I. Resnick. Limit theory for bilinear processes with heavy-tailed noise. Annals of Applied Probability, 6(4):1191–1210, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1035463328.

Dupuis:1998:LDP

[DR98] Paul Dupuis and Kavita Ramanan. Large deviation properties of data streams that share a buffer. Annals of Applied Probability, 8(4):1070–1129, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903374.

Dantzer:2002:FLS

[DR02] Jean-François Dantzer and Philippe Robert. Fluid limits of string valued Markov processes. Annals of Applied Probability, 12(3):860–889, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1031863172.

Durrett:2008:ODS

[DR08] Richard Durrett and Mateo Restrepo. One-dimensional stepping stone models, sardine genetics and Brownian local time. Annals of Applied Probability, 18(1):334–358, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1199890025.

Durrett:2009:CSE

[DR09] Rick Durrett and Daniel Remenik. Chaos in a spatial epidemic model. Annals of Applied Probability, 19(4):1656–1685, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1248700631.

Dolera:2010:RCL

[DR10] Emanuele Dolera and Eugenio Regazzini. The role of the central limit theorem in discovering sharp rates of convergence to equilibrium for the solution of the Kac equation. Annals of Applied Probability, 20(2):430–461, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1268143429.

DelMoral:2011:CIM

[DR11] Pierre Del Moral and Emmanuel Rio. Concentration inequalities for mean field particle models. Annals of Applied Probability, 21(3):1017–1052, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1307020390.

Daskalakis:2013:AFP

[DR13] Constantinos Daskalakis and Sebastien Roch. Alignment-free phylogenetic reconstruction: Sample complexity via a branching process analysis. Annals of Applied Probability, 23(2):693–721, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1360682027.

Drees:2000:WAT

[Dre00] Holger Drees. Weighted approximations of tail processes for βmixing random variables. Annals of Applied Probability, 10(4): 1274–1301, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487617.

Davies:2021:RTT

[DRR21] Sami Davies, Miklós Z. Rácz, and Cyrus Rashtchian. Reconstructing trees from traces. Annals of Applied Probability, 31 (6):2772–2810, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-6/Reconstructing-trees-from-traces/10.1214/21-
AAP1662.full.

Douc:2016:MSA

[DRS16] Randal Douc, François Roueff, and Tepmony Sim. The maximizing set of the asymptotic normalized log-likelihood for partially observed Markov chains. Annals of Applied Probability, 26(4):2357–2383, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1472745461.

DaiPra:2009:LPL

[DRST09] Paolo Dai Pra, Wolfgang J. Runggaldier, Elena Sartori, and Marco Tolotti. Large portfolio losses: a dynamic contagion model. Annals of Applied Probability, 19(1):347–394, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1235140342.

dosReis:2019:FWL

[dRST19] Gonçalo dos Reis, William Salkeld, and Julian Tugaut. Freidlin– Wentzell LDP in path space for McKean–Vlasov equations and the functional iterated logarithm law. Annals of Applied Probability, 29(3):1487–1540, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1550566836.

Durmus:2017:FLB

[DRVZ17] Alain Durmus, Gareth O. Roberts, Gilles Vilmart, and Konstantinos C. Zygalakis. Fast Langevin based algorithm for MCMC in high dimensions. Annals of Applied Probability, 27 (4):2195–2237, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1504080030.

Dick:2016:DBU

[DRZ16] Josef Dick, Daniel Rudolf, and Houying Zhu. Discrepancy bounds for uniformly ergodic Markov chain quasi-Monte Carlo. Annals of Applied Probability, 26(5):3178–3205, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1476884315.

Diaconis:1991:GBE

[DS91] Persi Diaconis and Daniel Stroock. Geometric bounds for eigenvalues of Markov chains. Annals of Applied Probability, 1(1):36– 61, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005980.

Durrett:1993:ACV

[DS93] Richard Durrett and Rinaldo Schinazi. Asymptotic critical value for a competition model. Annals of Applied Probability, 3(4): 1047–1066, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005272.

Delbaen:1995:EAC

[DS95] Freddy Delbaen and Walter Schachermayer. The existence of absolutely continuous local martingale measures. Annals of Applied Probability, 5(4):926–945, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004600.

Dietz:2005:LDC

[DS05] Zach Dietz and Sunder Sethuraman. Large deviations for a class of nonhomogeneous Markov chains. Annals of Applied Probability, 15(1A):421–486, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1106922333.

Duffie:2007:EIR

[DS07a] Darrell Duffie and Yeneng Sun. Existence of independent random matching. Annals of Applied Probability, 17(1):386–419, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1171377188.

Durrett:2007:WRS

[DS07b] Richard Durrett and Deena Schmidt. Waiting for regulatory sequences to appear. Annals of Applied Probability, 17(1):1–32, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1171377175.

Dalang:2015:QDP

[DS15] Robert C. Dalang and Albert N. Shiryaev. A quickest detection problem with an observation cost. Annals of Applied Probability, 25(3):1475–1512, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1427124134.

Dondl:2017:BSB

[DS17] Patrick W. Dondl and Michael Scheutzow. Ballistic and subballistic motion of interfaces in a field of random obstacles. Annals of Applied Probability, 27(5):3189–3200, October 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1509696044.

Davis:2018:CMC

[DS18] Erik Davis and Sunder Sethuraman. Consistency of modularity clustering on random geometric graphs. Annals of Applied Probability, 28(4):2003–2062, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1533780266.

Davis:2019:AGR

[DS19] Erik Davis and Sunder Sethuraman. Approximating geodesics via random points. Annals of Applied Probability, 29(3):1446– 1486, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1550566835.

Dhara:2022:LDU

[DS22] Souvik Dhara and Subhabrata Sen. Large deviation for uniform graphs with given degrees. Annals of Applied Probability, 32(3): 2327–2353, June 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
3/Large-deviation-for-uniform-graphs-with-given-degrees/
10.1214/21-AAP1745.full.

Diaconis:1993:CTR

[DSC93] Persi Diaconis and Laurent Saloff-Coste. Comparison theorems for reversible Markov chains. Annals of Applied Probability, 3 (3):696–730, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005359.

Diaconis:1996:LSI

[DSC96] P. Diaconis and L. Saloff-Coste. Logarithmic Sobolev inequalities for finite Markov chains. Annals of Applied Probability, 6(3):695– 750, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1034968224.

Diaconis:2006:SCO

[DSC06] Persi Diaconis and Laurent Saloff-Coste. Separation cut-offs for birth and death chains. Annals of Applied Probability, 16(4): 2098–2122, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1169065218.

deSaporta:2010:NMO

[dSDG10] Benoîte de Saporta, François Dufour, and Karen Gonzalez. Numerical method for optimal stopping of piecewise deterministic Markov processes. Annals of Applied Probability, 20(5): 1607–1637, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1282747395.

Duffie:1996:RVD

[DSS96] Darrell Duffie, Mark Schroder, and Costis Skiadas. Recursive valuation of defaultable securities and the timing of resolution of uncertainty. Annals of Applied Probability, 6(4):1075–1090, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1035463324.

Durrett:2009:WTP

[DSS09] Rick Durrett, Deena Schmidt, and Jason Schweinsberg. A waiting time problem arising from the study of multi-stage carcinogenesis. Annals of Applied Probability, 19(2):676–718, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1241702247.

Dupuis:2007:DIS

[DSW07] Paul Dupuis, Ali Devin Sezer, and Hui Wang. Dynamic importance sampling for queueing networks. Annals of Applied Probability, 17(4):1306–1346, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1186755241.

deSaporta:2005:TLD

[dSY05] Benoîte de Saporta and Jian-Feng Yao. Tail of a linear diffusion with Markov switching. Annals of Applied Probability, 15 (1B):992–1018, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1107271676.

Dupuis:2015:EAI

[DSZ15] Paul Dupuis, Konstantinos Spiliopoulos, and Xiang Zhou. Escaping from an attractor: Importance sampling and rest points I. Annals of Applied Probability, 25(5):2909–2958, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1438261056.

DelMoral:2005:BET

[DT05] Pierre Del Moral and Samy Tindel. A Berry–Esseen theorem for Feynman–Kac and interacting particle models. Annals of Applied Probability, 15(1B):941–962, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1107271673.

DelMoral:2018:SUP

[DT18] P. Del Moral and J. Tugaut. On the stability and the uniform propagation of chaos properties of Ensemble Kalman–Bucy filters. Annals of Applied Probability, 28(2):790–850, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1523433625.

duToit:2009:SSU

[dTP09] Jacques du Toit and Goran Peskir. Selling a stock at the ultimate maximum. Annals of Applied Probability, 19(3):983–1014, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1245071016.

Dombry:2022:HRV

[DTW22] Clément Dombry, Charles Tillier, and Olivier Wintenberger. Hidden regular variation for point processes and the single/ multiple large point heuristic. Annals of Applied Probability, 32(1):191–234, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-1/Hidden-regular-variation-for-point-processes-
and-the-single-multiple/10.1214/21-AAP1675.full.

Duffy:2016:ULC

[Duf16] James A. Duffy. A uniform law for convergence to the local times of linear fractional stable motions. Annals of Applied Probability, 26(1):45–72, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1452003234.

Durrett:1999:CRB

[Dur99a] Richard Durrett. Correction: “Robert Bättig 1968–1997”. Annals of Applied Probability, 9(4):??, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1029962874. See

[Dur99b].

Durrett:1999:RB

[Dur99b] Rick Durrett. Robert Bättig 1968–1997. Annals of Applied Probability, 9(2):??, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962741. See correction [Dur99a].

Durrett:2009:CSS

[Dur09] Rick Durrett. Coexistence in stochastic spatial models. Annals of Applied Probability, 19(2):477–496, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1241702238.

Durrett:2013:PGN

[Dur13] Rick Durrett. Population genetics of neutral mutations in exponentially growing cancer cell populations. Annals of Applied Probability, 23(1):230–250, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1359124385.

Deijfen:2016:WTI

[DvdH16] Maria Deijfen and Remco van der Hofstad. The winner takes it all. Annals of Applied Probability, 26(4):2419–2453, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1472745463.

Duffield:1998:LDI

[DW98] N. G. Duffield and W. Whitt. Large deviations of inverse processes with nonlinear scalings. Annals of Applied Probability, 8 (4):995–1026, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903372.

Dupuis:2005:DIS

[DW05a] Paul Dupuis and Hui Wang. Dynamic importance sampling for uniformly recurrent Markov chains. Annals of Applied Probability, 15(1A):1–38, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1106922319.

Dupuis:2005:CDC

[DW05b] Paul Dupuis and Hui Wang. On the convergence from discrete to continuous time in an optimal stopping problem. Annals of Applied Probability, 15(2):1339–1366, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1115137977.

Dong:2020:LDP

[DWZZ20] Zhao Dong, Jiang-Lun Wu, Rangrang Zhang, and Tusheng Zhang. Large deviation principles for first-order scalar conservation laws with stochastic forcing. Annals of Applied Probability, 30(1):324–367, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1582621226.

Dempster:1995:ICP

[DY95] M. A. H. Dempster and J. J. Ye. Impulse control of piecewise deterministic Markov processes. Annals of Applied Probability, 5(2):399–423, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004771.

Ding:2018:NSC

[DY18] Xiucai Ding and Fan Yang. A necessary and sufficient condition for edge universality at the largest singular values of covariance matrices. Annals of Applied Probability, 28(3):1679–1738, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1527840030.

Davis:1994:PSS

[DZ94] M. H. A. Davis and M. Zervos. A problem of singular stochastic control with discretionary stopping. Annals of Applied Probability, 4(1):226–240, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005209.

Duckworth:2001:MID

[DZ01] Kate Duckworth and Mihail Zervos. A model for investment decisions with switching costs. Annals of Applied Probability, 11(1):239–260, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/998926992.

Durrett:2015:CGS

[DZ15] Rick Durrett and Yuan Zhang. Coexistence of grass, saplings and trees in the Staver–Levin forest model. Annals of Applied Probability, 25(6):3434–3464, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1443703779.

DiNunno:2016:ASP

[DZ16] Giulia Di Nunno and Tusheng Zhang. Approximations of stochastic partial differential equations. Annals of Applied Probability, 26(3):1443–1466, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905008.

Eberle:2014:EBM

[Ebe14] Andreas Eberle. Error bounds for Metropolis–Hastings algorithms applied to perturbations of Gaussian measures in high dimensions. Annals of Applied Probability, 24(1):337–377, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1389278728.

Ettinger:2014:KBM

[EEH14] Boris Ettinger, Steven N. Evans, and Alexandru Hening. Killed Brownian motion with a prescribed lifetime distribution and models of default. Annals of Applied Probability, 24(1):1–33, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1389278717.

Ernst:2021:ABO

[EF21] Philip A. Ernst and Sandro Franceschi. Asymptotic behavior of the occupancy density for obliquely reflected Brownian motion in a half-plane and Martin boundary. Annals of Applied Probability, 31(6):2991–3016, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-6/Asymptotic-behavior-of-the-occupancy-density-
for-obliquely-reflected-Brownian/10.1214/21-AAP1681.
full.

Erhard:2021:SBR

[EFdS21] Dirk Erhard, Tertuliano Franco, and Diogo S. da Silva. The slow bond random walk and the snapping out Brownian motion. Annals of Applied Probability, 31(1):99–127, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/The-
slow-bond-random-walk-and-the-snapping-out-Brownian/
10.1214/20-AAP1584.full.

Etheridge:2017:BBM

[EFPS17] Alison Etheridge, Nic Freeman, Sarah Penington, and Daniel Straulino. Branching Brownian motion and selection in the spatial Λ-Fleming-Viot process. Annals of Applied Probability, 27 (5):2605–2645, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1509696030.

Elie:2007:KEG

[EFT07] Romuald Elie, Jean-David Fermanian, and Nizar Touzi. Kernel estimation of Greek weights by parameter randomization. Annals of Applied Probability, 17(4):1399–1423, August 2007. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1186755244.

Eldan:2018:ERG

[EG18] Ronen Eldan and Renan Gross. Exponential random graphs behave like mixtures of stochastic block models. Annals of Applied Probability, 28(6):3698–3735, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1538985633.

Eichinger:2022:MAT

[EGK22] Katharina Eichinger, Manuel V. Gnann, and Christian Kuehn. Multiscale analysis for traveling-pulse solutions to the stochastic FitzHugh–Nagumo equations. Annals of Applied Probability, 32(5):3229–3282, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-5/Multiscale-analysis-for-traveling-pulse-solutions-
to-the-stochastic-FitzHughNagumo/10.1214/21-AAP1759.
full.

Egloff:2005:MCA

[Egl05] Daniel Egloff. Monte Carlo algorithms for optimal stopping and statistical learning. Annals of Applied Probability, 15(2):1396– 1432, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1115137979.

Edelman:2016:BUR

[EGP16] Alan Edelman, A. Guionnet, and S. Péché. Beyond universality in random matrix theory. Annals of Applied Probability, 26(3):1659–1697, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1465905015.

Eichenauer-Herrmann:1995:PNG

[EH95] Jürgen Eichenauer-Herrmann. Pseudorandom number generation by nonlinear methods. International Statistical Review = Revue Internationale de Statistique, 63(2):247–255, August 1995. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823

(electronic). URL http://www.jstor.org/stable/1403620.

Ekstrom:2011:RTH

[EH11] Erik Ekström and David Hobson. Recovering a timehomogeneous stock price process from perpetual option prices. Annals of Applied Probability, 21(3):1102–1135, June 2011. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1307020393.

Embrechts:2016:BTD

[EHW16] Paul Embrechts, Marius Hofert, and Ruodu Wang. Bernoulli and tail-dependence compatibility. Annals of Applied Probability, 26(3):1636–1658, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1465905014.

Ettinger:2020:IFP

[EHW20] Boris Ettinger, Alexandru Hening, and Tak Kwong Wong. The inverse first passage time problem for killed Brownian motion. Annals of Applied Probability, 30(3):1251–1275, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-3/The-
inverse-first-passage-time-problem-for-killed-Brownian-
motion/10.1214/19-AAP1529.full.

Engelke:2017:RBM

[EI17] Sebastian Engelke and Jevgenijs Ivanovs. Robust bounds in multivariate extremes. Annals of Applied Probability, 27(6): 3706–3734, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1513328712.

Eichler:1995:ESP

[Eic95] Michael Eichler. Empirical spectral processes and their applications to stationary point processes. Annals of Applied Probability, 5(4):1161–1176, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004610.

Ekstrom:2016:IFP

[EJ16] Erik Ekström and Svante Janson. The inverse first-passage problem and optimal stopping. Annals of Applied Probability, 26 (5):3154–3177, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1476884314.

Ethier:1992:SDN

[EK92] S. N. Ethier and Thomas G. Kurtz. On the stationary distribution of the neutral diffusion model in population genetics. Annals of Applied Probability, 2(1):24–35, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005769.

ElKaroui:1994:DAP

[EK94] Nicole El Karoui and Ioannis Karatzas. Dynamic allocation problems in continuous time. Annals of Applied Probability, 4(2): 255–286, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005062.

Eder:2009:FPE

[EK09] Irmingard Eder and Claudia Klüppelberg. The first passage event for sums of dependent Lévy processes with applications to insurance risk. Annals of Applied Probability, 19(6):2047–2079, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1259158766.

Evans:2000:BTI

[EKPS00] William Evans, Claire Kenyon, Yuval Peres, and Leonard J. Schulman. Broadcasting on trees and the Ising model. Annals of Applied Probability, 10(2):410–433, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1019487349.

Erdos:2023:RCD

[EKR23] László Erdős, Torben Krüger, and David Renfrew. Randomly coupled differential equations with elliptic correlations. Annals of Applied Probability, 33(4):3098–3144, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-4/Randomly-
coupled-differential-equations-with-elliptic-correlations/
10.1214/22-AAP1886.full.

Egloff:2007:DLA

[EKT07] Daniel Egloff, Michael Kohler, and Nebojsa Todorovic. A dynamic look-ahead Monte Carlo algorithm for pricing Bermudan options. Annals of Applied Probability, 17(4):1138–1171, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1186755235.

ElKaroui:2009:CMS

[El 09] Noureddine El Karoui. Concentration of measure and spectra of random matrices: Applications to correlation matrices, elliptical distributions and beyond. Annals of Applied Probability, 19(6): 2362–2405, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1259158775.

Ethier:2010:MSM

[EL10] S. N. Ethier and Jiyeon Lee. A Markovian slot machine and Parrondo’s paradox. Annals of Applied Probability, 20(3):1098– 1125, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1276867308.

Einarsson:2016:CTB

[ELM^{+}16] Hafsteinn Einarsson, Johannes Lengler, Frank Mousset, Konstantinos Panagiotou, and Angelika Steger. Connectivity thresholds for bounded size rules. Annals of Applied Probability, 26 (5):3206–3250, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1476884316.

Ehrlacher:2022:AFB

[ELM22] Virginie Ehrlacher, Tony Lelièvre, and Pierre Monmarché. Adaptive force biasing algorithms: New convergence results and tensor approximations of the bias. Annals of Applied Probability, 32(5):3850–3888, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-5/Adaptive-force-biasing-algorithms--New-convergence-
results-and-tensor/10.1214/21-AAP1775.full.

ElKaroui:2017:MOR

[ELS17] Nicole El Karoui, Stéphane Loisel, and Yahia Salhi. Minimax optimality in robust detection of a disorder time in doublystochastic Poisson processes. Annals of Applied Probability, 27 (4):2515–2538, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1504080039.

Ellis:2010:ABF

[EMO10] Richard S. Ellis, Jonathan Machta, and Peter Tak-Hun Otto. Asymptotic behavior of the finite-size magnetization as a function of the speed of approach to criticality. Annals of Applied Probability, 20(6):2118–2161, December 2010. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1287494556.

Edelsbrunner:2018:RIP

[EN18] Herbert Edelsbrunner and Anton Nikitenko. Random inscribed polytopes have similar radius functions as Poisson–Delaunay mosaics. Annals of Applied Probability, 28(5):3215–3238, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1535443247.

Ellis:2005:APT

[EOT05] Richard S. Ellis, Peter T. Otto, and Hugo Touchette. Analysis of phase transitions in the mean-field Blume–Emery–Griffiths model. Annals of Applied Probability, 15(3):2203–2254, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1121433782.

Elskens:2010:DLM

[EP10] Yves Elskens and Etienne Pardoux. Diffusion limit for many particles in a periodic stochastic acceleration field. Annals of Applied Probability, 20(6):2022–2039, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1287494553.

Ernst:2022:QRT

[EP22] Philip A. Ernst and Goran Peskir. Quickest real-time detection of a Brownian coordinate drift. Annals of Applied Probability, 32(4):2652–2670, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-4/Quickest-real-time-detection-of-a-Brownian-coordinate-
drift/10.1214/21-AAP1742.full.

ElKaroui:2001:DMP

[EPQ01] N. El Karoui, S. Peng, and M. C. Quenez. A dynamic maximum principle for the optimization of recursive utilities under constraints. Annals of Applied Probability, 11(3):664–693, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345345.

Eberlein:2009:ETD

[EPS09] Ernst Eberlein, Antonis Papapantoleon, and Albert N. Shiryaev. Esscher transform and the duality principle for multidimensional semimartingales. Annals of Applied Probability, 19(5): 1944–1971, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1255699549.

Etheridge:2006:ASF

[EPW06] Alison Etheridge, Peter Pfaffelhuber, and Anton Wakolbinger. An approximate sampling formula under genetic hitchhiking. Annals of Applied Probability, 16(2):685–729, May 2006. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1151592248.

Ernst:2020:ORT

[EPZ20] P. A. Ernst, G. Peskir, and Q. Zhou. Optimal real-time detection of a drifting Brownian coordinate. Annals of Applied Probability, 30(3):1032–1065, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-3/Optimal-real-time-detection-of-a-drifting-Brownian-
coordinate/10.1214/19-AAP1522.full.

Eichelsbacher:2008:SMD

[ER08] Peter Eichelsbacher and Gesine Reinert. Stein’s method for discrete Gibbs measures. Annals of Applied Probability, 18(4):1588– 1618, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1216677133.

Evans:2010:DTM

[ER10] Steven N. Evans and Peter L. Ralph. Dynamics of the time to the most recent common ancestor in a large branching population. Annals of Applied Probability, 20(1):1–25, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1262962316.

ElEuch:2018:PHR

[ER18] Omar El Euch and Mathieu Rosenbaum. Perfect hedging in rough Heston models. Annals of Applied Probability, 28(6): 3813–3856, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1538985636.

Erhardsson:2000:CPA

[Erh00] Torkel Erhardsson. Compound Poisson approximation for counts of rare patterns in Markov chains and extreme sojourns in birthdeath chains. Annals of Applied Probability, 10(2):573–591, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1019487356.

Embrechts:1995:PDR

[ERY95] P. Embrechts, L. C. G. Rogers, and M. Yor. A proof of Dassios’ representation of the α-quantile of Brownian motion with drift. Annals of Applied Probability, 5(3):757–767, August 1995. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177004704.

Evans:2002:ESF

[ES02] Steven N. Evans and David Steinsaltz. Estimating some features of NK fitness landscapes. Annals of Applied Probability, 12(4): 1299–1321, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1037125864.

Embrechts:2003:RPH

[ES03] Paul Embrechts and Gennady Samorodnitsky. Ruin problem and how fast stochastic processes. Annals of Applied Probability, 13(1):1–36, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1042765661.

Eisenberg:1993:APT

[ESS93] Bennett Eisenberg, Gilbert Stengle, and Gilbert Strang. The asymptotic probability of a Tie for first place. Annals of Applied Probability, 3(3):731–745, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005360.

Enriquez:2013:QLF

[ESTZ13] Nathanaël Enriquez, Christophe Sabot, Laurent Tournier, and Olivier Zindy. Quenched limits for the fluctuations of transient random walks in random environment on Z. Annals of Applied Probability, 23(3):1148–1187, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1362684857.

Evans:2010:CBD

[ESU10] Steven N. Evans, Bernd Sturmfels, and Caroline Uhler. Commuting birth-and-death processes. Annals of Applied Probability, 20(1):238–266, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1262962323.

Escauriaza:2022:RES

[ESX22] Luis Escauriaza, Daniel C. Schwarz, and Hao Xing. Radner equilibrium and systems of quadratic BSDEs with discontinuous generators. Annals of Applied Probability, 32(5):3492–3536, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-5/Radner-
equilibrium-and-systems-of-quadratic-BSDEs-with-discontinuous-
generators/10.1214/21-AAP1765.full.

Ekeland:2005:TBP

[ET05] Ivar Ekeland and Erik Taflin. A theory of bond portfolios. Annals of Applied Probability, 15(2):1260–1305, May 2005. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1115137975.

Ekstrom:2009:BCB

[ET09] Erik Ekström and Johan Tysk. Bubbles, convexity and the Black–Scholes equation. Annals of Applied Probability, 19(4): 1369–1384, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1248700621.

Ekstrom:2011:BCS

[ET11] Erik Ekström and Johan Tysk. Boundary conditions for the single-factor term structure equation. Annals of Applied Probability, 21(1):332–350, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1292598037.

Ethier:1996:GSM

[Eth96] S. N. Ethier. A gambling system and a Markov chain. Annals of Applied Probability, 6(4):1248–1259, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1035463331.

Etheridge:2004:SEL

[Eth04] A. M. Etheridge. Survival and extinction in a locally regulated population. Annals of Applied Probability, 14(1):188–214, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1075828051.

Ekstrom:2006:VOS

[EV06] Erik Ekström and Stephane Villeneuve. On the value of optimal stopping games. Annals of Applied Probability, 16(3):1576–1596, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804992.

Etheridge:2012:SFV

[EV12] A. M. Etheridge and A. Véber. The spatial Λ-Fleming–Viot process on a large torus: Genealogies in the presence of recombination. Annals of Applied Probability, 22(6):2165–2209, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1353695951.

Evans:2001:SBG

[Eva01] Steven N. Evans. Stochastic billiards on general tables. Annals of Applied Probability, 11(2):419–437, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015345298.

Eibeck:2001:SPA

[EW01] Andreas Eibeck and Wolfgang Wagner. Stochastic particle approximations for Smoluchoski’s coagualtion equation. Annals of Applied Probability, 11(4):1137–1165, November 2001. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1015345398.

Eibeck:2003:SIP

[EW03] Andreas Eibeck and Wolfgang Wagner. Stochastic interacting particle systems and nonlinear kinetic equations. Annals of Applied Probability, 13(3):845–889, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1060202829.

Fedrizzi:2014:SSU

[Fed14] Ennio Fedrizzi. Stability of solitons under rapidly oscillating random perturbations of the initial conditions. Annals of Applied Probability, 24(2):616–651, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1394465367.

Fehrman:2023:LSR

[Feh23] Benjamin Fehrman. Large-scale regularity in stochastic homogenization with divergence-free drift. Annals of Applied Probability, 33(4):2559–2599, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-4/Large-scale-regularity-in-stochastic-
homogenization-with-divergence-free-drift/10.1214/22-
AAP1872.full.

Feng:2007:LDA

[Fen07] Shui Feng. Large deviations associated with Poisson–Dirichlet distribution and Ewens sampling formula. Annals of Applied Probability, 17(5/6):1570–1595, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1191419176.

Ferrari:2015:IEF

[Fer15] Giorgio Ferrari. On an integral equation for the free-boundary of stochastic, irreversible investment problems. Annals of Applied Probability, 25(1):150–176, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1418740182.

Feray:2020:CLT

[Fér20] Valentin Féray. Central limit theorems for patterns in multiset permutations and set partitions. Annals of Applied Probability, 30(1):287–323, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1582621225.

Foxall:2016:SCP

[FEvdD16] Eric Foxall, Roderick Edwards, and P. van den Driessche. Social contact processes and the partner model. Annals of Applied Probability, 26(3):1297–1328, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1465905004.

Ferrari:1994:NOP

[FF94] P. A. Ferrari and L. R. G. Fontes. The net output process of a system with infinitely many queues. Annals of Applied Probability, 4(4):1129–1144, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004907.

Feng:2012:STA

[FFK12] Jin Feng, Jean-Pierre Fouque, and Rohini Kumar. Small-time asymptotics for fast mean-reverting stochastic volatility models. Annals of Applied Probability, 22(4):1541–1575, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1344614203.

Ferrari:1993:CSP

[FFS93] Pablo A. Ferrari, Arnoldo Frigessi, and Roberto H. Schonmann. Convergence of some partially parallel Gibbs samplers with annealing. Annals of Applied Probability, 3(1):137–153, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005511.

Feng:2008:MDP

[FG08] Shui Feng and Fuqing Gao. Moderate deviations for Poisson– Dirichlet distribution. Annals of Applied Probability, 18(5): 1794–1824, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1225372950.

Federico:2018:VTS

[FG18] Salvatore Federico and Fausto Gozzi. Verification theorems for stochastic optimal control problems in Hilbert spaces by means of a generalized Dynkin formula. Annals of Applied Probability, 28(6):3558–3599, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1538985629.

Fang:2020:AEM

[FG20] Wei Fang and Michael B. Giles. Adaptive Euler–Maruyama method for SDEs with nonglobally Lipschitz drift. Annals of Applied Probability, 30(2):526–560, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1591603215.

Fisch:1993:MGH

[FGG93] Robert Fisch, Janko Gravner, and David Griffeath. Metastability in the Greenberg–Hastings model. Annals of Applied Probability, 3(4):935–967, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005268.

Fontes:2014:PTS

[FGG14] L. R. G. Fontes, R. J. Gava, and V. Gayrard. The K-process on a tree as a scaling limit of the GREM-like trap model. Annals of Applied Probability, 24(2):857–897, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1394465373.

Foo:2023:SPM

[FGLS23] Jasmine Foo, Einar Bjarki Gunnarsson, Kevin Leder, and Kathleen Storey. Spread of premalignant mutant clones and cancer initiation in multilayered tissue. Annals of Applied Probability, 33(1):299–343, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-1/Spread-of-premalignant-mutant-clones-and-cancer-
initiation-in-multilayered/10.1214/22-AAP1816.full.

Friz:2021:PAR

[FGP21] P. K. Friz, P. Gassiat, and P. Pigato. Precise asymptotics: Robust stochastic volatility models. Annals of Applied Probability, 31(2):896–940, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-2/Precise-asymptotics-Robust-stochastic-volatility-
models/10.1214/20-AAP1608.full.

Fox:1995:PSO

[FH95] Bennett L. Fox and George W. Heine. Probabilistic search with overrides. Annals of Applied Probability, 5(4):1087–1094, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004607.

Feng:1998:LDP

[FH98] Shui Feng and Fred M. Hoppe. Large deviation principles for some random combinatorial structures in population genetics and Brownian motion. Annals of Applied Probability, 8(4): 975–994, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903371.

Feng:2016:GCR

[FH16] Han Feng and David Hobson. Gambling in contests with random initial law. Annals of Applied Probability, 26(1):186–215, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1452003237.

Feng:2023:UAL

[FHH23] Xinwei Feng, Ying Hu, and Jianhui Huang. A unified approach to linear-quadratic-Gaussian mean-field team: Homogeneity, heterogeneity and quasi-exchangeability. Annals of Applied Probability, 33(4):2786–2823, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-4/A-unified-approach-to-linear-quadratic-
Gaussian-mean-field-team/10.1214/22-AAP1878.full.

Flandoli:2022:GWP

[FHLN22] Franco Flandoli, Martina Hofmanová, Dejun Luo, and Torstein Nilssen. Global well-posedness of the 3D Navier–Stokes equations perturbed by a deterministic vector field. Annals of Applied Probability, 32(4):2568–2586, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-4/Global-well-posedness-of-the-3D-NavierStokes-
equations-perturbed-by/10.1214/21-AAP1740.full.

Frigessi:1992:OSS

[FHY92] Arnoldo Frigessi, Chii-Ruey Hwang, and Laurent Younes. Optimal spectral structure of reversible stochastic matrices, Monte Carlo methods and the simulation of Markov random fields. Annals of Applied Probability, 2(3):610–628, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005652.

Fromm:2020:UMD

[FI20] Alexander Fromm and Peter Imkeller. Utility maximization via decoupling fields. Annals of Applied Probability, 30(6):2665– 2694, December 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
6/Utility-maximization-via-decoupling-fields/10.1214/
20-AAP1569.full.

Fontes:1998:SGR

[FIKP98] L. R. G. Fontes, M. Isopi, Y. Kohayakawa, and P. Picco. The spectral gap of the REM under Metropolis dynamics. Annals of Applied Probability, 8(3):917–943, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1028903457.

Fill:1991:EBC

[Fil91] James Allen Fill. Eigenvalue bounds on convergence to stationarity for nonreversible Markov chains, with an application to the exclusion process. Annals of Applied Probability, 1(1):62–87, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005981.

Fill:1998:IAP

[Fil98] James Allen Fill. An interruptible algorithm for perfect sampling via Markov chains. Annals of Applied Probability, 8(1):131–162, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1027961037.

Fill:2013:DCN

[Fil13] James Allen Fill. Distributional convergence for the number of symbol comparisons used by QuickSort. Annals of Applied Probability, 23(3):1129–1147, June 2013. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1362684856.

Fountoulakis:2022:DMR

[FIMS22] Nikolaos Fountoulakis, Tejas Iyer, Cécile Mailler, and Henning Sulzbach. Dynamical models for random simplicial complexes. Annals of Applied Probability, 32(4):2860–2913, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Dynamical-
models-for-random-simplicial-complexes/10.1214/21-AAP1752.
full.

Finnoff:1994:LLN

[Fin94] William Finnoff. Law of large numbers for a heterogeneous system of stochastic differential equations with strong local interaction and economic applications. Annals of Applied Probability, 4(2):494–528, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005070.

Fishman:1996:CSR

[Fis96] George S. Fishman. Coordinate selection rules for Gibbs sampling. Annals of Applied Probability, 6(2):444–465, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1034968139.

Fischer:2017:CBS

[Fis17] Markus Fischer. On the connection between symmetric N-player games and mean field games. Annals of Applied Probability, 27(2):757–810, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1495764366.

Fournier:2017:SPA

[FJ17] Nicolas Fournier and Benjamin Jourdain. Stochastic particle approximation of the Keller–Segel equation and two-dimensional generalization of Bessel processes. Annals of Applied Probability, 27(5):2807–2861, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1509696035.

Fukushima:2019:ZTL

[FJ19] Ryoki Fukushima and Stefan Junk. Zero temperature limit for the Brownian directed polymer among Poissonian disasters. Annals of Applied Probability, 29(6):3821–3860, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1578366327.

Fan:2022:TWE

[FJ22] Zhou Fan and Iain M. Johnstone. Tracy-Widom at each edge of real covariance and MANOVA estimators. Annals of Applied Probability, 32(4):2967–3003, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-4/TracyWidom-at-each-edge-of-real-covariance-
and-MANOVA-estimators/10.1214/21-AAP1754.full.

Foondun:2018:ARC

[FJL18] Mohammud Foondun, Mathew Joseph, and Shiu-Tang Li. An approximation result for a class of stochastic heat equations with colored noise. Annals of Applied Probability, 28(5):2855–2895, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1535443236.

Friesen:2020:SEE

[FJR20] Martin Friesen, Peng Jin, and Barbara Rüdiger. Stochastic equation and exponential ergodicity in Wasserstein distances for affine processes. Annals of Applied Probability, 30(5):2165–2195, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-5/Stochastic-
equation-and-exponential-ergodicity-in-Wasserstein-distances-
for-affine/10.1214/19-AAP1554.full.

Fannjiang:1999:TDM

[FK99a] Albert Fannjiang and Tomasz Komorowski. Turbulent diffusion in Markovian flows. Annals of Applied Probability, 9(3):591–610, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962805.

Fleischmann:1999:SDF

[FK99b] Klaus Fleischmann and Achim Klenke. Smooth density field of catalytic super-Brownian motion. Annals of Applied Probability, 9(2):298–318, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962743.

Frieze:1999:LSI

[FK99c] Alan Frieze and Ravi Kannan. Log-Sobolev inequalities and sampling from log-concave distributions. Annals of Applied Probability, 9(1):14–26, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962595.

Fannjiang:2000:FBM

[FK00a] Albert Fannjiang and Tomasz Komorowski. Fractional Brownian motions in a limit of turbulent transport. Annals of Applied Probability, 10(4):1100–1120, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487608.

Fleischmann:2000:BCS

[FK00b] Klaus Fleischmann and Achim Klenke. The biodiversity of catalytic super-Brownian motion. Annals of Applied Probability, 10 (4):1121–1136, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487609.

Fernholz:2010:OA

[FK10] Daniel Fernholz and Ioannis Karatzas. On optimal arbitrage. Annals of Applied Probability, 20(4):1179–1204, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1279638783.

Fernholz:2011:OAU

[FK11] Daniel Fernholz and Ioannis Karatzas. Optimal arbitrage under model uncertainty. Annals of Applied Probability, 21(6): 2191–2225, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1322057320.

Fill:2013:CIF

[FK13] James Allen Fill and Jonas Kahn. Comparison inequalities and fastest-mixing Markov chains. Annals of Applied Probability, 23 (5):1778–1816, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1377696298.

Forde:2016:LTO

[FK16] Martin Forde and Rohini Kumar. Large-time option pricing using the donsker-varadhan LDP-correlated stochastic volatility with stochastic interest rates and jumps. Annals of Applied Probability, 26(6):3699–3726, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1481792597.

Fang:2021:HDC

[FK21] Xiao Fang and Yuta Koike. High-dimensional central limit theorems by Stein’s method. Annals of Applied Probability, 31(4): 1660–1686, August 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
4/High-dimensional-central-limit-theorems-by-Steins-
method/10.1214/20-AAP1629.full.

Fang:2022:NEB

[FK22] Xiao Fang and Yuta Koike. New error bounds in multivariate normal approximations via exchangeable pairs with applications to Wishart matrices and fourth moment theorems. Annals of Applied Probability, 32(1):602–631, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/New-
error-bounds-in-multivariate-normal-approximations-via-
exchangeable-pairs/10.1214/21-AAP1690.full.

Fagin:2001:RWB

[FKK^{+}01] Ronald Fagin, Anna R. Karlin, Jon Kleinberg, Prabhakar Raghavan, Sridhar Rajagopalan, Ronitt Rubinfeld, Madhu Sudan, and Andrew Tomkins. Random walks with “back buttons”. Annals of Applied Probability, 11(3):810–862, August 2001. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1015345350.

Fountoulakis:2018:PTR

[FKKM18] Nikolaos Fountoulakis, Mihyun Kang, Christoph Koch, and Tamás Makai. A phase transition regarding the evolution of bootstrap processes in inhomogeneous random graphs. Annals of Applied Probability, 28(2):990–1051, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1523433630.

Ferrari:1996:PQS

[FKM96] P. A. Ferrari, H. Kesten, and S. Martínez. R-positivity, quasistationary distributions and ratio limit theorems for a class of probabilistic automata. Annals of Applied Probability, 6(2):577– 616, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968146.

Frieze:1994:CSL

[FKP94a] Alan Frieze, Ravi Kannan, and Nick Polson. Correction: “Sampling from Log-Concave Distributions”. Annals of Applied Probability, 4(4):1255, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004915. See [FKP94b].

Frieze:1994:SLC

[FKP94b] Alan Frieze, Ravi Kannan, and Nick Polson. Sampling from logconcave distributions. Annals of Applied Probability, 4(3):812– 837, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.

aoap/1177004973. See correction [FKP94a].

Foss:2023:PAP

[FKP23] Sergey Foss, Takis Konstantopoulos, and Artem Pyatkin. Probabilistic and analytical properties of the last passage percolation constant in a weighted random directed graph. Annals of Applied Probability, 33(2):731–753, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-2/Probabilistic-and-analytical-properties-
of-the-last-passage-percolation-constant/10.1214/22-
AAP1832.full.

Feigin:1996:PEM

[FKR96] Paul D. Feigin, Marie F. Kratz, and Sidney I. Resnick. Parameter estimation for moving averages with positive innovations. Annals of Applied Probability, 6(4):1157–1190, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1035463327.

Feige:2017:CSR

[FKR17] Uriel Feige, Michael Krivelevich, and Daniel Reichman. Contagious sets in random graphs. Annals of Applied Probability, 27 (5):2675–2697, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1509696032.

Fernholz:2018:VA

[FKR18] E. Robert Fernholz, Ioannis Karatzas, and Johannes Ruf. Volatility and arbitrage. Annals of Applied Probability, 28 (1):378–417, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1520046091.

Foss:1996:SPS

[FL96] Serguei Foss and Günter Last. Stability of polling systems with exhaustive service policies and state-dependent routing. Annals of Applied Probability, 6(1):116–137, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1034968068.

Frydman:2003:MLE

[FL03] Halina Frydman and Peter Lakner. Maximum likelihood estimation of hidden Markov processes. Annals of Applied Probability, 13(4):1296–1312, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1069786500.

Foo:2013:DCR

[FL13] Jasmine Foo and Kevin Leder. Dynamics of cancer recurrence. Annals of Applied Probability, 23(4):1437–1468, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1371834034.

Fang:2022:CCS

[FL22] Rongjuan Fang and Zenghu Li. Construction of continuous-state branching processes in varying environments. Annals of Applied Probability, 32(5):3645–3673, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Construction-of-continuous-state-branching-
processes-in-varying-environments/10.1214/21-AAP1770.
full.

Flatto:1997:WTD

[Fla97] L. Flatto. The waiting time distribution for the random order service M/M/1 queue. Annals of Applied Probability, 7(2):382– 409, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034625337.

Figueroa-Lopez:2019:CPD

[FLÓ19] José E. Figueroa-López and Sveinn Ólafsson. Change-point detection for Lévy processes. Annals of Applied Probability, 29(2):717–738, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1548298928.

Favaro:2013:CFG

[FLP13] Stefano Favaro, Antonio Lijoi, and Igor Prünster. Conditional formulae for Gibbs-type exchangeable random partitions. Annals of Applied Probability, 23(5):1721–1754, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1377696296.

Fan:2022:IRP

[FLR22a] Wai-Tong Louis Fan, Brandon Legried, and Sebastien Roch. An impossibility result for phylogeny reconstruction from k-mer counts. Annals of Applied Probability, 32(6):4893–4913, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/An-impossibility-
result-for-phylogeny-reconstruction-from-k-mer-counts/
10.1214/22-AAP1805.full.

Fredes:2022:CPM

[FLR22b] Luis Fredes, Amitai Linker, and Daniel Remenik. Coexistence for a population model with forest fire epidemics. Annals of Applied Probability, 32(5):4004–4037, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Coexistence-for-a-population-model-
with-forest-fire-epidemics/10.1214/22-AAP1780.full.

Fouque:1994:LTL

[FM94] Jean-Pierre Fouque and Ely Merzbach. A limit theorem for linear boundary value problems in random media. Annals of Applied Probability, 4(2):549–569, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005072.

Foley:2001:JSQ

[FM01] R. D. Foley and D. R. McDonald. Join the shortest queue: stability and exact asymptotics. Annals of Applied Probability, 11(3):569–607, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015345342.

Fournier:2004:MPD

[FM04] Nicolas Fournier and Sylvie Méléard. A microscopic probabilistic description of a locally regulated population and macroscopic approximations. Annals of Applied Probability, 14(4):1880–1919, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1099674082.

Foley:2005:BNE

[FM05a] Robert D. Foley and David R. McDonald. Bridges and networks: Exact asymptotics. Annals of Applied Probability, 15(1B):542– 586, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1107271660.

Foley:2005:LDM

[FM05b] Robert D. Foley and David R. McDonald. Large deviations of a modified Jackson network: Stability and rough asymptotics. Annals of Applied Probability, 15(1B):519–541, February 2005.

CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1107271659.

Fathi:2016:ERC

[FM16] Max Fathi and Jan Maas. Entropic Ricci curvature bounds for discrete interacting systems. Annals of Applied Probability, 26(3):1774–1806, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1465905019.

Fountoulakis:2018:LLN

[FM18] Nikolaos Fountoulakis and Tobias Müller. Law of large numbers for the largest component in a hyperbolic model of complex networks. Annals of Applied Probability, 28(1):607–650, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1520046096.

Fort:2008:OMS

[FMMP08] Gersende Fort, Sean Meyn, Eric Moulines, and Pierre Priouret. The ODE method for stability of skip-free Markov chains with applications to MCMC. Annals of Applied Probability, 18 (2):664–707, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1206018201.

Fernandez:2016:CQS

[FMN^{+}16] R. Fernandez, F. Manzo, F. R. Nardi, E. Scoppola, and J. Sohier. Conditioned, quasi-stationary, restricted measures and escape from metastable states. Annals of Applied Probability, 26(2):760–793, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1458651819.

Francos:1995:WLD

[FMP95] Joseph M. Francos, A. Zvi Meiri, and Boaz Porat. A Wold-like decomposition of two-dimensional discrete homogeneous random fields. Annals of Applied Probability, 5(1):248–260, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004839.

Fontes:2000:ADR

[FMP00] Luiz Renato Fontes, Pierre Mathieu, and Pierre Picco. On the averaged dynamics of the random field Curie–Weiss model. Annals of Applied Probability, 10(4):1212–1245, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1019487614.

Ferrari:2009:PTC

[FMP09] Pablo A. Ferrari, James B. Martin, and Leandro P. R. Pimentel. A phase transition for competition interfaces. Annals of Applied Probability, 19(1):281–317, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1235140340.

Fill:1996:DDR

[FMS96] James Allen Fill, Hosam M. Mahmoud, and Wojciech Szpankowski. On the distribution for the duration of a randomized leader election algorithm. Annals of Applied Probability, 6 (4):1260–1283, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1035463332.

Foss:2014:LRL

[FMS14] Sergey Foss, James B. Martin, and Philipp Schmidt. Longrange last-passage percolation on the line. Annals of Applied Probability, 24(1):198–234, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278724.

Fontes:1993:FPP

[FN93] Luiz Fontes and Charles M. Newman. First passage percolation for random colorings of Z^{d}. Annals of Applied Probability, 3(3):746–762, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005361. See correction [FN94].

Fontes:1994:CFP

[FN94] L. R. G. Fontes and Charles M. Newman. Correction: “First Passage Percolation for Random Colorings of Z^{d}”. Annals of Applied Probability, 4(1):254, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005211. See [FN93].

Franco:2017:LDE

[FN17] Tertuliano Franco and Adriana Neumann. Large deviations for the exclusion process with a slow bond. Annals of Applied Probability, 27(6):3547–3587, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1513328708.

Fontbona:2010:SVM

[Fon10] J. Fontbona. Stochastic vortex method for forced threedimensional Navier–Stokes equations and pathwise convergence rate. Annals of Applied Probability, 20(5):1761–1800, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1282747400.

Fournier:2000:ERS

[Fou00] Nicolas Fournier. Existence and regularity study for twodimensional Kac equation without cutoff by a probabilistic approach. Annals of Applied Probability, 10(2):434–462, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1019487350.

Fournier:2015:FEH

[Fou15] Nicolas Fournier. Finiteness of entropy for the homogeneous Boltzmann equation with measure initial condition. Annals of Applied Probability, 25(2):860–897, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1424355132.

Foucart:2022:PTC

[Fou22] Clément Foucart. A phase transition in the coming down from infinity of simple exchangeable fragmentation-coagulation processes. Annals of Applied Probability, 32(1):632–664, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/A-phase-
transition-in-the-coming-down-from-infinity-of/10.1214/
21-AAP1691.full.

Foxall:2016:CBP

[Fox16] Eric Foxall. Critical behaviour of the partner model. Annals of Applied Probability, 26(5):2824–2859, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1476884305.

Fill:1993:PFP

[FP93] James Allen Fill and Robin Pemantle. Percolation, first-passage percolation and covering times for Richardson’s model on the n-cube. Annals of Applied Probability, 3(2):593–629, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005440.

Fort:1995:CKA

[FP95a] Jean-Claude Fort and Gilles Pages. On the A.S. convergence of the Kohonen algorithm with a general neighborhood function. Annals of Applied Probability, 5(4):1177–1216, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177004611.

Frieze:1995:PAA

[FP95b] Alan Frieze and Boris G. Pittel. Probabilistic analysis of an algorithm in the theory of markets in indivisible goods. Annals of Applied Probability, 5(3):768–808, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004705.

Fuhrman:2015:RBS

[FP15] Marco Fuhrman and Huyên Pham. Randomized and backward SDE representation for optimal control of non-Markovian SDEs. Annals of Applied Probability, 25(4):2134–2167, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1432212439.

Frieze:2017:LVN

[FP17] Alan Frieze and Wesley Pegden. Looking for vertex number one. Annals of Applied Probability, 27(1):582–630, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1488790836.

Frieze:2018:DLA

[FP18] Alan Frieze and Wesley Pegden. Diffusion limited aggregation on the Boolean lattice. Annals of Applied Probability, 28(6): 3528–3557, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1538985628.

Forien:2022:HEM

[FP22] Raphaël Forien and Étienne Pardoux. Household epidemic models and McKean–Vlasov Poisson driven stochastic differential equations. Annals of Applied Probability, 32(2):1210–1233, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Household-
epidemic-models-and-McKeanVlasov--Poisson-driven-stochastic-
differential-equations/10.1214/21-AAP1706.full.

Forman:2018:UCL

[FPRW18] Noah Forman, Soumik Pal, Douglas Rizzolo, and Matthias Winkel. Uniform control of local times of spectrally positive stable processes. Annals of Applied Probability, 28(4):2592–2634, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1533780281.

Foss:2005:PEH

[FPZ05] Serguei Foss, Zbigniew Palmowski, and Stan Zachary. The probability of exceeding a high boundary on a random time interval for a heavy-tailed random walk. Annals of Applied Probability, 15(3):1936–1957, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433773.

Ferrari:1992:SAE

[FR92] P. A. Ferrari and K. Ravishankar. Shocks in asymmetric exclusion automata. Annals of Applied Probability, 2(4):928–941, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005581.

Fort:2005:SES

[FR05] G. Fort and G. O. Roberts. Subgeometric ergodicity of strong Markov processes. Annals of Applied Probability, 15(2):1565– 1589, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1115137986.

Faller:2011:ASM

[FR11] Andreas Faller and Ludger Rüschendorf. On approximative solutions of multistopping problems. Annals of Applied Probability, 21(5):1965–1993, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576614.

Francois:2002:GOE

[Fra02] Olivier François. Global optimization with exploration/selection algorithms and simulated annealing. Annals of Applied Probability, 12(1):248–271, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015961163.

Frieze:1991:LLM

[Fri91] Alan Frieze. On the length of the longest monotone subsequence in a random permutation. Annals of Applied Probability, 1(2): 301–305, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005939.

Frikha:2016:MLS

[Fri16] Noufel Frikha. Multi-level stochastic approximation algorithms. Annals of Applied Probability, 26(2):933–985, April 2016. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1458651825.

Foutel-Rodier:2021:ECU

[FRLS21] Félix Foutel-Rodier, Amaury Lambert, and Emmanuel Schertzer. Exchangeable coalescents, ultrametric spaces, nested intervalpartitions: A unifying approach. Annals of Applied Probability, 31(5):2046–2090, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-5/Exchangeable-coalescents-ultrametric-spaces-
nested-interval-partitions--A-unifying/10.1214/20-AAP1641.
full.

Fricker:1994:ASN

[FRST94] Christine Fricker, Philippe Robert, Ellen Saada, and Danielle Tibi. Analysis of some networks with interaction. Annals of Applied Probability, 4(4):1112–1128, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177004906.

Forman:2022:TPF

[FRSW22] Noah Forman, Douglas Rizzolo, Quan Shi, and Matthias Winkel. A two-parameter family of measure-valued diffusions with Poisson–Dirichlet stationary distributions. Annals of Applied Probability, 32(3):2211–2253, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-3/A-two-parameter-family-of-measure-
valued-diffusions-with-PoissonDirichlet/10.1214/21-AAP1732.
full.

Fricker:2003:DCL

[FRT03] Christine Fricker, Philippe Robert, and Danielle Tibi. A degenerate central limit theorem for single resource loss systems. Annals of Applied Probability, 13(2):561–575, May 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1050689594.

Faggionato:2014:UOD

[FRT14] A. Faggionato, C. Roberto, and C. Toninelli. Universality for one-dimensional hierarchical coalescence processes with double and triple merges. Annals of Applied Probability, 24(2):476–525, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1394465363.

Friedland:2004:CPE

[FRZ04] Shmuel Friedland, Brian Rider, and Ofer Zeitouni. Concentration of permanent estimators for certain large matrices. Annals of Applied Probability, 14(3):1559–1576, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1089736296.

Fleming:1999:OLT

[FS99] Wendell H. Fleming and Shuenn-Jyi Sheu. Optimal long term growth rate of expected utility of wealth. Annals of Applied Probability, 9(3):871–903, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1029962817.

Fleming:2002:RSC

[FS02] W. H. Fleming and S. J. Sheu. Risk-sensitive control and an optimal investment model II. Annals of Applied Probability, 12(2):

730–767, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1026915623.

Faure:2014:QSD

[FS14] Mathieu Faure and Sebastian J. Schreiber. Quasi-stationary distributions for randomly perturbed dynamical systems. Annals of Applied Probability, 24(2):553–598, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1394465365.

Fang:2016:PAT

[FS16] Xiao Fang and David Siegmund. Poisson approximation for two scan statistics with rates of convergence. Annals of Applied Probability, 26(4):2384–2418, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1472745462.

Finn:2022:CCF

[FS22] Thomas Finn and Alexandre Stauffer. Coexistence in competing first passage percolation with conversion. Annals of Applied Probability, 32(6):4459–4480, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-6/Coexistence-in-competing-first-passage-
percolation-with-conversion/10.1214/22-AAP1792.full.

Funaki:1995:GCR

[FSW95] T. Funaki, D. Surgailis, and W. A. Woyczynski. Gibbs–Cox random fields and Burgers turbulence. Annals of Applied Probability, 5(2):461–492, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004774.

Ferrari:2015:SLB

[FSW15] Patrik L. Ferrari, Herbert Spohn, and Thomas Weiss. Scaling limit for Brownian motions with one-sided collisions. Annals of Applied Probability, 25(3):1349–1382, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1427124131.

Fricker:2017:EEL

[FT17] Christine Fricker and Danielle Tibi. Equivalence of ensembles for large vehicle-sharing models. Annals of Applied Probability, 27(2):883–916, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1495764369.

Fahim:2011:PNM

[FTW11] Arash Fahim, Nizar Touzi, and Xavier Warin. A probabilistic numerical method for fully nonlinear parabolic PDEs. Annals of Applied Probability, 21(4):1322–1364, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1312818838.

Fuh:2004:UMR

[Fuh04] Cheng-Der Fuh. Uniform Markov renewal theory and ruin probabilities in Markov random walks. Annals of Applied Probability, 14(3):1202–1241, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1089736283.

Fukasawa:2011:DES

[Fuk11] Masaaki Fukasawa. Discretization error of stochastic integrals. Annals of Applied Probability, 21(4):1436–1465, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1312818841.

Fusai:2000:COA

[Fus00] Gianluca Fusai. Corridor options and arc-sine law. Annals of Applied Probability, 10(2):634–663, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487359.

Ferrari:2021:FAC

[FV21] Patrik L. Ferrari and Bálint Vető. Fluctuations of the Arctic curve in the tilings of the Aztec diamond on restricted domains. Annals of Applied Probability, 31(1):284–320, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/Fluctuations-
of-the-arctic-curve-in-the-tilings-of-the/10.1214/20-
AAP1590.full.

Freidlin:1999:RRP

[FW99a] Mark Freidlin and Matthias Weber. A remark on random perturbations of the nonlinear pendulum. Annals of Applied Probability, 9(3):611–628, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962806.

Fulman:1999:CEB

[FW99b] Jason Fulman and Elizabeth L. Wilmer. Comparing eigenvalue bounds for Markov chains: when does Poincaré beat Cheeger? Annals of Applied Probability, 9(1):1–13, February 1999. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1029962594.

Fu:2022:ABC

[FW22] Yingjia Fu and Ruth J. Williams. Asymptotic behavior of a critical fluid model for bandwidth sharing with general file size distributions. Annals of Applied Probability, 32(3):1862–1901, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Asymptotic-
behavior-of-a-critical-fluid-model-for-bandwidth-sharing/
10.1214/21-AAP1723.full.

Filipovic:2002:MTS

[FZ02] Damir Filipović and Jerzy Zabczyk. Markovian term structure models in discrete time. Annals of Applied Probability, 12(2): 710–729, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1026915622.

Foss:2003:MRT

[FZ03] Serguei Foss and Stan Zachary. The maximum on a random time interval of a random walk with long-tailed increments and negative drift. Annals of Applied Probability, 13(1):37–53, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1042765662.

Gapeev:2005:DPC

[Gap05] Pavel V. Gapeev. The disorder problem for compound Poisson processes with exponential jumps. Annals of Applied Probability, 15(1A):487–499, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1106922334.

Garet:2009:CFR

[Gar09] Olivier Garet. Capacitive flows on a 2d random net. Annals of Applied Probability, 19(2):641–660, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1241702245.

Gaudron:1998:SLC

[Gau98] Guillaume Gaudron. Scaling laws and convergence for the advection-diffusion equation. Annals of Applied Probability, 8 (3):649–663, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903445.

Greven:1993:VCS

[GdH93] Andreas Greven and Frank den Hollander. A variational characterization of the speed of a one-dimensional self- repellent random walk. Annals of Applied Probability, 3(4):1067–1099, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005273.

Gorham:2019:MSQ

[GDVM19] Jackson Gorham, Andrew B. Duncan, Sebastian J. Vollmer, and Lester Mackey. Measuring sample quality with diffusions. Annals of Applied Probability, 29(5):2884–2928, October 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1571385625.

Genovese:2023:MFR

[Gen23] Giuseppe Genovese. Minimax formula for the replica symmetric free energy of deep restricted Boltzmann machines. Annals of Applied Probability, 33(3):2324–2341, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-3/Minimax-formula-for-the-replica-symmetric-
free-energy-of-deep/10.1214/22-AAP1868.full.

Gerhold:2011:LSA

[Ger11] Stefan Gerhold. The Longstaff–Schwartz algorithm for Lévy models: Results on fast and slow convergence. Annals of Applied Probability, 21(2):589–608, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1300800982.

Gravner:1997:MTG

[GG97] Janko Gravner and David Griffeath. Multitype threshold growth: convergence to Poisson–Voronoi tessellations. Annals of Applied Probability, 7(3):615–647, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034801246.

Graham:2011:STR

[GG11] Benjamin Graham and Geoffrey Grimmett. Sharp thresholds for the random-cluster and Ising models. Annals of Applied Probability, 21(1):240–265, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1292598033.

Gamarnik:2013:RCS

[GG13a] David Gamarnik and David A. Goldberg. On the rate of convergence to stationarity of the M/M/N queue in the Halfin–Whitt regime. Annals of Applied Probability, 23(5):1879–1912, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1377696301.

Gamarnik:2013:SSG

[GG13b] David Gamarnik and David A. Goldberg. Steady-state GI/G/n queue in the Halfin–Whitt regime. Annals of Applied Probability, 23(6):2382–2419, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1382447692.

Galves:2013:KTD

[GGLO13] A. Galves, N. L. Garcia, E. Löcherbach, and E. Orlandi. Kalikow-type decomposition for multicolor infinite range particle systems. Annals of Applied Probability, 23(4):1629–1659, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1371834040.

Gonon:2023:ABR

[GGO23] Lukas Gonon, Lyudmila Grigoryeva, and Juan-Pablo Ortega. Approximation bounds for random neural networks and reservoir systems. Annals of Applied Probability, 33(1):28–69, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Approximation-
bounds-for-random-neural-networks-and-reservoir-systems/
10.1214/22-AAP1806.full.

Gelman:1997:WCO

[GGR97] A. Gelman, W. R. Gilks, and G. O. Roberts. Weak convergence and optimal scaling of random walk Metropolis algorithms. Annals of Applied Probability, 7(1):110–120, February 1997. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1034625254.

Gaier:2003:ARP

[GGS03] J. Gaier, P. Grandits, and W. Schachermayer. Asymptotic ruin probabilities and optimal investment. Annals of Applied Probability, 13(3):1054–1076, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1060202834.

Gamarnik:2005:ISF

[GH05] David Gamarnik and John J. Hasenbein. Instability in stochastic and fluid queueing networks. Annals of Applied Probability, 15 (3):1652–1690, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433764.

Gravner:2008:SCB

[GH08] Janko Gravner and Alexander E. Holroyd. Slow convergence in bootstrap percolation. Annals of Applied Probability, 18(3):909– 928, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1211819789.

Gnewuch:2021:DBC

[GH21] Michael Gnewuch and Nils Hebbinghaus. Discrepancy bounds for a class of negatively dependent random points including Latin hypercube samples. Annals of Applied Probability, 31(4):1944– 1965, August 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
4/Discrepancy-bounds-for-a-class-of-negatively-dependent-
random-points/10.1214/20-AAP1638.full.

Galeati:2022:RMS

[GH22] Lucio Galeati and Fabian A. Harang. Regularization of multiplicative SDEs through additive noise. Annals of Applied Probability, 32(5):3930–3963, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Regularization-of-multiplicative-SDEs-
through-additive-noise/10.1214/21-AAP1778.full.

Git:2007:EGR

[GHH07] Y. Git, J. W. Harris, and S. C. Harris. Exponential growth rates in a typed branching diffusion. Annals of Applied Probability, 17(2):609–653, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1174323258.

Glazebrook:2011:GNI

[GHK11] Kevin D. Glazebrook, David J. Hodge, and Chris Kirkbride. General notions of indexability for queueing control and asset management. Annals of Applied Probability, 21(3):876–907, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1307020386.

Guo:2003:CTC

[GHL03] Xianping Guo and Onésimo Hernández-Lerma. Continuoustime controlled Markov chains. Annals of Applied Probability, 13(1):363–388, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1042765671.

Galichon:2014:SCA

[GHLT14] A. Galichon, P. Henry-Labordère, and N. Touzi. A stochastic control approach to no-arbitrage bounds given marginals, with an application to lookback options. Annals of Applied Probability, 24(1):312–336, February 2014. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278727.

Glover:2013:TDB

[GHP13] Kristoffer Glover, Hardy Hulley, and Goran Peskir. Threedimensional Brownian motion and the golden ratio rule. Annals of Applied Probability, 23(3):895–922, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1362684849.

Gravner:2015:BPH

[GHPS15] Janko Gravner, Christopher Hoffman, James Pfeiffer, and David Sivakoff. Bootstrap percolation on the Hamming torus. Annals of Applied Probability, 25(1):287–323, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1418740187.

Gajrat:2003:LDA

[GHR03] Alexander Gajrat, Arie Hordijk, and Ad Ridder. Largedeviations analysis of the fluid approximation for a controllable tandem queue. Annals of Applied Probability, 13(4):1423–1448, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1069786504.

Gravner:2021:PBP

[GHS21] Janko Gravner, Alexander E. Holroyd, and David Sivakoff. Polluted bootstrap percolation in three dimensions. Annals of Applied Probability, 31(1):218–246, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-1/Polluted-bootstrap-percolation-in-
three-dimensions/10.1214/20-AAP1588.full.

Gnedin:2018:CSC

[GIMM18] Alexander Gnedin, Alexander Iksanov, Alexander Marynych, and Martin Möhle. The collision spectrum of Λ-coalescents. Annals of Applied Probability, 28(6):3857–3883, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1538985637.

Gnedin:2009:BSR

[GINR09] Alexander V. Gnedin, Alexander M. Iksanov, Pavlo Negadajlov, and Uwe Rösler. The Bernoulli sieve revisited. Annals of Applied Probability, 19(4):1634–1655, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1248700630.

Grigorova:2017:RBW

[GIO^{+}17] Miryana Grigorova, Peter Imkeller, Elias Offen, Youssef Ouknine, and Marie-Claire Quenez. Reflected BSDEs when the obstacle is not right-continuous and optimal stopping. Annals of Applied Probability, 27(5):3153–3188, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1509696043.

Guo:2018:RCD

[GJ18] Heng Guo and Mark Jerrum. Random cluster dynamics for the Ising model is rapidly mixing. Annals of Applied Probability, 28(2):1292–1313, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1523433636.

Georgiou:2015:SDS

[GJKS15] Nicos Georgiou, Mathew Joseph, Davar Khoshnevisan, and Shang-Yuan Shiu. Semi-discrete semi-linear parabolic SPDEs. Annals of Applied Probability, 25(5):2959–3006, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1438261057.

Gandolfi:1994:GLA

[GK94] Alberto Gandolfi and Harry Kesten. Greedy lattice animals II: Linear growth. Annals of Applied Probability, 4(1):76–107, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005201.

Glasserman:1995:LFP

[GK95a] Paul Glasserman and Shing-Gang Kou. Limits of first passage times to rare sets in regenerative processes. Annals of Applied Probability, 5(2):424–445, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004772.

Gnedin:1995:SGO

[GK95b] Alexander V. Gnedin and Ulrich Krengel. A stochastic game of optimal stopping and order selection. Annals of Applied Probability, 5(1):310–321, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004842.

Gnedin:1996:OSP

[GK96] Alexander V. Gnedin and Ulrich Krengel. Optimal selection problems based on exchangeable trials. Annals of Applied Probability, 6(3):862–882, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968230.

Gartner:2000:MAC

[GK00] Jürgen Gärtner and Wolfgang König. Moment asymptotics for the continuous parabolic Anderson model. Annals of Applied Probability, 10(1):192–217, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1019737669.

Goll:2003:CES

[GK03] Thomas Goll and Jan Kallsen. A complete explicit solution to the log-optimal portfolio problem. Annals of Applied Probability, 13(2):774–799, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1050689603.

Gromoll:2007:HTL

[GK07a] H. Christian Gromoll and Łukasz Kruk. Heavy traffic limit for a processor sharing queue with soft deadlines. Annals of Applied Probability, 17(3):1049–1101, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1179839182.

Guan:2007:SWM

[GK07b] Yongtao Guan and Stephen M. Krone. Small-world MCMC and convergence to multi-modal distributions: From slow mixing to fast mixing. Annals of Applied Probability, 17(1):284–304, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1171377185.

Gamarnik:2009:DSM

[GK09] David Gamarnik and Dmitriy Katz. On deciding stability of multiclass queueing networks under buffer priority scheduling policies. Annals of Applied Probability, 19(5):2008–2037, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1255699551.

Grubel:2016:FCL

[GK16] Rudolf Grübel and Zakhar Kabluchko. A functional central limit theorem for branching random walks, almost sure weak convergence and applications to random trees. Annals of Applied Probability, 26(6):3659–3698, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1481792596.

Gamarnik:2021:CPF

[GK21] David Gamarnik and Eren C. Kizildag. Computing the partition function of the Sherrington–Kirkpatrick model is hard on average. Annals of Applied Probability, 31(3):1474–1504, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-3/Computing-
the-partition-function-of-the-SherringtonKirkpatrick-
model-is-hard/10.1214/20-AAP1625.full.

Gravner:2023:TCP

[GK23] Janko Gravner and Brett Kolesnik. Transitive closure in a polluted environment. Annals of Applied Probability, 33(1):107–126, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Transitive-
closure-in-a-polluted-environment/10.1214/22-AAP1810.
full.

Grigorescu:2004:BDS

[GKS04] Ilie Grigorescu, Min Kang, and Timo Seppäläinen. Behavior dominated by slow particles in a disordered asymmetric exclusion process. Annals of Applied Probability, 14(3):1577–1602, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1089736297.

Gouet:2001:SMP

[GL01] Raúl Gouet and F. Javier López. Saturation in a makovian parking process. Annals of Applied Probability, 11(4):1116–1136, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1015345397.

Gobet:2014:ASO

[GL14] Emmanuel Gobet and Nicolas Landon. Almost sure optimal hedging strategy. Annals of Applied Probability, 24(4):1652– 1690, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1400073660.

Giacomin:2018:DWT

[GL18] Giambattista Giacomin and Hubert Lacoin. Disorder and wetting transition: The pinned harmonic crystal in dimension three or larger. Annals of Applied Probability, 28(1):577–606, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1520046095.

Guo:2021:PDO

[GL21] Ivan Guo and Grégoire Loeper. Path dependent optimal transport and model calibration on exotic derivatives. Annals of Applied Probability, 31(3):1232–1263, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-3/Path-dependent-optimal-transport-and-
model-calibration-on-exotic-derivatives/10.1214/20-AAP1617.
full.

Guo:2023:SLM

[GL23] Shuchen Guo and Dejun Luo. Scaling limit of moderately interacting particle systems with singular interaction and environmental noise. Annals of Applied Probability, 33(3):2066–2102, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/Scaling-
limit-of-moderately-interacting-particle-systems-with-
singular-interaction/10.1214/22-AAP1860.full.

Glazebrook:1993:IFC

[Gla93] K. D. Glazebrook. Indices for families of competing Markov decision processes with influence. Annals of Applied Probability, 3(4):1013–1032, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005270.

Goncalves:2017:NFO

[GLM17] Patrícia Gonçalves, Claudio Landim, and Aniura Milanés. Nonequilibrium fluctuations of one-dimensional boundary driven weakly asymmetric exclusion processes. Annals of Applied Probability, 27(1):140–177, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1488790825.

Gulikers:2018:IRR

[GLM18] Lennart Gulikers, Marc Lelarge, and Laurent Massoulié. An impossibility result for reconstruction in the degree-corrected stochastic block model. Annals of Applied Probability, 28(5): 3002–3027, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1535443240.

Grama:2023:KST

[GLP23] Ion Grama, Quansheng Liu, and Erwan Pin. A Kesten-stigum type theorem for a supercritical multitype branching process in a random environment. Annals of Applied Probability, 33(2): 1013–1051, April 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
2/A-KestenStigum-type-theorem-for-a-supercritical-multitype-
branching-process/10.1214/22-AAP1840.full.

Gravner:2018:LBC

[GLS18] Janko Gravner, Hanbaek Lyu, and David Sivakoff. Limiting behavior of 3-color excitable media on arbitrary graphs. Annals of Applied Probability, 28(6):3324–3357, December 2018. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1538985624.

Gobet:2005:RBM

[GLW05] Emmanuel Gobet, Jean-Philippe Lemor, and Xavier Warin. A regression-based Monte Carlo method to solve backward stochastic differential equations. Annals of Applied Probability, 15 (3):2172–2202, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433781.

Guillin:2022:UPL

[GLWZ22] Arnaud Guillin, Wei Liu, Liming Wu, and Chaoen Zhang. Uniform Poincaré and logarithmic Sobolev inequalities for mean field particle systems. Annals of Applied Probability, 32(3): 1590–1614, June 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
3/Uniform-Poincar%c3%a9-and-logarithmic-Sobolev-inequalities-
for-mean-field-particle/10.1214/21-AAP1707.full.

Graham:1995:DAR

[GM95] Carl Graham and Sylvie Meleard. Dynamic asymptotic results for a generalized star-shaped loss network. Annals of Applied Probability, 5(3):666–680, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004700.

Garet:2005:CTT

[GM05] Olivier Garet and Régine Marchand. Coexistence in two-type first-passage percolation models. Annals of Applied Probability, 15(1A):298–330, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1106922330.

Gouere:2008:CFP

[GM08] Jean-Baptiste Gouéré and Régine Marchand. Continuous firstpassage percolation and continuous greedy paths model: Linear growth. Annals of Applied Probability, 18(6):2300–2319, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1227708919.

Garet:2012:ASC

[GM12a] Olivier Garet and Régine Marchand. Asymptotic shape for the contact process in random environment. Annals of Applied Probability, 22(4):1362–1410, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1344614198.

Griffin:2012:PDR

[GM12b] Philip S. Griffin and Ross A. Maller. Path decomposition of ruinous behavior for a general Lévy insurance risk process. Annals of Applied Probability, 22(4):1411–1449, August 2012. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1344614199.

Gloria:2013:QVK

[GM13] Antoine Gloria and Jean-Christophe Mourrat. Quantitative version of the Kipnis–Varadhan theorem and Monte Carlo approximation of homogenized coefficients. Annals of Applied Probability, 23(4):1544–1583, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1371834038.

Gangopadhyay:2019:SAR

[GM19] Ujan Gangopadhyay and Krishanu Maulik. Stochastic approximation with random step sizes and urn models with random replacement matrices having finite mean. Annals of Applied Probability, 29(4):2033–2066, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1563869036.

Gassiat:2015:IER

[GMO15] Paul Gassiat, Aleksandar Mijatović, and Harald Oberhauser. An integral equation for Root’s barrier and the generation of Brownian increments. Annals of Applied Probability, 25(4):2039–2065, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1432212436.

Gong:2019:SRS

[GMŚ19] Ruoting Gong, Chenchen Mou, and Andrzej Świech. Stochastic representations for solutions to parabolic Dirichlet problems for nonlocal Bellman equations. Annals of Applied Probability, 29 (6):3271–3310, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1578366315.

Glaz:1991:TBA

[GN91] Joseph Glaz and Joseph I. Naus. Tight bounds and approximations for scan statistic probabilities for discrete data. Annals of Applied Probability, 1(2):306–318, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005940.

Goldstein:2017:GPT

[GNP17] Larry Goldstein, Ivan Nourdin, and Giovanni Peccati. Gaussian phase transitions and conic intrinsic volumes: Steining the Steiner formula. Annals of Applied Probability, 27(1):1–47, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1488790820.

Gantert:2023:MTS

[GNS23] Nina Gantert, Evita Nestoridi, and Dominik Schmid. Mixing times for the simple exclusion process with open boundaries. Annals of Applied Probability, 33(2):972–1012, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Mixing-
times-for-the-simple-exclusion-process-with-open-boundaries/
10.1214/22-AAP1839.full.

Ganesh:1998:LGP

[GO98] A. J. Ganesh and Neil O’Connell. The linear geodesic property is not generally preserved by a FIFO queue. Annals of Applied Probability, 8(1):98–111, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1027961035.

Graham:2000:LDE

[GO00] Carl Graham and Neil O’Connell. Large deviations at equilibrium for a large star-shaped loss network. Annals of Applied Probability, 10(1):104–122, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1019737666.

Gloria:2012:OEE

[GO12] Antoine Gloria and Felix Otto. An optimal error estimate in stochastic homogenization of discrete elliptic equations. Annals of Applied Probability, 22(1):1–28, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1328623694.

Guo:2019:CMM

[GO19] Gaoyue Guo and Jan Obłój. Computational methods for martingale optimal transport problems. Annals of Applied Probability, 29(6):3311–3347, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1578366316.

Goel:2006:ATB

[Goe06a] Sharad Goel. Analysis of top to bottom- k shuffles. Annals of Applied Probability, 16(1):30–55, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654280.

Goergen:2006:LVZ

[Goe06b] Laurent Goergen. Limit velocity and zero–one laws for diffusions in random environment. Annals of Applied Probability, 16 (3):1086–1123, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1159804976.

Goldie:1991:IRT

[Gol91] Charles M. Goldie. Implicit renewal theory and tails of solutions of random equations. Annals of Applied Probability, 1(1):126– 166, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005985.

Goldstein:2004:NAH

[Gol04] Larry Goldstein. Normal approximation for hierarchical structures. Annals of Applied Probability, 14(4):1950–1969, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1099674084.

Goldman:2010:PMV

[Gol10] André Goldman. The Palm measure and the Voronoi tessellation for the Ginibre process. Annals of Applied Probability, 20(1):90– 128, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1262962319.

Goldstein:2013:BEB

[Gol13] Larry Goldstein. A Berry–Esseen bound with applications to vertex degree counts in the Erdős–Rényi random graph. Annals of Applied Probability, 23(2):617–636, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1360682024.

Goldberg:2016:SOM

[Gol16] David A. Goldberg. Second-order Markov random fields for independent sets on the infinite Cayley tree. Annals of Applied Probability, 26(5):2626–2660, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1476884299.

Ganesh:2003:IRF

[GOP03] Ayalvadi Ganesh, Neil O’Connell, and Balaji Prabhakar. Invariant rate functions for discrete-time queues. Annals of Applied Probability, 13(2):446–474, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1050689588.

Gosselin:2001:ABA

[Gos01] Frèdèric Gosselin. Aysmptotic behavior of absorbing Markov chains conditional on nonabsorption for applications in conservation biology. Annals of Applied Probability, 11(1):261–284, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/998926993.

Gouere:2007:STS

[Gou07] Jean-Baptiste Gouéré. Shape of territories in some competing growth models. Annals of Applied Probability, 17(4):1273–1305, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1186755240.

Gouere:2009:SRS

[Gou09] Jean-Baptiste Gouéré. Subcritical regimes in some models of continuum percolation. Annals of Applied Probability, 19(4): 1292–1318, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1248700618.

Goldstein:2010:NAC

[GP10] Larry Goldstein and Mathew D. Penrose. Normal approximation for coverage models over binomial point processes. Annals of Applied Probability, 20(2):696–721, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1268143437.

Gaunt:2017:CSA

[GPR17] Robert E. Gaunt, Alastair M. Pickett, and Gesine Reinert. Chisquare approximation by Stein’s method with application to Pearson’s statistic. Annals of Applied Probability, 27(2):720–756, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1495764365.

Gromoll:2002:FLH

[GPW02] H. Christian Gromoll, Amber L. Puha, and Ruth J. Williams. The fluid limit of a heavily loaded processor sharing queue. Annals of Applied Probability, 12(3):797–859, August 2002. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1031863171.

Greven:2009:GCB

[GPW09] Andreas Greven, Lea Popovic, and Anita Winter. Genealogy of catalytic branching models. Annals of Applied Probability, 19(3):1232–1272, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1245071025.

Gao:2003:EDE

[GQ03] Fuqing Gao and Jeremy Quastel. Exponential decay of entropy in the random transposition and Bernoulli–Laplace models. Annals of Applied Probability, 13(4):1591–1600, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1069786512.

Goldstein:1997:SMZ

[GR97] Larry Goldstein and Gesine Reinert. Stein’s method and the zero bias transformation with application to simple random sampling. Annals of Applied Probability, 7(4):935–952, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1043862419.

Guo:2006:AOC

[GR06] Xianping Guo and Ulrich Rieder. Average optimality for continuous-time Markov decision processes in Polish spaces. Annals of Applied Probability, 16(2):730–756, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1151592249.

Gabetta:2008:CLT

[GR08] Ester Gabetta and Eugenio Regazzini. Central limit theorem for the solution of the Kac equation. Annals of Applied Probability, 18(6):2320–2336, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1227708920.

Graham:2009:IMC

[GR09] Carl Graham and Philippe Robert. Interacting multi-class transmissions in large stochastic networks. Annals of Applied Probability, 19(6):2334–2361, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1259158774.

Guasoni:2012:PRP

[GR12] Paolo Guasoni and Scott Robertson. Portfolios and risk premia for the long run. Annals of Applied Probability, 22(1):239–284, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1328623700.

Guasoni:2015:HAO

[GR15] Paolo Guasoni and Miklós Rásonyi. Hedging, arbitrage and optimality with superlinear frictions. Annals of Applied Probability, 25(4):2066–2095, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1432212437.

Gan:2021:SMP

[GR21] Han L. Gan and Nathan Ross. Stein’s method for the Poisson– Dirichlet distribution and the Ewens sampling formula, with applications to Wright–Fisher models. Annals of Applied Probability, 31(2):625–667, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/

issue-2/Steins-method-for-the-PoissonDirichlet-distribution-
and-the-Ewens-sampling/10.1214/20-AAP1600.full.

Graham:2009:RRM

[Gra09] Benjamin T. Graham. Rate of relaxation for a mean-field zerorange process. Annals of Applied Probability, 19(2):497–520, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1241702239.

Graham:2021:RPL

[Gra21] Carl Graham. Regenerative properties of the linear Hawkes process with unbounded memory. Annals of Applied Probability, 31 (6):2844–2863, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-6/Regenerative-properties-of-the-linear-Hawkes-
process-with-unbounded-memory/10.1214/21-AAP1664.full.

Grey:1994:RVT

[Gre94] D. R. Grey. Regular variation in the tail behaviour of solutions of random difference equations. Annals of Applied Probability, 4(1):169–183, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005205.

Griffin:2013:CEL

[Gri13] Philip S. Griffin. Convolution equivalent Lévy processes and first passage times. Annals of Applied Probability, 23(4):1506–1543, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1371834037.

Griffin:2016:SPB

[Gri16] Philip S. Griffin. Sample path behavior of a Lévy insurance risk process approaching ruin, under the Cramér–Lundberg and convolution equivalent conditions. Annals of Applied Probability, 26(1):360–401, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1452003242.

Goel:2005:MPR

[GRK05] Ashish Goel, Sanatan Rai, and Bhaskar Krishnamachari. Monotone properties of random geometric graphs have sharp thresholds. Annals of Applied Probability, 15(4):2535–2552, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1133965771.

Gromoll:2004:DAP

[Gro04] H. Christian Gromoll. Diffusion approximation for a processor sharing queue in heavy traffic. Annals of Applied Probability, 14(2):555–611, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1082737104.

Giraitis:2000:MLM

[GRS00] Liudas Giraitis, Peter M. Robinson, and Donatas Surgailis. A model for long memory conditional heteroscedasticity. Annals of Applied Probability, 10(3):1002–1024, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1019487516.

Gangopadhyay:2004:ROT

[GRS04] Sreela Gangopadhyay, Rahul Roy, and Anish Sarkar. Random oriented trees: a model of drainage networks. Annals of Applied Probability, 14(3):1242–1266, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1089736284.

Guasoni:2008:CPS

[GRS08] Paolo Guasoni, Miklós Rásonyi, and Walter Schachermayer. Consistent price systems and face-lifting pricing under transaction costs. Annals of Applied Probability, 18(2):491–520, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1206018195.

Gimbert:2016:VRG

[GRS^{+}16] Hugo Gimbert, Jérôme Renault, Sylvain Sorin, Xavier Venel, and Wieslaw Zielonka. On values of repeated games with signals. Annals of Applied Probability, 26(1):402–424, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1452003243.

Grubel:2009:SBS

[Grü09] Rudolf Grübel. On the silhouette of binary search trees. Annals of Applied Probability, 19(5):1781–1802, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1255699543.

Grubel:2014:STM

[Grü14] Rudolf Grübel. Search trees: Metric aspects and strong limit theorems. Annals of Applied Probability, 24(3):1269–1297, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1398258101.

Guillemin:2004:AAE

[GRZ04] Fabrice Guillemin, Philippe Robert, and Bert Zwart. AIMD algorithms and exponential functionals. Annals of Applied Probability, 14(1):90–117, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1075828048.

Grote:2002:AEF

[GS02] Mark N. Grote and Terence P. Speed. Approximate Ewens formulae for symmetric overdominance selection. Annals of Applied Probability, 12(2):637–663, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1026915619.

Grubel:2005:MPA

[GS05] Rudolf Grübel and Nikolče Stefanoski. Mixed Poisson approximation of node depth distributions in random binary search trees. Annals of Applied Probability, 15(1A):279–297, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1106922329.

Garnier:2009:CPW

[GS09] Josselin Garnier and Knut Sølna. Coupled paraxial wave equations in random media in the white-noise regime. Annals of Applied Probability, 19(1):318–346, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1235140341.

Guo:2011:DCT

[GS11] Xianping Guo and Xinyuan Song. Discounted continuous-time constrained Markov decision processes in Polish spaces. Annals of Applied Probability, 21(5):2016–2049, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1319576616.

Giles:2014:AMM

[GS14] Michael B. Giles and Lukasz Szpruch. Antithetic multilevel Monte Carlo estimation for multi-dimensional SDEs without Lévy area simulation. Annals of Applied Probability, 24(4):1585– 1620, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1400073658.

Gravner:2017:NSJ

[GS17] Janko Gravner and David Sivakoff. Nucleation scaling in jigsaw percolation. Annals of Applied Probability, 27(1):395–438, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1488790831.

Goldsheid:2018:REN

[GS18] Ilya Goldsheid and Sasha Sodin. Real eigenvalues in the non-Hermitian Anderson model. Annals of Applied Probability, 28 (5):3075–3093, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1535443242.

Gracar:2019:MSL

[GS19] Peter Gracar and Alexandre Stauffer. Multi-scale Lipschitz percolation of increasing events for Poisson random walks. Annals of Applied Probability, 29(1):376–433, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1544000432.

Gravner:2020:BPP

[GS20] Janko Gravner and David Sivakoff. Bootstrap percolation on the product of the two-dimensional lattice with a Hamming square. Annals of Applied Probability, 30(1):145–174, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1582621221.

Goldschmidt:2023:SLC

[GS23] Christina Goldschmidt and Robin Stephenson. The scaling limit of a critical random directed graph. Annals of Applied Probability, 33(3):2024–2065, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-3/The-scaling-limit-of-a-critical-random-directed-
graph/10.1214/22-AAP1859.full.

Giesecke:2013:DCL

[GSS13] Kay Giesecke, Konstantinos Spiliopoulos, and Richard B. Sowers. Default clustering in large portfolios: Typical events. Annals of Applied Probability, 23(1):348–385, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1359124389.

Grimmett:1998:DRG

[GSvdB98] Geoffrey R. Grimmett, Rinaldo B. Schinazi, and J. van den Berg. Dependent random graphs and spatial epidemics. Annals of Applied Probability, 8(2):317–336, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1028903529.

Griffiths:1999:AMG

[GT99] R. C. Griffiths and Simon Tavaré. The ages of mutations in gene trees. Annals of Applied Probability, 9(3):567–590, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1029962804.

Gu:2019:ALW

[GT19] Yu Gu and Li-Cheng Tsai. Another look into the Wong– Zakai theorem for stochastic heat equation. Annals of Applied Probability, 29(5):3037–3061, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1571385629.

Guo:2022:MDR

[GTW22] Chengkun Guo, Si Tang, and Ningxi Wei. On the minimal drift for recurrence in the frog model on d-ary trees. Annals of Applied Probability, 32(4):3004–3026, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-4/On-the-minimal-drift-for-recurrence-
in-the-frog-model/10.1214/21-AAP1755.full.

Gamarnik:2020:LBQ

[GTZ20] David Gamarnik, John N. Tsitsiklis, and Martin Zubeldia. A lower bound on the queueing delay in resource constrained load balancing. Annals of Applied Probability, 30(2):870–901, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1591603225.

Gu:2022:EAS

[Gu22] Chenlin Gu. An efficient algorithm for solving elliptic problems on percolation clusters. Annals of Applied Probability, 32(4): 2755–2810, August 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
4/An-efficient-algorithm-for-solving-elliptic-problems-
on-percolation-clusters/10.1214/21-AAP1748.full.

Guasoni:2002:OIT

[Gua02] Paolo Guasoni. Optimal investment with transaction costs and without semimartingales. Annals of Applied Probability, 12(4): 1227–1246, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1037125861.

Guerin:2003:SLE

[Gué03] Hélène Guérin. Solving Landau equation for some soft potentials through a probabilistic approach. Annals of Applied Probability, 13(2):515–539, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1050689592.

Gupta:2012:SMC

[Gup12] Ankit Gupta. Stochastic model for cell polarity. Annals of Applied Probability, 22(2):827–859, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1333372014.

Gurvich:2014:DMS

[Gur14] Itai Gurvich. Diffusion models and steady-state approximations for exponentially ergodic Markovian queues. Annals of Applied Probability, 24(6):2527–2559, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1409058039.

Guyon:2007:LTB

[Guy07] Julien Guyon. Limit theorems for bifurcating Markov chains. application to the detection of cellular aging. Annals of Applied Probability, 17(5/6):1538–1569, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1191419175.

Garavaglia:2020:LWC

[GvdHL20] Alessandro Garavaglia, Remco van der Hofstad, and Nelly Litvak. Local weak convergence for PageRank. Annals of Applied Probability, 30(1):40–79, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1582621219.

Ghiglietti:2017:CLT

[GVR17] Andrea Ghiglietti, Anand N. Vidyashankar, and William F. Rosenberger. Central limit theorem for an adaptive randomly reinforced urn model. Annals of Applied Probability, 27(5): 2956–3003, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1509696039.

Glynn:1991:DMQ

[GW91] Peter W. Glynn and Ward Whitt. Departures from many queues in series. Annals of Applied Probability, 1(4):546–572, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005838.

Glynn:1992:AVS

[GW92] Peter W. Glynn and Ward Whitt. The asymptotic validity of sequential stopping rules for stochastic simulations. Annals of Applied Probability, 2(1):180–198, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005777.

Gates:1993:MMS

[GW93] D. J. Gates and M. Westcott. Markov models of steady crystal growth. Annals of Applied Probability, 3(2):339–355, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005427.

Glasserman:1997:CIS

[GW97] Paul Glasserman and Yashan Wang. Counterexamples in importance sampling for large deviations probabilities. Annals of Applied Probability, 7(3):731–746, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034801251.

Glazebrook:2000:IBP

[GW00] K. D. Glazebrook and D. J. Wilkinson. Index-based policies for discounted multi-armed bandits on parallel machines. Annals of Applied Probability, 10(3):877–896, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1019487512.

Gromoll:2009:FLN

[GW09] H. Christian Gromoll and Ruth J. Williams. Fluid limits for networks with bandwidth sharing and general document size distributions. Annals of Applied Probability, 19(1):243–280, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1235140339.

Georgiou:2022:DDN

[GW22] Nicholas Georgiou and Andrew R. Wade. Deposition, diffusion, and nucleation on an interval. Annals of Applied Probability, 32 (6):4849–4892, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-6/Deposition-diffusion-and-nucleation-on-an-interval/
10.1214/22-AAP1804.full.

Glasserman:2004:NPV

[GY04] Paul Glasserman and Bin Yu. Number of paths versus number of basis functions in American option pricing. Annals of Applied Probability, 14(4):2090–2119, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1099674090.

Glynn:2000:MWQ

[GZ00] Peter W. Glynn and Assaf J. Zeevi. On the maximum workload of a queue fed by fractional Brownian motion. Annals of Applied Probability, 10(4):1084–1099, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1019487607.

Gamarnik:2006:VHT

[GZ06] David Gamarnik and Assaf Zeevi. Validity of heavy traffic steady-state approximations in generalized Jackson networks. Annals of Applied Probability, 16(1):56–90, February 2006. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1141654281.

Guo:2008:IPC

[GZ08] Xin Guo and Yan Zeng. Intensity process and compensator: a new filtration expansion approach and the Jeulin–Yor theorem. Annals of Applied Probability, 18(1):120–142, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1199890018.

Goldenshluger:2009:WOA

[GZ09] Alexander Goldenshluger and Assaf Zeevi. Woodroofe’s onearmed bandit problem revisited. Annals of Applied Probability, 19(4):1603–1633, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1248700629.

Guo:2015:MSH

[GZZ15] Wenjie Guo, Jianfeng Zhang, and Jia Zhuo. A monotone scheme for high-dimensional fully nonlinear PDEs. Annals of Applied Probability, 25(3):1540–1580, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1427124136.

Haas:2005:EFI

[Haa05] Bénédicte Haas. Equilibrium for fragmentation with immigration. Annals of Applied Probability, 15(3):1958–1996, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1121433774.

Haas:2010:ABS

[Haa10] Bénédicte Haas. Asymptotic behavior of solutions of the fragmentation equation with shattering: An approach via selfsimilar Markov processes. Annals of Applied Probability, 20 (2):382–429, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1268143428.

Haggstrom:1999:PCF

[Häg99] Olle Häggström. Positive correlations in the fuzzy Potts model. Annals of Applied Probability, 9(4):1149–1159, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1029962867.

Hajek:1996:BLI

[Haj96] Bruce Hajek. Balanced loads in infinite networks. Annals of Applied Probability, 6(1):48–75, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968065.

Hansen:2006:LAM

[Han06a] Niels Richard Hansen. Local alignment of Markov chains. Annals of Applied Probability, 16(3):1262–1296, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1159804981.

Hansen:2006:MRW

[Han06b] Niels Richard Hansen. The maximum of a random walk reflected at a general barrier. Annals of Applied Probability, 16(1):15–29, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1141654279.

Harrison:1998:HTA

[Har98] J. Michael Harrison. Heavy traffic analysis of a system with parallel servers: asymptotic optimality of discrete-review policies. Annals of Applied Probability, 8(3):822–848, August 1998. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1028903452.

Harrison:2000:BMO

[Har00] J. Michael Harrison. Brownian models of open processing networks: canonical representation of workload. Annals of Applied Probability, 10(1):75–103, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1019737665. See correc-

tions [Har03b, Har06].

Harrison:2003:BVB

[Har03a] J. Michael Harrison. A broader view of Brownian networks. Annals of Applied Probability, 13(3):1119–1150, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1060202837.

Harrison:2003:CBM

[Har03b] Michael J. Harrison. Correction: “Brownian models of open processing networks: canonical representation of workload”. Annals of Applied Probability, 13(1):390–393, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1042765673. See [Har00].

Harrison:2006:CBM

[Har06] J. Michael Harrison. Correction: “Brownian models of open processing networks: canonical representation of workload”. Annals of Applied Probability, 16(3):1703–1732, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1159804997. See [Har00].

Harper:2013:BSG

[Har13] Adam J. Harper. Bounds on the suprema of Gaussian processes, and omega results for the sum of a random multiplicative function. Annals of Applied Probability, 23(2):584–616, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1360682023.

Huo:2019:ZVM

[HD19] Ran Huo and Rick Durrett. The zealot voter model. Annals of Applied Probability, 29(5):3128–3154, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1571385631.

Hwang:2016:SPT

[HDI16] Sung Jin Hwang, Steven B. Damelin, and Alfred O. Hero III. Shortest path through random points. Annals of Applied Probability, 26(5):2791–2823, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1476884304.

Hu:2022:AOB

[HDP22] Yue Hu, Jing Dong, and Ohad Perry. Asymptotic optimality of the binomial-exhaustive policy for polling systems with large switchover times. Annals of Applied Probability, 32(6):4803– 4848, December 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
6/Asymptotic-optimality-of-the-binomial-exhaustive-policy-
for-polling-systems/10.1214/22-AAP1803.full.

He:2020:BEF

[He20] Yukun He. Bulk eigenvalue fluctuations of sparse random matrices. Annals of Applied Probability, 30(6):2846–2879, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-6/Bulk-
eigenvalue-fluctuations-of-sparse-random-matrices/10.
1214/20-AAP1575.full.

Heinrich:2005:LDE

[Hei05] Lothar Heinrich. Large deviations of the empirical volume fraction for stationary Poisson grain models. Annals of Applied Probability, 15(1A):392–420, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1106922332.

Henard:2015:FLC

[Hén15] Olivier Hénard. The fixation line in the Λ-coalescent. Annals of Applied Probability, 25(5):3007–3032, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1438261058.

Heydecker:2019:PCH

[Hey19] Daniel Heydecker. Pathwise convergence of the hard spheres Kac process. Annals of Applied Probability, 29(5):3062–3127, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1571385630.

Heydecker:2023:LDK

[Hey23] Daniel Heydecker. Large deviations of Kac’s conservative particle system and energy nonconserving solutions to the Boltzmann equation: a counterexample to the predicted rate function. Annals of Applied Probability, 33(3):1758–1826, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/Large-
deviations-of-Kacs-conservative-particle-system-and-
energy-nonconserving/10.1214/22-AAP1852.full.

Henderson:2008:ESO

[HH08] Vicky Henderson and David Hobson. An explicit solution for an optimal stopping/optimal control problem which models an asset sale. Annals of Applied Probability, 18(5):1681–1705, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1225372946.

Hoze:2018:SCF

[HH18] Nathanael Hoze and David Holcman. Stochastic coagulationfragmentation processes with a finite number of particles and applications. Annals of Applied Probability, 28(3):1449–1490, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1527840024.

Hammond:2019:CPI

[HH19] Alan Hammond and Milind Hegde. Critical point for infinite cycles in a random loop model on trees. Annals of Applied Probability, 29(4):2067–2088, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1563869037.

Holden:2009:AIM

[HHH09] Lars Holden, Ragnar Hauge, and Marit Holden. Adaptive independent Metropolis–Hastings. Annals of Applied Probability, 19(1):395–413, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1235140343.

Heydenreich:2017:SSS

[HHJ17] Markus Heydenreich, Tim Hulshof, and Joost Jorritsma. Structures in supercritical scale-free percolation. Annals of Applied Probability, 27(4):2569–2604, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1504080041.

Harris:2020:SMN

[HHK20] Simon C. Harris, Emma Horton, and Andreas E. Kyprianou. Stochastic methods for the neutron transport equation II: Almost sure growth. Annals of Applied Probability, 30(6):2815– 2845, December 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
6/Stochastic-methods-for-the-neutron-transport-equation-
II--Almost/10.1214/20-AAP1574.full.

Hirsch:2021:AWP

[HHK21] Christian Hirsch, Mark Holmes, and Victor Kleptsyn. Absence of WARM percolation in the very strong reinforcement regime. Annals of Applied Probability, 31(1):199–217, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/Absence-
of-WARM-percolation-in-the-very-strong-reinforcement-
regime/10.1214/20-AAP1587.full.

Hefter:2019:LEB

[HHMG19] Mario Hefter, André Herzwurm, and Thomas Müller-Gronbach. Lower error bounds for strong approximation of scalar SDEs with non-Lipschitzian coefficients. Annals of Applied Probability, 29(1):178–216, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1544000428.

Hwang:1993:AGD

[HHMS93] Chii-Ruey Hwang, Shu-Yin Hwang-Ma, and Shuenn-Jyi Sheu. Accelerating Gaussian diffusions. Annals of Applied Probability, 3(3):897–913, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005371.

Hwang:2005:AD

[HHMS05] Chii-Ruey Hwang, Shu-Yin Hwang-Ma, and Shuenn-Jyi Sheu. Accelerating diffusions. Annals of Applied Probability, 15(2): 1433–1444, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1115137980.

Hsing:1996:ETA

[HHR96] Tailen Hsing, Jürg Hüsler, and Rolf-Dieter Reiss. The extremes of a triangular array of normal random variables. Annals of Applied Probability, 6(2):671–686, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1034968149.

Hernandez-Hernandez:2015:ZSG

[HHSZ15] Daniel Hernandez-Hernandez, Robert S. Simon, and Mihail Zervos. A zero-sum game between a singular stochastic controller and a discretionary stopper. Annals of Applied Probability, 25(1):46–80, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1418740178.

Herrmann:2005:EPD

[HI05] Samuel Herrmann and Peter Imkeller. The exit problem for diffusions with time-periodic drift and stochastic resonance. Annals of Applied Probability, 15(1A):39–68, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1106922320.

Hofbauer:2009:TAR

[HI09] Josef Hofbauer and Lorens A. Imhof. Time averages, recurrence and transience in the stochastic replicator dynamics. Annals of Applied Probability, 19(4):1347–1368, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1248700620.

Hu:2005:UMI

[HIM05] Ying Hu, Peter Imkeller, and Matthias Müller. Utility maximization in incomplete markets. Annals of Applied Probability, 15(3):1691–1712, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433765.

Herrmann:2006:TTS

[HIP06] Samuel Herrmann, Peter Imkeller, and Dierk Peithmann. Transition times and stochastic resonance for multidimensional diffusions with time periodic drift: A large deviations approach. Annals of Applied Probability, 16(4):1851–1892, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1169065210.

Herrmann:2008:LDK

[HIP08] Samuel Herrmann, Peter Imkeller, and Dierk Peithmann. Large deviations and a Kramers’ type law for self-stabilizing diffusions. Annals of Applied Probability, 18(4):1379–1423, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1216677126.

Hosseini:2023:SGE

[HJ23] Bamdad Hosseini and James E. Johndrow. Spectral gaps and error estimates for infinite-dimensional Metropolis–Hastings with non-Gaussian priors. Annals of Applied Probability, 33(3):1827– 1873, June 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
3/Spectral-gaps-and-error-estimates-for-infinite-dimensional-
MetropolisHastings-with/10.1214/22-AAP1854.full.

Hoffman:2016:TRP

[HJJ16] Christopher Hoffman, Tobias Johnson, and Matthew Junge. From transience to recurrence with Poisson tree frogs. Annals of Applied Probability, 26(3):1620–1635, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905013.

Hutzenthaler:2012:SCE

[HJK12] Martin Hutzenthaler, Arnulf Jentzen, and Peter E. Kloeden. Strong convergence of an explicit numerical method for SDEs with nonglobally Lipschitz continuous coefficients. Annals of Applied Probability, 22(4):1611–1641, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1344614205.

Hutzenthaler:2013:DMM

[HJK13] Martin Hutzenthaler, Arnulf Jentzen, and Peter E. Kloeden. Divergence of the multilevel Monte Carlo Euler method for nonlinear stochastic differential equations. Annals of Applied Probability, 23(5):1913–1966, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1377696302.

Harris:2020:CSC

[HJR20] Simon C. Harris, Samuel G. G. Johnston, and Matthew I. Roberts. The coalescent structure of continuous-time Galton– Watson trees. Annals of Applied Probability, 30(3):1368–1414, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-3/The-
coalescent-structure-of-continuous-time-GaltonWatson-
trees/10.1214/19-AAP1532.full.

Hill:1992:SIO

[HK92] T. P. Hill and D. P. Kennedy. Sharp inequalities for optimal stopping with rewards based on ranks. Annals of Applied Probability, 2(2):503–517, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005713.

Hugonnier:2004:OIR

[HK04] Julien Hugonnier and Dmitry Kramkov. Optimal investment with random endowments in incomplete markets. Annals of Applied Probability, 14(2):845–864, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1082737114.

Hobson:2013:MFM

[HK13] David Hobson and Martin Klimmek. Maximizing functionals of the maximum in the Skorokhod embedding problem and an application to variance swaps. Annals of Applied Probability, 23 (5):2020–2052, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1377696305.

Hackmann:2016:ALP

[HK16] Daniel Hackmann and Alexey Kuznetsov. Approximating Lévy processes with completely monotone jumps. Annals of Applied Probability, 26(1):328–359, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1452003241.

He:2017:MES

[HK17] Yukun He and Antti Knowles. Mesoscopic eigenvalue statistics of Wigner matrices. Annals of Applied Probability, 27(3):1510– 1550, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1500451234.

Henning:2021:CLG

[HK21] Florian Henning and Christof Külske. Coexistence of localized Gibbs measures and delocalized gradient Gibbs measures on trees. Annals of Applied Probability, 31(5):2284–2310, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-5/Coexistence-
of-localized-Gibbs-measures-and-delocalized-gradient-
Gibbs-measures/10.1214/20-AAP1647.full.

Han:2022:BEB

[HK22] Qiyang Han and Kengo Kato. Berry–Esseen bounds for Chernoff-type nonstandard asymptotics in isotonic regression. Annals of Applied Probability, 32(2):1459–1498, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/BerryEsseen-
bounds-for-Chernoff-type-nonstandard-asymptotics-in-
isotonic-regression/10.1214/21-AAP1716.full.

Henning:2023:EGG

[HK23] Florian Henning and Christof Külske. Existence of gradient Gibbs measures on regular trees which are not translation invariant. Annals of Applied Probability, 33(4):3010–3038, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-4/Existence-
of-gradient-Gibbs-measures-on-regular-trees-which-are/
10.1214/22-AAP1883.full.

Hubalek:2006:VOH

[HKK06] Friedrich Hubalek, Jan Kallsen, and Leszek Krawczyk. Varianceoptimal hedging for processes with stationary independent increments. Annals of Applied Probability, 16(2):853–885, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1151592253.

Hsu:2019:MTE

[HKL^{+}19] Daniel Hsu, Aryeh Kontorovich, David A. Levin, Yuval Peres, Csaba Szepesvári, and Geoffrey Wolfer. Mixing time estimation in reversible Markov chains from a single sample path. Annals of Applied Probability, 29(4):2439–2480, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1563869047.

Horton:2020:SMN

[HKV20] Emma Horton, Andreas E. Kyprianou, and Denis Villemonais. Stochastic methods for the neutron transport equation i: Linear semigroup asymptotics. Annals of Applied Probability, 30 (6):2573–2612, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-6/Stochastic-methods-for-the-neutron-transport-
equation-I--Linear/10.1214/20-AAP1567.full.

Hunt:1997:OTR

[HL97] P. J. Hunt and C. N. Laws. Optimization via trunk reservation in single resource loss systems under heavy traffic. Annals of Applied Probability, 7(4):1058–1079, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1043862424.

Huebner:2000:AAK

[HL00] M. Huebner and S. Lototsky. Asymptotic analysis of a kernel estimator for parabolic SPDE’s with time-dependent coefficients. Annals of Applied Probability, 10(4):1246–1258, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1019487615.

Heunis:2001:USS

[HL01] Andrew J. Heunis and Vladimir M. Lucic. On uniqueness of solutions for the stochastic differential equations of nonlinear filtering. Annals of Applied Probability, 11(1):182–209, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
998926990.

Hambly:2017:SMV

[HL17] Ben Hambly and Sean Ledger. A stochastic McKean–Vlasov equation for absorbing diffusions on the half-line. Annals of Applied Probability, 27(5):2698–2752, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1509696033.

Holden:2020:LBT

[HL20] Nina Holden and Russell Lyons. Lower bounds for trace reconstruction. Annals of Applied Probability, 30(2):503–525, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/

1591603214. See erratum [HL22].

HellouindeMenibus:2021:ABO

[HL21] Benjamin Hellouin de Menibus and Yvan Le Borgne. Asymptotic behaviour of the one-dimensional “rock–paper–scissors” cyclic cellular automaton. Annals of Applied Probability, 31(5):2420– 2440, October 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
5/Asymptotic-behaviour-of-the-one-dimensional-rockpaperscissors-
cyclic-cellular-automaton/10.1214/20-AAP1651.full.

Holden:2022:ESL

[HL22] Nina Holden and Russell Lyons. Erratum to Lower bounds for trace reconstruction. Annals of Applied Probability, 32(4):3201– 3203, August 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
4/Erratum-to-Lower-bounds-for-trace-reconstruction/10.
1214/22-AAP1827.full. See [HL20].

Hult:2005:FLD

[HLMS05] Henrik Hult, Filip Lindskog, Thomas Mikosch, and Gennady Samorodnitsky. Functional large deviations for multivariate regularly varying random walks. Annals of Applied Probability, 15 (4):2651–2680, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965775.

Hachem:2007:DEC

[HLN07] Walid Hachem, Philippe Loubaton, and Jamal Najim. Deterministic equivalents for certain functionals of large random matrices. Annals of Applied Probability, 17(3):875–930, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1179839177.

Hachem:2008:CIT

[HLN08] Walid Hachem, Philippe Loubaton, and Jamal Najim. A CLT for information-theoretic statistics of Gram random matrices with a given variance profile. Annals of Applied Probability, 18(6): 2071–2130, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1227708913.

Hu:2016:RCA

[HLN16] Yaozhong Hu, Yanghui Liu, and David Nualart. Rate of convergence and asymptotic error distribution of Euler approximation schemes for fractional diffusions. Annals of Applied Probability, 26(2):1147–1207, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1458651830.

Hofmanova:2021:RPN

[HLN21a] Martina Hofmanová, James-Michael Leahy, and Torstein Nilssen. On a rough perturbation of the Navier–Stokes system and its vorticity formulation. Annals of Applied Probability, 31 (2):736–777, April 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
2/On-a-rough-perturbation-of-the-NavierStokes-system-
and-its/10.1214/20-AAP1603.full.

Hu:2021:CNS

[HLN21b] Yaozhong Hu, Yanghui Liu, and David Nualart. Crank–Nicolson scheme for stochastic differential equations driven by fractional Brownian motions. Annals of Applied Probability, 31(1):39–83, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/CrankNicolson-
scheme-for-stochastic-differential-equations-driven-
by-fractional-Brownian/10.1214/20-AAP1582.full.

Henry-Labordere:2016:MMM

[HLOST16] Pierre Henry-Labordère, Jan Obłój, Peter Spoida, and Nizar Touzi. The maximum maximum of a martingale with given n marginals. Annals of Applied Probability, 26(1):1–44, 02 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Hill:2022:STE

[HLR22] Max Hill, Brandon Legried, and Sebastien Roch. Species tree estimation under joint modeling of coalescence and duplication: Sample complexity of quartet methods. Annals of Applied Probability, 32(6):4681–4705, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-6/Species-tree-estimation-under-joint-
modeling-of-coalescence-and-duplication/10.1214/22-AAP1799.
full.

Hug:2016:SOP

[HLS16] Daniel Hug, Günter Last, and Matthias Schulte. Second-order properties and central limit theorems for geometric functionals of Boolean models. Annals of Applied Probability, 26(1):73–135, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1452003235.

Hambly:2019:MVE

[HLS19a] Ben Hambly, Sean Ledger, and Andreas Søjmark. A McKean– Vlasov equation with positive feedback and blow-ups. Annals of Applied Probability, 29(4):2338–2373, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1563869045.

Hwang:2019:LLT

[HLS19b] Jong Yun Hwang, Ji Oon Lee, and Kevin Schnelli. Local law and Tracy–Widom limit for sparse sample covariance matrices. Annals of Applied Probability, 29(5):3006–3036, October 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1571385628.

Henry-Labordere:2021:BDR

[HLT21] Pierre Henry-Labordère and Nizar Touzi. Branching diffusion representation for nonlinear Cauchy problems and Monte Carlo approximation. Annals of Applied Probability, 31(5):2350–2375, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-5/Branching-
diffusion-representation-for-nonlinear-Cauchy-problems-
and-Monte-Carlo/10.1214/20-AAP1649.full.

Henry-Labordere:2017:USS

[HLTT17] Pierre Henry-Labordère, Xiaolu Tan, and Nizar Touzi. Unbiased simulation of stochastic differential equations. Annals of Applied Probability, 27(6):3305–3341, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1513328702.

Han:2009:AIC

[HM09] Guangyue Han and Brian Marcus. Asymptotics of inputconstrained binary symmetric channel capacity. Annals of Applied Probability, 19(3):1063–1091, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1245071019.

Hegarty:2014:EAP

[HM14] Peter Hegarty and Anders Martinsson. On the existence of accessible paths in various models of fitness landscapes. Annals of Applied Probability, 24(4):1375–1395, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1400073652.

Hofmann:2000:SSC

[HMGR00] Norbert Hofmann, Thomas Müller-Gronbach, and Klaus Ritter. Step size control for the uniform approximation of systems of stochastic differential equations with additive noise. Annals of Applied Probability, 10(2):616–633, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487358.

Hermon:2020:SNM

[HMQS20] Jonathan Hermon, Ben Morris, Chuan Qin, and Allan Sly. The social network model on infinite graphs. Annals of Applied Probability, 30(2):902–935, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1591603226.

Huisinga:2004:PTM

[HMS04] Wilhelm Huisinga, Sean Meyn, and Christof Schütte. Phase transitions and metastability in Markovian and molecular systems. Annals of Applied Probability, 14(1):419–458, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1075828057.

Hartarsky:2022:CBS

[HMT22] Ivailo Hartarsky, Fabio Martinelli, and Cristina Toninelli. Coalescing and branching simple symmetric exclusion process. Annals of Applied Probability, 32(4):2841–2859, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Coalescing-
and-branching-simple-symmetric-exclusion-process/10.
1214/21-AAP1750.full.

Heiny:2021:PPC

[HMY21] Johannes Heiny, Thomas Mikosch, and Jorge Yslas. Point process convergence for the off-diagonal entries of sample covariance matrices. Annals of Applied Probability, 31(2):538–560, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/Point-
process-convergence-for-the-off-diagonal-entries-of-
sample/10.1214/20-AAP1597.full.

Hening:2018:CES

[HN18] Alexandru Hening and Dang H. Nguyen. Coexistence and extinction for stochastic Kolmogorov systems. Annals of Applied Probability, 28(3):1893–1942, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1527840035.

Hobson:2019:LCM

[HN19] David G. Hobson and Dominykas Norgilas. The left-curtain martingale coupling in the presence of atoms. Annals of Applied Probability, 29(3):1904–1928, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1550566846.

Hager:2022:MC

[HN22] Paul Hager and Eyal Neuman. The multiplicative chaos of H = 0 fractional Brownian fields. Annals of Applied Probability, 32(3): 2139–2179, June 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
3/The-multiplicative-chaos-of-H0-fractional-Brownian-
fields/10.1214/21-AAP1730.full.

Hughes:2013:RPM

[HNNZ13] Christopher Hughes, Joseph Najnudel, Ashkan Nikeghbali, and Dirk Zeindler. Random permutation matrices under the generalized Ewens measure. Annals of Applied Probability, 23(3):987– 1024, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1362684852.

Hata:2010:APM

[HNS10] Hiroaki Hata, Hideo Nagai, and Shuenn-Jyi Sheu. Asymptotics of the probability minimizing a “down-side” risk. Annals of Applied Probability, 20(1):52–89, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1262962318.

Hu:2011:MCB

[HNS11] Yaozhong Hu, David Nualart, and Xiaoming Song. Malliavin calculus for backward stochastic differential equations and application to numerical solutions. Annals of Applied Probability, 21(6):2379–2423, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1322057325.

Hening:2022:CDT

[HNS22] Alexandru Hening, Dang H. Nguyen, and Sebastian J. Schreiber. A classification of the dynamics of three-dimensional stochastic ecological systems. Annals of Applied Probability, 32(2):893–931, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/A-classification-
of-the-dynamics-of-three-dimensional-stochastic-ecological/
10.1214/21-AAP1699.full.

Hobson:1998:VMO

[Hob98] David G. Hobson. Volatility misspecification, option pricing and superreplication via coupling. Annals of Applied Probability, 8 (1):193–205, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1027961040.

Hobson:2016:MM

[Hob16] David Hobson. Mimicking martingales. Annals of Applied Probability, 26(4):2273–2303, August 2016. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1472745459.

Hoffman:2005:CRT

[Hof05] Christopher Hoffman. Coexistence for Richardson type competing spatial growth models. Annals of Applied Probability, 15 (1B):739–747, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1107271666.

Hoffman:2008:GFP

[Hof08] Christopher Hoffman. Geodesics in first passage percolation. Annals of Applied Probability, 18(5):1944–1969, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1225372957.

Hogan:1993:PCT

[Hog93] Michael Hogan. Problems in certain two-factor term structure models. Annals of Applied Probability, 3(2):576–581, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005438.

Holroyd:1998:EUI

[Hol98] Alexander E. Holroyd. Existence and uniqueness of infinite components in generic rigidity percolation. Annals of Applied Probability, 8(3):944–973, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903458.

Holroyd:2001:RPB

[Hol01] Alexander E. Holroyd. Rigidity percolation and boundary conditions. Annals of Applied Probability, 11(4):1063–1078, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345395.

Huang:1992:OCP

[HP92] Chi-Fu Huang and Henri Pages. Optimal consumption and portfolio policies with an infinite horizon: Existence and convergence. Annals of Applied Probability, 2(1):36–64, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005770.

Herczegh:2015:SPP

[HP15] Attila Herczegh and Vilmos Prokaj. Shadow price in the power utility case. Annals of Applied Probability, 25(5):2671–2707, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1438261051.

Hutzenthaler:2020:PCM

[HP20] Martin Hutzenthaler and Daniel Pieper. Propagation of chaos and the many-demes limit for weakly interacting diffusions in the sparse regime. Annals of Applied Probability, 30(5):2311–2354, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-5/Propagation-
of-chaos-and-the-many-demes-limit-for-weakly/10.1214/
20-AAP1559.full.

Hernandez:2023:MMG

[HP23] Camilo Hernández and Dylan Possamaï. Me, myself and I: a general theory of non-Markovian time-inconsistent stochastic control for sophisticated agents. Annals of Applied Probability, 33(2):1196–1258, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-2/Me-myself-and-I--A-general-theory-of-non/10.
1214/22-AAP1845.full.

Helmuth:2022:CDH

[HPP22] Tyler Helmuth, Will Perkins, and Samantha Petti. Correlation decay for hard spheres via Markov chains. Annals of Applied Probability, 32(3):2063–2082, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-3/Correlation-decay-for-hard-spheres-
via-Markov-chains/10.1214/21-AAP1728.full.

Husler:2003:CUN

[HPS03] J. Hüsler, V. Piterbarg, and O. Seleznjev. On convergence of the uniform norms for Gaussian processes and linear approximation problems. Annals of Applied Probability, 13(4):1615–1653, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1069786514.

Huzak:2004:RPD

[HPŠV04] Miljenko Huzak, Mihael Perman, Hrvoje Šikić, and Zoran Vondraček. Ruin probabilities and decompositions for general perturbed risk processes. Annals of Applied Probability, 14(3):1378– 1397, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1089736289.

Henderson:1995:IDT

[HPTvD95] W. Henderson, C. E. M. Pearce, P. G. Taylor, and N. M. van Dijk. Insensitivity in discrete-time generalized semi-Markov processes allowing multiple events and probabilistic service scheduling. Annals of Applied Probability, 5(1):78–96, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177004829.

Holst:1996:GSM

[HQR96] L. Holst, M. P. Quine, and J. Robinson. A general stochastic model for nucleation and linear growth. Annals of Applied Probability, 6(3):903–921, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968233.

Hall:1994:RBF

[HR94] Peter Hall and Rahul Roy. On the relationship between fractal dimension and fractal index for stationary stochastic processes. Annals of Applied Probability, 4(1):241–253, February 1994. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005210.

Hall:1997:ALT

[HR97] Peter Hall and Marc Raimondo. Approximating a line thrown at random onto a grid. Annals of Applied Probability, 7(3):648–665, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034801247.

Hobert:2004:MRA

[HR04] James P. Hobert and Christian P. Robert. A mixture representation of π with applications in Markov chain Monte Carlo and perfect sampling. Annals of Applied Probability, 14(3):1295–1305, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1089736286.

Heffernan:2007:LLR

[HR07] Janet E. Heffernan and Sidney I. Resnick. Limit laws for random vectors with an extreme component. Annals of Applied Probability, 17(2):537–571, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1174323256.

Heveling:2009:PVA

[HR09] Matthias Heveling and Matthias Reitzner. Poisson–Voronoi approximation. Annals of Applied Probability, 19(2):719–736, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1241702248.

Heath:1997:PBO

[HRS97] David Heath, Sidney Resnick, and Gennady Samorodnitsky. Patterns of buffer overflow in a class of queues with long memory in the input stream. Annals of Applied Probability, 7(4): 1021–1057, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1043862423.

Heath:1999:HSP

[HRS99] David Heath, Sidney Resnick, and Gennady Samorodnitsky. How system performance is affected by the interplay of averages in a fluid queue with long range dependence induced by heavy tails. Annals of Applied Probability, 9(2):352–375, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1029962746.

Hart:2008:ESS

[HRW08] Sergiu Hart, Yosef Rinott, and Benjamin Weiss. Evolutionarily stable strategies of random games, and the vertices of random polygons. Annals of Applied Probability, 18(1):259–287, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1199890023.

Heinricher:1993:OCRb

[HS93a] Arthur C. Heinricher and Richard H. Stockbridge. Optimal control and replacement with state-dependent failure rate: An invariant measure approach. Annals of Applied Probability, 3(2): 380–402, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005430.

Heinricher:1993:OCRa

[HS93b] Arthur C. Heinricher and Richard H. Stockbridge. Optimal control and replacement with state-dependent failure rate: Dynamic programming. Annals of Applied Probability, 3(2):364– 379, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005429.

Horvath:1999:LTQ

[HS99] Lajos Horváth and Qi-Man Shao. Limit theorems for quadratic forms with applications to Whittle’s estimate. Annals of Applied Probability, 9(1):146–187, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962600.

Hayes:2007:GLB

[HS07] Thomas P. Hayes and Alistair Sinclair. A general lower bound for mixing of single-site dynamics on graphs. Annals of Applied Probability, 17(3):931–952, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1179839178.

Hermon:2019:VAS

[HS19a] Jonathan Hermon and Justin Salez. A version of Aldous’ spectral-gap conjecture for the zero range process. Annals of Applied Probability, 29(4):2217–2229, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1563869041.

Herrmann:2019:RPH

[HS19b] Sebastian Herrmann and Florian Stebegg. Robust pricing and hedging around the globe. Annals of Applied Probability, 29 (6):3348–3386, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1578366317.

Hermon:2021:EDE

[HS21a] Jonathan Hermon and Justin Salez. Entropy dissipation estimates for inhomogeneous zero-range processes. Annals of Applied Probability, 31(5):2275–2283, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-5/Entropy-dissipation-estimates-for-
inhomogeneous-zero-range-processes/10.1214/20-AAP1646.
full.

Hermon:2021:IPH

[HS21b] Jonathan Hermon and Justin Salez. The interchange process on high-dimensional products. Annals of Applied Probability, 31(1): 84–98, February 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
1/The-interchange-process-on-high-dimensional-products/
10.1214/20-AAP1583.full.

Hermon:2023:MLS

[HS23] Jonathan Hermon and Justin Salez. Modified log-Sobolev inequalities for strong-Rayleigh measures. Annals of Applied Probability, 33(2):1301–1314, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-2/Modified-log-Sobolev-inequalities-
for-strong-Rayleigh-measures/10.1214/22-AAP1847.full.

Hotz:2013:SCL

[HSH^{+}13] Thomas Hotz, Sean Skwerer, Stephan Huckemann, Huiling Le, J. S. Marron, Jonathan C. Mattingly, Ezra Miller, James Nolen, Megan Owen, and Vic Patrangenaru. Sticky central limit theorems on open books. Annals of Applied Probability, 23(6): 2238–2258, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1382447687.

Hsing:1994:ADA

[Hsi94] Tailen Hsing. On the asymptotic distribution of the area outside a random convex Hull in a disk. Annals of Applied Probability, 4(2):478–493, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005069.

Heinrich:2006:CLT

[HSS06] Lothar Heinrich, Hendrik Schmidt, and Volker Schmidt. Central limit theorems for Poisson hyperplane tessellations. Annals of Applied Probability, 16(2):919–950, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1151592255.

Hilario:2023:PTP

[HSST23] Marcelo R. Hilário, Marcos Sá, Remy Sanchis, and Augusto Teixeira. Phase transition for percolation on a randomly stretched square lattice. Annals of Applied Probability, 33(4): 3145–3168, August 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
4/Phase-transition-for-percolation-on-a-randomly-stretched-
square-lattice/10.1214/22-AAP1887.full.

Hardin:1991:NRS

[HST91] Clyde D. Hardin, Jr., Gennady Samorodnitsky, and Murad S. Taqqu. Nonlinear regression of stable random variables. Annals of Applied Probability, 1(4):582–612, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005840.

Hiraoka:2018:LTP

[HST18] Yasuaki Hiraoka, Tomoyuki Shirai, and Khanh Duy Trinh. Limit theorems for persistence diagrams. Annals of Applied Probability, 28(5):2740–2780, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1535443233.

Hairer:2007:ASA

[HSV07] M. Hairer, A. M. Stuart, and J. Voss. Analysis of SPDEs arising in path sampling part II: The nonlinear case. Annals of Applied Probability, 17(5/6):1657–1706, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1191419180.

Hairer:2011:SCH

[HSV11] Martin Hairer, Andrew M. Stuart, and Jochen Voss. Sampling conditioned hypoelliptic diffusions. Annals of Applied Probability, 21(2):669–698, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1300800985.

Hairer:2014:SGM

[HSV14] Martin Hairer, Andrew M. Stuart, and Sebastian J. Vollmer. Spectral gaps for a Metropolis–Hastings algorithm in infinite dimensions. Annals of Applied Probability, 24(6):2455–2490, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1409058037.

Hu:2022:CSL

[HSX22] Ying Hu, Xiaomin Shi, and Zuo Quan Xu. Constrained stochastic LQ control with regime switching and application to portfolio selection. Annals of Applied Probability, 32(1):426–460, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Constrained-
stochastic-LQ-control-with-regime-switching-and-application-
to/10.1214/21-AAP1684.full.

Helmes:2017:CIM

[HSZ17] Kurt L. Helmes, Richard H. Stockbridge, and Chao Zhu. Continuous inventory models of diffusion type: Long-term average cost criterion. Annals of Applied Probability, 27(3):1831–1885, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1500451245.

Houdre:2012:RAF

[HT12] Christian Houdré and Zsolt Talata. On the rate of approximation in finite-alphabet longest increasing subsequence problems. Annals of Applied Probability, 22(6):2539–2559, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1353695961.

Huber:2004:PSU

[Hub04] Mark Huber. Perfect sampling using bounding chains. Annals of Applied Probability, 14(2):734–753, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1082737109.

Huber:2015:AAN

[Hub15] Mark Huber. Approximation algorithms for the normalizing constant of Gibbs distributions. Annals of Applied Probability, 25(2):974–985, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355135.

Harrison:1997:DCB

[HV97] J. Michael Harrison and Jan A. Van Mieghem. Dynamic control of Brownian networks: state space collapse and equivalent workload formulations. Annals of Applied Probability, 7(3):747–771, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034801252.

Hayes:2006:CSD

[HV06] Thomas P. Hayes and Eric Vigoda. Coupling with the stationary distribution and improved sampling for colorings and independent sets. Annals of Applied Probability, 16(3):1297–1318, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804982.

Harrison:1992:BMF

[HW92a] J. M. Harrison and R. J. Williams. Brownian models of feedforward queueing networks: Quasireversibility and product form solutions. Annals of Applied Probability, 2(2):263–293, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005704.

Holley:1992:MDS

[HW92b] Richard Holley and Edward C. Waymire. Multifractal dimensions and scaling exponents for strongly bounded random cascades. Annals of Applied Probability, 2(4):819–845, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005577.

Hu:1994:ERP

[HW94] Yiming Hu and W. A. Woyczynski. An extremal rearrangement property of statistical solutions of Burgers’ equation. Annals of Applied Probability, 4(3):838–858, August 1994. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004974.

Harrison:1996:MCQ

[HW96] J. M. Harrison and R. J. Williams. A multiclass closed queueing network with unconventional heavy traffic behavior. Annals of Applied Probability, 6(1):1–47, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968064.

Harrison:2005:WRG

[HW05] J. M. Harrison and R. J. Williams. Workload reduction of a generalized Brownian network. Annals of Applied Probability, 15 (4):2255–2295, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965763.

Hutzenthaler:2007:EBL

[HW07] M. Hutzenthaler and A. Wakolbinger. Ergodic behavior of locally regulated branching populations. Annals of Applied Probability, 17(2):474–501, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1174323254.

Hitczenko:2009:PTT

[HW09] Paweł Hitczenko and Jacek Wesołowski. Perpetuities with thin tails revisited. Annals of Applied Probability, 19(6):2080–2101, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1259158767.

Hitczenko:2010:PTT

[HW10] Paweł Hitczenko and Jacek Wesołowski. Perpetuities with thin tails revisited. Annals of Applied Probability, 20(3):1177, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1276867310.

Hwang:1996:LDC

[Hwa96] Hsien-Kuei Hwang. Large deviations for combinatorial distributions. I. Central limit theorems. Annals of Applied Probability, 6(1):297–319, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968075.

Hwang:1998:LDC

[Hwa98] Hsien-Kuei Hwang. Large deviations of combinatorial distributions. II. Local limit theorems. Annals of Applied Probability, 8(1):163–181, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1027961038.

Horst:2022:MSV

[HX22] Ulrich Horst and Wei Xu. The microstructure of stochastic volatility models with self-exciting jump dynamics. Annals of Applied Probability, 32(6):4568–4610, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/journals/annals-of-applied-
probability/volume-32/issue-6/The-microstructure-of-
stochastic-volatility-models-with-self-exciting-jump/
10.1214/22-AAP1796.full.

Horst:2022:PLU

[HXZ22] Ulrich Horst, Xiaonyu Xia, and Chao Zhou. Portfolio liquidation under factor uncertainty. Annals of Applied Probability, 32(1): 80–123, February 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
1/Portfolio-liquidation-under-factor-uncertainty/10.
1214/21-AAP1672.full.

Han:2020:ATS

[HYTC20] Ruijian Han, Rougang Ye, Chunxi Tan, and Kani Chen. Asymptotic theory of sparse Bradley–Terry model. Annals of Applied Probability, 30(5):2491–2515, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-5/Asymptotic-theory-of-sparse-BradleyTerry-
model/10.1214/20-AAP1564.full.

Isohatala:2023:DAM

[IH23] Jukka Isohätälä and William B. Haskell. A dynamic analytic method for risk-aware controlled martingale problems. Annals of Applied Probability, 33(3):1661–1700, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/A-dynamic-
analytic-method-for-risk-aware-controlled-martingale-
problems/10.1214/22-AAP1794.full.

Ichiba:2010:CBP

[IK10] Tomoyuki Ichiba and Ioannis Karatzas. On collisions of Brownian particles. Annals of Applied Probability, 20(3):951–977, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1276867303.

Ibrahimi:2015:SSR

[IKKM15] Morteza Ibrahimi, Yash Kanoria, Matt Kraning, and Andrea Montanari. The set of solutions of random XORSAT formulae. Annals of Applied Probability, 25(5):2743–2808, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1438261053.

Iyer:2015:LTS

[ILP15] Gautam Iyer, Nicholas Leger, and Robert L. Pego. Limit theorems for Smoluchowski dynamics associated with critical continuous-state branching processes. Annals of Applied Probability, 25(2):675–713, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355128.

Ivanoff:2002:RCS

[IM02] B. Gail Ivanoff and Ely Merzbach. Random censoring in setindexed survival analysis. Annals of Applied Probability, 12 (3):944–971, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1031863176.

Ivanoff:2010:ODC

[IM10] B. Gail Ivanoff and Ely Merzbach. Optimal detection of a change-set in a spatial Poisson process. Annals of Applied Probability, 20(2):640–659, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1268143435.

Imhof:2005:LRB

[Imh05] Lorens A. Imhof. The long-run behavior of the stochastic replicator dynamics. Annals of Applied Probability, 15(1B):1019–1045, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1107271677.

Iscoe:1993:CAS

[IMQ93] I. Iscoe, D. McDonald, and K. Qian. Capacity of ATM switches. Annals of Applied Probability, 3(2):277–295, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005425.

Ingrassia:1994:RCM

[Ing94] Salvatore Ingrassia. On the rate of convergence of the Metropolis algorithm and Gibbs sampler by geometric bounds. Annals of Applied Probability, 4(2):347–389, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005064.

Inoue:2002:ABP

[Ino02] Akihiko Inoue. Asymptotic behavior for partial autocorrelation functions of fractional ARIMA processes. Annals of Applied Probability, 12(4):1471–1491, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1037125870.

Imkeller:2013:DRN

[INPY13] Peter Imkeller, N. Sri Namachchivaya, Nicolas Perkowski, and Hoong C. Yeong. Dimensional reduction in nonlinear filtering: A homogenization approach. Annals of Applied Probability, 23 (6):2290–2326, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1382447689.

Ichiba:2011:HAM

[IPB^{+}11] Tomoyuki Ichiba, Vassilios Papathanakos, Adrian Banner, Ioannis Karatzas, and Robert Fernholz. Hybrid atlas models. Annals of Applied Probability, 21(2):609–644, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1300800983.

Ignatiouk-Robert:2000:LDJ

[IR00] Irina Ignatiouk-Robert. Large deviations of Jackson networks. Annals of Applied Probability, 10(3):962–1001, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1019487515.

Ignatiouk-Robert:2001:SPL

[IR01] Irina Ignatiouk-Robert. Sample path large deviations and convergence parameters. Annals of Applied Probability, 11(4): 1292–1329, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015345404.

Imkeller:2012:DQB

[IRR12] Peter Imkeller, Anthony Réveillac, and Anja Richter. Differentiability of quadratic BSDEs generated by continuous martingales. Annals of Applied Probability, 22(1):285–336, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1328623701.

Iyengar:2004:EPF

[IS04] Garud Iyengar and Karl Sigman. Exponential penalty function control of loss networks. Annals of Applied Probability, 14(4): 1698–1740, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674075.

Isham:1993:SME

[Ish93] Valerie Isham. Stochastic models for epidemics with special reference to AIDS. Annals of Applied Probability, 3(1):1–27, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005505.

Isham:1995:SMH

[Ish95] Valerie Isham. Stochastic models of host–macroparasite interaction. Annals of Applied Probability, 5(3):720–740, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177004702.

Igloi:1999:BSS

[IT99] E. Iglói and Gy. Terdik. Bilinear stochastic systems with fractional Brownian motion input. Annals of Applied Probability, 9(1):46–77, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962597.

Iyer:2012:NRG

[IT12] Srikanth K. Iyer and Debleena Thacker. Nonuniform random geometric graphs with location-dependent radii. Annals of Applied Probability, 22(5):2048–2066, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1350067993.

Iyer:2017:ANI

[IV17] Srikanth Iyer and Rahul Vaze. Achieving nonzero information velocity in wireless networks. Annals of Applied Probability, 27(1):48–64, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1488790821.

Ivanovs:2018:ZLP

[Iva18] Jevgenijs Ivanovs. Zooming in on a Lévy process at its supremum. Annals of Applied Probability, 28(2):912–940, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1523433628.

Jacka:2002:AOF

[Jac02] Saul Jacka. Avoiding the origin: a finite-fuel stochastic control problem. Annals of Applied Probability, 12(4):1378–1389, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1037125867.

Jaffard:2000:LWS

[Jaf00] Stéphane Jaffard. On lacunary wavelet series. Annals of Applied Probability, 10(1):313–329, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1019737675.

Jagers:1997:CPD

[Jag97] Peter Jagers. Coupling and population dependence in branching processes. Annals of Applied Probability, 7(2):281–298, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034625331.

Jagers:1999:BPD

[Jag99] Peter Jagers. Branching processes with dependence but homogeneous growth. Annals of Applied Probability, 9(4):1160–1174, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1029962868.

Jaillet:1993:CVT

[Jai93] Patrick Jaillet. Cube versus torus models and the Euclidean minimum spanning tree constant. Annals of Applied Probability, 3(2):582–592, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005439.

James:2010:LTL

[Jam10] Lancelot F. James. Lamperti-type laws. Annals of Applied Probability, 20(4):1303–1340, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1279638787.

Jankunas:2001:OCC

[Jan01] Andrius Jankunas. Optimal contingent claims. Annals of Applied Probability, 11(3):735–749, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1015345347.

Janson:2008:LCS

[Jan08] Svante Janson. The largest component in a subcritical random graph with a power law degree distribution. Annals of Applied Probability, 18(4):1651–1668, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1216677136.

Jarai:2017:PTS

[Jár17] Antal A. Járai. Phase transition in a sequential assignment problem on graphs. Annals of Applied Probability, 27(4):2098–2129, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1504080027.

Jaskiewicz:2007:AOR

[Jaś07] Anna Jaśkiewicz. Average optimality for risk-sensitive control with general state space. Annals of Applied Probability, 17(2):654–675, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1174323259.

Jacka:2007:DPM

[JB07] Saul Jacka and Abdelkarem Berkaoui. On the density of properly maximal claims in financial markets with transaction costs. Annals of Applied Probability, 17(2):716–740, April 2007. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1174323262.

Jaber:2021:WST

[JCLP21] Eduardo Abi Jaber, Christa Cuchiero, Martin Larsson, and Sergio Pulido. A weak solution theory for stochastic Volterra equations of convolution type. Annals of Applied Probability, 31(6): 2924–2952, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-6/A-weak-solution-theory-for-stochastic-Volterra-
equations-of-convolution/10.1214/21-AAP1667.full.

Jelenkovic:1999:AAM

[Jel99] Predrag R. Jelenković. Asymptotic approximation of the moveto-front search cost distribution and least-recently used caching fault probabilities. Annals of Applied Probability, 9(2):430–464, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962750.

Jiang:2004:ADL

[Jia04] Tiefeng Jiang. The asymptotic distributions of the largest entries of sample correlation matrices. Annals of Applied Probability, 14(2):865–880, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1082737115.

Jiang:2008:CTI

[Jia08] Long Jiang. Convexity, translation invariance and subadditivity for g-expectations and related risk measures. Annals of Applied Probability, 18(1):245–258, February 2008. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1199890022.

Jiang:2012:TVB

[Jia12] Yunjiang Jiang. Total variation bound for Kac’s random walk. Annals of Applied Probability, 22(4):1712–1727, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1344614209.

Jiang:2015:MTM

[Jia15] Yunjiang Jiang. Mixing time of Metropolis chain based on random transposition walk converging to multivariate Ewens distribution. Annals of Applied Probability, 25(3):1581–1615, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1427124137.

Jiang:2019:DSC

[Jia19] Tiefeng Jiang. Determinant of sample correlation matrix with application. Annals of Applied Probability, 29(3):1356–1397, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1550566833.

Jin:2019:LLC

[Jin19] Ke Jin. The length of the longest common subsequence of two independent Mallows permutations. Annals of Applied Probability, 29(3):1311–1355, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1550566832.

Jammalamadaka:2015:ADM

[JJ15] Sreenivasa Rao Jammalamadaka and Svante Janson. Asymptotic distribution of the maximum interpoint distance in a sample of random vectors with a spherically symmetric distribution. Annals of Applied Probability, 25(6):3571–3591, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1443703782.

Jia:2016:CSC

[JJQ16] Chen Jia, Da-Quan Jiang, and Min-Ping Qian. Cycle symmetries and circulation fluctuations for discrete-time and continuoustime Markov chains. Annals of Applied Probability, 26(4):2454– 2493, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1472745464.

Jahnel:2014:CNI

[JK14] Benedikt Jahnel and Christof Külske. A class of nonergodic interacting particle systems with unique invariant measure. Annals of Applied Probability, 24(6):2595–2643, December 2014. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1409058041.

Jahnel:2017:WRM

[JK17] Benedikt Jahnel and Christof Külske. The Widom–Rowlinson model under spin flip: Immediate loss and sharp recovery of quasilocality. Annals of Applied Probability, 27(6):3845–3892, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1513328715.

Jorritsma:2022:DEG

[JK22] Joost Jorritsma and Júlia Komjáthy. Distance evolutions in growing preferential attachment graphs. Annals of Applied Probability, 32(6):4356–4397, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-6/Distance-evolutions-in-growing-preferential-
attachment-graphs/10.1214/22-AAP1789.full.

Joyce:2002:GLA

[JKK02] Paul Joyce, Stephen M. Krone, and Thomas G. Kurtz. Gaussian limits associated with the Poisson–Dirichlet distribution and the Ewens sampling formula. Annals of Applied Probability, 12 (1):101–124, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015961157.

Joyce:2003:WCO

[JKK03] Paul Joyce, Stephen M. Krone, and Thomas G. Kurtz. When can one detect overdominant selection in the infinite-alleles model? Annals of Applied Probability, 13(1):181–212, January 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1042765666.

Jeanblanc:2007:MMM

[JKM07] Monique Jeanblanc, Susanne Klöppel, and Yoshio Miyahara. Minimal f^{q}-martingale measures for exponential Lévy processes. Annals of Applied Probability, 17(5/6):1615–1638, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1191419178.

Jansen:2015:LDC

[JKM15] Sabine Jansen, Wolfgang König, and Bernd Metzger. Large deviations for cluster size distributions in a continuous classical many-body system. Annals of Applied Probability, 25(2):930– 973, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1424355134.

Jara:2009:LTA

[JKO09] Milton Jara, Tomasz Komorowski, and Stefano Olla. Limit theorems for additive functionals of a Markov chain. Annals of Applied Probability, 19(6):2270–2300, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1259158772.

Jiao:2013:OIU

[JKP13] Ying Jiao, Idris Kharroubi, and Huyên Pham. Optimal investment under multiple defaults risk: A BSDE-decomposition approach. Annals of Applied Probability, 23(2):455–491, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1360682020.

Jagannath:2018:MCI

[JKS18] Aukosh Jagannath, Justin Ko, and Subhabrata Sen. Max κcut and the inhomogeneous Potts spin glass. Annals of Applied Probability, 28(3):1536–1572, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1527840026.

Jentzen:2011:ESN

[JKW11] Arnulf Jentzen, Peter Kloeden, and Georg Winkel. Efficient simulation of nonlinear parabolic SPDEs with additive noise. Annals of Applied Probability, 21(3):908–950, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1307020387.

Janson:2008:ANC

[JL08] Svante Janson and Malwina J. Luczak. Asymptotic normality of the k-core in random graphs. Annals of Applied Probability, 18(3):1085–1137, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1211819795.

Jourdain:2009:RAI

[JL09] Benjamin Jourdain and Jérôme Lelong. Robust adaptive importance sampling for normal random vectors. Annals of Applied Probability, 19(5):1687–1718, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1255699541.

Junge:2022:PSA

[JL22] Matthew Junge and Hanbaek Lyu. The phase structure of asymmetric ballistic annihilation. Annals of Applied Probability, 32 (5):3797–3816, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-5/The-phase-structure-of-asymmetric-ballistic-
annihilation/10.1214/21-AAP1773.full.

Jourdain:2015:OST

[JLM15] Benjamin Jourdain, Tony Lelièvre, and Blazej Miasojedow. Optimal scaling for the transient phase of the random walk Metropolis algorithm: The mean-field limit. Annals of Applied Probability, 25(4):2263–2300, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1432212442.

Jacob:2019:MCP

[JLM19] Emmanuel Jacob, Amitai Linker, and Peter Mörters. Metastability of the contact process on fast evolving scale-free networks. Annals of Applied Probability, 29(5):2654–2699, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1571385619.

Jagannath:2020:STT

[JLM20] Aukosh Jagannath, Patrick Lopatto, and Léo Miolane. Statistical thresholds for tensor PCA. Annals of Applied Probability, 30(4):1910–1933, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-4/Statistical-thresholds-for-tensor-PCA/10.1214/
19-AAP1547.full.

James:2008:DLF

[JLP08] Lancelot F. James, Antonio Lijoi, and Igor Prünster. Distributions of linear functionals of two parameter Poisson–Dirichlet random measures. Annals of Applied Probability, 18(2):521–551, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1206018196.

Jaber:2019:AVP

[JLP19] Eduardo Abi Jaber, Martin Larsson, and Sergio Pulido. Affine Volterra processes. Annals of Applied Probability, 29(5):3155– 3200, October 2019. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
euclid.aoap/1571385632.

Jakubenas:2003:SRP

[JLR03] Paulius Jakubėnas, Shlomo Levental, and Michał Ryznar. The super-replication problem via probabilistic methods. Annals of Applied Probability, 13(2):742–773, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1050689602.

Janson:2012:BPR

[JŁTV12] Svante Janson, Tomasz Łuczak, Tatyana Turova, and Thomas Vallier. Bootstrap percolation on the random graph G_{n,p}. Annals of Applied Probability, 22(5):1989–2047, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1350067992.

Jourdain:2021:CMD

[JLZ21] Benjamin Jourdain, Tony Lelièvre, and Pierre-André Zitt. Convergence of metadynamics: Discussion of the adiabatic hypothesis. Annals of Applied Probability, 31(5):2441–2477, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-5/Convergence-
of-metadynamics-Discussion-of-the-adiabatic-hypothesis/
10.1214/20-AAP1652.full.

Jensen:1991:PEF

[JM91] Jens Ledet Jensen and Jesper Moller. Pseudolikelihood for exponential family models of spatial point processes. Annals of Applied Probability, 1(3):445–461, August 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005877.

Jourdain:2002:AAP

[JM02] B. Jourdain and C. Martini. Approximation of American put prices by European prices via an embedding method. Annals of Applied Probability, 12(1):196–223, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015961161.

Jelenkovic:2003:ALP

[JM03] Predrag Jelenković and Petar Momčilović. Asymptotic loss probability in a finite buffer fluid queue with hetergeneous heavytailed on–off processes. Annals of Applied Probability, 13(2):576– 603, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1050689595.

Jourdain:2008:PCP

[JM08] Benjamin Jourdain and Florent Malrieu. Propagation of chaos and Poincaré inequalities for a system of particles interacting through their cdf. Annals of Applied Probability, 18(5): 1706–1736, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1225372947.

Johnstone:2012:FAT

[JM12] Iain M. Johnstone and Zongming Ma. Fast approach to the Tracy–Widom law at the edge of GOE and GUE. Annals of Applied Probability, 22(5):1962–1988, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1350067991.

Jacob:2015:SPA

[JM15] Emmanuel Jacob and Peter Mörters. Spatial preferential attachment networks: Power laws and clustering coefficients. Annals of Applied Probability, 25(2):632–662, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1424355126.

Jaber:2021:LQC

[JMP21] Eduardo Abi Jaber, Enzo Miller, and Huyên Pham. Linearquadratic control for a class of stochastic Volterra equations: Solvability and approximation. Annals of Applied Probability, 31(5):2244–2274, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-5/Linear-quadratic-control-for-a-class-of-stochastic-
Volterra-equations/10.1214/20-AAP1645.full.

Jacobsen:2009:IPR

[JMRS09] Martin Jacobsen, Thomas Mikosch, Jan Rosiński, and Gennady Samorodnitsky. Inverse problems for regular variation of linear filters, a cancellation property for σ-finite measures and identification of stable laws. Annals of Applied Probability, 19(1):210– 242, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1235140338.

Jeanblanc:2012:MVH

[JMSS12] Monique Jeanblanc, Michael Mania, Marina Santacroce, and Martin Schweizer. Mean-variance hedging via stochastic control and BSDEs for general semimartingales. Annals of Applied Probability, 22(6):2388–2428, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1353695957.

Jaramillo:2021:AFL

[JNP21] Arturo Jaramillo, Ivan Nourdin, and Giovanni Peccati. Approximation of fractional local times: Zero energy and derivatives. Annals of Applied Probability, 31(5):2143–2191, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-5/Approximation-
of-fractional-local-times-Zero-energy-and-derivatives/
10.1214/20-AAP1643.full.

Joffe:1993:NMB

[Jof93] A. Joffe. A new martingale in branching random walk. Annals of Applied Probability, 3(4):1145–1150, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005276.

Jones:1997:CMB

[Jon97] Owen Dafydd Jones. On the convergence of multitype branching processes with varying environments. Annals of Applied Probability, 7(3):772–801, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034801253.

Jonasson:2006:BRT

[Jon06a] Johan Jonasson. Biased random-to-top shuffling. Annals of Applied Probability, 16(2):1034–1058, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1151592258.

Jonasson:2006:OSM

[Jon06b] Johan Jonasson. The overhand shuffle mixes in Θ(n^{2} log n) steps. Annals of Applied Probability, 16(1):231–243, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1141654286.

Jordan:2002:ASC

[Jor02] Jonathan Jordan. Almost sure convergence for iterated functions of independent random variables. Annals of Applied Probability, 12(3):985–1000, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1031863178.

Joseph:2014:CSC

[Jos14] Adrien Joseph. The component sizes of a critical random graph with given degree sequence. Annals of Applied Probability, 24 (6):2560–2594, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1409058040.

Jourdain:2002:PCM

[Jou02] B. Jourdain. Probabilistic characteristics method for a onedimensional inviscid scalar conservation law. Annals of Applied Probability, 12(1):334–360, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015961167.

Johnson:2017:QDP

[JP17] Peter Johnson and Goran Peskir. Quickest detection problems for Bessel processes. Annals of Applied Probability, 27(2):1003– 1056, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1495764373.

Jarrow:2009:NAS

[JPS09] Robert A. Jarrow, Philip Protter, and Hasanjan Sayit. No arbitrage without semimartingales. Annals of Applied Probability, 19(2):596–616, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1241702243.

Jain:2022:FMO

[JPS^{+}22] Vishesh Jain, Natesh S. Pillai, Ashwin Sah, Mehtaab Sawhney, and Aaron Smith. Fast and memory-optimal dimension reduction using Kac’s walk. Annals of Applied Probability, 32(5):4038– 4064, October 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
5/Fast-and-memory-optimal-dimension-reduction-using-
Kacs-walk/10.1214/22-AAP1784.full.

Jones:1999:TFW

[JPV99] Brian D. Jones, Boris G. Pittel, and Joseph S. Verducci. Tree and forest weights and their application to nonuniform random graphs. Annals of Applied Probability, 9(1):197–215, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962602.

Jansons:1992:DBB

[JR92] Kalvis M. Jansons and L. C. G. Rogers. Decomposing the branching Brownian path. Annals of Applied Probability, 2 (4):973–986, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005584.

Jarner:2002:PCR

[JR02] Søren F. Jarner and Gareth O. Roberts. Polynomial convergence rates of Markov chains. Annals of Applied Probability, 12(1):224– 247, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1015961162.

Jacob:2014:WLA

[JR14] Pierre E. Jacob and Robin J. Ryder. The Wang–Landau algorithm reaches the flat histogram criterion in finite time. Annals of Applied Probability, 24(1):34–53, February 2014. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1389278718.

Jaisson:2015:LTN

[JR15] Thibault Jaisson and Mathieu Rosenbaum. Limit theorems for nearly unstable Hawkes processes. Annals of Applied Probability, 25(2):600–631, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355125.

Jaisson:2016:RFD

[JR16] Thibault Jaisson and Mathieu Rosenbaum. Rough fractional diffusions as scaling limits of nearly unstable heavy tailed Hawkes processes. Annals of Applied Probability, 26(5):2860–2882, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1476884306.

Ji:1996:CMS

[JS96] Chuanshu Ji and Lynne Seymour. A consistent model selection procedure for Markov random fields based on penalized pseudolikelihood. Annals of Applied Probability, 6(2):423–443, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968138.

Jones:2007:PLS

[JS07] Matthew O. Jones and Richard F. Serfozo. Poisson limits of sums of point processes and a particle-survivor model. Annals of Applied Probability, 17(1):265–283, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1171377184.

Jenkins:2010:ASF

[JS10] Paul A. Jenkins and Yun S. Song. An asymptotic sampling formula for the coalescent with recombination. Annals of Applied Probability, 20(3):1005–1028, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1276867305.

Jenkins:2012:PAE

[JS12] Paul A. Jenkins and Yun S. Song. Padé approximants and exact two-locus sampling distributions. Annals of Applied Probability, 22(2):576–607, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1333372008.

Jenkins:2017:ESW

[JS17] Paul A. Jenkins and Dario Spanò. Exact simulation of the Wright–Fisher diffusion. Annals of Applied Probability, 27(3): 1478–1509, June 2017. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
euclid.aoap/1500451229.

Jerrum:2004:EBP

[JSTV04] Mark Jerrum, Jung-Bae Son, Prasad Tetali, and Eric Vigoda. Elementary bounds on Poincaré and log-Sobolev constants for decomposable Markov chains. Annals of Applied Probability, 14 (4):1741–1765, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674076.

Junnila:2019:DLC

[JSW19] Janne Junnila, Eero Saksman, and Christian Webb. Decompositions of log-correlated fields with applications. Annals of Applied Probability, 29(6):3786–3820, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1578366326.

Junnila:2020:IMC

[JSW20] Janne Junnila, Eero Saksman, and Christian Webb. Imaginary multiplicative chaos: Moments, regularity and connections to the Ising model. Annals of Applied Probability, 30(5):2099–2164, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-5/Imaginary-
multiplicative-chaos--Moments-regularity-and-connections-
to-the/10.1214/19-AAP1553.full.

Janson:2003:VTP

[JT03] Svante Janson and Johan Tysk. Volatility time and properties of option prices. Annals of Applied Probability, 13(3):890–913, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1060202830.

Jacod:2010:DPV

[JT10] Jean Jacod and Viktor Todorov. Do price and volatility jump together? Annals of Applied Probability, 20(4):1425–1469, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1279638791.

Jacod:2018:LTI

[JT18] Jean Jacod and Viktor Todorov. Limit theorems for integrated local empirical characteristic exponents from noisy highfrequency data with application to volatility and jump activity estimation. Annals of Applied Probability, 28(1):511–576, February 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1520046094.

Janssen:2007:LTG

[JvL07] A. J. E. M. Janssen and J. S. H. van Leeuwaarden. On Lerch’s transcendent and the Gaussian random walk. Annals of Applied Probability, 17(2):421–439, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1174323252.

Janson:2021:PAV

[JW21] Svante Janson and Lutz Warnke. Preferential attachment without vertex growth: Emergence of the giant component. Annals of Applied Probability, 31(4):1523–1547, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/Preferential-
attachment-without-vertex-growth--Emergence-of-the-giant/
10.1214/20-AAP1610.full.

Jin:2014:LSD

[JWB^{+}14] Baisuo Jin, Chen Wang, Z. D. Bai, K. Krishnan Nair, and Matthew Harding. Limiting spectral distribution of a symmetrized auto-cross covariance matrix. Annals of Applied Probability, 24(3):1199–1225, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1398258099.

Jacka:2011:MTD

[JWW11] Saul Jacka, Jon Warren, and Peter Windridge. Minimizing the time to a decision. Annals of Applied Probability, 21(5): 1795–1826, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576609.

Jungel:2017:DBI

[JY17] Ansgar Jüngel and Wen Yue. Discrete beckner inequalities via the Bochner–Bakry–Emery approach for Markov chains. Annals of Applied Probability, 27(4):2238–2269, August 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1504080032.

James:2011:QC

[JZ11] Lancelot F. James and Zhiyuan Zhang. Quantile clocks. Annals of Applied Probability, 21(5):1627–1662, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1319576605.

Kabluchko:2010:SSG

[Kab10] Zakhar Kabluchko. Stationary systems of Gaussian processes. Annals of Applied Probability, 20(6):2295–2317, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1287494561.

Kabluchko:2012:DLH

[Kab12] Zakhar Kabluchko. Distribution of levels in high-dimensional random landscapes. Annals of Applied Probability, 22(1):337– 362, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1328623702.

Kahale:2008:ACP

[Kah08] Nabil Kahale. Analytic crossing probabilities for certain barriers by Brownian motion. Annals of Applied Probability, 18(4):1424– 1440, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1216677127.

Kallblad:2022:DPA

[Käl22] Sigrid Källblad. A dynamic programming approach to distribution-constrained optimal stopping. Annals of Applied Probability, 32(3):1902–1928, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-3/A-dynamic-programming-approach-to-
distribution-constrained-optimal-stopping/10.1214/21-
AAP1724.full.

Kargin:2007:LDI

[Kar07] Vladislav Kargin. A large deviation inequality for vector functions on finite reversible Markov chains. Annals of Applied Probability, 17(4):1202–1221, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1186755237.

Kardaras:2010:NIP

[Kar10a] Constantinos Kardaras. Numéraire-invariant preferences in financial modeling. Annals of Applied Probability, 20(5):1697– 1728, October 2010. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL http://projecteuclid.org/
euclid.aoap/1282747398.

Kargin:2010:PRM

[Kar10b] Vladislav Kargin. Products of random matrices: Dimension and growth in norm. Annals of Applied Probability, 20(3):890–906, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1276867301.

Kardaras:2013:CET

[Kar13] Constantinos Kardaras. On the closure in the Emery topology of semimartingale wealth-process sets. Annals of Applied Probability, 23(4):1355–1376, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1371834031.

Kardaras:2015:SBO

[Kar15] Constantinos Kardaras. On the stochastic behaviour of optional processes up to random times. Annals of Applied Probability, 25(2):429–464, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355119.

Kebaier:2005:SRE

[Keb05] Ahmed Kebaier. Statistical Romberg extrapolation: a new variance reduction method and applications to option pricing. Annals of Applied Probability, 15(4):2681–2705, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1133965776.

Keener:1994:QRL

[Kee94] Robert W. Keener. Quadrature routines for ladder variables. Annals of Applied Probability, 4(2):570–590, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005073.

Kelly:1991:LN

[Kel91] F. P. Kelly. Loss networks. Annals of Applied Probability, 1 (3):319–378, August 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005872.

Kella:1993:PTF

[Kel93] Offer Kella. Parallel and tandem fluid networks with dependent Levy inputs. Annals of Applied Probability, 3(3):682–695, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005358.

Kella:1996:SNF

[Kel96] Offer Kella. Stability and nonproduct form of stochastic fluid networks with Lévy inputs. Annals of Applied Probability, 6 (1):186–199, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968070.

Kelly:2013:UBR

[Kel13] Michael Kelly. Upper bound on the rate of adaptation in an asexual population. Annals of Applied Probability, 23(4):1377– 1408, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1371834032.

Kelly:2016:RPR

[Kel16] David Kelly. Rough path recursions and diffusion approximations. Annals of Applied Probability, 26(1):425–461, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1452003244.

Kendall:2011:GFP

[Ken11] Wilfrid S. Kendall. Geodesics and flows in a Poissonian city. Annals of Applied Probability, 21(3):801–842, June 2011. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1307020384.

Kersting:2012:ADL

[Ker12] Götz Kersting. The asymptotic distribution of the length of betacoalescent trees. Annals of Applied Probability, 22(5):2086–2107, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1350067995.

Kesten:1993:SCF

[Kes93] Harry Kesten. On the speed of convergence in first-passage percolation. Annals of Applied Probability, 3(2):296–338, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005426.

Kunsch:1995:HMR

[KGK95] Hans Kunsch, Stuart Geman, and Athanasios Kehagias. Hidden Markov random fields. Annals of Applied Probability, 5(3):577– 602, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004696.

Kifer:2006:EEBa

[Kif06a] Yuri Kifer. Error estimates for binomial approximations of game options. Annals of Applied Probability, 16(2):984–1033, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1151592257.

Kifer:2006:EEBb

[Kif06b] Yuri Kifer. Error estimates for binomial approximations of game options. Annals of Applied Probability, 16(4):2273–2275, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1169065226.

Karatzas:1996:PCC

[KK96] I. Karatzas and S. G. Kou. On the pricing of contingent claims under constraints. Annals of Applied Probability, 6(2):321–369, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968135.

Khasminskii:2001:LTB

[KK01] R. Z. Khasminskii and F. C. Klebaner. Long term behavior of solutions of the Lotka–Volterra system under small random perturbations. Annals of Applied Probability, 11(3):952–963, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345354.

Komorowski:2004:SLO

[KK04] Tomasz Komorowski and Grzegorz Krupa. On stationarity of Lagrangian observations of passive tracer velocity in a compressible environment. Annals of Applied Probability, 14(4):1666–1697, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1099674074.

Kang:2013:STS

[KK13] Hye-Won Kang and Thomas G. Kurtz. Separation of time-scales and model reduction for stochastic reaction networks. Annals of Applied Probability, 23(2):529–583, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1360682022.

Kang:2009:SSC

[KKLW09] W. N. Kang, F. P. Kelly, N. H. Lee, and R. J. Williams. State space collapse and diffusion approximation for a network operating under a fair bandwidth sharing policy. Annals of Applied Probability, 19(5):1719–1780, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1255699542.

Kluppelberg:2004:RPO

[KKM04] Claudia Klüppelberg, Andreas E. Kyprianou, and Ross A. Maller. Ruin probabilities and overshoots for general Lévy insurance risk processes. Annals of Applied Probability, 14(4): 1766–1801, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674077.

Khorunzhiy:2006:LTS

[KKM06] Oleksiy Khorunzhiy, Werner Kirsch, and Peter Müller. Lifshitz tails for spectra of Erdős–Rényi random graphs. Annals of Applied Probability, 16(1):295–309, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654288.

Kang:1995:SSM

[KKN95] Hyun-Chung Kang, Stephen M. Krone, and Claudia Neuhauser. Stepping-stone models with extinction and recolonization. Annals of Applied Probability, 5(4):1025–1060, November 1995. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177004605.

Kardaras:2015:SLM

[KKN15] Constantinos Kardaras, Dörte Kreher, and Ashkan Nikeghbali. Strict local martingales and bubbles. Annals of Applied Probability, 25(4):1827–1867, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1432212431.

Kuznetsov:2012:MLP

[KKP12] A. Kuznetsov, A. E. Kyprianou, and J. C. Pardo. Meromorphic Lévy processes and their fluctuation identities. Annals of Applied Probability, 22(3):1101–1135, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1337347540.

Kang:2014:CLT

[KKP14] Hye-Won Kang, Thomas G. Kurtz, and Lea Popovic. Central limit theorems and diffusion approximations for multiscale Markov chain models. Annals of Applied Probability, 24(2):721– 759, April 2014. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1394465370.

Kuznetsov:2011:WHM

[KKPvS11] A. Kuznetsov, A. E. Kyprianou, J. C. Pardo, and K. van Schaik. A Wiener–Hopf Monte Carlo simulation technique for Lévy processes. Annals of Applied Probability, 21(6):2171–2190, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1322057319.

Karlin:1991:SLT

[KL91] Samuel Karlin and Ming-Ying Leung. Some limit theorems on distributional patterns of balls in urns. Annals of Applied Probability, 1(4):513–538, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005836.

Kesten:1996:CLT

[KL96] Harry Kesten and Sungchul Lee. The central limit theorem for weighted minimal spanning trees on random points. Annals of Applied Probability, 6(2):495–527, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968141.

Khoshnevisan:1999:SCB

[KL99] Davar Khoshnevisan and Thomas M. Lewis. Stochastic calculus for Brownian motion on a Brownian fracture. Annals of Applied Probability, 9(3):629–667, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962807.

Klebaner:2001:AAE

[KL01] F. C. Klebaner and R. Liptser. Asymptotic analysis and extinction in a stochastic Lotka–Volterra model. Annals of Applied Probability, 11(4):1263–1291, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1015345403.

Kouritzin:2002:CMC

[KL02] Michael A. Kouritzin and Hongwei Long. Convergence of Markov chain approximations to stochastic reaction-diffusion equations. Annals of Applied Probability, 12(3):1039–1070, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1031863180.

Korolyuk:2004:ADA

[KL04] Vladimir S. Korolyuk and Nikolaos Limnios. Average and diffusion approximation of stochastic evolutionary systems in an asymptotic split state space. Annals of Applied Probability, 14 (1):489–516, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1075828059.

Klein:2003:FLL

[Kle03] Irene Klein. Free lunch for large financial markets with continuous price processes. Annals of Applied Probability, 13(4): 1494–1503, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1069786507.

Klein:2006:MFL

[Kle06] Irene Klein. Market free lunch and large financial markets. Annals of Applied Probability, 16(4):2055–2077, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1169065216.

Kliem:2019:RMS

[Kli19] Sandra Kliem. Right marker speeds of solutions to the KPP equation with noise. Annals of Applied Probability, 29(6):3637– 3694, December 2019. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
euclid.aoap/1578366323.

Kloeckner:2019:EBE

[Klo19] Benoît Kloeckner. Effective Berry–Esseen and concentration bounds for Markov chains with a spectral gap. Annals of Applied Probability, 29(3):1778–1807, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1550566842.

Kharroubi:2015:DTA

[KLP15] Idris Kharroubi, Nicolas Langrené, and Huyên Pham. Discrete time approximation of fully nonlinear HJB equations via BS-DEs with nonpositive jumps. Annals of Applied Probability, 25 (4):2301–2338, August 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1432212443.

Kind:1991:DAP

[KLR91] Paolo Kind, Robert Sh. Liptser, and Wolfgang J. Runggaldier. Diffusion approximation in past dependent models and applications to option pricing. Annals of Applied Probability, 1(3):379– 405, August 1991. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005873.

Kruk:2011:HTA

[KLRS11] Łukasz Kruk, John Lehoczky, Kavita Ramanan, and Steven Shreve. Heavy traffic analysis for EDF queues with reneging. Annals of Applied Probability, 21(2):484–545, April 2011. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1300800980.

Kogan:1995:SDB

[KLS95] Y. Kogan, R. Liptser, and M. Shenfild. State-dependent Benes buffer model with fast loading and output rates. Annals of Applied Probability, 5(1):97–120, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004830.

Kleptsina:1997:NFP

[KLS97] M. L. Kleptsina, R. Sh. Liptser, and A. P. Serebrovski. Nonlinear filtering problem with contamination. Annals of Applied Probability, 7(4):917–934, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1043862418.

Kruk:2006:ASS

[KLS06] Łukasz Kruk, John Lehoczky, and Steven Shreve. Accuracy of state space collapse for earliest-deadline-first queues. Annals of Applied Probability, 16(2):516–561, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1151592242.

Kuna:2011:NSC

[KLS11] Tobias Kuna, Joel L. Lebowitz, and Eugene R. Speer. Necessary and sufficient conditions for realizability of point processes. Annals of Applied Probability, 21(4):1253–1281, August 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1312818836.

Krichagina:1993:PCF

[KLST93] Elena V. Krichagina, Sheldon X. C. Lou, Suresh P. Sethi, and Michael I. Taksar. Production control in a failure-prone manufacturing system: Diffusion approximation and asymptotic optimality. Annals of Applied Probability, 3(2):421–453, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005432.

Kruk:2003:MIH

[KLSY03] Lukasz Kruk, John Lehoczky, Steven Shreve, and Shu-Ngai Yeung. Multiple-input heavy-traffic real-time queues. Annals of Applied Probability, 13(1):54–99, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1042765663.

Kruk:2004:EDF

[KLSY04] Łukasz Kruk, John Lehoczky, Steven Shreve, and Shu-Ngai Yeung. Earliest-deadline-first service in heavy-traffic acyclic networks. Annals of Applied Probability, 14(3):1306–1352, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1089736287.

Klebaner:1998:QSD

[KLZ98] Fima C. Klebaner, Justin Lazar, and Ofer Zeitouni. On the quasi-stationary distribution for some randomly perturbed transformations of an interval. Annals of Applied Probability, 8(1):300–315, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1027961045.

Kim:2023:MBB

[KLZ23] Yujin H. Kim, Eyal Lubetzky, and Ofer Zeitouni. The maximum of branching Brownian motion in. Annals of Applied Probability, 33(2):1315–1368, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-2/The-maximum-of-branching-Brownian-motion-in-
Rd/10.1214/22-AAP1848.full.

Kaspi:1995:LBM

[KM95] Haya Kaspi and Avi Mandelbaum. Levy bandits: Multi-armed bandits driven by Levy processes. Annals of Applied Probability, 5(2):541–565, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004777.

Kloppelberg:1996:GLF

[KM96] Claudia Klöppelberg and Thomas Mikosch. Gaussian limit fields for the integrated periodogram. Annals of Applied Probability, 6(3):969–991, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968236.

Kaspi:1998:MAB

[KM98] Haya Kaspi and Avishai Mandelbaum. Multi-armed bandits in discrete and continuous time. Annals of Applied Probability, 8 (4):1270–1290, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903380.

Kontoyiannis:2003:STL

[KM03] I. Kontoyiannis and S. P. Meyn. Spectral theory and limit theorems for geometrically ergodic Markov processes. Annals of Applied Probability, 13(1):304–362, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1042765670.

Kontoyiannis:2008:CEB

[KM08] Ioannis Kontoyiannis and Sean P. Meyn. Computable exponential bounds for screened estimation and simulation. Annals of Applied Probability, 18(4):1491–1518, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1216677129.

Kanoria:2011:MDT

[KM11] Yashodhan Kanoria and Andrea Montanari. Majority dynamics on trees and the dynamic cavity method. Annals of Applied Probability, 21(5):1694–1748, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1319576607.

Khare:2013:CAS

[KM13] Kshitij Khare and Nabanita Mukherjee. Convergence analysis of some multivariate Markov chains using stochastic monotonicity.

Annals of Applied Probability, 23(2):811–833, April 2013. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1360682030.

Kiwi:2018:SGR

[KM18] Marcos Kiwi and Dieter Mitsche. Spectral gap of random hyperbolic graphs and related parameters. Annals of Applied Probability, 28(2):941–989, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1523433629.

Kallsen:2010:USP

[KMK10a] J. Kallsen and J. Muhle-Karbe. On using shadow prices in portfolio optimization with transaction costs. Annals of Applied Probability, 20(4):1341–1358, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1279638788.

Kallsen:2010:UMM

[KMK10b] J. Kallsen and J. Muhle-Karbe. Utility maximization in models with conditionally independent increments. Annals of Applied Probability, 20(6):2162–2177, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1287494557.

Kim:2010:BPM

[KMPT10] Jeong Han Kim, Ravi Montenegro, Yuval Peres, and Prasad Tetali. A birthday paradox for Markov chains with an optimal bound for collision in the Pollard Rho algorithm for discrete logarithm. Annals of Applied Probability, 20(2):495–521, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1268143431.

Kabluchko:2017:GEE

[KMS17] Zakhar Kabluchko, Alexander Marynych, and Henning Sulzbach. General Edgeworth expansions with applications to profiles of random trees. Annals of Applied Probability, 27(6):3478–3524, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1513328706.

Kious:2022:FGG

[KMS22] Daniel Kious, Cécile Mailler, and Bruno Schapira. Finding geodesics on graphs using reinforcement learning. Annals of Applied Probability, 32(5):3889–3929, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Finding-geodesics-on-graphs-using-
reinforcement-learning/10.1214/21-AAP1777.full.

Karnam:2017:DAS

[KMZ17] Chandrasekhar Karnam, Jin Ma, and Jianfeng Zhang. Dynamic approaches for some time-inconsistent optimization problems. Annals of Applied Probability, 27(6):3435–3477, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1513328705.

Knessl:2000:AES

[Kne00] Charles Knessl. Asymptotic expansions for a stochastic model of queue storage. Annals of Applied Probability, 10(2):592–615, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1019487357.

Knispel:2012:ARU

[Kni12] Thomas Knispel. Asymptotics of robust utility maximization. Annals of Applied Probability, 22(1):172–212, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1328623698.

Kumar:2022:WPT

[KNRS22] Chaman Kumar, Neelima, Christoph Reisinger, and Wolfgang Stockinger. Well-posedness and tamed schemes for McKean– Vlasov equations with common noise. Annals of Applied Probability, 32(5):3283–3330, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Well-posedness-and-tamed-schemes-for-
McKeanVlasov-equations-with-common/10.1214/21-AAP1760.
full.

Karatzas:1992:RDD

[KO92] Ioannis Karatzas and Daniel L. Ocone. The resolvent of a degenerate diffusion on the plane, with application to partially observed stochastic control. Annals of Applied Probability, 2 (3):629–668, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005653.

Kolodziejek:2017:LTS

[Kol17] Bartosz Kolodziejek. Logarithmic tails of sums of products of positive random variables bounded by one. Annals of Applied Probability, 27(2):1171–1189, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1495764376.

Kolokoltsov:2022:QMF

[Kol22] Vassili N. Kolokoltsov. Quantum mean-field games. Annals of Applied Probability, 32(3):2254–2288, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-3/Quantum-mean-field-games/10.1214/21-
AAP1733.full.

Komlos:1991:PTS

[KOOS91] J. Komlos, A. Odlyzko, L. Ozarow, and L. A. Shepp. On the properties of a tree-structured server process. Annals of Applied Probability, 1(1):118–125, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005984.

Kotelenez:1992:FNR

[Kot92] Peter Kotelenez. Fluctuations in a nonlinear reaction-diffusion model. Annals of Applied Probability, 2(3):669–694, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005654.

Kotelenez:1995:SNS

[Kot95] Peter Kotelenez. A stochastic Navier–Stokes equation for the vorticity of a two-dimensional fluid. Annals of Applied Probability, 5(4):1126–1160, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004609.

Kirschenhofer:1996:NWD

[KP96] Peter Kirschenhofer and Helmut Prodinger. The number of winners in a discrete geometrically distributed sample. Annals of Applied Probability, 6(2):687–694, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968150.

Komorowski:1997:MGI

[KP97] Tomasz Komorowski and George Papanicolaou. Motion in a Gaussian incompressible flow. Annals of Applied Probability, 7 (1):229–264, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034625261.

Kirschenhofer:1998:ANW

[KP98] Peter Kirschenhofer and Helmut Prodinger. Addendum: “The number of winners in a discrete geometrically distributed sample”. Annals of Applied Probability, 8(2):647, May 1998. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1028903540.

Kyprianou:2003:POC

[KP03] A. E. Kyprianou and M. R. Pistorius. Perpetual options and canadization through fluctuation theory. Annals of Applied Probability, 13(3):1077–1098, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1060202835.

Kluppelberg:2004:TSD

[KP04a] Claudia Klüppelberg and Serguei Pergamenchtchikov. The tail of the stationary distribution of a random coefficient AR (q) model. Annals of Applied Probability, 14(2):971–1005, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1082737119.

Kratz:2004:RGM

[KP04b] Marie F. Kratz and Pierre Picco. A representation of Gibbs measure for the random energy model. Annals of Applied Probability, 14(2):651–677, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1082737106.

Kramkov:2016:SQB

[KP16] Dmitry Kramkov and Sergio Pulido. A system of quadratic BS-DEs arising in a price impact model. Annals of Applied Probability, 26(2):794–817, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1458651820.

Koltchinskii:2003:BGE

[KPL03] Vladimir Koltchinskii, Dmitriy Panchenko, and Fernando Lozano. Bounding the generalization error of convex combinations of classifiers: balancing the dimensionality and the margins. Annals of Applied Probability, 13(1):213–252, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1042765667.

Kyprianou:2010:EAT

[KPR10] A. E. Kyprianou, J. C. Pardo, and V. Rivero. Exact and asymptotic n-tuple laws at first and last passage. Annals of Applied Probability, 20(2):522–564, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1268143432.

Koehler:1998:EFE

[KPS98] J. R. Koehler, A. A. Puhalskii, and B. Simon. Estimating functions evaluated by simulation: a bayesian/analytic approach. Annals of Applied Probability, 8(4):1184–1215, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1028903377.

Krieger:2007:SSR

[KPSC07] Abba M. Krieger, Moshe Pollak, and Ester Samuel-Cahn. Select sets: Rank and file. Annals of Applied Probability, 17(1):360– 385, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1171377187.

Krieger:2010:ELM

[KPSC10] Abba M. Krieger, Moshe Pollak, and Ester Samuel-Cahn. Extreme(ly) mean(ingful): Sequential formation of a quality group. Annals of Applied Probability, 20(6):2261–2294, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1287494560.

Kanzieper:2016:WPL

[KPT^{+}16] Eugene Kanzieper, Mihail Poplavskyi, Carsten Timm, Roger Tribe, and Oleg Zaboronski. What is the probability that a large random matrix has no real eigenvalues? Annals of Applied Probability, 26(5):2733–2753, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1476884302.

Krishnan:2018:TWF

[KQ18] Arjun Krishnan and Jeremy Quastel. Tracy–Widom fluctuations for perturbations of the log-gamma polymer in intermediate disorder. Annals of Applied Probability, 28(6):3736–3764, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1538985634.

Kobylanski:2011:OMS

[KQRM11] Magdalena Kobylanski, Marie-Claire Quenez, and Elisabeth Rouy-Mironescu. Optimal multiple stopping time problem. Annals of Applied Probability, 21(4):1365–1399, August 2011. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1312818839.

Katehakis:1996:FSM

[KR96] Michael N. Katehakis and Uriel G. Rothblum. Finite state multiarmed bandit problems: sensitive-discount, average-reward and average-overtaking optimality. Annals of Applied Probability, 6 (3):1024–1034, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968239.

Kochman:2006:SPK

[KR06] Fred Kochman and Jim Reeds. A simple proof of Kaijser’s unique ergodicity result for hidden Markov α-chains. Annals of Applied Probability, 16(4):1805–1815, November 2006. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1169065208.

Kang:2010:FLM

[KR10] Weining Kang and Kavita Ramanan. Fluid limits of manyserver queues with reneging. Annals of Applied Probability, 20 (6):2204–2260, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1287494559.

Kaspi:2011:LLN

[KR11] Haya Kaspi and Kavita Ramanan. Law of large numbers limits for many-server queues. Annals of Applied Probability, 21(1):33– 114, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1292598028.

Kang:2012:AAS

[KR12a] Weining Kang and Kavita Ramanan. Asymptotic approximations for stationary distributions of many-server queues with abandonment. Annals of Applied Probability, 22(2):477–521, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1333372005.

Kardaras:2012:RMA

[KR12b] Constantinos Kardaras and Scott Robertson. Robust maximization of asymptotic growth. Annals of Applied Probability, 22 (4):1576–1610, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1344614204.

Kaspi:2013:SLM

[KR13] Haya Kaspi and Kavita Ramanan. SPDE limits of many-server queues. Annals of Applied Probability, 23(1):145–229, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1359124384.

Kang:2014:CSD

[KR14] Weining Kang and Kavita Ramanan. Characterization of stationary distributions of reflected diffusions. Annals of Applied Probability, 24(4):1329–1374, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1400073651.

Kortchemski:2019:CCC

[KR19] Igor Kortchemski and Loïc Richier. Condensation in critical Cauchy Bienaymé–Galton–Watson trees. Annals of Applied Probability, 29(3):1837–1877, June 2019. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1550566844.

Kardaras:2021:ERM

[KR21] Constantinos Kardaras and Scott Robertson. Ergodic robust maximization of asymptotic growth. Annals of Applied Probability, 31(4):1787–1819, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-4/Ergodic-robust-maximization-of-asymptotic-
growth/10.1214/20-AAP1634.full.

Klimek:2023:SFV

[KR23] Aleksander Klimek and Tommaso Cornelis Rosati. The spatial Λ-Fleming–Viot process in a random environment. Annals of Applied Probability, 33(3):2426–2492, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-3/The-spatial-%ce%9b-FlemingViot-process-
in-a-random-environment/10.1214/22-AAP1871.full.

Keller-Ressel:2015:EMA

[KRM15] Martin Keller-Ressel and Eberhard Mayerhofer. Exponential moments of affine processes. Annals of Applied Probability, 25(2):714–752, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355129.

Krone:1999:TSC

[Kro99] Stephen M. Krone. The two-stage contact process. Annals of Applied Probability, 9(2):331–351, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962745.

Kaushansky:2023:CTS

[KRSS23] Vadim Kaushansky, Christoph Reisinger, Mykhaylo Shkolnikov, and Zhuo Qun Song. Convergence of a time-stepping scheme to the free boundary in the supercooled Stefan problem. Annals of Applied Probability, 33(1):274–298, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Convergence-

of-a-time-stepping-scheme-to-the-free-boundary/10.1214/
22-AAP1815.full.

Keller-Ressel:2019:APB

[KRSW19] Martin Keller-Ressel, Thorsten Schmidt, and Robert Wardenga. Affine processes beyond stochastic continuity. Annals of Applied Probability, 29(6):3387–3437, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1578366318.

Kroese:1992:CPS

[KS92] Dirk P. Kroese and Volker Schmidt. A continuous polling system with general service times. Annals of Applied Probability, 2(4):906–927, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005580.

Kaj:1993:FPP

[KS93a] Ingemar Kaj and Paavo Salminen. On a first passage problem for branching Brownian motions. Annals of Applied Probability, 3(1):173–185, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005513.

Kook:1993:TST

[KS93b] Kwang Ho Kook and Richard F. Serfozo. Travel and sojourn times in stochastic networks. Annals of Applied Probability, 3 (1):228–252, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005517.

Kwiecinski:1996:SMD

[KS96] Andrzej Kwieciński and Ryszard Szekli. Some monotonicity and dependence properties of self-exciting point processes. Annals of Applied Probability, 6(4):1211–1231, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1035463329.

Karatzas:1999:CSD

[KS99a] Ioannis Karatzas and William D. Sudderth. Control and stopping of a diffusion process on an interval. Annals of Applied Probability, 9(1):188–196, February 1999. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962601.

Kramkov:1999:AEU

[KS99b] D. Kramkov and W. Schachermayer. The asymptotic elasticity of utility functions and optimal investment in incomplete markets. Annals of Applied Probability, 9(3):904–950, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1029962818.

Kramkov:2003:NSC

[KS03a] D. Kramkov and W. Schachermayer. Necessary and sufficient conditions in the problem of optimal investment in incomplete markets. Annals of Applied Probability, 13(4):1504–1516, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1069786508.

Kurkova:2003:MTJ

[KS03b] I. A. Kurkova and Y. M. Suhov. Malyshev’s theory and JSqueues. asymptotics of stationary probabilities. Annals of Applied Probability, 13(4):1313–1354, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1069786501.

Kouritzin:2005:RBP

[KS05] Michael A. Kouritzin and Wei Sun. Rates for branching particle approximations of continuous-discrete filters. Annals of Applied Probability, 15(4):2739–2772, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1133965778.

Kramkov:2006:TTD

[KS06a] Dmitry Kramkov and Mihai Sîrbu. On the two-times differentiability of the value functions in the problem of optimal investment in incomplete markets. Annals of Applied Probability, 16 (3):1352–1384, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1159804984.

Kramkov:2006:SAU

[KS06b] Dmitry Kramkov and Mihai Sîrbu. Sensitivity analysis of utilitybased prices and risk-tolerance wealth processes. Annals of Applied Probability, 16(4):2140–2194, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1169065220.

Khandwawala:2014:BPO

[KS14] Mustafa Khandwawala and Rajesh Sundaresan. Belief propagation for optimal edge cover in the random complete graph. Annals of Applied Probability, 24(6):2414–2454, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1409058036.

Karatzas:2016:DMM

[KS16] Ioannis Karatzas and Andrey Sarantsev. Diverse market models of competing Brownian particles with splits and mergers. Annals of Applied Probability, 26(3):1329–1361, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1465905005.

Kistler:2020:OFP

[KSS20] Nicola Kistler, Adrien Schertzer, and Marius A. Schmidt. Oriented first passage percolation in the mean field limit, 2. The extremal process. Annals of Applied Probability, 30(2):788–811, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1591603222.

Kroshnin:2021:SIB

[KSS21] Alexey Kroshnin, Vladimir Spokoiny, and Alexandra Suvorikova. Statistical inference for Bures–Wasserstein barycenters. Annals of Applied Probability, 31(3):1264–1298, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-3/Statistical-
inference-for-BuresWasserstein-barycenters/10.1214/20-
AAP1618.full.

Kroese:2004:SPT

[KST04] D. P. Kroese, W. R. W. Scheinhardt, and P. G. Taylor. Spectral properties of the tandem Jackson network, seen as a quasibirth-and-death process. Annals of Applied Probability, 14(4): 2057–2089, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674089.

Kunitomo:2003:VAE

[KT03] Naoto Kunitomo and Akihiko Takahashi. On validity of the asymptotic expansion approach in contingent claim analysis. Annals of Applied Probability, 13(3):914–952, August 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1060202831.

Konda:2004:CRL

[KT04] Vijay R. Konda and John N. Tsitsiklis. Convergence rate of linear two-time-scale stochastic approximation. Annals of Applied Probability, 14(2):796–819, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1082737112.

Kazi-Tani:2015:SOB

[KTPZ15] Nabil Kazi-Tani, Dylan Possamaï, and Chao Zhou. Second-order BSDEs with jumps: Formulation and uniqueness. Annals of Applied Probability, 25(5):2867–2908, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1438261055.

Kallblad:2017:OSE

[KTT17] Sigrid Källblad, Xiaolu Tan, and Nizar Touzi. Optimal Skorokhod embedding given full marginals and Azéma–Yor peacocks. Annals of Applied Probability, 27(2):686–719, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1495764364.

Kumar:2000:TSC

[Kum00] Sunil Kumar. Two-server closed networks in heavy traffic: diffusion limits and asymptotic optimality. Annals of Applied Probability, 10(3):930–961, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487514.

Kusuoka:1995:LTO

[Kus95] Shigeo Kusuoka. Limit theorem on option replication cost with transaction costs. Annals of Applied Probability, 5(1):198–221, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004836.

Kuznetsov:2010:WHF

[Kuz10] Alexey Kuznetsov. Wiener–Hopf factorization and distribution of extrema for a family of Lévy processes. Annals of Applied Probability, 20(5):1801–1830, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1282747401.

Klenke:2001:SAR

[KV01] Achim Klenke and Remco Van Der Hofstad. Self-attractive random plymers. Annals of Applied Probability, 11(4):1079–1115, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1015345396.

Kella:1991:QSV

[KW91] Offer Kella and Ward Whitt. Queues with server vacations and Levy processes with secondary jump input. Annals of Applied Probability, 1(1):104–117, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005983.

Kelly:2004:FMN

[KW04] F. P. Kelly and R. J. Williams. Fluid model for a network operating under a fair bandwidth-sharing policy. Annals of Applied Probability, 14(3):1055–1083, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1089736278.

Kang:2007:IPS

[KW07] W. Kang and R. J. Williams. An invariance principle for semimartingale reflecting Brownian motions in domains with piecewise smooth boundaries. Annals of Applied Probability, 17(2):741–779, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1174323263.

Kallianpur:1995:DAN

[KX95] G. Kallianpur and J. Xiong. Diffusion approximation of nuclear space-valued stochastic differential equations driven by Poisson random measures. Annals of Applied Probability, 5(2):493–517, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004775.

Kramkov:2022:OTP

[KX22] Dmitry Kramkov and Yan Xu. An optimal transport problem with backward martingale constraints motivated by insider trading. Annals of Applied Probability, 32(1):294–326, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/An-optimal-
transport-problem-with-backward-martingale-constraints-
motivated-by/10.1214/21-AAP1678.full.

Kella:2010:NFS

[KY10] Offer Kella and Marc Yor. A new formula for some linear stochastic equations with applications. Annals of Applied Probability, 20(2):367–381, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1268143427.

Klebaner:1994:EPC

[KZ94] Fima C. Klebaner and Ofer Zeitouni. The exit problem for a class of density-dependent branching systems. Annals of Applied Probability, 4(4):1188–1205, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004911.

Khare:2009:RCS

[KZ09] Kshitij Khare and Hua Zhou. Rates of convergence of some multivariate Markov chains with polynomial eigenfunctions. Annals of Applied Probability, 19(2):737–777, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1241702249.

Lachal:2003:SPD

[Lac03] Aimé Lachal. Some probability distributions in modeling DNA replication. Annals of Applied Probability, 13(3):1207–1230, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1060202839.

Lacoin:2017:RPT

[Lac17] Hubert Lacoin. The rounding of the phase transition for disordered pinning with stretched exponential tails. Annals of Applied Probability, 27(2):917–943, April 2017. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1495764370.

Lacker:2020:CCL

[Lac20] Daniel Lacker. On the convergence of closed-loop Nash equilibria to the mean field game limit. Annals of Applied Probability, 30(4):1693–1761, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-4/On-the-convergence-of-closed-loop-Nash-equilibria-
to-the/10.1214/19-AAP1541.full.

Lacoin:2022:URS

[Lac22] Hubert Lacoin. A universality result for subcritical complex Gaussian multiplicative chaos. Annals of Applied Probability, 32(1):269–293, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-1/A-universality-result-for-subcritical-complex-
Gaussian-multiplicative-chaos/10.1214/21-AAP1677.full.

Lalley:2000:RMS

[Lal00] Steven P. Lalley. On the rate of mixing for p-shuffles. Annals of Applied Probability, 10(4):1302–1321, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1019487618.

Lamberton:1998:EEB

[Lam98] Damien Lamberton. Error estimates for the binomial approximation of American put options. Annals of Applied Probability, 8(1):206–233, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1027961041.

Lambert:2005:BPL

[Lam05] Amaury Lambert. The branching process with logistic growth. Annals of Applied Probability, 15(2):1506–1535, May 2005. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1115137984.

Lanchier:2012:AMD

[Lan12] Nicolas Lanchier. The Axelrod model for the dissemination of culture revisited. Annals of Applied Probability, 22(2):860–880, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1333372015.

Lanchier:2015:EGL

[Lan15] N. Lanchier. Evolutionary games on the lattice: Payoffs affecting birth and death rates. Annals of Applied Probability, 25(3):1108– 1154, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1427124125.

Larsson:2014:FSS

[Lar14] Martin Larsson. Filtration shrinkage, strict local martingales and the Föllmer measure. Annals of Applied Probability, 24 (4):1739–1766, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1400073662.

Lasserre:2002:BMS

[Las02] Jean B. Lasserre. Bounds on measures satisfying moment conditions. Annals of Applied Probability, 12(3):1114–1137, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1031863183.

Lasserre:2004:BMS

[Las04] Jean B. Lasserre. Bounds on measures satisfying moment conditions. Annals of Applied Probability, 14(3):1603, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1089736298.

Lazrak:2004:GSD

[Laz04] Ali Lazrak. Generalized stochastic differential utility and preference for information. Annals of Applied Probability, 14(4): 2149–2175, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1099674092.

Luczak:1993:SRA

[LC93] Tomasz Luczak and Joel E. Cohen. Stratigraphy of a random acyclic directed graph: The size of trophic levels in the cascade model. Annals of Applied Probability, 3(2):403–420, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005431.

Leslie:2003:CMT

[LC03] David S. Leslie and E. J. Collins. Convergent multiple-timescales reinforcement learning algorithms in normal form games. Annals of Applied Probability, 13(4):1231–1251, November 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1069786497.

Louart:2022:CMR

[LC22] Cosme Louart and Romain Couillet. A concentration of measure and random matrix approach to large-dimensional robust statistics. Annals of Applied Probability, 32(6):4737–4762, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/A-concentration-
of-measure-and-random-matrix-approach-to-large/10.1214/
22-AAP1801.full.

Lionnet:2015:TDF

[LdRS15] Arnaud Lionnet, Gonçalo dos Reis, and Lukasz Szpruch. Time discretization of FBSDE with polynomial growth drivers and reaction-diffusion PDEs. Annals of Applied Probability, 25(5): 2563–2625, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1438261049.

Lionnet:2018:CQP

[LdRS18] Arnaud Lionnet, Goncalo dos Reis, and Lukasz Szpruch. Convergence and qualitative properties of modified explicit schemes for BSDEs with polynomial growth. Annals of Applied Probability, 28(4):2544–2591, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1533780280.

Le:2015:DPA

[Le15] Huiling Le. A diffusion process associated with Fréchet means. Annals of Applied Probability, 25(6):3033–3046, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1443703768.

Lee:1993:IGL

[Lee93] Sungchul Lee. An inequality for greedy lattice animals. Annals of Applied Probability, 3(4):1170–1188, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005278.

Lee:1997:CLT

[Lee97] Sungchul Lee. The central limit theorem for Euclidean minimal spanning trees I. Annals of Applied Probability, 7(4):996–1020, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1043862422.

Lefebvre:1991:FSP

[Lef91] Mario Lefebvre. Forcing a stochastic process to stay in or to leave a given region. Annals of Applied Probability, 1(1):167–172, February 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005986.

Lefebvre:2004:HPD

[Lef04] Mario Lefebvre. A homing problem for diffusion processes with control-dependent variance. Annals of Applied Probability, 14(2): 786–795, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1082737111.

Lehmann:2023:DEB

[Leh23] Tobias Lehmann. Darwinian evolution as Brownian motion on the simplex: a geometric perspective on stochastic replicator dynamics. Annals of Applied Probability, 33(1):344–375, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Darwinian-
evolution-as-Brownian-motion-on-the-simplex--A/10.1214/
22-AAP1817.full.

Leitner:2008:CPG

[Lei08] Johannes Leitner. Convex pricing by a generalized entropy penalty. Annals of Applied Probability, 18(2):620–631, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1206018199.

Lejay:2016:SBM

[Lej16] Antoine Lejay. The snapping out Brownian motion. Annals of Applied Probability, 26(3):1727–1742, June 2016. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905017.

Lezaud:1998:CTB

[Lez98] Pascal Lezaud. Chernoff-type bound for finite Markov chains. Annals of Applied Probability, 8(3):849–867, August 1998. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1028903453.

Li:2018:PCR

[Li18] Yao Li. On the polynomial convergence rate to nonequilibrium steady states. Annals of Applied Probability, 28(6):3765–3812, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1538985635.

Liggett:1995:SDT

[Lig95] Thomas M. Liggett. Survival of discrete time growth models, with applications to oriented percolation. Annals of Applied Probability, 5(3):613–636, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004698.

Limic:2001:LQH

[Lim01] Vlada Limic. A LIFO queue in heavy traffic. Annals of Applied Probability, 11(2):301–331, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1015345294.

Litterer:2012:HOR

[LL12] C. Litterer and T. Lyons. High order recombination and an application to cubature on Wiener space. Annals of Applied Probability, 22(4):1301–1327, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1344614196.

Loh:2013:SCL

[LL13] Po-Shen Loh and Eyal Lubetzky. Stochastic coalescence in logarithmic time. Annals of Applied Probability, 23(2):492–528, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1360682021.

Laslier:2017:RLC

[LL17] Benoît Laslier and Jean-François Laslier. Reinforcement learning from comparisons: Three alternatives are enough, two are not. Annals of Applied Probability, 27(5):2907–2925, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1509696037.

Labbe:2020:MTC

[LL20a] Cyril Labbé and Hubert Lacoin. Mixing time and cutoff for the weakly asymmetric simple exclusion process. Annals of Applied Probability, 30(4):1847–1883, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-4/Mixing-time-and-cutoff-for-the-weakly-
asymmetric-simple-exclusion/10.1214/19-AAP1545.full.

Lopes:2020:ETW

[LL20b] Fabio Lopes and Malwina Luczak. Extinction time for the weaker of two competing SIS epidemics. Annals of Applied Probability, 30(6):2880–2922, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-6/Extinction-time-for-the-weaker-of-
two-competing-SIS-epidemics/10.1214/20-AAP1576.full.

Lacker:2023:CLC

[LL23] Daniel Lacker and Luc Le Flem. Closed-loop convergence for mean field games with common noise. Annals of Applied Probability, 33(4):2681–2733, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-4/Closed-loop-convergence-for-mean-field-
games-with-common-noise/10.1214/22-AAP1876.full.

Louart:2018:RMA

[LLC18] Cosme Louart, Zhenyu Liao, and Romain Couillet. A random matrix approach to neural networks. Annals of Applied Probability, 28(2):1190–1248, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1523433634.

Li:2018:MDR

[LLLZ18] Xiaoou Li, Jingchen Liu, Jianfeng Lu, and Xiang Zhou. Moderate deviation for random elliptic PDE with small noise. Annals of Applied Probability, 28(5):2781–2813, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1535443234.

Liu:2008:ADB

[LLS08] Wei-Dong Liu, Zhengyan Lin, and Qi-Man Shao. The asymptotic distribution and Berry–Esseen bound of a new test for independence in high dimension with an application to stochastic optimization. Annals of Applied Probability, 18(6):2337–2366, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1227708921.

Lent:1996:TGS

[LM96] Janice Lent and Hosam M. Mahmoud. On tree-growing search strategies. Annals of Applied Probability, 6(4):1284–1302, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1035463333.

Lowe:2002:SRT

[LM02] Matthias Löwe and Heinrich Matzinger. Scenery reconstruction in two dimensions with many colors. Annals of Applied Probability, 12(4):1322–1347, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1037125865.

Luczak:2005:PTC

[LM05] Malwina J. Luczak and Colin McDiarmid. On the power of two choices: Balls and bins in continuous time. Annals of Applied Probability, 15(3):1733–1764, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1121433767.

Lejay:2006:SSO

[LM06] Antoine Lejay and Miguel Martinez. A scheme for simulating one-dimensional diffusion processes with discontinuous coefficients. Annals of Applied Probability, 16(1):107–139, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1141654283.

Liu:2009:ORI

[LM09] Yuping Liu and Jin Ma. Optimal reinsurance/investment problems for general insurance models. Annals of Applied Probability, 19(4):1495–1528, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1248700625.

Luczak:2015:BRR

[LM15] Malwina J. Luczak and Colin McDiarmid. Balanced routing of random calls. Annals of Applied Probability, 25(3):1279–1324, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1427124129.

Lyberopoulos:2021:CME

[LM21] Demetrios P. Lyberopoulos and Nikolaos D. Macheras. A characterization of martingale-equivalent mixed compound Poisson processes. Annals of Applied Probability, 31(2):778–805, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/A-characterization-
of-martingale-equivalent-mixed-compound-Poisson-processes/
10.1214/20-AAP1604.full.

Lund:1996:CEC

[LMT96] Robert B. Lund, Sean P. Meyn, and Richard L. Tweedie. Computable exponential convergence rates for stochastically ordered Markov processes. Annals of Applied Probability, 6(1):218–237, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968072.

Lember:2012:RCM

[LMT12] Jüri Lember, Heinrich Matzinger, and Felipe Torres. The rate of the convergence of the mean score in random sequence comparison. Annals of Applied Probability, 22(3):1046–1058, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1337347538.

Luczak:2005:SAS

[LN05] Malwina J. Luczak and James Norris. Strong approximation for the supermarket model. Annals of Applied Probability, 15 (3):2038–2061, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433776.

Lanchier:2006:SEM

[LN06a] N. Lanchier and C. Neuhauser. A spatially explicit model for competition among specialists and generalists in a heterogeneous environment. Annals of Applied Probability, 16(3):1385–1410, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804985.

Lanchier:2006:SSM

[LN06b] N. Lanchier and C. Neuhauser. Stochastic spatial models of host-pathogen and host-mutualist interactions I. Annals of Applied Probability, 16(1):448–474, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654294.

Lanchier:2009:SEN

[LN09] N. Lanchier and C. Neuhauser. Spatially explicit non-Mendelian diploid model. Annals of Applied Probability, 19(5):1880–1920, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1255699547.

Luczak:2013:AFV

[LN13] M. J. Luczak and J. R. Norris. Averaging over fast variables in the fluid limit for Markov chains: Application to the supermarket model with memory. Annals of Applied Probability, 23(3):957–986, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1362684851.

Last:2021:RCM

[LNS21] Günter Last, Franz Nestmann, and Matthias Schulte. The random connection model and functions of edge-marked Poisson processes: Second order properties and normal approximation. Annals of Applied Probability, 31(1):128–168, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/The-
random-connection-model-and-functions-of-edge-marked-
Poisson/10.1214/20-AAP1585.full.

LeGland:2004:SUA

[LO04] François Le Gland and Nadia Oudjane. Stability and uniform approximation of nonlinear filters using the Hilbert metric and application to particle filters. Annals of Applied Probability, 14 (1):144–187, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1075828050.

Leao:2013:WAW

[LO13] Dorival Leão and Alberto Ohashi. Weak approximations for Wiener functionals. Annals of Applied Probability, 23(4):1660– 1691, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.

aoap/1371834041. See corrigendum [LO17].

Leao:2017:CWA

[LO17] Dorival Leão and Alberto Ohashi. Corrigendum: “Weak approximations for Wiener functionals” [Ann. Appl. Probab. (2013) 23, 1660–1691]. Annals of Applied Probability, 27(2):1294–1295, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1495764380. See [LO13].

Loader:1992:BCP

[Loa92] Clive R. Loader. Boundary crossing probabilities for locally Poisson processes. Annals of Applied Probability, 2(1):199–228, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005778.

Loeffen:2008:OBS

[Loe08] R. L. Loeffen. On optimality of the barrier strategy in de Finetti’s dividend problem for spectrally negative Lévy processes. Annals of Applied Probability, 18(5):1669–1680, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1225372945.

Loeper:2018:OPL

[Loe18] Gregoire Loeper. Option pricing with linear market impact and nonlinear Black–Scholes equations. Annals of Applied Probability, 28(5):2664–2726, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1535443231.

Loh:1992:SMM

[Loh92] Wei-Liem Loh. Stein’s method and multinomial approximation. Annals of Applied Probability, 2(3):536–554, August 1992. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005648.

Lokka:2004:SPD

[LØP04] Arne Løkka, Bernt Øksendal, and Frank Proske. Stochastic partial differential equations driven by Lévy space-time white noise. Annals of Applied Probability, 14(3):1506–1528, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1089736294.

Lowe:1998:SCH

[Löw98] Matthias Löwe. On the storage capacity of Hopfield models with correlated patterns. Annals of Applied Probability, 8(4): 1216–1250, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903378.

Leon:2004:OHB

[LP04] Carlos A. León and François Perron. Optimal Hoeffding bounds for discrete reversible Markov chains. Annals of Applied Probability, 14(2):958–970, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1082737118.

Luschgy:2008:FQR

[LP08] Harald Luschgy and Gilles Pagès. Functional quantization rate and mean regularity of processes with an application to Lévy processes. Annals of Applied Probability, 18(2):427–469, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1206018193.

Lemaire:2010:URI

[LP10] Vincent Lemaire and Gilles Pagès. Unconstrained recursive importance sampling. Annals of Applied Probability, 20(3):1029– 1067, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1276867306.

Lambert:2013:CPP

[LP13a] Amaury Lambert and Lea Popovic. The coalescent point process of branching trees. Annals of Applied Probability, 23(1):99–144, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1359124383.

Laruelle:2013:RUM

[LP13b] Sophie Laruelle and Gilles Pagès. Randomized urn models revisited using stochastic approximation. Annals of Applied Probability, 23(4):1409–1436, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1371834033. See addendum

and corrigendum [LP17a].

Laruelle:2017:ACR

[LP17a] Sophie Laruelle and Gilles Pagès. Addendum and corrigendum to “Randomized urn models revisited using stochastic approximation”. Annals of Applied Probability, 27(2):1296–1298, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1495764381. See [LP13b].

Lux:2017:IFH

[LP17b] Thibaut Lux and Antonis Papapantoleon. Improved Fréchet– Hoeffding bounds on d-copulas and applications in model-free finance. Annals of Applied Probability, 27(6):3633–3671, 12 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Lucon:2021:PIN

[LP21] Eric Luçon and Christophe Poquet. Periodicity induced by noise and interaction in the kinetic mean-field FitzHugh–Nagumo model. Annals of Applied Probability, 31(2):561–593, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/Periodicity-
induced-by-noise-and-interaction-in-the-kinetic-mean/
10.1214/20-AAP1598.full.

Lambert:2023:SAG

[LP23] Gaultier Lambert and Elliot Paquette. Strong approximation of Gaussian β ensemble characteristic polynomials: The hyperbolic regime. Annals of Applied Probability, 33(1):549–612, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Strong-
approximation-of-Gaussian-%ce%b2-ensemble-characteristic-
polynomials--The/10.1214/22-AAP1823.full.

Lorig:2015:FDE

[LPP15] Matthew Lorig, Stefano Pagliarani, and Andrea Pascucci. A family of density expansions for Lévy-type processes. Annals of Applied Probability, 25(1):235–267, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1418740185.

Lambert:2021:CPH

[LPS21] Amaury Lambert, Verónica Miró Pina, and Emmanuel Schertzer. Chromosome painting: How recombination mixes ancestral colors. Annals of Applied Probability, 31(2):826–864, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/Chromosome-
painting-How-recombination-mixes-ancestral-colors/10.
1214/20-AAP1606.full.

Landim:2023:NSD

[LPSX23] Claudio Landim, Carlos G. Pacheco, Sunder Sethuraman, and Jianfei Xue. On a nonlinear SPDE derived from a hydrodynamic limit in a Sinai-type random environment. Annals of Applied Probability, 33(1):200–237, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-1/On-a-nonlinear-SPDE-derived-from-a-
hydrodynamic-limit-in/10.1214/22-AAP1813.full.

Lamberton:2004:WCT

[LPT04] Damien Lamberton, Gilles Pagès, and Pierre Tarrès. When can the two-armed bandit algorithm be trusted? Annals of Applied Probability, 14(3):1424–1454, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1089736291.

Lijoi:2008:BNE

[LPW08] Antonio Lijoi, Igor Prünster, and Stephen G. Walker. Bayesian nonparametric estimators derived from conditional Gibbs structures. Annals of Applied Probability, 18(4):1519–1547, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1216677130. See note [LPW14].

Lijoi:2014:NBN

[LPW14] Antonio Lijoi, Igor Prünster, and Stephen G. Walker. A note on “Bayesian nonparametric estimators derived from conditional Gibbs structures”. Annals of Applied Probability, 24(1):447–448, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1389278731. See [LPW08].

Last:2017:CFI

[LPZ17] Günter Last, Mathew D. Penrose, and Sergei Zuyev. On the capacity functional of the infinite cluster of a Boolean model. Annals of Applied Probability, 27(3):1678–1701, June 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1500451238.

Li:1999:ABI

[LR99] Deli Li and T. D. Rogers. Asymptotic behavior for iterated functions of random variables. Annals of Applied Probability, 9 (4):1175–1201, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962869.

Liu:2000:LTM

[LR00] Quansheng Liu and Alain Rouault. Limit theorems for Mandelbrot’s multiplicative cascades. Annals of Applied Probability, 10(1):218–239, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019737670;   http://
www.jstor.org/stable/2667193.

Li:2006:SSL

[LR06] Deli Li and Andrew Rosalsky. Some strong limit theorems for the largest entries of sample correlation matrices. Annals of Applied Probability, 16(1):423–447, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654293.

Litvak:2012:SAC

[LR12] Nelly Litvak and Philippe Robert. A scaling analysis of a cat and mouse Markov chain. Annals of Applied Probability, 22(2):792– 826, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1333372013.

Lindberg:2013:EDR

[LR13] Carl Lindberg and Holger Rootzén. Error distributions for random grid approximations of multidimensional stochastic integrals. Annals of Applied Probability, 23(2):834–857, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1360682031.

Li:2014:PEF

[LR14] Libo Li and Marek Rutkowski. Progressive enlargements of filtrations with pseudo-honest times. Annals of Applied Probability, 24(4):1509–1553, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1400073656.

Lipshutz:2018:DDS

[LR18] David Lipshutz and Kavita Ramanan. On directional derivatives of Skorokhod maps in convex polyhedral domains. Annals of Applied Probability, 28(2):688–750, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1523433623.

Lachieze-Rey:2019:NCN

[LR19] Raphaël Lachièze-Rey. Normal convergence of nonlocalised geometric functionals and shot-noise excursions. Annals of Applied Probability, 29(5):2613–2653, October 2019. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1571385618.

Lagarias:2021:BLM

[LR21] Jeffrey C. Lagarias and Brad Rodgers. Band-limited mimicry of point processes by point processes supported on a lattice. Annals of Applied Probability, 31(1):351–376, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/Band-
limited-mimicry-of-point-processes-by-point-processes-
supported/10.1214/20-AAP1592.full.

Leadbetter:1998:DTA

[LRdH98] M. Ross Leadbetter, Holger Rootzén, and Laurens de Haan. On the distribution of tail array sums for strongly mixing stationary sequences. Annals of Applied Probability, 8(3):868–885, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903454.

Lachieze-Rey:2015:RCR

[LRM15] Raphael Lachieze-Rey and Ilya Molchanov. Regularity conditions in the realisability problem with applications to point processes and random closed sets. Annals of Applied Probability, 25(1):116–149, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1418740181.

Lachieze-Rey:2017:NBE

[LRP17] Raphaël Lachièze-Rey and Giovanni Peccati. New Berry–Esseen bounds for functionals of binomial point processes. Annals of Applied Probability, 27(4):1992–2031, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1504080024.

Lachieze-Rey:2022:QTS

[LRPY22] Raphaël Lachièze-Rey, Giovanni Peccati, and Xiaochuan Yang. Quantitative two-scale stabilization on the Poisson space. Annals of Applied Probability, 32(4):3085–3145, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-4/Quantitative-

two-scale-stabilization-on-the-Poisson-space/10.1214/
21-AAP1768.full.

Latuszynski:2013:AGS

[ŁRR13] Krzysztof Łatuszyński, Gareth O. Roberts, and Jeffrey S. Rosenthal. Adaptive Gibbs samplers and related MCMC methods. Annals of Applied Probability, 23(1):66–98, February 2013. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1359124382.

Ley:2017:DBN

[LRS17] Christophe Ley, Gesine Reinert, and Yvik Swan. Distances between nested densities and a measure of the impact of the prior in Bayesian statistics. Annals of Applied Probability, 27(1):216–241, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1488790827.

Lachieze-Rey:2019:NAS

[LRSY19] Raphaël Lachièze-Rey, Matthias Schulte, and J. E. Yukich. Normal approximation for stabilizing functionals. Annals of Applied Probability, 29(2):931–993, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1548298934.

Latouche:2003:TMA

[LRT03] Guy Latouche, Marie-Ange Remiche, and Peter Taylor. Transient Markov arrival processes. Annals of Applied Probability, 13(2):628–640, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1050689597.

Lacker:2023:LWC

[LRW23] Daniel Lacker, Kavita Ramanan, and Ruoyu Wu. Local weak convergence for sparse networks of interacting processes. Annals of Applied Probability, 33(2):643–688, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Local-
weak-convergence-for-sparse-networks-of-interacting-
processes/10.1214/22-AAP1830.full.

Luen:2006:NPT

[LRZ06] Brad Luen, Kavita Ramanan, and Ilze Ziedins. Nonmonotonicity of phase transitions in a loss network with controls. Annals of Applied Probability, 16(3):1528–1562, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1159804990.

Levental:1995:NSC

[LS95] Shlomo Levental and Antolii V. Skorohod. A necessary and sufficient condition for absence of arbitrage with tame portfolios. Annals of Applied Probability, 5(4):906–925, November 1995. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177004599.

Levental:1997:PHO

[LS97] Shlomo Levental and Anatolii V. Skorohod. On the possibility of hedging options in the presence of transaction costs. Annals of Applied Probability, 7(2):410–443, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034625338.

Last:2000:SSF

[LS00] G. Last and R. Schassberger. On stationary stochastic flows and Palm probabilities of surface processes. Annals of Applied Probability, 10(2):463–492, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487351.

Lahiri:2009:BET

[LS09] S. N. Lahiri and S. Sun. A Berry–Esseen theorem for sample quantiles under weak dependence. Annals of Applied Probability, 19(1):108–126, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1235140334.

Lanchier:2013:FOD

[LS13] Nicolas Lanchier and Stylianos Scarlatos. Fixation in the onedimensional Axelrod model. Annals of Applied Probability, 23 (6):2538–2559, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1382447697.

Lucon:2014:MFL

[LS14] Eric Luçon and Wilhelm Stannat. Mean field limit for disordered diffusions with singular interactions. Annals of Applied Probability, 24(5):1946–1993, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1403812367.

Lang:2015:IGR

[LS15] Annika Lang and Christoph Schwab. Isotropic Gaussian random fields on the sphere: Regularity, fast simulation and stochastic partial differential equations. Annals of Applied Probability, 25 (6):3047–3094, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1443703769.

Lee:2016:TWD

[LS16a] Ji Oon Lee and Kevin Schnelli. Tracy–Widom distribution for the largest eigenvalue of real sample covariance matrices with general population. Annals of Applied Probability, 26(6): 3786–3839, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1481792600.

Lucon:2016:TGN

[LS16b] Eric Luçon and Wilhelm Stannat. Transition from Gaussian to non-Gaussian fluctuations for mean-field diffusions in spatial interaction. Annals of Applied Probability, 26(6):3840–3909, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1481792601.

Lalley:2017:CPR

[LS17a] Steven Lalley and Wei Su. Contact processes on random regular graphs. Annals of Applied Probability, 27(4):2061–2097, August 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1504080026.

Leobacher:2017:SOM

[LS17b] Gunther Leobacher and Michaela Szölgyenyi. A strong order 1/2 method for multidimensional SDEs with discontinuous drift. Annals of Applied Probability, 27(4):2383–2418, August 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1504080036. See

correction [LS19].

Leon:2018:HSF

[LS18] José R. León and Adeline Samson. Hypoelliptic stochastic FitzHugh–Nagumo neuronal model: Mixing, up-crossing and estimation of the spike rate. Annals of Applied Probability, 28 (4):2243–2274, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1533780272.

Leobacher:2019:CNS

[LS19] Gunther Leobacher and Michaela Szölgyenyi. Correction note: A strong order 1/2 method for multidimensional SDEs with discontinuous drift. Annals of Applied Probability, 29(5):3266–3269, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/

1571385635. See [LS17b].

Landon:2020:AMC

[LS20] Benjamin Landon and Philippe Sosoe. Applications of mesoscopic CLTs in random matrix theory. Annals of Applied Probability, 30(6):2769–2795, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-6/Applications-of-mesoscopic-CLTs-in-
random-matrix-theory/10.1214/20-AAP1572.full.

Li:2022:SFS

[LS22] Xue-Mei Li and Julian Sieber. Slow-fast systems with fractional environment and dynamics. Annals of Applied Probability, 32(5): 3964–4003, October 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
5/Slow-fast-systems-with-fractional-environment-and-
dynamics/10.1214/22-AAP1779.full.

Leimkuhler:2023:SRW

[LST23] Benedict Leimkuhler, Akash Sharma, and Michael V. Tretyakov. Simplest random walk for approximating Robin boundary value problems and ergodic limits of reflected diffusions. Annals of Applied Probability, 33(3):1904–1960, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/Simplest-
random-walk-for-approximating-Robin-boundary-value-problems-
and/10.1214/22-AAP1856.full.

Lange:1997:PSM

[LSZ97] Kenneth Lange, Terence P. Speed, and Hongyu Zhao. The Poisson-skip model of crossing-over. Annals of Applied Probability, 7(2):299–313, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034625332.

Lambert:2013:SLE

[LSZ13] Amaury Lambert, Florian Simatos, and Bert Zwart. Scaling limits via excursion theory: Interplay between Crump–Mode–Jagers branching processes and processor-sharing queues. Annals of Applied Probability, 23(6):2357–2381, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1382447691.

Litvak:2018:OSV

[LT18] Alexander E. Litvak and Konstantin Tikhomirov. Order statistics of vectors with dependent coordinates, and the Karhunen– Loève basis. Annals of Applied Probability, 28(4):2083–2104, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1533780268.

Liu:2019:FOE

[LT19] Yanghui Liu and Samy Tindel. First-order Euler scheme for SDEs driven by fractional Brownian motions: The rough case. Annals of Applied Probability, 29(2):758–826, April 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1548298930.

Lerouvillois:2022:HLI

[LT22] Vincent Lerouvillois and Fabio Toninelli. Hydrodynamic limit for a 2D interlaced particle process. Annals of Applied Probability, 32(1):167–190, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-1/Hydrodynamic-limit-for-a-2D-interlaced-
particle-process/10.1214/21-AAP1674.full.

Louidor:2014:WBM

[LTVR14] Oren Louidor, Ran Tessler, and Alexander Vandenberg-Rodes. The Williams–Bjerknes model on regular trees. Annals of Applied Probability, 24(5):1889–1917, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1403812365.

Ludena:2008:VMF

[Lud08] Carenne Ludeña. L^{p}-variations for multifractal fractional random walks. Annals of Applied Probability, 18(3):1138–1163, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1211819796.

Lamberton:2003:CPN

[LV03] Damien Lamberton and Stéphane Villeneuve. Critical price near maturity for an American option on a dividend-paying stock. Annals of Applied Probability, 13(2):800–815, May 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1050689604.

Limic:2010:VCL

[LV10] Vlada Limic and Stanislav Volkov. VRRW on complete-like graphs: Almost sure behavior. Annals of Applied Probability, 20 (6):2346–2388, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1287494563.

Litvak:2004:MTT

[LvZ04] Nelly Litvak and Willem R. van Zwet. On the minimal travel time needed to collect n items on a circle. Annals of Applied Probability, 14(2):881–902, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1082737116.

Lagarias:1992:PTS

[LW92] J. C. Lagarias and A. Weiss. The 3x + 1 problem: Two stochastic models. Annals of Applied Probability, 2(1):229–261, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005779.

Liu:2014:MSH

[LW14] Yunan Liu and Ward Whitt. Many-server heavy-traffic limit for queues with time-varying parameters. Annals of Applied Probability, 24(1):378–421, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278729.

Leite:2019:CLA

[LW19] Saul C. Leite and Ruth J. Williams. A constrained Langevin approximation for chemical reaction networks. Annals of Applied Probability, 29(3):1541–1608, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1550566837.

Lu:2022:E

[LW22] Jianfeng Lu and Lihan Wang. On explicit L^{2}-convergence rate estimate for piecewise deterministic Markov processes in MCMC algorithms. Annals of Applied Probability, 32(2):1333–1361, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/On-explicit-
L2-convergence-rate-estimate-for-piecewise-deterministic-
Markov/10.1214/21-AAP1710.full.

Liu:2014:CDE

[LX14] Jingchen Liu and Gongjun Xu. On the conditional distributions and the efficient simulations of exponential integrals of Gaussian random fields. Annals of Applied Probability, 24(4):1691–1738, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1400073661.

Li:2017:PCE

[LY17] Yao Li and Lai-Sang Young. Polynomial convergence to equilibrium for a system of interacting particles. Annals of Applied Probability, 27(1):65–90, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1488790822.

Lacoin:2022:MEB

[LY22] Hubert Lacoin and Shangjie Yang. Metastability for expanding bubbles on a sticky substrate. Annals of Applied Probability, 32 (5):3408–3449, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-5/Metastability-for-expanding-bubbles-on-a-sticky-
substrate/10.1214/21-AAP1763.full.

Li:2019:GCS

[LYZ19] Pei-Sen Li, Xu Yang, and Xiaowen Zhou. A general continuousstate nonlinear branching process. Annals of Applied Probability, 29(4):2523–2555, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1563869049.

Lyons:1998:NTG

[LZ98] Russell Lyons and Kevin Zumbrun. Normality of tree-growing search strategies. Annals of Applied Probability, 8(1):112–130, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1027961036.

Li:2006:CTM

[LZ06] Xun Li and Xun Yu Zhou. Continuous-time mean-variance efficiency: the 80% rule. Annals of Applied Probability, 16(4): 1751–1763, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1169065206.

Larsen:2013:UMU

[LŽ13] Kasper Larsen and Gordan Žitković. On utility maximization under convex portfolio constraints. Annals of Applied Probability, 23(2):665–692, April 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1360682026.

Li:2021:GCT

[LZL21] Jia Li, Congshan Zhang, and Yunxiao Liu. Glivenko-Cantelli theorems for integrated functionals of stochastic processes. Annals of Applied Probability, 31(4):1914–1943, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/GlivenkoCantelli-
theorems-for-integrated-functionals-of-stochastic-processes/
10.1214/20-AAP1637.full.

Maglaras:2000:DRP

[Mag00] Constantinos Maglaras. Discrete-review policies for scheduling stochastic networks: trajectory tracking and fluid-scale asymptotic optimality. Annals of Applied Probability, 10(3):897–929, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1019487513.

Mahmoud:1994:SLH

[Mah94] Hosam M. Mahmoud. A strong law for the height of random binary pyramids. Annals of Applied Probability, 4(3):923–932, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004977.

Majewski:2006:SPL

[Maj06] Kurt Majewski. Sample path large deviations for multiclass feedforward queueing networks in critical loading. Annals of Applied Probability, 16(4):1893–1924, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1169065211.

Malrieu:2003:CEG

[Mal03] Florent Malrieu. Convergence to equilibrium for granular media equations and their Euler schemes. Annals of Applied Probability, 13(2):540–560, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1050689593.

Mendoza-Arriaga:2014:TCC

[MAL14] Rafael Mendoza-Arriaga and Vadim Linetsky. Time-changed CIR default intensities with two-sided mean-reverting jumps. Annals of Applied Probability, 24(2):811–856, April 2014. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1394465372.

Mannion:1993:PRM

[Man93] David Mannion. Products of 2 × 2 random matrices. Annals of Applied Probability, 3(4):1189–1218, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005279.

Maansson:1999:PAC

[Mån99] Marianne Månsson. Poisson approximation in connection with clustering of random points. Annals of Applied Probability, 9(2): 465–492, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1029962751.

Marquez:1997:CRA

[Már97] David Márquez. Convergence rates for annealing diffusion processes. Annals of Applied Probability, 7(4):1118–1139, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1043862427.

Martin:2001:PPF

[Mar01] James B. Martin. Point processes in fast Jackson networks. Annals of Applied Probability, 11(3):650–663, August 2001. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1015345344.

Marchand:2002:SIT

[Mar02] R. Marchand. Strict inequalities for the time constant in first passage percolation. Annals of Applied Probability, 12(3):1001– 1038, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1031863179.

Marckert:2008:LPG

[Mar08] Jean-François Marckert. The lineage process in Galton–Watson trees and globally centered discrete snakes. Annals of Applied Probability, 18(1):209–244, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1199890021.

Martinsson:2016:UFP

[Mar16] Anders Martinsson. Unoriented first-passage percolation on the n-cube. Annals of Applied Probability, 26(5):2597–2625, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1476884298.

Marshall:2019:AMH

[Mar19] Nicholas F. Marshall. Approximating mixed Hölder functions using random samples. Annals of Applied Probability, 29(5): 2988–3005, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1571385627.

Mase:1995:CMP

[Mas95] Shigeru Mase. Consistency of the maximum pseudo-likelihood estimator of continuous state space Gibbsian processes. Annals of Applied Probability, 5(3):603–612, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004697.

Massoulie:2007:SPP

[Mas07] Laurent Massoulié. Structural properties of proportional fairness: Stability and insensitivity. Annals of Applied Probability, 17(3):809–839, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1179839175.

Matzinger:2005:RTC

[Mat05] Heinrich Matzinger. Reconstructing a two-color scenery by observing it along a simple random walk path. Annals of Applied Probability, 15(1B):778–819, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1107271668.

Mayberry:2009:GPC

[May09] John Mayberry. Gaussian perturbations of circle maps: a spectral approach. Annals of Applied Probability, 19(3):1143–1171, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1245071022.

McDiarmid:1995:MPB

[McD95] Colin McDiarmid. Minimal positions in a branching random walk. Annals of Applied Probability, 5(1):128–139, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004832.

McDonald:1999:AFP

[McD99] D. R. McDonald. Asymptotics of first passage times for random walk in an orthant. Annals of Applied Probability, 9(1):110–145, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962599.

Meyn:1994:SGJ

[MD94] S. P. Meyn and D. Down. Stability of generalized Jackson networks. Annals of Applied Probability, 4(1):124–148, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005203.

Miclo:2001:GIP

[MD01] L. Miclo and P. Del Moral. Genealogies and increasing propagation of chaos for Feynman–Kac and genetic models. Annals of Applied Probability, 11(4):1166–1198, November 2001. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1015345399.

Ma:2018:SEG

[MD18] Ruibo Ma and Rick Durrett. A simple evolutionary game arising from the study of the role of IGF-II in pancreatic cancer. Annals of Applied Probability, 28(5):2896–2921, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1535443237.

Meilijson:2009:ACD

[Mei09] Isaac Meilijson. On the adjustment coefficient, drawdowns and lundberg-type bounds for random walk. Annals of Applied Probability, 19(3):1015–1025, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1245071017.

Meka:2015:PTA

[Mek15] Raghu Meka. A polynomial time approximation scheme for computing the supremum of Gaussian processes. Annals of Applied Probability, 25(2):465–476, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1424355120.

Meleard:2000:TPV

[Mél00] Sylvie Méléard. A trajectorial proof of the vortex method for the two-dimensional Navier–Stokes equation. Annals of Applied Probability, 10(4):1197–1211, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487613.

Merlet:2007:CLT

[Mer07] Glenn Merlet. A central limit theorem for stochastic recursive sequences of topical operators. Annals of Applied Probability, 17(4):1347–1361, August 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1186755242.

Meyn:1995:TMQ

[Mey95] Sean P. Meyn. Transience of multiclass queueing networks via fluid limit models. Annals of Applied Probability, 5(4):946–957, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004601.

Meyn:2006:LDA

[Mey06] Sean P. Meyn. Large deviation asymptotics and control variates for simulating large functions. Annals of Applied Probability, 16(1):310–339, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1141654289.

Muller-Gronbach:2002:OUA

[MG02] Thomas Müller-Gronbach. The optimal uniform approximation of systems of stochastic differential equations. Annals of Applied Probability, 12(2):664–690, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1026915620.

Muller-Gronbach:2004:OPA

[MG04] Thomas Müller-Gronbach. Optimal pointwise approximation of SDEs based on Brownian motion at discrete points. Annals of Applied Probability, 14(4):1605–1642, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1099674072.

Mountford:2005:MSC

[MG05] Thomas Mountford and Hervé Guiol. The motion of a second class particle for the TASEP starting from a decreasing shock profile. Annals of Applied Probability, 15(2):1227–1259, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1115137974.

Muller-Gronbach:2023:SLE

[MGY23] Thomas Müller-Gronbach and Larisa Yaroslavtseva. Sharp lower error bounds for strong approximation of SDEs with discontinuous drift coefficient by coupling of noise. Annals of Applied Probability, 33(2):902–935, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-2/Sharp-lower-error-bounds-for-strong-
approximation-of-SDEs-with/10.1214/22-AAP1837.full.

McVinish:2021:FAS

[MH21] Ross McVinish and Liam Hodgkinson. Fast approximate simulation of finite long-range spin systems. Annals of Applied Probability, 31(3):1443–1473, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-3/Fast-approximate-simulation-of-finite-
long-range-spin-systems/10.1214/20-AAP1624.full.

Miclo:1996:PSD

[Mic96] Laurent Miclo. Sur les problèmes de sortie discrets inhomogènes. Annals of Applied Probability, 6(4):1112–1156, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1035463326.

Miclo:2002:ART

[Mic02] L. Miclo. About relaxation time of finite generalized Metropolis algorithms. Annals of Applied Probability, 12(4):1492–1515, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1037125871.

Miltersen:1994:ATT

[Mil94] Kristian R. Miltersen. An arbitrage theory of the term structure of interest rates. Annals of Applied Probability, 4(4):953–967, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004898.

Miyazawa:2004:HPM

[Miy04] Masakiyo Miyazawa. Hitting probabilities in a Markov additive process with linear movements and upward jumps: Applications to risk and queueing processes. Annals of Applied Probability, 14(2):1029–1054, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1082737121.

McKeague:2016:CED

[ML16] Ian W. McKeague and Bruce Levin. Convergence of empirical distributions in an interpretation of quantum mechanics. Annals of Applied Probability, 26(4):2540–2555, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1472745466.

Morina:2022:BFD

[MLNW22] Giulio Morina, Krzysztof Latuszy’nski, Piotr Nayar, and Alex Wendland. From the Bernoulli factory to a dice enterprise via perfect sampling of Markov chains. Annals of Applied Probability, 32(1):327–359, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-1/From-the-Bernoulli-factory-to-a-dice-
enterprise-via-perfect/10.1214/21-AAP1679.full.

Millet:2001:NSW

[MM01] Annie Millet and Pierre-Luc Morien. On a nonlinear stochastic wave equation in the plane: existence and uniqueness of the solution. Annals of Applied Probability, 11(3):922–951, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1015345353.

MacPhee:2003:CRW

[MM03] I. M. MacPhee and M. V. Menshikov. Critical random walks on two-dimensional complexes with applications to polling systems. Annals of Applied Probability, 13(4):1399–1422, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1069786503.

Mairesse:2007:RGB

[MM07] Jean Mairesse and Frédéric Mathéus. Randomly growing braid on three strands and the manta ray. Annals of Applied Probability, 17(2):502–536, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1174323255.

Mertikopoulos:2010:ERB

[MM10] Panayotis Mertikopoulos and Aris L. Moustakas. The emergence of rational behavior in the presence of stochastic perturbations. Annals of Applied Probability, 20(4):1359–1388, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1279638789.

McGoff:2022:GPC

[MMN22] Kevin McGoff, Sayan Mukherjee, and Andrew B. Nobel. Gibbs posterior convergence and the thermodynamic formalism. Annals of Applied Probability, 32(1):461–496, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Gibbs-
posterior-convergence-and-the-thermodynamic-formalism/
10.1214/21-AAP1685.full.

MacPhee:2007:MCM

[MMPP07] Iain MacPhee, Mikhail Menshikov, Dimitri Petritis, and Serguei Popov. A Markov chain model of a polling system with parameter regeneration. Annals of Applied Probability, 17(5/6): 1447–1473, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1191419172.

MacPhee:2008:PSP

[MMPP08] Iain MacPhee, Mikhail Menshikov, Dimitri Petritis, and Serguei Popov. Polling systems with parameter regeneration, the general case. Annals of Applied Probability, 18(6):2131–2155, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1227708914.

MacPhee:2006:PTR

[MMPV06] I. M. MacPhee, M. V. Menshikov, S. Popov, and S. Volkov. Periodicity in the transient regime of exhaustive polling systems.

Annals of Applied Probability, 16(4):1816–1850, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1169065209.

Majka:2020:NBS

[MMS20] Mateusz B. Majka, Aleksandar Mijatović, and Lukasz Szpruch. Nonasymptotic bounds for sampling algorithms without logconcavity. Annals of Applied Probability, 30(4):1534–1581, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/Nonasymptotic-
bounds-for-sampling-algorithms-without-log-concavity/
10.1214/19-AAP1535.full.

Moharrami:2021:PMP

[MMX21] Mehrdad Moharrami, Cristopher Moore, and Jiaming Xu. The planted matching problem: Phase transitions and exact results. Annals of Applied Probability, 31(6):2663–2720, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-6/The-
planted-matching-problem-Phase-transitions-and-exact-
results/10.1214/20-AAP1660.full.

Mathew:1997:IPC

[MN97] Thomas Mathew and Kenneth Nordström. Inequalities for the probability content of a rotated ellipse and related stochastic domination results. Annals of Applied Probability, 7(4): 1106–1117, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1043862426.

Mahmoud:2003:DDR

[MN03] Hosam M. Mahmoud and Ralph Neininger. Distribution of distances in random binary search trees. Annals of Applied Probability, 13(1):253–276, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1042765668.

Mourrat:2017:SLC

[MN17] Jean-Christophe Mourrat and James Nolen. Scaling limit of the corrector in stochastic homogenization. Annals of Applied Probability, 27(2):944–959, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1495764371.

Mueller:2022:SPM

[MN22] Carl Mueller and Eyal Neuman. Scaling properties of a moving polymer. Annals of Applied Probability, 32(6):4251–4278, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/Scaling-
properties-of-a-moving-polymer/10.1214/22-AAP1785.full.

Mandjes:2009:CSF

[MNG09] Michel Mandjes, Ilkka Norros, and Peter Glynn. On convergence to stationarity of fractional Brownian storage. Annals of Applied Probability, 19(4):1385–1403, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1248700622.

Morris:2014:MTC

[MNP14] Ben Morris, Weiyang Ning, and Yuval Peres. Mixing time of the card-cyclic-to-random shuffle. Annals of Applied Probability, 24 (5):1835–1849, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1403812363.

Mossel:2016:BPR

[MNS16] Elchanan Mossel, Joe Neeman, and Allan Sly. Belief propagation, robust reconstruction and optimal recovery of block models. Annals of Applied Probability, 26(4):2211–2256, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1472745457.

Monoyios:2022:DOC

[Mon22] Michael Monoyios. Duality for optimal consumption under no unbounded profit with bounded risk. Annals of Applied Probability, 32(5):3572–3613, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-32/issue-5/Duality-for-optimal-consumption-under-
no-unbounded-profit-with-bounded/10.1214/21-AAP1767.
full.

Morrow:1992:LDR

[Mor92] Gregory J. Morrow. Large deviation results for a class of Markov chains with application to an infinite alleles model of population genetics. Annals of Applied Probability, 2(4):857–905, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005579.

Mora:2005:NSC

[Mor05] Carlos M. Mora. Numerical solution of conservative finitedimensional stochastic Schrödinger equations. Annals of Applied Probability, 15(3):2144–2171, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1121433780.

Morris:2006:MTS

[Mor06] Ben Morris. The mixing time for simple exclusion. Annals of Applied Probability, 16(2):615–635, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1151592245.

Mossel:2001:RTB

[Mos01] Elchanan Mossel. Reconstruction on trees: Beating the second eigenvalue. Annals of Applied Probability, 11(1):285–300, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
998926994.

Mountford:2001:ERA

[Mou01] Tom Mountford. An extension of a result of andjel. Annals of Applied Probability, 11(2):405–418, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015345297.

Mourrat:2020:HJE

[Mou20] J.-C. Mourrat. Hamilton–Jacobi equations for finite-rank matrix inference. Annals of Applied Probability, 30(5):2234–2260, October 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-5/HamiltonJacobi-
equations-for-finite-rank-matrix-inference/10.1214/19-
AAP1556.full.

Moyal:2015:GBS

[Moy15] P. Moyal. A generalized backward scheme for solving nonmonotonic stochastic recursions. Annals of Applied Probability, 25(2):582–599, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355124.

Mountford:1995:WCD

[MP95] T. Mountford and B. Prabhakar. On the weak convergence of departures from an infinite series of ·/M/1 queues. Annals of Applied Probability, 5(1):121–127, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004831.

Mandelbaum:1998:SDS

[MP98] Avi Mandelbaum and Gennady Pats. State-dependent stochastic networks. Part I. Approximations and applications with continuous diffusion limits. Annals of Applied Probability, 8(2):569–646, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903539.

Madras:1999:ISF

[MP99] Neal Madras and Mauro Piccioni. Importance sampling for families of distributions. Annals of Applied Probability, 9(4): 1202–1225, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962870.

Mazza:2001:DFP

[MP01] Christian Mazza and Didier Piau. Dubins-freedman processes and RC filters. Annals of Applied Probability, 11(4):1330–1352, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1015345405.

Mossel:2003:IFT

[MP03] Elchanan Mossel and Yuval Peres. Information flow on trees. Annals of Applied Probability, 13(3):817–844, August 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1060202828.

Mokkadem:2006:CRA

[MP06] Abdelkader Mokkadem and Mariane Pelletier. Convergence rate and averaging of nonlinear two-time-scale stochastic approximation algorithms. Annals of Applied Probability, 16(3):1671–1702, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804996.

Mason:2009:ANP

[MP09] David M. Mason and Wolfgang Polonik. Asymptotic normality of plug-in level set estimates. Annals of Applied Probability, 19(3):1108–1142, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1245071021.

McSweeney:2014:SIB

[MP14] John K. McSweeney and Lea Popovic. Stochastically-induced bistability in chemical reaction systems. Annals of Applied Probability, 24(3):1226–1268, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1398258100.

Moyal:2017:IMQ

[MP17] Pascal Moyal and Ohad Perry. On the instability of matching queues. Annals of Applied Probability, 27(6):3385–3434, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1513328704.

Muller:2020:OCC

[MP20] Tobias Müller and Mathew D. Penrose. Optimal Cheeger cuts and bisections of random geometric graphs. Annals of Applied Probability, 30(3):1458–1483, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-30/issue-3/Optimal-Cheeger-cuts-and-bisections-
of-random-geometric-graphs/10.1214/19-AAP1534.full.

Mitsche:2021:LTC

[MP21] Dieter Mitsche and Mathew D. Penrose. Limit theory of combinatorial optimization for random geometric graphs. Annals of Applied Probability, 31(6):2721–2771, December 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-6/Limit-
theory-of-combinatorial-optimization-for-random-geometric-
graphs/10.1214/20-AAP1661.full.

Maillard:2022:IFC

[MP22a] Pascal Maillard and Elliot Paquette. Interval fragmentations with choice: Equidistribution and the evolution of tagged fragments. Annals of Applied Probability, 32(5):3537–3571, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-5/Interval-
fragmentations-with-choice--Equidistribution-and-the-
evolution-of/10.1214/21-AAP1766.full.

Motte:2022:MFM

[MP22b] Médéric Motte and Huyên Pham. Mean-field Markov decision processes with common noise and open-loop controls. Annals of Applied Probability, 32(2):1421–1458, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Mean-
field-Markov-decision-processes-with-common-noise-and-
open/10.1214/21-AAP1713.full.

Miller:2023:LSB

[MP23] Luke Miller and Helmut H. Pitters. Large-scale behaviour and hydrodynamic limit of beta coalescents. Annals of Applied Probability, 33(1):1–27, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-1/Large-scale-behaviour-and-hydrodynamic-
limit-of-beta-coalescents/10.1214/22-AAP1782.full.

Muirhead:2019:BAM

[MPdS19] S. Muirhead, R. Pymar, and R. S. dos Santos. The Bouchaud– Anderson model with double-exponential potential. Annals of Applied Probability, 29(1):264–325, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1544000430.

Matzinger:2017:RMS

[MPP17] Heinrich Matzinger, Angelica Pachon, and Serguei Popov. Reconstruction of a multidimensional scenery with a branching random walk. Annals of Applied Probability, 27(2):651–685, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1495764363.

Mattingly:2012:DLR

[MPS12] Jonathan C. Mattingly, Natesh S. Pillai, and Andrew M. Stuart. Diffusion limits of the random walk Metropolis algorithm in high dimensions. Annals of Applied Probability, 22(3):881–930, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1337347534.

Ma:2002:NMB

[MPST02] Jin Ma, Philip Protter, Jaime San Martín, and Soledad Torres. Numberical method for backward stochastic differential equations. Annals of Applied Probability, 12(1):302–316, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015961165.

Matoussi:2013:SOR

[MPZ13] Anis Matoussi, Dylan Possamai, and Chao Zhou. Second order reflected backward stochastic differential equations. Annals of Applied Probability, 23(6):2420–2457, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1382447693.

Matoussi:2021:CSS

[MPZ21] Anis Matoussi, Dylan Possamaï, and Chao Zhou. Corrigendum for “Second-order reflected backward stochastic differential equations” and “Second-order BSDEs with general reflection and game options under uncertainty”. Annals of Applied Probability, 31(3):1505–1522, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-31/
issue-3/Corrigendum-for-Second-order-reflected-backward-
stochastic-differential-equations-and/10.1214/20-AAP1622.
full.

Matthews:1993:ADN

[MR93] Peter C. Matthews and Andrew L. Rukhin. Asymptotic distribution of the normal sample range. Annals of Applied Probability, 3(2):454–466, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005433.

Malyshev:1994:PTL

[MR94a] Vadim Malyshev and Philippe Robert. Phase transition in a load sharing loss model. Annals of Applied Probability, 4(4): 1161–1176, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004909.

Meester:1994:UUO

[MR94b] Ronald Meester and Rahul Roy. Uniqueness of unbounded occupied and vacant components in Boolean models. Annals of Applied Probability, 4(3):933–951, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004978.

Meziares:2000:SBI

[MR00] Sophie Méziàres and Bernard Roynette. Study of a Brownian impulse. Annals of Applied Probability, 10(2):493–516, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1019487352.

Madras:2002:MCD

[MR02] Neal Madras and Dana Randall. Markov chain decomposition for convergence rate analysis. Annals of Applied Probability, 12(2): 581–606, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1026915617.

Mohamed:2005:PAS

[MR05] Hanène Mohamed and Philippe Robert. A probabilistic analysis of some tree algorithms. Annals of Applied Probability, 15(4): 2445–2471, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965768.

McDonald:2006:MFC

[MR06a] D. R. McDonald and J. Reynier. Mean field convergence of a model of multiple TCP connections through a buffer implementing RED. Annals of Applied Probability, 16(1):244–294, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1141654287.

Mossel:2006:LNP

[MR06b] Elchanan Mossel and Sébastien Roch. Learning nonsingular phylogenies and hidden Markov models. Annals of Applied Probability, 16(2):583–614, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1151592244.

Mora:2008:BPN

[MR08] Carlos M. Mora and Rolando Rebolledo. Basic properties of nonlinear stochastic Schrödinger equations driven by Brownian motions. Annals of Applied Probability, 18(2):591–619, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1206018198.

Mohamed:2010:DTA

[MR10] Hanène Mohamed and Philippe Robert. Dynamic tree algorithms. Annals of Applied Probability, 20(1):26–51, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1262962317.

Mossel:2012:PMC

[MR12] Elchanan Mossel and Sebastien Roch. Phylogenetic mixtures: Concentration of measure in the large-tree limit. Annals of Applied Probability, 22(6):2429–2459, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1353695958.

Mossel:2017:DBS

[MR17] Elchanan Mossel and Sebastien Roch. Distance-based species tree estimation under the coalescent: Information-theoretic trade-off between number of loci and sequence length. Annals of Applied Probability, 27(5):2926–2955, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1509696038.

Mikosch:2002:NTA

[MRRS02] Thomas Mikosch, Sidney Resnick, Holger Rootzén, and Alwin Stegeman. Is network traffic appriximated by stable Lévy motion or fractional Brownian motion? Annals of Applied Probability, 12(1):23–68, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015961155.

Mansfield:2001:LSS

[MRS01] Peter Mansfield, Svetlozar T. Rachev, and Gennady Samorodnitsky. Long strange segments of a stochastic process. Annals of Applied Probability, 11(3):878–921, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015345352.

Madaule:2016:GPF

[MRV16] Thomas Madaule, Rémi Rhodes, and Vincent Vargas. Glassy phase and freezing of log-correlated Gaussian potentials. Annals of Applied Probability, 26(2):643–690, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1458651816.

Marinucci:2021:NUF

[MRV21] Domenico Marinucci, Maurizia Rossi, and Anna Vidotto. Nonuniversal fluctuations of the empirical measure for isotropic stationary fields on S^{2} × R. Annals of Applied Probability, 31(5): 2311–2349, October 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
5/Non-universal-fluctuations-of-the-empirical-measure-
for-isotropic-stationary/10.1214/20-AAP1648.full.

Mendelson:2018:IBS

[MRW18] Shahar Mendelson, Holger Rauhut, and Rachel Ward. Improved bounds for sparse recovery from subsampled random convolutions. Annals of Applied Probability, 28(6):3491–3527, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1538985627.

Mahmoud:1991:DLR

[MS91] Hosam M. Mahmoud and R. T. Smythe. On the distribution of leaves in rooted subtrees of recursive trees. Annals of Applied Probability, 1(3):406–418, August 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005874.

Metry:1993:OPD

[MS93a] Magdy H. Metry and Allan R. Sampson. Orderings for positive dependence on multivariate empirical distributions. Annals of Applied Probability, 3(4):1241–1251, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005281.

Miyazawa:1993:LHD

[MS93b] Masakiyo Miyazawa and Volker Schmidt. On ladder height distributions of general risk processes. Annals of Applied Probability, 3(3):763–776, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005362.

Martin:1999:FJN

[MS99] J. B. Martin and Yu. M. Suhov. Fast Jackson networks. Annals of Applied Probability, 9(3):854–870, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962816.

Mikosch:2000:SND

[MS00] Thomas Mikosch and Gennady Samorodnitsky. The supremum of a negative drift random walk with dependent heavytailed steps. Annals of Applied Probability, 10(3):1025–1064, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1019487517.

Mania:2005:DEU

[MS05] Michael Mania and Martin Schweizer. Dynamic exponential utility indifference valuation. Annals of Applied Probability, 15 (3):2113–2143, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433779.

Marty:2011:GFW

[MS11] Renaud Marty and Knut Sølna. A general framework for waves in random media with long-range correlations. Annals of Applied Probability, 21(1):115–139, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1292598029.

Martinelli:2012:MTS

[MS12] Fabio Martinelli and Alistair Sinclair. Mixing time for the solidon-solid model. Annals of Applied Probability, 22(3):1136–1166, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1337347541.

Muller:2015:GAP

[MS15] Tobias Müller and Reto Spöhel. A geometric Achlioptas process. Annals of Applied Probability, 25(6):3295–3337, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1443703775.

Mangoubi:2018:RMG

[MS18] Oren Mangoubi and Aaron Smith. Rapid mixing of geodesic walks on manifolds with positive curvature. Annals of Applied Probability, 28(4):2501–2543, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1533780279.

Mostovyi:2020:OIC

[MS20] Oleksii Mostovyi and Mihai Sîrbu. Optimal investment and consumption with labor income in incomplete markets. Annals of Applied Probability, 30(2):747–787, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1591603221.

Mangoubi:2021:MHM

[MS21] Oren Mangoubi and Aaron Smith. Mixing of Hamiltonian Monte Carlo on strongly log-concave distributions: Continuous dynamics. Annals of Applied Probability, 31(5):2019–2045, October 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-5/Mixing-
of-Hamiltonian-Monte-Carlo-on-strongly-log-concave-distributions/
10.1214/20-AAP1640.full.

Mehri:2020:PCS

[MSSZ20] Sima Mehri, Michael Scheutzow, Wilhelm Stannat, and Bian Z. Zangeneh. Propagation of chaos for stochastic spatially structured neuronal networks with delay driven by jump diffusions. Annals of Applied Probability, 30(1):175–207, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1582621222.

Molchanov:1997:LSS

[MSW97] S. A. Molchanov, D. Surgailis, and W. A. Woyczynski. The largescale structure of the universe and quasi-Voronoi tessellation of shock fronts in forced Burgers turbulence in R^{d}. Annals of Applied Probability, 7(1):200–228, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034625260.

Monter:2019:DOR

[MSZ19] Sergio A. Almada Monter, Mykhaylo Shkolnikov, and Jiacheng Zhang. Dynamics of observables in rank-based models and performance of functionally generated portfolios. Annals of Applied Probability, 29(5):2849–2883, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1571385624.

Meyn:1994:CBG

[MT94a] Sean P. Meyn and R. L. Tweedie. Computable bounds for geometric convergence rates of Markov chains. Annals of Applied Probability, 4(4):981–1011, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004900.

Meyn:1994:SDC

[MT94b] Sean P. Meyn and R. L. Tweedie. State-dependent criteria for convergence of Markov chains. Annals of Applied Probability, 4(1):149–168, February 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005204.

Milstein:1999:SST

[MT99] G. N. Milstein and M. V. Tretyakov. Simulation of a space-time bounded diffusion. Annals of Applied Probability, 9(3):732–779, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962812.

Martinelli:2013:KCS

[MT13] F. Martinelli and C. Toninelli. Kinetically constrained spin models on trees. Annals of Applied Probability, 23(5):1967–1987, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1377696303.

Mukherjee:2022:CLT

[Muk22] Chiranjib Mukherjee. Central limit theorem for Gibbs measures on path spaces including long range and singular interactions and homogenization of the stochastic heat equation. Annals of Applied Probability, 32(3):2028–2062, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Central-
limit-theorem-for-Gibbs-measures-on-path-spaces-including/
10.1214/21-AAP1727.full.

Muller:2015:IGP

[Mül15] Sebastian Müller. Interacting growth processes and invariant percolation. Annals of Applied Probability, 25(1):268–286, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1418740186.

Muller:2018:SST

[Mül18] Marvin S. Müller. A stochastic Stefan-type problem under firstorder boundary conditions. Annals of Applied Probability, 28 (4):2335–2369, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1533780275.

Mossel:2006:LMC

[MV06] Elchanan Mossel and Eric Vigoda. Limitations of Markov chain Monte Carlo algorithms for Bayesian inference of phylogeny. Annals of Applied Probability, 16(4):2215–2234, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1169065222.

Marinucci:2016:HFA

[MV16] Domenico Marinucci and Sreekar Vadlamani. High-frequency asymptotics for Lipschitz-Killing curvatures of excursion sets on the sphere. Annals of Applied Probability, 26(1):462–506, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1452003245.

Mourrat:2018:SGR

[MV18] Jean-Christophe Mourrat and Daniel Valesin. Spatial Gibbs random graphs. Annals of Applied Probability, 28(2):751–789, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1523433624.

Mailler:2020:SAN

[MV20] Cécile Mailler and Denis Villemonais. Stochastic approximation on noncompact measure spaces and application to measurevalued Pólya processes. Annals of Applied Probability, 30(5): 2393–2438, October 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
5/Stochastic-approximation-on-noncompact-measure-spaces-
and-application-to-measure/10.1214/20-AAP1561.full.

Mandjes:2005:SPL

[MvU05] Michel Mandjes and Miranda van Uitert. Sample-path large deviations for tandem and priority queues with Gaussian inputs. Annals of Applied Probability, 15(2):1193–1226, May 2005. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1115137973.

Massey:1994:AMO

[MW94] William A. Massey and Ward Whitt. An analysis of the modified offered-load approximation for the nonstationary Erlang loss model. Annals of Applied Probability, 4(4):1145–1160, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004908.

Massey:1998:UAE

[MW98] William A. Massey and Ward Whitt. Uniform acceleration expansions for Markov chains with time-varying rates. Annals of Applied Probability, 8(4):1130–1155, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1028903375.

Mischaikow:2007:PVH

[MW07] Konstantin Mischaikow and Thomas Wanner. Probabilistic validation of homology computations for nodal domains. Annals of Applied Probability, 17(3):980–1018, June 2007. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1179839180.

Morvai:2008:UEB

[MW08] Gusztáv Morvai and Benjamin Weiss. On universal estimates for binary renewal processes. Annals of Applied Probability, 18 (5):1970–1992, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1225372958.

Mischaikow:2010:TGS

[MW10] Konstantin Mischaikow and Thomas Wanner. Topology-guided sampling of nonhomogeneous random processes. Annals of Applied Probability, 20(3):1068–1097, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1276867307.

Musacchio:2013:FLA

[MW13] John Musacchio and Jean Walrand. Fluid limits to analyze longterm flow rates of a stochastic network with ingress discarding. Annals of Applied Probability, 23(4):1318–1354, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1371834030.

Mayer-Wolf:2007:SRB

[MWZ07] Eddy Mayer-Wolf and Moshe Zakai. Some relations between mutual information and estimation error in Wiener space. Annals of Applied Probability, 17(3):1102–1116, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1179839183.

Ma:2015:WPF

[MWZZ15] Jin Ma, Zhen Wu, Detao Zhang, and Jianfeng Zhang. On wellposedness of forward-backward SDEs — a unified approach. Annals of Applied Probability, 25(4):2168–2214, August 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1432212440.

Malyshev:1996:CLC

[MY96] Vadim A. Malyshev and Andrei V. Yakovlev. Condensation in large closed Jackson networks. Annals of Applied Probability, 6(1):92–115, February 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968067.

Mykland:2000:CDH

[Myk00] Per Aslak Mykland. Conservative delta hedging. Annals of Applied Probability, 10(2):664–683, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487360.

Myneni:1992:PAO

[Myn92] Ravi Myneni. The pricing of the American option. Annals of Applied Probability, 2(1):1–23, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005768.

Marsaglia:1991:NCR

[MZ91] George Marsaglia and Arif Zaman. A new class of random number generators. Annals of Applied Probability, 1(3):462–480, August 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1177005878. See popular description in [Pet91]. See remarks in [EH95, TLC93] about the extremely bad lattice structure in high dimensions of the generators proposed in this paper.

Ma:2002:RTB

[MZ02] Jin Ma and Jianfeng Zhang. Representation theorems for backward stochastic differential equations. Annals of Applied Probability, 12(4):1390–1418, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1037125868.

Morrow:2005:SPL

[MZ05] G. J. Morrow and Y. Zhang. The sizes of the pioneering, lowest crossing and pivotal sites in critical percolation on the triangular lattice. Annals of Applied Probability, 15(3):1832–1886, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1121433771.

Maillard:2014:PMA

[MZ14] Pascal Maillard and Ofer Zeitouni. Performance of the Metropolis algorithm on a disordered tree: The Einstein relation. Annals of Applied Probability, 24(5):2070–2090, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1403812370.

Nagai:2012:DRM

[Nag12] Hideo Nagai. Downside risk minimization via a large deviations approach. Annals of Applied Probability, 22(2):608–669, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1333372009.

Nakashima:2011:ASC

[Nak11] Makoto Nakashima. Almost sure central limit theorem for branching random walks in random environment. Annals of Applied Probability, 21(1):351–373, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1292598038.

Narayanan:2016:RIP

[Nar16] Hariharan Narayanan. Randomized interior point methods for sampling and optimization. Annals of Applied Probability, 26 (1):597–641, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1452003248.

Neal:2006:MRR

[Nea06] Peter Neal. Multitype randomized Reed–Frost epidemics and epidemics upon random graphs. Annals of Applied Probability, 16(3):1166–1189, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1159804979.

Nguyen:1993:PNP

[Ngu93] Vien Nguyen. Processing networks with parallel and sequential tasks: Heavy traffic analysis and Brownian limits. Annals of Applied Probability, 3(1):28–55, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005506.

Nguyen:1994:TDF

[Ngu94] Vien Nguyen. The trouble with diversity: Fork-join networks with heterogeneous customer population. Annals of Applied Probability, 4(1):1–25, February 1994. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005198.

Nguyen:2017:ALE

[Ngu17] Hoi H. Nguyen. Asymptotic Lyapunov exponents for large random matrices. Annals of Applied Probability, 27(6):3672–3705, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1513328711.

Niu:1997:EVT

[Niu97] Xu-Feng Niu. Extreme value theory for a class of nonstationary time series with applications. Annals of Applied Probability, 7(2): 508–522, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1034625342.

Nam:2019:CCA

[NN19] Danny Nam and Evita Nestoridi. Cutoff for the cyclic adjacent transposition shuffle. Annals of Applied Probability, 29(6): 3861–3892, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1578366328.

Norris:1999:SCE

[Nor99] James R. Norris. Smoluchowski’s coagulation equation: uniqueness, nonuniqueness and a hydrodynamic limit for the stochastic coalescent. Annals of Applied Probability, 9(1):78–109, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962598.

Norris:2016:CEK

[Nor16] James Norris. A consistency estimate for Kac’s model of elastic collisions in a dilute gas. Annals of Applied Probability, 26(2):1029–1081, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1458651827.

Niemiro:1995:TES

[NP95] Wojciech Niemiro and Piotr Pokarowski. Tail events of some nonhomogeneous Markov chains. Annals of Applied Probability, 5(1):261–293, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004840.

Neuhauser:1999:ESV

[NP99] Claudia Neuhauser and Stephen W. Pacala. An explicitly spatial version of the Lotka–Volterra model with interspecific competition. Annals of Applied Probability, 9(4):1226–1259, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962871.

Nuzman:2001:RKH

[NP01] Carl J. Nuzman and H. Vincent Poor. Reproducing kernel Hilbert space methods for wide-sense self-similar processes. Annals of Applied Probability, 11(4):1199–1219, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1015345400.

Nagai:2002:RSD

[NP02] Hideo Nagai and Shige Peng. Risk-sensitive dynamic portfolio optimization with partial information on infinite time horizon. Annals of Applied Probability, 12(1):173–195, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1015961160.

Nacu:2005:FSN

[NP05]  Şerban Nacu and Yuval Peres. Fast simulation of new coins

from old. Annals of Applied Probability, 15(1A):93–115, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1106922322.

Neininger:2001:LLP

[NR01] Ralph Neininger and Ludger Rüschendorf. Limit laws for partial match queries in quadtrees. Annals of Applied Probability, 11(2): 452–469, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1015345300.

Neininger:2004:GLT

[NR04] Ralph Neininger and Ludger Rüschendorf. A general limit theorem for recursive algorithms and combinatorial structures. Annals of Applied Probability, 14(1):378–418, February 2004. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1075828056.

Neal:2006:OSP

[NR06] Peter Neal and Gareth Roberts. Optimal scaling for partially updating MCMC algorithms. Annals of Applied Probability, 16(2): 475–515, May 2006. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1151592241.

Natarovskii:2021:QSG

[NRS21] Viacheslav Natarovskii, Daniel Rudolf, and Björn Sprungk. Quantitative spectral gap estimate and Wasserstein contraction of simple slice sampling. Annals of Applied Probability, 31(2): 806–825, April 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
2/Quantitative-spectral-gap-estimate-and-Wasserstein-
contraction-of-simple-slice/10.1214/20-AAP1605.full.

Neal:2012:OSR

[NRY12] Peter Neal, Gareth Roberts, and Wai Kong Yuen. Optimal scaling of random walk Metropolis algorithms with discontinuous target densities. Annals of Applied Probability, 22(5):1880–1927, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1350067989.

Nadtochiy:2019:PSS

[NS19] Sergey Nadtochiy and Mykhaylo Shkolnikov. Particle systems with singular interaction through hitting times: Application in systemic risk modeling. Annals of Applied Probability, 29(1):89– 129, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1544000426.

Noack:2022:CMF

[NS22] Christian Noack and Philippe Sosoe. Central moments of the free energy of the stationary O’Connell–Yor polymer. Annals of Applied Probability, 32(5):3205–3228, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-5/Central-

moments-of-the-free-energy-of-the-stationary-OConnellYor/
10.1214/21-AAP1744.full.

Nutz:2020:CMF

[NST20] Marcel Nutz, Jaime San Martin, and Xiaowei Tan. Convergence to the mean field game limit: a case study. Annals of Applied Probability, 30(1):259–286, February 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/euclid.aoap/1582621224.

Ngo:2020:SIE

[NT20] Hoang-Long Ngo and Dai Taguchi. Semi-implicit Euler– Maruyama approximation for noncolliding particle systems. Annals of Applied Probability, 30(2):673–705, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1591603219.

Nummelin:2000:ECP

[Num00] Esa Nummelin. On the existence and convergence of price equilibria for random economies. Annals of Applied Probability, 10 (1):268–282, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019737673.

Nutz:2012:BEP

[Nut12] Marcel Nutz. The Bellman equation for power utility maximization with semimartingales. Annals of Applied Probability, 22(1):363–406, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1328623703.

Nutz:2013:RE

[Nut13] Marcel Nutz. Random G-expectations. Annals of Applied Probability, 23(5):1755–1777, October 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1377696297.

Ney:2003:HML

[NV03] P. E. Ney and A. N. Vidyashankar. Harmonic moments and large deviation rates for supercritical branching processes. Annals of Applied Probability, 13(2):475–489, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1050689589.

Ney:2004:LLT

[NV04] Peter E. Ney and Anand N. Vidyashankar. Local limit theory and large deviations for supercritical branching processes. Annals of Applied Probability, 14(3):1135–1166, August 2004. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1089736280. See

correction [NV06].

Ney:2006:CAL

[NV06] P. E. Ney and Anand N. Vidyashankar. Corrections and acknowledgment for “Local limit theory and large deviations for supercritical branching processes”. Annals of Applied Probability, 16(4):2272, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1169065225. See [NV04].

Nutz:2022:DOT

[NW22] Marcel Nutz and Ruodu Wang. The directional optimal transport. Annals of Applied Probability, 32(2):1400–1420, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/The-
directional-optimal-transport/10.1214/21-AAP1712.full.

Najim:2016:GFL

[NY16] Jamal Najim and Jianfeng Yao. Gaussian fluctuations for linear spectral statistics of large random covariance matrices. Annals of Applied Probability, 26(3):1837–1887, June 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905021.

Nguyen:2023:SAD

[NY23] Nhu Nguyen and George Yin. Stochastic approximation with discontinuous dynamics, differential inclusions, and applications. Annals of Applied Probability, 33(1):780–823, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Stochastic-
approximation-with-discontinuous-dynamics-differential-
inclusions-and-applications/10.1214/22-AAP1829.full.

Nutz:2015:OSU

[NZ15] Marcel Nutz and Jianfeng Zhang. Optimal stopping under adverse nonlinear expectation and related games. Annals of Applied Probability, 25(5):2503–2534, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1438261047.

Nutz:2020:COS

[NZ20] Marcel Nutz and Yuchong Zhang. Conditional optimal stopping: a time-inconsistent optimization. Annals of Applied Probability, 30(4):1669–1692, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-4/Conditional-optimal-stopping-A-time-inconsistent-
optimization/10.1214/19-AAP1540.full.

OCinneide:1991:PTD

[O’C91] Colm Art O’Cinneide. Phase-type distributions and majorization. Annals of Applied Probability, 1(2):219–227, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005935.

Oreshkin:2011:AEP

[OC11] Boris N. Oreshkin and Mark J. Coates. Analysis of error propagation in particle filters with approximation. Annals of Applied Probability, 21(6):2343–2378, December 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1322057324.

Olvera-Cravioto:2021:PBU

[OC21] Mariana Olvera-Cravioto. PageRank’s behavior under degree correlations. Annals of Applied Probability, 31(3):1403–1442, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-3/PageRanks-
behavior-under-degree-correlations/10.1214/20-AAP1623.
full.

Olvera-Cravioto:2011:THT

[OCBG11] Mariana Olvera-Cravioto, Jose Blanchet, and Peter Glynn. On the transition from heavy traffic to heavy tails for the M/G/1 queue: The regularly varying case. Annals of Applied Probability, 21(2):645–668, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1300800984.

Ohashi:2009:FTS

[Oha09] Alberto Ohashi. Fractional term structure models: No-arbitrage and consistency. Annals of Applied Probability, 19(4):1553–1580, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1248700627.

Oliveira:2009:CEK

[Oli09] Roberto Imbuzeiro Oliveira. On the convergence to equilibrium of Kac’s random walk on matrices. Annals of Applied Probability, 19(3):1200–1231, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1245071024.

Olofsson:1996:BPL

[Olo96] Peter Olofsson. Branching processes with local dependencies. Annals of Applied Probability, 6(1):238–268, February 1996. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1034968073.

ONeill:1997:EMR

[O’N97] Philip O’Neill. An epidemic model with removal-dependent infection rate. Annals of Applied Probability, 7(1):90–109, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034625253.

Onatski:2008:TWL

[Ona08] Alexei Onatski. The Tracy–Widom limit for the largest eigenvalues of singular complex Wishart matrices. Annals of Applied Probability, 18(2):470–490, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1206018194.

OCinneide:2000:NRT

[OP00] C. A. O’Cinneide and A. V. Pokrovskii. Nonuniform random transformations. Annals of Applied Probability, 10(4):1151–1181, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1019487611.

Ortmann:2016:EFR

[OQR16] Janosch Ortmann, Jeremy Quastel, and Daniel Remenik. Exact formulas for random growth with half-flat initial data. Annals of Applied Probability, 26(1):507–548, February 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1452003246.

Olesker-Taylor:2022:MLN

[OT22] Sam Olesker-Taylor. Metastability in loss networks with dynamic alternative routing. Annals of Applied Probability, 32(2): 1362–1399, April 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
2/Metastability-in-loss-networks-with-dynamic-alternative-
routing/10.1214/21-AAP1711.full.

Olesker-Taylor:2023:CRD

[OT23] Sam Olesker-Taylor. Cutoff for rewiring dynamics on perfect matchings. Annals of Applied Probability, 33(1):641–676, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Cutoff-
for-rewiring-dynamics-on-perfect-matchings/10.1214/22-
AAP1825.full.

Ott:2013:OSP

[Ott13] Curdin Ott. Optimal stopping problems for the maximum process with upper and lower caps. Annals of Applied Probability, 23 (6):2327–2356, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1382447690.

Oshima:2012:NEM

[OTV12] Kojiro Oshima, Josef Teichmann, and Dejan Velušček. A new extrapolation method for weak approximation schemes with applications. Annals of Applied Probability, 22(3):1008–1045, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1337347537.

Ou:1992:PBS

[OW92] Jihong Ou and Lawrence M. Wein. Performance bounds for scheduling queueing networks. Annals of Applied Probability, 2(2):460–480, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005711.

Owada:2018:LTB

[Owa18] Takashi Owada. Limit theorems for Betti numbers of extreme sample clouds with application to persistence barcodes. Annals of Applied Probability, 28(5):2814–2854, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1535443235.

Owada:2022:CPD

[Owa22] Takashi Owada. Convergence of persistence diagram in the sparse regime. Annals of Applied Probability, 32(6):4706–4736, December 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
6/Convergence-of-persistence-diagram-in-the-sparse-regime/
10.1214/22-AAP1800.full.

Owen:2002:UBO

[Owe02] M. P. Owen. Utility based optimal hedging in incomplete markets. Annals of Applied Probability, 12(2):691–709, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1026915621.

Ossiander:1997:SWF

[OWZ97] Mina Ossiander, Ed Waymire, and Qing Zhang. Some width function asymptotics for weighted trees. Annals of Applied Probability, 7(4):972–995, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1043862421.

Pal:2011:AMW

[Pal11] Soumik Pal. Analysis of market weights under volatilitystabilized market models. Annals of Applied Probability, 21 (3):1180–1213, June 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1307020395.

Panloup:2008:RCI

[Pan08] Fabien Panloup. Recursive computation of the invariant measure of a stochastic differential equation driven by a Lévy process. Annals of Applied Probability, 18(2):379–426, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1206018192.

Papangelou:1998:EOF

[Pap98] F. Papangelou. Elliptic and other functions in the large deviations behavior of the Wright–Fisher process. Annals of Applied Probability, 8(1):182–192, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1027961039.

Papangelou:2000:LDM

[Pap00] F. Papangelou. The large deviations of a multi-allele Wright– Fisher process mapped on the sphere. Annals of Applied Probability, 10(4):1259–1273, November 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1019487616.

Paulsen:2002:CLA

[Pau02] Jostein Paulsen. On Cramér-like asymptotics for risk processes with stochastic return on investments. Annals of Applied Probability, 12(4):1247–1260, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1037125862.

Penda:2022:CLT

[PD22] S. Valère Bitseki Penda and Jean-François Delmas. Central limit theorem for bifurcating Markov chains under pointwise ergodic conditions. Annals of Applied Probability, 32(5):3817–3849, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-5/Central-
limit-theorem-for-bifurcating-Markov-chains-under-pointwise-
ergodic/10.1214/21-AAP1774.full.

Penda:2014:DIM

[PDG14] S. Valère Bitseki Penda, Hacène Djellout, and Arnaud Guillin. Deviation inequalities, moderate deviations and some limit theorems for bifurcating Markov chains with application. Annals of Applied Probability, 24(1):235–291, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278725.

Pelletier:1998:WCR

[Pel98] Mariane Pelletier. Weak convergence rates for stochastic approximation with application to multiple targets and simulated annealing. Annals of Applied Probability, 8(1):10–44, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1027961032.

Peled:2010:RIP

[Pel10] Ron Peled. On rough isometries of Poisson processes on the line. Annals of Applied Probability, 20(2):462–494, April 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1268143430.

Pemantle:2009:SCN

[Pem09] Robin Pemantle. Search cost for a nearly optimal path in a binary tree. Annals of Applied Probability, 19(4):1273–1291, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1248700617.

Penrose:1993:SLB

[Pen93] Mathew D. Penrose. On the spread-out limit for bond and continuum percolation. Annals of Applied Probability, 3(1):253–276, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005518.

Penrose:1996:CPE

[Pen96] Mathew D. Penrose. Continuum percolation and Euclidean minimal spanning trees in high dimensions. Annals of Applied Probability, 6(2):528–544, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968142.

Penrose:1997:LER

[Pen97] Mathew D. Penrose. The longest edge of the random minimal spanning tree. Annals of Applied Probability, 7(2):340–361, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034625335.

Penrose:2000:VOP

[Pen00] Mathew D. Penrose. Vertex ordering and partitioning problems for random spatial graphs. Annals of Applied Probability, 10(2): 517–538, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1019487353.

Pene:2005:RCM

[Pèn05] Françoise Pène. Rate of convergence in the multidimensional central limit theorem for stationary processes. application to the Knudsen gas and to the Sinai billiard. Annals of Applied Probability, 15(4):2331–2392, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1133965765.

Penrose:2016:CSR

[Pen16] Mathew D. Penrose. Connectivity of soft random geometric graphs. Annals of Applied Probability, 26(2):986–1028, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1458651826.

Perfekt:1994:EBS

[Per94] Roland Perfekt. Extremal behaviour of stationary Markov chains with applications. Annals of Applied Probability, 4(2):529–548, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005071.

Perrut:2000:HLT

[Per00] Anne Perrut. Hydrodynamic limits for a two-species reactiondiffusion process. Annals of Applied Probability, 10(1):163–191, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1019737668.

Pergamenshchikov:2003:LTL

[Per03] S. Pergamenshchikov. Limit theorem for Leland’s strategy. Annals of Applied Probability, 13(3):1099–1118, August 2003. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1060202836.

Peskir:2014:QDH

[Pes14] Goran Peskir. Quickest detection of a hidden target and extremal surfaces. Annals of Applied Probability, 24(6):2340–2370, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1409058034.

Peskir:2019:COS

[Pes19] Goran Peskir. Continuity of the optimal stopping boundary for two-dimensional diffusions. Annals of Applied Probability, 29 (1):505–530, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1544000435.

Peterson:1991:NRN

[Pet91] Ivars Peterson. Numbers at random: Number theory supplies a superior random-number generator. Science News (Washington, DC), 140(19):300–301, November 9, 1991. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http://

www.jstor.org/stable/3975915.

Pan:2007:ADS

[PGZ07] Guang-Ming Pan, Mei-Hui Guo, and Wang Zhou. Asymptotic distributions of the signal-to-interference ratios of LMMSE detection in multiuser communications. Annals of Applied Probability, 17(1):181–206, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1171377181.

Pham:2002:MSR

[Pha02] Huyên Pham. Minimizing shortfall risk and applications to finance and insurance problems. Annals of Applied Probability, 12(1):143–172, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015961159.

Piau:1999:LDY

[Pia99] Didier Piau. Large deviations and Young measures for a Poissonian model of biphased material. Annals of Applied Probability, 9(3):706–718, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962810.

Piau:2005:IPC

[Pia05] Didier Piau. Invariance principle for the coverage rate of genomic physical mappings. Annals of Applied Probability, 15(4): 2553–2574, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1133965772.

Piau:2006:HCT

[Pia06] Didier Piau. Harmonic continuous-time branching moments. Annals of Applied Probability, 16(4):2078–2097, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1169065217.

Pinsky:1992:ECL

[Pin92] Mark A. Pinsky. Extremal character of the Lyapunov exponent of the stochastic harmonic oscillator. Annals of Applied Probability, 2(4):942–950, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005582.

Pittel:1992:LSS

[Pit92] Boris Pittel. On likely solutions of a stable marriage problem. Annals of Applied Probability, 2(2):358–401, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1177005708.

Pittel:1999:NCP

[Pit99] Boris Pittel. Normal convergence problem? two moments and a recurrence may be the clues. Annals of Applied Probability, 9 (4):1260–1302, November 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962872.

Pittel:2008:LCR

[Pit08] B. G. Pittel. On the largest component of a random graph with a subpower-law degree sequence in a subcritical phase. Annals of Applied Probability, 18(4):1636–1650, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1216677135.

Popovic:2004:AGC

[Pop04] Lea Popovic. Asymptotic genealogy of a critical branching process. Annals of Applied Probability, 14(4):2120–2148, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1099674091.

Povel:1995:WCC

[Pov95] Tobias Povel. On weak convergence of conditional survival measure of one-dimensional Brownian motion with a drift. Annals of Applied Probability, 5(1):222–238, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004837.

Pacheco:1996:MSM

[PP96] António Pacheco and N. U. Prabhu. A Markovian storage model. Annals of Applied Probability, 6(1):76–91, February 1996. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1034968066.

Panholzer:2004:STS

[PP04] Alois Panholzer and Helmut Prodinger. Spanning tree size in random binary search trees. Annals of Applied Probability, 14(2): 718–733, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1082737108.

Pal:2008:ODB

[PP08a] Soumik Pal and Jim Pitman. One-dimensional Brownian particle systems with rank-dependent drifts. Annals of Applied Probability, 18(6):2179–2207, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1227708916.

Peccati:2008:LQF

[PP08b] Giovanni Peccati and Igor Prünster. Linear and quadratic functionals of random hazard rates: An asymptotic analysis. Annals of Applied Probability, 18(5):1910–1943, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1225372956.

Pages:2012:EAD

[PP12] Gilles Pagès and Fabien Panloup. Ergodic approximation of the distribution of a stationary diffusion: Rate of convergence. Annals of Applied Probability, 22(3):1059–1100, June 2012. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1337347539.

Pfaffelhuber:2015:SLS

[PP15] Peter Pfaffelhuber and Lea Popovic. Scaling limits of spatial compartment models for chemical reaction networks. Annals of Applied Probability, 25(6):3162–3208, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1443703772.

Pages:2018:WML

[PP18] Gilles Pagès and Fabien Panloup. Weighted multilevel Langevin simulation of invariant measures. Annals of Applied Probability, 28(6):3358–3417, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1538985625.

Pang:2022:FLT

[PP22] Guodong Pang and Étienne Pardoux. Functional limit theorems for non-Markovian epidemic models. Annals of Applied Probability, 32(3):1615–1665, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-3/Functional-limit-theorems-for-non-Markovian-
epidemic-models/10.1214/21-AAP1717.full.

Pages:2023:ULA

[PP23] Gilles Pagès and Fabien Panloup. Unadjusted Langevin algorithm with multiplicative noise: Total variation and Wasserstein bounds. Annals of Applied Probability, 33(1):726–779, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Unadjusted-
Langevin-algorithm-with-multiplicative-noise--Total-
variation-and/10.1214/22-AAP1828.full.

Pichler:2018:DFM

[PPM18] Georg Pichler, Pablo Piantanida, and Gerald Matz. Dictator functions maximize mutual information. Annals of Applied Probability, 28(5):3094–3101, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1535443243.

Piatnitski:2022:LDM

[PPZ22] Andrey Piatnitski, Sergey Pirogov, and Elena Zhizhina. Large deviations for Markov jump processes in periodic and locally periodic environments. Annals of Applied Probability, 32(6):4611– 4641, December 2022. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-32/issue-
6/Large-deviations-for-Markov-jump-processes-in-periodic-
and-locally/10.1214/22-AAP1797.full.

Pham:2001:OPP

[PQ01] Huyên Pham and Marie-Claire Quenez. Optimal portfolio in partially observed stochastic volatility models. Annals of Applied Probability, 11(1):210–238, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/998926991.

Pestien:1994:ABL

[PR94] Victor Pestien and S. Ramakrishnan. Asymptotic behavior of large discrete-time cyclic queueing networks. Annals of Applied Probability, 4(2):591–606, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005074.

Palmowski:1998:SAF

[PR98a] Zbigniew Palmowski and Tomasz Rolski. The superposition of alternating on-off flows and a fluid model. Annals of Applied Probability, 8(2):524–540, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1028903537.

Paulauskas:1998:CPI

[PR98b] Vygantas Paulauskas and Svetlozar T. Rachev. Cointegrated processes with infinite variance innovations. Annals of Applied Probability, 8(3):775–792, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1028903450.

Puhalskii:2010:MSQ

[PR10] Anatolii A. Puhalskii and Josh E. Reed. On many-server queues in heavy traffic. Annals of Applied Probability, 20(1):129–195, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1262962320.

Pekoz:2013:DAR

[PRR13] Erol A. Peköz, Adrian Röllin, and Nathan Ross. Degree asymptotics with rates for preferential attachment random graphs. Annals of Applied Probability, 23(3):1188–1218, June 2013. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1362684858.

Przykucki:2023:PI

[PRS23] Michał Przykucki, Alexander Roberts, and Alex Scott. Parking on the integers. Annals of Applied Probability, 33(2):876–901, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Parking-
on-the-integers/10.1214/22-AAP1836.full.

Peres:2020:HFI

[PRSS20] Yuval Peres, Miklós Z. Rácz, Allan Sly, and Izabella Stuhl. How fragile are information cascades? Annals of Applied Probability, 30(6):2796–2814, December 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-30/
issue-6/How-fragile-are-information-cascades/10.1214/
20-AAP1573.full.

Pitt:1995:EAN

[PRW95] Loren D. Pitt, Raina Robeva, and Dao Yi Wang. An error analysis for the numerical calculation of certain random integrals: Part 1. Annals of Applied Probability, 5(1):171–197, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004835.

Pakes:1997:EDR

[PS97] A. G. Pakes and V. T. Stefanov. Explicit distributional results in pattern formation. Annals of Applied Probability, 7(3):666–678, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034801248.

Penrose:2005:EAR

[PS05] Mathew D. Penrose and Aidan Sudbury. Exact and approximate results for deposition and annihilation processes on graphs. Annals of Applied Probability, 15(1B):853–889, February 2005. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1107271670.

Pal:2014:CMB

[PS14] Soumik Pal and Mykhaylo Shkolnikov. Concentration of measure for Brownian particle systems interacting through their ranks. Annals of Applied Probability, 24(4):1482–1508, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1400073655.

Piiroinen:2016:FKF

[PS16] Petteri Piiroinen and Martin Simon. From Feynman–Kac formulae to numerical stochastic homogenization in electrical impedance tomography. Annals of Applied Probability, 26(5): 3001–3043, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1476884310.

Pillai:2017:KWS

[PS17a] Natesh S. Pillai and Aaron Smith. Kac’s walk on n-sphere mixes in n log n steps. Annals of Applied Probability, 27(1):631–650, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL https://projecteuclid.org/euclid.
aoap/1488790837.

Pistorius:2017:DDM

[PS17b] Martijn Pistorius and Mitja Stadje. On dynamic deviation measures and continuous-time portfolio optimization. Annals of Applied Probability, 27(6):3342–3384, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1513328703.

Poisat:2020:LTS

[PS20] Julien Poisat and François Simenhaus. A limit theorem for the survival probability of a simple random walk among power-law renewal obstacles. Annals of Applied Probability, 30(5):2030– 2068, October 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
5/A-limit-theorem-for-the-survival-probability-of-a-
simple/10.1214/19-AAP1551.full.

Pain:2022:CTH

[PS22] Michel Pain and Delphin Sénizergues. Correction terms for the height of weighted recursive trees. Annals of Applied Probability, 32(4):3027–3059, August 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-4/Correction-terms-for-the-height-of-weighted-
recursive-trees/10.1214/21-AAP1756.full.

Pillai:2012:OSD

[PST12] Natesh S. Pillai, Andrew M. Stuart, and Alexandre H. Thiéry. Optimal scaling and diffusion limits for the Langevin algorithm in high dimensions. Annals of Applied Probability, 22(6): 2320–2356, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1353695955.

Pfaffelhuber:2012:MRC

[PSW12] P. Pfaffelhuber, P. R. Staab, and A. Wakolbinger. Muller’s ratchet with compensatory mutations. Annals of Applied Probability, 22(5):2108–2132, October 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1350067996.

Perl:2015:EWM

[PSY15] Idan Perl, Arnab Sen, and Ariel Yadin. Extinction window of mean field branching annihilating random walk. Annals of Applied Probability, 25(6):3139–3161, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1443703771.

Podolskij:2014:LTN

[PSZ14] Mark Podolskij, Christian Schmidt, and Johanna F. Ziegel. Limit theorems for nondegenerate U-statistics of continuous semimartingales. Annals of Applied Probability, 24(6):2491– 2526, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1409058038.

Possamai:2015:WAS

[PT15] Dylan Possamaï and Xiaolu Tan. Weak approximation of secondorder BSDEs. Annals of Applied Probability, 25(5):2535–2562, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1438261048.

Poplavskyi:2017:DLR

[PTZ17] Mihail Poplavskyi, Roger Tribe, and Oleg Zaboronski. On the distribution of the largest real eigenvalue for the real Ginibre ensemble. Annals of Applied Probability, 27(3):1395–1413, June 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1500451226.

Possamai:2020:ZSP

[PTZ20] Dylan Possamaï, Nizar Touzi, and Jianfeng Zhang. Zero-sum path-dependent stochastic differential games in weak formulation. Annals of Applied Probability, 30(3):1415–1457, June 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-3/Zero-
sum-path-dependent-stochastic-differential-games-in-
weak-formulation/10.1214/19-AAP1533.full.

Puha:2015:DLS

[Puh15] Amber L. Puha. Diffusion limits for shortest remaining processing time queues under nonstandard spatial scaling. Annals of Applied Probability, 25(6):3381–3404, December 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1443703777.

Pulido:2014:FTA

[Pul14] Sergio Pulido. The fundamental theorem of asset pricing, the hedging problem and maximal claims in financial markets with short sales prohibitions. Annals of Applied Probability, 24(1):54– 75, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1389278719.

Podolskij:2020:EEE

[PVY20] Mark Podolskij, Bezirgen Veliyev, and Nakahiro Yoshida. Edgeworth expansion for Euler approximation of continuous diffusion processes. Annals of Applied Probability, 30(4):1971–2003, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/Edgeworth-
expansion-for-Euler-approximation-of-continuous-diffusion-
processes/10.1214/19-AAP1549.full.

Perman:1996:DBA

[PW96] Mihael Perman and Jon A. Wellner. On the distribution of Brownian areas. Annals of Applied Probability, 6(4):1091–1111, November 1996. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1035463325.

Puhalskii:1997:FLD

[PW97] Anatolii A. Puhalskii and Ward Whitt. Functional large deviation principles for first-passage-time processes. Annals of Applied Probability, 7(2):362–381, May 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1034625336.

Puha:2004:ISR

[PW04] Amber L. Puha and Ruth J. Williams. Invariant states and rates of convergence for a critical fluid model of a processor sharing queue. Annals of Applied Probability, 14(2):517–554, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1082737103.

Pang:2010:CQI

[PW10] Guodong Pang and Ward Whitt. Continuity of a queueing integral representation in the M_{1} topology. Annals of Applied Probability, 20(1):214–237, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1262962322.

Peigne:2021:RTD

[PW21] Marc Peigné and Wolfgang Woess. Recurrence of twodimensional queueing processes, and random walk exit times from the quadrant. Annals of Applied Probability, 31(6):2519– 2537, December 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-

6/Recurrence-of-two-dimensional-queueing-processes-and-
random-walk-exit/10.1214/20-AAP1654.full.

Peltola:2023:CPM

[PW23] Eveliina Peltola and Hao Wu. Crossing probabilities of multiple Ising interfaces. Annals of Applied Probability, 33(4):3169–3206, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-4/Crossing-
probabilities-of-multiple-Ising-interfaces/10.1214/22-
AAP1888.full.

Pollak:1998:NRR

[PY98] Moshe Pollak and Benjamin Yakir. A new representation for a renewal-theoretic constant appearing in asymptotic approximations of large deviations. Annals of Applied Probability, 8(3):749– 774, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1028903449.

Penrose:2001:CLT

[PY01] Mathew D. Penrose and J. E. Yukich. Central limit theorems for some graphs in computational geometry. Annals of Applied Probability, 11(4):1005–1041, November 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1015345393.

Penrose:2002:LTR

[PY02] Mathew D. Penrose and J. E. Yukich. Limit theory for random sequential packing and deposition. Annals of Applied Probability, 12(1):272–301, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1015961164.

Penrose:2003:WLL

[PY03] Mathew D. Penrose and J. E. Yukich. Weak laws of large numbers in geometric probability. Annals of Applied Probability, 13(1):277–303, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1042765669.

Penrose:2013:LTP

[PY13] Mathew D. Penrose and J. E. Yukich. Limit theory for point processes in manifolds. Annals of Applied Probability, 23(6):

2161–2211, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1382447685.

Pillai:2014:UCM

[PY14] Natesh S. Pillai and Jun Yin. Universality of covariance matrices. Annals of Applied Probability, 24(3):935–1001, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1398258093.

Podolskij:2016:EEF

[PY16] Mark Podolskij and Nakahiro Yoshida. Edgeworth expansion for functionals of continuous diffusion processes. Annals of Applied Probability, 26(6):3415–3455, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1481792589.

Pitman:2018:OSB

[PY18] Jim Pitman and Yuri Yakubovich. Ordered and size-biased frequencies in GEM and Gibbs’ models for species sampling. Annals of Applied Probability, 28(3):1793–1820, June 2018. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1527840032.

Pan:2008:CLT

[PZ08] G. M. Pan and W. Zhou. Central limit theorem for signalto-interference ratio of reduced rank linear receiver. Annals of Applied Probability, 18(3):1232–1270, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1211819800.

Pan:2011:CLT

[PZ11] G. M. Pan and W. Zhou. Central limit theorem for Hotelling’s T^{2} statistic under large dimension. Annals of Applied Probability, 21(5):1860–1910, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576611.

Popier:2019:SOB

[PZ19] Alexandre Popier and Chao Zhou. Second-order BSDE under monotonicity condition and liquidation problem under uncertainty. Annals of Applied Probability, 29(3):1685–1739, June 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1550566840.

Qin:2021:LSS

[QH21] Qian Qin and James P. Hobert. On the limitations of singlestep drift and minorization in Markov chain convergence analysis. Annals of Applied Probability, 31(4):1633–1659, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-4/On-the-
limitations-of-single-step-drift-and-minorization-in/
10.1214/20-AAP1628.full.

Qin:2022:WBM

[QH22] Qian Qin and James P. Hobert. Wasserstein-based methods for convergence complexity analysis of MCMC with applications. Annals of Applied Probability, 32(1):124–166, February 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-1/Wasserstein-
based-methods-for-convergence-complexity-analysis-of-
MCMC-with/10.1214/21-AAP1673.full.

Quine:1994:GBG

[QS94] M. P. Quine and W. Szczotka. Generalisations of the bienayme-Galton–Watson branching process via its representation as an embedded random walk. Annals of Applied Probability, 4(4): 1206–1222, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004912.

Qiu:2023:TFE

[QS23a] Jiaze Qiu and Subhabrata Sen. The TAP free energy for highdimensional linear regression. Annals of Applied Probability, 33 (4):2643–2680, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-4/The-TAP-free-energy-for-high-dimensional-linear-
regression/10.1214/22-AAP1874.full.

Quattropani:2023:MAP

[QS23b] Matteo Quattropani and Federico Sau. Mixing of the averaging process and its discrete dual on finite-dimensional geometries. Annals of Applied Probability, 33(2):936–971, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Mixing-
of-the-averaging-process-and-its-discrete-dual-on/10.
1214/22-AAP1838.full.

Qian:2018:RBS

[QX18] Zhongmin Qian and Mingyu Xu. Reflected backward stochastic differential equations with resistance. Annals of Applied Probability, 28(2):888–911, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1523433627.

Qiu:2023:SDG

[QZ23] Jinniao Qiu and Jing Zhang. Stochastic differential games with random coefficients and stochastic Hamilton–Jacobi–Bellman– Isaacs equations. Annals of Applied Probability, 33(2):689–730, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-2/Stochastic-
differential-games-with-random-coefficients-and-stochastic-
HamiltonJacobiBellmanIsaacs-equations/10.1214/22-AAP1831.
full.

Reed:2009:GGQ

[Ree09] Josh Reed. The g / GI / n queue in the Halfin–Whitt regime. Annals of Applied Probability, 19(6):2211–2269, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1259158771.

Reinert:1995:AEG

[Rei95] Gesine Reinert. The asymptotic evolution of the general stochastic epidemic. Annals of Applied Probability, 5(4):1061–1086, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004606.

Remenik:2008:CPD

[Rem08] Daniel Remenik. The contact process in a dynamic random environment. Annals of Applied Probability, 18(6):2392–2420, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1227708923.

Ren:2016:VSF

[Ren16] Zhenjie Ren. Viscosity solutions of fully nonlinear elliptic path dependent partial differential equations. Annals of Applied Probability, 26(6):3381–3414, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1481792588.

Reygner:2018:ELD

[Rey18] Julien Reygner. Equilibrium large deviations for mean-field systems with translation invariance. Annals of Applied Probability, 28(5):2922–2965, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1535443238.

Rhee:1993:MPS

[Rhe93] WanSoo T. Rhee. A matching problem and subadditive Euclidean functionals. Annals of Applied Probability, 3(3):794–801, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005364.

Rhee:1994:PAC

[Rhe94] WanSoo T. Rhee. Probabilistic analysis of a capacitated vehicle routing problem II. Annals of Applied Probability, 4(3):741–764, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004969.

Rhee:1995:RCC

[Rhe95] WanSoo T. Rhee. On rates of convergence for common subsequences and first passage time. Annals of Applied Probability, 5(1):44–48, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004827.

Rhee:2000:ODW

[Rhe00] WanSoo T. Rhee. Order of decay of the wasted space for a stochastic packing problem. Annals of Applied Probability, 10(2): 539–548, May 2000. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1019487354.

Richou:2011:NSB

[Ric11] Adrien Richou. Numerical simulation of BSDEs with drivers of quadratic growth. Annals of Applied Probability, 21(5): 1933–1964, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576613.

Rogers:1994:FMQ

[Rog94] L. C. G. Rogers. Fluid models in queueing theory and Wiener– Hopf factorization of Markov chains. Annals of Applied Probability, 4(2):390–413, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005065.

Roitershtein:2007:ODL

[Roi07] Alexander Roitershtein. One-dimensional linear recursions with Markov-dependent coefficients. Annals of Applied Probability, 17(2):572–608, April 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1174323257.

Rollin:2007:TPA

[Röl07] Adrian Röllin. Translated Poisson approximation using exchangeable pair couplings. Annals of Applied Probability, 17 (5/6):1596–1614, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1191419177.

Roos:1994:SMC

[Roo94] Małgorzata Roos. Stein’s method for compound Poisson approximation: The local approach. Annals of Applied Probability, 4 (4):1177–1187, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004910.

Rosenthal:1993:RCD

[Ros93] Jeffrey S. Rosenthal. Rates of convergence for data augmentation on finite sample spaces. Annals of Applied Probability, 3(3):819– 839, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177005366.

Rosenblatt:1995:PNG

[Ros95] Murray Rosenblatt. Prediction and non-Gaussian autoregressive stationary sequences. Annals of Applied Probability, 5(1):239– 247, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004838.

Resnick:1991:RUF

[RR91] Sidney I. Resnick and Rishin Roy. Random USC functions, maxstable processes and continuous choice. Annals of Applied Probability, 1(2):267–292, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005937.

Resnick:1994:SPM

[RR94] Sidney I. Resnick and Rishin Roy. Superextremal processes, max-stability and dynamic continuous choice. Annals of Applied Probability, 4(3):791–811, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004972.

Rinott:1997:CCR

[RR97a] Yosef Rinott and Vladimir Rotar. On coupling constructions and rates in the CLT for dependent summands with applications to the antivoter model and weighted U-statistics. Annals of Applied Probability, 7(4):1080–1105, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1043862425.

Rothmann:1997:LLN

[RR97b] Mark D. Rothmann and Ralph P. Russo. A law of large numbers on randomly deleted sets. Annals of Applied Probability, 7(1):170–182, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034625258.

Roberts:1998:CRG

[RR98a] Gareth O. Roberts and Jeffrey S. Rosenthal. On convergence rates of Gibbs samplers for uniform distributions. Annals of Applied Probability, 8(4):1291–1302, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1028903381.

Roberts:1998:TCP

[RR98b] Gareth O. Roberts and Jeffrey S. Rosenthal. Two convergence properties of hybrid samplers. Annals of Applied Probability, 8(2):397–407, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903533.

Resnick:2000:SSC

[RR00] Sidney Resnick and Holger Rootzén. Self-similar communication models and very heavy tails. Annals of Applied Probability, 10(3):753–778, August 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1019487509.

Ramanan:2003:FHT

[RR03] Kavita Ramanan and Martin I. Reiman. Fluid and heavy traffic diffusion limits for a generalized processor sharing model. Annals of Applied Probability, 13(1):100–139, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1042765664.

Roberts:2006:HRM

[RR06] Gareth O. Roberts and Jeffrey S. Rosenthal. Harris recurrence of Metropolis-within-Gibbs and trans-dimensional Markov chains. Annals of Applied Probability, 16(4):2123–2139, November 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1169065219.

Ramanan:2008:HTL

[RR08a] Kavita Ramanan and Martin I. Reiman. The heavy traffic limit of an unbalanced generalized processor sharing model. Annals of Applied Probability, 18(1):22–58, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1199890014.

Roberts:2008:VBM

[RR08b] Gareth O. Roberts and Jeffrey S. Rosenthal. Variance bounding Markov chains. Annals of Applied Probability, 18(3):1201–1214, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1211819798.

Roberts:2014:MMV

[RR14] Gareth O. Roberts and Jeffrey S. Rosenthal. Minimising MCMC variance via diffusion limits, with an application to simulated tempering. Annals of Applied Probability, 24(1):131–149, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1389278722.

Reinert:2019:ASD

[RR19] Gesine Reinert and Nathan Ross. Approximating stationary distributions of fast mixing Glauber dynamics, with applications to exponential random graphs. Annals of Applied Probability, 29(5):3201–3229, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1571385633.

Rogers:1991:EVH

[RS91] L. C. G. Rogers and S. E. Satchell. Estimating variance from high, low and closing prices. Annals of Applied Probability, 1 (4):504–512, November 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005835.

Roberts:1995:HRT

[RS95] G. O. Roberts and C. F. Shortland. The hazard rate tangent approximation for boundary hitting times. Annals of Applied Probability, 5(2):446–460, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004773.

Resnick:1998:TIE

[RS98] Sidney Resnick and Catalin Stărică. Tail index estimation for dependent data. Annals of Applied Probability, 8(4):1156–1183, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1028903376.

Ramanan:2001:LWD

[RS01] Kavita Ramanan and Alexander L. Stolyar. Largest weighted delay first scheduling: Large deviations and optimality. Annals of Applied Probability, 11(1):1–48, February 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/998926986.

Rabehasaina:2003:SAS

[RS03a] Landy Rabehasaina and Bruno Sericola. Stability analysis of second-order fluid flow models in a stationary ergodic environment. Annals of Applied Probability, 13(4):1449–1473, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1069786505.

Resnick:2003:LHP

[RS03b] Sidney Resnick and Gennady Samorodnitsky. Limits of on/off hierarchical product models for data transmission. Annals of Applied Probability, 13(4):1355–1398, November 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1069786502.

Rasonyi:2005:UMD

[RS05] Miklós Rásonyi and Lukasz Stettner. On utility maximization in discrete-time financial market models. Annals of Applied Probability, 15(2):1367–1395, May 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1115137978.

Rheinlander:2006:MEM

[RS06] Thorsten Rheinländer and Gallus Steiger. The minimal entropy martingale measure for general Barndorff–Nielsen/Shephard models. Annals of Applied Probability, 16(3):1319–1351, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804983.

Russo:2007:RPJ

[RS07] Ralph P. Russo and Nariankadu D. Shyamalkumar. Reading policies for joins: An asymptotic analysis. Annals of Applied Probability, 17(1):230–264, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1171377183.

Rider:2014:ELR

[RS14] Brian Rider and Christopher D. Sinclair. Extremal laws for the real Ginibre ensemble. Annals of Applied Probability, 24 (4):1621–1651, August 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1400073659.

Racz:2015:MSB

[RS15] Miklós Z. Rácz and Mykhaylo Shkolnikov. Multidimensional sticky Brownian motions as limits of exclusion processes. Annals of Applied Probability, 25(3):1155–1188, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1427124126.

Roberts:2018:ETC

[RŞ18] Matthew I. Roberts and Batı Şengül. Exceptional times of the critical dynamical Erdős–Rényi graph. Annals of Applied Probability, 28(4):2275–2308, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1533780273.

Racz:2022:CRG

[RS22] Miklós Z. Rácz and Anirudh Sridhar. Correlated randomly growing graphs. Annals of Applied Probability, 32(2):1058–1111, April 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-2/Correlated-
randomly-growing-graphs/10.1214/21-AAP1703.full.

Rosenzweig:2023:GTM

[RS23] Matthew Rosenzweig and Sylvia Serfaty. Global-in-time meanfield convergence for singular Riesz-type diffusive flows. Annals of Applied Probability, 33(2):754–798, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-2/Global-in-time-mean-field-convergence-
for-singular-Riesz-type/10.1214/22-AAP1833.full.

Rataj:2009:AWS

[RSM09] Jan Rataj, Evgeny Spodarev, and Daniel Meschenmoser. Approximations of the Wiener sausage and its curvature measures. Annals of Applied Probability, 19(5):1840–1859, October 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1255699545.

Roy:2016:HLD

[RSS16] Rahul Roy, Kumarjit Saha, and Anish Sarkar. Hack’s law in a drainage network model: a Brownian web approach. Annals of Applied Probability, 26(3):1807–1836, June 2016. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1465905020.

Resnick:1999:HMC

[RSX99] Sidney Resnick, Gennady Samorodnitsky, and Fang Xue. How misleading can sample ACFs of stable MAs be? (very!). Annals of Applied Probability, 9(3):797–817, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1029962814.

Rhee:1992:MRS

[RT92] WanSoo T. Rhee and Michel Talagrand. Matching random subsets of the cube with a tight control on one coordinate. Annals of Applied Probability, 2(3):695–713, August 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005655.

Rhee:1996:PRI

[RT96] WanSoo T. Rhee and Michel Talagrand. Packing random intervals. Annals of Applied Probability, 6(2):572–576, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1034968145.

Rosenbaum:2014:AOD

[RT14] Mathieu Rosenbaum and Peter Tankov. Asymptotically optimal discretization of hedging strategies with jumps. Annals of Applied Probability, 24(3):1002–1048, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1398258094.

Reed:2015:DVH

[RT15] Josh Reed and Rishi Talreja. Distribution-valued heavy-traffic limits for the G/GI/∞ queue. Annals of Applied Probability, 25(3):1420–1474, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

Ruschendorf:2008:COS

[RU08] Ludger Rüschendorf and Mikhail A. Urusov. On a class of optimal stopping problems for diffusions with discontinuous coefficients. Annals of Applied Probability, 18(3):847–878, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1211819787.

Robert:2015:SAR

[RV15] Philippe Robert and Amandine Véber. A stochastic analysis of resource sharing with logarithmic weights. Annals of Applied Probability, 25(5):2626–2670, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1438261050.

Rebeschini:2015:CLP

[RvH15] Patrick Rebeschini and Ramon van Handel. Can local particle filters beat the curse of dimensionality? Annals of Applied Probability, 25(5):2809–2866, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1438261054.

Rucinski:1997:RGP

[RW97] A. Ruciński and N. C. Wormald. Random graph processes with maximum degree 2. Annals of Applied Probability, 7(1):183–199, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034625259.

Riordan:2012:APP

[RW12] Oliver Riordan and Lutz Warnke. Achlioptas process phase transitions are continuous. Annals of Applied Probability, 22 (4):1450–1464, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1344614200.

Ruggiero:2013:ADP

[RWF13] Matteo Ruggiero, Stephen G. Walker, and Stefano Favaro. Alpha-diversity processes and normalized inverse-Gaussian diffusions. Annals of Applied Probability, 23(1):386–425, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1359124390.

Ritter:1995:MIA

[RWW95] Klaus Ritter, Grzegorz W. Wasilkowski, and Henryk Wozniakowski. Multivariate integration and approximation for random fields satisfying Sacks–Ylvisaker conditions. Annals of Applied Probability, 5(2):518–540, May 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004776.

Redmond:1994:LTR

[RY94] C. Redmond and J. E. Yukich. Limit theorems and rates of convergence for Euclidean functionals. Annals of Applied Probability, 4(4):1057–1073, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177004902.

Radin:2013:PTE

[RY13] Charles Radin and Mei Yin. Phase transitions in exponential random graphs. Annals of Applied Probability, 23(6):2458–2471, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1382447694.

Rockner:2021:CST

[RYZ21] Michael Röckner, Huanyu Yang, and Rongchan Zhu. Conservative stochastic two-dimensional Cahn–Hilliard equation. Annals of Applied Probability, 31(3):1336–1375, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/journals/annals-of-applied-
probability/volume-31/issue-3/Conservative-stochastic-
two-dimensional-CahnHilliard-equation/10.1214/20-AAP1620.
full.

Rogers:1999:SAO

[RZ99] L. C. G. Rogers and O. Zane. Saddlepoint approximations to option prices. Annals of Applied Probability, 9(2):493–503, May 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962752.

Rogers:2008:ECH

[RZ08] L. C. G. Rogers and Fanyin Zhou. Estimating correlation from high, low, opening and closing prices. Annals of Applied Probability, 18(2):813–823, April 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1206018204.

Rodosthenous:2018:BOC

[RZ18] Neofytos Rodosthenous and Hongzhong Zhang. Beating the omega clock: An optimal stopping problem with random timehorizon under spectrally negative Lévy models. Annals of Applied Probability, 28(4):2105–2140, August 2018. CODEN ????

ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1533780269.

Rollin:2023:DMV

[RZ23] Adrian Röllin and Zhuo-Song Zhang. Dense multigraphonvalued stochastic processes and edge-changing dynamics in the configuration model. Annals of Applied Probability, 33(4):3207– 3239, August 2023. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-33/issue-
4/Dense-multigraphon-valued-stochastic-processes-and-
edge-changing-dynamics-in/10.1214/22-AAP1889.full.

Sabanis:2016:EAV

[Sab16] Sotirios Sabanis. Euler approximations with varying coefficients: The case of superlinearly growing diffusion coefficients. Annals of Applied Probability, 26(4):2083–2105, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1472745452.

Sadowsky:1996:MCE

[Sad96] John S. Sadowsky. On Monte Carlo estimation of large deviations probabilities. Annals of Applied Probability, 6(2):399–422, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034968137.

Sadeghi:1998:ABS

[Sad98] Ali A. Sadeghi. Asymptotic behavior of self-organizing maps with nonuniform stimuli distribution. Annals of Applied Probability, 8(1):281–299, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1027961044.

Sandberg:2008:NSH

[San08] Oskar Sandberg. Neighbor selection and hitting probability in small-world graphs. Annals of Applied Probability, 18(5): 1771–1793, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1225372949.

Santos:2010:CPS

[San10] Manuel S. Santos. Consistency properties of a simulation-based estimator for dynamic processes. Annals of Applied Probability, 20(1):196–213, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1262962321.

Sasada:2018:GKF

[Sas18] Makiko Sasada. On the Green–Kubo formula and the gradient condition on currents. Annals of Applied Probability, 28 (5):2727–2739, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1535443232.

Sankararaman:2019:IQN

[SBF19] Abishek Sankararaman, François Baccelli, and Sergey Foss. Interference queueing networks on grids. Annals of Applied Probability, 29(5):2929–2987, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https:

//projecteuclid.org/euclid.aoap/1571385626.

Schweizer:1992:MVH

[Sch92] Martin Schweizer. Mean-variance hedging for general claims. Annals of Applied Probability, 2(1):171–179, February 1992. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005776.

Schinazi:1997:PPH

[Sch97a] Rinaldo B. Schinazi. Predator-prey and host-parasite spatial stochastic models. Annals of Applied Probability, 7(1):1–9, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1034625250.

Schmidli:1997:ERT

[Sch97b] H. Schmidli. An extension to the renewal theorem and an application to risk theory. Annals of Applied Probability, 7(1):121– 133, February 1997. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1034625255.

Schachermayer:2001:OII

[Sch01] Walter Schachermayer. Optimal investment in incomplete markets when wealth may become negative. Annals of Applied Probability, 11(3):694–734, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1015345346.

Schmidli:2002:MRP

[Sch02] Hanspeter Schmidli. On minimizing the ruin probability by investment and reinsurance. Annals of Applied Probability, 12 (3):890–907, August 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1031863173.

Schied:2004:NPP

[Sch04] Alexander Schied. On the Neyman–Pearson problem for lawinvariant risk measures and robust utility functionals. Annals of Applied Probability, 14(3):1398–1423, August 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1089736290.

Schinazi:2005:MEA

[Sch05a] Rinaldo B. Schinazi. Mass extinctions: an alternative to the Allee effect. Annals of Applied Probability, 15(1B):984–991, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1107271675.

Schuhmacher:2005:UBS

[Sch05b] Dominic Schuhmacher. Upper bounds for spatial point process approximations. Annals of Applied Probability, 15(1B):615–651, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1107271662.

Schopp:2010:FLT

[Sch10] Eva-Maria Schopp. A functional limit theorem for the profile of b-ary trees. Annals of Applied Probability, 20(3):907–950, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1276867302.

Schied:2013:CPF

[Sch13] Alexander Schied. A control problem with fuel constraint and Dawson–Watanabe superprocesses. Annals of Applied Probability, 23(6):2472–2499, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1382447695.

Schapira:2021:LSV

[Sch21] Bruno Schapira. Localization on 5 sites for vertex reinforced random walks: Towards a characterization. Annals of Applied Probability, 31(4):1774–1786, August 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-4/Localization-on-5-sites-for-vertex-
reinforced-random-walks/10.1214/20-AAP1633.full.

Saloff-Coste:2010:TIM

[SCZ10] L. Saloff-Coste and J. Zúñiga. Time inhomogeneous Markov chains with wave-like behavior. Annals of Applied Probability, 20(5):1831–1853, October 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1282747402.

Schweinsberg:2005:RPA

[SD05] Jason Schweinsberg and Rick Durrett. Random partitions approximating the coalescence of lineages during a selective sweep. Annals of Applied Probability, 15(3):1591–1651, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1121433763.

Seeger:2020:ASV

[See20] Benjamin Seeger. Approximation schemes for viscosity solutions of fully nonlinear stochastic partial differential equations. Annals of Applied Probability, 30(4):1784–1823, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/Approximation-
schemes-for-viscosity-solutions-of-fully-nonlinear-stochastic-
partial/10.1214/19-AAP1543.full.

Seierstad:2009:CLT

[Sei09] Taral Guldahl Seierstad. A central limit theorem via differential equations. Annals of Applied Probability, 19(2):661–675, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1241702246.

Sellke:1995:HMI

[Sel95] Thomas M. Sellke. How many IID samples does it take to see all the balls in a Box? Annals of Applied Probability, 5(1):294–309, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004841.

Seppalainen:1997:ISI

[Sep97a] Timo Seppäläinen. Increasing sequences of independent points on the planar lattice. Annals of Applied Probability, 7(4): 886–898, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1043862416.

Seppalainen:1997:SLQ

[Sep97b] Timo Seppäläinen. A scaling limit for queues in series. Annals of Applied Probability, 7(4):855–872, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1043862414.

Serinko:1994:CAL

[Ser94] Regis J. Serinko. A consistent approach to least squares estimation of correlation dimension in weak Bernoulli dynamical systems. Annals of Applied Probability, 4(4):1234–1254, November 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004914.

Sezer:2010:WDP

[Sez10] Semih Onur Sezer. On the Wiener disorder problem. Annals of Applied Probability, 20(4):1537–1566, August 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1279638794.

Sun:2016:ALU

[SFR16] Wen Sun, Mathieu Feuillet, and Philippe Robert. Analysis of large unreliable stochastic networks. Annals of Applied Probability, 26(5):2959–3000, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1476884309.

Shaikhet:2015:NCN

[Sha15] Gennady Shaikhet. Necessary condition for null controllability in many-server heavy traffic. Annals of Applied Probability, 25 (1):406–428, February 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1418740190.

Shapira:2020:KCM

[Sha20] Assaf Shapira. Kinetically constrained models with random constraints. Annals of Applied Probability, 30(2):987–1006, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1591603228.

Shen:2002:APC

[She02] Lian Shen. Asymptotic properties of certain anisotropic walks in random media. Annals of Applied Probability, 12(2):477–510, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1026915612.

Shkolnikov:2011:CPS

[Shk11] Mykhaylo Shkolnikov. Competing particle systems evolving by interacting Lévy processes. Annals of Applied Probability, 21 (5):1911–1932, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1319576612.

Silvestrov:1996:RRG

[Sil96] Dmitrii S. Silvestrov. Recurrence relations for generalized hitting times for semi-Markov processes. Annals of Applied Probability, 6(2):617–649, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968147.

Sagitov:2005:CES

[SJ05] Serik Sagitov and Peter Jagers. The coalescent effective size of age-structured populations. Annals of Applied Probability, 15 (3):1778–1797, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433769.

Sly:2008:UTT

[Sly08] Allan Sly. Uniqueness thresholds on trees versus graphs. Annals of Applied Probability, 18(5):1897–1909, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1225372955.

Smith:2014:GSS

[Smi14] Aaron Smith. A Gibbs sampler on the n-simplex. Annals of Applied Probability, 24(1):114–130, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278721.

Shepp:1993:ROR

[SS93] Larry Shepp and A. N. Shiryaev. The Russian option: Reduced regret. Annals of Applied Probability, 3(3):631–640, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005355.

Shreve:1994:OIC

[SS94] S. E. Shreve and H. M. Soner. Optimal investment and consumption with transaction costs. Annals of Applied Probability, 4(3):609–692, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004966.

Steel:2006:VDT

[SS06] M. A. Steel and L. A. Székely. On the variational distance of two trees. Annals of Applied Probability, 16(3):1563–1575, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804991.

Sturm:2008:VMH

[SS08] Anja Sturm and Jan Swart. Voter models with heterozygosity selection. Annals of Applied Probability, 18(1):59–99, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1199890015.

Shah:2012:RSA

[SS12] Devavrat Shah and Jinwoo Shin. Randomized scheduling algorithm for queueing networks. Annals of Applied Probability, 22 (1):128–171, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1328623697.

Sturm:2015:PSC

[SS15] Anja Sturm and Jan M. Swart. A particle system with cooperative branching and coalescence. Annals of Applied Probability, 25(3):1616–1649, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1427124138.

Shneer:2018:SCD

[SS18] Seva Shneer and Alexander Stolyar. Stability conditions for a discrete-time decentralised medium access algorithm. Annals of Applied Probability, 28(6):3600–3628, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1538985630.

Schlichting:2019:PLS

[SS19] André Schlichting and Martin Slowik. Poincaré and logarithmic Sobolev constants for metastable Markov chains via capacitary inequalities. Annals of Applied Probability, 29(6):3438–3488, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1578366319.

Soner:1995:TNN

[SSC95] H. M. Soner, S. E. Shreve, and J. Cvitanic. There is no nontrivial hedging portfolio for option pricing with transaction costs. Annals of Applied Probability, 5(2):327–355, May 1995. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177004767.

Schumacher:2018:GCT

[SSV18] Christoph Schumacher, Fabian Schwarzenberger, and Ivan Veselić. Glivenko–Cantelli theory, Ornstein–Weiss quasi-tilings, and uniform ergodic theorems for distribution-valued fields over amenable groups. Annals of Applied Probability, 28(4):2417– 2450, August 2018. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1533780277.

Siegmund-Schultze:2006:IGT

[SSW06] Rainer Siegmund-Schultze and Wolfgang Wagner. Induced gelation in a two-site spatial coagulation model. Annals of Applied Probability, 16(1):370–402, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1141654291.

Spencer:2014:QFI

[SSX14] Joel Spencer, Madhu Sudan, and Kuang Xu. Queuing with future information. Annals of Applied Probability, 24(5):2091– 2142, October 2014. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL http://projecteuclid.org/
euclid.aoap/1403812371.

Sun:2019:EIB

[SSY19] Rongfeng Sun, Jan M. Swart, and Jinjiong Yu. Equilibrium interfaces of biased voter models. Annals of Applied Probability, 29(4):2556–2593, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1563869050.

Schonmann:1998:LMF

[ST98] Roberto H. Schonmann and Nelson I. Tanaka. Lack of monotonicity in ferromagnetic Ising model phase diagrams. Annals of Applied Probability, 8(1):234–245, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1027961042.

Silvestrov:2004:LTM

[ST04] Dmitrii S. Silvestrov and Jozef L. Teugels. Limit theorems for mixed max–sum processes with renewal stopping. Annals of Applied Probability, 14(4):1838–1868, November 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1099674080.

Simatos:2010:SHS

[ST10] Florian Simatos and Danielle Tibi. Spatial homogenization in a stochastic network with mobility. Annals of Applied Probability, 20(1):312–355, February 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1262962325.

Szpruch:2021:AMS

[ST21] Lukasz Szpruch and Alvin Tse. Antithetic multilevel sampling method for nonlinear functionals of measure. Annals of Applied Probability, 31(3):1100–1139, June 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-31/issue-3/Antithetic-multilevel-sampling-method-
for-nonlinear-functionals-of-measure/10.1214/20-AAP1614.
full.

Stadje:1997:ABS

[Sta97] Wolfgang Stadje. Asymptotic behavior of some interactive population flow models. Annals of Applied Probability, 7(4):837–854, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1043862413.

Stacey:2003:PIP

[Sta03] Alan Stacey. Partial immunization processes. Annals of Applied Probability, 13(2):669–690, May 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1050689599.

Stauffer:2015:STP

[Sta15] Alexandre Stauffer. Space-time percolation and detection by mobile nodes. Annals of Applied Probability, 25(5):2416–2461, October 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1438261045.

Stein:1993:APC

[Ste93] Michael L. Stein. Asymptotic properties of centered systematic sampling for predicting integrals of spatial processes. Annals of Applied Probability, 3(3):874–880, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177005369.

Stein:1995:PIS

[Ste95] Michael L. Stein. Predicting integrals of stochastic processes. Annals of Applied Probability, 5(1):158–170, February 1995. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177004834.

Stein:1999:PRF

[Ste99] Michael L. Stein. Predicting random fields with increasing dense observations. Annals of Applied Probability, 9(1):242–273, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1029962604.

Stolyar:2003:CEE

[Sto03] Alexander L. Stolyar. Control of end-to-end delay tails in a multiclass network: LWDF discipline optimality. Annals of Applied Probability, 13(3):1151–1206, August 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1060202838.

Stolyar:2004:MSG

[Sto04] Alexander L. Stolyar. MaxWeight scheduling in a generalized switch: State space collapse and workload minimization in heavy traffic. Annals of Applied Probability, 14(1):1–53, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1075828046.

Strasser:2005:CAF

[Str05] Eva Strasser. Characterization of arbitrage-free markets. Annals of Applied Probability, 15(1A):116–124, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1106922323.

Szpruch:2019:IMP

[STT19] Lukasz Szpruch, Shuren Tan, and Alvin Tse. Iterative multilevel particle approximation for McKean–Vlasov SDEs. Annals of Applied Probability, 29(4):2230–2265, August 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1563869042.

Soner:2013:DFS

[STZ13] H. Mete Soner, Nizar Touzi, and Jianfeng Zhang. Dual formulation of second order target problems. Annals of Applied Probability, 23(1):308–347, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1359124388.

Shah:2014:QPF

[STZ14] D. Shah, J. N. Tsitsiklis, and Y. Zhong. Qualitative properties of α-fair policies in bandwidth-sharing networks. Annals of Applied Probability, 24(1):76–113, February 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1389278720.

Sudbury:2008:EDR

[Sud08] Aidan Sudbury. The expected duration of random sequential adsorption. Annals of Applied Probability, 18(1):100–108, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1199890016.

Schwerer:1994:BMC

[SV94] Elizabeth Schwerer and Jan A. Van Mieghem. Brownian models of closed queueing networks: Explicit solutions for balanced three-station systems. Annals of Applied Probability, 4(2):448– 477, May 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005068.

Solan:2002:UVR

[SV02] Eilon Solan and Nicolas Vieille. Uniform value in recursive games. Annals of Applied Probability, 12(4):1185–1201, November 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1037125859.

Saksman:2010:EAM

[SV10] Eero Saksman and Matti Vihola. On the ergodicity of the adaptive Metropolis algorithm on unbounded domains. Annals of Applied Probability, 20(6):2178–2203, December 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1287494558.

Shah:2012:SNM

[SW12] Devavrat Shah and Damon Wischik. Switched networks with maximum weight policies: Fluid approximation and multiplicative state space collapse. Annals of Applied Probability, 22(1):70– 127, February 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1328623696.

Swart:2018:RRS

[Swa18] Jan M. Swart. Rigorous results for the Stigler–Luckock model for the evolution of an order book. Annals of Applied Probability, 28(3):1491–1535, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1527840025.

Shah:2014:OQS

[SWZ14] D. Shah, N. S. Walton, and Y. Zhong. Optimal queue-size scaling in switched networks. Annals of Applied Probability, 24(6): 2207–2245, December 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1409058031.

Schnelli:2023:CRT

[SX23] Kevin Schnelli and Yuanyuan Xu. Convergence rate to the Tracy–Widom laws for the largest eigenvalue of sample covariance matrices. Annals of Applied Probability, 33(1):677–725, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Convergence-
rate-to-the-TracyWidom-laws-for-the-largest-eigenvalue/
10.1214/22-AAP1826.full.

Sun:2021:ISL

[SXY21] Jingrui Sun, Jie Xiong, and Jiongmin Yong. Indefinite stochastic linear-quadratic optimal control problems with random coefficients: Closed-loop representation of open-loop optimal controls. Annals of Applied Probability, 31(1):460–499, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/Indefinite-
stochastic-linear-quadratic-optimal-control-problems-
with-random-coefficients/10.1214/20-AAP1595.full.

Sigman:1994:FMI

[SY94] Karl Sigman and David D. Yao. Finite moments for inventory processes. Annals of Applied Probability, 4(3):765–778, August 1994. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177004970.

Serfozo:1998:MNP

[SY98] Richard F. Serfozo and Bingyi Yang. Markov network processes with string transitions. Annals of Applied Probability, 8(3):793– 821, August 1998. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1028903451.

Stolyar:2013:SLF

[SY13] Alexander L. Stolyar and Elena Yudovina. Systems with large flexible server pools: Instability of “natural” load balancing. Annals of Applied Probability, 23(5):2099–2138, October 2013. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1377696307.

Schulte:2023:RMN

[SY23] Matthias Schulte and J. E. Yukich. Rates of multivariate normal approximation for statistics in geometric probability. Annals of Applied Probability, 33(1):507–548, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Rates-
of-multivariate-normal-approximation-for-statistics-
in-geometric-probability/10.1214/22-AAP1822.full.

Song:2020:HDL

[SYY20] Jian Song, Jianfeng Yao, and Wangjun Yuan. High-dimensional limits of eigenvalue distributions for general Wishart process. Annals of Applied Probability, 30(4):1642–1668, August 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-30/issue-4/High-
dimensional-limits-of-eigenvalue-distributions-for-general-
Wishart-process/10.1214/19-AAP1539.full.

Saidi:2006:SGE

[SZ06] Youssef Saïdi and Jean-Michel Zakoïan. Stationarity and geometric ergodicity of a class of nonlinear ARCH models. Annals of Applied Probability, 16(4):2256–2271, November 2006. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1169065224.

Sly:2017:GDC

[SZ17] Allan Sly and Yumeng Zhang. The Glauber dynamics of colorings on trees is rapidly mixing throughout the nonreconstruction regime. Annals of Applied Probability, 27(5):2646–2674, October 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1509696031.

Shao:2021:CTM

[SZZ21] Qi-Man Shao, Mengchen Zhang, and Zhuo-Song Zhang. Cramértype moderate deviation theorems for nonnormal approximation. Annals of Applied Probability, 31(1):247–283, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/Cram%
c3%a9r-type-moderate-deviation-theorems-for-nonnormal-
approximation/10.1214/20-AAP1589.full.

Shen:2023:SAP

[SZZ23] Hao Shen, Rongchan Zhu, and Xiangchan Zhu. An SPDE approach to perturbation theory of. Annals of Applied Probability, 33(4):2600–2642, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-4/An-SPDE-approach-to-perturbation-theory-of-%
ce%a624--Asymptoticity/10.1214/22-AAP1873.full.

Taflin:2011:GIB

[Taf11] Erik Taflin. Generalized integrands and bond portfolios: Pitfalls and counter examples. Annals of Applied Probability, 21(1):266– 282, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1292598034.

Takacs:1992:RWP

[Tak92] Lajos Takacs. Random walk processes and their applications in order statistics. Annals of Applied Probability, 2(2):435–459, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005710.

Takacs:1993:DIA

[Tak93] Lajos Takacs. On the distribution of the integral of the absolute value of the Brownian motion. Annals of Applied Probability, 3(1):186–197, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005514.

Takacs:1995:LTB

[Tak95] Lajos Takacs. On the local time of the Brownian motion. Annals of Applied Probability, 5(3):741–756, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004703.

Takacs:1996:GAS

[Tak96] Lajos Takács. On a generalization of the arc-sine law. Annals of Applied Probability, 6(3):1035–1040, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968240.

Talagrand:1992:MRS

[Tal92] Michel Talagrand. Matching random samples in many dimensions. Annals of Applied Probability, 2(4):846–856, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005578.

Tan:2014:DTP

[Tan14] Xiaolu Tan. Discrete-time probabilistic approximation of pathdependent stochastic control problems. Annals of Applied Probability, 24(5):1803–1834, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1403812362.

Tremblay:2023:EEN

[TBUA23] Nicolas Tremblay, Simon Barthelmé, Konstantin Usevich, and Pierre-Olivier Amblard. Extended L-ensembles: a new representation for determinantal point processes. Annals of Applied Probability, 33(1):613–640, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-1/Extended-L-ensembles--A-new-representation-
for-determinantal-point/10.1214/22-AAP1824.full.

Tadic:2017:ABS

[TD17] Vladislav B. Tadić and Arnaud Doucet. Asymptotic bias of stochastic gradient search. Annals of Applied Probability, 27 (6):3255–3304, December 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1513328701.

Teixeira:2009:UIC

[Tei09] Augusto Teixeira. On the uniqueness of the infinite cluster of the vacant set of random interlacements. Annals of Applied Probability, 19(1):454–466, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1235140345.

Tierney:1998:NMH

[Tie98] Luke Tierney. A note on Metropolis–Hastings kernels for general state spaces. Annals of Applied Probability, 8(1):1–9, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1027961031.

Takemura:2002:ETE

[TK02] Akimichi Takemura and Satoshi Kuriki. On the equivalence of the tube and Euler characteristic methods for the distribution of the maximum of Gaussian fields over piecewise smooth domains. Annals of Applied Probability, 12(2):768–796, May 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1026915624.

Tanaka:2009:OAM

[TKH09] Hideyuki Tanaka and Arturo Kohatsu-Higa. An operator approach for Markov chain weak approximations with an application to infinite activity Lévy driven SDEs. Annals of Applied Probability, 19(3):1026–1062, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1245071018.

Tezuka:1993:LSA

[TLC93] Shu Tezuka, Pierre L’Ecuyer, and Raymond Couture. On the lattice structure of the add-with-carry and subtract-with-borrow random number generators. ACM Transactions on Modeling and Computer Simulation, 3(4):315–331, October 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See remark in [EH95, page 248], and [MZ91] for the original work analyzed in this paper.

Todorov:2019:NSV

[Tod19] Viktor Todorov. Nonparametric spot volatility from options. Annals of Applied Probability, 29(6):3590–3636, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1578366322.

Tomasevic:2021:NMV

[Tom21] Milica Tomasević. A new McKean–Vlasov stochastic interpretation of the parabolic-parabolic Keller–Segel model: The twodimensional case. Annals of Applied Probability, 31(1):432–459, February 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-1/A-new-
McKeanVlasov-stochastic-interpretation-of-the-parabolic-
parabolic-KellerSegel/10.1214/20-AAP1594.full.

Toninelli:2008:DPM

[Ton08] Fabio Lucio Toninelli. Disordered pinning models and copolymers: Beyond annealed bounds. Annals of Applied Probability, 18(4):1569–1587, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1216677132.

Toomey:2002:LDP

[Too02] Fergal Toomey. Large deviations of products of random topical operators. Annals of Applied Probability, 12(1):317–333, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015961166.

Torrisi:2016:GAN

[Tor16] Giovanni Luca Torrisi. Gaussian approximation of nonlinear Hawkes processes. Annals of Applied Probability, 26(4):2106– 2140, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1472745453.

Touzi:2000:OID

[Tou00] Nizar Touzi. Optimal insurance demand under marked point processes shocks. Annals of Applied Probability, 10(1):283–312, February 2000. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1019737674.

Touboul:2014:PCN

[Tou14] Jonathan Touboul. Propagation of chaos in neural fields. Annals of Applied Probability, 24(3):1298–1328, June 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1398258102. See erratum

[Tou18].

Touboul:2018:EPC

[Tou18] Jonathan Touboul. Erratum: “Propagation of chaos in neural fields” [Ann. Appl. Probab. 24 (2014) 1298–1328]. Annals of Applied Probability, 28(5):3287–3289, October 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1535443249. See [Tou14].

Tran:2023:MFZ

[Tra23] Hong Quan Tran. The mean-field zero-range process with unbounded monotone rates: Mixing time, cutoff, and Poincaré constant. Annals of Applied Probability, 33(3):1732–1757, June 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-3/The-
mean-field-zero-range-process-with-unbounded-monotone-
rates/10.1214/22-AAP1851.full.

Trillos:2015:DSM

[Tri15] Camilo Andrés García Trillos. A decreasing step method for strongly oscillating stochastic models. Annals of Applied Probability, 25(2):986–1029, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1424355136.

Tsitsiklis:1994:SPG

[Tsi94] John N. Tsitsiklis. A short proof of the Gittins Index Theorem. Annals of Applied Probability, 4(1):194–199, February 1994. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177005207.

Tameling:2019:EOT

[TSM19] Carla Tameling, Max Sommerfeld, and Axel Munk. Empirical optimal transport on countable metric spaces: Distributional limits and statistical applications. Annals of Applied Probability, 29(5):2744–2781, October 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/euclid.aoap/1571385621.

Tsukuda:2019:PAE

[Tsu19] Koji Tsukuda. On Poisson approximations for the Ewens sampling formula when the mutation parameter grows with the sample size. Annals of Applied Probability, 29(2):1188–1232, April 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1548298939.

Todorov:2011:LTP

[TT11] Viktor Todorov and George Tauchen. Limit theorems for power variations of pure-jump processes with application to activity estimation. Annals of Applied Probability, 21(2):546–588, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1300800981.

Teixeira:2013:RIA

[TT13] Augusto Teixeira and Johan Tykesson. Random interlacements and amenability. Annals of Applied Probability, 23(3):923–956, June 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1362684850.

Todorov:2014:LTE

[TT14] Viktor Todorov and George Tauchen. Limit theorems for the empirical distribution function of scaled increments of Itô semimartingales at high frequencies. Annals of Applied Probability, 24(5):1850–1888, October 2014. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1403812364.

Turova:2018:PTO

[Tur18] Tatyana S. Turova. Phase transitions in the one-dimensional Coulomb gas ensembles. Annals of Applied Probability, 28 (2):1249–1291, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1523433635.

Talay:2003:SPM

[TV03] Denis Talay and Olivier Vaillant. A stochastic particle method with random weights for the computation of statistical solutions of McKean–Vlasov equations. Annals of Applied Probability, 13(1):140–180, January 2003. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1042765665.

Tarres:2012:ETA

[TV12] Pierre Tarrès and Pierre Vandekerkhove. On ergodic twoarmed bandits. Annals of Applied Probability, 22(2):457–476, April 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1333372004.

Thacker:2018:BAM

[TV18] Debleena Thacker and Stanislav Volkov. Border aggregation model. Annals of Applied Probability, 28(3):1604–1633, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1527840028.

Tong:2012:ESC

[TvH12] Xin Thomson Tong and Ramon van Handel. Ergodicity and stability of the conditional distributions of nondegenerate Markov chains. Annals of Applied Probability, 22(4):1495–1540, August 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1344614202.

Tetali:2012:PTM

[TVVY12] Prasad Tetali, Juan C. Vera, Eric Vigoda, and Linji Yang. Phase transition for the mixing time of the Glauber dynamics for coloring regular trees. Annals of Applied Probability, 22(6):2210–2239, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1353695952.

Tracy:2007:NBE

[TW07] Craig A. Tracy and Harold Widom. Nonintersecting Brownian excursions. Annals of Applied Probability, 17(3):953–979, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1179839179.

Talreja:2009:HTL

[TW09] Rishi Talreja and Ward Whitt. Heavy-traffic limits for waiting times in many-server queues with abandonment. Annals of Applied Probability, 19(6):2137–2175, December 2009. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1259158769.

Talagrand:1993:ISE

[TY93] M. Talagrand and J. E. Yukich. The integrability of the square exponential transportation cost. Annals of Applied Probability, 3(4):1100–1111, November 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005274.

Terasawa:2011:SPL

[TY11] Yutaka Terasawa and Nobuo Yoshida. Stochastic power law fluids: Existence and uniqueness of weak solutions. Annals of Applied Probability, 21(5):1827–1859, October 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1319576610.

Takahashi:2016:WAA

[TY16] Akihiko Takahashi and Toshihiro Yamada. A weak approximation with asymptotic expansion and multidimensional Malliavin weights. Annals of Applied Probability, 26(2):818–856, April 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1458651821.

Verzelen:2015:CDS

[VAC15] Nicolas Verzelen and Ery Arias-Castro. Community detection in sparse random networks. Annals of Applied Probability, 25 (6):3465–3510, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1443703780.

vandenBerg:2011:SPT

[vdB11] J. van den Berg. Sharpness of the percolation transition in the two-dimensional contact process. Annals of Applied Probability, 21(1):374–395, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1292598039.

vandenBerg:1993:ITC

[vdBK93] J. van den Berg and H. Kesten. Inequalities for the time constant in first-passage percolation. Annals of Applied Probability, 3(1):56–80, February 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005507.

vandenBerg:2017:TDV

[vdBN17] Jacob van den Berg and Pierre Nolin. Two-dimensional volume-frozen percolation: Exceptional scales. Annals of Applied Probability, 27(1):91–108, February 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1488790823.

vanderHofstad:2016:SRP

[vdHHKR16] Remco van der Hofstad, Mark Holmes, Alexey Kuznetsov, and Wioletta Ruszel. Strongly reinforced Pólya urns with graphbased competition. Annals of Applied Probability, 26(4):2494– 2539, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1472745465.

vanderHofstad:2008:WAS

[vdHMS08] Remco van der Hofstad, Peter Mörters, and Nadia Sidorova. Weak and almost sure limits for the parabolic Anderson model with heavy tailed potentials. Annals of Applied Probability, 18 (6):2450–2494, December 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1227708925.

vanderHoorn:2018:TDD

[vdHOC18] Pim van der Hoorn and Mariana Olvera-Cravioto. Typical distances in the directed configuration model. Annals of Applied Probability, 28(3):1739–1792, June 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1527840031.

vanderHofstad:2021:OSS

[vdHvLS21] Remco van der Hofstad, Johan S. H. van Leeuwaarden, and Clara Stegehuis. Optimal subgraph structures in scale-free configuration models. Annals of Applied Probability, 31(2):501–537, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/Optimal-
subgraph-structures-in-scale-free-configuration-models/
10.1214/20-AAP1580.full.

vanEnter:2008:NRG

[vEK08] Aernout C. D. van Enter and Christof Külske. Nonexistence of random gradient Gibbs measures in continuous interface models in d =2. Annals of Applied Probability, 18(1):109–119, February 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1199890017.

Verloop:2016:AOP

[Ver16] I. M. Verloop. Asymptotically optimal priority policies for indexable and nonindexable restless bandits. Annals of Applied Probability, 26(4):1947–1995, August 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1472745449.

VanDenBerg:1995:TIC

[VG95] J. Van Den Berg and A. Gandolfi. A triangle inequality for covariances of binary FKG random variables. Annals of Applied Probability, 5(1):322–326, February 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1177004843.

vanHandel:2009:UOH

[vH09] Ramon van Handel. Uniform observability of hidden Markov models and filter stability for unstable signals. Annals of Applied Probability, 19(3):1172–1199, June 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1245071023.

Vitale:1991:EAR

[Vit91] Richard A. Vitale. Expected absolute random determinants and zonoids. Annals of Applied Probability, 1(2):293–300, May 1991.

CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005938.

Viveros:2023:DPV

[Viv23] Roberto Viveros. Directed polymer for very heavy tailed random walks. Annals of Applied Probability, 33(1):490–506, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-33/issue-1/Directed-
polymer-for-very-heavy-tailed-random-walks/10.1214/22-
AAP1821.full.

Vogrinc:2021:COS

[VK21] Jure Vogrinc and Wilfrid S. Kendall. Counterexamples for optimal scaling of Metropolis–Hastings chains with rough target densities. Annals of Applied Probability, 31(2):972–1019, April 2021. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-31/issue-2/Counterexamples-
for-optimal-scaling-of-MetropolisHastings-chains-with-
rough-target/10.1214/20-AAP1612.full.

vanMieghem:1995:DSC

[vM95] Jan A. van Mieghem. Dynamic scheduling with convex delay costs: The generalized c|mu rule. Annals of Applied Probability, 5(3):809–833, August 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004706.

Vath:2008:MSS

[VPV08] Vathana Ly Vath, Huyên Pham, and Stéphane Villeneuve. A mixed singular/switching control problem for a dividend policy with reversible technology investment. Annals of Applied Probability, 18(3):1164–1200, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1211819797.

Vysotsky:2008:CSM

[Vys08] Vladislav V. Vysotsky. Clustering in a stochastic model of onedimensional gas. Annals of Applied Probability, 18(3):1026–1058, June 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1211819793.

Viens:2019:MAF

[VZ19] Frederi Viens and Jianfeng Zhang. A martingale approach for fractional Brownian motions and related path dependent PDEs. Annals of Applied Probability, 29(6):3489–3540, December 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1578366320.

Wagner:2005:EPS

[Wag05] Wolfgang Wagner. Explosion phenomena in stochastic coagulation– fragmentation models. Annals of Applied Probability, 15(3): 2081–2112, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433778.

Walton:2009:PFR

[Wal09] N. S. Walton. Proportional fairness and its relationship with multi-class queueing networks. Annals of Applied Probability, 19 (6):2301–2333, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1259158773.

Wang:2023:CWD

[Wan23] Feng-Yu Wang. Convergence in Wasserstein distance for empirical measures of semilinear SPDEs. Annals of Applied Probability, 33(1):70–84, February 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-1/Convergence-in-Wasserstein-distance-for-empirical-
measures-of-semilinear-SPDEs/10.1214/22-AAP1807.full.

Weber:1992:GIM

[Web92] Richard Weber. On the Gittins index for multiarmed bandits. Annals of Applied Probability, 2(4):1024–1033, November 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1177005588.

Weber:2001:CNO

[Web01a] Matthias Weber. Correction note: “On occupation time functionals for diffusion processes and birth-and-death processes on graphs” [ Ann. Appl. Probab. 11 (2001), no. 2, 544–567; 1 843 057]. Annals of Applied Probability, 11(3):1003, August 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/

1015345356. See [Web01b].

Weber:2001:OTF

[Web01b] Matthias Weber. On occupation time functionals for diffusion processes and birth-and-death processes on graphs. Annals of Applied Probability, 11(2):544–567, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1015345303. See correction

[Web01a].

Weerasinghe:1998:SOS

[Wee98] Ananda P. N. Weerasinghe. Singular optimal strategies for investment with transaction costs. Annals of Applied Probability, 8(4):1312–1330, November 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1028903383.

Wahlberg:1993:PSM

[WH93] B. Wahlberg and E. J. Hannan. Parametric signal modelling using Laguerre filters. Annals of Applied Probability, 3(2):467– 496, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005434.

Watanabe:2017:FLA

[WH17] Shun Watanabe and Masahito Hayashi. Finite-length analysis on tail probability for Markov chain and application to simple hypothesis testing. Annals of Applied Probability, 27(2):811–845, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1495764367.

Whittle:1991:NNI

[Whi91] P. Whittle. Neural nets and implicit inference. Annals of Applied Probability, 1(2):173–188, May 1991. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1177005932.

Whiteley:2013:SPS

[Whi13] Nick Whiteley. Stability properties of some particle filters. Annals of Applied Probability, 23(6):2500–2537, December 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1382447696.

Walker:2007:FVP

[WHN07] Stephen G. Walker, Spyridon J. Hatjispyros, and Theodoros Nicoleris. A Fleming–Viot process and Bayesian nonparametrics. Annals of Applied Probability, 17(1):67–80, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1171377177.

Wiktorsson:2001:JCF

[Wik01] Magnus Wiktorsson. Joint characteristic function and simultaneous simulation of iterated Itô integrals for multiple independent Brownian motions. Annals of Applied Probability, 11(2):470–487, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345301.

Wilson:2004:MTL

[Wil04] David Bruce Wilson. Mixing times of lozenge tiling and card shuffling Markov chains. Annals of Applied Probability, 14 (1):274–325, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1075828054.

Windisch:2008:DDC

[Win08] David Windisch. On the disconnection of a discrete cylinder by a biased random walk. Annals of Applied Probability, 18 (4):1441–1490, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1216677128.

Wischik:2001:SPL

[Wis01] Damon J. Wischik. Sample path large deviations for queues with many inputs. Annals of Applied Probability, 11(2):379–404, May 2001. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015345296.

Wang:2015:SLE

[WJB^{+}15] Chen Wang, Baisuo Jin, Z. D. Bai, K. Krishnan Nair, and Matthew Harding. Strong limit of the extreme eigenvalues of a symmetrized auto-cross covariance matrix. Annals of Applied Probability, 25(6):3624–3683, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1443703784.

Wang:2019:TPQ

[WKRS19] Andi Q. Wang, Martin Kolb, Gareth O. Roberts, and David Steinsaltz. Theoretical properties of quasi-stationary Monte Carlo methods. Annals of Applied Probability, 29(1):434–457, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1544000433.

Whiteley:2016:PSN

[WL16] Nick Whiteley and Anthony Lee. Perfect sampling for nonhomogeneous Markov chains and hidden Markov models. Annals of Applied Probability, 26(5):3044–3077, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1476884311.

Weissman:2004:UPR

[WM04] Tsachy Weissman and Neri Merhav. Universal prediction of random binary sequences in a noisy environment. Annals of Applied Probability, 14(1):54–89, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1075828047.

Wood:2012:UCL

[Woo12] Philip Matchett Wood. Universality and the circular law for sparse random matrices. Annals of Applied Probability, 22 (3):1266–1300, June 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1337347545.

Wormald:1995:DER

[Wor95] Nicholas C. Wormald. Differential equations for random processes and random graphs. Annals of Applied Probability, 5(4): 1217–1235, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177004612.

Wang:2021:REM

[WPRS21] Andi Q. Wang, Murray Pollock, Gareth O. Roberts, and David Steinsaltz. Regeneration-enriched Markov processes with application to Monte Carlo. Annals of Applied Probability, 31(2): 703–735, April 2021. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-31/issue-
2/Regeneration-enriched-Markov-processes-with-application-
to-Monte-Carlo/10.1214/20-AAP1602.full.

Woodard:2009:CRM

[WSH09] Dawn B. Woodard, Scott C. Schmidler, and Mark Huber. Conditions for rapid mixing of parallel and simulated tempering on multimodal distributions. Annals of Applied Probability, 19(2):617–640, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1241702244.

Wu:2009:RET

[Wu09] Wei Biao Wu. Recursive estimation of time-average variance constants. Annals of Applied Probability, 19(4):1529–1552, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1248700626.

Wuthrich:2006:HMR

[Wüt06] Mario V. Wüthrich. A heteropolymer in a medium with random droplets. Annals of Applied Probability, 16(3):1653–1670, August 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1159804995.

Wang:2020:TII

[WW20] Neng-Yi Wang and Liming Wu. Transport-information inequalities for Markov chains. Annals of Applied Probability, 30(3): 1276–1320, June 2020. CODEN ???? ISSN 1050-5164 (print),

2168-8737 (electronic). URL https://projecteuclid.org/
journals/annals-of-applied-probability/volume-30/issue-
3/Transport-information-inequalities-for-Markov-chains/
10.1214/19-AAP1530.full.

Wu:2023:TCU

[WXY23] Yihong Wu, Jiaming Xu, and Sophie H. Yu. Testing correlation of unlabeled random graphs. Annals of Applied Probability, 33(4):2519–2558, August 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-33/
issue-4/Testing-correlation-of-unlabeled-random-graphs/
10.1214/22-AAP1786.full.

Wyner:1999:MRW

[Wyn99] Abraham J. Wyner. More on recurrence and waiting times. Annals of Applied Probability, 9(3):780–796, August 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1029962813.

Wu:2020:VSP

[WZ20] Cong Wu and Jianfeng Zhang. Viscosity solutions to parabolic master equations and McKean–Vlasov SDEs with closed-loop controls. Annals of Applied Probability, 30(2):936–986, April 2020. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/euclid.aoap/
1591603227.

Xia:1997:UFD

[Xia97] Aihua Xia. On using the first difference in the Stein–Chen method. Annals of Applied Probability, 7(4):899–916, November 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1043862417.

Xu:1992:DMOa

[XS92a] Gan-Lin Xu and Steven E. Shreve. A duality method for optimal consumption and investment under short- selling prohibition. I. General market coefficients. Annals of Applied Probability, 2(1):87–112, February 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005772.

Xu:1992:DMOb

[XS92b] Gan-Lin Xu and Steven E. Shreve. A duality method for optimal consumption and investment under short-selling prohibition. II. Constant market coefficients. Annals of Applied Probability, 2(2):314–328, May 1992. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005706.

Xu:2018:UPC

[Xu18] Liping Xu. Uniqueness and propagation of chaos for the Boltzmann equation with moderately soft potentials. Annals of Applied Probability, 28(2):1136–1189, April 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1523433633.

Xu:2019:ASL

[Xu19] Lihu Xu. Approximation of stable law in wasserstein-1 distance by Stein’s method. Annals of Applied Probability, 29(1):458–504, February 2019. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1544000434.

Xu:2013:OSU

[XZ13] Zuo Quan Xu and Xun Yu Zhou. Optimal stopping under probability distortion. Annals of Applied Probability, 23(1):251–282, February 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1359124386.

Yogeshwaran:2015:TRC

[YA15] D. Yogeshwaran and Robert J. Adler. On the topology of random complexes built over stationary point processes. Annals of Applied Probability, 25(6):3338–3380, December 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1443703776.

Yamada:1995:DAO

[Yam95] Keigo Yamada. Diffusion approximation for open statedependent queueing networks in the heavy traffic situation. Annals of Applied Probability, 5(4):958–982, November 1995. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1177004602.

Yan:2005:AEM

[Yan05] Liqing Yan. Asymptotic error for the Milstein scheme for SDEs driven by continuous semimartingales. Annals of Applied Probability, 15(4):2706–2738, November 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1133965777.

Yan:2020:NLD

[Yan20] Jun Yan. Nonlinear large deviations: Beyond the hypercube. Annals of Applied Probability, 30(2):812–846, April 2020. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1591603223.

Yao:1997:IRP

[Yao97] Yi-Ching Yao. On independence of k-record processes: Ignatov’s theorem revisited. Annals of Applied Probability, 7(3):815–821, August 1997. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1034801255.

Yatracos:2009:ADC

[Yat09] Yannis G. Yatracos. The asymptotic distribution of a clusterindex for i.i.d. normal random variables. Annals of Applied Probability, 19(2):585–595, April 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1241702242.

Ycart:1993:PPE

[Yca93] Bernard Ycart. The philosophers’ process: An ergodic reversible nearest particle system. Annals of Applied Probability, 3(2):356– 363, May 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005428.

Yu:2010:ABR

[YEC10] Feng Yu, Alison Etheridge, and Charles Cuthbertson. Asymptotic behavior of the rate of adaptation. Annals of Applied Probability, 20(3):978–1004, June 2010. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1276867304.

Yamada:1993:LTF

[YH93] Keigo Yamada and Sung Jo Hong. A limit theorem for functionals of jumps of birth and death processes under heavy traffic condition. Annals of Applied Probability, 3(3):840–862, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1177005367.

Yin:2022:PEE

[YM22] Yanqing Yin and Yanyuan Ma. Properties of eigenvalues and eigenvectors of large-dimensional sample correlation matrices. Annals of Applied Probability, 32(6):4763–4802, December 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-6/Properties-

of-eigenvalues-and-eigenvectors-of-large-dimensional-
sample-correlation/10.1214/22-AAP1802.full.

Yoshida:2008:CLT

[Yos08] Nobuo Yoshida. Central limit theorem for branching random walks in random environment. Annals of Applied Probability, 18(4):1619–1635, August 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1216677134.

Yoshida:2012:SST

[Yos12] Nobuo Yoshida. Stochastic shear thickening fluids: Strong convergence of the Galerkin approximation and the energy equality. Annals of Applied Probability, 22(3):1215–1242, June 2012. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1337347543.

Yang:2023:CRM

[YR23] Jun Yang and Jeffrey S. Rosenthal. Complexity results for MCMC derived from quantitative bounds. Annals of Applied Probability, 33(2):1259–1300, April 2023. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://

projecteuclid.org/journals/annals-of-applied-probability/
volume-33/issue-2/Complexity-results-for-MCMC-derived-
from-quantitative-bounds/10.1214/22-AAP1846.full.

Yin:2016:AQE

[YRF16] Mei Yin, Alessandro Rinaldo, and Sukhada Fadnavis. Asymptotic quantization of exponential random graphs. Annals of Applied Probability, 26(6):3251–3285, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1481792585.

Yao:1996:LDI

[YS96] Yi-Ching Yao and Gordon Simons. A large-dimensional independent and identically distributed property for nearest neighbor counts in Poisson processes. Annals of Applied Probability, 6(2):561–571, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968144.

Yasodharan:2022:LDM

[YS22] Sarath Yasodharan and Rajesh Sundaresan. Large deviations of mean-field interacting particle systems in a fast varying environment. Annals of Applied Probability, 32(3):1666–1704, June 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL https://projecteuclid.org/journals/
annals-of-applied-probability/volume-32/issue-3/Large-
deviations-of-mean-field-interacting-particle-systems-
in-a/10.1214/21-AAP1718.full.

Yu:2007:SDD

[Yu07a] Feng Yu. Stationary distribution for dioecious branching particle systems with rapid stirring. Annals of Applied Probability, 17 (5/6):1508–1537, October 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1197909501.

Yu:2007:SDM

[Yu07b] Feng Yu. Stationary distributions of a model of sympatric speciation. Annals of Applied Probability, 17(3):840–874, June 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1179839176.

Yu:2015:UMA

[Yu15] Xiang Yu. Utility maximization with addictive consumption habit formation in incomplete semimartingale markets. Annals of Applied Probability, 25(3):1383–1419, June 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1427124132.

Yu:2017:OCU

[Yu17a] Xiang Yu. Optimal consumption under habit formation in markets with transaction costs and random endowments. Annals of Applied Probability, 27(2):960–1002, April 2017. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1495764372.

Yu:2017:UCC

[Yu17b] Yaming Yu. On the unique crossing conjecture of Diaconis and Perlman on convolutions of gamma random variables. Annals of Applied Probability, 27(6):3893–3910, December 2017. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

https://projecteuclid.org/euclid.aoap/1513328716.

Yukich:1996:ETS

[Yuk96] J. E. Yukich. Ergodic theorems for some classical problems in combinatorial optimization. Annals of Applied Probability, 6 (3):1006–1023, August 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1034968238.

Yukich:1999:ALM

[Yuk99] J. E. Yukich. Asymptotics for the length of a minimal triangulation on a random sample. Annals of Applied Probability, 9(1):27–45, February 1999. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1029962596.

Yukich:2015:SOS

[Yuk15] J. E. Yukich. Surface order scaling in stochastic geometry. Annals of Applied Probability, 25(1):177–210, February 2015. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1418740183.

Yun:1998:EIH

[Yun98] Seokhoon Yun. The extremal index of a higher-order stationary Markov chain. Annals of Applied Probability, 8(2):408–437, May 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1028903534.

Yakovlev:2009:RFM

[YY09] Andrei Y. Yakovlev and Nikolay M. Yanev. Relative frequencies in multitype branching processes. Annals of Applied Probability, 19(1):1–14, February 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1235140330.

Ye:2018:JDA

[YY18] Heng-Qing Ye and David D. Yao. Justifying diffusion approximations for multiclass queueing networks under a moment condition. Annals of Applied Probability, 28(6):3652–3697, December 2018. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1538985632.

Yao:2022:EDL

[YY22] Jianfeng Yao and Wangjun Yuan. On eigenvalue distributions of large autocovariance matrices. Annals of Applied Probability, 32(5):3450–3491, October 2022. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL https://projecteuclid.

org/journals/annals-of-applied-probability/volume-32/
issue-5/On-eigenvalue-distributions-of-large-autocovariance-
matrices/10.1214/21-AAP1764.full.

Yin:2007:SPM

[YZ07] George Yin and Hanqin Zhang. Singularly perturbed Markov chains: Limit results and applications. Annals of Applied Probability, 17(1):207–229, February 2007. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1171377182.

Yilmaz:2019:NHO

[YZ19] Atilla Yilmaz and Ofer Zeitouni. Nonconvex homogenization for one-dimensional controlled random walks in random potential. Annals of Applied Probability, 29(1):36–88, February 2019. CO-DEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL

http://projecteuclid.org/euclid.aoap/1544000425.

Zwart:2005:SAH

[ZBD05] Bert Zwart, Sem Borst, and Krzystof Debicki. Subexponential asymptotics of hybrid fluid and ruin models. Annals of Applied Probability, 15(1A):500–517, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1106922335.

Zwart:2004:EAF

[ZBM04] Bert Zwart, Sem Borst, and Michel Mandjes. Exact asymptotics for fluid queues fed by multiple heavy-tailed on–off flows. Annals of Applied Probability, 14(2):903–957, May 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1082737117.

Zahle:2005:SSM

[ZCD05] Iljana Zähle, J. Theodore Cox, and Richard Durrett. The stepping stone model. II: Genealogies and the infinite sites model. Annals of Applied Probability, 15(1B):671–699, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1107271664.

Zhang:2011:DLL

[ZDZ11] Jiheng Zhang, J. G. Dai, and Bert Zwart. Diffusion limits of limited processor sharing queues. Annals of Applied Probability, 21(2):745–799, April 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1300800987.

Zerner:1998:DDG

[Zer98] Martin P. W. Zerner. Directional decay of the Green’s function for a random nonnegative potential on Z^{d}. Annals of Applied Probability, 8(1):246–280, February 1998. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1027961043.

Zhang:1995:LTM

[Zha95] Yu Zhang. A limit theorem for matching random sequences allowing deletions. Annals of Applied Probability, 5(4):1236–1240, November 1995. CODEN ???? ISSN 1050-5164 (print), 2168-

8737 (electronic). URL http://projecteuclid.org/euclid.
aoap/1177004613.

Zhang:2004:NSB

[Zha04] Jianfeng Zhang. A numerical scheme for BSDEs. Annals of Applied Probability, 14(1):459–488, February 2004. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1075828058.

Zhang:2005:RSB

[Zha05] Jianfeng Zhang. Representation of solutions to BSDEs associated with a degenerate FSDE. Annals of Applied Probability, 15(3):1798–1831, August 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1121433770.

Zhang:2012:SFD

[Zha12] Xicheng Zhang. Stochastic functional differential equations driven by Lévy processes and quasi-linear partial integrodifferential equations. Annals of Applied Probability, 22(6): 2505–2538, December 2012. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1353695960.

Zhang:2016:CLT

[Zha16a] Li-Xin Zhang. Central limit theorems of a recursive stochastic algorithm with applications to adaptive designs. Annals of Applied Probability, 26(6):3630–3658, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1481792595.

Zhang:2016:LAR

[Zha16b] Tusheng Zhang. Lattice approximations of reflected stochastic partial differential equations driven by space-time white noise. Annals of Applied Probability, 26(6):3602–3629, December 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic).

URL http://projecteuclid.org/euclid.aoap/1481792594.

Zhang:2016:SDE

[Zha16c] Xicheng Zhang. Stochastic differential equations with Sobolev diffusion and singular drift and applications. Annals of Applied Probability, 26(5):2697–2732, October 2016. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1476884301.

Zhang:2006:ATS

[ZHC06] Li-X. Zhang, Feifang Hu, and Siu Hung Cheung. Asymptotic theorems of sequential estimation-adjusted urn models. Annals of Applied Probability, 16(1):340–369, February 2006. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:

//projecteuclid.org/euclid.aoap/1141654290.

Zhu:2015:LDM

[Zhu15] Lingjiong Zhu. Large deviations for Markovian nonlinear Hawkes processes. Annals of Applied Probability, 25(2):548–581, April 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1424355123.

Zitkovic:2005:UMS

[Žit05] Gordan Žitković. Utility maximization with a stochastic clock and an unbounded random endowment. Annals of Applied Probability, 15(1B):748–777, February 2005. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1107271667.

Zitkovic:2009:DCS

[Žit09] Gordan Žitković. A dual characterization of self-generation and exponential forward performances. Annals of Applied Probability, 19(6):2176–2210, December 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http:/

/projecteuclid.org/euclid.aoap/1259158770.

Zhang:2015:RCS

[ZRH15] Hongzhong Zhang, Neofytos Rodosthenous, and Olympia Hadjiliadis. Robustness of the N-CUSUM stopping rule in a Wiener disorder problem. Annals of Applied Probability, 25(6): 3405–3433, December 2015. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1443703778.

Zucca:2009:IFP

[ZS09] Cristina Zucca and Laura Sacerdote. On the inverse firstpassage-time problem for a Wiener process. Annals of Applied Probability, 19(4):1319–1346, August 2009. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1248700619.

Zhao:2008:MA

[ZW08] Ou Zhao and Michael Woodroofe. On martingale approximations. Annals of Applied Probability, 18(5):1831–1847, October 2008. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1225372952.

Zhang:1996:CLT

[ZY96] Q. Zhang and G. Yin. A central limit theorem for singularly perturbed nonstationary finite state Markov chains. Annals of Applied Probability, 6(2):650–670, May 1996. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://

projecteuclid.org/euclid.aoap/1034968148.

Zachary:2002:RHK

[ZZ02] Stan Zachary and Ilze Ziedins. A refinement of the Hunt–Kurtz theory of large loss networks, with an application to virtual partitioning. Annals of Applied Probability, 12(1):1–22, February 2002. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1015961154.